A Complete Bibliography of Publications in the SIAM Journal on Applied Mathematics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 2.30

Title word cross-reference

#4348 [1574]. #5412 [2405]. #6819 [2450].

(m, 2) [1171]. (m − 3) [1767]. (m − 4) [1767]. (s, S) [2363, 1355, 90]. (σ, S) [586]. 0 [1328]. 1 [1316, 2337, 3314, 1328]. 1/f [2885]. 2π [2379]. 2 × 2 [3170]. 3 [659]. A [1637]. a > 0 [1404]. AD + DA′ = −C [554]. α [1878, 2144]. α sin ζ = ζ [958]. AX + XB = −Q [552]. AX + XB = C [1042, 337, 1384]. AX − XB = −C [866]. AX − XB = C [1183, 1495, 47]. AX − Y B = C [1495]. AX = XB [840]. AXB + CXD = E [1507]. AY + Y B = −Q [889]. AZ + B − ZCZ − ZD = 0 [1323]. B_{n}

[661]. β [2965, 3408]. β = −1 [1289]. C ; [985, 1900, 1901, 1472, 728]. C^{1} [3267]. D ; [1637, 158, 3534, 1089]. D/M/1 [1483]. DA ; [1637]. V̈ > 0 [1051]. ẍ + f(x, ẋ) + g(x) = 0 ; [91]. δ [2272]. E^{n} [185]. ϵ [402, 1117]. F ; [1155, 3513, 2089]. F (i + 1, k + 1, n + 1) = ; q^{k+1F} (i, k + 1, n) − q^{k}F (i, k, n) [1397]. ; f(x) = a [743]. ; \frac{d^{2}y}{dt^{2}} + w^{2}y = ε( ∑_{i=1}^{n} b_{i} cos 2a_{i}t)y [1375]. G ; [662, 2980]. G/M/m [2836]. γ = 0 [1289]. ; GF (2^{s}) [579]. GF (3 · 2^{n} + 1) [1466]. GI/G/1 ; [279]. H [1548, 2164, 2402]. I [330]. ; ∫_{0}^{1} ⋯ ∫_{0}^{1}(x_{1}^{2} + ⋯ + x_{k}^{2})^{1/2}dx_{1} ⋯ dx_{k} [1310]. ; J [530]. J_{1}^{2}(ζ) − J_{0}(ζ)J_{2}(ζ) = 0 [2784]. J_{m} ; [881]. K [110, 104, 184, 672, 52, 2337, 832]. ; L [428, 133, 984]. l^{1} [2169]. L^{2}(−∞, ∞) ; [2057]. L^{∞} [2307]. L_{2} [2338]. l_{p} ; [398, 1233, 471]. λ [2453, 3120]. λ − ω [1854].

LR [388]. L^{1} [3171]. Lu + λb(x)u = g(x, u) [720]. M [977]. M/D/1 [1201]. M/G/1 [891, 2342, 448]. M/G/2 [2399]. M/G/∞ [11]. R^{2} [1572]. R^{3} [3550]. ML [1637]. mth [1284]. M × M [337]. μ [1680, 1864]. M|G|Î [914]. N [2262, 1158, 1097, 1252, 2317, 238, 734, 828, 2240, 735, 72, 2680, 3421, 220, 248, 540, 1529, 1405, 207, 534, 553, 1684, 2417, 2852, 1598, 931, 1225, 1068]. n + 1 [255]. ∇^{2}u(x) = f(x) + k(x)u(x) [2344]. NP [1508]. n × n [1284, 337]. O(1) [3344]. O(2) [2916]. O(4) [1365]. ω [1739, 3120]. ω > 0 [1404]. {sgn(f(t_{i}) + x_{i})} [2089]. P [975, 690, 1923, 1775, 1776, 417, 2417]. p = 2 [1814]. p = 8m + 1 [1189]. p = 8m − 1 [986]. ∂_{y}u + ∂_{x}\sqrt{1 − u^{2}} = 0 [3296]. ∂u/∂t = −u · Δu [940]. ϕ^{4} [2068]. ϕ_{x}^{3}ϕ_{xx} = ϕ_{ỹỹ} [3501]. P N [1410]. Q [1407, 1585, 417]. Q_{ν−1/2}^{−μ}(cos θ) [2009]. R [1726, 3026, 1135]. R^{3} [2912, 3205]. R^{n} [843, 870, 1298]. RC [67]. ℜ^{3} [3477]. S [1742]. S^{T X} + XS = −Q [552]. S_{t} = [(βS + γ)^{−2}S_{x}]_{x} + α(βS + γ)_{x}^{−2S} [1989]. σ, S) [1802]. SO(3) [1698]. T [881, 985, 1571, 1155, 1432]. u′′(t) + a(t)f(u)g(u′) = 0 [21]. u_{0} [163]. V [2358]. ε [1404]. x′′ + A(t)g(x) = f(t) [843]. x′′ + a(t)x^{2n−1} = 0 [196]. x − ε(1 − x^{2n+2})x + x^{2n+1} = εa cos ωt [1404]. x − y [382]. XA + BX = C [270]. XAX = X [1113]. X → X − CXC [1151]. y′(x) = my(x − d(x)) [1160]. y′(x) = y(x − d(x)) [715]. y′′ + λp(x)y = 0, y(a) = y(b) = O [753]. y′′′ + py = 0 [1143]. y′′ + yF (y^{2}, x) = 0 [902]. y^{′2} − 2p(x)y = −p(x)q(x) [755]. Y B − AY = F [61]. ∥ϕ^{3}∥ ≤ c∥∇ϕ∥^{3} [728].

-1 [446]. -actin [3513]. -Approximation [398]. -arcs [3267]. -Ary [553]. -Baecklund [1171]. -Bifurcation [2453]. -Body [659]. -Cells [2965, 3408]. -centers [1775]. -Circulant [1135]. -closure [2980]. -Complete [1508]. -Complexity [1472].

-Components [832]. -convergence [3171]. -Criteria [471]. -Cube [248, 1684]. -Cyclic [881]. -D [3314]. -Densities [1680, 1864]. -Designs [1571]. -Determinants [614]. -Dimensional [2262, 1158, 828, 1598, 2680]. -Distributions [1155]. -Edge-Connectivity [1068]. -Entropy [402]. -Excited [238]. -Fold [1529]. -Function [1548, 2164, 2272, 2402]. -Functions [662]. -Interpolation [1878, 2144]. -Junction [2417]. -limits [1739]. -lines [1900]. -Matrices [1585, 1410, 690, 1407]. -Matrix [1637]. -medians [1776]. -Membrane [133]. -of- [2337]. -Partite [672]. -Person [540, 1097, 734, 735, 1405]. -Point [1252]. -Policies [2363, 1355]. -Ports [207]. -Positive [163]. -Sequence [977]. -Shaped [428]. -Soliton [3421]. -Space [158, 1233, 1089, 72]. -Spaces [220, 534]. -states [2358]. -Statistic [1432]. -Symmetry [2916]. -Tournaments [931, 1225]. -Transformations [330]. -Trees [52]. -varieties [1901]. -Vessel [2852]. -Wave [3534].

0-1 [1430, 1398].

1 [3030]. 1-1 [1169]. 1-Factors [602]. 11 [37]. 13 [1494, 139]. 14 [38, 144, 138, 184]. 15 [253, 534]. 17 [667]. 19 [854]. 1st [1931].

2-Variate [741]. 20 [1101, 792, 896]. 22 [988]. 24 [2405]. 25 [1156]. 27 [1232]. 28 [1361]. 2nd [1931].

3-Commodity [374]. 30 [1479]. 31

[1512, 1585, 2450, 1574]. 32 [1546, 1677]. 35 [1721]. 37 [1789, 1788]. 38 [1885]. 39 [1932].

4 [189]. 42 [2044, 2100]. 44 [2244, 2252]. 45 [2309]. 47 [2539, 2405]. 48 [2639, 2562].

50 [2731]. 51 [2903]. 52 [3043]. 53 [3183].

54 [1574]. 55 [2450].

80m [1885]. 83j [2100]. 85d [2244]. 86j [2309]. 88d [2539]. 89a [2562]. 89c [2639].

92d [2903]. 93h [3043]. 94h [3183].

ABC [2978]. ABC’s [3350]. Abel [3331, 3114, 171]. Abelian [2031, 104]. Abnormal [1827]. Above [473, 238]. Abscissa [764]. Absorbed [2594]. Absorbing [3224, 3062, 3093]. Absorption [2143, 3481, 107]. Abstract [1287]. AC [3453]. Accelerated [518]. Accelerating [2801]. Acceleration [1785]. Accelerators [2150, 2026, 2027]. Accommodation [3023]. Accompanied [2444]. According [78]. Accuracy [1271, 221, 2002, 131, 2005, 31]. Accurate [1961, 3493, 2108, 3191]. Ackerberg [1941]. Acknowledgment [2244]. Acoustic [2798, 3013, 2308, 3553, 3463, 3479, 3158, 3186, 2830, 3130, 2753, 2998, 2462, 2536, 1238, 3473, 2222, 2343, 3233, 3297]. Acoustical [3452]. Acoustics [2482]. Acquisition [2731, 2681]. Across [2021, 2104, 1734]. actin [3513]. Acting [1203, 1300]. Activation [3318]. Active [67, 807, 2073, 2133, 3135]. Activity [1790, 2965, 3459, 2116, 3408]. Actuators [3307]. Adapting [1544]. Adaption [85]. Adaptive [3000, 2491, 2671]. Addenda [1885]. Addendum [1787, 2405, 140, 973]. Adding [2732]. Addition [1000, 1048]. Additive [3484, 1266, 981]. Adiabatic [2899, 3512, 3365, 2027, 2062]. Adjacency [429, 241, 55, 1328]. Adjacent [1462]. Adjoint [655, 2264, 323]. Adjointness [576, 486]. Adjustment [1737]. Admittance [782]. Admitting [2774]. Adsorption [2461, 2037]. Advection [1720, 2863]. Advection-Diffusion [2863]. Aerodynamic [3542]. Aerodynamics [3070, 3501]. Aeroelastic [3103]. Affect [3041]. Affine [3244, 2434, 1298]. After [3493, 1193]. Again [2719]. Age [3458, 2551, 1706, 1676, 2152, 2833, 3488, 3053, 1477, 2257, 3041, 2415]. Age-Dependent [3458, 1676, 2152, 2833, 1477]. Age-Structured [2551, 3488, 3053, 2415]. Aggregates [476]. Aggregation [2935, 3350]. Aging [1916]. AIDS [2832, 2918, 3236, 3041]. Air [2766]. Airfoil [567, 259, 1514, 3070]. Airfoils [3211]. Airy [242]. Akers [530]. al [3176]. Albert [639]. Albert-Nelson [639]. Algebra [1001, 1772]. Algebraic [129, 2924, 1741, 769, 845, 545, 2579, 121, 1654]. Algebras [1001, 1002, 998, 1003, 1685, 996, 1005, 1006, 994, 1000, 1004, 999, 997, 995]. Algorithm [1532, 1086, 2774, 2197, 1836, 505, 1188, 1380, 2112, 638, 985, 884, 1797, 1963, 1607, 283, 1550, 734, 860, 2689, 2690, 732]. Algorithmic [1573, 1775, 1776]. Algorithms [539, 1534, 1837, 3134, 1214, 1287, 1458, 253, 136, 2693, 2867, 2638, 2070, 3381, 1633, 1048, 214, 2656, 2780, 2168]. Aligned [3549]. Alignment [2574, 2529, 2972, 2292]. Alleles [1502]. Allelic [1671]. Allen [1784, 3283]. Allocating [1914]. Allocation [149]. Allow [878]. Allowed [2747]. Alloy [3467, 3307, 2571]. Alloys [2293, 2302]. Almost [750, 871, 1704, 1630, 2718, 484, 762, 1043, 2012, 380, 266, 768, 2362, 1610, 797]. Along [1918, 2189, 1999, 868]. Alpha [2909]. Alphabetical [781]. also [579]. Alt [3480]. Alternate [71]. Alternated [124]. Alternating [1293, 1715, 2447]. Ambient [1362]. American [1914]. Among [2970, 1543, 2642]. Ampère [3449]. Ampère/Transport [3449]. Amplification [1461]. Amplifiers [3259]. Amplitude [1732, 509, 2110, 2234, 1522, 1346, 3344]. Amplitudes [2470, 2639]. Analogue [1958]. Analogues [2780]. Analysing [3029]. Analysis [3419, 2919, 3144, 2383, 2339, 831, 2465, 760, 3250, 1911, 2523, 3387, 1634, 1740, 2807, 1221, 1258, 3505, 2461, 2158, 2985, 1077, 1222, 3319, 2887, 2550, 3101, 2134, 2663, 2750, 2828, 3086, 3124, 2924, 3100, 2755, 2352, 94, 2801, 3368, 2917, 2216, 2484, 1918, 3044, 117, 3128, 2590, 2870, 3008, 308, 2418, 3555, 3509, 3530, 1667, 1959, 2603, 2577, 3072, 2837, 3414, 2819, 1478, 2886, 3524, 2900, 3492, 2693, 1878, 2979, 1564, 2674, 2278, 2301, 2706, 2194, 2237, 2342, 2792, 2869, 3396, 1127, 3204, 3222, 3028, 2148, 2636, 2000, 2254, 3068, 1846, 2846, 2778, 2664, 590, 2008, 3381, 2528, 2824]. Analysis [2884, 2217, 561, 2749, 2945, 2321, 2597, 1024, 1862, 2940, 588, 2417, 3177, 601, 2001, 1710, 3560, 2711, 2844, 2821, 679, 1100, 2628, 3217, 2629, 2898, 2806, 1618, 1603, 3332, 2582, 274, 3408, 2659, 2499, 2562, 2469, 2742, 3448, 1763, 2279, 1544, 2433, 2285, 2286, 3069, 2146, 2003, 2358, 1931]. Analytic [211, 294, 202, 188, 3399, 228, 3229, 2002, 255, 2367, 520, 498, 2276, 2212, 507, 1289, 687]. Analytical [958, 2272, 1809, 1403, 1404, 3560, 2894, 3462, 3004, 2745]. Analytically [1193, 2497]. Analyzing [352]. Angle [2086]. Angles [2420, 2421, 2623]. Angular [168, 1295, 1703, 793]. Anisotropic [3407, 582, 945, 2221, 2936, 3190, 561, 944, 2688, 284, 3145, 3287]. Annular [2873]. Annulus [2206]. Anomalies [406, 2279, 2375]. Anomalous [3173]. Antenna [3489, 214]. Anti [1624, 426]. Anti-Hermitian [426]. Anti-Plane [1624]. Antiplane [3476]. Antisymmetric [3221, 538]. Any [3424]. Aperture [2709, 2462, 3575]. Apex [3549]. Appearing [1529]. Appell [96]. Appl [1789, 854, 2244, 988, 1721, 2731, 1546, 1512, 2639, 2252, 3043, 667, 1479, 38, 2539, 1885, 253, 1932, 534, 2309, 1585, 1101, 2405, 2044, 144, 2903, 1232, 2450, 3183, 1788, 1156, 792, 896, 1361, 1574, 1677, 2100, 2562, 138, 184]. Applicability [888]. Applicable [1172].

Application [2505, 3018, 2838, 2906, 1833, 1079, 2947, 2199, 519, 3538, 894, 3435, 1188, 2131, 484, 3417, 646, 255, 2819, 35, 1635, 2141, 246, 2580, 769, 1874, 955, 2784, 3397, 2431, 2432, 2791, 1607, 742, 2230, 587, 1075, 550, 182, 463, 214, 736, 388, 3240, 960, 758, 3216, 3465, 800, 828, 1602, 2995, 2240, 1884, 350, 2541, 1964, 2507, 1746]. Applications [2166, 2339, 3325, 1001, 1002, 211, 206, 871, 1558, 1961, 291, 202, 998, 307, 2030, 1411, 1713, 1293, 1557, 57, 203, 1397, 3328, 207, 2958, 568, 3118, 913, 2583, 1003, 204, 208, 2182, 3403, 2009, 3083, 209, 2214, 49, 996, 711, 1005, 1006, 994, 105, 1000, 1486, 288, 210, 163, 632, 2314, 1004, 999, 739, 997, 1806, 205, 1263, 756, 1540, 2259, 1062, 3070, 256, 45, 1680, 3437, 1674, 636, 2033, 222, 740, 428, 3501, 15, 190, 995, 1306, 212, 201, 213, 2544]. Applied [921, 41, 900, 2055, 542]. Apportionment [1555, 1766]. Approach [1192, 322, 130, 69, 897, 1398, 1464, 716, 2328, 2325, 2875, 3532, 2553, 1484, 676, 2741, 2530, 1908, 2135, 3460, 1133, 2218, 1770, 2616, 2446, 1272, 1095, 1891, 2022, 581, 1037, 2710, 205, 3427, 1476, 2993, 3177, 1734, 2894, 507, 513, 1482, 2640, 245, 1775, 1776, 2799, 3295, 3004]. Approaches [3042]. Appropriate [1545]. Approximate [1131, 1413, 1092, 1142, 461, 2112, 3238, 1498, 1279, 900, 2686, 1058, 710, 31, 2913]. Approximately [1413, 941]. Approximating [1464, 334, 1816, 3477, 2344, 23, 3205]. Approximation [460, 3574, 3293, 3321, 2192, 3404, 3037, 2763, 523, 3456, 1145, 3329, 3548, 3386, 3338, 99, 308, 925, 2002, 2038, 2067, 2814, 3417, 1614, 3478, 3260, 1097, 620, 402, 2599, 2707, 3038, 2790, 2836, 3570, 1575, 2165, 3254, 2005, 1786, 1607, 910, 3001, 3306, 1263, 1129, 398, 3474, 2332, 1332, 235, 1839, 819, 3378, 1018, 1618, 3185, 59, 438, 2385, 2456]. Approximation-Theoretic [2599]. Approximations [3561, 2472, 2471, 3182, 157, 2812, 2436, 2135, 2190, 2733, 1970, 955, 862, 133, 3110, 1114, 3500, 1009, 1120, 3491, 586, 2365, 1392]. Aquifers [1965]. Arbitrarily [2838, 1067, 870, 1878]. Arbitrary [1351, 2092, 397, 1201, 1353, 2422, 2977, 377, 393, 3007, 3436, 2498, 32, 126, 868, 1161, 2912, 477]. Arclength [643]. Arcs [1816, 504, 1282, 3267]. Area [3109, 3142]. Arguments [475]. Aris [2244, 2157]. Arises [2381]. Arising [969, 3019, 1423, 2495, 650, 1834, 3173, 1609, 836, 2190, 2123, 3024, 2948, 2061, 3054, 2088, 3337, 2492, 718, 1359, 2904, 1102, 3415, 277, 2672, 2410, 2851, 407, 213, 387]. Arranged [3048]. Array [3553, 1504, 3373, 1206, 1829]. Arrays [248, 2967, 397, 1697]. Arrhenius [1719, 1821, 1832, 1861, 1922]. Arrivals [2199, 2227, 2252, 2296]. Arrow [349]. Arterial [2516]. Artificial [3495]. Ary [553]. Aspects [2394, 1264, 1410, 1662, 1414, 2450, 1573, 3126]. Assisted [2856, 2719]. Associated [703, 1108, 97, 82, 75, 3228, 1741, 1869, 454, 1268, 2288, 1786, 386, 892, 1470, 1020, 2478, 2727, 2810]. Association [369]. Asymmetric [1265, 3392, 315, 3413]. Asymptotes [1690]. Asymptotic [319, 515, 1953, 907, 3489, 2547, 2383, 2339, 233, 2534, 3125, 141, 2501, 2500, 2807, 1969, 244, 3361, 3505, 803, 3159, 1077, 1222, 1103, 2754, 1988, 1399, 890, 851, 1898, 2540, 1642, 1730, 3212, 864, 1704, 3509, 3530, 3077, 2942, 1, 1174, 2067, 3475, 2481, 2600, 2762, 259, 1091, 2436, 3414, 3354, 2741, 1224, 3527, 613, 827, 1880, 633, 1968, 2154, 2716, 1337, 2123, 3542, 6, 1453, 2029, 2194, 2237, 558, 1027, 3243, 3303, 3447, 2306, 2342, 2793, 3136, 3294, 3571, 947, 2405, 443, 2385, 1970, 1539, 1846, 1567, 3469, 3496, 2008, 2649, 2098, 2139, 312, 2943]. Asymptotic [2884, 2082, 2003, 1742, 1786, 189, 3110, 1774, 2321, 2597, 3500, 2670, 563, 1565, 2087, 1157, 2704, 1327, 783, 1129, 1921, 1331, 1580, 1709, 425, 2618, 387, 2332, 1865, 2417, 2672, 1254, 1439, 2001, 1164, 699, 794, 393, 1965, 1641, 1625, 2844, 3181, 1864, 586, 2334, 394, 1208, 580, 489, 719, 896, 917, 1603, 3332, 1771, 1400, 2740, 2437, 15, 190, 1976, 2659, 867, 973, 3358, 3349, 1080, 2742, 630, 1675, 771, 2541, 2090, 3213, 2507]. asymptotic-numerical [3213]. Asymptotical [3326]. Asymptotically [1366, 380, 2208, 2808, 2408, 588, 2036, 3096, 868]. Asymptotics [3576, 3325, 2567, 3438, 3271, 2439, 1825, 2694, 3097, 2103, 3502, 2860, 2267, 1154, 3137, 3340, 3003, 2861, 2913, 3341]. Atmosphere [360, 1514, 1238]. Atmospheric [2782]. Atom [1365]. Atomic [996, 3207]. Attack [3503]. Attenuation [3013, 134]. Attracting [3394]. Attraction [1303, 433]. Attractivity [3470]. Attractor [3506]. Attractors [2956, 1823, 2717]. Attributes [478]. augmented [1995, 2204, 2380]. Augmenting [2295]. Auto [1452]. Auto- [1452]. Autocatalysis [3517]. Automata [3583, 318]. Automatic [2355]. Automorphism [1608, 66]. Automorphisms [1605, 467]. Autonne [584]. Autonomous [881, 431, 838, 988, 430, 611, 2040]. Auxiliary [1301, 271]. Available [787]. Average [1876, 2754, 3152, 2870, 1680, 1864]. Averaged [3371]. Averages [2262, 2386, 2956]. Averaging [2838, 3086, 899, 1700, 1294, 1974, 2864, 1448, 1794, 1885, 2863, 3547, 338, 432, 2193, 2263, 1674]. Avoiding [1177, 55]. Axes [2692]. Axi [1591]. Axi-Symmetric [1591]. Axial [1640, 3527, 273, 3542, 2749]. Axially [2156, 317, 1796, 255, 1879, 409]. Axiom [981]. Axiomatic [551]. Axis [1902]. Axisymmetric [575, 2119, 3564, 125, 2660, 3098, 965, 178, 2175, 20, 2934, 2659, 137, 1995]. Axon [1980, 2373, 2631]. Axons [1930].

Back [290, 2719]. Background [2522].

Backprojection [2689]. Backpropagation [2522]. Backward [2573]. Bacterial [1491, 3299]. Baecklund [1171]. Baker [3166]. Balance [3480, 2953, 3043, 3170]. Balanced [646, 1318, 1606]. Ball [1041]. Balloon [3425]. Balls [3550, 2846]. Banach [721, 727, 148, 491, 147]. Band [1595, 1076, 3577, 3345, 3233, 3297, 2195, 1335, 1162, 1232, 2364, 3578, 941]. Band-gap [3233, 3297]. Band-Limited [2195, 2364, 941]. Band-Pass [1076]. Bandlimit [913]. Bandlimited [1309, 402, 2035, 2242, 2598, 2448]. Bands [3016, 2913]. Bandwidth [1596, 1185]. Banzhaf [1653]. Bar [361]. Bargaining [1139]. Barotropic [2658]. Barriers [1734]. Bars [754, 944, 982]. Bartlett [2005]. Based [2383, 2434, 3210, 297, 1545, 1570, 2638, 2070, 1590, 1048, 3544, 3545, 2168]. Bases [298]. Basic [903, 3486, 2146, 824]. Basins [3494]. Basis [2870, 2538, 2957]. Batch [2199, 2398]. Batchelor [3447]. Bathtub [2488]. Bathymetry [3038]. Bayes [1617]. Bayesian [322]. BCH [612]. Be [2987]. Beach [2439, 2608, 2781, 1547, 3502, 2300, 2001]. Beam [1779, 3139, 2985, 1836, 3105, 2283, 3582, 2026, 2994, 1755, 1788, 2093, 2549]. Beams [2663, 1022, 2027, 3093, 951, 134, 3126]. Bearing [2134]. Bearings [3426]. Bearings-Only [3426]. Bechhofer [2236]. Beck [1752]. Before [1461]. Behavior [3428, 3412, 141, 1969, 3361, 1984, 2398, 2663, 1628, 2162, 1827, 3348, 3307, 1143, 3077, 1174, 901, 651, 966, 3040, 923, 1956, 1224, 1450, 1337, 1623, 3351, 2668, 1453, 2303, 711, 3068, 1567, 3496, 268, 2649, 3565, 3514, 2670, 2087, 2187, 783, 1709, 2747, 3061, 2618, 1367, 470, 699, 683, 2334, 394, 580, 1018, 282, 1855, 852, 853, 2224, 2437, 3358, 3349, 684, 1336, 3179, 2385, 3003]. Behaviors [1730, 1774]. Behaviour [2408, 3065]. Behind [2218]. Belated [802].

Bell [1284]. Bellman [399]. Belousov [1777, 1824, 1226]. Belt [1784]. BEM [2993]. Ben [534]. Ben-Israel [534]. Bénard [1732, 711, 2247]. Bending [1308, 2549]. Beneath [2766]. Benjamin [3579]. Bergman [1624]. Bessel [233, 75, 1973, 2598, 213]. Bessel-Type [2598, 213]. Best [18, 3282, 1338, 473, 1843, 477, 1670, 2337]. best- [2337]. Beta [1933, 556]. Between [1255, 1859, 1265, 1210, 1424, 381, 2127, 2169, 3286, 3445, 593, 2126, 2953, 1155, 3546, 916, 3195, 279, 1371, 1264, 249, 1537, 1270, 3285, 1762, 489, 3332, 2409, 499, 1581, 3406, 1670, 3043, 1113, 3039, 2518, 648, 1739, 2623]. Beyond [2424, 3242]. Bezoutian [801]. Bi [2839]. Bi-Cherenkov [2839]. Bias [2704]. Biased [3372, 3210, 249, 2667]. Bifurcating [3092, 2039, 1890]. Bifurcation [2916, 2356, 2565, 2966, 1847, 3387, 1634, 1322, 2378, 2158, 2581, 1782, 3538, 1145, 3368, 1557, 720, 2357, 1121, 1856, 2126, 2099, 2184, 2310, 2453, 2125, 2253, 1741, 2474, 2937, 2701, 2918, 2046, 2979, 2493, 1760, 1928, 2172, 2586, 1797, 1622, 2254, 3196, 108, 2060, 1469, 3355, 3268, 2177, 2248, 2513, 2816, 2903, 3130, 2749, 2683, 2377, 2322, 1531, 1810, 2584, 2116, 1726, 2128, 2297, 1736, 1756, 3560, 1647, 2083, 2058, 2628, 2454, 2588, 2566, 2720, 2323, 1572, 1975, 2499, 3025, 1931]. Bifurcations [2324, 1616, 2173, 2298, 3229, 2406, 1448, 3537, 2172, 3026, 2288, 1526, 2509, 3095, 2063, 2938, 3106, 2091, 1967, 2658, 2478, 3122, 3302]. Biharmonic [595, 654, 1533, 1561, 1845, 2440]. Bilateral [1845, 1087]. Bilinear [889]. Billet [3336]. Billets [3198]. Billiards [3261]. Bimodal [2158, 2177]. Binary [3467, 1063, 873, 2293, 2302, 816, 1750, 2979, 579, 281, 767, 863, 2699, 369, 2106, 2562, 2469, 2571]. Binomial [1578, 322, 1529, 1848, 1431, 780, 2280, 1497]. Biochemical [1322, 1874, 1469, 601]. Biological [2574, 1265, 3368, 1266, 1838, 2301, 2528, 1264, 2772, 1263, 1262, 2191].

Biology [2529, 1880, 1493]. biomembranes [1897, 1896, 1964]. Biorthogonal [1533, 1561, 3230]. Bipartite [457, 287]. Biperiodic [3321]. Bipolar [2586, 573]. Birth [2419, 2032, 2299, 2208, 2336, 2036, 1768]. Bistable [2592, 2041]. Bistatic [2651]. Bivariate [1789, 1769, 1614, 722]. Black [2833, 2606]. Black-Footed [2833]. Black-Oil [2606]. Blasius [222]. Blasting [2382]. Blobs [620]. Bloch [3374]. Block [934, 1695, 2010, 646, 647, 1029, 1513, 45, 1762]. Block-Diagonalization [45]. Block-Toeplitz [1695]. Blocking [765, 403, 2592]. Blocky [3281]. Blow [2424, 3085]. Blow-Up [3085, 2424]. Blowing [1673]. Blowoff [497]. Blowup [1911, 3009, 3350]. Bluff [1209]. Blunt [3398]. Blunt-Nosed [3398]. Blurred [3281, 2404, 2539]. Blurring [3483]. Boards [1431]. BOD [1476]. Bodies [2192, 1986, 1844, 796]. Body [515, 1927, 2072, 2156, 2416, 2907, 2524, 3528, 575, 979, 314, 304, 1985, 1091, 1588, 1796, 2583, 659, 191, 1879, 1968, 3238, 2218, 2188, 2645, 1203, 182, 1129, 3189, 3398, 1111, 1220, 2543, 3213]. Body-body [2543]. Body-Wall [2524]. Bogdanov [2940]. Bogoliubov [41]. Bogoljubov [603]. Boiling [2145]. Boltzmann [2927, 1296, 1604, 78, 2636, 2768, 2370, 542]. Bonded [30, 2440]. Bonferroni [1338, 1843]. Bonhoeffer [2582]. Boolean [2162, 359, 695, 194, 924, 529]. Bordered [1257, 1379, 243]. Born [2522]. Borne [1425]. Bose [365]. Bose-Chaudhuri-Hocquenghem [365]. Both [2573]. Bouncing [1041]. Bound [1341, 1562, 472, 1455, 3048, 2005, 1767, 2788, 3347, 941]. Boundaries [2841, 1428, 168, 2495, 3286, 3445, 834, 2021, 2104, 3375, 2697]. Boundary [1255, 1436, 3539, 2963, 3561, 2049, 146, 2494, 1996, 462, 976, 939, 1269, 3361, 3037, 240, 1423, 2291, 494, 1708, 674, 839, 1092, 1312, 1103, 2750, 2925, 890, 851, 2221, 97, 694, 650, 2702, 157, 2570, 1351, 2856, 593, 864, 3062, 125, 666, 1676, 1834, 3455, 2245, 3173, 356, 3107, 723, 2558, 1168, 1484, 2568, 153, 2458, 1337, 2123, 2423, 1144, 2771, 2948, 624, 964, 713, 965, 3554, 705, 497, 987, 1863, 2706, 849, 389, 1942, 954, 1252, 1770, 348, 443, 3457, 74, 2000, 3068, 1383, 2317, 2564, 590, 749, 648, 2171, 511, 1036, 528, 2230, 1813, 3258, 1213]. Boundary [1845, 421, 2462, 1662, 2657, 2430, 968, 3147, 193, 2643, 3174, 1206, 92, 1099, 1187, 707, 88, 1840, 2389, 344, 1112, 2566, 2813, 2327, 2109, 1231, 436, 3462, 407, 3187, 2878, 1137, 2365, 2366, 2062, 1808, 1246, 2047, 868, 793, 1080, 918, 630, 2912, 323, 486, 1182, 2777, 2541, 2828, 2555, 2556, 3154, 3344, 1937, 3525, 2285, 2286, 3069, 2544, 2507]. Boundary-Layer [1269, 443, 511, 1036, 1187]. Boundary-Value [1436, 2291, 2000, 3068]. Bounded [1160, 305, 16, 948, 3433, 721, 394, 3549, 1161, 2912]. Boundedness [2500, 430, 843, 1376, 91]. Bounding [1305, 1615, 199, 2431, 2432]. Bounds [3430, 844, 1803, 2457, 2610, 1034, 1141, 804, 3018, 871, 939, 615, 1160, 2307, 1119, 612, 1007, 2403, 895, 925, 406, 2074, 912, 113, 64, 2089, 753, 2316, 1338, 524, 1066, 1567, 1372, 3075, 3320, 3081, 936, 1026, 1213, 1845, 1587, 2280, 2215, 414, 1656, 177, 730, 614, 764, 88, 216, 983, 33, 1200, 1618, 1843, 1408, 1402, 32, 102, 2511]. Boussinesq [443, 3441]. Box [2229, 3171]. Bracing [1727, 1728]. Brain [3583, 3581]. Brain-Cell [3583]. Branches [1928]. Branching [3454, 2560, 2630, 1191, 1874, 1875, 2634, 3540, 197]. Branching/Chain [2268]. Breakdown [2843, 1834, 1230]. Breaking [2937, 2268, 2620]. Breakup [3584, 3491]. Breather [2914, 2068]. Breathers [3537, 3066]. Bridge [2715]. Bridges [3419]. Broadcast [1860]. Broadwell [3171].

Broken [3080]. Brownian [2115, 1620, 1948, 2304, 2314, 1446]. Brussellator [2140]. Bubble [2872, 3283]. Bubbles [3572]. Bubbly [3245, 3246]. Buckled [2698, 1740, 932]. Buckley [2713, 2853, 3366]. Buckling [146, 1085, 901, 2741, 106, 790, 2175, 1755, 1788, 3487]. Bueckner [38]. Buffer [2398, 891, 1161, 3179]. Buffered [3437]. Building [1568, 2284]. Bulk [2610, 2084, 2444, 3126]. Bullets [2753, 2998]. Bump [1544]. Bunched [2027]. Buoyant [2778]. Burgers [2121, 3156, 2696, 3273]. Burner [2147, 1917]. Burner-Stabilized [1917]. Burning [2019]. Bursting [3364, 1718, 2326, 2965, 3095, 3459, 3015, 2851, 2582, 3408]. Business [3056].

Cahn [3022, 3051, 3283]. Calcium [3309, 3437]. Calculating [3323, 2186, 2661]. Calculation [3018, 716, 592, 1005, 1006, 2044, 1990, 2757, 1153]. Calculations [2113, 1955, 451]. Calculus [1013, 924]. Campbell [1962]. Can [2253]. Cancer [3052]. Cannibalization [644]. Cannibalize [1972]. Canonical [861, 1110, 2009, 997, 398, 879]. Cantor [3261]. Cap [2494, 2318]. Capacitance [1988]. Capacity [3270, 1022, 2693, 1631, 2691, 3396, 368, 3002, 3340, 3030, 3341]. Capillarity [3251]. Capillary [2157, 3140, 3333, 2244]. Caps [146, 2175, 1349]. Capture [3208, 2838, 2074]. Caratheodory [1028]. Carburization [3098]. Cardinality [738]. Carleman [2229]. Carlitz [1319]. Carlo [307, 1702, 62, 35, 2385, 144]. Carrier [2909, 1425]. Carrier-Borne [1425]. Carrying [1847, 1022]. Cascade [885]. Cascading [2647, 2099, 2063]. Case [1911, 3416, 1258, 902, 2277, 2484, 2927, 2692, 367, 2626, 2374, 3088, 3087, 3439, 3245, 3560, 2295, 973, 1763]. Cases [2534].

Casimir [3515]. Casing [30]. Catalyst [2943, 2514]. Catalysts [2881]. Catalytic [3094, 1652, 1867, 638, 1893, 3061]. Cauchy [637, 211, 1558, 547, 2381, 2211, 1563, 2865, 944, 177, 730, 707, 395]. Causal [3282]. Causality [2464, 3362]. Caused [2881, 2518, 2682]. Caustic [1858, 2163]. Caustics [2662, 3159, 2299]. Cave [447]. Cavity [3322]. Cell [2383, 3042, 2789, 3556, 3480, 3583, 3210, 2953, 2351, 2275, 2242, 1331, 2014, 2573, 2538, 3043]. Cells [760, 3513, 2413, 2965, 3408, 542]. Cellular [2647, 2893, 3520, 3583, 2979, 2177, 2205, 1828, 2788, 1710, 3052, 2894, 3298]. Censoring [473]. Census [3053]. Center [2601, 2074, 2542, 2733, 2589, 3355]. Center-Unstable [2601]. Centerline [949]. Centers [2357, 1775]. Central [2016, 697, 102]. Centralized [260]. Centre [2765, 2206]. Ceramic [2739]. Ceramics [3322]. Certain [1578, 1532, 244, 130, 392, 494, 1063, 2419, 1621, 1388, 1293, 1472, 1128, 1193, 192, 297, 243, 2733, 2214, 6, 313, 752, 2520, 1172, 379, 2743, 142, 555, 2212, 580, 1632, 1200, 917, 1400, 60, 1336, 1182]. Chain [642, 508, 1505, 2268, 2664, 681, 2968, 3317, 3540, 710, 3462, 3490]. Chain-Branching [2268, 3540]. Chain-Branching/Chain-Breaking [2268]. Chain-Breaking [2268]. Chains [2271, 3338, 2081, 3079, 865, 2961, 3243, 2736, 2850, 3469, 352]. Chambered [1742]. chance [1405]. chance-constrained [1405]. Change [2095, 939, 2832, 1708, 429]. Changes [2815, 1022, 3351]. Changing [2702, 3306, 3017]. Channel [1413, 887, 1077, 2801, 1619, 2194, 2237, 1343, 2697, 3065, 393, 2410, 1601, 2105, 2742]. Channels [3270, 2951, 2952, 2933, 1193, 1631, 2177, 2248, 2765, 368, 3356, 159, 3191]. Chaos [2695, 2700, 3064, 3227, 2969, 2160, 2406, 3518, 2121, 3433, 3431]. Chaotic [3117, 3107, 1794, 1885, 2668, 2073, 2408, 2040, 2851, 2978]. Character [343].

Characteristic [2355, 1502, 1609, 83, 658, 2580, 2021, 2104, 354, 474]. Characteristics [2288, 293, 3536]. Characterization [1789, 1769, 3151, 1309, 2435, 1848, 2491, 2211, 1337, 690, 1653, 859, 1602, 1983, 521, 1900, 1901]. Characterizations [935, 2401, 1433, 1654]. Characterize [121]. Characterized [1301]. Characterizing [2477]. Charge [2687, 979, 1395, 1521, 2909, 1574, 1204]. Charged [1169, 3371, 2722, 1897, 1896, 1964]. Charges [3169]. Chaudhuri [365]. Chebyshev [133]. Chemical [2519, 560, 1103, 650, 1145, 1196, 1867, 2298, 2882, 3414, 1720, 2443, 1757, 1729, 1815, 1102, 283, 121, 2209, 451, 2444]. Chemicals [962, 1094, 1095]. Chemistry [2835]. Chemoattractant [3480]. Chemostat [2250, 2269, 3077, 2612, 2900, 3024, 2124, 2669, 3496, 3531, 3552, 3390, 3585, 1913, 2883, 3358, 3349]. Chemostat-Related [2669]. Chemostat-Type [3390, 3585]. Chemotactic [2953, 3503, 3043]. Chemotaxis [1491, 3299]. Cherenkov [2839]. Chi [439, 2005]. Chi-Square [439, 2005]. Chiral [3428, 3561]. Chirality [3574, 3561]. Choice [1266, 1542, 1236]. Cholesteric [1030]. Christoffel [565]. Chromatic [669, 1165]. Chromatography [2876, 3054]. Circle [2211]. Circuit [1416, 2929]. Circuits [272, 2641, 2586]. Circulant [1135, 1318]. Circulants [967]. Circular [119, 3019, 262, 89, 1085, 1988, 304, 1150, 1308, 2510, 2521, 106, 623, 31, 1282, 2828]. Circulation [3235, 1273]. Citizen [1590]. Clamped [2828]. Class [1317, 269, 2465, 2596, 663, 1283, 26, 1979, 1160, 2551, 2250, 1309, 2085, 2483, 946, 1076, 3014, 2917, 415, 612, 1178, 2481, 2600, 62, 237, 2895, 1609, 1458, 1275, 827, 278, 1382, 1348, 3172, 2716, 2123, 3178, 2948, 2733, 3450, 2061, 3162, 806, 2264, 725, 862, 1357, 115, 975, 563, 2808, 301, 1327, 1171, 860, 2212, 1261, 1107, 2334, 957, 1318, 1766, 543, 2109, 222, 2989, 3408, 1137, 2479, 691, 1463, 2541]. Classes [430, 1239, 3110, 2559, 354, 368, 101, 159]. Classical [3533, 3411, 2586, 701, 3289, 1220]. Classification [833, 2209, 1517, 2525]. Classifier [2531]. Cliff [2382]. Climate [2125]. Climatic [2095]. Clinical [2236]. Clique [844, 1048]. Clock [2744]. Clock-Regulated [2744]. Close [622]. Closed [3125, 2807, 2767, 617, 3236, 2644, 2497, 2321, 2449, 2559, 2645, 3431, 3383, 821, 3129]. Closed-Form [2767, 2645]. Closely [3388, 1874]. Closure [3482, 2980]. Cluster [1564, 1745, 1900, 1901]. Clustering [3149]. Clusters [832]. Cnoidal [2219]. Coalition [1283]. Coalitional [1405]. Coalitions [403]. coating [3300]. Cochlea [1326, 1825, 1325, 2270, 2475]. Cochran [1361, 1211]. Cocycles [1324]. Code [1189, 372]. Codes [269, 887, 770, 612, 1608, 516, 1661, 781, 1631, 986, 1619, 365, 579, 1434, 1767, 187, 1762, 920, 13, 140]. Codimension [2124]. Coding [1202, 1374, 256]. Coefficient [2609, 2142, 2754, 1704, 1945, 3524, 2962, 3481, 2805, 1497]. Coefficients [1592, 1634, 26, 1411, 942, 635, 61, 2038, 2895, 2517, 613, 1850, 1932, 2407, 805, 2362, 3420, 1162, 1232, 3091, 1905, 1134, 331, 70, 251, 687]. Coexistence [3359, 3563, 2233, 3391, 2612, 2124, 2650, 1913]. Coexisting [2974]. Coherence [2153]. Coherent [847, 564, 1956, 2720]. Cold [2458]. Coleman [1653]. Collapse [3350]. Collected [1545]. Collective [3136, 3294, 2758]. College [1590]. Collision [2049, 1296, 3519]. Collisionless [3245]. Collisions [1724, 1703]. Colorations [1251]. Colored [2652, 2489, 411]. Coloring [495, 1282]. Column [923, 3187]. Columns [2698]. Combat [1725]. Combinations [1909, 2353, 1901]. Combinatorial [478, 602, 458, 1684, 333, 423, 2209]. Combined [1859, 106, 3251]. Combines [2760]. Combustible [3535, 2396, 1853, 2059]. Combustion [2700, 3064, 2359, 2291, 2814, 2892, 1924, 2513, 2816, 2903, 3130, 1665, 2042, 3063, 2473, 2485, 1817, 3418, 2981, 1907, 2834, 2891]. Combustion-Driven [3130]. Comma [372]. Comma-Free [372]. Comments [38, 1629, 1812]. Committees [1959]. Commodity [374]. Common [2709, 3338, 532, 2015, 1402]. Communication [660, 1860, 1631, 2972, 1285]. Communications [69, 3580]. Commutation [1321]. Commute [219]. Commuting [750, 2031, 916]. Compact [484, 2477]. Compaction [3494]. Compactness [1577]. Comparing [2228]. Comparison [3042, 755, 937, 566, 903, 784, 2552, 2190, 1250, 632, 3031, 1123, 1201, 571, 2665, 1099, 1674, 2191]. Compartmental [2515]. Competing [3392, 2395, 1730, 2277, 2110, 1678, 1623, 2723, 2361, 2025]. Competition [3504, 3359, 3123, 2233, 2611, 3119, 3077, 2612, 1467, 1954, 3024, 2317, 3546, 1527, 1264, 3531, 3390, 3585, 1967, 2852, 3108, 2883, 3358, 1739]. Competition-Diffusion [3119]. Competition-Mediated [3359]. Competitive [3252, 3563, 1717, 2670, 2650, 1060, 1488, 1913, 3393]. Competitors [2900]. Complement [1061, 1052, 1216]. Complementarity [1429, 1385, 859]. Complementary [669, 583, 1165, 1954, 1015, 910, 860]. Complete [1465, 452, 672, 495, 595, 1508, 1016, 3, 1494, 287, 1336, 2057]. Completeness [2859]. Completions [1822]. Complex [2424, 2097, 1108, 2131, 2202, 870, 2701, 2114, 2779, 2827, 330, 1797, 2840, 2860, 442, 3362, 1468, 451, 2711, 852, 853, 438]. Complex-Symmetric [1108].

Complex-Valued [2840]. Complexes [121]. Complexity [1316, 1472, 1509, 1596, 2316, 3487]. Compliance [1026]. Compliant [2697]. Component [3018]. Components [832, 3304]. Composite [3018, 3416, 971, 2486, 1021, 2132, 967, 414, 9, 3049, 3183]. Composites [2457, 3569, 2714, 3075, 3320, 1057, 2680, 2829]. Compound [1096]. Compounding [935]. Compressed [3527]. Compressible [2698, 2134, 170, 882, 2823, 2974, 849, 3199, 2111]. Compression [3133, 2062]. Compressional [1072, 3247]. Compressive [2843]. Compressor [2749]. Computable [3406]. Computation [3092, 3010, 1978, 3104, 1838, 218, 2264, 2128, 2297, 1109, 1572]. Computational [1001, 1002, 998, 3509, 1003, 996, 2865, 1005, 1006, 994, 1000, 1004, 999, 997, 995]. Computationally [2638]. Computations [3272, 112, 3185, 2934, 2375]. Computed [2770, 2070, 2931, 2168]. Computer [772, 781, 2719, 108, 1229]. Computer-Implemented [772]. Computerized [2137]. Computers [1005, 1006]. Computing [1795, 1997, 657, 751, 1509, 1468, 735]. Concave [176, 532]. Concentrated [3527, 3048]. Concentrating [3209]. Concentric [1581]. Concept [417, 3410, 2640]. Concepts [1283]. Concerning [1310, 2002, 1132, 1064, 1714, 1376, 144, 377, 103]. Conclusions [1326]. Condensation [2370]. Condensed [2700, 2359, 2892, 2082, 3063, 2834, 2891]. Condensed-Phase [2892]. Condition [3539, 2963, 549, 3173, 3034, 1572, 1693, 3525, 2544]. Conditional [1702, 3030]. Conditionally [3261, 697, 1445, 1130, 1541]. Conditioned [2198, 2086, 2449, 2559]. Conditioning [2572]. Conditions [408, 1141, 1996, 2360, 453, 3558, 97, 444, 593, 864, 483, 3062, 261, 762, 2558, 3535, 1505, 1019, 2708, 3554, 2096, 2080, 389, 1942, 768, 3457, 1714, 2317, 1943, 3158, 175, 1841, 1124, 271, 2462, 1445, 1360, 3357, 1835, 3462, 2037, 2149, 90, 369, 196, 2912, 850, 323, 771, 2777, 2828, 3154]. Condorcet [1542, 1650]. Conductance [3047]. Conductances [1868]. Conducting [119, 979, 3388, 3286, 3445, 1067, 191, 3320, 605]. Conduction [1683, 2360, 942, 2731, 2681, 1315, 2143, 3385, 2230, 974, 1964]. Conductive [2769, 2963, 3505, 3257]. Conductivities [3576, 2518]. Conductivity [3430, 3435, 2896, 2486, 3554, 3550, 2680, 2431, 2432]. Conductor [971, 259]. Conduits [2857]. Cone [2596, 972, 2283, 1577, 3582, 1521, 825, 2093]. Cone-Beam [2093]. Cone-Compactness [1577]. Cone-Convex [972]. Cone-Convexity [1577]. Cone-Efficiency [1577]. Cones [860, 3007, 717, 355]. Confidence [1887]. Configurations [423, 1518]. Confined [1863, 1735, 1937]. Congressional [1766]. Congruence [1110, 1175, 879]. Congruence-Normal [1110]. Conic [623]. Conical [1582, 754, 20, 1292, 3549]. Conjecture [2619, 776, 472, 1926, 1073, 481, 2028]. Conjugate [1141, 453, 485, 848, 1159, 3332, 879]. Conjugate-Normal [879]. Connected [248, 299, 694, 2506, 3581, 66, 2912]. Connectedness [1687, 1472, 28]. Connecting [2210, 1044, 3503, 796]. Connection [381, 2159, 2860, 1762]. Connections [3010, 1537]. Connectivity [1611, 1050, 548, 1068]. Consensus [1692]. Consequences [3546, 2479]. Conservation [2811, 2368, 3101, 3248, 2926, 3172, 2944, 3215, 3162, 3035, 3283, 3296]. Conservative [3395, 1691, 3529]. Conserved [3009, 3085]. Considerations [811, 2282]. Considered [1685]. Consistency [1049, 2278]. Consistent [1234, 698, 1650]. Constant [456, 3018, 2142, 1411, 3192, 2895, 2145, 1863, 1786, 2132, 296, 3317, 859, 13, 140, 3179, 1814, 1937]. Constant-Sum [296]. Constants [1666]. Constitution [2260]. Constitutive [1017]. Constrained [3558, 261, 670, 175, 587, 960, 518, 1153, 2675, 1693, 2433, 1405]. Constraint [560, 2050, 1301]. Constraints [58, 252, 953, 1215, 176, 434, 1986, 234, 1026, 2675]. Construct [519]. Construction [465, 3555, 145, 646, 845, 1185, 1523]. Constructive [1370, 2357, 1116]. Consumption [2396, 1060]. Consumptions [1488]. Contact [1740, 1624, 2910, 901, 932, 3385, 3444, 3285, 792, 678, 2623]. Contacts [2970]. Contained [192]. Container [1732, 173]. Containers [905]. Containing [262, 89, 97, 2152, 2680, 344, 2498, 2238]. Containment [2568, 1550]. Contaminant [3271, 2765, 3191]. Contaminated [2060]. Contamination [2847]. Contemporaneously [1881]. Content [584, 891]. Context [2191]. Continuation [211, 3229, 3177, 2212, 2640]. Continued [706]. Continuity [1195, 1074, 1454, 1159, 373, 740, 378]. Continuous [1192, 2173, 2501, 2500, 546, 2435, 2986, 635, 2298, 3248, 3348, 2002, 2692, 1705, 502, 620, 2823, 2363, 1467, 1954, 1252, 3381, 1841, 362, 2938, 235, 989, 1156, 3378, 3108, 609, 2498, 231, 1638, 2504]. Continuous-Time [1192]. Continuum [2967, 2935, 3443, 1025, 1009, 3459]. Contour [161]. Contractions [2375]. Contractive [789]. Contrast [3404, 3505, 3411, 3242]. Contributions [1992]. Control [1644, 871, 1595, 998, 1620, 2484, 2310, 2453, 3367, 2152, 2112, 2088, 1710, 3052, 428, 2224, 2200]. Controllability [1031]. Controlled [2880, 3438, 2626, 2492, 2723, 1446, 3560, 2555, 2556, 2556]. Controversy [2976, 2394]. Convection [1859, 1732, 3085, 3286, 3445, 3073, 2186, 2778, 2512, 2290, 2509, 2048, 2862, 3131, 3189, 1511, 2369, 3216, 2444, 1288, 1586]. Convection-Diffusion [2186, 3131, 3216]. Convective [2049, 3424, 1782, 1530, 2064]. Convergence [2166, 1811, 3484, 157, 3454, 747, 3338, 2896, 549, 1243, 2164, 1043, 1340, 2542, 56, 2693, 3170, 1533, 1561, 480, 1915, 725, 2661, 3471, 3394, 3575, 154, 1244, 992, 1680, 1783, 3171]. Convergent [2542, 3197]. Converging [2526]. Converse [761, 364]. Conversion [2759]. Converter [3094, 282]. Converters [3061]. Convex [1283, 304, 972, 1276, 165, 334, 626, 472, 620, 1816, 1027, 1240, 342, 1124, 2675, 1618, 12, 1306, 1901]. Convexity [25, 966, 1577, 1440]. Convexity-Preserving [25]. Convolution [1836, 1529, 2031, 2983, 442, 664, 410, 39, 150]. Convolutional [1608, 1762]. Convolutions [1466]. Cooling [956]. Cooperating [2334]. Cooperative [1248, 169, 1274]. Coordinate [1918]. Coordinates [517, 302, 3343, 1939, 2047, 573]. Corduneanu [513]. Core [540, 1640, 1985, 2950, 3174, 3058]. Cored [3199]. Cores [1463]. Corner [870, 1583, 1395, 1574]. Corners [3046]. Correcting [1943, 187]. Correction [2135]. Corrections [3049, 3183]. corrector [3254]. Correlated [1753, 1716, 2746, 1881]. Correlation [1910, 725, 281, 1881]. Correlations [1452]. Correspondence [758]. Corrigendum [2731, 3477]. Corrosion [3451]. Coset [83]. Cosserat [1437]. Cost [904, 1579, 2972, 1290, 1474]. Costas [2588]. Costs [1914, 1579, 677, 2088, 419, 679, 1290]. Couette [435]. Countable [1475]. Counter [1371, 639]. Counter-Rotating [1371]. Counterexample [599, 1073]. Counterflow [3045, 3418]. Counterflowing [3332]. Counterion [2370]. Counterpropagating [3163]. Counterrotating [3058]. Counting [673, 689, 1202, 1482]. Coupled [3449, 3221, 2735, 2815, 2663, 2298, 2201, 2326, 2684, 2967, 3079, 3040, 2012, 3532, 836, 2535, 2413, 2736, 2256, 2664, 3276, 2683, 2772, 1757, 1815, 1088, 2938, 601, 3290, 2058, 2315, 3266, 3291]. Couplers [2505]. Coupling [2928, 3519, 1729, 2222]. Court [1914]. Covariance [452, 1737, 1759, 246, 857, 1412, 1751, 477]. Covariances [929]. Coverage [1958, 2655]. Covering [904, 93, 1767, 490]. Coverings [617]. Covers [1569, 738, 318]. Crab [1544]. Crack [3218, 3401, 3521, 1583, 1952, 1807, 344]. Cracked [1100]. Cracks [572, 1203, 2456]. Cramer [473]. created [2227, 2252, 2296]. Creation [2443]. Creep [2576, 1023, 1540]. Creeping [933]. Criteria [471, 2346, 1750, 806, 779, 3394, 717, 1197, 1299]. Criterion [1147, 390, 876]. Critical [1501, 1302, 1044, 2104, 2842, 1587, 1010, 1756, 3211, 2678, 990, 3357, 2888, 740]. Criticality [2613, 1922, 2497, 1451, 2511]. Critically [3288, 3126]. Cross [1413, 2470, 2639, 1759, 2886, 1521, 393, 3007, 1346, 1452, 3487]. Cross-Correlations [1452]. Cross-Covariance [1759]. Cross-Section [1521]. Cross-Validation [2886]. Crossed [134]. Crossing [2776]. Crossings [3, 1494, 2579, 103]. Cryptography [1293]. Crystal [2946, 2875, 3072, 1134, 1231]. Crystalline [3214, 2487, 3308]. Crystallographic [2714]. Crystals [3577, 363, 1030, 3486, 2185, 3297]. CSTR [3560]. Cube [248, 2319, 1496, 1684]. Cubic [2329, 800, 2091]. Culture [3108]. Cultures [1467, 1954]. Cumulative [803, 1706]. Curl [739]. Current [2985, 3128, 3105, 3323, 2929, 2770, 2818, 3373, 3451, 2931]. Current-Limiting [2929]. Currents [2318, 3169]. Curvature [2648, 252, 2936, 3039, 643, 3295]. curvature-dependent [3295]. Curve [2283, 3471, 2569]. Curved [1640, 3371, 2114, 692, 2176, 1709, 915, 3356, 1093].

Curves [2350, 2488, 800, 3432]. Cusp [2312]. Cusped [2299]. Cusps [2357]. Customers [893]. Cut [1296, 2819]. Cut-Grow [2819]. Cut-Off [1296]. Cutoff [1669]. Cuts [861]. Cutting [903, 1458]. CVD [2959]. Cycle [2011, 2383, 1826, 2216, 2275, 1729, 2503]. Cycles [2669, 1459, 1800, 1324, 3394, 1123]. Cyclic [881, 269, 612, 1593, 579, 2492, 2989, 2224, 2737]. Cyclical [3439]. Cylinder [258, 2411, 582, 1150, 1378, 2858, 3247, 2585, 645, 2616, 785, 30, 344, 324, 476, 1771, 137]. Cylinders [3388, 1209, 2665, 126, 1581]. Cylindrical [1732, 360, 3447, 561, 3491, 1331, 31, 1377, 3213]. Cylindrically [3555]. Cytology [1528].

D [38, 3314]. D. [37, 605]. Dam [1904]. Damage [3556]. Damped [980, 1126, 631, 2508, 3288, 1333, 1071, 3240, 400, 2720]. Damped-Driven [2720]. Damping [937, 500, 2754, 2795, 3507, 559, 591, 1243, 335, 1420]. Darcy [3492]. Data [2505, 227, 2537, 3281, 3455, 1545, 1613, 2823, 3562, 2792, 3498, 3028, 2597, 3174, 3053, 707, 1285, 1185, 3314]. Data-Handling [2792, 2597]. Davidon [434, 309]. Davies [3185]. DC [3372, 812]. Deadlines [2991]. Death [2419, 2684, 2208, 2336, 2036, 3053, 2883]. Decay [1791, 2271, 2335, 2427, 2754, 2795, 1635, 2229, 3517, 2998, 598, 1461]. Decaying [1640]. Decentral [3276]. Decimation [763]. Decision [82, 1475, 740, 853]. Deck [2102]. Decomposition [3484, 143, 2854, 3374, 3532, 1379, 1330, 1851, 1274, 287, 691, 1675]. Decompositions [3330, 1216]. Deconvolution [2307, 2265, 2435]. Deconvolutions [3478]. Decoupling [2575]. decreasing [2385]. Deducing [1430, 1850, 1932]. Deep [1904, 3089]. Defect [3463, 2915]. Defects [2964, 2627]. Defence [2499]. Deficient [3486]. Defined [998, 3229, 181, 742]. Defining [483].

Definite [1695, 1630, 626, 884, 1426, 640, 1679, 774]. Definiteness [408]. Definition [2211, 71]. Definitions [210]. Deflagration [2848, 3146]. deflagrations [2090]. Deflections [829]. Deformation [3033, 1944]. Deformations [89, 3509, 868, 3296]. Deformed [1712]. Deforming [773]. Degenerate [676, 2513, 2087, 1055]. Degree [3103, 1330, 1487, 3196, 541, 1485, 1814]. Degrees [1050]. Delay [2919, 1811, 2565, 3241, 2501, 2500, 2178, 1160, 1686, 3124, 3123, 2007, 3235, 1557, 3076, 677, 807, 3440, 875, 3547, 1714, 1376, 2205, 352, 1197, 3358, 1638]. Delay-Differential [2178, 1197]. Delay-Induced [3241]. Delayed [3364, 3077, 3222, 3470, 3122]. Delays [1278, 673, 2346, 2612, 3547, 2734, 3349]. Delta [965, 1157, 1400]. Demand [1753, 1802, 1460]. Demands [419, 2227, 2252, 2296]. Demodulating [1193]. Demographic [2832]. Denoted [661]. Dense [870, 3501]. Densities [1789, 1769, 3323, 2653, 1698, 115, 1680, 1864, 898, 3304]. Density [1722, 2982, 1362, 2164, 2521, 1994, 68, 741, 2816, 2903, 1786, 134, 3273, 410, 2868, 1231, 2415, 2238]. Density-Dependent [2415]. Density-Potential [1722]. Departure [1201]. Dependence [2411, 546, 3128, 2823, 722, 1252, 3039, 950]. Dependencies [3546]. Dependent [2919, 3458, 1803, 141, 3159, 143, 2755, 1443, 3274, 1676, 1350, 3310, 3034, 2152, 1450, 2674, 2575, 2833, 2342, 2354, 3136, 3294, 2616, 1597, 3420, 361, 2857, 2734, 1723, 1247, 1353, 569, 3169, 1768, 2995, 636, 1477, 2257, 3041, 2415, 2191, 1179, 1906, 3179, 3275, 3154, 3295]. Depending [239]. Deposition [3414, 578, 2813]. Depth [1504]. Depths [3151]. Derivable [301]. Derivation [551, 1413, 2889, 2599, 1155, 1742, 1786, 2753, 3472, 3421]. Derivative [3338, 2558, 389, 1942]. Derivatives [921, 1131, 250, 405, 1994, 219, 1565, 767, 550, 706, 1680, 1864]. Derived [2403, 2685, 2161]. Desalination [375]. Described [3340, 3341]. Describing [2368, 2250, 2245, 1280, 2644, 2020]. Description [2765, 2894, 3226]. Design [478, 1458, 1504, 1104, 1010, 214, 2531, 332, 3266, 1763]. Designated [681]. Designing [1471]. Designs [2663, 1571, 646, 1019]. Desorption [3530]. Detecting [682]. Detection [1900, 1901, 1424, 3423, 1827, 2387, 1698, 1048, 468, 2955, 2479]. Determinable [721, 727, 159]. Determinancy [112]. Determinant [549, 1285]. Determinantal [1473]. Determinants [615, 614]. Determination [188, 2265, 1920, 3583, 49, 3420, 852, 853]. Determined [2692, 1550]. Determining [3425, 942, 2004, 1503, 2729, 1071, 155, 510, 2805, 3053, 1346]. Deterministic [1649, 1721, 2927, 959, 1178, 1744, 1445, 358, 2334, 1130, 1541]. Detonation [2648, 2942, 3146, 3039, 2129, 3357, 3318]. Detonations [2796, 2008, 2090]. Development [2662, 1538, 2312, 2300, 1877, 1808, 2642]. Developments [1307]. Deviating [475]. Deviation [2806]. Deviations [2306, 2165, 3469, 1444]. Device [2572, 2217, 2171, 2704, 7, 2628, 2629, 2787]. Devices [2929, 3074, 3417, 3266]. DFFE [3088]. Diagonal [225, 741, 115, 1440]. Diagonalizable [3331]. Diagonalization [246, 3114, 45, 1175]. Diagonalizing [2830]. Diagram [1332]. Diagrams [881, 349, 984]. Diameter [1202]. Diaphragms [288]. Dielectric [1169, 2416, 2907, 3018, 2071, 89, 3463, 2858, 3526, 2132, 951, 1300, 100, 3233, 3297, 3525]. Difeomorphisms [1572]. Difference [1271, 441, 729, 3338, 2162, 676, 2207, 1442, 603, 2518, 1915, 2000, 3068, 3547, 2364, 2654, 1522, 386, 1883, 3503, 924, 2375, 2285, 2286, 3069]. Differences [1283, 1713]. Different [2013, 3496, 100]. Differentiability [723, 1454, 717]. Differential [2609, 3010, 980, 1126, 431, 911, 976, 2142, 2775, 2774, 500, 2178, 453, 291, 2773, 2013, 2261, 3009, 3085, 777, 1278, 1686, 135, 1411, 188, 2735, 537, 1312, 519, 874, 902, 851, 2633, 1621, 635, 2917, 415, 1704, 483, 1121, 1031, 61, 535, 1128, 2038, 761, 723, 3367, 1700, 1717, 237, 265, 1344, 1484, 380, 3260, 2031, 413, 1049, 613, 3524, 340, 2480, 2983, 41, 2733, 624, 964, 2141, 6, 733, 611, 753, 558, 1314, 266, 1570, 1104, 843, 1172, 655, 806, 1797, 2043, 746, 2000, 3068, 218, 379, 2088, 1552, 2890, 1259, 862, 1357, 312, 749, 335, 779, 1823]. Differential [1249, 391, 1162, 1232, 271, 96, 2734, 3176, 386, 421, 2393, 555, 618, 908, 1476, 2645, 756, 498, 1102, 2040, 1883, 2584, 2259, 848, 2128, 1439, 475, 778, 1099, 470, 990, 819, 562, 45, 216, 957, 394, 622, 70, 745, 580, 1725, 1415, 1197, 917, 1292, 229, 1346, 3370, 1449, 1677, 436, 577, 1420, 171, 2479, 2883, 693, 1675, 213, 3122, 1299, 2285, 2286, 3069, 1818]. Differential-Difference [2000, 3068, 386, 1883, 2285, 2286, 3069]. Differentials [77]. Differentiation [1369]. Difficulty [2458]. Diffraction [319, 515, 1592, 2071, 118, 119, 890, 2540, 1067, 3377, 2467, 805, 1237, 692, 3525, 3526, 1090, 301, 127, 3007, 3126, 821, 1771, 54, 245, 574, 867, 973, 1267, 3448, 630, 1291]. Diffuse [1719]. Diffused [1039, 3481]. Diffusion [2880, 3312, 3284, 2465, 2360, 2411, 2804, 2501, 3416, 3163, 1969, 1890, 1849, 2933, 1979, 2773, 2155, 696, 1852, 1833, 2085, 3319, 1207, 1354, 1535, 3456, 1035, 1525, 1672, 2120, 2803, 2570, 1559, 1628, 1758, 1826, 1799, 2233, 3386, 2751, 2856, 3274, 3583, 2881, 1676, 1945, 1830, 3173, 3119, 3379, 2333, 2189, 3073, 1530, 2325, 2895, 1329, 1591, 1599, 2018, 2350, 2517, 3524, 1880, 2190, 2722, 3024, 2592, 2441, 2061, 3519, 3045, 2627, 2303, 3236, 1929, 2161, 2790, 3570, 3571, 1720, 2316, 2996, 2863, 1915, 2051, 3457, 2242, 3337, 2186, 1851, 2649, 2644, 2466, 2133, 2288, 2492, 1742, 2021, 2022]. Diffusion [2104, 2314, 2788, 2230, 2206, 3001, 1888, 2087, 321, 2591, 1247, 2490, 3121, 2862, 3131, 1934, 2041, 3431, 2128, 1058, 601, 2267, 2569, 2666, 3050, 3216, 1734, 2444, 3120, 983, 1200, 1603, 1671, 1855, 2037, 2149, 2055, 407, 2794, 375, 2180, 1783, 1390, 1391, 1392, 3148, 2555]. Diffusion-Convection [2862]. Diffusion-Coupled [601]. Diffusion-Like [983]. Diffusion-Reaction [3274, 2788, 2230, 2128]. Diffusional [2646]. Diffusional/Thermal [2646]. Diffusions [2194, 2237, 2487, 1891, 3138, 1263, 2374, 3181, 3096, 2742, 3304, 2385]. Diffusive [2178, 1649, 1721, 3157, 2593, 2372, 3192, 1821, 1832, 1861, 3481, 1511, 1288]. Diffusivity [3231]. Diffussive [1729]. Digital [2870, 3228, 2884, 1497, 1886]. Digraph [52]. Digraphs [172, 349, 1219, 1692]. Dihedral [2420, 2421]. Dilations [2746]. Dilemma [685]. Dilute [2979, 2699, 2562, 2469]. Dilution [1612]. Dimension [1519, 1660, 2515, 2693, 389]. Dimensional [515, 1436, 1592, 2429, 2637, 2173, 3270, 462, 1795, 1997, 3506, 567, 2796, 2378, 3326, 3484, 766, 3499, 3446, 3409, 3282, 3480, 3424, 3100, 3577, 3368, 3483, 2969, 2439, 44, 798, 2262, 2273, 3400, 1170, 1302, 948, 2406, 835, 1158, 2762, 1091, 1241, 2895, 2602, 1956, 3253, 1866, 1991, 882, 1908, 2154, 1842, 3343, 2506, 2304, 2827, 2194, 2237, 2818, 688, 1101, 2218, 3280, 2303, 1334, 2341, 2502, 2150, 3032, 2108, 2270, 2653, 1691, 2475, 2990, 3376, 749, 3381, 530, 2683, 1751, 2203, 2968, 3491, 726, 2490, 3135, 1912, 2447, 2374, 3088, 3087, 2672, 1134, 2445, 2536]. Dimensional [2680, 2665, 828, 2444, 393, 3572, 2617, 506, 2628, 2806, 3363, 1598, 1220, 2582, 2891, 2957, 1137, 757, 1639, 316, 2742, 2780, 2279, 2375, 3393, 3297, 1995, 2204].

Dimensionality [2092]. Dimensions

[3202, 2630, 769, 530, 3380, 2358]. Dimpling [2175]. Diodes [2667]. Dipole [3574, 1067]. Dirac [965, 1157, 1400]. Direct [2924, 2710, 3088, 2297, 796, 995, 2456]. Direct-Formulation [3088]. Directed [52, 747, 82, 28, 1573]. Directional [2945, 2699, 2571, 3299]. Director [2855]. Dirichlet [168, 2831, 1007, 3287, 741, 2462]. Dirichlet-to-Neumann [3287]. Disappearance [2497, 2511]. Disbursement [1304]. Disc [286]. Discharge [1423, 1834, 2958, 3194]. Discipline [105]. Disciplines [527]. Disconjugacy [185]. Disconjugate [1143]. Discontinuities [1592, 1125, 878, 3142, 2617]. Discontinuity [3548]. Discontinuous [3101, 3248, 3417, 2728, 598, 3471, 3575, 533, 1925]. Discount [959]. Discounting [1304]. Discrete [1935, 3182, 3227, 1546, 1506, 880, 3338, 1571, 2281, 3152, 959, 2969, 2967, 798, 2759, 2612, 1167, 3228, 1599, 1223, 620, 2413, 2636, 3481, 166, 3197, 3135, 992, 1470, 2091, 2877, 1517, 3462, 1510, 1518, 1306, 1402, 3349, 2239, 1185]. Discrete-Time [3338, 2969, 798, 2759, 3228]. discretization [3299]. Discussion [2920, 2544]. Disease [2083]. Diseases [3252]. Dishonest [130]. Disjoint [1330, 1487, 1364, 1045, 1194]. Disk [119, 1988, 2213, 1938]. Disks [3019, 2213, 1371, 489]. Dislocations [573]. Disorder [3032]. Disordered [2260]. Disorders [3290]. Dispersal [2986, 1489, 2552, 2064]. Dispersed [2875]. Dispersion [2567, 1940, 2593, 2131, 2231, 3537, 962, 2857, 2765, 2654, 3356, 2369, 3191]. Dispersive [118, 2923, 2864, 3006, 2634, 3362, 2829, 79, 3059]. Displacement [895, 2065, 361, 759, 2605]. Displacements [1046, 324]. Dissection [1626]. Dissipation [2776, 2024, 2658]. Dissipative [206, 2663, 2664, 2915, 3142, 2390, 3148]. dissociation [1897, 1896, 1964]. Dissolution [3234]. Distance [887, 1706, 2922, 29, 1320, 365]. Distances [1265, 1109, 3406]. Distinct [354]. Distinguishable [941]. distortion [2456]. Distributed [589, 2795, 117, 807, 962, 1802, 818, 675, 2280, 856, 3358]. Distribution [1578, 935, 1803, 3125, 3016, 2789, 305, 971, 1548, 979, 2483, 1848, 1464, 2809, 2319, 1059, 2521, 1140, 891, 2211, 1948, 834, 2954, 780, 1117, 1560, 3080, 1716, 1691, 473, 675, 3, 1494, 2251, 2321, 2449, 2559, 1565, 1476, 2259, 1971, 2503, 533, 1432, 331, 3181, 1602, 50, 556, 1983, 2014, 2573, 476, 499, 1161, 3301, 2421]. Distribution-Free [1803]. Distributional [2195, 205, 449, 53, 8, 110, 17, 104, 138, 184]. Distributions [883, 211, 206, 202, 299, 1529, 880, 3328, 2991, 75, 1421, 1705, 3513, 1191, 1345, 2275, 2535, 1074, 2079, 2214, 2401, 439, 722, 1987, 2136, 2520, 1155, 210, 2477, 3339, 279, 1841, 249, 1617, 1499, 2175, 571, 1659, 545, 2598, 989, 1156, 2868, 1933, 2075, 1976, 60, 2745]. Disturbance [3446, 3409, 3070, 1735]. Disturbances [1595, 1322, 1206]. Divergence [2609]. Divisibility [1629]. Divisible [2435, 75, 1074]. Division [2383, 2275, 213]. Divisors [881, 1151]. DNA [2789, 2553]. DO [1476]. Dogs [2833]. Domain [2154, 950, 1701, 3306, 1256, 594, 3285, 438, 2912, 2375]. Domains [3368, 1294, 3573, 1027, 2447, 1303, 825, 2548, 1408]. Dominance [2954]. Dominated [1235]. Dominating [1819]. Doniach [3145]. Dopant [2487]. Doping [3105]. Double [2126, 1980, 3518, 1511, 1432, 1288, 2631, 2478]. Double-Diffusive [1511, 1288]. Double-Pulse [2631]. Doublet [2062]. Doubling [2625, 3106]. Doublings [2160, 3107]. Doubly [1549]. Doubly-Reflected [1549]. Down [1538]. Drag [2016, 3382, 3004, 3200, 3168]. Drazin [1411, 1369, 1419, 1380, 1513, 1433]. Drift [1691]. Driven [2751, 2652, 2309, 2266, 2076, 2818, 2673, 2587, 3130, 2048, 3317, 3106, 2720, 3291]. Drop [1817, 2409, 2623]. Drug [3438]. Drying [2949, 2948]. Dual [1086, 1215, 366, 2493, 752, 2336, 2226, 1434, 522, 263, 1897]. Duality [906, 109, 1233, 1116, 502, 278, 93, 217, 40, 398, 621, 148, 2675, 231]. Due [2803, 3345, 3535, 1922, 987, 2825, 2533, 2001]. Duels [401, 1387]. Duffing [2379, 1761, 108, 391, 2620, 2024, 2305]. During [3530, 2632]. Dyed [2999]. Dynamic [3419, 2534, 1532, 2115, 897, 2162, 3348, 959, 1178, 1744, 2464, 1382, 3492, 179, 3209, 3196, 1272, 1924, 570, 214, 1810, 915, 949, 1511, 1755, 1788, 1355, 358, 2656, 3191, 126, 1463, 3246]. Dynamical [2277, 3354, 1044, 1814, 2733, 2489, 1536, 2683, 3197, 1367, 2091, 2479, 51]. Dynamics [3458, 3042, 2601, 3315, 3022, 2893, 2500, 3389, 3023, 3533, 2495, 3557, 3227, 1730, 1557, 3391, 3119, 3102, 2814, 3198, 3532, 3461, 1657, 3103, 3537, 2918, 2855, 3140, 3402, 3403, 2961, 2412, 2527, 1832, 1922, 2673, 3482, 1227, 2087, 2187, 2915, 2408, 1662, 3148, 1247, 2909, 3551, 3317, 3439, 2672, 3567, 186, 2315, 543, 1477, 1894, 2257, 3041, 2323, 2498, 3129, 2883, 1754, 3226, 3084, 3295]. Dynamics-II [186].

Earth [2279, 2576]. EBK [2861]. Echelette [3448]. Ecological [1791, 1638]. Ecology [2649]. Economic [1884]. Economics [600, 854]. Economy [1556]. Ecosystems [2277]. Eddy [3447, 3451]. Edge [515, 669, 495, 1533, 1561, 1487, 1364, 3502, 511, 1068, 1246, 245, 126, 868, 1819]. Edges [805, 1237, 496, 1487, 510]. Effect [3090, 2817, 3127, 3220, 3299, 2648, 1207, 1354, 221, 2398, 2910, 3194, 3535, 1022, 3180, 1637, 131, 3519, 572, 2697, 3370, 3046, 2740, 2878, 3300]. Effective [2610, 2486, 1337, 2996, 3242, 3231, 2132, 3536, 3091, 2829]. Effectiveness [3422]. Effects [2841, 3495, 2565, 3504, 3141, 3163, 1890, 2515, 2115, 160, 1366, 1396, 1512, 2593, 3034, 2293, 2867, 3322, 2165, 1714, 3320, 2943, 1785, 2247, 598, 2064, 3234, 1756, 1840, 2676, 3223, 1492, 2571, 2959, 3511, 2385]. Efficiencies [473]. Efficiency [1441, 505, 649, 1577, 2138, 2822, 3578]. Efficient [2638, 3324, 2808, 588]. Eigenfrequencies [3326, 2828]. Eigenfunction [2728, 1977]. Eigenfunctions [1286, 2600, 739, 699, 794, 488]. Eigenmodes [3115]. Eigenshape [2819]. Eigenstrain [2188]. Eigensystem [3232]. Eigensystems [1108]. Eigenvalue [1804, 3325, 130, 509, 519, 97, 2127, 720, 1217, 784, 2126, 10, 1648, 1637, 836, 429, 753, 1027, 1138, 1622, 133, 1024, 3055, 3240, 1992, 2451, 3012, 3011, 1408, 793, 2912, 488]. Eigenvalues [1199, 27, 808, 1141, 3224, 2099, 2770, 10, 912, 1760, 1418, 2793, 1874, 1875, 2264, 950, 1891, 2144, 142, 952, 1134, 3049, 699, 33, 3161, 2478, 3183]. Eigenvectors [750, 912, 1749]. Eigenvibrations [1351]. Elastic [1046, 3005, 3090, 1778, 2457, 3092, 1984, 262, 2265, 509, 3507, 582, 945, 1351, 2982, 3206, 1072, 666, 2131, 3555, 901, 3247, 1017, 1016, 341, 2771, 2714, 2309, 2266, 2188, 645, 1952, 790, 3075, 3186, 3568, 3427, 1203, 1902, 1010, 915, 414, 829, 30, 1261, 1977, 344, 1100, 2440, 2677, 573, 137, 3406, 1291, 1377]. Elastic-Plastic [3005, 3090, 1984, 1016]. Elastica [1740, 932, 332]. Elasticity [1436, 37, 3034, 2, 2154, 80, 1015, 3258, 698, 3251, 215, 284, 1995, 2204]. Elasto [3144, 192]. Elasto-Plastic-Microstructure [3144]. Elasto-Plasticity [192]. Elastodynamic [2859, 759, 2456]. Elastoplastic [3412, 3225]. Elastostatics [2221, 3521, 344]. Electing [1959]. Election [1650]. Electoral [1590]. Electric [760, 2928, 2071, 1076, 1067, 2387, 2770, 2492, 1897, 1896, 1964, 1331, 2931]. Electric-Thermal [2928]. Electrical [3265, 979, 3128, 1078, 2641, 3554, 900, 2930, 2965, 2689, 3408]. Electro [1630, 2562, 2469]. Electro-Epitaxial [2562, 2469]. Electro-Mechanical [1630]. Electrode [2930, 2931]. Electrodes [2867]. Electrodiffusion [2445]. Electrodynamics [1220]. Electrolytes [1169]. Electromagnetic [3428, 3561, 119, 2197, 2858, 1796, 3257, 739, 2753, 3278, 3378, 3007, 2617, 1204, 501, 2279]. electromagnetically [2594]. Electron [2092]. Electroneutrality [2445]. Electropainting [2101]. Electrostatic [1927, 2416, 2907, 3388, 259, 191, 2150, 1349]. Element [3164, 1273, 3477, 3088, 3087, 3205]. Element/Spectral [3477, 3205]. Elementary [881, 1636, 381, 1151, 1044, 336, 204, 2744, 512]. Elements [3037, 207, 1685, 1011, 2193, 1440, 2549]. Eliminating [451]. Elimination [1573]. Ellipsoid [1820]. Ellipsoidal [2136, 2241, 1070]. Elliptic [3164, 1996, 2928, 939, 3019, 2219, 1399, 2170, 700, 869, 3541, 1172, 576, 1905, 177, 216, 687, 2544]. Elucidation [1481]. Embedded [1300]. Embedding [1347]. Emission [3324, 107, 3175]. Empirical [571]. Emptiness [1223]. Emulation [3503]. End [3220, 3034, 645, 2850, 361, 1093, 324, 276, 2571]. End-Displacement [361]. End-Resonance [3034]. Ended [1258]. Endemic [2551]. Endpoint [2657, 618]. Energy [1566, 3515, 2795, 314, 3218, 2092, 2074, 1952, 1015, 3427, 1247, 1353, 1300, 3318]. Energy-Casimir [3515]. Energy-Dependent [1247, 1353]. English [1914]. Enhanced [3073]. Enskog [2618]. Entire [10]. Entrance [2510]. Entries [952]. Entropies [989, 1156]. Entropy [2181, 402, 2599, 2675]. Entry [1635]. Enumeration [411]. Envelope [49, 1147, 1936]. Envelopes [1450].

Environment [3319, 2876, 3328, 1502, 1492]. Environment-Man [3319]. Environments [1459, 2200]. Enzyme [1530]. Ephaptic [1930]. Epicenter [1902]. Epicentral [1902]. Epicentral-Axis [1902]. Epidemic [1649, 1721, 1872, 3223, 1425]. Epidemics [2551, 2085, 3319, 3422]. Epitaxial [2562, 2469]. Epitrochoids [49]. Epoch [1201]. Equal [3527, 2392, 1239, 3142, 757]. Equality [434]. Equally [3016]. Equation [2695, 3579, 2198, 2011, 536, 2769, 2609, 1811, 980, 1126, 2997, 2472, 2471, 3250, 1040, 3022, 1277, 1683, 937, 729, 1969, 2424, 2923, 2774, 3434, 3019, 2178, 2496, 81, 453, 84, 2130, 3009, 3485, 2155, 777, 958, 1686, 2379, 3446, 3409, 1852, 2094, 2735, 494, 839, 1312, 2754, 1142, 3227, 1105, 1393, 1642, 2007, 317, 3224, 3235, 3288, 3454, 1396, 1512, 2216, 382, 1801, 203, 864, 2927, 1704, 2936, 1121, 1333, 61, 1604, 1945, 2560, 1981, 878, 1989, 2814, 2351, 1831, 2920, 2381, 255, 3116, 78, 1435, 1224, 1113, 3143, 1071, 2121, 3524, 840, 823, 836, 633, 3521, 2889]. Equation [1409, 2632, 2207, 2853, 3366, 2190, 1442, 273, 2779, 1144, 713, 337, 733, 1453, 1454, 1563, 389, 2719, 3038, 752, 2865, 843, 454, 889, 1575, 1042, 769, 3032, 2051, 2636, 3156, 1186, 1455, 1957, 1539, 2392, 1260, 1095, 1384, 3373, 108, 1851, 2840, 47, 1627, 2098, 2139, 718, 391, 1323, 1973, 43, 814, 1507, 1075, 2826, 2006, 1359, 1888, 2087, 424, 3306, 386, 3115, 421, 1404, 618, 908, 1476, 1102, 783, 2805, 2862, 1709, 1883, 2618, 463, 73, 123, 2422, 2977, 736, 1905, 122, 968, 418, 2584, 1332, 1934, 2643, 940, 1256, 1254, 2370, 3070, 730, 2024]. Equation [3441, 990, 1470, 1625, 707, 2410, 983, 270, 554, 394, 2688, 3436, 242, 743, 400, 2720, 229, 222, 1449, 1677, 577, 1082, 3421, 171, 2440, 395, 3283, 2255, 2437, 91, 3193, 1829, 2658, 2562, 2469, 1289, 2375, 542, 3432, 1390, 3122, 3300, 3298, 1299, 3154, 3314, 1523, 2144, 2614, 3264, 3204]. Equational [2733]. Equations [1271, 3312, 3574, 3539, 714, 3412, 2376, 2963, 2916, 3561, 3010, 2997, 969, 3321, 141, 431, 911, 1996, 976, 2411, 3449, 2928, 921, 1131, 1413, 3262, 663, 118, 2142, 2775, 500, 631, 1047, 291, 3484, 2013, 3085, 1278, 135, 1411, 188, 547, 1833, 1535, 537, 129, 874, 2795, 3480, 817, 1399, 2170, 2925, 902, 851, 2633, 1621, 2767, 2924, 1672, 1388, 3014, 700, 1628, 2352, 1826, 2110, 2917, 559, 2201, 415, 1315, 2162, 2927, 483, 1895, 591, 3210, 2372, 535, 1128, 2942, 366, 461, 1980, 1174, 2373, 2038, 761, 970, 1043, 1495, 2953, 2202, 1427, 367, 1170, 1302, 2325]. Equations [723, 3367, 2481, 2762, 1700, 2012, 1717, 237, 2895, 265, 10, 1344, 1484, 676, 1329, 1591, 380, 3260, 413, 1903, 1381, 1049, 2018, 35, 613, 827, 340, 1982, 836, 3573, 2480, 2823, 2983, 2713, 1635, 2716, 2874, 3024, 2962, 3343, 625, 41, 603, 2232, 624, 964, 1500, 2141, 2668, 6, 1250, 544, 611, 849, 558, 1314, 266, 2818, 1570, 1426, 752, 2161, 2482, 2694, 2793, 3580, 1172, 357, 443, 2580, 768, 806, 1797, 2043, 746, 1915, 3457, 524, 1147, 2000, 3068, 1486, 1567, 218, 379, 3547, 2088, 1552, 3337, 288, 3026, 2890, 2778, 1259, 1376, 862, 1357, 312, 749, 1385]. Equations [2217, 335, 561, 779, 1823, 396, 1249, 528, 2860, 1245, 2230, 1162, 1232, 2757, 271, 2753, 552, 2734, 563, 598, 1522, 2808, 386, 1327, 2657, 2393, 555, 3575, 2645, 1120, 3491, 756, 498, 154, 2040, 286, 2276, 3551, 2768, 1612, 3147, 1414, 2450, 2686, 848, 2128, 1058, 1439, 3050, 450, 812, 3472, 475, 778, 1099, 470, 174, 1688, 819, 922, 562, 45, 88, 216, 1190, 468, 2083, 957, 263, 2628, 622, 70, 745, 3279, 580, 2629, 1725, 1415, 1197, 1200, 917, 1603, 1292, 1598, 1346, 991, 3370, 2582, 1420, 2479, 1212, 2631, 687, 1675, 213, 2541, 2452, 3299]. equations [3051, 3043, 3303, 2285, 2286, 3069, 2861, 2913, 1403, 2358, 3148, 2507]. equi [2432]. equi-sized [2432]. Equidiffusional [2846]. Equilibrated [645]. Equilibria [1687, 2026, 1841, 3050, 327, 2901, 735, 3122]. Equilibrium [1778, 560, 2164, 1705, 3417, 2460, 1097, 2229, 1500, 480, 2455, 1807, 1891, 1817, 283, 19, 2557, 451, 2062, 1783, 3171]. Equivalence [1458, 1385, 23]. Equivalent [437, 667, 2043, 492, 404, 12, 3525]. Equivariant [2463, 1745]. Erdélyi [189]. Ergodic [1053, 2503]. Ergodicity [2203, 448]. Erlang [1802, 3003]. Errata [854, 1546, 1512, 2639, 3043, 1479, 2539, 1885, 253, 2450, 1361, 1677, 139, 138, 184]. Erratum [1789, 988, 1721, 2252, 667, 1932, 2309, 1585, 1101, 2044, 2903, 1232, 1494, 3585, 3183, 1788, 1156, 792, 896, 2100, 2562]. Error [921, 3484, 2307, 2403, 3524, 2182, 2693, 64, 2089, 1666, 2998, 182, 187, 983, 2531, 101, 1618, 102]. Error-Correcting [187]. Errors [1994]. Escape [2328, 2165]. Essential [576]. Estimate [1234, 1663, 588]. Estimates [921, 1131, 3484, 267, 367, 297, 2182, 1635, 1172, 14, 2165, 1666, 1565, 1659]. Estimating [729, 2386]. Estimation [18, 3338, 1827, 2052, 1081, 2337, 471, 811, 2338, 733, 1133, 3324, 1698, 1963, 1037, 2006, 2734, 1617, 182, 387, 410, 518, 1680, 264, 420, 86, 2834, 2780, 477, 3292]. Estimations [2922, 14, 3091]. estimator [1670]. Estimators [1953, 2181, 649, 1887, 62, 1994, 144, 473, 1864, 63]. Euclid [1607]. Euclidean [534, 220, 44, 646, 1517, 2295]. Euler [2920, 3343, 1239, 2358, 3491]. Evaluation [1961, 3159, 303, 3388, 1797, 944, 1663, 3177, 565, 2402, 2548, 2640]. Evanescent [2114]. Even [1197]. Every [495]. Everyday [995]. Evidence [1855]. Evolution [3516, 1833, 2795, 1105, 2924, 2712, 2560, 1381, 3365, 2589, 3099, 3517, 2666, 2665, 3120, 3285, 2738]. Evolutionary [2649, 2820, 1109]. Evolutions [1951, 2580, 3313]. Exact [2505, 566, 2219, 958, 1142, 237, 3154, 3136, 3294, 1993, 1664, 2743, 1359, 2753, 2370, 2020, 101, 3193, 867, 973]. Exactly [1989, 3529]. Examining [3236]. Example [2424, 2418, 416]. Examples [1984, 2941, 2031]. Exceptional [1218].

Excess [1248]. Excesses [1283]. Exchange [2092, 2289, 1062]. Excitable [2756, 1718, 3235, 2593, 2234, 2326, 3277, 1599, 1871, 2371, 2413, 2575, 2851]. Excitation [1764, 2279]. Excitations [1957, 3433]. Excitatory [3149]. Excited [923, 3518, 238]. Exclusion [3252, 3563, 2182]. Exercise [2158]. Exhibiting [1311, 3476, 1770, 1469, 3095, 1756, 2389, 1137]. Existence [1141, 3449, 2424, 2783, 81, 453, 1119, 1105, 1351, 3067, 3056, 356, 1981, 2373, 1777, 762, 3443, 367, 2159, 1651, 1050, 1226, 2458, 2278, 339, 954, 1487, 1138, 2636, 524, 590, 1701, 1249, 814, 2187, 2473, 3121, 1129, 1057, 2618, 19, 3362, 1343, 812, 529, 819, 2410, 2931, 1394, 3468, 1639, 737]. Existence/Stability [2278]. Existing [3087]. Exit [2652, 3339, 1536, 2710, 2672, 1734, 3181, 1889, 1976]. Expanding [2205, 3370]. Expansion [3125, 2516, 766, 3480, 2260, 880, 1106, 2608, 1319, 64, 1438, 1389, 741, 2098, 2139, 115, 2654, 1524, 5, 100, 1977, 3289, 2957, 2180]. Expansions [1169, 2011, 233, 244, 1988, 2540, 2767, 3160, 1295, 3212, 2752, 1, 1421, 970, 2600, 2760, 3354, 2195, 3243, 3571, 2728, 2183, 74, 1552, 440, 3376, 189, 1331, 425, 888, 3169, 1965, 2888, 3012, 630, 2196]. Experiment [1763]. Experimental [2265, 1458, 3562]. Experiments [1081, 991]. Explicit [1758, 1373, 3536, 918]. Exploding [3516]. Exploitative [1954]. Explosion [1863, 1937]. Explosions [3540]. Explosive [2589]. Exponent [2339, 1690, 3395, 2587, 2990, 2490, 446, 3291]. Exponential [2721, 935, 871, 2427, 2795, 1753, 2991, 1339, 1479, 2664, 1698, 779, 2860, 448, 656, 5, 1432, 1497, 1857, 2861, 2913]. Exponentially [649]. Exponents [2637, 3260, 2743]. Exposed [1733]. Expression [24]. Expressions [538, 198, 1157, 1958, 1400]. Extend [2181]. Extended [2226, 2822, 2877, 231, 852, 853]. Extending [1417]. Extension [3515, 641, 434, 2599, 1429, 200, 1600, 199, 1933, 420, 242, 652, 2675, 1650, 1229]. Extensions [487, 1916, 1211, 1361, 771]. Exterior [1436, 2769, 1996, 2154, 501]. External [1708, 1604, 2900, 1070, 3111, 1137]. Extinction [2797, 3269, 1719, 2143, 1527, 2443, 2624, 1971, 2894, 2794]. Extinguished [2527]. Extrapolation [166, 362]. Extremal [1518]. Extreme [2319, 257, 473]. Extreme-Value [473]. extremely [1937]. Extremum [882]. Extrinsic [3127, 3220]. Eye [1544].

F [46]. Faces [1084, 1684, 100]. Facets [1569]. Factor [959]. Factorial [2791]. Factorials [1713]. Factoring [1655]. Factorisation [2320]. Factorization [2721, 1393, 2078, 1772, 388, 5, 898]. Factorizations [1373]. Factors [2721, 345, 602, 1203, 3223, 3544]. Failure [3208, 2823, 2413, 1580]. Failures [634]. Falkner [823, 454, 2643, 1289]. False [2970, 2454, 2588, 856]. Families [3092, 484, 3478, 379, 144, 796, 63, 826]. Family [49, 800, 1282]. Fan [1836, 2994]. Fan-Beam [1836]. Far [1920, 3224, 2439, 3543]. Farkas [825]. Fast [1532, 2505, 3494, 1545, 2122, 3381, 3095, 3459, 2569, 3304]. Faster [2922]. Fault [924]. Feasibility [2886]. Feasible [1163, 1066]. Features [984]. Feedback [2414, 1883, 1367, 1710]. Fermat [551]. Ferrets [2833]. Ferromagnet [2463]. Fertility [1768, 2415]. Few [1487]. Fewest [445]. Fiber [2621, 3217]. Fibers [2575, 3259]. fibre [3296]. fibre-reinforced [3296]. Fibres [3221]. Fickian [3192, 2350]. Field [3576, 1927, 2416, 2907, 2016, 2797, 3163, 3201, 1920, 3224, 1918, 2439, 2608, 2910, 1067, 2038, 1831, 1091, 1466, 208, 191, 3573, 1968, 2412, 2527, 2502, 3543, 1664, 2644, 2442, 1658, 321, 3536, 2645, 605, 1054, 1300, 2557, 1204, 2740, 2878, 1655, 2180, 2780, 501, 3202, 1897, 1896, 1964, 2614, 1226]. Field-Effect [2910]. Field-High [2442]. Fields [760, 3159, 3388, 3452, 363, 3323, 2800, 2874, 1842, 2855, 2506, 3045, 1891, 701, 409, 3273, 3378, 920, 759, 316]. Filament [3141, 1640]. File [195]. Files [478, 333]. Film [2123]. Films [3466, 2875, 3402, 3403, 2905, 1754]. Filter [1978, 1836, 422, 3228, 2822]. Filtered [2112]. Filtering [2519, 1961, 116, 733, 2194, 2237, 2374, 1482, 2742]. Filters [2534, 1076, 2870, 1886]. Filtration [3418, 2981]. Fin [956]. Finding [2131, 986, 924, 2656]. Fine [3385]. Fingering [3347]. fingers [3214]. Finite [3164, 3005, 3270, 3506, 2817, 3037, 1732, 765, 2199, 2398, 89, 747, 3338, 1938, 44, 798, 1266, 1273, 2759, 646, 1466, 676, 2031, 2867, 230, 179, 2089, 2143, 2188, 3477, 2399, 2691, 3396, 1664, 2587, 47, 927, 815, 698, 1522, 974, 1587, 2175, 3410, 3088, 3087, 656, 1134, 1355, 358, 710, 3002, 3340, 920, 251, 2549, 1771, 852, 853, 2548, 39, 449, 364, 1655, 3030, 3058, 318, 2375, 542, 1995, 2204, 3205, 3341, 2504]. Finite-Capacity [3340, 3341]. Finite-Difference [3338, 2375]. Finite-Dimensional [3270, 44, 798]. Finite-Horizon [2759]. Finite-Source [230]. Finite-State [179, 2587]. Fire [1960, 2988]. Firing [1387, 2988]. First [1592, 1978, 3009, 3085, 135, 2420, 2735, 839, 3014, 1880, 1223, 2353, 753, 2673, 1127, 2653, 598, 2185, 2805, 2938, 1360, 3049, 3183, 88, 826]. First-Order [135, 2735, 2673, 2805, 2938, 3049, 3183]. First-Passage [2653]. Fisher [3023, 1567, 1263]. Fisk [733]. Fission [578]. Fitzhugh [1982, 2161, 1883]. Five [2619]. Fixed [751, 484, 1202, 235, 3096, 738, 868]. Fixed-Point [484, 3096]. Flame [2797, 2459, 2496, 2147, 2458, 2846, 2034, 2045, 2082, 2177, 2248, 2646, 1786, 1785, 1853, 1917, 1999, 2624, 2844, 1835, 2023, 2348].

Flames [2647, 2893, 2158, 1207, 1354, 3520, 2395, 2347, 2814, 2845, 2268, 2441, 2412, 2527, 3045, 2177, 1742, 2205, 2176, 1828, 2786, 2274, 2428, 2894, 1892, 2222]. Flat [1235, 511, 1395, 1093, 809, 1574]. Flattening [1080]. Flexible [1970]. Flexural [341]. Flexure [262, 1093, 190]. Flicker [1867]. Flickering [3520, 1867]. Flight [3542]. Floating [35]. Flooding [2873]. Floods [2484]. Flow [1255, 2173, 3504, 2951, 2952, 2607, 2072, 2156, 320, 2368, 2797, 1131, 567, 3453, 2933, 2783, 575, 680, 1077, 1235, 1538, 129, 3086, 1642, 3100, 170, 1307, 304, 1643, 2313, 2801, 3286, 894, 1362, 1989, 2873, 435, 2425, 3072, 773, 835, 259, 2510, 1345, 2976, 823, 2823, 2713, 1879, 1635, 2312, 3492, 2367, 2679, 2827, 1533, 1561, 447, 3230, 2218, 2960, 3045, 3346, 3566, 2696, 785, 3479, 1582, 2661, 2644, 2784, 1209, 2824, 3355, 1371, 2749, 2765, 511, 1036, 1514, 2526, 3088, 888, 3147, 2643, 3060, 3188, 2697, 374, 532, 3211, 3245, 2503, 3398, 2678, 610, 828, 1966, 1187]. Flow [1965, 2111, 2410, 3572, 1601, 809, 3017, 489, 719, 896, 2705, 1231, 2105, 3057, 2891, 2934, 3549, 1246, 1581, 2959, 2978, 3059, 2289, 3246, 3295]. Flows [2662, 412, 599, 2866, 3284, 3559, 1795, 1997, 2763, 2815, 1437, 2056, 1782, 3564, 716, 2782, 3073, 2873, 1358, 3476, 882, 22, 2123, 933, 2076, 2314, 3231, 1184, 3384, 301, 2422, 1054, 1270, 223, 3216, 2444, 1394, 543, 3363, 1474, 1137, 3405, 3213, 1793, 2003, 2992]. Fluctuating [3243, 1492, 1746, 1747]. Fluctuations [3056, 1869, 1731, 1800, 3465, 2294]. Fluid [1847, 919, 546, 1538, 2056, 170, 2131, 2872, 3535, 3247, 1657, 3342, 3450, 2218, 2475, 1227, 3186, 3231, 739, 2187, 3199, 1662, 2977, 3216, 20, 2111, 3357, 2821, 2868, 1877, 3089, 1208, 543, 173, 905, 2677, 3305, 2053, 2100, 1229, 3179, 3246]. Fluid-Loaded [2131]. Fluid-Solid [3450, 3186]. Fluids [2887, 1437, 2974, 933, 2824, 2816, 2903, 1813, 1242]. Fluorescent [2999]. Flux [3139, 3101, 3345, 3248, 3366, 3111]. Fluxons [3104]. FM [3581]. Focused [3139]. Focusing [2299, 3204]. Fokker [3361, 3303, 3250, 1260, 1891]. Fold [1529]. Folding [3219, 2292, 2053, 2100]. Food [2335, 2968]. Footed [2833]. Foraging [2429]. Force [2016, 383, 835, 2668, 3207, 1300, 2557]. Forced [2756, 1126, 1859, 1366, 2761, 1043, 2603, 2849, 1158, 3116, 1794, 1885, 2850, 1259, 3036, 123, 3189, 1190, 3332, 3302]. Forced-Convection [1859]. Forces [921, 89, 1604, 1203]. Forcing [2926, 1435, 917, 2658]. Forest [1364]. Forest-Realizable [1364]. Fork [2790]. Form [225, 1341, 1063, 2767, 1702, 2864, 1110, 300, 2644, 3517, 2645, 791, 3383, 842, 879, 788, 542, 2609]. Formal [570, 1439]. Formalism [2272, 585]. Formation [3262, 1196, 3157, 3368, 3345, 3311, 3584, 1858, 2311, 2528, 2998, 2267, 3184, 2146]. Formed [3198]. Forms [299, 1987, 479, 693]. Formula [2210, 1319, 1545, 1000, 200, 2093, 1515]. Formulae [96]. Formulas [1061, 2771, 597, 892, 59, 17]. Formulated [3449]. Formulation [2519, 494, 2162, 308, 2935, 1325, 1963, 2422, 2977, 3088, 2617, 1998]. Formulations [3087]. Formulism [531, 782]. Forsythe [886]. Forward [1844, 2573]. Foundation [1520, 829, 2635]. Four [2203, 468, 411]. Four-Colored [411]. Four-Mode [2203]. Fourier [1546, 189, 1935, 305, 3553, 1506, 1, 1421, 1698, 2477, 2282, 449]. Fourth [1271, 453, 1121, 806, 1138, 541, 848, 730, 216, 1190]. Fourth-Degree [541]. fractal [3301]. Fraction [3583, 3045]. Fractional [903, 114, 1345, 3517, 550]. Fractions [3045, 706]. Fracture [3218, 2381, 785]. Fractured [696, 3086]. Frame [1188, 1380]. Framework [3177]. Frameworks [1727, 1728]. Frederiks [3468]. Fredholm [2600, 2762, 814]. Free [1803, 808, 2049, 2933, 2291, 674, 979, 2555, 2556, 2570, 3548, 2856, 2216, 2010, 2245, 3102, 3107, 2568, 2312, 2585, 1539, 178, 2879, 30, 707, 1394, 2566, 2109, 3057, 372, 2878, 3549, 2062]. Free-Convective [2049]. Free-Discontinuity [3548]. Freedom [3103, 3196, 2989]. Freezing [3495, 2622]. French [2024]. Frequencies [1034, 3510, 3195]. Frequency [2798, 3428, 2011, 3399, 1820, 1985, 2333, 2914, 78, 1825, 2318, 3450, 2585, 2575, 688, 1101, 2117, 3526, 3376, 2442, 3502, 681, 2968, 2294, 3007, 1267, 1377, 3525, 2799]. Friction [3133]. Friendship [877]. Frobenius [2778, 632]. Front [3192, 2189, 2814, 2458, 1929, 1665, 1853, 1999, 2267, 3050, 2981, 3295]. Fronts [2349, 3372, 3467, 2803, 3021, 3118, 3400, 2624, 2146]. Frost [2622, 3132]. Frustration [3241]. Fuel [1665]. Fuels [2441, 2034]. Fujii [1073]. Fully [1019, 1039, 982]. Function [907, 27, 233, 1457, 1548, 2153, 2355, 3248, 3229, 2164, 1276, 334, 1427, 1168, 2602, 2272, 3366, 2367, 965, 2089, 658, 741, 2208, 955, 300, 1646, 2644, 936, 181, 281, 1386, 1565, 321, 1331, 1881, 464, 3071, 532, 565, 474, 377, 706, 2402, 2688, 3279, 2806, 2238, 3301, 2421]. Functional [141, 431, 777, 3182, 2260, 2917, 535, 1128, 1144, 611, 266, 2580, 862, 1357, 2791, 396, 1359, 2734, 990, 2083, 394, 3267, 1299]. Functional-Differential [611]. Functionals [1799, 3386, 3440, 3324, 427, 2675]. Functions [3029, 1616, 258, 211, 206, 3018, 294, 2219, 202, 465, 2250, 1309, 2265, 381, 228, 1051, 3329, 770, 1295, 1305, 972, 203, 1696, 2281, 483, 351, 207, 564, 75, 484, 869, 1542, 1340, 626, 595, 1339, 1479, 602, 83, 204, 208, 152, 1615, 3377, 2195, 1838, 654, 1850, 1932, 209, 402, 1248, 662, 2035, 2520, 330, 2242, 2476, 3496, 210, 34, 458, 857, 1124, 742, 999, 96, 205, 1679, 1157, 425, 446, 545, 2598, 512, 1159, 533, 5, 2212, 764, 529, 2448, 704, 2885, 295, 171, 2440, 1400, 2957, 2883, 941, 1236, 1452, 212, 201, 213, 1185, 2057, 3206]. functions [1814]. Fundamental [2186, 2217, 1122]. Further [2922, 1363]. fuzzy [1900, 1901].

Gain [2078]. Gained [2625]. Gains [1474]. Galaxies [1722, 1272]. Galaxy [76]. Galerkin [2243, 2583, 2760, 2733]. Gallery [2706]. Galloping [2541]. Gambling [1304]. Game [540, 1472, 1167, 1576, 1097, 658, 300, 169, 1139, 474]. Games [1417, 403, 953, 1248, 1096, 1274, 619, 260, 296, 1699, 734, 852, 853, 735, 1463, 1405]. Gamma [1716, 1933, 556]. Gap [170, 3578, 3233, 3297]. Gaps [3577, 2400]. Gas [2798, 766, 1423, 2134, 2854, 1834, 2814, 2092, 3333, 3482, 1272, 610, 828, 1966, 1112, 2033, 1231, 1944, 3187, 1735]. Gasdynamic [3285]. Gaseous [2120, 1853, 2042]. Gases [3417, 3501]. Gasless [3063]. Gated [2971, 3497, 3137]. Gauss [929, 1670, 2054, 1037, 1751, 652, 477]. Gauss-Markoff [1670]. Gaussian [1789, 1769, 3293, 616, 2596, 887, 2307, 3483, 2038, 1842, 1619, 3482, 1263, 24, 368, 410, 556, 3126, 103, 499]. Gauzes [1867]. Gavrilov [3560]. Gegenbauer [115, 2504]. Gels [2722]. General [2721, 2928, 1047, 2250, 117, 482, 297, 1348, 3402, 3396, 3570, 3571, 1915, 3496, 1004, 293, 1037, 1521, 2040, 1360, 3177, 1656, 328, 707, 3383, 31, 973, 2883, 573, 2810, 477, 1670, 3301, 1299, 3170, 2511, 2431, 2623]. Generalised [2782]. Generalization [2500, 1611, 1052, 993, 209, 705, 1597, 530, 46, 734, 639, 1518]. Generalizations [1136, 878, 2827, 1324, 87, 2344, 999]. Generalized [1578, 1198, 2505, 976, 211, 206, 810, 202, 1061, 1713, 641, 946, 317, 1753, 880, 1848, 203, 351, 207, 455, 145, 3132, 255, 1974, 1319, 1140, 204, 208, 152, 1257, 1363, 1183, 627, 633, 2196, 625, 1743, 2232, 209, 780, 2493, 3353, 1314, 628, 2488, 330, 224, 1486, 1253, 1218, 210, 34, 647, 525, 943, 1029, 1114, 1038, 1251, 205, 417, 1679, 1845, 3278, 550, 618, 442, 664, 236, 724, 621, 695, 123, 963, 736, 938, 493, 2598, 193, 199, 1882, 2448, 892, 451, 989, 1156, 2361, 373, 1541, 86, 295, 1518, 438, 150, 212, 201, 213, 3405, 1783, 1390]. generalized [3331, 824]. Generated [294, 3452, 1772, 1552, 2400, 1514, 1749]. Generating [3029, 1084, 880, 2355, 1381, 1838, 2208, 5]. Generation [3372, 3442, 52, 2576, 3109, 1842, 3346, 955, 1658, 2705, 3462, 1748, 757]. Generator [1605]. Generators [1293]. Generic [3483, 1058]. Genetic [3067, 1685, 2670, 1263]. Genetics [1609, 2061, 1603]. Geodesics [1054]. Geographically [1971]. Geological [2053, 2100]. Geometric [935, 583, 1356, 2712, 109, 2325, 1301, 1816, 1770, 910, 3313, 283, 398, 621, 3308, 1998]. Geometrical [265, 2506, 2371, 3204, 2288, 2860, 1966, 3126, 3226]. Geometrics [646]. Geometries [757]. Geometry [220, 3265, 2817, 1690, 382, 3311, 2486, 1022, 1453, 1821, 534, 1353, 2828]. Geophysical [2273, 3238]. Geothermal [3218]. Gersgorin [615]. GF [1434]. GI [3030]. GI/M/1 [3030]. Ginzburg [3145, 3008, 3275, 3534]. Given [1737, 2355, 1613]. Gives [625]. Giving [2767]. Glacier [2577]. Glass [1712]. Glassy [2120, 2555]. Global [3412, 2698, 2501, 2500, 3516, 2843, 3557, 1649, 1721, 3507, 2755, 2352, 3391, 3077, 2125, 633, 3178, 3519, 1928, 786, 1714, 3496, 2940, 2618, 2993, 3390, 3585, 917, 2883, 3358, 3349, 1638, 3470, 2385]. Glow [1423]. Godunov [3170]. Goes [683]. Golf [800]. Golub [886]. Gomory [903, 732]. Gompertz [2351]. Gonorrhea [3252]. Gordon [3104, 203, 1333, 2914, 1563, 2840, 418]. Görtler [2289]. Goursat [1563]. Governed [1435]. Governing [2645]. Graded [1712, 3578]. Graded-Band-Gap [3578]. Gradient [2841, 3010, 2333, 3253, 42, 963, 518]. Gradients [3480, 485]. Gradostat [2852]. Gradual [1281]. Granular [2942, 3133, 2526, 2678, 3465, 2934]. Graph [844, 69, 129, 703, 642, 289, 1605, 429, 813, 241, 1330, 1251, 2972, 2747, 55, 256, 287, 295, 66, 691]. Graph-Theoretic [69]. Graph-Theoretical [241]. Graph-Valued [2747]. Graphic [684]. Graphical [1822, 1032, 1485]. Graphs [1465, 669, 672, 1611, 310, 1313, 1165, 747, 1472, 82, 495, 602, 457, 132, 1050, 28, 1707, 1239, 832, 1324, 3, 1494, 359, 548, 1573, 521, 738, 1819, 1693]. Grating [3448]. Gravitational [2412, 2527, 321, 1703]. Gravity [3062, 1869, 2300, 1394, 3185, 3549]. Gravity-Wave [1869]. Gray [770, 2882]. Greater [959]. Green [2265, 869, 1427, 1168, 34, 2644, 321, 618, 1331, 193, 2688, 3279, 2440, 426]. Greitzer [3506]. Griffin [38, 37]. Griffith [572, 1807, 1203]. Gross [1723]. Gross-Substitute [1723]. Groundwater [3088, 3059]. Group [1370, 887, 2774, 3429, 1903, 2918, 1589, 1765, 1365, 1619, 604, 1004, 1888, 2393, 940, 3266, 66, 2499, 2737]. Group-Invariant [2393]. Group-Testing [604]. Group-Theoretic [3266]. Groups [1293, 1608, 996, 2477, 701, 1107, 2725, 1175]. Grow [2819]. Growth [2919, 3504, 2946, 715, 1160, 2007, 3077, 113, 2443, 3347, 3552, 2723, 1768, 3052, 2562, 2469, 60, 1814]. GTD [2009]. Guckenheimer [3560]. Guide [1125]. Guided [3044, 1651]. Guides [566, 1516]. Gunn [2817, 3127, 3220]. Gyroscopic [969].

H [605]. Habitat [2429]. Habitats [2429, 1489]. Hadamard [467]. Haemofilters [2906]. Hahn [1571, 597]. Half [2170, 945, 1624, 3555, 3150, 648, 2059, 1902, 77, 1377]. Half-Order [77]. Half-Plane [2170]. Half-Space [945, 1624, 3555, 648, 1902, 1377]. Halfplane [127]. Halfspace [2739, 2688]. Hamilton [1612]. Hamiltonian [2802, 2899, 240, 642, 2939, 3196, 2380]. Hamiltonians [1906]. Hammerstein [1531]. Handicapping [800]. Handling [2792, 2597]. Hanging [339]. Hankel [330, 442, 664, 53, 60, 104]. Haploid [1502]. Hard [1267]. Hardening [3476, 1182]. Harmonic [3569, 2308, 1350, 2718, 3257, 3477, 1597, 2587, 268, 756, 2340, 1088, 2584, 2666, 1543, 1008, 3300, 2343, 3205]. Harmonics [1571, 1000, 1070]. Harmonizable [2225]. Harvested [3391]. Harvesting [3458, 3167, 1746, 1747]. Having [1374, 3551]. Hazard [568]. Head [3001]. Head-of-the-Line [3001]. Heat [84, 227, 2155, 942, 971, 2731, 2681, 839, 3424, 1315, 3307, 35, 2143, 1268, 3385, 2230, 1888, 974, 968, 707, 683, 3332, 2220, 3111, 757, 3193]. Heat/Mass [3424]. Heated [286, 719, 896]. Heating [2739, 3006, 3322, 3047]. Heave [2622, 3132]. Heaving [2585]. Heavy [3125, 582, 2837, 2792, 2869, 3097, 3396, 3028, 2449, 2559, 3137, 3003]. Heavy-Traffic [2837, 2792]. Heiles [2345]. Hele [2312, 3384]. Helium [1227]. Helmholtz [2100, 2769, 3300, 3224, 739, 3184, 2688, 3264, 2053]. Hénon [2345]. Hénon-Heiles [2345]. Hereditary [1366, 1396, 1512, 622]. Herglotz [3206]. Hermite [1427, 1140, 1212]. Hermitian [731, 903, 1426, 2638, 426]. Heteroclinic [3010, 3538, 2372, 2956, 2717]. Heterogeneous [2472, 2471, 3272, 3429, 2832, 2947, 1362, 1118, 2552, 2317, 3536, 2901, 1377, 2431, 2432]. Heterosexual [3236]. Hexagon [2397]. Hexagonal [1669]. Hidden [682]. Hierarchical [1471]. Hierarchies [3018]. Hierarchy [3522, 3523]. High [3164, 2798, 3141, 1269, 3453, 3404, 3505, 3275, 3411, 1974, 78, 2318, 2835, 3526, 3242, 3026, 2442, 2816, 2903, 1371, 391, 3502, 2968, 2799, 2374, 1726, 3060, 3188, 3578, 3318, 3007, 1377, 3525]. High-Contrast [3411, 3242]. High-Density [2816, 2903]. High-Frequency [2798, 2318, 3502, 2968, 3007, 2799]. High-kappa [3275]. High-Order [3164, 3141]. High-Speed [2835]. Higher [2472, 381, 2969, 805, 2778, 1342, 432, 687]. Higher-Dimensional [2969]. Highest [1228]. Highly [2460, 2554, 2530, 2108, 3320, 2390]. Hilbert [885, 309, 3292, 298, 2320, 2617, 503, 531, 850]. Hill [1801, 1409, 2255, 2437]. Hilliard [3022, 3051]. Hinshelwood [2037, 2149]. Hirsch [776, 481]. History [3557, 2702]. Hitting [2304]. HIV [2832, 3369, 2918, 3223, 3041]. HIV/AIDS [2832, 2918, 3041]. Hocquenghem [365]. Hodograph [2633]. Hohlfeld [3176]. Hole [262, 89]. Hollow [3527, 1878]. Hollowed [178]. Holomorphic [425]. Homoclinic [2695, 2493, 3026, 2479]. Homogeneity [2005]. Homogeneous [3561, 2331, 2164, 1842, 360, 3450, 2073, 2447, 2430, 1783, 1814]. Homogenization [3569, 3407, 3160, 3374, 3553, 2926, 2875, 3310, 2685, 2874, 1694, 3375, 1839, 3363]. Homogenized [3049, 3183]. Homological [1517]. Homology [2155]. Homosexual [3223]. Homotopic [833]. Hopf [2903, 2320, 2721, 3330, 2324, 2356, 2565, 2966, 1847, 2201, 2125, 2474, 3025, 2493, 1373, 2513, 2816, 2377, 2116, 2128, 2297, 1967, 2083, 1572]. Hopfield [3470]. Hopper [2934]. Hoppers [2526]. Horizon [2759, 1355]. Horizontal [3286, 1378, 1504]. Horizontally [2279]. Horseshoe [1544]. Hot [1893]. Howson [734]. Huffman [1406, 1743]. Hulls [1306]. Human [563, 1894]. Huntington [1555]. Hurwitz [526]. Hybrid [1592, 36, 2583, 2760, 3213]. Hydrodynamic [2848, 3074, 2827, 2150, 3514, 167, 1955, 1054, 156, 1349]. Hydrodynamics [1870, 3382, 2544]. Hydroelastic [1940, 1825]. Hydrogen [1365]. Hydrophone [2226, 1829]. Hydrostatics [1030]. Hyperbolic [1271, 2607, 3265, 118, 135, 817, 2560, 2545, 2546, 3162, 2890, 485, 3035, 2805, 2686, 1688, 3290, 1122, 2479, 2525]. Hyperbolicity [3005, 1315, 2713, 3343]. Hypercube [861]. Hypercycle [2207]. Hypergeometric [1589, 2476, 999, 96]. Hypergraphic [1616]. Hyperplane [3551]. Hyperplanes [72]. Hypersonic [575, 1235, 3398]. Hyperspherical [268, 2680]. Hysteresis [3412, 3354, 2683, 3106]. Hysteresis-Type [3412]. Hysteretic [3495].

Ideal [2185, 1057]. Identical [2324, 3048]. Identifiability [2964, 1537]. Identification [2609, 2866, 873, 1781, 3524, 1600, 3174, 1568, 2284, 2549]. Identifying [1945]. Identities [397]. Identity [1473, 1806]. Idle [279]. Ignition [2590, 3535, 1719, 1832, 2396, 1733, 1893, 2059, 3046, 3111]. IHR [1916]. IHRA [1916]. II [1176, 2452, 2642, 2966, 2952, 2543, 1519, 1391, 1901, 1265, 1728, 3453, 405, 1732, 3085, 1354, 1326, 3203, 3201, 3407, 3124, 2556, 2343, 3008, 3352, 3297, 2660, 2296, 3461, 2204, 1431, 68, 1110, 1814, 3403, 3083, 2545, 1561, 1776, 1937, 2096, 2237, 3523, 3294, 3571, 2502, 1006, 1475, 3526, 2285, 1539, 1095, 1747, 2168, 473, 2432, 1897, 2511, 2861, 2485, 1403, 952, 1327, 1488, 398, 2421, 3246, 596, 186, 3341, 853, 2623, 378, 2504, 190, 2992]. III [1660, 1392, 1481, 2546, 2286, 1368, 1964, 2913, 2380, 21]. Ill [2198, 2086, 3478, 3460, 3471, 3254]. Ill-Conditioned [2198, 2086]. Ill-Posed [3478, 3460, 3471, 3254]. Image [3312, 3182, 3105, 2558, 2404, 2539, 2886, 1545, 1942, 1070]. Images [2198, 2985, 3282, 3388, 3483, 3281, 2964, 3027, 316]. Imaginary [233, 3510, 2478]. Imaging [3263, 3323, 2867, 3451, 2689, 2617]. Imbedded [2279]. Imbedding [894, 429, 293, 528, 2548]. Immersed [2825, 3217]. Immiscible [2605]. Impact [418]. Impedance [2516, 3265, 1920, 3128, 2867, 2930, 2689, 2912, 3525]. Impenetrable [1985]. Imperfect [2126, 2310, 852, 853, 1975]. Imperfection [1890, 2842]. Imperfections [146, 2973]. Impinging [1944]. Implementation [1963]. Implemented [772]. Implications [2841, 594, 491]. Importance [1702]. Important [1829]. Imposed [3536]. Improbability [1703]. Improved [539, 3249, 1351, 2690]. Improvement [7]. Improving [31]. Impulse [2360, 2731, 2681, 183, 1980, 1981]. Impulsive [563, 809, 3305]. Impulsively [1018, 1808]. In-Plane [3075]. Incidence [496, 155, 510]. Incident [2156, 1072, 1378, 2608, 1796, 3247, 3410, 3126]. Incipient [2778]. Inclined [1538]. Inclusion [1072, 2660, 2182, 3521, 1807, 2680, 2659]. Inclusion-Exclusion [2182]. Inclusions [572]. Incoherent [786]. Incomplete [2540, 3455, 646]. Incompletely [1657]. Incompressible [3484, 3247, 1205, 2679, 3231, 719, 896, 2677, 1995, 2204]. Inconsistent [1384]. Incorporate [2599]. Incorporating [2612]. Increasing [2517]. Increment [1549]. Indentation [137]. Independence [228, 1117, 427]. Independent [799, 1548, 1549, 2172, 3028, 1404, 2280, 545, 50]. Independently [589]. Index [469, 1558, 317, 2159, 2211, 56, 68, 3395, 744, 1083, 1513, 1653, 1433, 977]. Indicator [1612, 111]. Indices [1031, 3210, 1712]. Induced [2695, 2173, 760, 3241, 2985, 3105, 1869, 2318, 2249, 3431]. Inducible [158, 1089]. Induction [1078, 1863, 2442, 3115, 3285, 1182]. Inelastic [3434, 2050, 1017]. Inequalities [1430, 1128, 153, 1987, 80, 640, 458, 975, 2215, 216, 983, 490]. Inequality [1636, 808, 1007, 434, 1205, 944, 712, 728, 377, 1306]. Inertia [993]. Infection [3369, 3236, 3041]. Infection-Age-Dependent [3041]. Infectious [2083]. Infectivity [3041]. Infiltration [2604, 2550]. Infinite [2516, 1417, 2971, 262, 1419, 3464, 2925, 591, 549, 1427, 1887, 835, 2510, 3253, 278, 1065, 1629, 1928, 645, 2683, 1751, 550, 1331, 194, 1256, 819, 809, 324, 489, 3137, 3340, 137, 318, 1336, 3341, 3004, 3200]. Infinite- [3340, 3341]. Infinite-Dimensional [3253, 2683]. Infinite-Server [2971, 3137]. Infinitely [2435, 75, 1074, 1904]. Infinity [1736]. Inflated [868]. Inflows [799, 1904]. Influence [3556, 384, 1994, 321, 3572]. Information [616, 3270, 478, 2030, 1464, 592, 787, 1053, 1033, 333, 852, 853, 1402, 2433]. Infrared [2706]. Infrequently [2303]. Inhibiting [1652]. Inhibition [3079, 2682]. Inhibition-Produced [3079]. Inhibitor [2900]. Inhibitory [2234, 3015]. Inhomogeneities [2117]. Inhomogeneity [3345]. Inhomogeneous [3539, 808, 1034, 1002, 851, 1624, 2221, 3277, 2084, 1873, 3510, 2148, 2073, 73, 1244, 2536, 3175, 3309, 3473, 1391]. Initial [2523, 3361, 1423, 839, 2925, 851, 1676, 1168, 3535, 1547, 2563, 2311, 2823, 558, 1127, 2636, 1943, 1094, 1095, 3546, 3397, 3036, 1809, 1662, 2246, 1445, 1058, 88, 3187, 2238, 918, 2541, 2507]. Initial-Boundary [3361, 2925, 1676, 88, 2541]. Initial-Boundary-Value [3187]. Initial-Value [2523, 1547]. Initiated [2813]. Initiation [2441, 3357]. Injection [987, 2410]. Inner [1005, 655, 2098, 493]. Input [3128, 272, 1960, 3570, 3571, 1480, 3531, 2821, 2868, 1946, 639, 2224, 1161, 1805, 3179]. Input-Output [272, 1161, 1805]. Inputs [2398, 1223]. Inspections [634]. Instabilities [2110, 3107, 2150, 1522, 2699, 2571, 2642].

Instability [3155, 431, 3127, 3123, 894, 2484, 2718, 2145, 2646, 2945, 2340, 2678, 2682, 3184, 1674, 2053, 2100, 3405, 3344, 2380]. Instant [1201]. Instantaneous [321, 3549]. Insulation [3535]. Integer [1370, 1457, 1047, 1116, 1466, 1214, 1455]. Integers [1063, 742, 791]. Integrability [736]. Integrable [2939, 613, 3327, 917]. Integral [2198, 536, 714, 2376, 3561, 211, 3265, 663, 3019, 1457, 2130, 3009, 3085, 2735, 494, 97, 1388, 3014, 3274, 366, 461, 1831, 2381, 2481, 2600, 2762, 1294, 2436, 10, 2272, 2031, 1903, 827, 836, 3521, 2983, 2716, 625, 2009, 2232, 713, 965, 1453, 662, 2482, 2793, 3570, 597, 2476, 288, 1095, 3373, 718, 2314, 181, 2230, 814, 2006, 234, 1327, 2657, 24, 1709, 73, 802, 2422, 2977, 122, 3560, 2263, 468, 2402, 263, 708, 826, 3340, 3341, 3363, 1082, 2548, 2640, 759, 1829, 542, 2452, 3331, 3114]. Integrals [637, 258, 1780, 1961, 244, 3159, 2540, 1, 3109, 189, 396, 863, 1679, 2332, 2993, 3177]. Integrate [1960, 2988]. Integrate-and-Fire [1960, 2988]. Integrated [3139, 2809, 2884]. Integrating [446, 3544, 3545]. Integration [608, 613, 1674, 15, 190]. Integrations [3215]. Integrators [3529]. Integro [2773, 908, 3503, 2584, 745, 1818]. Integro-Difference [3503]. Integro-Differential [2773, 908, 2584, 745, 1818]. Integrodifferential [1557, 3513]. Intensities [3575]. Intensity [1203, 1206]. Interacting [2789, 2258, 2724, 2487, 2064, 2069, 134, 679, 1492]. Interaction [1255, 1890, 1269, 2359, 2459, 575, 2755, 1559, 2632, 2392, 3186, 1270, 1300, 3050, 1967, 2642, 2614]. Interactions [2524, 1791, 2179, 1489, 3461, 3581, 1738, 3158, 2442, 3130, 2020, 2543]. Interactive [3199]. Intercepts [1690]. Interest [1382, 458]. Interface [3262, 494, 2550, 3461, 2518, 3320, 3376, 1709, 3093, 951, 323, 2992]. Interfaces [2495, 3119, 2979, 1237, 3375].

Interfacial [3385]. Interference [376]. Interior [921, 2004, 2911, 963, 2365]. Internal [1221, 2302, 2502, 275, 3578, 1238, 3089, 344, 3161, 757]. Interoutput [279]. Interpolated [2538]. Interpolation [2281, 99, 3401, 1878, 2035, 1498, 2225, 2727, 2810, 2144]. Interpretation [306, 1702, 596, 325]. Interpreted [1233]. Intersection [946, 490, 3057]. Interval [288, 2400, 1647]. Intervals [1887, 606, 855, 2400, 2655, 499]. Intricate [2698]. Intriguing [3439]. Intrinsic [3149, 1966, 2885]. Introduced [1206]. Introduction [691]. Invariance [2899, 483, 3512, 611, 1386, 1171, 940, 3363]. Invariant [3244, 2871, 2775, 2776, 2435, 894, 203, 1717, 208, 1597, 3261, 473, 293, 528, 701, 739, 2393, 3313, 3551, 2259, 819, 355, 2548, 2075, 2479, 850]. Inventory [539, 81, 2363, 419, 1802, 742, 1352, 1355, 586, 679, 90, 2200, 1460]. Inverse [460, 3239, 3430, 3325, 1779, 810, 2197, 961, 2331, 2261, 3263, 1061, 657, 1411, 1369, 1419, 1852, 2094, 2307, 2731, 2681, 1052, 2537, 3113, 3255, 1682, 641, 946, 694, 1503, 635, 2223, 2308, 3014, 2896, 3128, 3455, 3379, 2910, 145, 2895, 1991, 1648, 1908, 1380, 1379, 3225, 3257, 1409, 1613, 625, 2506, 2771, 3020, 2599, 3550, 3238, 1389, 1104, 3054, 3562, 3541, 2616, 3498, 1412, 1513, 1666, 2113, 3258, 417, 2462, 3278, 3410, 963, 3055, 3240, 938, 960, 1934, 1433, 2536, 1470, 2617, 263, 1135, 373, 2240, 2330, 3508, 3490, 2912, 2777, 2343, 3314, 3292, 824, 2456, 2504]. Inverses [1257, 1363, 1132, 1183, 627, 224, 1486, 1253, 34, 916, 647, 525, 1029, 1114, 1038, 695, 872, 795, 837]. Inversion [934, 2505, 2519, 3326, 351, 1188, 2262, 2273, 1866, 1545, 2983, 2771, 337, 1664, 2108, 2522, 900, 955, 2466, 943, 3176, 2561, 442, 2730, 989, 1156, 2402, 2093, 2957, 17]. Inversions [2709]. Inverted [3433]. Inverter [2908]. Inverting [884]. Investigation [255, 1375, 108, 792, 678, 2511].

Investigations [3523]. Investment [151]. Inviscid [1221, 2872, 3584, 1956, 882, 1270]. Involve [912]. Involving [3330, 258, 2735, 1216, 1092, 1624, 1, 932, 367, 2346, 662, 3243, 2476, 189, 2704, 2730, 263, 2499, 573]. Invulnerability [672]. Invulnerable [1611]. Ion [2951, 2952, 3583]. Ionic [2722, 2557]. Iron [3021]. Iron-Rich [3021]. Irreducibility [913, 912]. Irreducible [881]. Irregular [348, 2448]. Island [2987]. Islet [3459]. Isochronous [905]. Isola [1742]. Isolas [2032, 2357, 2140]. Isolation [3422, 2578, 2626]. Isometries [44, 872, 493, 842]. Isomorphism [295]. Isoperimetric [808, 1510, 1408]. Isospectral [2521]. Isothermal [2514, 2037]. Isotropic [3476, 2937, 1011, 1658, 3568, 1902, 344]. Isotropy [2434]. Israel [534]. Issue [1265, 1266, 1264, 1263, 1262]. Iteration [519, 2078, 1217]. Iterative [135, 180, 897, 382, 1243, 1570, 1486, 578, 1659, 892]. Ito [768, 2743, 2808, 1212]. IV [1421, 1404].

J [1789, 854, 2244, 37, 988, 1721, 2731, 1546, 1512, 2639, 2252, 3043, 667, 1479, 38, 2539, 1885, 253, 1932, 534, 2309, 1585, 1101, 2405, 2044, 144, 2903, 189, 1232, 1494, 2450, 3183, 1788, 1156, 792, 896, 1361, 1574, 1677, 139, 2100, 2562, 138, 184]. J. [144, 1574]. Jacobi [317, 1000, 87]. Jacobians [16]. Jacobson [865]. Jet [2523, 2119, 3491, 1944]. Jets [3584, 3102, 2764, 1754]. Job [875, 685, 530]. Join [2790]. Joint [953, 1987]. Joints [2663]. Jordan [2402]. Josephson [3104, 2668, 2073, 2133]. Jump [1741, 1922, 2029, 2306, 2683, 1865]. Jumps [2673]. Junction [3104, 162, 2073, 2133, 2417]. Justification [2351].

Kakutani [751]. Kalman [2822]. Kamke [2361]. kappa [3275]. Kármán [1793, 2003]. Kasami [1073]. KdV [2424, 2174, 2392, 3273]. Keller [2828]. Kelvin [2100, 3507, 3184, 2053]. Kepler [3289]. Kernel [3330, 2009, 169, 1139]. Kernels [2721, 714, 1780, 1108, 1106, 10, 2232, 246]. Killing [2061]. Kind [233, 1529, 3014, 356, 2079, 2185, 2422, 2977, 325]. Kinematic [3476]. Kinetic [2461, 2927, 962, 992, 3245, 3246, 2877, 3370]. Kinetics [2841, 1119, 3274, 2351, 2268, 1719, 1821, 1832, 1861, 709, 3149, 277, 3142, 601, 2267]. kink [2642]. Kirchhoff [3020, 3287, 2456, 1336]. Klein [203, 2840, 418]. Knapsack [1163, 1569, 1066]. Knowledge [3239, 2599]. Known [3150]. Knudsen [1231]. Kolmogorov [3303, 3405]. Korn [1205]. Korteweg [2767, 3116, 3143, 2632, 1075, 2826, 3436, 3421]. Kosterlitz [3566]. Kounias [1843]. Kramer [2727, 2810]. Kramer-Type [2727]. Kramers [3250, 2022]. Kronecker [911, 1787, 424]. Krylov [41, 603]. Kuhn [771]. Kulkarni [2236]. Kuramoto [2601, 3298, 2719]. Kutta [2808].

L.P. [1233]. Labelled [2228]. Lag [1388, 746, 394]. Lagerstrom [680, 2679, 1627, 1254, 1949]. Lagrange [686, 58, 880, 2920, 1191, 1319, 487, 2079, 2401, 1499, 5, 2727, 2810]. Lagrangian [2782, 1995, 2204, 2463, 2047]. Laguerre [166, 362]. Lamé [3420, 325]. Lamella [3140]. Laminar [2714, 497, 2218, 1036, 2410, 1601, 2105]. Lanchester [1725]. Lanchester-Type [1725]. Landau [3145, 3008, 3275, 3534]. Landscapes [2820]. Langmuir [1895, 2037, 2149]. Laplace [2997, 138, 460, 1390, 84, 1106, 836, 752, 2865, 769, 1268, 900, 955, 387, 122, 8, 17]. Large [3425, 926, 262, 2789, 1354, 509, 3086, 1628, 3271, 2902, 3509, 1790, 2234, 2012, 3532, 2937, 2724, 1438, 2306, 2165, 1186, 1875, 3469, 3514, 2321, 2449, 2400, 2559, 1815, 1020, 197, 2806, 2255, 868, 793, 3405]. Large-Scale [3086, 3405]. Larger [2052]. Largest [2319]. Laser [2985, 1108, 2245, 3105, 3106]. Laser-Beam-Induced [2985, 3105]. Lasers [2984]. Last [239]. Latent [683]. Lateral [2234, 754]. Lateral-Inhibitory [2234]. Laterally [3450]. Lattice [804, 2923, 3499, 3583, 2319, 3032, 818, 712, 2656, 2780, 2345]. Lattices [1148, 3290]. Laundry [2949]. Laura [3439]. Law [1136, 2789, 2313, 3248, 197, 1028, 3296, 3170]. Lawrence [3145]. Laws [2811, 2368, 2926, 3354, 1017, 3172, 2944, 3215, 3162, 3035, 1062, 1336]. Layer [1255, 1269, 1235, 1708, 666, 2423, 497, 987, 2706, 849, 2766, 1770, 348, 443, 3526, 3156, 3068, 511, 1036, 3568, 2591, 3147, 1206, 92, 1187, 2389, 3264, 1231, 1808, 1246, 868, 1080, 630, 3473, 3525]. Layered [3325, 2384, 2495, 3452, 2273, 3400, 3242, 3031, 2053, 2100, 2279, 3145, 2992]. Layers [3284, 2049, 2932, 3249, 890, 3274, 2118, 2657, 2428, 3540, 1877, 2806, 3161, 1137, 793, 3344]. LBIC [2964]. Lead [2099, 2253, 1738]. Leading [3023, 2098, 2139]. Leakage [3535]. Lean [2441]. Least [904, 657, 649, 811, 2491, 429, 1570, 1384, 3081, 98, 1037, 588, 800, 1154, 86, 1886]. Least-Squares [588, 1886]. Lee [516]. Leibniz [550]. Lemke [1429, 734, 860]. Lemma [584, 2564, 2561, 2402]. Lemmas [825]. Length [1876, 1077, 336, 781, 816, 1631, 815, 1587]. Leonard [3392, 2276]. Leontief [776]. Leontovich [3539]. L’Équation [2024]. Leslie [1946]. Leukemia [3423]. Level [2434, 3118, 346, 3460, 347, 1556]. Levels [3269]. Leverett [2713, 2853, 3366]. Levinson [2780]. levitated [2594]. Lewis [1597, 1574]. Lie [1001, 1002, 2774, 998, 1003, 996, 1005, 1006, 994, 1000, 2477, 1004, 999, 997, 1175, 3266, 995]. Lienard [937, 1043, 633, 1376, 3095, 123, 2614]. Life [3557, 564, 1617]. Lifted [3290]. Lifting [1643]. Ligands [2788]. Light [2875, 2999, 2441, 1712, 3481, 3061, 3468]. Light-Nematic [3468]. Light-Off [3061]. Lighthill [517, 302]. Like [2411, 2767, 2840, 2020, 983, 1855]. Likelihood [2181, 1081, 2724, 1659, 1663, 24]. Limit [2011, 3208, 3364, 2933, 3533, 3160, 1826, 2216, 439, 611, 2669, 1218, 1459, 1800, 1729, 1123, 1499, 697, 3441, 797, 3383, 1515, 102, 3246]. Limitations [2537]. Limited [2086, 131, 2195, 3222, 1335, 2364, 2462, 941]. Limiting [3416, 2991, 2929, 1345, 1623, 3339, 2370, 1971, 683, 3108, 2358]. Limits [3284, 2534, 2517, 1083, 3275, 1739]. Liñán [2348]. Lindemann [2854]. Lindstedt [1523]. Line [3239, 1611, 2925, 1643, 349, 3150, 807, 1011, 2073, 2133, 238, 3001, 321, 1048]. Linear [1014, 2210, 1192, 2804, 1356, 1221, 18, 2142, 294, 1195, 1047, 453, 240, 2329, 1923, 1411, 2435, 129, 228, 3123, 1898, 1398, 267, 946, 3423, 2169, 899, 1315, 99, 3062, 591, 1031, 207, 61, 308, 2231, 811, 2333, 1059, 1593, 114, 247, 822, 1115, 2481, 176, 272, 83, 434, 1344, 502, 1345, 1003, 875, 2874, 625, 41, 2867, 544, 1454, 1133, 3395, 1426, 1993, 1429, 1338, 768, 2043, 3324, 1455, 1486, 1539, 1383, 1253, 2743, 2362, 1094, 1095, 2186, 2990, 473, 936, 1004, 830, 1162, 1232, 975, 1037, 1075, 1114, 2006, 1038, 1213, 321, 974, 618, 1451, 2490, 492]. Linear [1610, 427, 464, 697, 963, 1414, 2450, 2686, 848, 758, 1439, 328, 859, 199, 1087, 730, 594, 470, 377, 819, 3318, 45, 404, 7, 186, 70, 433, 3508, 652, 3363, 1415, 1618, 991, 436, 1843, 231, 428, 3266, 159, 13, 140, 4, 59, 139, 226, 684, 1805, 2605, 51, 1554, 477, 3382, 1670, 3524, 2544, 824, 1900]. Linearised [2439]. Linearity [842, 2106]. Linearizable [2633]. linearization [3287]. Linearized [3151, 1296, 2608, 2771, 1030].

Linearly [158, 1089, 3162, 518]. Lines [160, 3083, 2304, 1054, 918, 2279, 2541, 1900, 3082]. Linkage [2664]. Linked [1008]. Liouville [874, 27, 1613, 654, 348, 454, 1875, 749, 2727, 2810, 2777]. Liouville-Type [874]. Liquid [2848, 2523, 363, 3584, 2875, 3072, 1030, 2585, 2518, 1227, 809, 3017, 1944, 2409, 2562, 2469, 32, 1754, 2623]. Load [2837, 1022, 106, 1010, 2503, 1122, 126]. Load-Preferential [2837]. Loaded [2131, 772, 1018]. Loading [885, 645, 1955]. Loadings [1093]. Loads [1909, 3527, 2353, 2309, 2266, 2890, 2492]. Local [1722, 3507, 2941, 2897, 3352, 2814, 3353, 3582, 954, 576, 2117, 2943, 818, 2514, 2445, 2995, 2566, 2415, 2358]. Localization [2143, 2989, 3462, 3380]. Localized [1322, 2763, 3520, 3115, 3093, 3011, 3487]. Locally [2795, 1240]. Locate [257]. Located [1203]. Location [250, 405, 326, 471, 297, 1490, 3339, 1009, 1123, 1445, 656, 1775, 1776]. Locational [480]. Lock [2454, 2588, 2955]. Locked [2418, 2141, 2454, 2955]. Locking [838, 988, 2914, 2737, 3511]. Locking-In [988, 838]. Log [24, 274]. Log-Periodic [274]. Logarithmic [1, 2599, 2183, 189, 2330, 3012]. Logarithms [353]. Logical [198]. Logistic [2007, 1717, 1442]. Long [1413, 3127, 3445, 3461, 1311, 3099, 2664, 1844, 1681, 2699, 1093, 3065, 3441, 2015]. Long-Channel [3065]. Long-Wave [2699]. Long-Wavelength [3445]. Longitudinal [2516, 623]. Look [1941]. Loop [3538, 2418, 1633, 3431, 2454, 1336]. Loops [2141, 496, 2588, 2955]. Lorentz [203, 208]. Lorentz-Invariant [203, 208]. Lorenz [2202, 3026, 1726]. Loss [3005, 1502, 1033, 3500, 3003]. Lossless [343, 918]. Lossy [3113, 1334, 3510, 1332]. Lost [2625]. Lotka [2501, 2233, 2406, 3440, 2671, 3551, 1783, 3393]. Love [3439, 3020, 3287, 759]. Low [3428, 3561, 2524, 680, 3399, 3424, 1642, 314, 1820, 1985, 2510, 1825, 3333, 3450, 2585, 986, 2194, 2237, 2117, 3376, 2968, 187, 1206, 3060, 3188, 1231, 2677, 1267, 2742, 2543, 3213]. Low- [3333, 2968]. Low-Chirality [3561]. Low-Frequency [3399, 1820, 3450, 2585]. Lower [804, 1119, 1562, 1007, 1767, 614, 764, 2511]. Lowest [1804, 1656, 1408]. LSC [881]. Lubrication [3262, 2134, 1112]. Lubrication-Type [3262]. Lumping [1145, 3370]. Lyapunov [1636, 2339, 2637, 1040, 1277, 465, 1051, 1128, 3260, 3440, 3395, 2587, 2743, 1368, 2990, 3291, 2490].

M [3030]. M. [605]. Macaulay [1518]. Macdonald [1457]. Mach [2146]. Machine [875, 1868]. Machines [3080]. MacMahon [1079]. macromolecular [1897, 1896, 1964]. Macroscopic [3565]. Magma [2576]. Magnetic [363, 1067, 3323, 1831, 2800, 238, 605, 1054]. Magneto [1054]. Magneto-Hydrodynamic [1054]. Magnetogasdynamic [320]. Magnetohydrodynamics [2800, 3232]. Magnetohydrostatic [409]. Magnetoresistance [387]. Magnitude [2282]. Maintainability [2224]. Maintained [100]. Major [241]. Majority [296]. Malik [3360]. Man [3319]. Man- [3319]. Managed [3537]. Maneuvering [2468]. Manifold [2601, 2733, 2765]. Manifolds [1294, 2436, 2463, 2263, 708]. Many [2277, 2792, 3028, 2597]. Map [2831, 2406, 3537, 3290, 3287]. Mapping [2142, 651, 2819, 565, 235, 1925, 2538]. Mappings [2871, 2602, 2408, 1298]. Maps [3569, 3166, 2666, 1647, 2091, 411]. Marginal [1789, 1769, 2130]. Marin [1843]. Marketing [1556]. Markoff [1670]. Markov [3396, 929, 3179, 143, 3338, 2081, 292, 508, 114, 2759, 3243, 2306, 928, 2587, 3469, 2054, 181, 1037, 681, 1751, 1904, 2747, 704, 352, 710, 652, 639, 2224, 477]. Markov-Modulated [3396]. Markovian [799, 2328, 1223, 2673]. Markovian-Independent [799]. Marshak [1846]. Martingale [1192, 2553, 1133, 1482]. Mass [2906, 3424, 2603, 2849, 987, 1094, 1095, 1268, 3503, 2879, 3283]. Mass-Spring [2879]. Massless [2547]. Master [1079, 1496]. Matched [1988, 3212, 1331, 2365]. Matching [1180, 3249, 457, 2972]. Matchings [1633]. Material [3005, 3509, 3133, 2486, 2589, 3526, 2132, 1709, 3525]. Materials [3412, 1002, 3404, 1315, 2942, 3535, 1311, 1348, 2999, 1021, 3047, 3465, 3296]. Math [1789, 854, 2244, 988, 1721, 2731, 1546, 1512, 2639, 2252, 3043, 667, 1479, 38, 2539, 1885, 253, 1932, 534, 2309, 1585, 1101, 2405, 2044, 144, 2903, 1232, 2450, 3183, 1788, 1156, 792, 896, 1361, 1574, 1677, 2100, 2562, 138, 184]. Mathematical [3419, 1520, 831, 2871, 1911, 2519, 3453, 680, 2250, 2269, 1326, 2984, 2621, 3044, 2847, 2577, 1489, 1831, 1467, 1325, 80, 3054, 3580, 2148, 76, 962, 217, 40, 3334, 2382, 3578, 2282, 1710, 174, 2606, 2883, 771, 1896]. Mathematics [1265, 1266, 1264, 1263, 1262]. Mathieu [1801, 1071, 463, 1332, 400, 3432]. Matrices [1297, 1040, 343, 469, 731, 1401, 750, 326, 615, 285, 1416, 459, 1188, 36, 689, 661, 25, 1152, 1609, 1648, 1275, 467, 1363, 1132, 1065, 1379, 627, 56, 68, 1110, 884, 1585, 2638, 224, 2446, 311, 640, 219, 479, 87, 1410, 1412, 647, 525, 943, 1029, 1513, 975, 690, 142, 952, 695, 194, 354, 1699, 960, 193, 1318, 1346, 1654, 355, 879, 795, 1407, 426, 350]. Matricial [1031, 898]. Matrix [2721, 3330, 934, 1180, 1804, 526, 801, 911, 1277, 1558, 1419, 1216, 641, 1696, 183, 61, 1495, 602, 1113, 1257, 912, 1637, 840, 67, 243, 429, 653, 253, 136, 337, 1250, 1426, 889, 1664, 1042, 744, 1373, 1284, 47, 458, 1324, 1083, 1323, 329, 1507, 552, 424, 417, 55, 277, 48, 967, 1699, 938, 2320, 512, 71, 270, 1135, 1153, 1211, 1361, 295, 3266, 671].

Matrix-Valued [1696]. Matroid [813]. Matroids [1032]. Matter [290]. Max [58, 57, 1474]. Max-Min [58, 57]. Maxima [686]. Maximal [1661, 3231, 2938, 223, 614]. Maximally [248]. Maximization [1441, 422, 434, 1285]. Maximizing [456, 3577]. Maximum [1180, 1803, 2181, 1081, 607, 1276, 626, 787, 909, 1637, 68, 3172, 2599, 2724, 3080, 1444, 1540, 1659, 1663, 1055, 774, 395, 66, 757]. Maximum-Likelihood [1659, 1663]. Maxwell [3574, 3539, 2963, 3321, 3573, 3205, 3477, 2753]. May [2099, 2205, 3041, 3392, 2276]. McShane [802]. Mean [907, 1803, 3163, 1978, 3202, 116, 2052, 811, 2038, 2652, 1539, 725, 3048, 2442, 724, 101, 2614]. Mean-field [3202, 2614]. Mean-Square [811]. Meandering [2105]. Means [1052, 3433, 166, 5, 1292]. Measurability [1038]. Measure [912, 1925]. Measure-Preserving [1925]. Measured [2194, 2237, 1201, 2742]. Measurement [3430, 2235, 2867, 3550, 1963]. Measurements [1920, 2403, 2387]. Measures [2241]. Measuring [3180, 1504]. Mechanical [1630, 2935, 3532, 1500, 1986, 2528, 1580]. Mechanics [831, 919, 3443, 2381, 1025, 739, 997, 1009, 1229]. Mechanism [2854, 2347, 3209, 2045]. Mechanisms [3584, 1491, 2209]. Mechanoreceptor [2685]. Media [3428, 3256, 412, 599, 3325, 2866, 3293, 2472, 2471, 1876, 2368, 1773, 2567, 1961, 3505, 2815, 3452, 2260, 3456, 3235, 3271, 2593, 2231, 2273, 3463, 3411, 1989, 3475, 1599, 2889, 1205, 3333, 3450, 1871, 2371, 3006, 2108, 3099, 3242, 1268, 3158, 2791, 2654, 3031, 756, 3135, 3234, 783, 3362, 414, 2369, 3175, 2440, 1711, 2605, 3233, 3297, 2431, 2432, 2799, 2992]. medians [1776]. Mediated [3359]. Medical [853]. Medium [3539, 3315, 2607, 3018, 3416, 2384, 2595, 696, 2604, 3013, 1584, 3224, 2234, 3277, 1873, 3225, 2713, 1334, 3510, 3242, 962, 1094, 1095, 3568, 2473, 2485, 726, 820, 3060, 3189, 3188, 1300, 127, 2536, 2829, 3049, 2020, 3309, 2240, 3279, 1998, 3183]. Mel′nikov [2602, 2390]. Mellin [263, 636, 138, 8]. Melnikov [3302]. Melting [164, 32]. Members [1580]. Membrane [808, 2951, 2952, 2494, 2843, 1500, 1960, 133, 92, 327, 557]. Membranes [1718, 2521, 33, 2851, 868, 1080]. Memory [2565, 1315, 3307, 1311, 2322]. Memoryless [913]. Merged [1235]. Merits [1914]. Meromorphic [343, 1340, 10]. Metal [2910, 2740, 2594]. Metaphor [3365]. Metastable [2303, 3161, 3283]. Method [686, 1310, 3425, 3244, 431, 1277, 1795, 1997, 956, 1978, 3515, 2906, 517, 135, 2763, 519, 1125, 2750, 3388, 817, 890, 3078, 3134, 1621, 1624, 946, 2924, 1503, 2308, 3423, 302, 3553, 3454, 3274, 382, 899, 2927, 1217, 2355, 117, 2896, 61, 399, 3555, 1215, 1243, 1059, 192, 2558, 1025, 1294, 1974, 434, 265, 42, 413, 2864, 3524, 385, 3573, 2182, 2983, 309, 2287, 41, 603, 1878, 665, 713, 1816, 2599, 1939, 313, 1942, 3477, 2865, 1375, 1429, 2863, 93, 1383, 3209, 2208, 485, 2044, 1990, 2661, 3397, 293, 1075, 3176, 2654, 698, 1806, 3394, 3199, 1123, 3427, 1120]. Method [182, 1612, 3088, 1810, 1865, 992, 2297, 668, 2263, 518, 2111, 352, 420, 31, 2240, 2327, 1843, 852, 2548, 2055, 3544, 3545, 1392, 2828, 2343, 3170, 3205, 3213, 2144]. Methods [3164, 3489, 1555, 1001, 1002, 18, 2838, 998, 2261, 307, 2181, 3182, 455, 1003, 3521, 1842, 1745, 2141, 996, 2029, 1005, 1006, 994, 1000, 2108, 1486, 3451, 312, 2344, 1004, 999, 2230, 997, 2364, 2808, 1809, 167, 1600, 3035, 432, 2758, 2246, 963, 388, 358, 1766, 1517, 2740, 995, 1995, 2204, 3254, 1746, 1747]. Metric [434, 516]. Metrics [2169]. MHD [1358, 1184, 282]. Micro [1467]. Micro-Organisms [1467]. Microbial [3504]. Microcirculation [1280]. Microelectrodes [3373].

Microenvironment [3583]. Microgeometries [2457]. Microorganisms [1954]. Microscope [3207]. Microstrip [1988]. Microstructure [3090, 3144, 1025, 2290]. Microstructured [3414]. Microstructures [2533]. Microwave [2739, 3006, 3322, 3047]. Microwaves [3423]. Middle [2427]. Midgap [3463]. Migration [3210, 3333]. Mild [3038]. Mild-Slope [3038]. Min [58, 57, 1474]. Min-Cost [1474]. Miner [306, 596]. Mineral [2066]. Mines [2387]. Minima [261, 175, 587]. Minimal [1569, 2871, 505, 183, 1509, 254, 1320, 496, 155, 833, 1087, 1173, 1153, 369]. Minimality [2225]. Minimally [1219, 852, 853]. Minimax [787, 909, 1115, 1339, 1479, 3474, 264]. Minimization [165, 482, 1596, 858, 2904, 2385]. Minimizations [1532]. minimizers [3267]. Minimizing [653, 742, 504]. Minimum [1210, 1579, 403, 626, 2922, 1418, 365, 2400, 510, 1010, 728, 377, 1290, 101, 2323, 738, 1185]. Minimum-Weight [1010]. Mining [2066]. Miroshin [2167]. Miscible [3100]. Mite [2755]. Mixed [2811, 799, 2368, 527, 1047, 125, 666, 1613, 1144, 713, 3215, 2230, 968, 344, 2109]. Mixed-Integer [1047]. Mixing [1077, 3351, 3413, 2203, 3347, 3540]. Mixture [893, 2979, 3045, 1853, 1659, 1663, 2699]. Mixture-Fraction [3045]. Mixtures [1629, 3385, 1617]. Mobility [2242]. Modal [2114, 606, 855]. Mode [1738, 3130, 1090, 2203, 984, 1656, 2989, 2068]. Model [2919, 2383, 831, 3241, 3506, 1322, 3434, 3533, 2773, 3104, 2515, 680, 2250, 2269, 3557, 1649, 1721, 2986, 2414, 3563, 1326, 3101, 3202, 3392, 3124, 1645, 1642, 2755, 2233, 2213, 2751, 2611, 3583, 3435, 2621, 3067, 3008, 3534, 3077, 2847, 2882, 1777, 1280, 3443, 3132, 2612, 2425, 3074, 3072, 3513, 2125, 3414, 3535, 3518, 1345, 2152, 1919, 893, 1226, 2458, 2207, 2017, 1825, 2685, 2900, 2918, 3492, 2679, 149, 2229, 3333, 1325, 1564, 1500, 112, 2301, 1098, 3054, 3038, 3236, 2833, 2790, 3571, 3486, 2150, 2669, 3222, 1874, 2270, 1714, 2475, 962, 685, 2649, 3514, 1627, 1469, 2528]. Model [1924, 3481, 2788, 530, 1590, 1036, 1774, 1037, 2670, 563, 2064, 578, 1681, 3459, 3149, 2747, 2382, 1367, 3503, 2417, 1726, 1768, 1254, 3578, 1511, 1756, 3567, 3465, 195, 3065, 2711, 1288, 1967, 3108, 3015, 2305, 151, 1477, 1894, 2851, 1671, 2606, 1461, 2498, 1425, 1884, 1748, 2883, 3358, 3349, 2053, 2100, 2562, 2469, 3003, 1556, 1544, 2023, 2541, 3171, 1670, 3275, 3233, 1896]. Modeled [2671, 2549]. Modeling [3042, 2607, 1930, 2985, 3319, 2984, 3008, 3107, 2553, 3537, 3422, 2278, 3522, 2492, 1883, 3088, 3087, 3216, 2083, 1449, 1677, 2740, 2878, 3511]. Modelling [2085, 3248, 2171, 2930, 2584, 3191]. Models [3419, 929, 3090, 3144, 2465, 3125, 3141, 2360, 2501, 2500, 3252, 3123, 3269, 1557, 1918, 2630, 2160, 3218, 1273, 1178, 2935, 1489, 1599, 2275, 3440, 1872, 2535, 2363, 3351, 1960, 3522, 3523, 3190, 2921, 3097, 2488, 3209, 2054, 2830, 2943, 3355, 1786, 1751, 2203, 1263, 2965, 3390, 3585, 2665, 1355, 2901, 1949, 2334, 2931, 1946, 2257, 652, 3002, 3408, 477, 3470, 3145, 3179]. Modern [1307, 1725]. Modes [1669, 1890, 3463, 3450, 2513, 238, 2891, 2642]. Modified [233, 519, 943, 3055, 1810, 1884]. Modulated [2508, 1525, 1193, 2287, 3396]. Modulation [1873, 3472]. modulational [3344]. Modulations [2174]. Modules [26]. Moduli [3075, 414]. Modulus [2610]. Mole [2835]. Molecular [2933, 3533, 1481, 307, 2164, 996, 1720]. Molecules [1838]. Moment [2210, 3395, 3482, 3397, 2654, 3291, 420]. Moment-Difference [2654]. Moments [1578, 1951, 2420, 1549, 2214, 2488, 2580, 1383, 1372, 3169, 2794, 3030]. Momentum [1295, 1703]. Monge [3449]. Monochromatic [495]. Monomolecular [2806]. Monotone [1069, 1745, 3575]. Monotonicity [1199, 810, 1945, 2520, 3071]. Monte [307, 1702, 62, 35, 2385, 144]. Moore [3506, 1052, 1379, 916]. Morphological [2870, 2699, 2571]. Morphologically [2403]. Morphology [2871]. Morphometric [1528]. Morrison [1574]. Mortality [2415]. Mortell [1926, 1812]. Mosaic [3081]. MOSFETs [3065]. Motility [3504]. Motion [2016, 3262, 1640, 3203, 3407, 3466, 1620, 582, 2936, 2781, 3371, 3461, 2404, 2539, 1869, 1948, 2304, 2585, 2960, 2073, 3199, 1902, 2977, 1445, 1446, 915, 3308, 1111, 1130, 1541, 173, 2635, 3161, 1808, 3058, 3299]. Motion-Blurred [2404, 2539]. Motion-by-Curvature [2936]. Motion-Induced [1869]. Motions [2173, 546, 1794, 1885, 20, 1208]. Motivated [210]. Movement [3156]. Moving [1428, 2153, 2702, 2870, 3173, 2872, 841, 834, 2423, 2948, 2766, 3158, 1447, 1899, 2430, 3169, 3289, 2327, 1204, 1711, 2047, 793]. Moving-Average [2870]. MR [2244, 2639, 3043, 2539, 1885, 2309, 2405, 2903, 2450, 3183, 1574, 2100, 2562]. Muller [1767]. Multi [2046, 1090, 1868, 1639]. Multi-Machine [1868]. Multi-Mode [1090]. Multi-Parameter [2046]. Multi-Phase [1639]. Multichain [292]. Multicommodity [532, 679]. Multicomponent [2876, 644]. Multidegree [2989]. Multidegree-of-Freedom [2989]. Multidimensional [3157, 1106, 430, 909, 1948, 3083, 3082, 3457, 3444, 2579, 610, 3318, 1212]. Multigroup [1501, 2613, 395]. Multijunction [2787]. Multilane [3522, 3523]. Multilayer [3280, 2341, 2502]. Multimode [3242]. Multinomial [604]. Multiphase [3355, 1244]. Multiple [2574, 1778, 1519, 1660, 478, 2529, 3563, 2170, 2925, 650, 1035, 2570, 1918, 3021, 2099, 1981, 2760, 3116, 1982, 2918, 1719, 1760, 496, 2172, 524, 1622, 288, 2661, 1731, 1800, 2034, 3110, 2559, 510, 2972, 2643, 2445, 1164, 2888, 2787, 3226, 3545, 3122]. multiple-delayed [3122]. Multiple-Group [2918]. Multiple-Scale [2170, 2925]. Multiple-Valued [478]. Multiplexing [3581]. Multiplication [2253, 653, 533]. Multiplications [653]. Multiplicative [1799]. Multiplicities [1005, 1006]. Multiplicity [1317, 1779, 2395, 811, 2938, 2658]. Multiplier [58, 638]. Multipliers [487]. Multiply [881, 2912]. Multipoint [528]. Multipole [1851]. Multiprocessing [406]. Multiproduct [151]. Multiprogramming [3029]. Multistability [3269]. Multistage [151]. Multistep [3274, 3146, 2246]. Multitime [1943]. Multitype [2837, 3000]. Multivalued [537]. Multivariable [1552]. Multivariate [439, 2259, 1933, 652, 1763]. Multiwave [1125]. Mumford [3182, 3267]. Mutants [2014, 2573]. Mutation [1919, 1173, 2573]. Mutual [616, 2153, 592]. Mutualists [2361].

Nagumo [3227, 1982, 2161]. Narrow [2951, 2952, 2709, 170]. Narrowly [3139]. Nash [1097]. Natural [3425, 1859, 2048, 3418]. Natural- [1859]. Nature [2086, 2202]. Navier [3484, 547, 2823, 1635, 1186, 154]. Near [3208, 3364, 3159, 2127, 3435, 2608, 2126, 2542, 2545, 2546, 2846, 2513, 2816, 2903, 511, 1036, 1806, 2634, 2977, 2993, 3177, 2894, 2058, 2888, 3057, 2640, 2525]. Near- [2846]. Near-Field [2608]. Near-Identity [1806]. Near-Singular [2640, 3177]. Near-Threshold [3364]. Near-Wall [3435]. Nearly [1413, 2527, 2863, 2392, 2748, 2825, 3327]. Necessary [3558, 670, 653, 2080, 768]. Negative [1578, 2379, 1529, 317, 1848, 780, 1475, 1372]. Negligibility [1013]. neighborhood [1931]. Neighbourhood [2213]. Nelson [639]. Nematic [3468]. Neoperiodic [2512]. Neosteady [2512]. Nernst [3335, 3334]. Nerve [760, 1683, 2360, 3221, 1980, 1981, 2373, 1960, 2757, 2116, 2631]. nervous [1964]. Nested [535]. Net [2738]. Nets [660, 1790, 1856]. Network [3241, 3404, 2729, 2831, 36, 3333, 900, 2884, 3110, 681, 2671, 3149, 922, 3316, 350, 1606, 1668, 3470, 1775, 1776]. Networks [3453, 1473, 1579, 1069, 117, 2630, 1860, 822, 1478, 22, 67, 2827, 569, 1468, 328, 812, 121, 1749, 1840, 1290, 2295, 2737, 1336, 1556]. Neumann [2831, 3287, 313, 2317]. Neural [3241, 2414, 1790, 2160, 2684, 2671, 3149, 1883, 1840, 2738, 3316, 3470]. Neurodynamics [1520]. Neuronal [1856, 3194, 1748]. Neurons [2988]. Neutral [1259, 990]. Neutron [1428, 2613, 382, 834, 1353, 1451, 1587, 395, 1711, 542]. Nevanlinna [1696]. Newcomb [466]. Newton [519, 3394]. Newtonian [2100, 1235, 657, 2824, 1703, 2053]. Niche [2179]. Ninomiya [1073]. NLS [2720]. No [2458, 300, 3393, 1789, 854, 2244, 37, 988, 1721, 2731, 1546, 1512, 2639, 2252, 3043, 667, 1479, 38, 2539, 1885, 253, 1932, 534, 2309, 1585, 1101, 2405, 2044, 144, 2903, 189, 1232, 1494, 2450, 3183, 1788, 1156, 792, 896, 1361, 1574, 1677, 139, 2100, 2562, 138, 184]. Node [1918, 1336]. Noise [2637, 3293, 3270, 1595, 1424, 1682, 2418, 422, 2652, 2194, 2237, 2489, 2165, 2990, 1963, 1162, 1232, 2374, 2822, 2340, 187, 3181, 2885, 103, 2742, 2385, 3291]. Noiseless [1374]. Noisy [3281, 401, 3165, 3397, 2294, 2738]. Non [2328, 3192, 2939, 2031, 2264, 2824, 1075, 19]. Non-Abelian [2031]. Non-Fickian [3192]. Non-Hamiltonian [2939]. Non-Markovian [2328]. Non-Newtonian [2824]. Non-Product [19]. Non-Self-Adjoint [2264]. Nonadaptive [3429]. Nonadiabatic [2459, 2496, 2412, 2527, 2894, 1892]. Nonadjacent [791]. Nonadjoint [486]. Nonanticipative [183]. Nonautonomous [779, 470]. Noncentral [439]. Noncircular [949, 3487]. Nonclassical [3134, 2247]. Noncommutative [3330]. Nonconservative [3216]. Nonconstant [3502]. Nonconvex [1198, 876]. Noncooperative [1097, 734]. Noncritical [367]. Nondifferentiable [1420]. Nondiscrete [1387]. Nondispersive [2020]. Nondissipative [1054]. Nonempty [1463]. Nonexistence [317, 1318, 920]. Nongeneric [3538]. Nonhomogeneous [1911, 2887, 153, 2889, 1842, 341, 3050, 1415, 2224, 1425]. Nonidentically [818, 2280]. Nonisolated [504]. Nonisothermal [2607, 2149]. Nonisotropic [3548]. Nonlinear [2695, 3144, 2642, 980, 1126, 2997, 2494, 2893, 2848, 1969, 956, 1961, 500, 631, 1978, 2496, 3249, 2508, 2838, 3515, 2013, 2155, 777, 1686, 942, 1852, 2735, 1649, 1721, 2947, 509, 2151, 2550, 1119, 1092, 3255, 519, 897, 1579, 1399, 2170, 2925, 748, 3124, 851, 2924, 650, 1035, 1525, 3329, 1388, 157, 1628, 16, 3368, 160, 2484, 415, 430, 930, 720, 1217, 1069, 383, 591, 1121, 1308, 1945, 2560, 1243, 784, 2967, 3079, 2453, 970, 1831, 822, 723, 3228, 2425, 923, 265, 487, 676, 42, 1873, 3260, 2152, 1381, 1450, 2864, 677, 827, 113, 1872, 2480]. Nonlinear [2190, 2722, 2112, 2962, 3512, 2287, 2467, 2641, 2733, 2944, 6, 733, 1422, 558, 2143, 3162, 2532, 711, 947, 2405, 1015, 1172, 3204, 443, 2043, 3259, 1138, 2051, 3457, 524, 1875, 2254, 1957, 1846, 1567, 218, 1691, 2088, 590, 1259, 2054, 2840, 416, 749, 2528, 2442, 2943, 1385, 335, 779, 570, 1828, 1264, 98, 1249, 2945, 2132, 1973, 338, 815, 1359, 1888, 563, 598, 2064, 234, 1024, 1813, 2087, 2187, 698, 2634, 421, 1809, 1404, 1327, 555, 1600, 2645, 1120, 1451, 498, 2514, 2040, 2374, 2822, 2862, 2618, 569, 2768, 3093, 1905, 148, 1360, 1768, 2129, 1058, 1256]. Nonlinear [2697, 1511, 812, 1755, 1788, 92, 475, 518, 1641, 562, 1288, 1290, 3089, 2305, 717, 1112, 708, 1477, 1894, 3508, 332, 513, 1200, 1598, 229, 282, 2498, 1855, 577, 1420, 2989, 147, 3187, 239, 79, 375, 3193, 2562, 2469, 557, 868, 793, 3226, 2541, 1391, 3170, 1523, 1403, 2380, 1818]. Nonlinearities [817, 1688]. Nonlinearly [1778, 2058]. Nonlinearly-Coupled [2058]. Nonlocal [3516, 3548, 3443, 3143, 3554]. Nonmonotone [674]. Nonmonotonicity [1139]. Nonnegative [408, 1275, 912, 1065, 56, 68, 1426, 1455, 1699, 554, 477]. Nonnegativity [526]. Nonnormal [1837]. Nonoccurrence [1531]. Nonohmic [3047]. Nonoscillation [902, 746, 1190]. Nonparametric [1827, 1994, 1490, 606, 1680, 1864]. Nonpatchy [2429]. Nonperishable [1352]. Nonranked [1667]. nonreflecting [3154]. Nonresonance [762]. Nonscattering [2107]. Nonseparable [2860, 742, 2861, 2913]. Nonsingular [881]. Nonsteady [2510, 2624]. Nonstrictly [2607]. Nonsymmetric [714, 2127, 2126, 3091, 2658]. Nontrivial [1868, 3468]. Nonuniform [3387, 2427, 2231, 2060, 2943, 2497, 757, 2511]. Nonuniformity [3512]. Nonuniformly [3357]. Nonunique [1601]. Nonuniqueness [933, 3258]. Nonzero [1408]. Norm [983]. Normal [907, 1040, 2807, 1222, 1614, 2135, 1110, 2921, 1659, 1414, 2450, 1122, 879]. Normality [1864]. Normally [1378, 2608, 2479]. Normed [876]. Norms [1297]. Norton [2028]. Nosed [3398]. Note [1789, 1769, 1578, 801, 258, 731, 371, 517, 547, 1579, 2119, 94, 3152, 399, 689, 661, 1350, 1089, 162, 1381, 2311, 865, 627, 1223, 1615, 2889, 41, 1589, 722, 1456, 2708, 2029, 1583, 1952, 2518, 2183, 1962, 1011, 1893, 1268, 2840, 144, 2825, 189, 2826, 23, 3176, 424, 974, 968, 450, 778, 1949, 1977, 886, 825, 1983, 743, 489, 652, 1574, 2635, 1225, 1843, 39, 1400, 1068, 671, 2345]. Notes [3232]. Notions [1916]. Novel [2308, 3070, 2343]. Noyes [1226]. Nozzle [3564]. NRZ [3580]. Nuclear [1119, 1247, 1367, 1449, 1677]. Nucleation [3022]. Nuclei [1283]. Nucleolus [1417, 658, 858, 1139, 474, 1405]. Nucleoluses [1096]. Null [3151]. Number [1163, 2339, 2637, 804, 1269, 680, 1642, 158, 1089, 2299, 653, 2367, 496, 3080, 1186, 1066, 1455, 2990, 1239, 1371, 553, 271, 510, 712, 2490, 3410, 3060, 3071, 3188, 859, 2014, 2498, 2677, 941, 3213, 3004, 3200]. Numbers [669, 2524, 2789, 1529, 3424, 1165, 3338, 2510, 2867, 197, 2901, 2543]. Numerical [2811, 3010, 1996, 2796, 2946, 3299, 37, 3446, 2223, 3386, 2917, 3229, 3555, 2660, 3367, 3133, 2895, 1995, 2204, 659, 2, 836, 3521, 1842, 3215, 313, 480, 2960, 3523, 3562, 1127, 3209, 520, 698, 1809, 167, 3035, 1663, 2128, 1674, 3462, 2068, 1554, 3213, 2380, 1747, 1964]. Numerically [2695]. Nusselt [3071]. Nutrient [2250, 1467, 3390, 3585]. Nutrients [1954]. Nyquist [2922].

Object [3430, 1899, 682, 120]. Objective [532, 1618, 378]. Objects [3429, 2151]. Oblate [1927, 2072, 2156, 2907, 1968]. Oblique [2558, 1113, 1942, 329, 671]. Observability [971]. Observation [239]. Observations [2004, 2386, 297, 1482]. Observed [1706, 2303]. Observers [2534]. Obstacle [3561, 1920, 835, 3543, 3498, 3526, 3186]. Obstacles [2107]. Obtaining [904, 1659, 1107]. Occupancy [385, 105]. Occurrence [2969]. Occurring [1989, 2938, 1285]. Ocean [1273, 1869, 1238]. Oceanic [2782]. Odd [1619]. ODEs [3499]. Odyssey [2986]. Off [1296, 2206, 3061]. Off-Centre [2206]. Offshore [1955]. Oil [2484, 2065, 2606]. O’Malley [1941]. On-Surface [3158]. One [3430, 2429, 3092, 1519, 1660, 469, 2796, 2050, 3480, 3100, 2213, 2801, 2179, 1918, 959, 3400, 334, 948, 1241, 2895, 1866, 1991, 3103, 1908, 2827, 3550, 2194, 2237, 389, 2818, 688, 1101, 3280, 2303, 1334, 2341, 2502, 2850, 3566, 3032, 3196, 1691, 3381, 3491, 726, 1912, 2447, 2374, 3087, 1134, 2445, 2665, 2444, 707, 1967, 2617, 2628, 2806, 1220, 757, 1639]. One- [2796, 2665]. One-Dimensional [2429, 3480, 3100, 3400, 948, 1241, 2895, 1866, 1991, 1908, 2827, 2194, 2237, 2818, 688, 1101, 3280, 2303, 1334, 2341, 2502, 3032, 1691, 3381, 3491, 726, 1912, 2447, 2374, 3087, 1134, 2445, 2444, 2617, 2628, 2806, 1220, 757]. One-Parameter [3092]. One-Phase [707]. One-Prey [2179, 1967]. One-Sided [2050, 3566]. Only [3426]. Onset [3220, 3286]. Open [1258, 566, 2114, 3261, 1656, 276]. Open-Ended [1258]. Opened [1807, 344]. Operation [1631]. Operational [413]. Operations [922]. Operator [919, 871, 1624, 3224, 1296, 3478, 2996, 1486, 1891, 739, 2561, 618, 1181, 3131, 2279, 3114, 2975]. Operator-Valued [2279]. Operators [702, 1176, 452, 206, 294, 519, 635, 1630, 869, 2600, 2031, 1759, 1257, 1772, 1993, 576, 1253, 2264, 163, 632, 298, 830, 1114, 1038, 1531, 2259, 594, 1107, 1992, 1055, 39, 438, 684, 3331]. Oppenheimer [3383]. Opposite [901, 3527]. Oppositely [489]. Optic [1998]. Optical [2505, 1258, 422, 3562, 3580, 3259, 3474, 2531, 3266, 3448]. Optics [2325, 2860]. Optimal [2166, 1014, 3489, 3458, 2429, 2457, 2610, 1195, 81, 3557, 1620, 2651, 1313, 1076, 2078, 437, 667, 873, 2065, 677, 841, 781, 3343, 2970, 1133, 2194, 2237, 2088, 1279, 419, 3075, 2226, 1698, 1746, 1747, 1447, 3565, 2066, 914, 2006, 3488, 1026, 1673, 682, 3347, 1304, 1060, 1488, 1445, 1446, 644, 2656, 3096, 792, 678, 856, 1618, 1626, 428, 1460, 2742, 788]. Optimality [1339, 1479, 1406, 1019, 2363, 1475, 1355, 876, 717, 90]. Optimally [1611]. Optimization [2243, 956, 926, 660, 670, 42, 786, 1009, 236, 1360]. Optimizing [3429].

Optimum [1750, 161, 3211, 332]. Options [3152, 1460]. Orbit [1107]. Orbital [491]. Orbits [314, 2372, 3026, 3197, 1129, 796, 1647, 1674, 2479, 3393, 1931]. Order [1271, 2167, 3164, 1592, 3576, 844, 883, 980, 1126, 233, 2472, 1136, 3141, 939, 1457, 453, 777, 135, 1923, 2735, 1312, 2427, 748, 902, 267, 635, 415, 1820, 1121, 1128, 261, 3228, 2837, 1974, 265, 1553, 1549, 1866, 3478, 1505, 613, 340, 2123, 933, 2009, 1564, 6, 246, 805, 1104, 2673, 2520, 2488, 1127, 806, 1138, 379, 124, 2778, 335, 3517, 175, 115, 1767, 1249, 391, 1124, 1162, 1232, 1342, 598, 2118, 618, 432, 2805, 2938, 984, 848, 1360, 1439, 1882, 3049, 730, 77, 778, 470, 1688, 892, 562, 216, 983, 1190, 580, 1415, 1197, 917, 229]. Order [1449, 1677, 577, 1420, 2726, 66, 687, 1299, 1403, 3183]. Ordered [297]. Ordering [1014, 875, 2241, 419, 922, 586]. Orderings [158, 1089, 2241]. Ordinary [2775, 2774, 188, 537, 483, 1121, 723, 1717, 1484, 340, 624, 964, 1797, 379, 749, 1823, 555, 1439, 778, 1099, 470, 45, 622, 70, 2479]. Ordinate [992]. Ores [3021]. Organisms [1467, 2954]. organization [1814]. Organized [1790]. Organizing [3355, 195]. Orientation [363]. Orientational [3513]. Oriented [1067, 922, 521]. Ornstein [2809, 2821]. Orr [2975]. Orthogonal [2397, 1358, 3532, 1152, 1113, 74, 2336, 1054]. Orthogonality [522]. Orthonormal [1295, 2057]. Orthotropic [3075]. Oscillating [3559, 567, 2426, 3542, 1913, 1668, 3405]. Oscillation [980, 937, 500, 631, 777, 748, 591, 2326, 843, 806, 746, 3033, 1415, 1197, 229]. Oscillations [2356, 2966, 3241, 2178, 1035, 1525, 899, 1938, 1243, 2872, 2603, 2849, 2012, 1435, 2937, 1872, 1794, 1885, 315, 2407, 1422, 1259, 416, 2073, 2133, 3130, 178, 2968, 2591, 1404, 123, 2533, 905, 1449, 1677, 3187, 2541, 2285, 3069, 1403, 3302]. Oscillator [2339, 1761, 2761, 1350, 2684, 2718, 1158, 688, 1101, 2586, 1597, 2587, 2620, 2340, 1523]. Oscillators [2324, 2508, 2838, 2684, 2967, 3079, 3040, 3103, 3512, 3581, 2532, 2736, 2850, 2772, 1757, 1729, 1815, 756, 1088, 2294, 2305, 2058, 2315, 1008, 3316, 3291]. Oscillatory [2173, 2802, 2899, 3249, 715, 1160, 1145, 1196, 2258, 2926, 2814, 1299, 2864, 1226, 2874, 1960, 2124, 2443, 555, 3121, 3240, 2001, 3436, 1554, 3025]. Oseen [1879, 2367, 3004, 3200, 3168]. Osmosis [2048, 375]. Other [3252, 2736, 2734, 1228]. Out-of-Plane [915]. Outer [1006, 2139]. Outgassing [2847]. Output [272, 2226, 1201, 1483, 1161, 1805, 3179]. Outputs [1829]. overall [2431, 2432]. Overconvergence [2608]. Overdamped [3317]. Overflow [1161]. Overhead [2541]. Overlap [1701, 1958]. Overtaking [2869]. Owl [2986]. Oxidation [3021, 2293, 2302, 2578, 2626]. Oxidation-Reduction [3021]. Oxide [2910, 2740].

Packet [1873]. Packing [2213, 516, 2400]. Padé [1837, 3397, 1607]. Pair [836, 1268, 1300, 3058, 2478]. Pairs [1546, 1506, 485, 144, 63, 1179]. Paley [209]. Pancreatic [2965, 3459, 3408]. Paper [37, 38, 534, 1812, 144, 1574, 3183]. Papoulis [2196]. Parabolic [3164, 3293, 2471, 258, 2192, 1142, 2357, 2812, 2326, 1174, 356, 676, 1657, 3524, 544, 3054, 3562, 2996, 1575, 1774, 2128, 88, 1055, 1246]. Paraboloid [719, 896]. Paraboloidal [443, 1122]. Paradox [3360, 2885]. Paradoxical [2547, 2682]. Parallel [2880, 3425, 1859, 1930, 1258, 3221, 112, 2575, 3001, 1580, 2409, 2745, 2227, 2252, 2296, 2623]. Parallel-Plane [1258]. Parameter [3092, 141, 1092, 1103, 97, 117, 2310, 367, 10, 1781, 2046, 1963, 421, 990]. Parameter-Dependent [141]. Parameterized [2466]. Parameters [1737, 926, 188, 593, 471, 297, 2474, 2172, 473, 2734, 1659, 1445, 2879, 3279]. Parametric [1064, 226]. Parametrically [923, 3518]. Parametrization [1695]. Paraxial [2472, 2889]. Pareto [1316, 589]. Parity [859, 1497]. Park [2987]. Parsimonious [2103]. Part [2951, 2952, 2524, 2543, 1594, 1732, 3555, 2660, 3443, 1025, 1144, 1533, 1561, 2194, 2237, 2473, 2485, 852, 853, 2659, 3085, 3378]. Partial [3239, 2609, 3010, 2142, 291, 2013, 3009, 3085, 2030, 2735, 2307, 519, 874, 851, 2633, 44, 2038, 237, 1484, 413, 1910, 2733, 558, 1570, 2088, 2890, 271, 1522, 2645, 2040, 872, 216, 3316, 1292, 1675]. Partially [3270, 2825]. Participants [1798]. Particle [3139, 3078, 3454, 2927, 3461, 2785, 204, 2150, 2840, 2026, 2027, 2019, 321, 3199, 2909, 2422, 2977, 2345]. Particle-Like [2840]. Particles [3371, 2652, 2314, 2857, 1958, 134]. Particular [663]. Particulate [1196]. Partite [672]. Partition [1566, 2694, 300, 2806]. Partitioned [326, 615, 1363, 627, 938]. Partitioning [1313, 906]. Partitions [72]. Pass [1076]. Passage [2565, 1978, 3364, 3493, 1880, 2353, 1922, 688, 1101, 2653, 3268, 1756, 3025]. Passages [681]. Passive [207, 204, 885]. Passivity [782]. Past [2156, 320, 575, 1235, 773, 835, 1879, 2367, 447, 2422, 2643, 3060, 3189, 3188, 809, 719, 896, 251, 1231, 1246, 3213]. Path [1876, 1086, 2272, 816, 2314, 256]. Path-Integral [2272]. Pathological [1984]. Paths [1044, 2701, 253, 136, 1539, 1550, 55, 2656]. Pattern [1920, 444, 3368, 3311, 3066, 2528, 1184]. Patterning [3079]. Patterns [2324, 363, 3128, 2703, 2126, 798, 1599, 1590, 3488, 833, 2711, 2737, 51]. Payments [540]. PDE [3506, 939, 3101]. Peakedness [2241]. Pearson [1421]. Peclet [3424, 3060, 3071, 3188]. Peculiarities [1821]. Penalties [875]. Penalty [455].

Pencil [3240]. Pencils [731, 1787]. Pendant [2228]. Pendula [3317]. Pendulum [221, 2426, 3433, 3226]. Penetrable [1985]. Penrose [1052, 1379, 916]. Penultimate [638]. Perfect [3336, 516, 920]. Perfectly [2260, 1067, 2645]. Perfectly-Plastic [2645]. Performance [2354]. Period [2011, 2625, 2160, 3107, 3194, 2253, 1863, 1928, 2732, 3106]. Period-Adding [2732]. Periodic [3458, 2039, 969, 2797, 2567, 1890, 3372, 2773, 3485, 1686, 838, 988, 2269, 2379, 2085, 2414, 1898, 1035, 1388, 3374, 3553, 1366, 1396, 1557, 1512, 1443, 2179, 1704, 345, 3328, 3056, 3008, 2372, 763, 2967, 2099, 2189, 3073, 356, 3411, 484, 762, 2914, 367, 2425, 2892, 3310, 2486, 2272, 1435, 1329, 1591, 380, 2253, 1442, 2779, 2668, 339, 1314, 266, 357, 2996, 2863, 1738, 218, 1021, 3026, 3373, 2288, 3219, 2248, 2513, 2816, 2903, 335, 1249, 3231, 234, 2654, 1723, 2438, 2408, 1404, 3536, 432, 1129, 1931, 820, 1921, 3531, 833, 1905, 984, 1726, 2369, 450, 2829]. Periodic [3049, 3472, 796, 778, 797, 2444, 3120, 3089, 70, 826, 1598, 577, 1420, 3462, 274, 2055, 2479, 737, 2239, 3432, 3393, 3233, 3297, 3581, 2913, 1403, 3183, 1818]. Periodical [2133]. Periodically [3400, 2746, 2377, 3031, 3106]. Periodicity [26, 1043, 544]. Perishable [1352]. Perishable/Nonperishable [1352]. Permanent [3517]. Permanents [359]. Permeability [3333]. Permutation [846, 1347, 359]. Permutation-Embedding [1347]. Permutohedron [1462]. Perona [3360]. Persistence [2832, 3056]. Person [540, 1097, 734, 852, 853, 735, 1405]. Perturbation [1169, 3539, 760, 3250, 3335, 1740, 1978, 3480, 817, 2216, 399, 2328, 970, 1170, 1302, 2583, 2760, 3532, 3165, 1389, 1585, 3395, 2218, 1947, 2183, 2000, 3068, 440, 1701, 2217, 2022, 598, 182, 2965, 1181, 601, 1641, 1949, 186, 1112, 2629, 3289, 284, 274, 1407, 2794, 2180, 1229, 3544, 3545, 2285, 2286, 3069, 1746, 2146, 1897]. Perturbation-Galerkin [2760].

Perturbation/Proper [3532].

Perturbations

[1278, 3013, 1312, 700, 3067, 2067, 2939, 1049, 2172, 1526, 990, 978, 2717, 1461, 2878, 3011]. Perturbed [3164, 2376, 969, 1594, 939, 2838, 3104, 838, 988, 1898, 2081, 3555, 2245, 2174, 2914, 2762, 1484, 3143, 2563, 153, 1337, 1422, 3243, 1480, 2564, 2060, 2171, 1536, 814, 2377, 1950, 1809, 2118, 2657, 432, 3357, 2451, 708, 3436, 2898, 3012, 737, 2452, 3303, 3304]. Perturbing [3288, 433]. Petrarch [3439]. Petroleum [2606]. Petrov [435]. pharmaceuticals [2555, 2556]. Phase [2841, 2721, 3576, 2457, 27, 2607, 2700, 2368, 3387, 2359, 2508, 2838, 3493, 2495, 3086, 2924, 3100, 2854, 2876, 899, 2418, 1989, 651, 966, 2914, 2873, 3040, 2425, 2892, 3461, 3518, 2713, 3342, 2141, 2736, 786, 3566, 3259, 1094, 1095, 200, 2082, 3063, 1912, 3216, 2282, 707, 2454, 2955, 3436, 2582, 2834, 1639, 2562, 2469, 2737]. Phase-Locked [2141, 2955]. Phase-Locking [2737]. Phase-Plane [2582]. Phase-Sensitive [3259]. Phase-Shift-Averaging [899]. Phaselocking [2333]. Phases [2974, 2518]. Phenomena [1847, 2235, 1979, 1718, 2395, 1741, 2592, 2029, 2736, 2824, 1865, 2041]. Phenomenological [238]. Phenomenon [3212, 2732, 2497, 92, 2053, 2100, 868]. Photonic [3577, 3297]. Phyllotaxis [2819]. Physical [584, 2879, 1966]. Physically [210]. Physics [336, 204, 701, 2768]. Pick [1696]. Picture [1862]. Piecewise [899, 822, 2895, 1429, 928, 2430, 1298, 1618]. Piecewise-Linear [899, 822]. Piezoelectric [3528]. Pitchfork [3268]. Pitfalls [1524]. Pivot [1287]. Pivotal [922]. Placement [289]. Planar [657, 3521, 3537, 1707, 949, 1485, 287, 521, 66]. Planck [3250, 3335, 3361, 3303, 1260, 1891, 3334]. Plane [2880, 319, 2465, 1566, 1258, 2933, 37, 1538, 3013, 1437, 2056, 2170, 890, 1624, 252, 1072, 403, 2781, 192, 435, 1796, 3247, 3027, 1458, 2, 933, 1027, 3075, 2008, 2176, 1917, 3426, 238, 2438, 2634, 1057, 2977, 915, 1966, 3120, 215, 489, 284, 2582, 1408, 54, 867, 973, 3448]. Planes [903, 765, 2409, 2623]. Planktonic [1774]. Plans [1060, 1488]. Plants [2819]. Plasma [3435, 2245, 3098, 2568, 2060]. Plasmas [3515, 2150, 2026, 2027]. Plastic [3005, 3090, 3144, 1984, 1016, 785, 2645, 1018, 982]. Plasticity [192, 3486, 1023, 3465]. Plate [262, 1235, 2131, 835, 3020, 106, 2748, 2825, 511, 1395, 3251, 1093, 809, 1574, 1246, 3473, 2828, 3287]. Platelet [2935]. Plates [2843, 1085, 1351, 1308, 2741, 341, 1172, 950, 3451, 1792, 3033, 1100, 126]. Player [1417]. Plexus [2324]. Plug [2147]. Poincaré [1523]. Point [320, 3208, 3364, 240, 180, 694, 157, 2881, 3248, 2372, 484, 1302, 340, 3280, 2627, 954, 1252, 14, 2254, 108, 2513, 1036, 1550, 193, 1159, 470, 1122, 876, 717, 2058, 3096, 740, 1204, 2365, 2366, 2286]. Point-to-Periodic [2372]. Pointed [3572]. Pointless [1012]. Points [514, 1141, 804, 168, 453, 158, 2357, 751, 2619, 1044, 1089, 1484, 1097, 1797, 2816, 2903, 2104, 712, 19, 848, 2128, 2297, 1439, 235, 476, 1725, 2726]. Pointwise [2427, 1007, 177, 88, 216]. Poiseuille [435]. Poisons [2293]. Poisson [3361, 948, 3335, 2199, 2398, 384, 2362, 3334, 2370, 1602, 1983, 856, 2375]. Poisson-Distributed [856]. Poisson-Vlasov [948]. Poissonized [2806]. Pol [2011, 3485, 2216, 2761, 1435, 2779, 1957, 2098, 2139, 2582]. Polar [178, 2175]. Polarization [2651, 3566]. Polarized [2071, 89, 2279, 1897, 1896, 1964]. Polars [1198]. Poles [2097, 520]. Policies [1178, 1744, 2363, 1475, 1355, 644, 586, 90]. Policy [1195, 2065, 419, 1446]. Pollution [3087, 3059]. Polycrystal [3486]. polycrystalline [2980]. Polycrystals [2610]. Polydisperse [2844]. Polygenic [1705]. Polygonal [2973, 1550]. Polygons [565]. Polyhedra [846, 870]. Polyhedral [2158, 1276, 3573, 2786]. Polyhedron [1084, 481]. Polyhedrons [2331]. Polymatrix [953]. Polymer [2484, 2875]. Polymer-Dispersed [2875]. Polymeric [2722]. Polymerization [2002, 2084, 2747]. Polymers [2120, 2803, 3530, 2555]. Polymorphisms [1919]. Polynomial [1690, 1393, 1305, 626, 1427, 1614, 1553, 2211, 1615, 2035, 2488, 2638, 1664, 2070, 2168, 133, 541, 236, 764, 743, 609, 774]. Polynomials [26, 1076, 317, 1571, 2397, 257, 2491, 1000, 1284, 597, 2336, 115, 166, 362, 46, 331, 1655, 1212]. Polyominoes [516]. Polytope [1569, 1690]. Polytopes [472, 1496]. Polytropes [3155]. Pool [2145]. Pooled [2443]. Poorly [3286, 3445]. Popov [513]. Population [2919, 3458, 3042, 1811, 2500, 3389, 2773, 2832, 3557, 2789, 3078, 1396, 1557, 1512, 1676, 3067, 1489, 1224, 2152, 1880, 2061, 3236, 3222, 1493, 43, 2064, 2087, 2672, 1768, 3053, 1971, 1477, 1894, 1603, 2498, 2415]. Populations [3556, 2258, 2953, 1502, 1972, 2552, 2954, 927, 3488, 1815, 2901, 2014, 2573, 3223, 2988, 3043]. Porosity [2815]. Porous [412, 599, 2607, 2368, 1773, 2567, 2147, 2815, 2801, 3271, 1989, 2713, 962, 1094, 1095, 1268, 2943, 2473, 2485, 3234, 783, 3060, 3189, 3188, 2369, 2410, 1601, 2605, 2992]. Port [503, 531]. Ports [207, 885, 850]. Posed [3478, 3460, 3471, 3254]. Positive [408, 2465, 206, 937, 1780, 294, 1457, 523, 2613, 1695, 626, 722, 884, 357, 1475, 311, 640, 163, 632, 1451, 1710, 1602, 774, 684]. Positively [16]. Positivity [1632]. Possess [1632]. Possessing [826]. Possible [2160]. Possibly [929, 1751]. Post [901, 790]. Post-Buckling [790]. Postbuckling [923]. Posynomial [1166]. Potential [1169, 3239, 1927, 2072, 1722, 3019, 2331, 657, 317, 3310, 259, 1091, 255, 2272, 3150, 2976, 1968, 447, 2960, 2996, 3195, 1891, 2113, 1331, 1734, 2330, 2255, 15, 190]. Potentials [2776, 1296, 3417, 1017, 3278, 100, 2913]. Power [2011, 3453, 1690, 2426, 2761, 3229, 3194, 1025, 397, 2542, 1869, 2535, 2407, 1334, 1480, 1284, 2492, 1590, 1882, 1868, 1497, 2813, 2594]. Powers [1465, 523, 1548, 56, 68, 1303]. Practical [1830, 1963, 1662]. Prairie [2833]. Prandtl [3424, 3447]. Pre [1834]. Pre-Breakdown [1834]. Precedence [875]. Precise [1991]. Precision [2867]. Predation [2833, 3546]. Predator [2271, 3124, 2755, 1559, 2352, 1366, 1443, 2179, 3391, 2372, 1170, 1302, 1489, 2159, 2474, 3178, 1714, 1774, 3567, 1967, 2499]. Predator-Prey [2271, 3124, 2755, 1559, 2352, 1366, 1443, 3391, 2372, 1170, 1302, 2159, 2474, 3178, 3567, 2499]. Predators [2277, 1678, 2025]. Prediction [1549, 131, 2364, 1725, 251, 1482]. Predictor [1087, 3254]. predictor-corrector [3254]. Predictors [239]. Preface [3153, 994, 1262, 201]. Preference [1265]. Preferential [2837, 2441]. Prefix [1661]. Prefix-Synchronized [1661]. Premixed [2814, 2412, 2527, 1917, 2428, 2844, 2348]. Prescribed [1050, 142, 952, 425, 707, 324, 51]. Presence [1594, 3286, 1149, 1150, 1720, 605, 856, 3511]. Presentation [831]. Preserve [247]. Preserving [25, 1925]. Pressure [1131, 3272, 3452, 3034, 3492, 2145, 275, 3306, 2175, 2001, 344]. Pressure-Step [3034]. Pressures [868]. Preventing [3422]. Previously [962]. Prey [2271, 3124, 2755, 1559, 2352, 1366, 1443, 2179, 3391, 2372, 1170, 1302, 2159, 2474, 3178, 1714, 2317, 1774, 3567, 1967, 2499]. Prey-Predator [1714]. Preys [2277]. Prices [1060, 1488]. Primal [1356]. Primary [2324]. Principal [2381]. Principle [551, 3528, 2947, 882, 1570, 1015, 1540, 758, 2725, 1650, 687, 3382]. Principles [2615, 1528, 3172, 2599, 1055]. Prior [787, 2599, 14, 1405]. Priori [895, 3150, 2182, 1213, 216, 983]. Priorities [629]. Priority [2244, 677, 893, 376, 230]. Priors [3429]. Prison [3369]. Probabilistic [306, 1645, 2275, 596, 195]. Probabilistically [2837]. Probabilities [322, 1421, 3261, 3500]. Probability [637, 880, 1464, 3328, 741, 2477, 1802, 1527, 115, 675, 1476, 2280, 545, 410, 264, 710, 1161]. Probe [2265]. Probes [3323]. Probing [2595]. Problem [1163, 2811, 3430, 3539, 3489, 2572, 3325, 1779, 3250, 2607, 2505, 2494, 2804, 3449, 1441, 168, 322, 3361, 3453, 2331, 3326, 3263, 2687, 1423, 657, 1210, 547, 942, 1852, 2094, 2291, 2731, 2681, 1752, 2529, 509, 3113, 660, 674, 839, 2134, 979, 1312, 69, 874, 1103, 2613, 1682, 1398, 1562, 2223, 2308, 314, 304, 906, 2856, 747, 2086, 593, 720, 1696, 125, 666, 1676, 1007, 1121, 1834, 466, 3307, 2896, 3128, 2245, 895, 2439, 3379, 1059, 2002, 356, 761, 3107, 1508, 1168, 2895, 659, 1547, 1648, 2568, 29, 1880, 3225, 836, 3257, 1409, 1613, 1635, 2423, 2771, 3020]. Problem [2545, 2546, 1456, 3550, 1861, 1438, 2194, 2237, 447, 3280, 1104, 3054, 3413, 954, 2482, 3562, 2399, 3080, 3136, 3294, 3541, 858, 1429, 348, 454, 1138, 2636, 1066, 790, 1689, 900, 2466, 1701, 749, 1802, 3339, 1385, 3397, 2171, 133, 1536, 2022, 2247, 161, 1936, 2113, 238, 2006, 1813, 2710, 1352, 3258, 1213, 2904, 1845, 421, 2704, 2462, 3251, 3427, 1353, 1540, 182, 1129, 2758, 1537, 3415, 463, 73, 277, 3055, 3240, 418, 1934, 2672, 1058, 374, 1756, 107, 1111, 2536, 2730, 481, 859, 2282, 156, 177, 730, 796, 1734, 794, 1470, 1641, 88, 1992, 586, 2617]. Problem [344, 490, 1112, 3217, 2330, 2898, 332, 3289, 1346, 1220, 295, 3490, 407, 395, 2878, 2794, 190, 2025, 375, 1889, 1285, 557, 793, 2238, 2742, 918, 2912, 213, 1182, 2777, 1390, 1391, 1392, 2343, 2525, 3314, 2345, 2980, 1931]. Problems [3330, 3164, 1436, 3495, 3428, 2243, 2769, 462, 976, 539, 600, 854, 904, 560, 926, 3516, 1430, 3019, 3037, 130, 227, 2906, 2013, 240, 3085, 37, 1923, 3446, 3409, 1833, 494, 2537, 1092, 3255, 2483, 519, 897, 2925, 851, 1624, 2221, 946, 694, 650, 2702, 157, 1351, 3548, 894, 1217, 82, 3455, 2328, 3173, 784, 3118, 2910, 1273, 932, 192, 1115, 1339, 1479, 3000, 1301, 3034, 1484, 1991, 3478, 1458, 2, 2563, 153, 1657, 204, 1741, 1908, 278, 3460, 385, 1064, 3521, 2154, 1337, 2123, 2962, 1144, 2009, 2506, 2948, 624, 964, 713, 965, 2599, 1533, 1561, 733, 705, 313]. Problems [480, 1563, 558, 1027, 3230, 2188, 1252, 2865, 711, 1127, 3486, 1770, 1947, 1475, 2863, 74, 2000, 2254, 3068, 93, 1383, 1943, 440, 2839, 2564, 1268, 2186, 590, 3376, 144, 648, 2290, 528, 2230, 1666, 338, 1024, 698, 3471, 1009, 3036, 1809, 3278, 1662, 167, 1451, 283, 182, 1057, 1958, 2246, 968, 193, 3444, 9, 2292, 3033, 1261, 1099, 358, 3378, 707, 683, 562, 2389, 63, 717, 2655, 2451, 215, 2566, 2327, 2033, 2109, 284, 436, 231, 740, 1349, 2037, 2149, 407, 1510, 3012, 3011, 15, 190, 1639, 2047, 32, 573, 2780, 630, 2727, 2810, 323, 486, 2541, 488, 1391, 1392]. problems [2555, 2556, 1995, 2204, 2525, 1775, 1776, 3254, 2285, 2286, 3069, 2456, 2507]. Procedure [2243, 3249, 1866, 1659, 1181, 1100, 2431]. Procedures [873, 85, 1589, 101]. Process [1849, 143, 3456, 3160, 1753, 2809, 2081, 2590, 3180, 2653, 2587, 2362, 1444, 1963, 2492, 3081, 181, 1201, 2747, 1446, 1710, 1839, 2821, 3332, 853, 1482, 2794, 499]. Processes [2167, 1317, 2465, 616, 600, 854, 2596, 1849, 2461, 803, 3558, 2434, 1309, 1923, 2419, 1035, 1196, 1525, 1799, 3386, 2759, 1191, 2065, 1549, 2112, 2948, 2674, 246, 1053, 1551, 2061, 179, 2303, 2035, 2306, 1475, 928, 1335, 1691, 124, 2208, 2336, 1698, 1646, 857, 1033, 2746, 1263, 1102, 427, 464, 697, 24, 2225, 2036, 2666, 2503, 1087, 1882, 704, 197, 2868, 1321, 2238, 3511]. Processing [2404, 2539, 101, 1829]. Processor [3125, 2807, 2535, 2691, 2921, 3097, 2251, 2321, 2449, 2559, 3001, 3030]. Processor-Shared [2691, 2921, 3097]. Processor-Sharing [3125, 2807, 2251, 2321, 2449, 2559, 3030]. Processors [437, 667]. Produced [1046, 3079, 3323]. Product [2516, 961, 655, 1352, 578, 19, 493]. Products [514, 911, 202, 172, 1548, 28, 1065, 1117, 424, 359, 194, 548, 50, 556, 995, 325]. Profile [1594, 3105]. Profiles [2460, 3474, 3301]. Program [280, 12]. Programming [583, 1198, 1532, 1356, 600, 854, 1398, 946, 959, 109, 1233, 1059, 165, 1116, 1178, 114, 487, 278, 1064, 179, 1240, 342, 217, 40, 910, 283, 724, 398, 621, 214, 963, 148, 668, 358, 876, 717, 2656, 231, 428, 147, 12, 95, 4, 59, 139, 226, 771, 824]. Programming-Type [214]. Programs [1166, 1370, 1356, 1195, 16, 292, 308, 1215, 1744, 176, 1301, 502, 492, 398, 199, 1618, 428, 232, 378]. Progressing [585, 1204]. Progressive [1311]. Projection [946, 963, 518, 2055]. Projections [2198, 2692]. Projective [168, 765, 403]. Projectors [1113]. Prolate [2281, 3188, 2239]. Prolongation [520]. Proof [1642, 1340, 2619, 435, 985, 241, 481, 2675, 90]. Proofs [161]. Propagating [2841, 2349, 3452, 2703, 2189, 2845, 3142, 2267]. Propagation [2880, 3293, 1594, 3372, 3208, 2459, 2496, 2887, 2750, 1503, 3553, 2347, 2812, 3218, 1193, 1241, 3034, 2121, 2114, 341, 2413, 1952, 947, 2405, 3242, 2034, 2045, 2082, 2177, 2248, 1665, 1785, 1853, 815, 1359, 726, 756, 3362, 2041, 2624, 2829, 1238, 1835, 1907, 1516, 3309, 2834, 2799]. Propellant [2646]. Propellants [2848]. Proper [1441, 1401, 3532, 512]. Properties [1286, 1953, 1578, 412, 599, 1126, 3335, 1558, 887, 26, 3404, 285, 803, 1309, 1535, 1713, 930, 1397, 3210, 2812, 289, 1188, 3411, 145, 2935, 255, 297, 1003, 1140, 2853, 1850, 1932, 1056, 439, 1454, 112, 2793, 1993, 1253, 855, 3479, 725, 2492, 2765, 618, 3334, 2744, 187, 512, 1171, 3308, 1966, 529, 3015, 1886, 931, 1225, 355, 21, 426, 488]. Property [649, 241, 1607, 555, 194, 2370, 859, 1632, 395, 318]. Proportional [3560]. Proportional-Integral [3560]. Proportioning [1556, 1606, 1668]. Proportions [1663]. Proposal [111]. Propulsion [2016]. Prospection [2771]. Protein [2788]. Protein-Bound [2788]. Protosequence [2292]. Proximity [3083, 2015]. PS [3002]. Pseudo [961, 1124, 2561, 1697]. Pseudo-Convex [1124]. Pseudo-Inverse [961]. Pseudo-Inversion [2561]. Pseudo-Random [1697]. Pseudoinverse [371, 3478, 329, 587, 967, 71, 671]. Pseudoinverses [408, 311, 775, 1515]. Pseudolocal [3237]. Pseudoplastic [1813]. Pseudospectra [2975, 3131]. Pseudosymmetric [22, 132]. Psi [706]. Pulsatile [2945]. Pulsating [2359, 2459, 2189, 1665, 1853, 1999]. Pulsations [3107, 2248, 1917]. Pulsatory [2483]. Pulse [3117, 3315, 1930, 2384, 2595, 2307, 3235, 1982, 3242, 725, 3276, 2772, 815, 2456, 2631]. Pulse-Coupled [3276, 2772]. Pulsed [1067, 282]. Pulses [2178, 3221, 2435, 3442, 3452, 2703, 3259, 3327]. Punch [1583]. Pure [3034, 1474, 2478]. Pure-End-Condition [3034]. Purely [2800, 3510]. Pursuit [252].

Quadratic [1198, 3516, 299, 2329, 1215, 1276, 626, 1987, 986, 2208, 479, 1024, 398, 48, 3240, 2036, 1189, 668, 774, 1906, 2525]. Quadratically [2675, 2433]. Quadrics [646]. Qualification [670]. Qualitative [3335, 2814, 1348, 2141, 1924, 3334]. Quantal [1081]. Quantitative [2636, 3451]. Quantization [2166, 3312, 2861]. Quantizers [2166]. Quantum [1001, 3533, 3074, 3417, 997, 3578, 995, 1906]. Quantum-Classical [3533]. Quarantine [3369]. Quarter [2465]. Quasi [2804, 3505, 2540, 3056, 2189, 1276, 2920, 2892, 2253, 3524, 3581, 544, 2865, 1952, 1240, 342, 3381, 2248, 2513, 2816, 2903, 432, 730, 2037, 3432]. Quasi-Continuous [3381]. Quasi-Convex [1276, 1240, 342]. Quasi-Euler [2920]. Quasi-Isothermal [2037]. Quasi-Linear [2804, 544, 730, 3524]. Quasi-Periodic [3056, 2189, 2892, 2253, 2248, 2513, 2816, 2903, 432, 3432, 3581]. Quasi-Reversibility [2865]. Quasi-Static [3505, 1952]. Quasi-Uniform [2540]. Quasiisotropic [2256]. Quasilinear [1105]. Quasiperiodic [1801, 2058]. Questionnaire [1118]. Queue [2971, 2199, 2398, 1753, 891, 1382, 2399, 2836, 2921, 3097, 3396, 279, 914, 448, 11, 1201, 1483, 2821, 3002, 3137, 3030]. Queueing [527, 2991, 3497, 2837, 893, 376, 112, 2342, 2790, 3110, 581, 2745]. Queues [3029, 1715, 1910, 677, 230, 2354, 2691, 2869, 3570, 3571, 1218, 3001, 2744, 2868, 3340, 2227, 2252, 2296, 3341]. Quotients [1548, 1117].

R [38]. R. [37]. Radar [3575, 1829]. Radial [84, 227, 766, 89, 46, 828]. Radially [1712, 3420, 3017]. Radiation [2243, 1784, 1222, 3442, 2483, 3556, 1682, 2708, 2995, 1998]. Radiation-Resisted [1222]. Radiative [2347, 360, 273, 718, 463]. Radius [1169, 626, 1767]. Radon [3353, 2995, 1545, 2911, 1857, 3380]. Rain [2503]. Rain-Flow-Cycle [2503]. Rainfall [2604]. Random [907, 883, 1804, 2339, 3293, 3504, 2235, 1961, 1951, 2595, 1142, 2260, 3456, 3329, 2213, 3338, 2319, 2355, 3210, 3328, 2560, 3076, 75, 2718, 2038, 3109, 3475, 1909, 1700, 1241, 346, 1478, 1648, 35, 1842, 1074, 1056, 2353, 2309, 2266, 3395, 3280, 2303, 1334, 1720, 2341, 2502, 14, 1764, 1383, 1716, 3337, 3099, 1459, 2890, 2791, 818, 1234, 581, 814, 675, 3, 281, 1494, 2400, 2973, 1038, 3350, 1476, 726, 756, 571, 1912, 2280, 2447, 2430, 464, 697, 1958, 545, 2193, 347, 1148, 107, 829, 1303, 1749, 264, 2885, 957, 50, 556, 2655, 2015, 2717, 636, 1543, 476, 1884]. Random [1492, 2180, 1179, 1402, 316, 2238, 2780, 1697, 488, 1390, 1391, 1392, 2799, 1390, 1391]. Randomly [1594, 871, 2384, 310, 3452, 1898, 3497, 2386, 2012, 841, 2035, 1480, 3048, 1899, 1536, 2200, 1746, 1747]. Randomly-Spaced [2386]. Randomness [1749, 3223]. Range [3461, 1311, 594, 2682, 2531]. Range-Domain [594]. Ranges [1368]. Rank [1180, 883]. Ranking [322, 2169, 1798]. Ranks [1029]. Rao [473, 1983]. Rapid [2673, 1720, 2945, 3503, 60, 1937, 3069, 2285]. Rapidly [3155, 1545, 3243, 3036, 1492]. Rapidly-Fluctuating [1492]. Rate [2269, 1207, 1354, 2754, 2007, 3338, 3152, 2328, 1502, 891, 3039, 1952, 1666, 914, 1023, 3347, 1768, 3053, 1971, 2821, 1018]. Rates [3016, 1382, 3236, 3496, 2036, 1680, 2883, 32, 3179]. Ratio [384, 422, 2619, 1418, 1716, 2374, 24]. Rational [460, 2505, 1188, 1664, 955]. Ratios [187]. Ray [1258, 566, 2887, 1125, 2273, 2287, 2122, 1238, 2111, 2957, 276, 3448, 3256]. Ray-Optical [1258]. Rayleigh [2752, 439, 249, 156]. Rays [2114, 2467, 1712, 2888]. RC [807]. rd [1767]. Reactant [2396]. Reacting [2798, 3017, 2033, 2635, 2146]. Reaction [3312, 2411, 3163, 1890, 1322, 1979, 2773, 1833, 2085, 3319, 1535, 2395, 1672, 2570, 1628, 1758, 1826, 3274, 2881, 3021, 1830, 2333, 2189, 1777, 1530, 2325, 3146, 1824, 1329, 1591, 2018, 1226, 3024, 2592, 2589, 3519, 2626, 2835, 1929, 2161, 1720, 2316, 3457, 1874, 3337, 2133, 2288, 1469, 2045, 1665, 2788, 2443, 2230, 2206, 2591, 3121, 2428, 2686, 2938, 2041, 2128, 1058, 601, 2569, 2666, 3050, 3120, 1200, 1461, 1855, 2037, 2149, 3148, 2992]. Reaction-Controlled [2626]. Reaction-Diffusion [3312, 2411, 3163, 1890, 1979, 2773, 1833, 2085, 3319, 1535, 1672, 2570, 1628, 1758, 1826, 1830, 2333, 2189, 2325, 1329, 1591, 2018, 3024, 2592, 3519, 1929, 2161, 2316, 2133, 2288, 2206, 2591, 3121, 2041, 1058, 2666, 3050, 3120, 1200, 1855, 2037, 2149, 3148]. Reaction-Hyperbolic [2686]. Reactions [1911, 2148, 2274, 2444]. Reactive [2550, 2815, 2347, 2942, 1719, 1821, 1832, 1861, 1733, 1870, 3357, 2992]. Reactive-Diffuse [1719]. Reactive-Diffusive [1821, 1832, 1861]. Reactor [3504, 1119, 1103, 2613, 650, 834, 1247, 1102, 1367, 277, 1449, 1677, 2959]. Reactors [1145, 1652, 1867, 2298, 2460, 2554, 2530, 2938]. Real [206, 1690, 2701, 1970, 354, 2402, 3291]. Real-Power [1690]. Real-Variable [1970]. Reality [3383]. Realizability [1765, 1693, 850]. Realizable [1364, 1045, 1194]. Realization [183, 10, 496]. Realizations [798, 1487, 1364, 155, 833, 404]. Realizing [1330]. Receptor [1491]. Reciprocal [2167, 3320, 916]. Reciprocals [1303]. Recognition [444, 3415]. Recognizing [3197]. Recombination [1256]. Reconstruction [2198, 2329, 2030, 1836, 2153, 2896, 3105, 2283, 3150, 1545, 2077, 3554, 2089, 3112, 3207, 1862, 2879, 3279, 2093]. Recourse [232, 378]. Recovery [3013, 2484, 3281, 2615, 2902, 2065, 3481]. Recruitment [3222]. Rectangular [1297, 1413, 1401, 1727, 1728, 772, 2486, 1480, 288, 1792, 2857, 3111, 2828]. Rectification [3087]. Rectilinear [1508, 29, 1320]. Recurrence [1616, 1397, 892, 1632, 1884]. Recurrences [392]. Recurrent [2414, 1883]. Recursion [385, 2694, 96]. Recursions [1553]. Recursive [2534, 18, 1837, 588]. Recycling [3390, 3585]. Redox [2349]. Reduced [255, 3498, 1575, 2638]. Reduces [3480]. Reducibility [1171]. Reducible [2465, 2257]. Reducing [1166, 2997, 176]. Reduction [452, 280, 2774, 1106, 2081, 3021, 2870, 1004, 271, 3138, 507, 995]. Reductions [3134]. Redundant [634]. Reed [1767]. Reentry [2080, 1098]. Reference [221, 743]. Refined [3500]. Refinement [3185]. Reflected [2866, 3277, 1549, 1948, 3481]. Reflecting [3386]. Reflection [1566, 1258, 3151, 2384, 3113, 1584, 1645, 2962, 1334, 1480, 2341, 2502, 2661, 1281, 820, 3093, 951, 2617, 276, 687]. Reflective [2812]. Reflector [3151]. Reflectors [1587]. Refracting [2341, 2502]. Refraction [1566, 3536]. Refractive [3280, 1712]. Refractory [3194]. Regime [1235, 3526, 3525, 2348]. Region [1423, 2308, 3391, 367, 2154, 1292, 2343]. Regions [1996, 1071, 3230, 1039, 3147, 1270, 400, 2437]. Regression [929, 1848, 2277, 709, 1751, 588, 652, 2106]. Regular [672, 538, 2183, 2973]. Regular-Polygonal [2973]. Regularity [412, 599, 615, 3443, 1646, 464]. Regularization [2923, 3460, 3471, 3254]. Regularizations [2641]. Regularized [2936, 482]. Regularly [446]. Regulated [2744]. Reinforced [3337, 3350, 3296]. Reiterated [3565]. Related [2198, 291, 3159, 2291, 2395, 1076, 930, 784, 1144, 2488, 2669, 298, 2290, 1774, 2006, 432, 1965, 1497, 1568, 2284]. Relation [1722, 1210, 3492, 3039, 1155, 916, 648, 695, 1321]. Relations [1616, 1265, 2859, 1940, 2131, 272, 1113, 93, 2476, 3320, 23, 522, 122, 892, 1632]. Relationship [1424, 593, 2953, 835, 1319, 3043]. Relationships [1670, 2169, 906, 3195, 279, 249, 1543]. Relative [1914, 3040, 2138, 3081, 2998]. Relatively [2682]. Relaxation [2356, 2966, 2178, 2803, 1435, 3035, 3149, 2315]. Relaxations [1370]. Relay [445]. Release [3389, 3438, 1952, 2555, 2556]. Relevance [1182]. Relevant [375]. Reliability [1953, 3100, 94, 568, 2309, 2266, 634, 1617, 1468].

Reliable [1471]. Remainder [638]. Remark [37, 38, 2480, 2232, 2766, 712]. Remarks [1269, 202, 1214, 1781, 534, 2650, 1699, 2025, 426]. Removal [3496, 1048, 32]. Rendez [3118]. Rendez-Vous [3118]. Renewal [600, 854, 292, 114, 2112, 1551, 3570, 3571, 936, 1245]. Renormalization [3566]. Renormalized [3048]. Repairable [634]. Repairman [3000, 3080]. Repairmen [2399]. Repeated [2219]. Repetitive [2116]. Replacement [539, 927]. Replacement-Stocking [539]. report [2337]. reporting [2337]. Representation [1457, 1063, 3235, 899, 595, 152, 1032, 2136, 965, 1772, 1045, 1194, 597, 2477, 298, 1033, 2497, 1023, 1524, 791, 1433, 2579, 1204, 316, 3082]. Representations [637, 2871, 1001, 202, 1003, 3570, 1004, 3471, 938, 3340, 759, 995, 1670, 3341]. Representatives [1107]. Representing [2919, 1605]. Repression [2515]. Reproduction [2901]. Reservoir [3218, 2606]. Reservoirs [799, 3272, 3086]. Residuals [2230]. Residue [26, 2199, 986, 1189]. Residues [2402]. Resistance [285, 3385]. Resisted [1222]. Resistive [2769, 1473]. Resistivity [2910]. Resistor [2831]. Resistors [2729]. Resolution [3128, 7, 2995]. Resolvents [206]. Resonance [2756, 2565, 2095, 2802, 2899, 3249, 3493, 1761, 2308, 3323, 1158, 2939, 3034, 2074, 2944, 2096, 2080, 688, 1101, 1098, 1770, 1947, 1689, 3130, 391, 2620, 338, 1950, 2389, 2058, 2366, 1941, 2343]. Resonances [2761, 2285]. Resonant [1764, 3510, 2816, 2903, 3130, 2584, 3187, 2978, 3170, 2642]. Resonantly [2069]. Resource [2179, 3108, 1746, 1747]. Resources [2066, 3500]. Respect [723]. Response [2494, 2731, 2681, 1708, 1072, 183, 1081, 3077, 3247, 162, 1960, 1764, 3496, 312, 361, 2251, 2449, 2559, 2883, 1735]. Response-Time [2251, 2449, 2559]. Responses [3013, 1742, 2658]. Restricted [1737, 314, 304, 346, 587, 182, 1129, 796, 1931]. Restriction [1551, 995]. Restrictions [1414, 2450, 1111]. Result [3464]. Resultant [840]. Resultants [257, 866, 1183]. Results [431, 1833, 1326, 1642, 1945, 2002, 2660, 1596, 1700, 128, 457, 472, 1363, 1348, 2743, 606, 1376, 3397, 219, 1164, 1211, 1361, 1568, 2284, 3462, 103, 2745, 1747, 1964]. Retarded [1190]. Retrieval [478, 333, 2282]. Retrograde [304, 796]. Revealed [634]. Reversal [2873, 1146, 3147]. Reverse [1136, 2704, 2667, 375]. Reversed [1166, 823, 888]. Reversibility [2785, 2865]. Review [2363]. Revised [253, 136, 668]. Revisited [3401, 2348]. Revolution [1927, 2072, 2156, 2907, 1588, 1796, 191, 1879, 1968, 719, 896, 868, 1080]. Reynolds [2524, 2543, 1269, 680, 1642, 2510, 2367, 3213, 1186, 1371, 1625, 2677, 3004, 3200]. Rhombical [33]. Rib [2741]. Rib-Stiffened [2741]. Riccati [1393, 1250, 357, 736, 450]. Riccati-Type [450]. Rich [2347, 3021]. Richardson [399]. Ricocheting [42]. Riemann [2811, 2607, 3446, 3409, 2525, 2545, 2546, 77, 2617, 2033]. Riff [508]. Right [637, 635]. Right-Cauchy [637]. Rigid [2662, 1072, 2960, 1986, 572, 1807, 2645, 1291, 3473]. Rigid-Smooth [1291]. Rigidities [982]. Rigidity [944]. Ring [26, 3235]. Rings [3387, 1935, 901, 2908, 2747, 1242]. Rings-Allowed [2747]. Risk [1386, 3223]. RNA [2292]. Robin [979, 1861]. Robustness [1117, 2165]. Rock [3218, 2053, 2100]. Rod [1970, 815]. Rods [2937, 1712, 915, 949, 126, 3487]. Role [1496, 80]. Roll [1538, 2080, 1098]. Rolling [2825]. Rook [1431]. Room [845]. Roosbroeck [2278]. Root [1341, 326, 2131]. Root-Finding [2131]. Root-Location [326]. Roots [386, 743]. Ropes [2455]. Rotary [3542]. Rotating [2916, 3155, 1782, 1672, 3445, 2882, 3038, 1970, 785, 2908, 2512, 1371, 3135, 286, 1238, 1877, 489, 1998]. Rotation [2339, 2637, 2990, 2490, 605, 1020]. Rotational [2056, 1698]. Rotationally [2039, 739]. Rotatory [2937]. Rough [3464, 1645, 3375, 3301]. Rounded [3572]. Rounding [1994]. Route [481, 358]. Routes [2700, 2816, 2903]. Routing [2398]. Rubin [1983]. Rubinow [2828]. Rugged [2820]. Ruin [3136, 3294, 2758]. Rule [306, 143, 550, 2236, 3142, 596]. Rules [1430, 1914]. Rumor [2388]. Runaway [1652, 2590, 2554, 2845, 2249, 2148, 3285]. Runge [2808]. Runs [1551]. Rupture [275].

S [37, 38]. Saddle [180, 2719, 470, 876, 717]. Saddle-Point [470, 876, 717]. Saffman [3214]. Saint [3528]. Salesman [1456, 1328]. Same [1184]. Sample [1803, 2210, 353, 3098, 768, 725, 1646, 1162, 1232, 1881, 464]. Sampled [871, 2035]. Sampling [3016, 2329, 305, 3423, 3152, 1702, 2195, 2077, 64, 1498, 1962, 927, 1234, 2994, 2598, 2448, 2727, 2810, 3380, 2196]. Sand [831]. Sans [2024]. Satellite [131, 1674]. satisficing [1405]. Satisfy [171]. Satisfying [1051]. Saturated [1094, 1095, 3060, 3189, 3188]. Saturation [3492, 2557]. Scalar [3139, 2926, 1588, 2853, 3096, 2688, 3468, 54, 3233]. Scale [3282, 3086, 2170, 2925, 345, 1790, 3532, 297, 2241, 3385, 2995, 2075, 3405]. Scale-Invariant [2075]. Scales [2013, 1493, 1279, 1552, 1921, 1164, 2888, 3545, 3292]. Scaling [1534, 1221, 3354, 1344]. Scanning [7]. Scattered [2859]. Scatterers [1592]. Scattering [3239, 3428, 3539, 3256, 3561, 2866, 1876, 2505, 1519, 1660, 343, 2192, 1961, 3434, 2197, 566, 2261, 2151, 3464, 3399, 1125, 2097, 2651, 2308, 3014, 1820, 1985, 2470, 2639, 2752, 2982, 3076, 2875, 2858, 1588, 1796, 255, 3257, 3238, 1389, 2616, 3543, 3498, 2117, 3186, 3376, 3195, 2748, 1844, 2462, 3410, 794, 2240, 3508, 1346, 2107, 2677, 3448, 531, 3473, 3300, 2343]. Schauder [484]. Scheduling [1382, 530, 1009, 2215]. Scheme [480, 1570, 2443, 338]. Schemes [1271, 280, 3146, 3575, 3380]. Schrödinger [2695, 3239, 3204, 3032, 2144, 1470]. Schur [1061, 1052, 1216, 1695, 2780]. Schwartz [209, 533]. Schwarz [1007, 944, 565]. Schwarzschild [390]. Sciences [1265, 1266, 1264, 1263, 1262]. Scientific [3425, 1836]. Scoring [1236]. Scott [2882]. Screen [319, 890, 54]. Scroggs [371]. Scroggs-Odell [371]. Search [1177, 787, 909, 1167, 1115, 1576, 841, 781, 816, 2970, 1899, 2066, 682, 1445, 792, 678, 856, 1130, 1541, 2068]. Searching [1447]. Seasonal [1904]. Second [2167, 1592, 3576, 980, 1126, 939, 777, 1312, 2427, 748, 902, 635, 415, 1820, 1128, 356, 261, 3228, 265, 1553, 1549, 1866, 3478, 613, 340, 933, 2009, 2079, 6, 246, 1104, 124, 335, 175, 115, 1249, 1124, 1162, 1232, 2118, 618, 2422, 2977, 1439, 1882, 778, 470, 1688, 562, 1415, 917, 229, 577, 1420, 2726, 1028, 325, 1299, 1403]. Second-Kind [2977]. Second-Order [2167, 3576, 3228, 3478, 124, 115, 2118, 1882]. Secondary [2378, 1085, 1856, 2184, 2622, 3132, 2065, 2253, 1760, 1469, 2248, 1792, 1213, 1531]. Section [1413, 1521, 393, 1346, 3487]. Sections [2470, 2639, 623, 3007]. Sectorial [3230]. Sedimentary [3494]. Sedimentation [3248, 3348, 2785, 3048]. Segel [3480]. Segmentation [3244, 3182]. Segregation [3461]. Seismic [3151, 2122]. Seismology [3113]. Selection [1836, 3253, 1919, 3066, 914, 1671]. Selective [3351, 2337]. Self [3535, 3180, 645, 576, 1957, 2264, 1434, 217, 40, 55, 195, 2988, 1137, 323]. Self-Adjoint [323]. Self-Adjointness [576]. Self-Avoiding [55]. Self-Dual [1434]. Self-Duality [217, 40]. Self-Equilibrated [645]. Self-Excitations [1957].

Self-Ignition [3535]. Self-Organizing [195]. Self-similar [3180]. Self-Similarity [1137]. Self-Sustained [2988]. Semi [2925, 143, 835, 2858, 2510, 645, 2839, 181, 704, 809, 324, 639]. Semi-Cherenkov [2839]. Semi-Cylinder [2858]. Semi-Infinite [2925, 835, 2510, 645, 809, 324]. Semi-Markov [143, 181, 704, 639]. Semiconductor [2572, 3372, 2984, 2927, 2910, 2964, 3105, 3074, 3417, 2278, 2818, 2908, 2217, 2171, 2704, 2768, 3578, 2667, 2628, 2629, 2740, 2787]. Semiconductors [2817, 3127, 3220, 2985, 3514]. Semidefinite [311]. Semidynamical [514]. Semigeostrophic [3449]. Semilinear [2928, 1951, 1833, 839, 1174, 2507]. Semimonotonically [153]. Semiordered [981]. Semipassivity [782]. Semisimple [1003]. Sense [247, 2195, 1133, 2448, 1321, 2196]. Sensing [2220]. Sensitive [2460, 2554, 2530, 3259, 1018]. Sensitivities [3500]. Sensitivity [3546, 205, 1023, 3161]. Sensory [2685]. Separability [705]. Separable [714, 1618]. Separated [1208]. Separation [2130, 1365, 987, 1036, 1187]. Separator [1707]. Separatrix [2776]. Sequence [2574, 2529, 3107, 1822, 2733, 2089, 496, 1330, 1915, 681, 155, 510, 1189, 977]. Sequences [673, 763, 1347, 83, 1487, 628, 1364, 1045, 1194, 281, 767, 863, 1173, 1485, 2191, 1179, 2239]. Sequencing [437, 667, 875]. Sequency [770]. Sequential [1398, 2298, 482, 787, 272, 1631, 927, 2045, 2274, 2236, 388, 2938]. Series [2011, 1310, 2219, 2329, 2767, 2426, 2761, 2298, 116, 3229, 1421, 2603, 2542, 152, 2535, 1850, 1932, 2407, 2214, 1533, 1561, 3230, 752, 268, 2477, 47, 2791, 1234, 166, 362, 522, 550, 2938, 656, 251, 2957, 1763, 1229, 3303, 2431, 2432]. Server [2971, 376, 2354, 1218, 3137]. Servers [2199, 2398]. Service [1644, 2971, 1753, 1715, 2837, 3000, 891, 3396, 914, 3137, 3030]. Service-Adaptive [3000]. Service-Rate [914]. Set [1687, 2871, 1195, 847, 1605, 1276, 1427, 3460, 1139, 819, 679, 740, 95, 1808]. Set-Up [679]. Sets [441, 560, 765, 135, 2434, 1918, 3118, 2692, 1301, 611, 3261, 1550, 991, 1819]. Settler [3101]. Several [58, 3013, 2277, 2474]. Sexually [3252]. Seymour [1926, 1812]. Shabat [2505, 1389]. Shading [3575]. Shadow [3526]. Shadowing [2564]. Shah [3182, 3267]. Shallow [1131, 3480, 2812, 3038, 2175, 1122, 173]. Shanks [1340]. Shannon [2329]. Shape [3425, 1351, 3307, 2010, 2870, 3027, 3575, 2422, 2977, 2992]. Shape-from-Shading [3575]. Shaped [1878, 1583, 428]. Shapes [2044, 1990, 3406]. Shared [2691, 2921, 3097, 3500]. Sharing [3125, 2807, 2398, 2179, 2251, 2321, 2449, 2559, 3001, 629, 3030]. Sharp [2803]. Shaw [2312, 3384]. Shear [2662, 3559, 2581, 3286, 3345, 3509, 3072, 3476, 3075, 1209, 2824, 2314, 321, 2697, 3033, 1261]. Shear-Flow [3072]. Sheared [2010]. Sheet [89]. Sheets [2932, 3219, 2764, 2642]. Shell [1782, 275, 1122]. Shells [921, 790, 561, 2509, 31, 15, 190]. Shielded [3323]. Shift [2508, 2838, 899, 3436]. Shifts [1056, 3068, 3069]. Shock [1255, 1269, 1222, 3157, 1858, 2542, 2800, 2249, 3156, 1870, 2118, 3285, 3383, 2877, 2146]. Shock-Induced [2249]. Shock-Wave [3383]. Shocks [2751, 2299, 2311, 2813]. Shooting [3425, 978]. Shop [685, 530, 2215]. Short [1149, 1150, 3038, 312, 1281, 1681, 2682, 1154]. Short-Range [2682]. Shortage [1802]. Shortcomings [3087]. Shorted [702, 1176]. Shortening [2569]. Shortest [1086, 642, 253, 136, 481, 358]. Shot [2709, 1646]. Should [2987]. Shtokalo [1375]. Shuffling [508]. Shutdown [581].

Shuttle [1644]. Si [2901]. SIAM

[1789, 854, 2244, 37, 988, 1721, 2731, 1546, 1512, 2639, 2252, 3043, 667, 1479, 38, 2539, 1885, 253, 1932, 534, 2309, 1585, 1101, 2405, 2044, 144, 2903, 189, 1232, 1494, 2450, 3183, 1788, 1156, 792, 896, 1361, 1574, 1677, 139, 2100, 2562, 138, 184]. sic [1729]. Side [540]. Side-Payments [540]. Sided [2050, 3566]. Sidepayments [1248]. Sides [901]. Sieve [2902]. Signal [2615, 2902, 422, 14, 2374, 187, 2579, 101, 86, 1829]. Signal-To-Noise [187, 422, 2374]. Signaling [2922]. Signals [3016, 2329, 1424, 1193, 786, 3381, 2364, 159]. Signed [1692]. Signs [1490]. Silent [1387]. Silicon [2487, 2627, 2578, 2626]. Similar [1275, 1958, 3180]. Similarity [2581, 2190, 2423, 1870, 3070, 3501, 1137, 1639]. Simon [466]. Simple [1002, 3558, 2426, 1959, 1340, 2849, 2619, 2125, 2530, 1155, 3081, 1096, 2443, 1009, 3552, 3465, 610, 2675, 3421]. Simplex [804, 712, 668]. Simplexes [617]. Simplices [1626]. Simplification [340, 561, 198]. Simplified [2942, 3072]. Simulated [3138]. Simulating [2759]. Simulation [3335, 2418, 3133, 2960, 2531, 2606]. Simulations [2946, 3199]. Simultaneous [452, 1106, 99, 246, 1487, 2292]. Sine [3104, 1333, 2914, 1563]. Sine-Gordon [3104, 1333, 2914, 1563]. Single [3086, 3210, 1982, 875, 1467, 2482, 2354, 1218, 3108, 1477]. Single-Machine [875]. Single-Nutrient [1467]. Singular [1297, 929, 3539, 2356, 2966, 514, 760, 3250, 2494, 663, 1978, 3533, 227, 1411, 1312, 700, 1834, 3067, 2328, 461, 2184, 2067, 1831, 3532, 2563, 1450, 1049, 1880, 1379, 3573, 3165, 2716, 2467, 2779, 2232, 2172, 1947, 576, 2183, 2000, 2285, 2286, 3068, 3069, 288, 440, 268, 1384, 1701, 2217, 1526, 2022, 1751, 1813, 2187, 1809, 182, 1414, 2450, 2993, 3177, 1641, 978, 1949, 1112, 2566, 2629, 1440, 2037, 2149, 2548, 2640, 775, 2878, 2365, 2366, 2794]. Singularities [1, 1869, 1850, 1932, 2463, 189, 1070, 3384, 155, 3289, 325]. Singularity [3262, 3584, 243, 2545, 2546, 987, 1583, 3204, 2518, 1395, 1521, 533, 3184, 1574, 2525]. Singularly [3164, 2376, 2452, 969, 3104, 2245, 2762, 1484, 3143, 2563, 1337, 3243, 2564, 2171, 1809, 2118, 2657, 2451, 2898, 3012, 3303, 3304]. Singularly-Perturbed [1809]. SIS [3563, 3351]. Sitar [2050]. Sites [3389, 1897, 1896, 1964]. Sivashinsky [2601, 3298, 2719]. Size [3430, 2515, 2611, 3367, 2275, 3320, 3058, 2745]. Size-Structured [2611]. sized [2432]. Skan [823, 454, 2643, 1289]. Skin [2711]. Skinning [3530]. Slab [1428, 125, 834, 1453, 1587, 164, 506]. Slacked [165]. Slender [2416, 2524, 2543, 2192, 575, 2660, 3102, 1091, 1588, 1796, 2583, 191, 1879, 1521, 3189, 2764, 2659]. Slicings [832]. Slider [2134]. Slip [1235]. Slip-Dominated [1235]. Slipless [137]. Slips [3518]. Slit [1651]. Slits [574]. Slope [3038, 3502]. Sloping [2439, 2608, 1547, 2300, 2001]. Sloshing [2584]. Slow [2565, 3208, 3364, 3493, 2326, 3494, 773, 2074, 3025, 2679, 1922, 3268, 3459, 1921, 2977, 1756, 2569, 3015, 3304, 1735]. Slowly [2802, 2899, 2508, 1584, 1525, 3311, 2310, 3040, 1741, 2299, 2287, 688, 1101, 3095, 1992, 1516, 2385]. Slowly-Varying [2310, 1516]. Slowness [3151]. Small [1169, 500, 3446, 3409, 1092, 817, 1503, 403, 2418, 367, 1382, 3238, 3395, 3045, 2165, 3068, 1539, 1731, 1800, 2857, 421, 2490, 2822, 2340, 3070, 1688, 990, 3181, 3161, 2238, 3069, 1403]. Small-Disturbance [3070]. Small-Noise [3181]. Smooth [3186, 1586, 1291]. Smoother [3096]. Smoothing [1192, 483, 166, 362, 3313, 3096, 609, 1482]. Smoulder [2880]. Snakes [1465]. Snyder [3383]. Sobolev [1862, 728]. Social [1265, 1266, 1542, 2954, 1264, 1263, 1262, 1236]. Soft [2804, 3399, 1820, 2190]. Soil [2604, 2284]. Soils [3400, 2622]. Sojourn [2596, 2869, 3030]. Solid [2291, 582, 3450, 2218, 2518, 2396, 1733, 3186, 2646, 2019, 1665, 30, 3111]. Solid-Liquid [2518]. Solidification [3467, 2979, 2945, 2699, 2676, 2571]. Solids [2841, 3476, 2816, 2903, 3046]. Solitary [1724, 3579, 2817, 3127, 2219, 1979, 3442, 1777, 3116, 3143, 2299, 1855, 1228]. Solitary-Wave [3579]. Soliton [2632, 1936, 2826, 2020, 1855, 3421]. Soliton-Like [2020, 1855]. Solitons [3475, 2392]. Solute [1773, 2157, 2567, 3467, 2244]. Solution [3330, 1996, 1277, 168, 729, 1283, 3019, 1047, 227, 37, 3485, 135, 657, 1210, 1649, 1721, 392, 3399, 1103, 2170, 2925, 946, 2223, 3454, 2426, 3274, 2201, 593, 864, 3307, 772, 1059, 2660, 3367, 2381, 2481, 2858, 237, 2895, 1991, 3527, 2, 2542, 827, 1880, 385, 1064, 866, 836, 3573, 2632, 2983, 2190, 2716, 2123, 1325, 337, 339, 1453, 1454, 1863, 313, 1027, 1570, 3230, 3413, 752, 3562, 858, 889, 769, 93, 2270, 2475, 1260, 288, 2186, 300, 3373, 47, 2230, 814, 2826, 238, 2006, 552, 563, 578, 698, 3306, 522, 1845, 2704, 1327, 2862, 1709, 73, 3147]. Solution [1414, 2450, 2672, 1058, 1254, 3216, 2292, 1641, 1625, 1122, 2389, 489, 719, 896, 284, 853, 3490, 95, 3193, 1246, 1829, 973, 573, 1080, 2238, 213, 1554, 3299, 1995, 2204, 1897]. Solutions [1163, 2811, 3579, 3428, 2039, 969, 1779, 1683, 2411, 2804, 2928, 904, 3163, 1969, 3516, 939, 2775, 631, 903, 2843, 715, 1160, 958, 1686, 575, 838, 988, 2379, 3094, 3499, 2085, 2414, 1535, 1119, 2391, 1092, 129, 519, 2427, 2613, 1399, 851, 1142, 2767, 650, 1035, 1196, 1672, 2258, 1388, 304, 2570, 1628, 1758, 1826, 2782, 1366, 1396, 1557, 1512, 2761, 430, 203, 720, 1704, 3067, 3056, 1830, 461, 2967, 1980, 1174, 3118, 356, 1981, 1777, 762, 1043, 1427, 367, 1115, 2159, 2600, 2762, 259, 1091, 255, 2510, 3034, 3116, 1224, 1329, 1591, 1113, 380, 2253, 35, 113, 823]. Solutions [1226, 3521, 1016, 1968, 2154, 1442, 2112, 3165, 1337, 2779, 625, 2423, 624, 964, 1719, 389, 558, 1314, 1426, 954, 1929, 2161, 3136, 3294, 1172, 357, 443, 1138, 2051, 2636, 524, 1186, 1066, 1455, 1622, 1567, 218, 379, 1691, 1279, 3026, 1870, 1384, 1851, 2649, 1376, 2840, 3514, 2512, 335, 2344, 1249, 2443, 975, 1359, 2753, 598, 2187, 3471, 3536, 2393, 3575, 2645, 1451, 1102, 154, 783, 3035, 1057, 48, 122, 2584, 2643, 1726, 1256, 1439, 2370, 3070, 859, 450, 812, 3472, 475, 778, 797, 1925, 2020, 819, 2410, 88, 216, 983, 1601, 3383, 554, 186, 957]. Solutions [2877, 31, 3436, 580, 1725, 2720, 1200, 821, 1292, 1598, 2109, 222, 852, 853, 577, 1420, 3421, 3468, 3501, 2787, 3283, 91, 1639, 737, 2658, 867, 973, 21, 126, 557, 2631, 1289, 918, 687, 1390, 1391, 1405, 3025, 2525, 2090, 1523, 2358, 1818, 1392]. Solvable [2774, 1989, 1681]. Solving [462, 1621, 2308, 382, 2927, 1217, 192, 265, 1908, 713, 480, 3397, 283, 2036, 2327, 2343]. Some [1286, 714, 935, 1636, 808, 1001, 976, 211, 600, 854, 527, 540, 1269, 26, 202, 307, 546, 1833, 2537, 1713, 1399, 2221, 1642, 82, 1188, 366, 568, 3118, 2002, 970, 3584, 1528, 1339, 1479, 2762, 1700, 128, 176, 1214, 457, 676, 1287, 472, 2031, 1994, 627, 1629, 1765, 2506, 2827, 1117, 480, 1585, 1993, 1127, 3486, 2394, 1486, 597, 2476, 1376, 312, 640, 3397, 163, 279, 219, 1324, 1410, 1412, 370, 338, 1009, 1662, 1451, 1540, 1537, 1699, 122, 968, 938, 1414, 2450, 512, 1164, 331, 683, 256, 491, 636, 2257, 1211, 1361, 1197, 1886, 991, 931, 1225, 1568]. Some [2284, 740, 1349, 1407, 2025, 21, 103, 196, 426, 104, 1670]. Sommerfeld [3278, 2975]. Sonar [1829]. Sonic [766, 2119, 3564, 3070]. Sorbing [962, 1094, 1095]. Sorption [2517]. Sorting [370]. Sound [1594, 78, 3346, 3112, 2705]. Source [320, 766, 2094, 2483, 2153, 3248, 1504, 230, 3162, 2399, 238, 321, 2730, 3279, 2813]. Sources [2283, 2792, 3028, 1374, 2597, 2137]. Souriau [1188, 1380]. Souriau-Frame [1188, 1380]. Space [118, 3151, 2687, 3282, 1119, 945, 1624, 2924, 2223, 158, 3555, 1233, 2164, 1089, 1685, 309, 389, 2143, 3222, 2636, 648, 298, 1862, 2059, 1902, 1881, 148, 358, 1598, 147, 72, 1377, 2057]. Space-Homogeneous [2164]. Space-Limited [3222]. Space-Time [118, 1119, 1881]. Spaced [3016, 3388, 2386, 1874]. Spaces [220, 3265, 703, 2169, 99, 44, 455, 534, 655, 19, 721, 727, 493, 876, 491]. Spacings [353]. Span [2427]. Spanier [144]. Spanning [1330]. Sparse [991]. Sparsity [922]. Sparsity-Oriented [922]. Spatial [2773, 3499, 3124, 3227, 3368, 1599, 1635, 3204, 2133, 1256, 3487]. Spatially [1911, 2235, 1790, 2084, 2914, 3535, 1450, 2552, 3366, 2148, 2317, 2073, 2654, 1244, 3472, 2720]. Spatio [2937]. Spatio-Temporal [2937]. Spatiotemporal [3518]. Special [2534, 1265, 1266, 1264, 1263, 1262, 565]. Species [2986, 3392, 1730, 2277, 2611, 1623, 1720, 2317, 1264, 2723, 451, 2334, 1477, 1492, 1931, 1739]. Specified [1010, 1703]. Spectra [2419, 1869, 3327, 2879, 1882, 3233, 3297]. Spectral [1317, 3016, 2181, 2537, 2078, 3411, 2338, 1478, 1132, 3477, 1253, 3479, 2203, 872, 1135, 355, 795, 837, 898, 3205, 2913]. Spectroscopy [996]. Spectrum [3559, 3507, 384, 3194, 1234, 277]. Speech [3415]. Speed [1594, 2648, 1503, 3192, 2835, 3048, 3112]. Speeding [2656]. Speeds [974, 793]. Sphere [1169, 1149, 626, 773, 3527, 2367, 3346, 541, 605, 3060, 2677, 774, 2594, 3004, 3168]. Spheres [998, 3399, 2960, 3048, 1300, 2432]. Spherical [146, 2494, 3326, 1782, 2262, 2318, 2674, 1821, 1000, 275, 2624, 100, 2813, 1292, 1349, 868, 2375, 1377]. Spherically [3112]. Spheroid [3188]. Spheroidal [1072, 2281, 2239]. Spike [1748]. Spikes [2851]. Spinning [2848, 2359, 2459, 2042, 1907]. Spins [2683].

Spiral [1722, 1672, 1830, 1854, 2018, 2371, 2950, 1272, 3063, 3121]. Spline [2077]. Splines [2329]. Split [2780]. Splitting [2099, 2616, 3279]. Splittings [1401]. spontaneous [2146]. Spray [3198, 2844]. Spray-Formed [3198]. Spraying [3336]. Spread [2085, 3369, 2083, 3059]. Spreading [2388]. Spring [2603, 2849, 2879]. Spring-Mass [2603, 2849]. Springs [2547]. Spurious [3035]. Square [1341, 2809, 811, 439, 744, 2005, 1083, 619, 101, 3046]. Squares [657, 649, 2491, 1570, 845, 1384, 98, 1037, 588, 800, 1154, 86, 1886]. St. [190]. stabilities [2289]. Stability [2811, 2210, 2349, 1126, 431, 2848, 3122, 3241, 1441, 2501, 2500, 2523, 1595, 3127, 3515, 1784, 3104, 3272, 2147, 3558, 2551, 971, 1833, 1535, 2550, 2815, 2391, 897, 3507, 1782, 1898, 673, 1826, 2917, 363, 1333, 3128, 3067, 3298, 2453, 1981, 1489, 2346, 435, 2892, 1700, 128, 1294, 3116, 1071, 1275, 153, 2152, 2474, 633, 2974, 3178, 2918, 2278, 6, 1053, 544, 789, 3395, 1929, 2532, 1375, 768, 3259, 1622, 1147, 2696, 2846, 1714, 2317, 2362, 2890, 2664, 1357, 2288, 2943, 561, 779, 2176, 570, 1785, 2749, 1162, 1232, 2438, 2485, 167, 2514, 1610, 154, 2274, 2428, 1921]. Stability [2526, 120, 1332, 915, 949, 3390, 3585, 156, 764, 3418, 990, 3318, 2901, 1818, 1892, 1868, 3089, 3217, 491, 743, 400, 513, 1632, 2720, 917, 2409, 2623, 1572, 2437, 2415, 364, 2562, 2469, 196, 1638, 2631, 1805, 2605, 2348, 2992]. Stabilization [2427, 2750, 2313, 3242, 3479]. Stabilized [1917, 520]. Stable [1687, 3094, 657, 1923, 650, 1706, 2233, 3056, 2967, 2012, 1919, 1908, 1957, 3135, 2225, 764, 1647, 433, 1461]. Stacking [2709]. Stage [2919, 673, 1475, 3222]. Stage-Structured [2919]. Stagnation [1708]. Standard [2864]. Standing [2378, 3518, 2133]. Stars [2574, 76]. Start [1094, 1095]. State [2919, 881, 3412, 2572, 3335, 3127, 1210, 143, 666, 3348, 1178, 1744, 891, 2275, 179, 3039, 2342, 2354, 1738, 2587, 3373, 1590, 2734, 3334, 2514, 154, 1934, 2678, 358, 1967, 2628, 2934, 2787, 1783, 1390, 1391, 1392, 3025, 2358]. State-Dependence [3039]. State-Dependent [2919, 2342, 2354, 2734]. States [1778, 2698, 1791, 2271, 2335, 1145, 2127, 2233, 2317, 2943, 1731, 1792, 2509, 681, 2206, 2445, 2738, 1461, 2905, 1606, 1668, 3298, 2358]. Static [3505, 3354, 2506, 2766, 1952]. Stationarity [247]. Stationarizing [1056]. Stationary [456, 3515, 2234, 3116, 2535, 2278, 1053, 2161, 1015, 1260, 200, 2060, 298, 2746, 541, 1120, 833, 1087, 1882, 2821, 2868, 586, 679, 710, 1321, 251, 499]. Statistic [1432]. Statistical [1520, 1773, 444, 289, 1956, 3292, 2492, 1602, 420, 1975, 488]. Statistics [883, 2384, 2712, 2520, 2488, 2502, 124, 1537, 636]. Status [2547]. STD [3563, 3351, 3422]. Steady [2572, 3335, 320, 1890, 3127, 766, 1210, 1437, 2127, 2801, 666, 3348, 2099, 1178, 1744, 2510, 891, 2275, 1738, 2317, 2008, 3373, 2943, 1731, 2019, 2206, 2358, 3334, 2514, 154, 1934, 2445, 1966, 1967, 1394, 2628, 2934, 2787, 3058, 1668, 1390, 1391, 1392, 3025, 2090]. Steady-State [3335, 3348, 891, 2275, 1738, 3373, 3334, 2628, 2934, 2787, 2358]. Steel [3098]. Steep [3185]. Stefan [2223, 1438, 707, 683, 1639]. Stein [1151]. Steiner [1562, 505, 2619, 1509, 1508, 254, 29, 1320, 2295, 2690]. Stekloff [699, 1408]. stem [2146]. Step [3367, 3034, 1733, 181, 1157, 1400]. Stepwise [124]. Stereological [1528, 1887]. Stereology [2420, 2421]. Stewart [1612]. Stieltjes [3018]. Stiff [1621, 3035, 1554]. Stiffened [2741]. Stiffness [2379]. Stikker [3193]. Stirred [2298, 2938]. Stochastic [1317, 536, 2210, 2095, 600, 854, 1780, 729, 871, 1595, 1961, 130, 1424, 2134, 1682, 2260, 3269, 593, 2418, 1233, 925, 25, 1059, 3109, 3367, 1576, 237, 2464, 1609, 3260, 1132, 2853, 3366, 246, 733, 3280, 2532, 1127, 2580, 1915, 1335, 2088, 2446, 2362, 419, 2990, 1444, 3339, 1527, 1033, 396, 1590, 2614, 1263, 908, 1600, 756, 2490, 492, 1610, 2215, 802, 1445, 1087, 2305, 151, 428, 407, 232, 378, 795, 1425, 1748, 2163, 1460, 4, 59, 139, 1889, 2238, 2385]. Stochastically [1950]. Stocking [539]. Stoichiometric [3045, 174]. Stoichiometry [1835]. Stokes [2156, 3484, 547, 3212, 2823, 1635, 1533, 1561, 3230, 1186, 1582, 2661, 154, 2422, 888, 1882, 3572, 3549, 759, 1581]. Stopping [2236]. Storage [1223]. Strain [2841, 1207, 1354, 3142]. Strain-Gradient [2841]. Strained [517, 302, 3274, 1939, 2644]. Straining [1354]. Strains [3005, 3563, 698, 1020]. Strange [1823]. Strategic [1667, 2017]. Strategies [677, 2552, 3422]. Strategy [3503, 852, 853]. Stratified [2616, 2108, 1514, 3362, 3089, 1242]. Stratonovich [2808]. Streamline [1184]. Streamlines [2978]. Streams [1476, 3332]. Streetwalker [685]. Strength [1300]. Stress [921, 945, 2751, 2856, 2010, 563, 2645, 1203, 286, 3465, 284, 3549]. Stress-Assisted [2856]. Stress-Driven [2751]. Stress-Field [2645]. Stress-Free [2010, 3549]. Stressed [1019]. Stretching [2660, 1011, 1013, 2643, 3305, 2659, 1080]. Striations [2060]. Strict [261, 2713]. Strictly [1593]. String [1034, 2050, 509, 383, 339, 829, 2635, 793]. Strings [1778, 2664, 2015]. Strip [3371, 3427, 1203, 1977, 344, 1408, 867, 973, 1267]. Stripping [3568, 3264]. Strips [574, 126, 2913]. Strong [575, 1222, 878, 2542, 1386, 1270, 197, 1135, 2058, 2037, 1055, 3011, 3171]. Strong-Interaction [575]. Strongly [2508, 2838, 2864, 2716, 2945, 2667, 2315, 2149]. Structural [903, 3100, 3532, 3158, 1024, 1568, 2284]. Structural/Mechanical [3532]. Structure [2695, 2756, 2698, 2349, 2796, 227, 1935, 1481, 3124, 1706, 3483, 508, 1214, 1471, 56, 68, 2800, 1017, 2632, 2367, 722, 2674, 2835, 3204, 1021, 2846, 3565, 169, 1684, 2826, 238, 1386, 2274, 1726, 3578, 2024, 2606, 2498, 274, 2055, 477, 1900, 3233, 3297, 824]. Structured [2919, 2551, 3557, 946, 2611, 1475, 3488, 3053, 1971, 199, 2415]. Structures [3321, 1265, 2773, 2750, 3374, 1266, 1717, 1956, 2309, 2266, 3045, 981, 2133, 2205, 2842, 2973, 1026, 3536, 1955, 3290, 1018, 1517, 2224, 3148]. Studies [3534, 659, 2380]. Study [2811, 3335, 2804, 3453, 3446, 1836, 2199, 2484, 3229, 1831, 1219, 2190, 2141, 2719, 3447, 3491, 3334, 2276, 2624, 2579, 1491, 3184, 3187]. Studying [1243]. Sturm [27, 1613, 348, 454, 1875, 749, 2727, 2810, 2777]. Sub [2175]. Sub-Buckling [2175]. Subadditive [419]. Subcharacteristic [1146]. Subclasses [142]. Subcodes [1189]. Subcommunities [2361]. Subcritical [2747, 2323]. Subdefinite [479]. Subfunctions [529]. Subgraphs [495, 691]. Subgroups [995]. Subharmonic [3493, 2391, 108, 3106]. subharmonics [3302]. Subject [363, 1502, 1986]. Subjected [89, 361, 126]. Sublinear [777]. Submanifold [304]. Submatrices [241]. Submerged [1149, 1150, 1378, 2748]. Subsamples [267]. Subsequences [2015]. Subset [1663]. Subsets [721, 727]. Subsonic [567, 3564, 2976]. Subspace [1737]. Subspaces [220, 534]. Substance [2562, 2469, 3511]. Substance-Locking [3511]. Substitute [1723]. Substitution [776, 318]. Substrate [1280]. substructure [1900, 1901]. Subsurface [3087]. Subsystem [3095]. Subsystems [3459]. Subtrees [553]. Success [1551]. Successive [157, 14, 1120, 1392]. Sufficiency [161]. Sufficient [3558, 549, 261, 670, 1505, 768, 1537, 1693]. Suggested [2360, 101]. Suitable [444]. Sum [397, 675, 296, 2280, 377, 852, 853]. Summation [597]. Summing [3012]. Sums [589, 2219, 523, 1560, 1497].

Superconducting [3387, 3569, 3202, 3466, 2905]. Superconductors [3203, 3201, 3407, 3008, 3534, 3190, 3145]. Supercritical [3116]. Superheating [3201]. Superlattices [3372]. Superlinear [902, 1531]. Supersaturated [1196]. Supersonic [3343, 1514, 2642, 3344]. Supersonic-Transonic [1514]. Supervision [444]. Supply [1460]. Support [305, 2426, 1827]. Supported [2357]. Supporting [1935]. Supports [238, 2240]. Suppression [3442, 3433]. Sure [871, 1424, 2718, 2012, 768, 2362, 1610]. Surface [2880, 1046, 2004, 979, 1645, 582, 1920, 1149, 1150, 1378, 3555, 2892, 1651, 3414, 585, 2293, 2312, 2318, 2708, 2585, 2076, 2766, 1011, 1733, 1893, 3158, 3373, 2757, 1090, 1817, 2001, 2444, 1238, 3285, 3184, 1394, 2764, 1944, 3057, 3305, 3549, 757, 1182, 3301]. Surfaces [3092, 1316, 3464, 1690, 345, 2010, 275, 3313, 77, 20]. Surfactant [3572]. Surges [2577, 2929]. Survival [2552]. Suspension [3419, 2661, 2715]. Suspensions [2115, 2347, 3048, 2432]. Sustainable [607, 3167]. Sustained [2802, 2245, 2074, 2096, 2080, 1689, 2988]. Swelling [2556]. Swelling-controlled [2556]. Swirling [2784, 1793]. Switch [1634, 2301]. Switch-Type [1634]. Switchback [2183]. Switching [2807, 2921, 3097, 2088, 1239, 3002]. Symmetric [2039, 731, 3262, 3221, 1108, 2127, 317, 2126, 255, 538, 397, 1549, 1591, 1505, 502, 1448, 1879, 2208, 640, 553, 3112, 23, 2842, 621, 3240, 409, 2036, 764, 3017, 1153, 2905, 3058]. Symmetrical [1203, 100]. Symmetrically [1203]. Symmetries [2737]. Symmetrization [1199]. Symmetrized [1181]. Symmetry [2916, 3037, 3326, 3134, 2937, 273, 3365, 2714, 2620, 2973, 2263, 2989].

Symmetry-Breaking [2937]. Synaptic [3389]. Synchronization [3276, 2772, 3316]. Synchronized [1661]. Synchronous [2199, 2398, 2112]. Synchrony [3149]. Synthesis [2505, 415, 36, 2553, 272, 67, 3381, 142, 328, 2737, 350, 51]. Synthetic [3575]. System [2811, 1953, 2756, 3125, 2607, 2807, 3361, 3085, 2735, 2085, 3319, 2258, 3157, 2755, 2352, 3391, 2326, 2333, 2189, 2603, 2849, 2346, 948, 3228, 3369, 2012, 1717, 1344, 204, 2474, 113, 1223, 376, 2480, 3024, 625, 2278, 1821, 1832, 1861, 3519, 753, 3395, 3477, 1929, 2161, 2342, 2792, 3028, 1218, 2088, 3337, 749, 2492, 1385, 581, 1245, 1774, 2683, 2206, 2251, 2321, 2449, 2559, 2597, 2187, 2915, 386, 2671, 3327, 1247, 498, 3121, 2940, 3095, 1580, 2768, 2879, 3055, 2041, 177, 195, 679, 2083, 186, 745, 1855, 3468, 2499, 918, 1556, 2745, 2057, 3205]. Systems [2210, 2637, 3010, 514, 2039, 2997, 969, 3335, 343, 1634, 3163, 2519, 871, 527, 1890, 1595, 478, 2802, 2899, 3249, 1047, 1979, 998, 240, 838, 988, 847, 3499, 1411, 188, 2435, 2947, 1535, 2427, 2613, 3507, 3388, 1898, 2570, 1628, 1940, 1758, 2110, 1366, 1557, 1443, 899, 430, 930, 1695, 591, 1031, 2969, 61, 2991, 1630, 1830, 2418, 3311, 3192, 3119, 3497, 461, 2684, 3051, 1593, 762, 1667, 1959, 247, 2159, 3532, 2939, 3461, 2464, 1044, 1854, 776, 340, 1448, 2138, 2112, 3165, 3178, 2592, 2733, 2944, 3215, 1719, 1922, 2724, 2413, 3243, 1104, 2489, 711, 2316, 655]. Systems [2165, 1875, 3196, 2743, 2088, 2362, 2256, 2990, 1986, 981, 312, 2133, 2288, 1731, 1800, 1536, 634, 2497, 2230, 271, 2968, 234, 1723, 424, 2634, 2591, 2118, 3197, 1600, 2490, 3334, 1610, 1304, 432, 1921, 2040, 2822, 3240, 1244, 333, 984, 3216, 2390, 3472, 470, 797, 1164, 2263, 3120, 45, 404, 644, 1868, 2091, 3437, 433, 708, 826, 513, 1632, 436, 1703, 2989, 2224, 1492, 737, 1638, 1805, 51, 782, 2745, 1554, 1783, 3393, 1814, 2511, 2380, 1818]. Szasz [438].

Table [1546, 1506]. Tables [1837]. Tail [3023]. Tails [2299, 3436]. Takens [2940]. Tandem [2869, 3570, 3571]. Tangential [2016]. Tank [2298, 3038, 2938]. Tapering [2621]. Target [1177, 841, 1447, 1445]. Targets [856, 1130, 1541]. Tasks [995]. Taxis [3350]. Taylor [2244, 1932, 3214, 1310, 2157, 1850, 711, 3356]. Technique [462, 904, 608, 1305, 2131, 165, 482, 2583, 2760, 1071, 1615, 769, 2390, 451, 978, 1082]. Techniques [180, 894, 2270, 2965, 199, 2212, 1018, 2806]. Technology [456]. telegraph [2507]. Telegraphic [2637, 2990]. Temperature [971, 2755, 2929, 2460, 152, 1863, 1733, 1367, 1937]. Temperature-Dependent [2755]. Temperatures [2004, 100, 757]. Tempered [211]. Temporal [2519, 2937]. Temporally [3362]. Tension [2076, 3184, 2764]. Tensor [172, 1274, 701]. Tensors [225, 1342]. Term [2094, 1435, 2098, 2139]. Terminal [1644, 532, 562]. Termination [112]. Terms [408, 2329, 2926, 208, 3162, 2183, 3169, 3216, 917]. Ternary [372]. Territorial [2986]. Tertiary [2484]. Test [2434, 1939]. Testing [1589, 1765, 604, 2005]. Tests [3429, 1490]. th [1923, 1767, 984]. Their [3325, 3018, 3037, 291, 1833, 303, 3328, 83, 3116, 1972, 1850, 2214, 2401, 3546, 2791, 219, 2746, 3500, 1062, 3418, 2738, 488, 52, 1932, 163, 1499, 3302]. Theorem [929, 1534, 58, 1935, 305, 1079, 3113, 2199, 993, 748, 902, 317, 356, 1078, 484, 913, 654, 2046, 209, 789, 1498, 1962, 1707, 1374, 2054, 98, 1751, 739, 3258, 1343, 2448, 989, 1156, 45, 877, 2361, 886, 1211, 231, 1518, 1028, 102, 2810, 477]. Theorems [980, 211, 540, 81, 326, 615, 3484, 546, 1105, 972, 784, 1458, 132, 2190, 1765, 1250, 1585, 1218, 1357, 632, 818, 1324, 370, 1499, 697, 2598, 3362, 1099, 1674, 652, 1407, 364, 65, 501, 2727, 104]. Theoretic [847, 69, 1464, 1564, 2599, 3266, 2433]. Theoretical [2796, 241, 2766, 1662, 2624, 1491, 2538]. Theories [208, 2027, 3031, 2678, 2014]. Theorists [995]. Theory [686, 319, 515, 1001, 206, 755, 937, 2095, 2519, 1773, 2157, 566, 998, 1077, 1423, 2115, 1103, 890, 2097, 3124, 3123, 3538, 1108, 650, 1196, 3014, 1643, 1351, 57, 1985, 2470, 2639, 1315, 2357, 1333, 959, 399, 109, 308, 1118, 2958, 822, 2583, 128, 255, 1431, 2065, 204, 1903, 1049, 2123, 1467, 3402, 3020, 3542, 1787, 80, 2724, 2371, 447, 3447, 2306, 3580, 947, 2405, 1015, 1172, 348, 1539, 3547, 3469, 3337, 3242, 76, 962, 2594, 900, 3167, 2466, 416, 632, 2082, 718, 570, 1786, 1828, 1410, 293, 999, 739, 1813, 578, 974, 1404, 301, 1600, 498, 1537]. Theory [387, 960, 2259, 992, 758, 1511, 1439, 3245, 533, 92, 1109, 393, 1238, 2410, 468, 2877, 1112, 1208, 2566, 1361, 2220, 1771, 1975, 1285, 1463, 276, 1182, 2023, 2541, 2244, 2337, 3287, 2090, 2146, 1403, 3246, 3302, 2507]. thereof [1901]. Thermal [1911, 2928, 3263, 945, 1867, 3286, 3345, 2590, 3107, 3417, 2845, 2249, 2145, 1863, 2518, 2148, 2778, 2512, 2943, 2646, 3147, 3285, 1735, 1937]. Thermo [162]. Thermodynamics [1028]. Thermoelastic [2752, 666]. Thermoelasticity [2470, 2639, 3444]. Thermoelectric [3307]. Thermoplastic [3509]. Thermostat [2492]. Thermostat-Controlled [2492]. Thermostatics [2221]. Thermosyphon [3431, 3129]. Thick [125, 3527, 275, 178]. Thickness [921, 3387, 3466]. Thin [3284, 1927, 2072, 2156, 2907, 3141, 2933, 2843, 3466, 259, 1968, 2154, 2123, 3402, 3403, 2707, 1172, 3526, 790, 3219, 1844, 3251, 1514, 2428, 3398, 30, 3033, 2764, 2905, 1754, 3300, 2828, 3525]. Thin-Filament [3141]. Third [233, 2123, 379, 892, 983, 580, 1449, 1677]. Third-Order [2123, 1677]. Thouless [3566]. Three [515, 2173, 2368, 1997, 2378, 3326, 766, 3282, 3202, 3392, 314, 304, 2273, 2406, 882, 2713, 3343, 1144, 954, 2150, 1194, 769, 2108, 981, 1264, 2826, 2968, 182, 1129, 1111, 2665, 796, 393, 2582, 2957, 1137, 2780, 2375, 3393, 2345, 1739, 2204]. Three-Body [314, 304, 1129, 1111]. Three-Dimensional [515, 2173, 1997, 2378, 3326, 766, 3282, 2273, 2406, 882, 3343, 2150, 2108, 2968, 2665, 393, 2582, 2957, 1137, 2780, 2375, 3393, 2204]. Three-Part [1144]. three-particle [2345]. Three-Phase [2368, 2713]. Three-Point [954]. Three-Soliton [2826]. Threshold [2360, 3364, 1979, 2757, 3468]. Thresholds [2551]. Thrusters [3435]. Time [2919, 1192, 3125, 2501, 2500, 118, 1978, 2776, 2013, 3159, 2435, 1119, 3179, 2702, 1388, 2308, 1753, 1628, 1557, 1443, 2179, 3274, 3338, 116, 3271, 930, 2969, 207, 1350, 798, 3076, 1860, 2386, 2759, 2346, 3228, 3310, 2956, 3034, 2785, 162, 1591, 3257, 2674, 1438, 2143, 3477, 2869, 3136, 3294, 2616, 2863, 746, 2051, 1493, 1567, 1597, 3547, 1279, 1552, 3099, 3514, 361, 279, 1234, 2857, 3231, 166, 362, 2251, 2321, 2449, 2559, 2734, 1723, 205, 1201, 1921, 1881, 1580, 569, 1810, 3169, 1164, 2444, 819, 2995, 513, 2738, 251, 1055, 364, 1976, 1906, 629, 1638, 941, 3030]. Time [1805, 1763, 782, 850, 2745, 3300, 2343, 3275, 3154, 3205]. Time- [2734]. Time-Asymptotic [1567]. Time-Dependent [3159, 1443, 3274, 1350, 3310, 3034, 2674, 3136, 3294, 1597, 2857, 569, 3169, 2995, 1906, 3179, 3275]. Time-Harmonic [2308, 3257, 3477, 3300, 2343, 3205]. Time-Invariant [2776, 819, 850]. Time-Limited [941]. Time-Periodic [1591, 2863, 3231, 2444]. Time-Related [930]. Time-Varying [207, 3547, 205, 513, 1805, 782, 850]. Timed [3497]. Times [2596, 2807, 2971, 2652, 2353, 2921, 3097, 3396, 2165, 312, 3181, 3002, 3137]. Timing [3556, 406]. Timoshenko [1788, 1755]. Tip [3401, 3207]. Tippe [3084]. Tissue [2804, 1280, 2190, 2665]. Titchmarsh [913].

TM [1258]. Toda [2923, 3490]. Toeplitz [1695, 934, 884, 1772, 2638, 1412]. Tomography [3256, 3265, 3151, 2030, 2086, 3128, 2941, 2897, 2770, 3237, 3353, 3582, 3562, 3324, 2070, 2137, 2994, 2930, 3175, 2931, 2957, 3352, 2168]. top [3084]. Topics [86]. Topological [285, 3227, 99, 455, 665, 3290]. Topologies [1478, 336, 1749]. Tori [1738, 178, 2479]. Toroidal [2509]. Torques [3317]. Torsion [73, 982]. Torsional [3141, 1938, 944, 1540, 982]. Torso [563]. Tortuosity [3583]. Torus [2783, 2202, 1518]. Total [526, 459, 1165]. Totally [2341, 2502]. Touchings [346, 347]. Tournament [1798]. Tournaments [172, 931, 1225]. Tours [1328]. Towed [120]. Trace [3526, 2240]. Traceable [310]. Tracer [2665]. Traces [2982]. Tracing [2273, 2122, 1712]. Tracking [3426]. Tracks [1447, 2909]. Traction [2982, 1057, 215]. Traffic [2837, 3522, 3523, 2792, 2869, 3396, 3028, 3137]. Trailing [511]. Trains [3117, 2373, 2914, 725]. Trait [3253]. Traits [1705]. Trajectory [111]. Transcendental [958, 381, 386]. Transfer [3424, 2347, 3307, 2892, 67, 2999, 360, 273, 718, 463, 1868, 3332]. Transform [460, 1935, 305, 97, 1546, 1506, 1106, 568, 2009, 3353, 1389, 662, 1993, 2270, 1268, 900, 955, 2911, 664, 387, 2282, 2995, 1857, 2957, 449, 3380, 2504]. Transform-arising [387]. Transformation [225, 904, 1849, 1151, 442, 53, 110, 60, 17, 184, 150, 212]. Transformations [2633, 25, 1340, 913, 1152, 1110, 849, 330, 1368, 997, 1888, 1806, 1171, 879, 2106, 8, 104, 138]. Transforms [84, 351, 1664, 2598, 263, 636]. Transient [2899, 3249, 1758, 3307, 2958, 711, 1567, 361, 2300, 2417, 3065, 2868, 2629, 3332, 3003, 2279, 1937]. transients [1735]. Transistor [2910, 2586, 2740]. Transistors [2878]. Transit [314]. Transition [1761, 2942, 1741, 2845, 3342, 3566, 2205, 2118, 1912, 2036, 1586, 2033, 2323, 3432, 3304, 3468]. Transitions [3387, 2127, 2126, 2406, 2736, 2185, 1810]. Translation [2871, 830]. Translation-Invariant [2871]. Transmission [1930, 160, 3475, 2918, 2482, 3028, 951, 1134, 918, 2279, 2541, 3314]. Transmitted [3252]. Transonic [1255, 3446, 3409, 1307, 1643, 2313, 716, 2976, 2394, 1514, 3070, 3211, 3501]. Transport [2997, 1876, 1428, 3449, 1773, 2157, 3434, 3404, 3505, 1501, 2906, 2613, 3456, 382, 3271, 2927, 3210, 1280, 2874, 273, 1453, 1454, 1539, 1094, 1095, 1268, 144, 1120, 1353, 1451, 1587, 3234, 1181, 1709, 1244, 2533, 1839, 506, 63, 395, 1711, 3511, 542, 2244, 3299]. Transportation [972, 1009]. Transversal [161]. Transverse [2603, 2849, 1022, 2800, 2467, 1184, 30, 3033]. Transversely [3139, 2937, 3568, 1902]. Trapezoidal [173]. Trapped [2409, 2623]. Trapping [1930, 3467, 2939]. Traps [1148]. Traveling [3064, 3163, 3372, 3023, 1979, 3499, 1758, 3067, 2372, 2373, 1777, 1824, 2864, 2458, 1456, 1929, 1999, 2040, 1328, 2686, 3216, 828, 3418, 2981, 3437, 2898]. Travelling [2178, 2773, 2483, 2159, 3519, 2288, 3517, 2715, 2473, 2485, 3052]. Treated [1356]. Treatment [847, 1895, 2272, 3026]. Tree [1562, 1508, 813, 1045, 1194, 553]. Treelike [875]. Trees [2574, 52, 505, 1509, 254, 1406, 1202, 2228, 781, 816, 1750, 1320, 1743, 1330, 1324, 1173, 445, 2690]. Trench [1581]. Trenches [1582]. Trends [204]. Tri [1671]. Tri-Allelic [1671]. Trial [94]. Trials [2236]. Triangles [2819]. Triangular [647, 525, 1513]. Tridiagonal [142, 3240, 350]. Tridiagonalization [388]. Trigonometric [3397]. Trigonometrical [522]. Triorthogonal [372]. Triple [2127, 2102, 263]. Triply [66]. Tristable [3050]. Truncated [1529]. Truncation [64]. Trusses [1016]. Tube [2510, 3140, 2784, 1601]. Tubelike [1080]. Tubes [1847, 2173, 2958, 901, 3130]. Tubular [1145, 2460, 2554, 2530]. Tucker [771]. Tulin [3185]. Tunnel [567]. Turbulence [1956, 1230, 1011, 2203, 2246, 3273]. Turbulent [1255, 1795, 1997, 1269, 3484, 1077, 3166]. Turing [3123]. Turning [1484, 340, 3280, 2254, 1439, 2365, 2366, 2726, 2286]. Turnpike [959]. Two [1436, 1592, 2637, 2457, 3125, 2607, 1644, 2700, 462, 2360, 3018, 1795, 1430, 567, 3484, 240, 1984, 2420, 3499, 3446, 3409, 3399, 1103, 3424, 694, 3100, 157, 2223, 3577, 3368, 2179, 2876, 3483, 2201, 2630, 2439, 1715, 2227, 2252, 2296, 1989, 2873, 3040, 1170, 1302, 1489, 2346, 437, 667, 2425, 835, 2762, 1091, 2602, 1956, 132, 893, 28, 1382, 2480, 2632, 2154, 1842, 1954, 2900, 2506, 3066, 2304, 246, 2124, 1418, 2218, 2399, 3570, 3571, 1720, 1364, 3222, 1694, 2270, 2653, 2475, 440, 2990, 3376, 144, 749, 1239, 2082, 1786, 634, 530, 3001, 2203, 96, 1352, 3031, 2490]. Two [1203, 3135, 2274, 2428, 571, 24, 1367, 3088, 193, 1810, 2672, 223, 2536, 3216, 2981, 3572, 1967, 3015, 506, 63, 2015, 2315, 489, 3363, 1220, 3332, 852, 853, 1349, 2409, 2891, 2726, 1460, 1402, 65, 316, 2742, 1556, 2279, 2745, 3380, 3291, 2358, 2623, 3297, 1995]. Two-Body [1220]. Two-Class [1382]. Two-Component [3018]. Two-Dimensional [1436, 1592, 2637, 462, 1795, 567, 3484, 3499, 3446, 3409, 3424, 3577, 3368, 3483, 2439, 1170, 1302, 835, 2762, 1091, 2602, 1956, 2154, 1842, 2506, 2304, 2218, 2270, 2653, 2475, 2990, 3376, 749, 530, 2203, 2490, 3135, 3088, 2672, 2536, 3572, 506, 3363, 2891, 316, 2742, 2279, 3297, 1995]. Two-Front [2981]. Two-Graphs [1239]. Two-Level [1556]. Two-Parameter [1103]. Two-Person [852, 853]. Two-Phase [2457, 2607, 3100, 1989, 2873, 2425, 2082, 3216]. Two-Point [240, 694, 157, 193].

Two-Predator [2179, 1967]. Two-Product [1352]. Two-Soliton [2632]. Two-Stage [3222]. Two-Temperature [1367]. Two-Terminal [1644]. Two-Time [1810]. Two-Unit [634]. Two-Variable [440]. Two-Way [223]. Type [2811, 1317, 3412, 2368, 1634, 3262, 3023, 3203, 3201, 3407, 874, 94, 1895, 2958, 1128, 1421, 1614, 1431, 3440, 3215, 733, 1957, 1567, 749, 473, 1412, 814, 908, 214, 2598, 3390, 3585, 450, 470, 3120, 2305, 1725, 2727, 213, 1554, 3008]. Type-II [3008]. Types [893, 1694, 2240].

Uhlenbeck [2809, 2821]. Ultimate [1070]. Unbiased [18, 297, 249, 1670]. Unbounded [2429, 560, 2258, 727]. Unbuffered [3500]. Uncertain [1148]. Uncertainty [2615, 2725, 12, 95]. Unconfined [1965]. Unconstrained [165, 482, 621]. Underdamped [3250]. Underdeveloped [1556]. Understanding [3208]. Underwater [1594, 2066]. Undetermined [61]. Unequal [3576, 2383]. Unfinished [3340, 3341]. Unfolding [2124, 3560]. Uniaxial [3133, 2463, 127]. Unidirectional [2979]. Unidirectionally [119]. Unified [2135, 1037, 2799]. Uniform [319, 515, 2156, 1413, 244, 1457, 1791, 2271, 2335, 2971, 2427, 890, 2540, 1704, 259, 1091, 3535, 1968, 2154, 1027, 1970, 3471, 605, 1439, 1238, 1965, 2389, 3137, 1771, 2365, 630]. Uniformity [1939]. Uniformly [869, 1240, 675, 2410, 433, 1204]. Unilateral [1219, 1986]. Unilaterally [2635]. Unimodular [1416, 1654]. Unimodularity [459]. Uninterrupted [792, 678]. Unique [2599]. Uniquely [2692]. Uniqueness [2783, 81, 2331, 3151, 3113, 674, 1119, 3464, 3455, 1945, 3257, 933, 624, 964, 3550, 954, 3541, 1138, 814, 1184, 2187, 974, 3410, 3362, 3308, 812, 529, 2931, 2330, 1568, 2284, 2549, 1639, 501, 3314]. Unisignants [303]. Unit [861, 2211, 1496, 857, 634, 541, 2400, 1157, 619, 1400]. Unit-Step [1157, 1400].

Unitarity [1346]. Unitary [1152, 1110, 879]. Unknown [3430, 3270, 188, 942, 2094, 116, 1945, 2137, 2730]. Unmyelinated [1930]. Unperturbed [1302]. Unrenormalizable [208]. Unseparated [1208]. Unstable [2601, 3315, 3559, 2796, 2581, 3162, 711, 2676]. Unsteady [2049, 2946, 3446, 3409, 1280, 3199, 1187, 393, 1965, 3285]. Unstirred [2084, 3024, 3531]. Unstressed [262]. Unusual [1001, 3173, 106]. Upper [844, 804, 1119, 2170, 1007, 925, 472, 1455, 3048, 199, 2511]. Upstream [2423]. Uptake [2250, 2350, 2788]. Uranium [3021]. Urine [3209]. Usage [3125, 2807, 2921, 3097, 2449, 2559, 3003]. Use [2519, 880, 257, 1903, 14, 1414, 2450]. Used [98]. Useful [608]. Using [3312, 1061, 267, 3423, 3554, 1631, 3562, 1375, 3498, 1664, 3259, 2270, 1552, 3481, 1004, 362, 522, 3491, 2402, 792, 678]. Utilities [1374]. Utility [456, 1386].

V [3068]. Vacancy [3081]. Vaccination [3488]. Vacuo [162]. Vacuo-Thermo-Junction [162]. Vacuum [766]. Valid [1430]. Validation [2886]. Validity [200, 1524]. Value [1436, 907, 2494, 462, 976, 2523, 3361, 3037, 240, 1423, 2291, 494, 839, 1092, 1103, 2925, 851, 2221, 694, 650, 157, 1351, 2319, 116, 593, 125, 666, 1676, 1834, 356, 2381, 1168, 1484, 1547, 2563, 1379, 1337, 2123, 1144, 2771, 624, 964, 713, 965, 705, 558, 954, 1252, 1127, 2636, 74, 2000, 3068, 1383, 2564, 3546, 590, 749, 473, 648, 2171, 528, 2230, 1813, 3258, 1213, 3036, 1845, 421, 1809, 1662, 2246, 968, 728, 1058, 1099, 562, 88, 344, 1112, 1618, 436, 407, 3187, 2238, 918, 788, 323, 486, 1182, 2541, 2285, 2286, 3069, 2507]. Valued [478, 1696, 2840, 2747, 2279]. Values [1297, 257, 723, 1168, 1502, 1609, 541, 354, 512, 1440, 2323, 1402]. Vapor [3414]. Varga [37]. Variable [907, 921, 2754, 2007, 635, 864, 3367, 434, 2152, 781, 1631, 2835, 1970, 2522, 440, 3108, 2995]. Variable-Length [781, 1631]. Variable-Yield [3108]. Variables [2223, 255, 1365, 1716, 312, 818, 675, 271, 96, 2280, 1881, 545, 1303, 3015, 50, 556, 2402, 636, 1543, 2498, 369, 1179, 1402]. Variance [1887, 3138, 1881, 1958, 507]. Variances [925, 2005]. Variate [741]. Variates [589, 1548, 2355, 1614]. Variation [665, 2670, 877]. Variational [3244, 3321, 919, 2947, 3182, 1895, 772, 1484, 3048, 2344, 3415, 9, 693, 788, 3382]. Variations [1636, 1503, 1362]. varieties [1901]. Various [1903, 1750]. Varying [2802, 2899, 2508, 1584, 3466, 207, 3311, 2310, 1502, 1741, 2299, 891, 688, 1101, 3547, 830, 2765, 205, 3095, 446, 2821, 1992, 1516, 513, 2200, 1805, 782, 850]. Vector [2166, 1737, 1441, 2596, 2197, 1312, 703, 772, 1064, 1685, 2874, 1851, 2118, 2193, 1882]. Vector-Processes [1882]. Vectors [1609, 1560, 986, 512, 1882, 710, 3545]. Vehicles [2468]. Vehicular [3522, 3523]. Velocimeter [2220]. Velocity [1131, 3272, 1640, 1708, 864, 1866, 1991, 3039, 1054, 2730, 2676, 2834, 2180]. Venant [3528, 190]. Version [2728, 1154]. Vertex [1521, 1573, 445]. Vertical [1859, 221, 383, 2748, 2825, 2177, 1877, 3057, 2959]. Vertices [1462, 2228, 805, 1237, 1328, 445]. Very [3453, 2802, 2899, 575, 170, 2510, 312, 2449, 2559]. Vesicle [2044, 1990]. Vessel [2852]. VI [3069]. Via [27, 1198, 1393, 304, 1827, 2386, 3460, 3374, 3338, 3067, 2159, 2385, 3254, 3279]. Vibrating [1779, 2050, 2750, 339, 950, 2664, 3055]. Vibration [221, 754, 2828]. Vibrational [384]. Vibrations [1970, 30]. Vicinity [1709]. View [2102]. Virtual [1025]. Viscoelastic [3412, 3141, 2887, 2803, 3157, 3102, 2696, 361, 815]. Viscoelasticity [895, 955, 1213, 758]. Viscometer [2958]. Viscoplastic [2581, 3476]. Viscosity [2841, 2815, 3118, 2475, 3575, 3017]. Viscous [2811, 680, 766, 546, 1538, 2056, 2754, 170, 2781, 2660, 1358, 773, 835, 3247, 2679, 301, 3347, 2977, 1270, 20, 2111, 809, 719, 896, 2677, 3305, 2659]. Vlasov [948, 3361]. Voigt [3507]. Voltage [3372, 3453]. Volterra [1512, 2198, 2376, 2452, 2501, 2500, 2178, 2735, 2233, 1396, 2406, 827, 3440, 3254, 2431, 2432, 43, 2671, 1327, 908, 1524, 3551, 1783, 3393]. Volterra-Type [3440]. Volume [3583]. Volumetric [3313]. Vortex [2932, 1640, 3553, 1209, 3058, 1242, 2642]. Vortical [2763, 1640, 3346, 2705]. Vortices [3202, 3203, 3407, 3466, 3190, 1206, 1242]. Vorticity [2658]. Voter [1841]. Voting [1667, 1959, 2017, 2138, 1841, 1590]. Vous [3118]. Vries [2767, 3116, 3143, 2632, 1075, 2826, 3436, 3421]. Vries-Like [2767].

Waals [2974, 3342]. Waiting [3125, 1753, 2051, 279, 2321, 2745]. Waiting-Time [2051, 2321]. Wakes [1209]. Walk [883, 346, 35, 581, 571, 1912, 2447, 2430, 347, 1148, 107, 1884]. Walks [75, 1074, 3337, 2820, 3350]. Wall [2524, 2801, 3435, 3133, 2423, 2518, 3552, 2977, 3057]. Walls [2662, 1859, 2801]. Walsh [770, 1452]. Waltman [65]. Warfare [1725]. Warm [2026]. Warm-Beam [2026]. Washout [2269, 3531]. Water [1131, 3044, 2812, 2766, 1094, 1095, 2748, 3087]. Wave [1724, 1255, 3579, 3293, 2472, 2471, 2411, 1413, 3163, 2071, 1594, 1269, 3127, 2496, 566, 3499, 2094, 2887, 3013, 1125, 2754, 2750, 2925, 2924, 1672, 1758, 3288, 559, 2484, 1072, 864, 2281, 3067, 2812, 3534, 1830, 2781, 2373, 1777, 970, 878, 1858, 2159, 1588, 1796, 1241, 255, 3034, 3247, 3116, 585, 1873, 1450, 2299, 1869, 2962, 341, 2592, 389, 1929, 2161, 947, 2405, 2616, 1575, 3526, 1230, 1957, 3242, 2133, 2442, 3355, 2860, 1090, 1359, 563, 598, 974, 726, 756, 820, 3362, 2699, 2829, 1261, 3441, 1238, 3285, 3318, 1154, 1516, 3383, 2877, 2995, 2688, 3279, 2898, 1598, 1204]. Wave [54, 3448, 2541, 3300, 2828, 3154, 3314, 2861, 2913, 2799]. Wave-Blocking [2592]. Wave-Guides [1516]. Wave-Like [2411]. Wave-Splitting [3279]. Wavefield [2522]. Wavefields [2859]. Waveform [1104]. Waveforms [2265]. Wavefronts [3099]. Waveguide [1258, 1504, 1878, 1480, 288, 3266, 276]. Waveguides [1669, 772, 2114, 2706, 1656]. Wavelength [3445]. Wavelet [3381, 3380]. Wavenumbers [1669]. Waves [1046, 2798, 2916, 3064, 1566, 1722, 1221, 2178, 2817, 2359, 2459, 3023, 2219, 2378, 1979, 2773, 1222, 1538, 1584, 3078, 1645, 1867, 2308, 3553, 1940, 2982, 1149, 1150, 1378, 3062, 3044, 2131, 2372, 3555, 2231, 2608, 2234, 3277, 2189, 2882, 3411, 2425, 2892, 3146, 1651, 2858, 1824, 3518, 1311, 1854, 1547, 585, 3143, 2864, 2542, 2018, 1982, 2800, 3257, 2114, 2287, 2467, 3519, 1871, 2371, 3162, 2766, 3038, 1846, 1021, 2908, 1272, 2288, 2748, 2825, 2042, 1999, 3517, 2715, 3502, 1281, 1729, 2069, 1681, 2438, 2634, 2473, 2485, 3031, 3121, 3135, 2040, 3410, 2300, 2686, 623, 2129, 1343, 2697]. Waves [3216, 2001, 610, 828, 3418, 393, 3052, 3120, 2981, 1907, 1041, 3007, 3309, 3437, 3089, 2240, 3126, 1855, 3185, 79, 1228, 574, 867, 973, 1291, 1377, 2343, 3344]. Wavetrains [2174]. Way [223]. Wazewski [665]. Weak [3449, 559, 261, 878, 1720, 1915, 3032, 872, 2813, 837]. Weakly [3144, 3249, 2508, 2151, 2550, 3255, 1399, 2170, 2925, 851, 3192, 2967, 1873, 3143, 2652, 2112, 2962, 3581, 3162, 2303, 218, 2943, 234, 1088, 2129, 2697, 2536, 3216, 2541, 2507]. Weber [3377]. Webs [2335]. Wedge [2071, 1012, 1067, 1583, 1090, 1261, 821, 1808, 1291]. Wedge-Shaped [1583]. Wedges [2558]. Weierstrass [664, 212]. Weighing [1318]. Weight [986, 1026, 1010, 13, 140]. Weighted [649, 82, 2079, 2401, 1490, 2230, 724, 445, 775, 1515, 738]. Weighted-Signs [1490]. Weighting [798, 833, 251, 609, 51]. Weights [2491, 296, 775]. Wetting [3400, 2904]. Where [1637]. Which [878, 2381, 912, 625, 579, 1203, 872, 1768, 171, 247, 312, 219, 121, 2240, 3002, 355, 941]. Whirling [509, 128, 793]. Whispering [2706]. Whispering-Gallery [2706]. White [3293]. Whitham [2923, 3580]. Whole [648]. Whose [10, 1445, 3216, 1632]. Wide [3016, 1595, 247, 1133, 1162, 1232, 1321]. Wide-Band [1595]. Wide-Sense [1133]. Width [2411]. Wiener [209, 2721, 3330, 1961, 1849, 2201, 1373, 2431, 2432, 2314, 1524, 1600, 2320]. Wigner [1993]. Wilkins [956]. Wind [567, 321]. Wings [3542]. Wire [2707, 692]. Wires [1867, 2145]. Within [134]. Without [540, 444, 1366, 1243, 1489, 3354, 1248, 496, 927, 335, 1033, 433, 2024, 717]. WKB [2836, 1075]. Wordlength [131]. Work [3340, 3341]. Wright [1263].

X [3256, 2957]. X-Ray [2957, 3256].

Yield [3108]. Yields [607]. Young [1972].

Zakai [1309, 3454]. Zakharov [2505, 1389]. Zernike [46]. Zero [469, 1848, 2211, 2579, 683, 1725, 852, 853, 103, 2478]. Zero-Crossings [103]. Zero-Sum [852, 853]. Zeros [250, 405, 2784, 656, 499]. Zhabotinskii [1461, 1226]. Zhabotinskiĭ [1777, 1824]. Zonal [268]. Zone [2835, 2442, 3347].

References

Erdelyi:1956:AEF

[1] A. Erdélyi. Asymptotic expansions of Fourier integrals involving logarithmic singularities. Journal of the Society for Industrial and Applied Mathematics, 4 (1):38–47, March 1956. CODEN JSI-MAV. ISSN 0368-4245 (print), 1095-712X (electronic). See note [189].

Griffin:1963:NSP

[2] D. S. Griffin and R. S. Varga. Numerical solution of plane elasticity problems. Journal of the Society for Industrial and Applied Mathematics, 11 (4):1046–1062, December 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic). See remark [37].

Moon:1965:DCR

[3] J. W. Moon. On the distribution of crossings in random complete graphs. Journal of the Society for Industrial and Applied Mathematics, 13(2):506– 510, June 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic). See erratum [1494].

Williams:1965:SLP

[4] A. C. Williams. On stochastic linear programming. Journal of the Society for Industrial and Applied Mathematics, 13(4):927–940, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic). See errata [139].

Sack:1966:FLE

[5] R. A. Sack. Factorization of Lagrange’s expansion by means of exponential generating functions. SIAM Journal on Applied Mathematics, 14(1):1–15, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jones:1966:ASC

[6] John Jones, Jr. On the asymptotic stability of certain second order nonlinear differential equations. SIAM Journal on Applied Mathematics, 14 (1):16–22, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1966:IRL

[7] Harvey A. Smith. Improvement of the resolution of a linear scanning device. SIAM Journal on Applied Mathematics, 14(1):23–40, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1966:DLM

[8] A. H. Zemanian. The distributional Laplace and Mellin transformations. SIAM Journal on Applied Mathematics, 14(1):41–59, January 1966. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [138].

Sanchez:1966:CVP

[9] David A. Sanchez. On composite variational problems. SIAM Journal on Applied Mathematics, 14(1):60–64, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goodwin:1966:REI

[10] Bruce E. Goodwin. On the realization of the eigenvalues of integral equations whose kernels are entire or meromorphic in the eigenvalue parameter. SIAM Journal on Applied Mathematics, 14(1):65–85, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Newell:1966:Q

[11] G. F. Newell. The M/G/∞ queue. SIAM Journal on Applied Mathematics, 14(1):86–88, January 1966. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wets:1966:PUUa

[12] R. J. B. Wets. Programming under uncertainty: The equivalent convex program. SIAM Journal on Applied Mathematics, 14(1):89–105, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Weiss:1966:LCC

[13] E. Weiss. Linear codes of constant weight. SIAM Journal on Applied Mathematics, 14(1):106–111, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kozin:1966:UPE

[14] C. H. Kozin and G. R. Cooper. The use of prior estimates in successive point estimations of a random signal. SIAM Journal on Applied Mathematics, 14 (1):112–130, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Westbrook:1966:AAI

[15] D. R. Westbrook. Applications of asymptotic integration to potential problems for shells. SIAM Journal on Applied Mathematics, 14(1):131– 146, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cottle:1966:NPP

[16] Richard W. Cottle. Nonlinear programs with positively bounded jacobians. SIAM Journal on Applied Mathematics, 14(1):147–158, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1966:IFD

[17] A. H. Zemanian. Inversion formulas for the distributional Laplace transformation. SIAM Journal on Applied Mathematics, 14(1):159–166, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Blum:1966:BLU

[18] Marvin Blum. Best linear unbiased estimation by recursive methods. SIAM Journal on Applied Mathematics, 14 (1):167–180, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ponstein:1966:EEP

[19] J. Ponstein. Existence of equilibrium points in non-product spaces. SIAM Journal on Applied Mathematics, 14 (1):181–190, January 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schwiderski:1966:AVF

[20] E. W. Schwiderski, H. J. Lugt, and P. Ugincius. Axisymmetric viscous fluid motions around conical surfaces. SIAM Journal on Applied Mathematics, 14(1):191–208, January 1966. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wong:1966:SPS

[21] James S. W. Wong. Some properties of solutions of u′′(t) + a(t)f(u)g(u′) = 0. III. SIAM Journal on Applied Mathematics, 14(2):209–214, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gupta:1966:FPN

[22] Ram Prakash Gupta. On flows in pseudosymmetric networks. SIAM Journal on Applied Mathematics, 14 (2):215–225, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moon:1966:NAS

[23] J. W. Moon. A note on approximating symmetric relations by equivalence relations. SIAM Journal on Applied Mathematics, 14(2):226–227, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pitcher:1966:IEL

[24] Tom S. Pitcher. An integral expression for the log likelihood ratio of two Gaussian processes. SIAM Journal on Applied Mathematics, 14(2):228–233, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eppen:1966:CPT

[25] Gary D. Eppen. Convexity-preserving transformations with stochastic matrices. SIAM Journal on Applied Mathematics, 14(2):234–236, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bollman:1966:SPP

[26] D. Bollman. Some periodicity properties of modules over the ring of polynomials with coefficients in a residue class ring. SIAM Journal on Applied Mathematics, 14(2):237–241, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bailey:1966:SLE

[27] Paul B. Bailey. Sturm–Liouville eigenvalues via a phase function. SIAM Journal on Applied Mathematics, 14 (2):242–249, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Harary:1966:CPT

[28] Frank Harary and Charles A. Trauth, Jr. Connectedness of products of two directed graphs. SIAM Journal on Applied Mathematics, 14(2):250–254, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hanan:1966:SPR

[29] M. Hanan. On Steiner’s problem with rectilinear distance. SIAM Journal on Applied Mathematics, 14(2):255–265, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sann:1966:FTE

[30] Robert I. Sann and Bernard W. Shaffer. Free transverse elastic vibrations of a solid cylinder bonded to a thin casing. SIAM Journal on Applied Mathematics, 14(2):266–285, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stafford:1966:GMI

[31] Roger O. Stafford and Kwan Rim. A general method of improving the accuracy of approximate solutions of circular cylindrical shells. SIAM Journal on Applied Mathematics, 14(2):286–305, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wu:1966:BMP

[32] Ting-Shu Wu and Bruno A. Boley. Bounds in melting problems with arbitrary rates of liquid removal. SIAM Journal on Applied Mathematics, 14 (2):306–323, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stadter:1966:BER

[33] James T. Stadter. Bounds to eigenvalues of rhombical membranes. SIAM Journal on Applied Mathematics, 14 (2):324–341, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Loud:1966:GIG

[34] W. S. Loud. Generalized inverses and generalized Green’s functions. SIAM Journal on Applied Mathematics, 14 (2):342–369, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Haji-Sheikh:1966:FRW

[35] A. Haji-Sheikh and E. M. Sparrow. The floating random walk and its application to Monte Carlo solutions of heat equations. SIAM Journal on Applied Mathematics, 14(2):370–389, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Duffin:1966:NST

[36] R. J. Duffin, D. Hazony, and N. Morrison. Network synthesis through hybrid matrices. SIAM Journal on Applied Mathematics, 14(2):390–413, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bueckner:1966:RPD

[37] Hans F. Bueckner. Remark on a paper by D. S. Griffin and R. S. Varga “Numerical Solution of Plane Elasticity Problems” [J. SIAM 11 (1963), no. 4, 1046–1062]. SIAM Journal on Applied Mathematics, 14(2):414, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [2].

Griffi:1966:CBR

[38] D. S. Griffi and R. S. Varga. Comments on Bueckner’s “Remark on a Paper by D. S. Griffin and R. S. Varga” [SIAM J. Appl. Math. 14 (1966), no. 2, 414– 414]. SIAM Journal on Applied Mathematics, 14(2):415–416, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wallen:1966:NFC

[39] Lawrence J. Wallen. A note on finite convolution operators. SIAM Journal on Applied Mathematics, 14(3):417– 419, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mond:1966:SDM

[40] Bertram Mond and Richard W. Cottle. Self-duality in mathematical programming. SIAM Journal on Applied Mathematics, 14(3):420–423, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huston:1966:NKB

[41] Ronald L. Huston and Louis F. Doty. Note on the Krylov–Bogoliubov method applied to linear differential equations. SIAM Journal on Applied Mathematics, 14(3):424–428, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Greenstadt:1966:RGM

[42] J. L. Greenstadt. A ricocheting gradient method for nonlinear optimization. SIAM Journal on Applied Mathematics, 14(3):429–445, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miller:1966:VPE

[43] R. K. Miller. On Volterra’s population equation. SIAM Journal on Applied Mathematics, 14(3):446–452, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Erdelyi:1966:PIF

[44] Ivan Erdélyi. On partial isometries in finite-dimensional Euclidean spaces. SIAM Journal on Applied Mathematics, 14(3):453–467, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sibuya:1966:BDT

[45] Yasutaka Sibuya. A block-diagonalization theorem for systems of linear ordinary differential equations and its applications. SIAM Journal on Applied Mathematics, 14(3):468–475, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Myrick:1966:GRP

[46] Donal R. Myrick. A generalization of the radial polynomials of F. zernike. SIAM Journal on Applied Mathematics, 14(3):476–489, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ma:1966:FSS

[47] Er-Chieh Ma. A finite series solution of the matrix equation AX − XB = C. SIAM Journal on Applied Mathematics, 14(3):490–495, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Potter:1966:MQS

[48] James E. Potter. Matrix quadratic solutions. SIAM Journal on Applied Mathematics, 14(3):496–501, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jeffrey:1966:DEF

[49] A. Jeffrey. On the determination of the envelope of a family of epitrochoids with applications. SIAM Journal on Applied Mathematics, 14(3):502–510, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Springer:1966:DPI

[50] M. D. Springer and W. E. Thompson. The distribution of products of independent random variables. SIAM Journal on Applied Mathematics, 14 (3):511–526, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Youla:1966:SLD

[51] D. C. Youla. The synthesis of linear dynamical systems from prescribed weighting patterns. SIAM Journal on Applied Mathematics, 14(3):527–549, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1966:DTD

[52] Wai-Kai Chen. On directed trees and directed k-trees of a digraph and their generation. SIAM Journal on Applied Mathematics, 14(3):550–560, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1966:DHT

[53] A. H. Zemanian. A distributional Hankel transformation. SIAM Journal on Applied Mathematics, 14(3):561–576, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wolfe:1966:DSW

[54] Peter Wolfe. Diffraction of a scalar wave by a plane screen. SIAM Journal on Applied Mathematics, 14(3):577– 599, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ponstein:1966:SAP

[55] J. Ponstein. Self-avoiding paths and the adjacency matrix of a graph. SIAM Journal on Applied Mathematics, 14 (3):600–609, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Heap:1966:SPNa

[56] B. R. Heap and M. S. Lynn. The structure of powers of nonnegative matrices: I. the index of convergence. SIAM Journal on Applied Mathematics, 14(3):610–639, May 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Danskin:1966:TMM

[57] John M. Danskin. The theory of maxmin, with applications. SIAM Journal on Applied Mathematics, 14(4):641– 664, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bram:1966:LMT

[58] Joseph Bram. The Lagrange multiplier theorem for max-min with several constraints. SIAM Journal on Applied Mathematics, 14(4):665–667, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Williams:1966:AFS

[59] A. C. Williams. Approximation formulas for stochastic linear programming. SIAM Journal on Applied Mathematics, 14(4):668–677, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1966:HTC

[60] A. H. Zemanian. The Hankel transformation of certain distributions of rapid growth. SIAM Journal on Applied Mathematics, 14(4):678–690, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dou:1966:MUC

[61] A. Dou. Method of undetermined coefficients in linear differential systems and the matrix equation Y B−AY = F. SIAM Journal on Applied Mathematics, 14(4):691–696, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gelbard:1966:NCM

[62] E. M. Gelbard, L. A. Ondis, II, and J. Spanier. A new class of Monte Carlo estimators. SIAM Journal on Applied Mathematics, 14(4):697–701, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Spanier:1966:TPF

[63] J. Spanier. Two pairs of families of estimators for transport problems. SIAM Journal on Applied Mathematics, 14(4):702–713, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See note [144].

Jagerman:1966:BTE

[64] D. Jagerman. Bounds for truncation error of the sampling expansion. SIAM Journal on Applied Mathematics, 14(4):714–723, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wong:1966:TTW

[65] James S. W. Wong. On two theorems of waltman. SIAM Journal on Applied Mathematics, 14(4):724–728, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Weinberg:1966:MOA

[66] Louis Weinberg. On the maximum order of the automorphism group of a planar triply connected graph. SIAM Journal on Applied Mathematics, 14 (4):729–738, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hazony:1966:TMS

[67] D. Hazony and R. D. Joseph. Transfer matrix synthesis with active RC networks. SIAM Journal on Applied Mathematics, 14(4):739–761, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Heap:1966:SPNb

[68] B. R. Heap and M. S. Lynn. The structure of powers of nonnegative matrices: II. the index of maximum density. SIAM Journal on Applied Mathematics, 14(4):762–777, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chartrand:1966:GTA

[69] Gary Chartrand. A graph-theoretic approach to a communications problem. SIAM Journal on Applied Mathematics, 14(4):778–781, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stevens:1966:LOD

[70] Karl K. Stevens. On linear ordinary differential equations with periodic coefficients. SIAM Journal on Applied Mathematics, 14(4):782–795, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Scroggs:1966:ADP

[71] James E. Scroggs and Patrick L. Odell. An alternate definition of a pseudoinverse of a matrix. SIAM Journal on Applied Mathematics, 14(4):796–810, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Winder:1966:PSH

[72] R. O. Winder. Partitions of N-space by hyperplanes. SIAM Journal on Applied Mathematics, 14(4):811–818, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ponter:1966:IES

[73] A. R. S. Ponter. An integral equation solution of the inhomogeneous torsion problem. SIAM Journal on Applied Mathematics, 14(4):819–830, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lanczos:1966:BVP

[74] C. Lanczos. Boundary value problems and orthogonal expansions. SIAM Journal on Applied Mathematics, 14 (4):831–863, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Feller:1966:IDD

[75] William Feller. Infinitely divisible distributions and Bessel functions associated with random walks. SIAM Journal on Applied Mathematics, 14(4):864– 875, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lin:1966:MTG

[76] C. C. Lin. On the mathematical theory of a galaxy of stars. SIAM Journal on Applied Mathematics, 14(4):876–921, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schiffer:1966:HOD

[77] Menahem Schiffer. Half-order differentials on Riemann surfaces. SIAM Journal on Applied Mathematics, 14 (4):922–934, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grad:1966:HFS

[78] Harold Grad. High frequency sound according to the Boltzmann equation. SIAM Journal on Applied Mathematics, 14(4):935–955, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Whitham:1966:NDW

[79] G. B. Whitham. Nonlinear dispersive waves. SIAM Journal on Applied Mathematics, 14(4):956–958, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

John:1966:RIM

[80] F. John. The role of inequalities in the mathematical theory of elasticity. SIAM Journal on Applied Mathematics, 14(4):959, July 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boylan:1966:EUT

[81] Edward S. Boylan. Existence and uniqueness theorems for the optimal inventory equation. SIAM Journal on Applied Mathematics, 14(5):961–969, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Deuel:1966:SDP

[82] D. R. Deuel and A. Gill. Some decision problems associated with weighted, directed graphs. SIAM Journal on Applied Mathematics, 14(5):970–979, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gold:1966:CLS

[83] Robert Gold. Characteristic linear sequences and their coset functions. SIAM Journal on Applied Mathematics, 14(5):980–985, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bragg:1966:RHE

[84] Louis R. Bragg. The radial heat equation and Laplace transforms. SIAM Journal on Applied Mathematics, 14 (5):986–993, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hansalik:1966:AP

[85] William Hansalik. Adaption procedures. SIAM Journal on Applied Mathematics, 14(5):994–997, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Swerling:1966:TGL

[86] P. Swerling. Topics in generalized least squares signal estimation. SIAM Journal on Applied Mathematics, 14(5): 998–1031, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maybee:1966:NGJ

[87] John S. Maybee. New generalizations of Jacobi matrices. SIAM Journal on Applied Mathematics, 14(5): 1032–1037, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sigillito:1966:PBS

[88] V. G. Sigillito. Pointwise bounds for solutions of the first initial-boundary value problem for parabolic equations. SIAM Journal on Applied Mathematics, 14(5):1038–1056, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chaudhry:1966:FDA

[89] H. R. Chaudhry. Finite deformations around a hole in a circular sheet containing a polarized dielectric subjected to radial forces. SIAM Journal on Applied Mathematics, 14(5): 1057–1066, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Veinott:1966:OIP

[90] Arthur F. Veinott, Jr. On the optimality of (s, S) inventory policies: New conditions and a new proof. SIAM Journal on Applied Mathematics, 14(5):1067–1083, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Willett:1966:BSE

[91] D. W. Willett and J. S. W. Wong. The boundedness of solutions of the equation ẍ+f(x, ẋ)+g(x) = 0. SIAM Journal on Applied Mathematics, 14(5): 1084–1098, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schechter:1966:BLP

[92] Murray Schechter. A boundary layer phenomenon in nonlinear membrane theory. SIAM Journal on Applied Mathematics, 14(5):1099–1114, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lawler:1966:CPD

[93] Eugene L. Lawler. Covering problems: Duality relations and a new method of solution. SIAM Journal on Applied Mathematics, 14(5): 1115–1132, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cox:1966:NAT

[94] D. R. Cox. A note on the analysis of a type of reliability trial. SIAM Journal on Applied Mathematics, 14(5): 1133–1142, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wets:1966:PUUb

[95] Roger Wets. Programming under uncertainty: The solution set. SIAM Journal on Applied Mathematics, 14 (5):1143–1151, September 1966. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mullen:1966:DRF

[96] James A. Mullen. The differential recursion formulae for Appell’s hypergeometric functions of two variables. SIAM Journal on Applied Mathematics, 14(5):1152–1163, September 1966.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1966:ITA

[97] Donald S. Cohen. An integral transform associated with boundary conditions containing an eigenvalue parameter. SIAM Journal on Applied Mathematics, 14(5):1164–1175, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meeter:1966:TUN

[98] Duane A. Meeter. On a theorem used in nonlinear least squares. SIAM Journal on Applied Mathematics, 14(5): 1176–1179, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Deutsch:1966:SIA

[99] Frank Deutsch. Simultaneous interpolation and approximation in topological linear spaces. SIAM Journal on Applied Mathematics, 14(5): 1180–1190, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Singh:1966:SES

[100] Harbhajan Singh. Symmetrical expansion of a spherical dielectric with faces maintained at different temperatures and potentials. SIAM Journal on Applied Mathematics, 14(5): 1191–1198, September 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Swerling:1966:CSP

[101] P. Swerling. Classes of signal processing procedures suggested by exact minimum mean square error procedures.

SIAM Journal on Applied Mathematics, 14(6):1199–1224, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zahl:1966:BCL

[102] Samuel Zahl. Bounds for the central limit theorem error. SIAM Journal on Applied Mathematics, 14(6): 1225–1245, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wong:1966:SRC

[103] E. Wong. Some results concerning the zero-crossings of Gaussian noise. SIAM Journal on Applied Mathematics, 14(6):1246–1254, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1966:SAT

[104] A. H. Zemanian. Some Abelian theorems for the distributional Hankel and K transformations. SIAM Journal on Applied Mathematics, 14(6): 1255–1265, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Konheim:1966:ODA

[105] Alan G. Konheim and Benjamin Weiss. An occupancy discipline and applications. SIAM Journal on Applied Mathematics, 14(6):1266–1274, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keller:1966:BCP

[106] Herbert B. Keller. Buckling of a circular plate under an unusual combined load. SIAM Journal on Applied Mathematics, 14(6):1275–1292, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rudemo:1966:AEP

[107] Mats Rudemo. On an absorption and emission problem for random walk. SIAM Journal on Applied Mathematics, 14(6):1293–1297, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ludeke:1966:CIS

[108] C. A. Ludeke and J. E. Cornett. A computer investigation of a subharmonic bifurcation point in the Duffing equation. SIAM Journal on Applied Mathematics, 14(6):1298–1306, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Duffin:1966:DTG

[109] R. J. Duffin and Elmor L. Peterson. Duality theory for geometric programming. SIAM Journal on Applied Mathematics, 14(6):1307–1349, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1966:DT

[110] A. H. Zemanian. A distributional K transformation. SIAM Journal on Applied Mathematics, 14(6):1350–1365, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [184].

Zaroodny:1966:TIP

[111] Serge J. Zaroodny. Trajectory indicator — A proposal. SIAM Journal on Applied Mathematics, 14(6): 1366–1389, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Karp:1966:PMP

[112] Richard M. Karp and Raymond E. Miller. Properties of a model for parallel computations: Determinancy, termination, queueing. SIAM Journal on Applied Mathematics, 14(6): 1390–1411, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hastings:1966:BGS

[113] S. P. Hastings. Bounds on the growth of solutions of a nonlinear system. SIAM Journal on Applied Mathematics, 14(6):1412–1417, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fox:1966:MRP

[114] Bennett Fox. Markov renewal programming by linear fractional programming. SIAM Journal on Applied Mathematics, 14(6):1418–1432, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McFadden:1966:DEG

[115] J. A. McFadden. A diagonal expansion in Gegenbauer polynomials for a class of second-order probability densities. SIAM Journal on Applied Mathematics, 14(6):1433–1436, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Davisson:1966:FTS

[116] Lee D. Davisson. Filtering of time series with unknown mean value. SIAM Journal on Applied Mathematics, 14 (6):1437–1445, November 1966. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dicker:1966:ADP

[117] Daniel Dicker. Analysis of distributed parameter networks — A general method. SIAM Journal on Applied Mathematics, 14(6):1446–1453, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bleistein:1966:STD

[118] Norman Bleistein and Robert M. Lewis. Space-time diffraction for dispersive hyperbolic equations. SIAM Journal on Applied Mathematics, 14 (6):1454–1470, November 1966. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boersma:1966:EDU

[119] J. Boersma. Electromagnetic diffraction by a unidirectionally conducting circular disk. SIAM Journal on Applied Mathematics, 14(6):1471–1495, November 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reid:1967:STO

[120] W. P. Reid. Stability of a towed object. SIAM Journal on Applied Mathematics, 15(1):1–12, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sellers:1967:ACW

[121] Peter Sellers. Algebraic complexes which characterize chemical networks. SIAM Journal on Applied Mathematics, 15(1):13–68, January 1967. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ranger:1967:SIR

[122] K. B. Ranger. Some integral relations for solutions of Laplace’s equation. SIAM Journal on Applied Mathematics, 15(1):69–74, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ponzo:1967:FOG

[123] Peter J. Ponzo. Forced oscillations of the generalized Lienard equation. SIAM Journal on Applied Mathematics, 15(1):75–87, January 1967. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Leneman:1967:SOS

[124] Oscar A. Z. Leneman. Second-order statistics of alternated stepwise processes. SIAM Journal on Applied Mathematics, 15(1):88–97, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dhaliwal:1967:AMB

[125] R. S. Dhaliwal. An axisymmetric mixed boundary value problem for a thick slab. SIAM Journal on Applied Mathematics, 15(1):98–106, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wu:1967:SPS

[126] Chien-Heng Wu and Robert Plunkett. On the solutions of plates, strips, rods, and cylinders subjected to arbitrary dynamic edge load. SIAM Journal on Applied Mathematics, 15 (1):107–119, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rulf:1967:DHU

[127] Benjamin Rulf. Diffraction by a halfplane in a uniaxial medium. SIAM Journal on Applied Mathematics, 15 (1):120–127, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Genin:1967:SNS

[128] J. Genin and J. S. Maybee. Some new stability results in the theory of whirling. SIAM Journal on Applied Mathematics, 15(1):128–135, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1967:FGS

[129] Wai kai Chen. On flow graph solutions of linear algebraic equations. SIAM Journal on Applied Mathematics, 15 (1):136–142, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boyce:1967:DAC

[130] William E. Boyce. A “dishonest” approach to certain stochastic eigenvalue problems. SIAM Journal on Applied Mathematics, 15(1):143–152, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Holick:1967:ELW

[131] A. Holick. Effect of limited wordlength on satellite prediction accuracy. SIAM Journal on Applied Mathematics, 15 (1):153–167, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gupta:1967:TTP

[132] Ram Prakash Gupta. Two theorems on pseudosymmetric graphs. SIAM Journal on Applied Mathematics, 15 (1):168–171, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mason:1967:CPA

[133] J. C. Mason. Chebyshev polynomial approximations for the L-membrane eigenvalue problem. SIAM Journal on Applied Mathematics, 15(1):172– 186, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rehm:1967:DAW

[134] Ronald G. Rehm. Density attenuation within interacting crossed beams of particles. SIAM Journal on Applied Mathematics, 15(1):187–192, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bush:1967:IMS

[135] S. F. Bush. An iterative method for the solution of sets of first-order hyperbolic differential equations. SIAM Journal on Applied Mathematics, 15 (1):193–206, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hu:1967:RMA

[136] T. C. Hu. Revised matrix algorithms for shortest paths. SIAM Journal on Applied Mathematics, 15(1):207–218, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [253].

Yau:1967:ASI

[137] W. F. Yau. Axisymmetric slipless indentation of an infinite, elastic cylinder. SIAM Journal on Applied Mathematics, 15(1):219–227, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1967:EDL

[138] A. H. Zemanian. Errata: “The Distributional Laplace and Mellin Transformations” [SIAM J. Appl. Math. 14 (1966), no. 1, 41–59]. SIAM Journal on Applied Mathematics, 15(1):227, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [8].

Williams:1967:ESL

[139] A. C. Williams. Errata: “On Stochastic Linear Programming” [J. SIAM 13 (1965), no. 4, 927–940]. SIAM Journal on Applied Mathematics, 15(1):228, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [4].

Weiss:1967:ALC

[140] Edwin Weiss. Addendum: Linear codes of constant weight. SIAM Journal on Applied Mathematics, 15 (1):229–230, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barnard:1967:ABP

[141] R. D. Barnard. On the asymptotic behavior of parameter-dependent functional equations. SIAM Journal on Applied Mathematics, 15(2):231–240, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Navot:1967:SCS

[142] Israel Navot. The synthesis of certain subclasses of tridiagonal matrices with prescribed eigenvalues. SIAM Journal on Applied Mathematics, 15 (2):241–251, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cinlar:1967:DSM

[143] E. Cinlar. Decomposition of a semi-Markov process under a state dependent rule. SIAM Journal on Applied Mathematics, 15(2):252–263, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

MacMillan:1967:NPJ

[144] D. B. MacMillan. Note on a paper by J. Spanier concerning Monte Carlo estimators “Two Pairs of Families of Estimators for Transport Problems” [SIAM J. Appl. Math. 14 (1966), no. 4, 702– 713]. SIAM Journal on Applied Mathematics, 15(2):264–268, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [63].

Fisher:1967:CPG

[145] Armen G. Fisher. On construction and properties of the generalized inverse. SIAM Journal on Applied Mathematics, 15(2):269–272, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bauer:1967:BIB

[146] Frances Bauer, Herbert B. Keller, and Edward L. Reiss. Boundary imperfections in the buckling of spherical caps. SIAM Journal on Applied Mathematics, 15(2):273–283, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Varaiya:1967:NPB

[147] P. P. Varaiya. Nonlinear programming in Banach space. SIAM Journal on Applied Mathematics, 15(2):284–293, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ritter:1967:DNP

[148] K. Ritter. Duality for nonlinear programming in a Banach space. SIAM Journal on Applied Mathematics, 15 (2):294–302, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Iglehart:1967:AM

[149] D. L. Iglehart and A. P. Lalchandani. An allocation model. SIAM Journal on Applied Mathematics, 15 (2):303–323, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1967:GCT

[150] A. H. Zemanian. A generalized convolution transformation. SIAM Journal on Applied Mathematics, 15(2):324– 346, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sreedharan:1967:SMM

[151] V. P. Sreedharan and H. H. Wein. A stochastic, multistage, multiproduct investment model. SIAM Journal on Applied Mathematics, 15(2):347–358, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Haimo:1967:SRG

[152] Deborah Tepper Haimo. Series representation of generalized temperature functions. SIAM Journal on Applied Mathematics, 15(2):359–367, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gustafson:1967:SIS

[153] Karl Gustafson. Stability inequalities for semimonotonically perturbed nonhomogeneous boundary problems. SIAM Journal on Applied Mathematics, 15(2):368–391, March 1967. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Payne:1967:SSN

[154] L. E. Payne. On the stability of solutions of the Navier–Stokes equations and convergence to steady state. SIAM Journal on Applied Mathematics, 15 (2):392–405, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Owens:1967:DMS

[155] A. B. Owens and H. M. Trent. On determining minimal singularities for the realizations of an incidence sequence. SIAM Journal on Applied Mathematics, 15(2):406–418, March 1967. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sattinger:1967:RPH

[156] D. H. Sattinger. On the Rayleigh problem in hydrodynamic stability. SIAM Journal on Applied Mathematics, 15 (2):419–425, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Coles:1967:CSA

[157] W. J. Coles and T. L. Sherman. Convergence of successive approximations for nonlinear two-point boundary value problems. SIAM Journal on Applied Mathematics, 15(2):426–433, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cover:1967:NLI

[158] Thomas M. Cover. The number of linearly inducible orderings of points in dspace. SIAM Journal on Applied Mathematics, 15(2):434–439, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Varaiya:1967:DCS

[159] P. P. Varaiya. On determinable classes of signals and linear channels. SIAM Journal on Applied Mathematics, 15 (2):440–449, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cumberbatch:1967:NET

[160] E. Cumberbatch. Nonlinear effects in transmission lines. SIAM Journal on Applied Mathematics, 15(2):450–463, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miele:1967:SPP

[161] Angelo Miele and David G. Hull. Sufficiency proofs for the problem of the optimum transversal contour. SIAM Journal on Applied Mathematics, 15 (2):464–477, March 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gowar:1967:NRT

[162] N. W. Gowar. A note on the response time of the vacuo-thermo-junction. SIAM Journal on Applied Mathematics, 15(3):479–483, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Marek:1967:POS

[163] Ivo Marek. u_{0}-positive operators and some of their applications. SIAM Journal on Applied Mathematics, 15 (3):484–494, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rose:1967:MS

[164] Milton E. Rose. On the melting of a slab. SIAM Journal on Applied Mathematics, 15(3):495–504, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fiacco:1967:SUM

[165] Anthony V. Fiacco and Garth P. Mc-Cormick. The slacked unconstrained minimization technique for convex programming. SIAM Journal on Applied Mathematics, 15(3):505–515, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1967:SET

[166] Norman Morrison. Smoothing and extrapolation of time series by means of discrete Laguerre polynomials. SIAM Journal on Applied Mathematics, 15 (3):516–538, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Osborne:1967:NMH

[167] M. R. Osborne. Numerical methods for hydrodynamic stability problems. SIAM Journal on Applied Mathematics, 15(3):539–557, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Benveniste:1967:PSD

[168] Jacques E. Benveniste. Projective solution of the Dirichlet problem for boundaries with angular points. SIAM Journal on Applied Mathematics, 15 (3):558–568, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maschler:1967:SKC

[169] Michael Maschler and Bezalel Peleg. The structure of the kernel of a cooperative game. SIAM Journal on Applied Mathematics, 15(3):569–604, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cole:1967:FVC

[170] J. D. Cole, H. B. Keller, and P. G. Saffman. The flow of a viscous compressible fluid through a very narrow gap. SIAM Journal on Applied Mathematics, 15(3):605–617, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vein:1967:FWS

[171] P. R. Vein. Functions which satisfy Abel’s differential equation. SIAM Journal on Applied Mathematics, 15 (3):618–623, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Capobianco:1967:TTP

[172] Michael F. Capobianco. Tournaments and tensor products of digraphs. SIAM Journal on Applied Mathematics, 15 (3):624–626, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Troesch:1967:FMS

[173] B. A. Troesch. Fluid motion in a shallow trapezoidal container. SIAM Journal on Applied Mathematics, 15 (3):627–636, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sellers:1967:SEM

[174] Peter H. Sellers. Stoichiometric equations are mathematical equations. SIAM Journal on Applied Mathematics, 15(3):637–640, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McCormick:1967:SOC

[175] Garth P. McCormick. Second order conditions for constrained minima. SIAM Journal on Applied Mathematics, 15(3):641–652, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Geoffrion:1967:RCP

[176] Arthur M. Geoffrion. Reducing concave programs with some linear constraints. SIAM Journal on Applied Mathematics, 15(3):653–664, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schaefer:1967:PBC

[177] Philip W. Schaefer. Pointwise bounds in the Cauchy problem for an elliptic system. SIAM Journal on Applied Mathematics, 15(3):665–677, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McGill:1967:PAF

[178] David J. McGill and Kenneth H. Lenzen. Polar axisymmetric free oscillations of thick hollowed tori. SIAM Journal on Applied Mathematics, 15 (3):678–692, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Karp:1967:FSP

[179] Richard M. Karp and Michael Held. Finite-state processes and dynamic programming. SIAM Journal on Applied Mathematics, 15(3):693–718, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Butz:1967:ISP

[180] Arthur R. Butz. Iterative saddle point techniques. SIAM Journal on Applied Mathematics, 15(3):719–725, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McLean:1967:ISF

[181] R. A. McLean and M. F. Neuts. The integral of a step function defined on a semi-Markov process. SIAM Journal on Applied Mathematics, 15(3):726– 737, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Perko:1967:MEE

[182] Lawrence M. Perko. A method of error estimation in singular perturbation problems with application to the restricted three body problem. SIAM Journal on Applied Mathematics, 15(3):738–753, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Desoer:1967:MRN

[183] C. A. Desoer and P. Varaiya. The minimal realization of a nonanticipative impulse response matrix. SIAM Journal on Applied Mathematics, 15 (3):754–764, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1967:EDT

[184] A. H. Zemanian. Errata: “A Distributional K Transformation” [SIAM J. Appl. Math. 14 (1966), no. 6, 1350– 1365]. SIAM Journal on Applied Mathematics, 15(3):765, May 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [110].

Kreith:1967:D

[185] Kurt Kreith. Disconjugacy in E^{n}. SIAM Journal on Applied Mathematics, 15(4):767–770, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Soong:1967:PSL

[186] T. T. Soong and D. A. Prelewicz. On perturbation solutions to linear system dynamics-II. SIAM Journal on Applied Mathematics, 15(4):771–774, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Posner:1967:PEC

[187] Edward C. Posner. Properties of error-correcting codes at low signal-to-noise ratios. SIAM Journal on Applied Mathematics, 15(4):775–798, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cannon:1967:DUP

[188] J. R. Cannon and D. L. Filmer. The determination of unknown parameters in analytic systems of ordinary differential equations. SIAM Journal on Applied Mathematics, 15(4):799–809, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McKenna:1967:NAE

[189] James McKenna. Note on “Asymptotic Expansions of Fourier Integrals Involving Logarithmic Singularities” [J. SIAM 4 (1956), no. 1, 38–47], by A. Erdélyi. SIAM Journal on Applied Mathematics, 15(4):810–812, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1].

Westbrook:1967:AAI

[190] D. R. Westbrook. Applications of asymptotic integration to potential problems in shells. II: The flexure problem of St. Venant. SIAM Journal on Applied Mathematics, 15(4):813–823, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Handelsman:1967:EFA

[191] Richard A. Handelsman and Joseph B. Keller. The electrostatic field around a slender conducting body of revolution.

SIAM Journal on Applied Mathematics, 15(4):824–841, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Francis:1967:MSC

[192] Philip H. Francis and Kwan Rim. A method of solving certain plane problems of contained elasto-plasticity. SIAM Journal on Applied Mathematics, 15(4):842–855, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reid:1967:GGM

[193] William T. Reid. Generalized Green’s matrices for two-point boundary problems. SIAM Journal on Applied Mathematics, 15(4):856–870, July 1967. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pullman:1967:PIP

[194] N. J. Pullman. A property of infinite products of Boolean matrices. SIAM Journal on Applied Mathematics, 15(4):871–873, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schay:1967:PMS

[195] Geza Schay, Jr. and Francis W. Dauer. A probabilistic model of a selforganizing file system. SIAM Journal on Applied Mathematics, 15(4):874– 888, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wong:1967:SSC

[196] James S. W. Wong. Some stability conditions for x′′ + a(t)x^{2n−1} = 0.

SIAM Journal on Applied Mathematics, 15(4):889–892, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Siddiqui:1967:SLL

[197] M. M. Siddiqui. Strong law of large numbers for branching processes. SIAM Journal on Applied Mathematics, 15(4):893–897, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mayoh:1967:SLE

[198] B. H. Mayoh. On the simplification of logical expressions. SIAM Journal on Applied Mathematics, 15(4):898–905, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sakarovitch:1967:EGU

[199] M. Sakarovitch and R. Saigal. An extension of generalized upper bounding techniques for structured linear programs. SIAM Journal on Applied Mathematics, 15(4):906–914, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ludwig:1967:EVS

[200] Donald Ludwig. An extension of the validity of the stationary phase formula. SIAM Journal on Applied Mathematics, 15(4):915–923, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1967:PAG

[201] A. H. Zemanian. Preface (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15(4):925–927, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bremermann:1967:SRA

[202] H. J. Bremermann. Some remarks on analytic representations and products of distributions (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15(4):929– 943, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

DeJager:1967:LIS

[203] E. M. De Jager. The lorentz-invariant solutions of the Klein–Gordon equation (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15(4):944–963, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Guttinger:1967:GFE

[204] W. Guttinger. Generalized functions in elementary particle physics and passive system theory: Recent trends and problems (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15(4):964–1000, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Newcomb:1967:DAT

[205] R. W. Newcomb and B. D. O. Anderson. A distributional approach to time-varying sensitivity (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15 (4):1001–1010, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Beltrami:1967:DOP

[206] E. J. Beltrami. Dissipative operators, positive real resolvents and the theory of distributions (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15 (4):1011–1017, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dolezal:1967:LPP

[207] Vaclav Dolezal. On linear passive nports with time-varying elements (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15(4):1018–1029, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Guttinger:1967:UFT

[208] W. Guttinger and E. Pfaffelhuber. Unrenormalizable field theories in terms of lorentz-invariant generalized functions (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15(4):1030–1045, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jaffe:1967:GPW

[209] Arthur M. Jaffe. A generalization of the Paley–Wiener–Schwartz theorem (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15(4):1046–1047, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Liverman:1967:PMD

[210] T. P. G. Liverman. Physically motivated definitions of distributions (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15(4):1048–1076, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Beltrami:1967:CIT

[211] E. J. Beltrami and M. R. Wohlers. The Cauchy integral of tempered distributions and some theorems on analytic continuation (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15 (4):1077–1087, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1967:GWT

[212] A. H. Zemanian. A generalized Weierstrass transformation (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15 (4):1088–1105, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1967:SDP

[213] A. H. Zemanian. A solution of a division problem arising from Bessel-type differential equations (in the applications of generalized functions). SIAM Journal on Applied Mathematics, 15 (4):1106–1111, July 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pruzan:1967:ADP

[214] P. M. Pruzan and B. Munch-Andersen. On the application of dynamic programming-type algorithms to antenna design. SIAM Journal on Applied Mathematics, 15(5):1113–1129, September 1967. CODEN SMJMAP.

ISSN 0036-1399 (print), 1095-712X (electronic).

Stippes:1967:TPP

[215] M. Stippes and D. Shadman. Traction problems in plane elasticity. SIAM Journal on Applied Mathematics, 15 (5):1130–1135, September 1967. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sigillito:1967:PIP

[216] V. G. Sigillito. A priori inequalities and pointwise bounds for solutions of fourth order elliptic partial differential equations. SIAM Journal on Applied Mathematics, 15(5):1136–1155, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mehndiratta:1967:SDM

[217] S. L. Mehndiratta. Self-duality in mathematical programming. SIAM Journal on Applied Mathematics, 15 (5):1156–1157, September 1967. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lazer:1967:CPS

[218] A. C. Lazer. On the computation of periodic solutions of weakly nonlinear differential equations. SIAM Journal on Applied Mathematics, 15(5): 1158–1170, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Martin:1967:SRM

[219] J. F. P. Martin. Some results on matrices which commute with their derivatives. SIAM Journal on Applied Mathematics, 15(5):1171–1183, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ben-Israel:1967:GSE

[220] Adi Ben-Israel. On the geometry of subspaces in Euclidean n-spaces. SIAM Journal on Applied Mathematics, 15(5):1184–1198, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See remarks [534].

Caughey:1967:EVA

[221] T. K. Caughey. Effect of vibration on the accuracy of a vertical reference pendulum. SIAM Journal on Applied Mathematics, 15(5):1199–1208, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Treve:1967:CSB

[222] Yvain M. Treve. On a class of solutions of the Blasius equation with applications. SIAM Journal on Applied Mathematics, 15(5):1209–1227, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rothschild:1967:MTW

[223] B. Rothschild and A. Whinston. Maximal two-way flows. SIAM Journal on Applied Mathematics, 15(5): 1228–1238, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Langenhop:1967:GIM

[224] C. E. Langenhop. On generalized inverses of matrices. SIAM Journal on Applied Mathematics, 15(5): 1239–1246, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Aberth:1967:TTD

[225] Oliver Aberth. The transformation of tensors into diagonal form. SIAM Journal on Applied Mathematics, 15(5): 1247–1252, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Willner:1967:PLP

[226] Leopold B. Willner. On parametric linear programming. SIAM Journal on Applied Mathematics, 15(5): 1253–1257, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bragg:1967:SSR

[227] L. R. Bragg. On the solution structure of radial heat problems with singular data. SIAM Journal on Applied Mathematics, 15(5):1258–1271, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1967:LIA

[228] Chi-Tsong Chen. Linear independence of analytic functions. SIAM Journal on Applied Mathematics, 15(5): 1272–1274, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tomastik:1967:ONS

[229] E. C. Tomastik. Oscillation of a nonlinear second order differential equation. SIAM Journal on Applied Mathematics, 15(5):1275–1277, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jaiswal:1967:FSP

[230] N. K. Jaiswal and K. Thiruvengadam. Finite-source priority queues. SIAM Journal on Applied Mathematics, 15 (5):1278–1293, September 1967. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tyndall:1967:EDT

[231] William F. Tyndall. An extended duality theorem for continuous linear programming problems. SIAM Journal on Applied Mathematics, 15(5): 1294–1298, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Walkup:1967:SPR

[232] David W. Walkup and Roger J.- B. Wets. Stochastic programs with recourse. SIAM Journal on Applied Mathematics, 15(5):1299–1314, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Balogh:1967:AEM

[233] Charles B. Balogh. Asymptotic expansions of the modified Bessel function of the third kind of imaginary order. SIAM Journal on Applied Mathematics, 15(5):1315–1323, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Musa:1967:ICW

[234] Samuel A. Musa. Integral constraints in weakly nonlinear periodic systems. SIAM Journal on Applied Mathematics, 15(5):1324–1327, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Scarf:1967:AFP

[235] Herbert Scarf. The approximation of fixed points of a continuous mapping. SIAM Journal on Applied Mathematics, 15(5):1328–1343, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Passy:1967:GPO

[236] U. Passy and D. J. Wilde. Generalized polynomial optimization. SIAM Journal on Applied Mathematics, 15(5): 1344–1356, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gibson:1967:ESC

[237] Walter E. Gibson. An exact solution for a class of stochastic partial differential equations. SIAM Journal on Applied Mathematics, 15(5): 1357–1362, September 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morgan:1967:SPP

[238] R. C. Morgan, F. C. Karal, and S. N. Karp. Solution to the phenomenological problem of a magnetic line source above a plane structure that supports N-excited modes. SIAM Journal on Applied Mathematics, 15(6): 1363–1377, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Weiner:1967:NPD

[239] Howard J. Weiner. Nonlinear predictors depending on the last observation. SIAM Journal on Applied Mathematics, 15(6):1378–1384, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bucy:1967:TPB

[240] R. S. Bucy. Two-point boundary value problems of linear Hamiltonian systems. SIAM Journal on Applied Mathematics, 15(6):1385–1389, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kamae:1967:GTP

[241] T. Kamae. A graph-theoretical proof of an adjacency property of major submatrices. SIAM Journal on Applied Mathematics, 15(6):1390–1399, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Swanson:1967:EAE

[242] C. A. Swanson and V. B. Headley. An extension of Airy’s equation. SIAM Journal on Applied Mathematics, 15 (6):1400–1412, November 1967. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hearon:1967:SCB

[243] John Z. Hearon. On the singularity of a certain bordered matrix. SIAM Journal on Applied Mathematics, 15(6): 1413–1421, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bleistein:1967:UAE

[244] Norman Bleistein and Richard A. Handelsman. Uniform asymptotic expansions of certain integrals. SIAM Journal on Applied Mathematics, 15(6): 1422–1433, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wolfe:1967:NAE

[245] Peter Wolfe. A new approach to edge diffraction. SIAM Journal on Applied Mathematics, 15(6):1434–1469, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kadota:1967:SDT

[246] T. T. Kadota. Simultaneous diagonalization of two covariance kernels and application to second order stochastic processes. SIAM Journal on Applied Mathematics, 15(6): 1470–1480, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Franaszek:1967:LSW

[247] P. A. Franaszek. On linear systems which preserve wide sense stationarity. SIAM Journal on Applied Mathematics, 15(6):1481–1484, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bernstein:1967:MCA

[248] A. J. Bernstein. Maximally connected arrays on the n-cube. SIAM Journal on Applied Mathematics, 15(6): 1485–1489, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miller:1967:RBB

[249] K. S. Miller and Harold Sackrowitz. Relationships between biased and unbiased Rayleigh distributions. SIAM Journal on Applied Mathematics, 15 (6):1490–1495, November 1967. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Booth:1967:LZD

[250] R. S. Booth. Location of zeros of derivatives. SIAM Journal on Applied Mathematics, 15(6):1496–1501, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Trench:1967:WCP

[251] William F. Trench. Weighting coefficients for the prediction of stationary time series from the finite past. SIAM Journal on Applied Mathematics, 15(6):1502–1510, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cockayne:1967:PPC

[252] Ernest Cockayne. Plane pursuit with curvature constraints. SIAM Journal on Applied Mathematics, 15(6): 1511–1516, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hu:1967:ERM

[253] T. C. Hu. Errata: “Revised Matrix Algorithms for Shortest Paths” [SIAM J. Appl. Math. 15 (1967), no. 1, 207– 218]. SIAM Journal on Applied Mathematics, 15(6):1517, November 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [136].

Gilbert:1968:SMT

[254] E. N. Gilbert and H. O. Pollak. Steiner minimal trees. SIAM Journal on Applied Mathematics, 16(1):1–29, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gilbert:1968:IAP

[255] Robert P. Gilbert. An investigation of the analytic properties of solutions to the generalized axially symmetric, reduced wave equation in n + 1 variables, with an application to the theory of potential scattering. SIAM Journal on Applied Mathematics, 16 (1):30–50, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shirakawa:1968:CPG

[256] Isao Shirakawa, Tadao Kasami, and Hiroshi Ozaki. On the coding of the path in a graph and some applications. SIAM Journal on Applied Mathematics, 16(1):51–61, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Epstein:1968:URL

[257] M. P. Epstein. The use of resultants to locate extreme values of polynomials. SIAM Journal on Applied Mathematics, 16(1):62–70, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barr:1968:NII

[258] G. E. Barr. A note on integrals involving parabolic cylinder functions. SIAM Journal on Applied Mathematics, 16 (1):71–74, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Geer:1968:UAS

[259] James F. Geer and Joseph B. Keller. Uniform asymptotic solutions for potential flow around a thin airfoil and the electrostatic potential about a thin conductor. SIAM Journal on Applied Mathematics, 16(1):75–101, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peleg:1968:CG

[260] Bezalel Peleg. Centralized games. SIAM Journal on Applied Mathematics, 16(1):102–104, January 1968. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fiacco:1968:SOS

[261] Anthony V. Fiacco. Second order sufficient conditions for weak and strict constrained minima. SIAM Journal on Applied Mathematics, 16(1):105– 108, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burniston:1968:FIE

[262] E. E. Burniston. On the flexure of an infinite elastic plate containing a large unstressed circular hole. SIAM Journal on Applied Mathematics, 16 (1):109–125, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Srivastav:1968:DTI

[263] R. P. Srivastav and K. S. Parihar. Dual and triple integral equations involving inverse mellin transforms. SIAM Journal on Applied Mathematics, 16 (1):126–133, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Skibinsky:1968:MER

[264] Morris Skibinsky. Minimax estimation of a random probability. SIAM Journal on Applied Mathematics, 16 (1):134–145, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goldwyn:1968:GMS

[265] Roger M. Goldwyn and L. de L. Sloan. A geometrical method of solving second order nonlinear differential equations. SIAM Journal on Applied Mathematics, 16(1):146–154, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kemper:1968:APF

[266] Gene A. Kemper. Almost periodic functional differential equations. SIAM Journal on Applied Mathematics, 16 (1):155–161, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chu:1968:LOE

[267] J. T. Chu and K. Ya’coub. Linear order estimates using subsamples. SIAM Journal on Applied Mathematics, 16 (1):162–166, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lo:1968:SBZ

[268] Chi Yeung Lo. Singular behavior of zonal hyperspherical harmonic series. SIAM Journal on Applied Mathematics, 16(1):167–180, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Anderson:1968:NCC

[269] Douglas R. Anderson. A new class of cyclic codes. SIAM Journal on Applied Mathematics, 16(1):181–197, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1968:ME

[270] R. A. Smith. Matrix equation XA + BX = C. SIAM Journal on Applied Mathematics, 16(1):198–201, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moran:1968:RNV

[271] M. J. Moran and R. A. Gaggioli. Reduction of the number of variables in systems of partial differential equations, with auxiliary conditions. SIAM Journal on Applied Mathematics, 16 (1):202–215, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gill:1968:SLS

[272] Arthur Gill. Synthesis of linear sequential circuits from input-output relations. SIAM Journal on Applied Mathematics, 16(1):216–227, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hunt:1968:TER

[273] G. E. Hunt. The transport equation of radiative transfer with axial symmetry. SIAM Journal on Applied Mathematics, 16(1):228–237, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Varon:1968:PAL

[274] D. Varon, R. B. Kieburtz, and B. R-S. Cheo. A perturbation analysis of a log-periodic structure. SIAM Journal on Applied Mathematics, 16 (1):238–254, January 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Makky:1968:RST

[275] Sadia M. Makky. Rupture surfaces of a thick spherical shell under internal pressure. SIAM Journal on Applied Mathematics, 16(2):255–267, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yee:1968:RTR

[276] H. Y. Yee, L. B. Felsen, and J. B. Keller. Ray theory of reflection from the open end of a waveguide. SIAM Journal on Applied Mathematics, 16 (2):268–300, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Porsching:1968:SMA

[277] T. A. Porsching. On the spectrum of a matrix arising from a problem in reactor kinetics. SIAM Journal on Applied Mathematics, 16(2):301–317, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hanson:1968:DCI

[278] M. A. Hanson. Duality for a class of infinite programming problems. SIAM Journal on Applied Mathematics, 16 (2):318–323, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Marshall:1968:SRB

[279] K. T. Marshall. Some relationships between the distributions of waiting time, idle time and interoutput time in the GI/G/1 queue. SIAM Journal on Applied Mathematics, 16(2):324–327, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Basu:1968:RPS

[280] S. K. Basu. On reduction of program schemes. SIAM Journal on Applied Mathematics, 16(2):328–339, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moon:1968:CFR

[281] J. W. Moon and L. Moser. On the correlation function of random binary sequences. SIAM Journal on Applied Mathematics, 16(2):340–343, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Treve:1968:NBP

[282] Yvain M. Treve. On the nonlinear behavior of a pulsed MHD converter. SIAM Journal on Applied Mathematics, 16(2):344–362, March 1968. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Passy:1968:GPA

[283] U. Passy and D. J. Wilde. A geometric programming algorithm for solving chemical equilibrium problems. SIAM Journal on Applied Mathematics, 16 (2):363–373, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tu:1968:PSP

[284] Yih o Tu. Perturbation solution of plane stress problems in anisotropic elasticity. SIAM Journal on Applied Mathematics, 16(2):374–386, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brown:1968:TPR

[285] David P. Brown. Topological properties of resistance matrices. SIAM Journal on Applied Mathematics, 16 (2):387–394, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pfeffer:1968:SEH

[286] W. F. Pfeffer. On the stress equations for a heated rotating disc. SIAM Journal on Applied Mathematics, 16 (2):395–407, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shirakawa:1968:PDC

[287] Isao Shirakawa, Hiromitsu Takahashi, and Hiroshi Ozaki. On the planar decomposition of a complete bipartite graph. SIAM Journal on Applied Mathematics, 16(2):408–416, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lewin:1968:SSI

[288] L. Lewin. The solution of singular integral equations over a multiple interval and applications to multiple diaphragms in rectangular waveguide. SIAM Journal on Applied Mathematics, 16(2):417–438, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Donath:1968:SPP

[289] W. E. Donath. Statistical properties of the placement of a graph. SIAM Journal on Applied Mathematics, 16 (2):439–457, March 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Anonymous:1968:BM

[290] Anonymous. Back matter. SIAM Journal on Applied Mathematics, 16(3):??, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bragg:1968:RPD

[291] L. R. Bragg and J. W. Dettman. Related partial differential equations and their applications. SIAM Journal on Applied Mathematics, 16(3):459–467, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Denardo:1968:MMR

[292] E. V. Denardo and B. L. Fox. Multichain Markov renewal programs. SIAM Journal on Applied Mathematics, 16 (3):468–487, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1968:GTC

[293] G. H. Meyer. On a general theory of characteristics and the method of invariant imbedding. SIAM Journal on Applied Mathematics, 16(3):488–509, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boehme:1968:PLO

[294] T. K. Boehme and R. E. Powell. Positive linear operators generated by analytic functions. SIAM Journal on Applied Mathematics, 16(3):510–519, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Turner:1968:GMF

[295] James Turner. Generalized matrix functions and the graph isomorphism problem. SIAM Journal on Applied Mathematics, 16(3):520–526, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peleg:1968:WCS

[296] Bezalel Peleg. On weights of constantsum majority games. SIAM Journal on Applied Mathematics, 16(3):527–532, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Govindarajulu:1968:CGP

[297] Zakkula Govindarajulu. Certain general properties of unbiased estimates of location and scale parameters based on ordered observations. SIAM Journal on Applied Mathematics, 16(3):533–551, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Masry:1968:BHS

[298] E. Masry, K. Steiglitz, and B. Liu. Bases in Hilbert space related to the representation of stationary operators. SIAM Journal on Applied Mathematics, 16(3):552–562, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bresters:1968:DCQ

[299] D. W. Bresters. On distributions connected with quadratic forms. SIAM Journal on Applied Mathematics, 16 (3):563–581, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lucas:1968:GPF

[300] W. F. Lucas. A game in partition function form with no solution. SIAM Journal on Applied Mathematics, 16(3):582–585, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Olmstead:1968:CVF

[301] W. E. Olmstead. A class of viscous flows derivable from diffraction theory. SIAM Journal on Applied Mathematics, 16(3):586–595, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Comstock:1968:LMS

[302] Craig Comstock. On Lighthill’s method of strained coordinates. SIAM Journal on Applied Mathematics, 16(3):596– 602, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1968:UTE

[303] Wai-Kai Chen. On unisignants and their evaluation. SIAM Journal on Applied Mathematics, 16(3):603–619, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Conley:1968:RCS

[304] C. C. Conley. The retrograde circular solutions of the restricted threebody problem via a submanifold convex to the flow. SIAM Journal on Applied Mathematics, 16(3):620–625, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Campbell:1968:STF

[305] L. L. Campbell. Sampling theorem for the Fourier transform of a distribution with bounded support. SIAM Journal on Applied Mathematics, 16 (3):626–636, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Birnbaum:1968:PIM

[306] Z. W. Birnbaum and Sam C. Saunders. A probabilistic interpretation of Miner’s rule. SIAM Journal on Applied Mathematics, 16(3):637–652, May 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bunker:1968:SMA

[307] Don L. Bunker. Some molecular applications of Monte Carlo methods. SIAM Journal on Applied Mathematics, 16(4):653–661, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Duffin:1968:FLP

[308] R. J. Duffin and L. A. Karlovitz. Formulation of linear programs in analysis. I: Approximation theory. SIAM Journal on Applied Mathematics, 16 (4):662–675, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Horwitz:1968:DMH

[309] Lawrence B. Horwitz and Philip E. Sarachik. Davidon’s method in Hilbert space. SIAM Journal on Applied Mathematics, 16(4):676–695, July 1968. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chartrand:1968:RTG

[310] Gary Chartrand and Hudson V. Kronk. Randomly traceable graphs. SIAM Journal on Applied Mathematics, 16 (4):696–700, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lewis:1968:PPS

[311] T. O. Lewis and T. G. Newman. Pseudoinverses of positive semidefinite matrices. SIAM Journal on Applied Mathematics, 16(4):701–703, July 1968. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

MacMillan:1968:AMS

[312] D. B. MacMillan. Asymptotic methods for systems of differential equations in which some variables have very short response times. SIAM Journal on Applied Mathematics, 16(4):704–722, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Katsanis:1968:NMS

[313] Theodore Katsanis. A numerical method for the solution of certain Neumann problems. SIAM Journal on Applied Mathematics, 16(4):723–731, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Conley:1968:LET

[314] C. C. Conley. Low energy transit orbits in the restricted three-body problem. SIAM Journal on Applied Mathematics, 16(4):732–746, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Howard:1968:AO

[315] James E. Howard. Asymmetric oscillations. SIAM Journal on Applied Mathematics, 16(4):747–755, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wong:1968:TDR

[316] E. Wong. Two-dimensional random fields and representation of images. SIAM Journal on Applied Mathematics, 16(4):756–770, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Colton:1968:JPN

[317] David Colton. Jacobi polynomials of negative index and a nonexistence theorem for the generalized axially symmetric potential equation. SIAM Journal on Applied Mathematics, 16 (4):771–776, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yeh:1968:FCS

[318] Raymond T. Yeh. Finite covers with substitution property on infinite automata. SIAM Journal on Applied Mathematics, 16(4):777–782, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ahluwalia:1968:UAT

[319] D. S. Ahluwalia, R. M. Lewis, and J. Boersma. Uniform asymptotic theory of diffraction by a plane screen. SIAM Journal on Applied Mathematics, 16(4):783–807, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bazer:1968:SMF

[320] J. Bazer and David H. Y. Yen. Steady magnetogasdynamic flow past a point source. SIAM Journal on Applied Mathematics, 16(4):808–833, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neuringer:1968:GFI

[321] Joseph L. Neuringer. Green’s function for an instantaneous line particle source diffusion in a gravitational field and under the influence of a linear shear wind. SIAM Journal on Applied Mathematics, 16(4):834–842, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bland:1968:BAP

[322] R. P. Bland and T. L. Bratcher. A Bayesian approach to the problem of ranking binomial probabilities. SIAM Journal on Applied Mathematics, 16(4):843–850, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zettl:1968:ASA

[323] Anton Zettl. Adjoint and self-adjoint boundary value problems with interface conditions. SIAM Journal on Applied Mathematics, 16(4):851–859, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Swan:1968:SIC

[324] G. W. Swan. The semi-infinite cylinder with prescribed end displacements. SIAM Journal on Applied Mathematics, 16(4):860–881, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zucker:1968:ISL

[325] Robert D. Zucker. Interpretation of the singularities in Lamé products of the second kind. SIAM Journal on Applied Mathematics, 16(4):882–887, July 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brenner:1968:NRL

[326] J. L. Brenner. New root-location theorems for partitioned matrices. SIAM Journal on Applied Mathematics, 16 (5):889–898, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shapiro:1968:ME

[327] Norman Z. Shapiro and Lloyd S. Shapley. On membrane equilibria. SIAM Journal on Applied Mathematics, 16 (5):899–923, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Saeks:1968:SGL

[328] R. Saeks. Synthesis of general linear networks. SIAM Journal on Applied Mathematics, 16(5):924–930, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Milne:1968:OMP

[329] R. D. Milne. An oblique matrix pseudoinverse. SIAM Journal on Applied Mathematics, 16(5):931–944, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Koh:1968:CHT

[330] E. L. Koh and A. H. Zemanian. The complex Hankel and I-transformations of generalized functions. SIAM Journal on Applied Mathematics, 16(5): 945–957, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shehadeh:1968:DCS

[331] Nazmi M. Shehadeh. On the distribution of the coefficients of some polynomials. SIAM Journal on Applied Mathematics, 16(5):958–963, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tadjbakhsh:1968:ODP

[332] I. Tadjbakhsh. An optimum design problem for the nonlinear elastica. SIAM Journal on Applied Mathematics, 16(5):964–972, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ray-Chaudhuri:1968:CIR

[333] D. K. Ray-Chaudhuri. Combinatorial information retrieval systems for files. SIAM Journal on Applied Mathematics, 16(5):973–992, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Folkman:1968:AOC

[334] Jon Folkman and Norman Shapiro. Approximating one convex function by another. SIAM Journal on Applied Mathematics, 16(5):993–997, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Marlin:1968:PSS

[335] J. A. Marlin and D. F. Ullrich. Periodic solutions of second order nonlinear differential equations without damping. SIAM Journal on Applied Mathematics, 16(5):998–1010, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gudder:1968:ELT

[336] Stanley P. Gudder. Elementary length topologies in physics. SIAM Journal on Applied Mathematics, 16(5): 1011–1019, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jameson:1968:SEI

[337] Antony Jameson. Solution of the equation AX + XB = C by inversion of an M × M or N × N matrix. SIAM Journal on Applied Mathematics, 16(5): 1020–1023, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1968:ASS

[338] J. A. Morrison. An averaging scheme for some nonlinear resonance problems. SIAM Journal on Applied Mathematics, 16(5):1024–1047, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Johnson:1968:EPS

[339] Harry L. Johnson. The existence of a periodic solution of a vibrating hanging string. SIAM Journal on Applied Mathematics, 16(5):1048–1058, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hanson:1968:SSO

[340] Richard J. Hanson. Simplification of second order systems of ordinary differential equations with a turning point. SIAM Journal on Applied Mathematics, 16(5):1059–1080, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hurd:1968:FWP

[341] A. E. Hurd and A. Robinson. On flexural wave propagation in nonhomogeneous elastic plates. SIAM Journal on Applied Mathematics, 16(5): 1081–1089, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Luenberger:1968:QCP

[342] David G. Luenberger. Quasi-convex programming. SIAM Journal on Applied Mathematics, 16(5):1090–1095, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Beltrami:1968:MCS

[343] E. J. Beltrami. On the meromorphic character of scattering matrices for lossless systems. SIAM Journal on Applied Mathematics, 16(5): 1096–1100, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Srivastav:1968:MBV

[344] R. P. Srivastav and Prem Narain. A mixed boundary value problem of elastostatics for an isotropic elastic cylinder containing a strip crack opened by internal pressure. SIAM Journal on Applied Mathematics, 16(5):

1101–1118, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Diliberto:1968:SFP

[345] Stephen P. Diliberto. Scale factors for periodic surfaces. SIAM Journal on Applied Mathematics, 16(5): 1119–1122, September 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gray:1968:LTR

[346] A. H. Gray, Jr. Level touchings in a restricted random walk. SIAM Journal on Applied Mathematics, 16(6): 1123–1129, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosenstock:1968:LTR

[347] Herbert B. Rosenstock. Level touchings in a random walk. SIAM Journal on Applied Mathematics, 16(6): 1130–1131, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kotorynski:1968:ISL

[348] W. P. Kotorynski. On an irregular Sturm–Liouville problem in boundary layer theory. SIAM Journal on Applied Mathematics, 16(6):1132–1140, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Geller:1968:ADL

[349] Dennis P. Geller and Frank Harary. Arrow diagrams are line digraphs. SIAM Journal on Applied Mathematics, 16(6):1141–1145, November 1968.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yarlagadda:1968:ATM

[350] R. Yarlagadda. An application of tridiagonal matrices to network synthesis. SIAM Journal on Applied Mathematics, 16(6):1146–1162, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ditzian:1968:ITG

[351] Z. Ditzian. Inversion of transforms of generalized functions. SIAM Journal on Applied Mathematics, 16(6): 1163–1172, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sklansky:1968:DMA

[352] J. Sklansky and K. R. Kaplan. The delay method of analyzing Markov chains. SIAM Journal on Applied Mathematics, 16(6):1173–1183, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Blumenthal:1968:LSS

[353] Saul Blumenthal. Logarithms of sample spacings. SIAM Journal on Applied Mathematics, 16(6):1184–1191, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Radke:1968:CMD

[354] C. E. Radke. Classes of matrices with distinct, real characteristic values. SIAM Journal on Applied Mathematics, 16(6):1192–1207, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vandergraft:1968:SPM

[355] James S. Vandergraft. Spectral properties of matrices which have invariant cones. SIAM Journal on Applied Mathematics, 16(6):1208–1222, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Farlow:1968:ETP

[356] S. J. Farlow. An existence theorem for periodic solutions of a parabolic boundary value problem of the second kind. SIAM Journal on Applied Mathematics, 16(6):1223–1226, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kotin:1968:PPS

[357] Leon Kotin. On positive and periodic solutions of Riccati equations. SIAM Journal on Applied Mathematics, 16(6):1227–1231, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shapiro:1968:SRM

[358] Jeremy F. Shapiro. Shortest route methods for finite state space deterministic dynamic programming problems. SIAM Journal on Applied Mathematics, 16(6):1232–1250, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pereira:1968:BPP

[359] J. M. S. Simoes Pereira. Boolean permanents, permutation graphs and products. SIAM Journal on Applied Mathematics, 16(6):1251–1254, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hunt:1968:RTH

[360] G. E. Hunt. Radiative transfer in a homogeneous cylindrical atmosphere. SIAM Journal on Applied Mathematics, 16(6):1255–1265, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Magrab:1968:TRV

[361] Edward B. Magrab and Donald J. Ochsner. Transient response of a viscoelastic bar subjected to a time dependent end-displacement. SIAM Journal on Applied Mathematics, 16(6): 1266–1279, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1968:SEC

[362] Norman Morrison. Smoothing and extrapolation of continuous time series using Laguerre polynomials. SIAM Journal on Applied Mathematics, 16 (6):1280–1304, November 1968. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dafermos:1968:SOP

[363] Constantine M. Dafermos. Stability of orientation patterns of liquid crystals subject to magnetic fields. SIAM Journal on Applied Mathematics, 16(6): 1305–1318, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Weiss:1968:CTF

[364] Leonard Weiss. Converse theorems for finite time stability. SIAM Journal on Applied Mathematics, 16(6): 1319–1324, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lum:1968:MDB

[365] Vincent Lum and R. T. Chien. On the minimum distance of bose-chaudhuri-hocquenghem codes. SIAM Journal on Applied Mathematics, 16(6): 1325–1337, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Erdelyi:1968:SDI

[366] A. Erdélyi. Some dual integral equations. SIAM Journal on Applied Mathematics, 16(6):1338–1340, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Freedman:1968:EER

[367] H. I. Freedman. Estimates on the existence region for periodic solutions of equations involving a small parameter. I: The noncritical case. SIAM Journal on Applied Mathematics, 16(6): 1341–1349, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Root:1968:CCG

[368] W. L. Root and P. P. Varaiya. Capacity of classes of Gaussian channels. SIAM Journal on Applied Mathematics, 16(6):1350–1393, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Walkup:1968:MCA

[369] D. W. Walkup. Minimal conditions for association of binary variables. SIAM Journal on Applied Mathematics, 16(6):1394–1403, November 1968. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morris:1969:STS

[370] Robert Morris. Some theorems on sorting. SIAM Journal on Applied Mathematics, 17(1):1–6, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boullion:1969:NSO

[371] T. L. Boullion and P. L. Odell. A note on the scroggs-odell pseudoinverse. SIAM Journal on Applied Mathematics, 17(1):7–10, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:1969:CFT

[372] Chia C. Wang. Comma-free ternary triorthogonal code. SIAM Journal on Applied Mathematics, 17(1):11–32, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stewart:1969:CGI

[373] G. W. Stewart. On the continuity of the generalized inverse. SIAM Journal on Applied Mathematics, 17 (1):33–45, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rothfarb:1969:CFP

[374] B. Rothfarb and I. T. Frisch. On the 3-commodity flow problem. SIAM Journal on Applied Mathematics, 17 (1):46–58, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Williams:1969:NDP

[375] F. A. Williams. A nonlinear diffusion problem relevant to desalination by reverse osmosis. SIAM Journal on Applied Mathematics, 17(1):59–73, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Heyman:1969:PQS

[376] Daniel P. Heyman. A priority queueing system with server interference. SIAM Journal on Applied Mathematics, 17 (1):74–82, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shapiro:1969:ICM

[377] N. Z. Shapiro. An inequality concerning the minimum of the sum of an arbitrary function and a linear function. SIAM Journal on Applied Mathematics, 17(1):83–97, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Walkup:1969:SPR

[378] David W. Walkup and Roger J.-B. Wets. Stochastic programs with recourse II: On the continuity of the objective. SIAM Journal on Applied Mathematics, 17(1):98–103, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Legatos:1969:CFS

[379] G. G. Legatos and V. A. Staikos. On certain families of solutions of third order ordinary differential equations. SIAM Journal on Applied Mathematics, 17(1):104–108, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grimmer:1969:AAP

[380] R. C. Grimmer. Asymptotically almost periodic solutions of differential equations. SIAM Journal on Applied Mathematics, 17(1):109–115, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carlson:1969:CBE

[381] B. C. Carlson. A connection between elementary functions and higher transcendental functions. SIAM Journal on Applied Mathematics, 17(1):116– 148, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Davis:1969:IMS

[382] J. A. Davis and L. A. Hageman. An iterative method for solving the neutron transport equation in x − y geometry. SIAM Journal on Applied Mathematics, 17(1):149–171, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dickey:1969:NSU

[383] R. W. Dickey. The nonlinear string under a vertical force. SIAM Journal on Applied Mathematics, 17(1): 172–178, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Duffin:1969:IPR

[384] R. J. Duffin. The influence of Poisson’s ratio on the vibrational spectrum. SIAM Journal on Applied Mathematics, 17(1):179–191, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hardy:1969:RMS

[385] William C. Hardy. A recursion method of solution for occupancy problems. SIAM Journal on Applied Mathematics, 17(1):192–197, January 1969. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Noonburg:1969:RTE

[386] V. W. Noonburg. Roots of a transcendental equation associated with a system of differential-difference equations. SIAM Journal on Applied Mathematics, 17(1):198–205, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pollak:1969:AEL

[387] H. O. Pollak and Lowell Schoenfeld. The asymptotic estimation of a Laplace transform-arising in the theory of magnetoresistance. SIAM Journal on Applied Mathematics, 17(1):206–211, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rainey:1969:AFS

[388] J. W. Rainey and G. J. Habetler. An application of the LR factorization to sequential tridiagonalization methods. SIAM Journal on Applied Mathematics, 17(1):212–221, January 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keast:1969:SWE

[389] P. Keast. On the solutions of the wave equation in one space dimension under derivative boundary conditions. SIAM Journal on Applied Mathematics, 17 (2):223–230, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosencrans:1969:SC

[390] Steven Rosencrans. On Schwarzschild’s criterion. SIAM Journal on Applied Mathematics, 17(2):231–239, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meissner:1969:HOR

[391] Loren P. Meissner. High order resonance for Duffing’s differential equation. SIAM Journal on Applied Mathematics, 17(2):240–250, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carlitz:1969:SCR

[392] L. Carlitz. Solution of certain recurrences. SIAM Journal on Applied Mathematics, 17(2):251–259, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shen:1969:ATU

[393] M. C. Shen. Asymptotic theory of unsteady three-dimensional waves in a channel of arbitrary cross section. SIAM Journal on Applied Mathematics, 17(2):260–271, March 1969. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stephan:1969:ABF

[394] Bruce H. Stephan. Asymptotic behavior of a functional differential equation with bounded lag. SIAM Journal on Applied Mathematics, 17(2):272–279, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:1969:MPC

[395] P. K. C. Wang. A maximum property of Cauchy’s problem for the multigroup neutron transport equation. SIAM Journal on Applied Mathematics, 17 (2):280–286, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McShane:1969:SIS

[396] E. J. McShane. Stochastic integrals and stochastic functional equations. SIAM Journal on Applied Mathematics, 17 (2):287–306, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gould:1969:PSI

[397] H. W. Gould. Power sum identities for arbitrary symmetric arrays. SIAM Journal on Applied Mathematics, 17 (2):307–316, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peterson:1969:GPD

[398] Elmor L. Peterson and Joseph G. Ecker. Geometric programming: Duality in quadratic programming and l_{p}- approximation II (canonical programs). SIAM Journal on Applied Mathematics, 17(2):317–340, March 1969. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Drobnies:1969:NMB

[399] Saul I. Drobnies. A note on a method of Bellman and Richardson in perturbation theory. SIAM Journal on Applied Mathematics, 17(2):341–342, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Taylor:1969:SRD

[400] James H. Taylor and Kumpati S. Narendra. Stability regions for the damped Mathieu equation. SIAM Journal on Applied Mathematics, 17 (2):343–352, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fox:1969:ND

[401] Martin Fox and George S. Kimeldorf. Noisy duels. SIAM Journal on Applied Mathematics, 17(2):353–361, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jagerman:1969:EAB

[402] D. Jagerman. ϵ-entropy and approximation of bandlimited functions. SIAM Journal on Applied Mathematics, 17 (2):362–377, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

DiPaola:1969:MBC

[403] Jane W. Di Paola. On minimum blocking coalitions in small projective plane games. SIAM Journal on Applied Mathematics, 17(2):378–392, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Silverman:1969:ERL

[404] L. M. Silverman and H. E. Meadows. Equivalent realizations of linear systems. SIAM Journal on Applied Mathematics, 17(2):393–408, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Booth:1969:LZD

[405] R. S. Booth. Location of zeros of derivatives. II. SIAM Journal on Applied Mathematics, 17(2):409–415, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Graham:1969:BMT

[406] R. L. Graham. Bounds on multiprocessing timing anomalies. SIAM Journal on Applied Mathematics, 17 (2):416–429, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wagner:1969:BVP

[407] N. R. Wagner. On a boundary value problem arising in stochastic diffusion problems. SIAM Journal on Applied Mathematics, 17(2):430–433, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Albert:1969:CPN

[408] Arthur Albert. Conditions for positive and nonnegative definiteness in terms of pseudoinverses. SIAM Journal on Applied Mathematics, 17(2):434–440, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ranger:1969:ASM

[409] K. B. Ranger. On axially symmetric magnetohydrostatic fields. SIAM Journal on Applied Mathematics, 17 (2):441–446, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schwartz:1969:EGC

[410] S. C. Schwartz. On the estimation of a Gaussian convolution probability density. SIAM Journal on Applied Mathematics, 17(2):447–453, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tutte:1969:EFC

[411] W. T. Tutte. On the enumeration of four-colored maps. SIAM Journal on Applied Mathematics, 17(2):454–460, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Aronson:1969:RPF

[412] D. G. Aronson. Regularity properties of flows through porous media. SIAM Journal on Applied Mathematics, 17 (2):461–467, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gutterman:1969:OMP

[413] M. Gutterman. An operational method in partial differential equations. SIAM Journal on Applied Mathematics, 17 (2):468–493, March 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rubenfeld:1969:BEM

[414] Lester A. Rubenfeld and Joseph B. Keller. Bounds on elastic moduli of composite media. SIAM Journal on Applied Mathematics, 17(3):495–510, May 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dasarathy:1969:SCS

[415] B. V. Dasarathy and P. Srinivasan. On the synthesis of a class of second order nonlinear differential equations. SIAM Journal on Applied Mathematics, 17(3):511–515, May 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Macki:1969:ETN

[416] Jack W. Macki. An example in the theory of nonlinear oscillations. SIAM Journal on Applied Mathematics, 17(3):516–519, May 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Newman:1969:CGI

[417] T. G. Newman and P. L. Odell. On the concept of A p-q generalized inverse of a matrix. SIAM Journal on Applied Mathematics, 17(3):520–525, May 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reiss:1969:IPK

[418] Edward L. Reiss. The impact problem for the Klein–Gordon equation. SIAM Journal on Applied Mathematics, 17(3):526–542, May 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lippman:1969:OIP

[419] Steven A. Lippman. Optimal inventory policy with subadditive ordering costs and stochastic demands. SIAM Journal on Applied Mathematics, 17(3):543–559, May 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Soong:1969:EMM

[420] T. T. Soong. An extension of the moment method in statistical estimation. SIAM Journal on Applied Mathematics, 17(3):560–568, May 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

OMalley:1969:BVP

[421] R. E. O’Malley, Jr. On a boundary value problem for a nonlinear differential equation with a small parameter. SIAM Journal on Applied Mathematics, 17(3):569–581, May 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Falk:1969:MSN

[422] James E. Falk. Maximization of signal-to-noise ratio in an optical filter. SIAM Journal on Applied Mathematics, 17(3):582–592, May 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ryser:1969:CC

[423] H. J. Ryser. Combinatorial configurations. SIAM Journal on Applied Mathematics, 17(3):593–602, May 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neudecker:1969:NKM

[424] H. Neudecker. A note on Kronecker matrix products and matrix equation systems. SIAM Journal on Applied Mathematics, 17(3):603–606, May 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pittnauer:1969:HFP

[425] Franz Pittnauer. Holomorphic functions with prescribed asymptotic expansions. SIAM Journal on Applied Mathematics, 17(3):607–613, May 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wong:1969:SRH

[426] James S. W. Wong. Some remarks on Hermitian and anti-Hermitian properties of Green’s matrices. SIAM Journal on Applied Mathematics, 17(4):615– 623, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pierre:1969:ILF

[427] Percy A. Pierre. On the independence of linear functionals of linear processes. SIAM Journal on Applied Mathematics, 17(4):624–637, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

VanSlyke:1969:SLP

[428] R. M. Van Slyke and Roger Wets. Lshaped linear programs with applications to optimal control and stochastic programming. SIAM Journal on Applied Mathematics, 17(4):638–663, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hoffman:1969:CLE

[429] Alan J. Hoffman. The change in the least eigenvalue of the adjacency matrix of a graph under imbedding. SIAM Journal on Applied Mathematics, 17(4):664–671, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

DeFigueiredo:1969:BSC

[430] Rui J. P. De Figueiredo and Chieng-Yi Chang. On the boundedness of solutions of classes of multidimensional nonlinear autonomous systems. SIAM Journal on Applied Mathematics, 17(4):672–680, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barnea:1969:MNR

[431] Daniel I. Barnea. A method and new results for stability and instability of autonomous functional differential equations. SIAM Journal on Applied Mathematics, 17(4):681–697, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Perko:1969:HOA

[432] Lawrence M. Perko. Higher order averaging and related methods for perturbed periodic and quasi-periodic systems. SIAM Journal on Applied Mathematics, 17(4):698–724, July 1969. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Strauss:1969:PUS

[433] Aaron Strauss and James A. Yorke. Perturbing uniformly stable linear systems with and without attraction. SIAM Journal on Applied Mathematics, 17(4):725–738, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goldfarb:1969:EDV

[434] Donald Goldfarb. Extension of Davidon’s variable metric method to maximization under linear inequality and equality constraints. SIAM Journal on Applied Mathematics, 17(4):739–764, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gallagher:1969:PPS

[435] A. P. Gallagher. On a proof by Petrov of the stability of plane Couette flow and plane Poiseuille flow. SIAM Journal on Applied Mathematics, 17(4):765–768, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tucker:1969:BVP

[436] D. H. Tucker. Boundary value problems for linear differential systems. SIAM Journal on Applied Mathematics, 17(4):769–783, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fujii:1969:OST

[437] M. Fujii, T. Kasami, and K. Ninomiya. Optimal sequencing of two equivalent processors. SIAM Journal on Applied Mathematics, 17(4):784–789, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [667].

Wood:1969:GSO

[438] B. Wood. Generalized szasz operators for the approximation in the complex domain. SIAM Journal on Applied Mathematics, 17(4):790–801, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jensen:1969:LPN

[439] D. R. Jensen. Limit properties of noncentral multivariate Rayleigh and chisquare distributions. SIAM Journal on Applied Mathematics, 17(4):802–814, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lick:1969:TVE

[440] Wilbert Lick. Two-variable expansions and singular perturbation problems. SIAM Journal on Applied Mathematics, 17(4):815–825, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baumert:1969:DS

[441] L. D. Baumert. Difference sets. SIAM Journal on Applied Mathematics, 17 (4):826–833, July 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pandey:1969:CIG

[442] J. N. Pandey. Complex inversion for the generalized Hankel convolution transformation. SIAM Journal on Applied Mathematics, 17(5):835–848, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kotorynski:1969:ASN

[443] W. P. Kotorynski. Asymptotic solutions of the nonlinear Boussinesq equations in a paraboloidal boundary-layer. SIAM Journal on Applied Mathematics, 17(5):849–871, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cooper:1969:SCS

[444] David B. Cooper and Ralph J. Schwarz. On suitable conditions for statistical pattern recognition without supervision. SIAM Journal on Applied Mathematics, 17(5):872–896, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shein:1969:VWT

[445] N. P. Shein and I. T. Frisch. Vertex weighted trees with fewest relay vertices. SIAM Journal on Applied Mathematics, 17(5):897–903, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Polfeldt:1969:IRV

[446] Thomas Polfeldt. Integrating regularly varying functions with exponent -1. SIAM Journal on Applied Mathematics, 17(5):904–908, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kelman:1969:FPC

[447] Robert B. Kelman. Flow past a cave: A problem in potential theory. SIAM Journal on Applied Mathematics, 17 (5):909–920, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neuts:1969:EEQ

[448] Marcel F. Neuts and Jozef L. Teugels. Exponential ergodicity of the M/G/1 queue. SIAM Journal on Applied Mathematics, 17(5):921–929, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Warmbrod:1969:DFF

[449] G. K. Warmbrod. The distributional finite Fourier transform. SIAM Journal on Applied Mathematics, 17(5):

930–956, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sanchez:1969:NPS

[450] David A. Sanchez. A note on periodic solutions of Riccati-type equations. SIAM Journal on Applied Mathematics, 17(5):957–959, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shapiro:1969:GTE

[451] N. Z. Shapiro. A generalized technique for eliminating species in complex chemical equilibrium calculations. SIAM Journal on Applied Mathematics, 17(5):960–971, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baker:1969:CSR

[452] Charles R. Baker. Complete simultaneous reduction of covariance operators. SIAM Journal on Applied Mathematics, 17(5):972–983, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bradley:1969:CEC

[453] John S. Bradley. Conditions for the existence of conjugate points for a fourth order linear differential equation. SIAM Journal on Applied Mathematics, 17(5):984–991, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kotorynski:1969:SLP

[454] W. P. Kotorynski. A Sturm–Liouville problem associated with the Falkner– Skan equation. SIAM Journal on Applied Mathematics, 17(5):992–995, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fiacco:1969:GPM

[455] A. V. Fiacco and A. P. Jones. Generalized penalty methods in topological spaces. SIAM Journal on Applied Mathematics, 17(5):996–1000, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Beals:1969:MSU

[456] Richard Beals and Tjalling C. Koopmans. Maximizing stationary utility in a constant technology. SIAM Journal on Applied Mathematics, 17(5): 1001–1015, September 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Graham:1969:SRM

[457] R. L. Graham and L. H. Harper. Some results on matching in bipartite graphs. SIAM Journal on Applied Mathematics, 17(6):1017–1022, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Marcus:1969:IMF

[458] Marvin Marcus. Inequalities for matrix functions of combinatorial interest. SIAM Journal on Applied Mathematics, 17(6):1023–1031, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chandrasekaran:1969:TUM

[459] R. Chandrasekaran. Total unimodularity of matrices. SIAM Journal on Applied Mathematics, 17(6):1032–1034, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Akin:1969:RAI

[460] J. E. Akin and J. Counts. On rational approximation to the inverse Laplace transform. SIAM Journal on Applied Mathematics, 17(6):1035–1040, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Erdogan:1969:ASS

[461] F. Erdogan. Approximate solutions of systems of singular integral equations. SIAM Journal on Applied Mathematics, 17(6):1041–1059, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bebernes:1969:TST

[462] J. W. Bebernes and Russell Wilhelmsen. A technique for solving twodimensional boundary value problems. SIAM Journal on Applied Mathematics, 17(6):1060–1064, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pomraning:1969:AME

[463] G. C. Pomraning. An application of the Mathieu equation to a radiative transfer problem. SIAM Journal on Applied Mathematics, 17(6): 1065–1069, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pierre:1969:SFR

[464] Percy A. Pierre. The sample function regularity of linear random processes.

SIAM Journal on Applied Mathematics, 17(6):1070–1077, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burton:1969:CLF

[465] T. A. Burton. On the construction of Lyapunov functions. SIAM Journal on Applied Mathematics, 17(6): 1078–1085, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dillon:1969:SNP

[466] J. F. Dillon and D. P. Roselle. Simon Newcomb’s problem. SIAM Journal on Applied Mathematics, 17(6): 1086–1093, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hall:1969:AHM

[467] Marshall Hall, Jr. Automorphisms of Hadamard matrices. SIAM Journal on Applied Mathematics, 17(6): 1094–1101, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Slepian:1969:FIE

[468] D. Slepian and T. T. Kadota. Four integral equations of detection theory. SIAM Journal on Applied Mathematics, 17(6):1102–1117, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ben-Israel:1969:MIZ

[469] Adi Ben-Israel. On matrices of index zero or one. SIAM Journal on Applied Mathematics, 17(6):1118–1121, November 1969. CODEN SMJMAP.

ISSN 0036-1399 (print), 1095-712X (electronic).

Schuur:1969:SPT

[470] J. D. Schuur. Saddle-point type behavior for nonautonomous second order linear systems of ordinary differential equations. SIAM Journal on Applied Mathematics, 17(6):1122–1129, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ekblom:1969:CEL

[471] H. Ekblom and S. Henriksson. L_{p}- criteria for the estimation of location parameters. SIAM Journal on Applied Mathematics, 17(6):1130–1141, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grunbaum:1969:SRU

[472] Branko Grunbaum. Some results on the upper bound conjecture for convex polytopes. SIAM Journal on Applied Mathematics, 17(6):1142–1149, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mann:1969:CRE

[473] Nancy R. Mann. Cramer–Rao efficiencies of best linear invariant estimators of parameters of the extreme-value distribution under type II censoring from above. SIAM Journal on Applied Mathematics, 17(6):1150–1162, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schmeidler:1969:NCF

[474] David Schmeidler. The nucleolus of a characteristic function game. SIAM Journal on Applied Mathematics, 17 (6):1163–1170, November 1969. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schmitt:1969:SND

[475] Klaus Schmitt. On solutions of nonlinear differential equations with deviating arguments. SIAM Journal on Applied Mathematics, 17(6):1171–1176, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tate:1969:ARD

[476] C. Tate. Aggregates in a random distribution of points in a cylinder. SIAM Journal on Applied Mathematics, 17(6):1177–1189, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zyskind:1969:BLE

[477] George Zyskind and Frank B. Martin. On best linear estimation and a general Gauss–Markov theorem in linear models with arbitrary nonnegative covariance structure. SIAM Journal on Applied Mathematics, 17(6): 1190–1202, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bose:1969:DCI

[478] R. C. Bose and Gary G. Koch. The design of combinatorial information retrieval systems for files with multiplevalued attributes. SIAM Journal on Applied Mathematics, 17(6):1203–1214, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Martos:1969:SMQ

[479] Bela Martos. Subdefinite matrices and quadratic forms. SIAM Journal on Applied Mathematics, 17(6): 1215–1223, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Katz:1969:CNS

[480] I. Norman Katz. On the convergence of a numerical scheme for solving some locational equilibrium problems. SIAM Journal on Applied Mathematics, 17(6):1224–1231, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Saigal:1969:PHC

[481] Romesh Saigal. A proof of the Hirsch conjecture on the polyhedron of the shortest route problem. SIAM Journal on Applied Mathematics, 17(6): 1232–1238, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fiacco:1969:GRS

[482] Anthony V. Fiacco. A general regularized sequential unconstrained minimization technique. SIAM Journal on Applied Mathematics, 17(6):1239–1245, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Denman:1969:ICO

[483] Harry H. Denman and William J. Larkin, III. Invariance conditions on ordinary differential equations defining smoothing functions. SIAM Journal on Applied Mathematics, 17(6): 1246–1257, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fink:1969:CFA

[484] A. M. Fink. Compact families of almost periodic functions and an application of the Schauder fixed-point theorem. SIAM Journal on Applied Mathematics, 17(6):1258–1262, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Luenberger:1969:HPM

[485] David G. Luenberger. Hyperbolic pairs in the method of conjugate gradients. SIAM Journal on Applied Mathematics, 17(6):1263–1267, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zettl:1969:ANB

[486] Anton Zettl. Adjointness in nonadjoint boundary value problems. SIAM Journal on Applied Mathematics, 17(6): 1268–1279, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gould:1969:ELM

[487] F. J. Gould. Extensions of Lagrange multipliers in nonlinear programming. SIAM Journal on Applied Mathematics, 17(6):1280–1297, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

vanderLinde:1969:ERE

[488] Reinhoud H. van der Linde. Eigenfunctions of random eigenvalue problems and their statistical properties. SIAM Journal on Applied Mathematics, 17(6):1298–1304, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tam:1969:NAS

[489] K. Kuen Tam. A note on the asymptotic solution of the flow between two oppositely rotating infinite plane disks. SIAM Journal on Applied Mathematics, 17(6):1305–1310, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stanton:1969:IIC

[490] R. G. Stanton and J. G. Kalbfleisch. Intersection inequalities for the covering problem. SIAM Journal on Applied Mathematics, 17(6):1311–1316, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stokes:1969:SIO

[491] Arnold P. Stokes. Some implications of orbital stability in Banach spaces. SIAM Journal on Applied Mathematics, 17(6):1317–1325, November 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Parikh:1970:ESL

[492] Shailendra C. Parikh. Equivalent stochastic linear programs. SIAM Journal on Applied Mathematics, 18 (1):1–5, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ratz:1970:IGI

[493] Jurg Ratz. On isometries of generalized inner product spaces. SIAM Journal on Applied Mathematics, 18(1):6–9, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carr:1970:IEF

[494] R. E. Carr. Integral equation formulation for certain interface boundary value problems. SIAM Journal on Applied Mathematics, 18(1):10–18, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Folkman:1970:GMC

[495] Jon Folkman. Graphs with monochromatic complete subgraphs in every edge coloring. SIAM Journal on Applied Mathematics, 18(1):19–24, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kleitman:1970:MNM

[496] Daniel J. Kleitman. Minimal number of multiple edges in realization of an incidence sequence without loops. SIAM Journal on Applied Mathematics, 18 (1):25–28, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kassoy:1970:LBL

[497] David R. Kassoy. On laminar boundary layer blowoff. SIAM Journal on Applied Mathematics, 18(1):29–40, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Parente:1970:NDE

[498] R. B. Parente. Nonlinear differential equations and analytic system theory. SIAM Journal on Applied Mathematics, 18(1):41–66, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wong:1970:DIB

[499] E. Wong. The distribution of intervals between zeros for a stationary Gaussian process. SIAM Journal on Applied Mathematics, 18(1):67–73, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bobisud:1970:OND

[500] L. E. Bobisud. Oscillation of nonlinear differential equations with small nonlinear damping. SIAM Journal on Applied Mathematics, 18(1):74–76, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yee:1970:UTE

[501] Kane S. Yee. Uniqueness theorems for an exterior electromagnetic field. SIAM Journal on Applied Mathematics, 18 (1):77–83, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grinold:1970:SDC

[502] Richard C. Grinold. Symmetric duality for continuous linear programs. SIAM Journal on Applied Mathematics, 18 (1):84–97, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1970:HP

[503] A. H. Zemanian. The Hilbert port. SIAM Journal on Applied Mathematics, 18(1):98–138, January 1970. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Milnes:1970:NMA

[504] Harold Willis Milnes and S. K. Hildebrand. Nonisolated minimizing arcs. SIAM Journal on Applied Mathematics, 18(1):139–149, January 1970. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cockayne:1970:EAS

[505] E. J. Cockayne. On the efficiency of the algorithm for Steiner minimal trees. SIAM Journal on Applied Mathematics, 18(1):150–159, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sotoodehnia:1970:TDT

[506] A. Sotoodehnia and R. C. Erdmann. Two-dimensional transport in a slab. SIAM Journal on Applied Mathematics, 18(1):160–171, January 1970. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Spanier:1970:AAV

[507] Jerome Spanier. An analytic approach to variance reduction. SIAM Journal on Applied Mathematics, 18 (1):172–190, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Doner:1970:MCS

[508] J. R. Doner and V. R. R. Uppuluri. A Markov chain structure for riff shuffling. SIAM Journal on Applied Mathematics, 18(1):191–209, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Caughey:1970:LAW

[509] T. K. Caughey. Large amplitude whirling of an elastic string — A nonlinear eigenvalue problem. SIAM Journal on Applied Mathematics, 18 (1):210–237, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Owens:1970:DMN

[510] Alvin B. Owens. On determining the minimum number of multiple edges for an incidence sequence. SIAM Journal on Applied Mathematics, 18 (1):238–240, January 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Messiter:1970:BLF

[511] A. F. Messiter. Boundary-layer flow near the trailing edge of a flat plate. SIAM Journal on Applied Mathematics, 18(1):241–257, January 1970. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reid:1970:SEP

[512] William T. Reid. Some elementary properties of proper values and proper vectors of matrix functions. SIAM Journal on Applied Mathematics, 18 (2):259–266, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Taylor:1970:CPA

[513] James H. Taylor and Kumpati S. Narendra. The Corduneanu–Popov approach to the stability of nonlinear time-varying systems. SIAM Journal on Applied Mathematics, 18(2):267– 281, March 1970. CODEN SMJMAP.

ISSN 0036-1399 (print), 1095-712X (electronic).

Bajaj:1970:SPP

[514] Prem N. Bajaj. Singular points in products of semidynamical systems. SIAM Journal on Applied Mathematics, 18 (2):282–286, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ahluwalia:1970:UAT

[515] Daljit S. Ahluwalia. Uniform asymptotic theory of diffraction by the edge of a three-dimensional body. SIAM Journal on Applied Mathematics, 18 (2):287–301, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Golomb:1970:PCL

[516] Solomon W. Golomb and Lloyd R. Welch. Perfect codes in the Lee metric and the packing of polyominoes. SIAM Journal on Applied Mathematics, 18 (2):302–317, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burnside:1970:NLM

[517] R. R. Burnside. A note on Lighthill’s method of strained coordinates. SIAM Journal on Applied Mathematics, 18 (2):318–321, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shanno:1970:AGP

[518] D. F. Shanno. An accelerated gradient projection method for linearly constrained nonlinear estimation. SIAM Journal on Applied Mathematics, 18 (2):322–334, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1970:AMN

[519] Y. M. Chen and P. L. Christiansen. Application of a modified Newton’s iteration method to construct solutions of eigenvalue problems of nonlinear partial differential operators. SIAM Journal on Applied Mathematics, 18 (2):335–345, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miller:1970:SNA

[520] Keith Miller. Stabilized numerical analytic prolongation with poles. SIAM Journal on Applied Mathematics, 18 (2):346–363, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ulrich:1970:CPO

[521] John Wade Ulrich. A characterization of planar oriented graphs. SIAM Journal on Applied Mathematics, 18 (2):364–371, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Noble:1970:SDT

[522] B. Noble and J. R. Whiteman. Solution of dual trigonometrical series using orthogonality relations. SIAM Journal on Applied Mathematics, 18 (2):372–379, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cantor:1970:APS

[523] David G. Cantor and John W. Evans. On approximation by positive sums of powers. SIAM Journal on Applied Mathematics, 18(2):380–388, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Laetsch:1970:EBM

[524] Theodore Laetsch. Existence and bounds for multiple solutions of nonlinear equations. SIAM Journal on Applied Mathematics, 18(2):389–400, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1970:GIT

[525] Carl D. Meyer, Jr. Generalized inverses of triangular matrices. SIAM Journal on Applied Mathematics, 18 (2):401–406, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Asner:1970:TNH

[526] Bernard A. Asner, Jr. On the total nonnegativity of the Hurwitz matrix. SIAM Journal on Applied Mathematics, 18(2):407–414, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bhat:1970:SQS

[527] U. Narayan Bhat. Some queueing systems with mixed disciplines. SIAM Journal on Applied Mathematics, 18 (2):415–432, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1970:IIE

[528] G. H. Meyer. The invariant imbedding equations for multipoint boundary value problems. SIAM Journal on Applied Mathematics, 18(2):433–453, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schwartz:1970:EUP

[529] Ralph E. Schwartz. Existence and uniqueness properties of subfunctions of Boolean functions. SIAM Journal on Applied Mathematics, 18(2):454–461, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mensch:1970:GAT

[530] Gerhard Mensch. Generalization of Akers’ two-dimensional job shop scheduling model to J dimensions. SIAM Journal on Applied Mathematics, 18 (2):462–466, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1970:SFH

[531] A. H. Zemanian. A scattering formulism for the Hilbert port. SIAM Journal on Applied Mathematics, 18 (2):467–488, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rothfarb:1970:CTM

[532] B. Rothfarb and I. T. Frisch. Common terminal multicommodity flow with a concave objective function. SIAM Journal on Applied Mathematics, 18 (2):489–502, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sabharwal:1970:MSF

[533] C. L. Sabharwal. Multiplication of singularity functions by discontinuous functions in Schwartz distribution theory. SIAM Journal on Applied Mathematics, 18(2):503–510, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Katz:1970:RPB

[534] Irving Jack Katz. Remarks on a paper of Ben-Israel: “On the Geometry of Subspaces in Euclidean n-Spaces” [SIAM J. Appl. Math. 15 (1967), no. 5, 1184–1198]. SIAM Journal on Applied Mathematics, 18(2):511–513, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [220].

Dunkel:1970:NFD

[535] Gregory M. Dunkel. On nested functional differential equations. SIAM Journal on Applied Mathematics, 18 (2):514–525, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Anderson:1970:SIE

[536] Marshall W. Anderson. A stochastic integral equation. SIAM Journal on Applied Mathematics, 18(2):526–532, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cellina:1970:MDE

[537] Arrigo Cellina. Multivalued differential equations and ordinary differential equations. SIAM Journal on Applied Mathematics, 18(2):533–538, March 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gill:1970:SAR

[538] A. Gill. Symmetric and antisymmetric regular expressions. SIAM Journal on Applied Mathematics, 18(3):539–557, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bell:1970:IAI

[539] Colin E. Bell. Improved algorithms for inventory and replacement-stocking problems. SIAM Journal on Applied Mathematics, 18(3):558–566, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Billera:1970:STC

[540] Louis J. Billera. Some theorems on the core of an n-person game without side-payments. SIAM Journal on Applied Mathematics, 18(3):567–579, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morris:1970:SVF

[541] William L. Morris. On the stationary values of a fourth-degree polynomial on the unit sphere. SIAM Journal on Applied Mathematics, 18(3):580–583, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zischka:1970:IFB

[542] K. A. Zischka. On the integral form of the Boltzmann equation of neutron transport as applied to finite cells. SIAM Journal on Applied Mathematics, 18(3):584–595, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Suryanarayan:1970:CFF

[543] E. R. Suryanarayan. On a class of flows in fluid dynamics. SIAM Journal on Applied Mathematics, 18(3):596–600, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kallina:1970:PSL

[544] Carl Kallina. Periodicity and stability for linear and quasi-linear parabolic equations. SIAM Journal on Applied Mathematics, 18(3):601–613, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Prasad:1970:PDA

[545] Ram D. Prasad. Probability distributions of algebraic functions of independent random variables. SIAM Journal on Applied Mathematics, 18 (3):614–626, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cannon:1970:SCD

[546] J. R. Cannon and George H. Knightly. Some continuous dependence theorems for viscous fluid motions. SIAM Journal on Applied Mathematics, 18 (3):627–640, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cannon:1970:NCP

[547] J. R. Cannon and George H. Knightly. A note on the Cauchy problem for the Navier–Stokes equations. SIAM Journal on Applied Mathematics, 18 (3):641–644, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reid:1970:CPG

[548] K. B. Reid. Connectivity in products of graphs. SIAM Journal on Applied Mathematics, 18(3):645–651, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eaves:1970:SCC

[549] R. E. Eaves. A sufficient condition for the convergence of an infinite determinant. SIAM Journal on Applied Mathematics, 18(3):652–657, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Osler:1970:LRF

[550] Thomas J. Osler. Leibniz rule for fractional derivatives generalized and an application to infinite series. SIAM Journal on Applied Mathematics, 18 (3):658–674, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Aronsson:1970:ADF

[551] Gunnar Aronsson. Axiomatic derivation of Fermat’s principle. SIAM Journal on Applied Mathematics, 18 (3):675–681, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Muller:1970:SME

[552] P. Chr. Muller. Solution of the matrix equations AX + XB = −Q and S^{T X} + XS = −Q. SIAM Journal on Applied Mathematics, 18(3):682–687, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matula:1970:NSS

[553] David W. Matula. On the number of subtrees of a symmetric n-ary tree. SIAM Journal on Applied Mathematics, 18(3):688–703, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Snyders:1970:NSE

[554] J. Snyders and M. Zakai. On nonnegative solutions of the equation AD + DA′ = −C. SIAM Journal on Applied Mathematics, 18(3):704–714, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Onose:1970:OPC

[555] Hiroshi Onose. Oscillatory property of certain nonlinear ordinary differential equations. SIAM Journal on Applied Mathematics, 18(3):715–719, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Springer:1970:DPB

[556] M. D. Springer and W. E. Thompson. The distribution of products of beta, gamma and Gaussian random variables. SIAM Journal on Applied Mathematics, 18(4):721–737, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wu:1970:SNM

[557] Chien-Heng Wu. On the solutions of a nonlinear membrane problem. SIAM Journal on Applied Mathematics, 18 (4):738–747, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keller:1970:ASI

[558] Joseph B. Keller and Stanley Kogelman. Asymptotic solutions of initial value problems for nonlinear partial differential equations. SIAM Journal on Applied Mathematics, 18(4):748–758, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dafermos:1970:WEW

[559] Constantine M. Dafermos. Wave equations with weak damping. SIAM Journal on Applied Mathematics, 18 (4):759–767, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bigelow:1970:CEP

[560] James H. Bigelow, James C. Dehaven, and Norman Z. Shapiro. Chemical equilibrium problems with unbounded constraint sets. SIAM Journal on Applied Mathematics, 18(4):768–775, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Martin:1970:SES

[561] R. E. Martin and Dan D. Drew. On the simplification of the equations for the stability analysis of anisotropic cylindrical shells. SIAM Journal on Applied Mathematics, 18(4):776–782, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shreve:1970:TVP

[562] Warren E. Shreve. Terminal value problems for second order nonlinear differential equations. SIAM Journal on Applied Mathematics, 18(4):783– 791, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Murray:1970:ASC

[563] J. D. Murray and A. B. Tayler. An asymptotic solution of a class of nonlinear wave equations: A model for the human torso under impulsive stress.

SIAM Journal on Applied Mathematics, 18(4):792–809, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Esary:1970:CLF

[564] J. D. Esary and A. W. Marshall. Coherent life functions. SIAM Journal on Applied Mathematics, 18(4):810–814, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sanday:1970:ESC

[565] S. C. Sanday. Evaluation of the Schwarz–Christoffel mapping function for special polygons. SIAM Journal on Applied Mathematics, 18(4):815–817, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bowman:1970:CRT

[566] John J. Bowman. Comparison of ray theory with exact theory for scattering by open wave guides. SIAM Journal on Applied Mathematics, 18(4):818–829, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bland:1970:TDO

[567] Samuel R. Bland. The two-dimensional oscillating airfoil in a wind tunnel in subsonic flow. SIAM Journal on Applied Mathematics, 18(4):830–848, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Esary:1970:SRA

[568] J. D. Esary, A. W. Marshall, and F. Proschan. Some reliability applications of the hazard transform. SIAM Journal on Applied Mathematics, 18 (4):849–860, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Porsching:1970:TDN

[569] T. A. Porsching. Time-dependent nonlinear networks. SIAM Journal on Applied Mathematics, 18(4):861–871, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1970:NDS

[570] Bernard J. Matkowsky. Nonlinear dynamic stability: A formal theory. SIAM Journal on Applied Mathematics, 18(4):872–883, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Percus:1970:RWC

[571] Ora Engelberg Percus and Jerome K. Percus. Random walk and the comparison of two empirical distributions. SIAM Journal on Applied Mathematics, 18(4):884–886, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lowengrub:1970:ERI

[572] M. Lowengrub and R. P. Srivastav. Effect of rigid inclusions in Griffith cracks. SIAM Journal on Applied Mathematics, 18(4):887–893, June 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wu:1970:GSB

[573] T. T. Wu and J. L. Nowinski. A general solution in bipolar coordinates to problems involving elastic dislocations. SIAM Journal on Applied Mathematics, 19(1):1–19, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wolfe:1970:DWS

[574] Peter Wolfe. The diffraction of waves by slits and strips. SIAM Journal on Applied Mathematics, 19(1):20–32, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bush:1970:HSI

[575] William B. Bush. Hypersonic stronginteraction solutions for flow past a very slender axisymmetric body. SIAM Journal on Applied Mathematics, 19(1):33–50, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kreith:1970:LES

[576] Kurt Kreith. Local essential selfadjointness for singular elliptic operators. SIAM Journal on Applied Mathematics, 19(1):51–55, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Utz:1970:PSN

[577] W. R. Utz. Periodic solutions of a nonlinear second order differential equation. SIAM Journal on Applied Mathematics, 19(1):56–59, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nelson:1970:TIS

[578] Paul Nelson, Jr. and T. S. Kress. Theory and iterative solution of a model for fission product deposition.

SIAM Journal on Applied Mathematics, 19(1):60–74, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

MacWilliams:1970:BCC

[579] F. J. MacWilliams. On binary cyclic codes which are also cyclic codes over GF (2^{s}). SIAM Journal on Applied Mathematics, 19(1):75–95, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Swick:1970:ABS

[580] K. E. Swick. Asymptotic behavior of the solutions of certain third order differential equations. SIAM Journal on Applied Mathematics, 19(1):96–102, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Milch:1970:RWA

[581] Paul R. Milch and Mark H. Waggoner. A random walk approach to a shutdown queueing system. SIAM Journal on Applied Mathematics, 19(1):103– 115, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Clements:1970:MHC

[582] D. L. Clements. On the motion of a heavy cylinder over the surface of an anisotropic elastic solid. SIAM Journal on Applied Mathematics, 19(1):116– 124, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Avriel:1970:CGP

[583] M. Avriel and A. C. Williams. Complementary geometric programming.

SIAM Journal on Applied Mathematics, 19(1):125–141, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gillies:1970:PCA

[584] A. W. Gillies. The physical content of Autonne’s lemma. SIAM Journal on Applied Mathematics, 19(1):142–143, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grimshaw:1970:PWF

[585] R. Grimshaw. A progressing wave formalism for surface waves. SIAM Journal on Applied Mathematics, 19 (1):144–154, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sivazlian:1970:AAO

[586] B. D. Sivazlian. Asymptotic approximations to ordering policies in a stationary (σ, S) inventory problem. SIAM Journal on Applied Mathematics, 19(1):155–166, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Minamide:1970:RPA

[587] Nariyasu Minamide and Kahei Nakamura. A restricted pseudoinverse and its application to constrained minima. SIAM Journal on Applied Mathematics, 19(1):167–177, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Richardson:1970:RRA

[588] Henry R. Richardson. Recursive regression analysis when the least-squares estimate is not asymptotically efficient.

SIAM Journal on Applied Mathematics, 19(1):178–190, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Blum:1970:SID

[589] M. Blum. On the sums of independently distributed Pareto variates. SIAM Journal on Applied Mathematics, 19(1):191–198, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Locker:1970:EAN

[590] John Locker. An existence analysis for nonlinear boundary value problems. SIAM Journal on Applied Mathematics, 19(1):199–207, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dickey:1970:ISN

[591] R. W. Dickey. Infinite systems of nonlinear oscillation equations with linear damping. SIAM Journal on Applied Mathematics, 19(1):208–214, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Duncan:1970:CMI

[592] Tyrone E. Duncan. On the calculation of mutual information. SIAM Journal on Applied Mathematics, 19(1):215– 220, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Day:1970:RBS

[593] William B. Day and William E. Boyce. On the relationship between the solution of a stochastic boundary value problem and parameters in the boundary conditions. SIAM Journal on Applied Mathematics, 19(1):221–234, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schroder:1970:RDI

[594] Johann Schroder. Range-domain implications for linear operators. SIAM Journal on Applied Mathematics, 19 (1):235–242, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fosdick:1970:CRB

[595] Roger L. Fosdick. On the complete representation of biharmonic functions. SIAM Journal on Applied Mathematics, 19(1):243–250, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Saunders:1970:PIM

[596] Sam C. Saunders. A probabilistic interpretation of Miner’s rule. II. SIAM Journal on Applied Mathematics, 19(1):251–265, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lee:1970:IRS

[597] P. A. Lee. An integral representation and some summation formulas for the Hahn polynomials. SIAM Journal on Applied Mathematics, 19(2): 266–272, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Murray:1970:PED

[598] J. D. Murray. Perturbation effects on the decay of discontinuous solutions of nonlinear first order wave equations. SIAM Journal on Applied Mathematics, 19(2):273–298, September 1970.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Aronson:1970:RPF

[599] D. G. Aronson. Regularity properties of flows through porous media: A counterexample. SIAM Journal on Applied Mathematics, 19(2):299–307, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bereanu:1970:RPS

[600] Bernard Bereanu. Renewal processes and some stochastic programming problems in economics. SIAM Journal on Applied Mathematics, 19 (2):308–322, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [854].

Ross:1970:PAD

[601] L. W. Ross. Perturbation analysis of diffusion-coupled biochemical reaction kinetics. SIAM Journal on Applied Mathematics, 19(2):323–329, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gibson:1970:CMF

[602] P. M. Gibson. Combinatorial matrix functions and 1-factors of graphs. SIAM Journal on Applied Mathematics, 19(2):330–333, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huston:1970:KBM

[603] Ronald L. Huston. Krylov–Bogoljubov method for difference equations. SIAM Journal on Applied Mathematics, 19 (2):334–339, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kumar:1970:MGT

[604] S. Kumar. Multinomial group-testing. SIAM Journal on Applied Mathematics, 19(2):340–350, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ranger:1970:MHD

[605] K. B. Ranger. M. H. D. rotation of a conducting sphere in the presence of a uniform magnetic field. SIAM Journal on Applied Mathematics, 19 (2):351–355, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lientz:1970:RNM

[606] B. P. Lientz. Results on nonparametric modal intervals. SIAM Journal on Applied Mathematics, 19(2): 356–366, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dunkel:1970:MSY

[607] Gregory M. Dunkel. Maximum sustainable yields. SIAM Journal on Applied Mathematics, 19(2):367–378, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Booth:1970:UIT

[608] Dexter J. Booth. A useful integration technique. SIAM Journal on Applied Mathematics, 19(2):379–381, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Trench:1970:CSP

[609] William F. Trench. Continuous smoothing with polynomial weighting. SIAM Journal on Applied Mathematics, 19(2):382–389, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schindler:1970:SWM

[610] G. M. Schindler. Simple waves in multidimensional gas flow. SIAM Journal on Applied Mathematics, 19(2): 390–407, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kappel:1970:ILS

[611] F. Kappel. The invariance of limit sets for autonomous functional-differential equations. SIAM Journal on Applied Mathematics, 19(2):408–419, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Delsarte:1970:BBC

[612] P. Delsarte. BCH bounds for a class of cyclic codes. SIAM Journal on Applied Mathematics, 19(2):420–429, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hallam:1970:AIS

[613] Thomas G. Hallam. Asymptotic integration of second order differential equations with integrable coefficients. SIAM Journal on Applied Mathematics, 19(2):430–439, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schmidt:1970:LBM

[614] K. W. Schmidt. Lower bounds for maximal (0,1)-determinants. SIAM Journal on Applied Mathematics, 19(2): 440–442, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brenner:1970:GTR

[615] J. L. Brenner. Gersgorin theorems, regularity theorems, and bounds for determinants of partitioned matrices. SIAM Journal on Applied Mathematics, 19 (2):443–450, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baker:1970:MIG

[616] Charles R. Baker. Mutual information for Gaussian processes. SIAM Journal on Applied Mathematics, 19 (2):451–458, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kannai:1970:CCS

[617] Yakar Kannai. On closed coverings of simplexes. SIAM Journal on Applied Mathematics, 19(2):459–461, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Packard:1970:EPG

[618] R. W. Packard. Endpoint properties of the Green’s operator of a generalized second order linear differential equation. SIAM Journal on Applied Mathematics, 19(2):462–472, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Parthasarathy:1970:GUS

[619] T. Parthasarathy. On games over the unit square. SIAM Journal on Applied Mathematics, 19(2):473–476, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hodes:1970:DAC

[620] Louis Hodes. Discrete approximation of continuous convex blobs. SIAM Journal on Applied Mathematics, 19 (2):477–485, September 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peterson:1970:SDG

[621] Elmor L. Peterson. Symmetric duality for generalized unconstrained geometric programming. SIAM Journal on Applied Mathematics, 19(3): 487–526, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stephan:1970:HEC

[622] Bruce H. Stephan. On hereditary equations close to ordinary differential equations. SIAM Journal on Applied Mathematics, 19(3):527–531, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rogge:1970:LWC

[623] Thomas R. Rogge. Longitudinal waves in circular conic sections. SIAM Journal on Applied Mathematics, 19(3): 532–541, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jackson:1970:USB

[624] Lloyd Jackson and Gene Klaasen. Uniqueness of solutions of boundary value problems for ordinary differential equations. SIAM Journal on Applied Mathematics, 19(3):542–546, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hurt:1970:GIW

[625] Michiel F. Hurt and Carter Waid. A generalized inverse which gives all the integral solutions to a system of linear equations. SIAM Journal on Applied Mathematics, 19(3):547–550, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Forsythe:1970:MMP

[626] George E. Forsythe. The maximum and minimum of a positive definite quadratic polynomial on a sphere are convex functions of the radius. SIAM Journal on Applied Mathematics, 19 (3):551–554, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Harwood:1970:NGI

[627] W. R. Harwood, V. Lovass-Nagy, and D. L. Powers. A note on the generalized inverses of some partitioned matrices. SIAM Journal on Applied Mathematics, 19(3):555–559, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knoshaug:1970:GS

[628] Clayton Knoshaug. Generalized sequences. SIAM Journal on Applied Mathematics, 19(3):560–565, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wolff:1970:TSP

[629] Ronald W. Wolff. Time sharing with priorities. SIAM Journal on Applied Mathematics, 19(3):566–574, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zauderer:1970:BLU

[630] Erich Zauderer. Boundary layer and uniform asymptotic expansions for diffraction problems. SIAM Journal on Applied Mathematics, 19(3): 575–600, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bobisud:1970:OSD

[631] L. E. Bobisud. Oscillation of solutions of damped nonlinear equations. SIAM Journal on Applied Mathematics, 19 (3):601–606, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Marek:1970:FTP

[632] Ivo Marek. Frobenius theory of positive operators: Comparison theorems and applications. SIAM Journal on Applied Mathematics, 19(3):607–628, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Heidel:1970:GAS

[633] J. W. Heidel. Global asymptotic stability of a generalized Lienard equation. SIAM Journal on Applied Mathematics, 19(3):629–636, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mazumdar:1970:RTU

[634] M. Mazumdar. Reliability of two-unit redundant repairable systems when failures are revealed by inspections. SIAM Journal on Applied Mathematics, 19(4):637–647, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohoon:1970:CRI

[635] D. K. Cohoon. A continuous right inverse for second order differential operators with variable coefficients. SIAM Journal on Applied Mathematics, 19 (4):648–657, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Subrahmaniam:1970:SAM

[636] K. Subrahmaniam. On some applications of mellin transforms to statistics: Dependent random variables. SIAM Journal on Applied Mathematics, 19 (4):658–662, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baker:1970:PRR

[637] James D. Baker. Probability representations with right-cauchy integrals. SIAM Journal on Applied Mathematics, 19(4):663–667, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Householder:1970:PRA

[638] Alston S. Householder. On the penultimate remainder algorithm and the catalytic multiplier. SIAM Journal on Applied Mathematics, 19(4):668–671, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vandewiele:1970:GAN

[639] Georges Vandewiele. A generalization of the albert-nelson counter with semi-Markov input. SIAM Journal on Applied Mathematics, 19(4):672–678, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Man:1970:SIP

[640] F. T. Man. Some inequalities for positive definite symmetric matrices. SIAM Journal on Applied Mathematics, 19 (4):679–681, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cline:1970:EGI

[641] Randall E. Cline and T. N. E. Greville. An extension of the generalized inverse of a matrix. SIAM Journal on Applied Mathematics, 19(4): 682–688, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Christofides:1970:SHC

[642] Nicos Christofides. The shortest Hamiltonian chain of a graph. SIAM Journal on Applied Mathematics, 19(4): 689–696, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Robertson:1970:CA

[643] H. G. Robertson. Curvature and arclength. SIAM Journal on Applied Mathematics, 19(4):697–699, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Simon:1970:OCP

[644] R. M. Simon. Optimal cannibalization policies for multicomponent systems. SIAM Journal on Applied Mathematics, 19(4):700–711, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klemm:1970:SIE

[645] James L. Klemm and Robert Wm. Little. The semi-infinite elastic cylinder under self-equilibrated end loading. SIAM Journal on Applied Mathematics, 19(4):712–729, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ghosh:1970:QFE

[646] Sakti P. Ghosh. On quadrics in finite Euclidean geometrics and application to construction of balanced incomplete block designs. SIAM Journal on Applied Mathematics, 19(4): 730–740, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1970:GIB

[647] Carl D. Meyer, Jr. Generalized inverses of block triangular matrices. SIAM Journal on Applied Mathematics, 19 (4):741–750, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maria:1970:RBB

[648] N. L. Maria. Relation between boundary value problems in a half-space and the whole space. SIAM Journal on Applied Mathematics, 19(4):751–768, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cooke:1970:EPE

[649] Charlie H. Cooke. An efficiency property of the exponentially weighted least squares estimators. SIAM Journal on Applied Mathematics, 19(4):769–777, December 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1971:MSS

[650] Donald S. Cohen. Multiple stable solutions of nonlinear boundary value problems arising in chemical reactor theory. SIAM Journal on Applied Mathematics, 20(1):1–13, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ford:1971:MPB

[651] Wayne T. Ford. A mapping in phase behavior. SIAM Journal on Applied Mathematics, 20(1):14–23, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tan:1971:NEG

[652] W. Y. Tan. Note on an extension of the Gauss–Markov theorems to multivariate linear regression models. SIAM Journal on Applied Mathematics, 20 (1):24–29, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hopcroft:1971:MNM

[653] J. E. Hopcroft and L. R. Kerr. On minimizing the number of multiplications necessary for matrix multiplication. SIAM Journal on Applied Mathematics, 20(1):30–36, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huilgol:1971:LTB

[654] R. R. Huilgol. On Liouville’s theorem for biharmonic functions. SIAM Journal on Applied Mathematics, 20 (1):37–39, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Krall:1971:ADS

[655] Allan M. Krall. Adjoint differential systems in inner product spaces. SIAM Journal on Applied Mathematics, 20 (1):40–43, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rhoten:1971:LZF

[656] R. P. Rhoten and J. K. Aggarwal. The location of zeros of finite exponential series. SIAM Journal on Applied Mathematics, 20(1):44–50, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cabayan:1971:CSL

[657] Hrair S. Cabayan and Geneva G. Belford. On computing a stable least squares solution to the inverse problem for a planar Newtonian potential. SIAM Journal on Applied Mathematics, 20(1):51–61, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kohlberg:1971:NCF

[658] Elon Kohlberg. On the nucleolus of a characteristic function game. SIAM Journal on Applied Mathematics, 20 (1):62–66, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Greenspan:1971:NSB

[659] Donald Greenspan. Numerical studies of the 3-body problem. SIAM Journal on Applied Mathematics, 20 (1):67–78, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chan:1971:OPC

[660] S. Y. Chan and W. S. Chou. An optimization problem in communication nets. SIAM Journal on Applied Mathematics, 20(1):79–86, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eberlein:1971:NMD

[661] P. J. Eberlein. A note on the matrices denoted B_{n}. SIAM Journal on Applied Mathematics, 20(1):87–92, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kesarwani:1971:ITI

[662] Roop Narain Kesarwani. On an integral transform involving G-functions. SIAM Journal on Applied Mathematics, 20(1):93–98, January 1971. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bierman:1971:PCS

[663] Gerald J. Bierman. A particular class of singular integral equations. SIAM Journal on Applied Mathematics, 20 (1):99–109, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pandey:1971:GWH

[664] J. N. Pandey. The generalized Weierstrass Hankel convolution transform. SIAM Journal on Applied Mathematics, 20(1):110–123, January 1971. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jackson:1971:VTM

[665] Lloyd K. Jackson and Gene Klaasen. A variation of the topological method of wazewski. SIAM Journal on Applied Mathematics, 20(1):124–130, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dhaliwal:1971:SST

[666] Ranjit S. Dhaliwal. The steady state thermoelastic mixed boundary value problem for the elastic layer. SIAM Journal on Applied Mathematics, 20 (1):131–140, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fujii:1971:EOS

[667] M. Fujii, T. Kasami, and K. Ninomiya. Erratum: “Optimal Sequencing of Two Equivalent Processors” [SIAM J. Appl. Math. 17 (1969), no. 4, 784–789]. SIAM Journal on Applied Mathematics, 20(1):141, January 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [437].

Rusin:1971:RSM

[668] M. H. Rusin. A revised simplex method for quadratic programming. SIAM Journal on Applied Mathematics, 20 (2):143–160, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Alavi:1971:CGE

[669] Y. Alavi and M. Behzad. Complementary graphs and edge chromatic numbers. SIAM Journal on Applied Mathematics, 20(2):161–163, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gould:1971:NSQ

[670] F. J. Gould and Jon W. Tolle. A necessary and sufficient qualification for constrained optimization. SIAM Journal on Applied Mathematics, 20 (2):164–172, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ward:1971:NOM

[671] J. F. Ward, T. L. Boullion, and T. O. Lewis. A note on the oblique matrix pseudoinverse. SIAM Journal on Applied Mathematics, 20(2):173–175, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boesch:1971:IRC

[672] F. T. Boesch and A. Felzer. On the invulnerability of the regular complete k-partite graphs. SIAM Journal on Applied Mathematics, 20(2):176–182, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohn:1971:SCS

[673] M. Cohn, S. Even, S. W. Golomb, and A. Lempel. The stability of counting sequences under stage delays. SIAM Journal on Applied Mathematics, 20 (2):183–188, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chan:1971:UNF

[674] C. Y. Chan. Uniqueness of a nonmonotone free boundary problem. SIAM Journal on Applied Mathematics, 20 (2):189–194, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mitra:1971:PDS

[675] Samarendra Kumar Mitra. On the probability distribution of the sum of uniformly distributed random variables. SIAM Journal on Applied Mathematics, 20(2):195–198, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Graveleau:1971:FDA

[676] J. L. Graveleau and P. Jamet. A finite difference approach to some degenerate nonlinear parabolic equations. SIAM Journal on Applied Mathematics, 20 (2):199–223, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Haji:1971:OSP

[677] Rasoul Haji and G. F. Newell. Optimal strategies for priority queues with nonlinear costs of delay. SIAM Journal on Applied Mathematics, 20 (2):224–240, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stone:1971:OSU

[678] Lawrence D. Stone and Jerome A. Stanshine. Optimal search using uninterrupted contact investigation. SIAM Journal on Applied Mathematics, 20 (2):241–263, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [792].

Sivazlian:1971:SAM

[679] B. D. Sivazlian. Stationary analysis of a multicommodity inventory system with interacting set-up costs. SIAM Journal on Applied Mathematics, 20 (2):264–278, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bush:1971:LMM

[680] William B. Bush. On the Lagerstrom mathematical model for viscous flow at low Reynolds number. SIAM Journal on Applied Mathematics, 20 (2):279–287, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mizuki:1971:FPT

[681] Mikiso Mizuki. Frequency of passages through a designated sequence or network of states of a Markov chain. SIAM Journal on Applied Mathematics, 20 (2):288–297, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Onaga:1971:OSD

[682] Kenji Onaga. Optimal search for detecting a hidden object. SIAM Journal on Applied Mathematics, 20 (2):298–318, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sherman:1971:LBS

[683] B. Sherman. Limiting behavior in some Stefan problems as the latent heat goes to zero. SIAM Journal on Applied Mathematics, 20(2):319–327, March 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wood:1971:GBP

[684] Bruce Wood. Graphic behavior of positive linear operators. SIAM Journal on Applied Mathematics, 20(3):329–335, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lippman:1971:SDJ

[685] Steven A. Lippman and Sheldon M. Ross. The streetwalker’s dilemma: A job shop model. SIAM Journal on Applied Mathematics, 20(3):336–342, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Afriat:1971:TMM

[686] S. N. Afriat. Theory of maxima and the method of Lagrange. SIAM Journal on Applied Mathematics, 20(3):343–357, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yu:1971:RPS

[687] Chung Ling Yu. Reflection principle for solutions of higher order elliptic equations with analytic coefficients. SIAM Journal on Applied Mathematics, 20(3):358–363, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kevorkian:1971:PTR

[688] J. Kevorkian. Passage through resonance for a one-dimensional oscillator with slowly varying frequency. SIAM Journal on Applied Mathematics, 20(3):364–373, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [1101].

Duvall:1971:NCM

[689] P. F. Duvall, Jr. and P. W. Harley, III. A note on counting matrices. SIAM Journal on Applied Mathematics, 20(3):374–377, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Murty:1971:CM

[690] Katta G. Murty. On a characterization of P -matrices. SIAM Journal on Applied Mathematics, 20(3):378–384, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Williams:1971:DGI

[691] T. W. Williams and L. M. Maxwell. The decomposition of a graph and the introduction of a new class of subgraphs. SIAM Journal on Applied Mathematics, 20(3):385–389, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keller:1971:DCW

[692] Joseph B. Keller and Daljit S. Ahluwalia. Diffraction by a curved wire. SIAM Journal on Applied Mathematics, 20(3):390–405, May 1971.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zarghamee:1971:DVF

[693] Mehdi S. Zarghamee. Differential of variational forms. SIAM Journal on Applied Mathematics, 20(3):406–412, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1971:IPC

[694] Donald S. Cohen and Erik B. Hansen. Inverse problems connected with twopoint boundary value problems. SIAM Journal on Applied Mathematics, 20 (3):413–425, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Plemmons:1971:GIB

[695] R. J. Plemmons. Generalized inverses of Boolean relation matrices. SIAM Journal on Applied Mathematics, 20(3):426–433, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cannon:1971:DFM

[696] J. R. Cannon and G. H. Meyer. On diffusion in a fractured medium. SIAM Journal on Applied Mathematics, 20 (3):434–448, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pierre:1971:CLT

[697] Percy A. Pierre. Central limit theorems for conditionally linear random processes. SIAM Journal on Applied Mathematics, 20(3):449–461, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nemat-Nasser:1971:CNM

[698] S. Nemat-Nasser and H. D. Shatoff. A consistent numerical method for the solution of nonlinear elasticity problems at finite strains. SIAM Journal on Applied Mathematics, 20(3):462–481, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shamma:1971:ABS

[699] S. E. Shamma. Asymptotic behavior of Stekloff eigenvalues and eigenfunctions. SIAM Journal on Applied Mathematics, 20(3):482–490, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Comstock:1971:SPE

[700] Craig Comstock. Singular perturbations of elliptic equations. I. SIAM Journal on Applied Mathematics, 20 (3):491–502, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miller:1971:ITF

[701] Willard Miller, Jr. Invariant tensor fields in physics and the classical groups. SIAM Journal on Applied Mathematics, 20(3):503–519, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Anderson:1971:SO

[702] William N. Anderson, Jr. Shorted operators. SIAM Journal on Applied Mathematics, 20(3):520–525, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1971:VSA

[703] Wai kai Chen. On vector spaces associated with a graph. SIAM Journal on Applied Mathematics, 20(3):526–529, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Serfozo:1971:FSM

[704] Richard F. Serfozo. Functions of semi-Markov processes. SIAM Journal on Applied Mathematics, 20(3):530–535, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Karp:1971:GSB

[705] S. N. Karp and S. E. Shamma. A generalization of separability in boundary value problems. SIAM Journal on Applied Mathematics, 20(3):536–546, May 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shenton:1971:CFP

[706] L. R. Shenton and K. O. Bowman. Continued fractions for the psi function and its derivatives. SIAM Journal on Applied Mathematics, 20(4):547–554, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://www.jstor.org/
stable/2099856.

Sherman:1971:GOP

[707] B. Sherman. General one-phase Stefan problems and free boundary problems for the heat equation with Cauchy data prescribed on the free boundary. SIAM Journal on Applied Mathematics, 20(4):555–570, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Suber:1971:IMP

[708] H. H. Suber. Integral manifolds for perturbed nonlinear systems. SIAM Journal on Applied Mathematics, 20 (4):571–589, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Krishnamurthy:1971:KR

[709] L. Krishnamurthy and F. A. Williams. Kinetics and regression. SIAM Journal on Applied Mathematics, 20(4):590– 611, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1971:ASP

[710] J. L. Smith. Approximate stationary probability vectors of a finite Markov chain. SIAM Journal on Applied Mathematics, 20(4):612–618, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kogelman:1971:TBU

[711] Stanley Kogelman and Joseph B. Keller. Transient behavior of unstable nonlinear systems with applications to the Bénard and Taylor problems. SIAM Journal on Applied Mathematics, 20(4):619–637, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Padberg:1971:RIN

[712] Manfred W. Padberg. A remark on “An Inequality for the Number of Lattice Points in a Simplex”. SIAM Journal on Applied Mathematics, 20 (4):638–641, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jain:1971:IEM

[713] D. L. Jain and R. P. Kanwal. An integral equation method for solving mixed boundary value problems. SIAM Journal on Applied Mathematics, 20 (4):642–658, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Anderson:1971:SIE

[714] Brian D. O. Anderson and Thomas Kailath. Some integral equations with nonsymmetric separable kernels. SIAM Journal on Applied Mathematics, 20 (4):659–669, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Buchanan:1971:GOS

[715] J. Buchanan. Growth of oscillatory solutions of y′(x) = y(x − d(x)). SIAM Journal on Applied Mathematics, 20(4):670–676, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Crown:1971:NAC

[716] J. C. Crown. A new approach to the calculation of transonic flows. SIAM Journal on Applied Mathematics, 20 (4):677–697, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sposito:1971:SPO

[717] Vince A. Sposito and H. T. David. Saddle-point optimality criteria of nonlinear programming problems over cones without differentiability. SIAM Journal on Applied Mathematics, 20 (4):698–702, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Martin:1971:NIE

[718] Benjamin J. Martin. On a new integral equation arising in the theory of radiative transfer. SIAM Journal on Applied Mathematics, 20(4):703–713, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tam:1971:ASV

[719] K. K. Tam. On the asymptotic solution of viscous incompressible flow past a heated paraboloid of revolution. SIAM Journal on Applied Mathematics, 20(4):714–721, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [896].

Dean:1971:BSN

[720] Edward T. Dean and Paul L. Chambre. On the bifurcation of solutions of the nonlinear eigenvalue problem Lu + λb(x)u = g(x, u). SIAM Journal on Applied Mathematics, 20(4):722–734, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rajput:1971:BDS

[721] Balram S. Rajput. Bounded determinable subsets of Banach spaces. SIAM Journal on Applied Mathematics, 20(4):735–748, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jensen:1971:NPD

[722] D. R. Jensen. A note on positive dependence and the structure of bivariate distributions. SIAM Journal on Applied Mathematics, 20(4):749–753, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gaines:1971:DRB

[723] Robert Gaines. Differentiability with respect to boundary values for nonlinear ordinary differential equations. SIAM Journal on Applied Mathematics, 20(4):754–762, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Passy:1971:GWM

[724] U. Passy. Generalized weighted mean programming. SIAM Journal on Applied Mathematics, 20(4):763–778, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lugannani:1971:CPS

[725] Robert Lugannani. Convergence properties of the sample mean and sample correlation for a class of pulse trains. SIAM Journal on Applied Mathematics, 21(1):1–12, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Papanicolaou:1971:WPO

[726] George C. Papanicolaou. Wave propagation in a one-dimensional random medium. SIAM Journal on Applied Mathematics, 21(1):13–18, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rajput:1971:UDS

[727] Balram S. Rajput. Unbounded determinable subsets of Banach spaces. SIAM Journal on Applied Mathematics, 21(1):19–29, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosen:1971:MVS

[728] Gerald Rosen. Minimum value for c in the Sobolev inequality ∥ϕ^{3}∥ ≤ c∥∇ϕ∥^{3}. SIAM Journal on Applied Mathematics, 21(1):30–32, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berk:1971:ESS

[729] Kenneth N. Berk. Estimating the solution of a stochastic difference equation. SIAM Journal on Applied Mathematics, 21(1):33–43, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schaefer:1971:BCP

[730] Philip W. Schaefer. Bounds in the Cauchy problem for a fourth order quasi-linear equation. SIAM Journal on Applied Mathematics, 21(1):44–50, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berman:1971:NPH

[731] Abraham Berman and Adi Ben-Israel. A note on pencils of Hermitian or symmetric matrices. SIAM Journal on Applied Mathematics, 21(1):51–54, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jeroslow:1971:AG

[732] R. G. Jeroslow and K. O. Kortanek. On an algorithm of Gomory. SIAM Journal on Applied Mathematics, 21(1):55–60, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kallianpur:1971:SDE

[733] G. Kallianpur and C. Striebel. A stochastic differential equation of Fisk type for estimation and nonlinear filtering problems. SIAM Journal on Applied Mathematics, 21(1):61–72, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosenmuller:1971:GLH

[734] J. Rosenmuller. On a generalization of the Lemke–Howson algorithm to noncooperative N-person games. SIAM Journal on Applied Mathematics, 21(1):73–79, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wilson:1971:CEP

[735] Robert Wilson. Computing equilibria of N-person games. SIAM Journal on Applied Mathematics, 21(1):80–87, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rahman:1971:IAG

[736] M. Rahman. On the integrability and application of the generalized Riccati equation. SIAM Journal on Applied Mathematics, 21(1):88–94, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Winston:1971:EPS

[737] Elliot Winston. The existence of periodic solutions of perturbed periodic systems. SIAM Journal on Applied Mathematics, 21(1):95–103, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

White:1971:MCF

[738] Lee J. White. Minimum covers of fixed cardinality in weighted graphs. SIAM Journal on Applied Mathematics, 21(1):104–113, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moses:1971:ECO

[739] Harry E. Moses. Eigenfunctions of the curl operator, rotationally invariant Helmholtz theorem, and applications to electromagnetic theory and fluid mechanics. SIAM Journal on Applied Mathematics, 21(1):114–144, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ueda:1971:CCP

[740] Takao Ueda. The continuity of a critical point set with applications to some decision problems. SIAM Journal on Applied Mathematics, 21(1):145–154, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lee:1971:DEV

[741] P. A. Lee. A diagonal expansion for the 2-variate Dirichlet probability density function. SIAM Journal on Applied Mathematics, 21(1):155–165, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miller:1971:MNF

[742] Bruce L. Miller. On minimizing nonseparable functions defined on the integers with an inventory application. SIAM Journal on Applied Mathematics, 21(1):166–185, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tallis:1971:NRP

[743] G. M. Tallis and G. Gordon. A note on the roots of the polynomial equation f(x) = a with reference to stability. SIAM Journal on Applied Mathematics, 21(1):186–190, July 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Langenhop:1971:ISM

[744] C. E. Langenhop. On the index of a square matrix. SIAM Journal on Applied Mathematics, 21(2): 191–194, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Suhadolc:1971:SID

[745] Anton Suhadolc. On a system of integro-differential equations. SIAM Journal on Applied Mathematics, 21 (2):195–206, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kung:1971:OND

[746] George C. T. Kung. Oscillation and nonoscillation of differential equations with a time lag. SIAM Journal on Applied Mathematics, 21(2):207–213, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dachs:1971:CPF

[747] Jose N. W. Dachs. A convergence problem on finite directed graphs. SIAM Journal on Applied Mathematics, 21 (2):214–220, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chiou:1971:SON

[748] Kuo-Liang Chiou. A second order nonlinear oscillation theorem. SIAM Journal on Applied Mathematics, 21 (2):221–224, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Macki:1971:NBV

[749] Jack W. Macki and Paul Waltman. A nonlinear boundary value problem of Sturm–Liouville type for a twodimensional system of ordinary differential equations. SIAM Journal on Applied Mathematics, 21(2):225–231, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bernstein:1971:AEA

[750] Allen R. Bernstein. Almost eigenvectors for almost commuting matrices. SIAM Journal on Applied Mathematics, 21(2):232–235, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eaves:1971:CKF

[751] B. Curtis Eaves. Computing Kakutani fixed points. SIAM Journal on Applied Mathematics, 21(2):236–244, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kiyono:1971:SLE

[752] Takeshi Kiyono and Masaaki Shimasaki. On the solution of Laplace’s equation by certain dual series equations. SIAM Journal on Applied Mathematics, 21(2):245–257, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ke:1971:BFE

[753] William Oo Kian Ke. Bounds for the first eigenvalue of the differential system y′′ + λp(x)y = 0, y(a) = y(b) = O. SIAM Journal on Applied Mathematics, 21(2):258–264, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McKenna:1971:LVC

[754] James McKenna. On the lateral vibration of conical bars. SIAM Journal on Applied Mathematics, 21(2): 265–278, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Benson:1971:CT

[755] Donald C. Benson. Comparison theory for y^{′2} − 2p(x)y = −p(x)q(x). SIAM Journal on Applied Mathematics, 21 (2):279–286, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Papanicolau:1971:SDE

[756] George Papanicolau and Joseph B. Keller. Stochastic differential equations with applications to random harmonic oscillators and wave propagation in random media. SIAM Journal on Applied Mathematics, 21(2): 287–305, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wilkins:1971:MIT

[757] J. Ernest Wilkins, Jr. Maximum internal temperatures in one-dimensional geometries with nonuniform heat generation and equal surface temperatures. SIAM Journal on Applied Mathematics, 21(2):306–320, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rizzo:1971:ACP

[758] F. J. Rizzo and D. J. Shippy. An application of the correspondence principle of linear viscoelasticity theory. SIAM Journal on Applied Mathematics, 21 (2):321–330, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wheeler:1971:IRS

[759] Lewis Wheeler. On the integral representations of Stokes and Love for elastodynamic displacement fields. SIAM Journal on Applied Mathematics, 21 (2):331–338, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barcilon:1971:SPA

[760] Victor Barcilon, Julian D. Cole, and Robert S. Eisenberg. A singular perturbation analysis of induced electric fields in nerve cells. SIAM Journal on Applied Mathematics, 21(2): 339–354, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fink:1971:CPD

[761] A. M. Fink. A converse problem in differential equations. SIAM Journal on Applied Mathematics, 21(2): 355–361, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fink:1971:NCE

[762] A. M. Fink and G. Seifert. Nonresonance conditions for the existence of almost periodic solutions of almost periodic systems. SIAM Journal on Applied Mathematics, 21(2):362–366, September 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Duvall:1971:DPS

[763] Paul F. Duvall and Jack C. Mortick. Decimation of periodic sequences. SIAM Journal on Applied Mathematics, 21(3):367–372, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schrack:1971:LBA

[764] Gunther Friedemann Schrack. Lower bounds to the abscissa of stability of a stable polynomial from symmetric functions. SIAM Journal on Applied Mathematics, 21(3):373–379, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bruen:1971:BSF

[765] A. Bruen. Blocking sets in finite projective planes. SIAM Journal on Applied Mathematics, 21(3):380–392, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bush:1971:TDS

[766] William B. Bush and Robert Rosen. The three-dimensional steady radial expansion of a viscous gas from a sonic source into a vacuum. SIAM Journal on Applied Mathematics, 21 (3):393–406, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nathanson:1971:DBS

[767] Melvyn B. Nathanson. Derivatives of binary sequences. SIAM Journal on Applied Mathematics, 21(3): 407–412, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kozin:1971:NSC

[768] F. Kozin and S. Prodromou. Necessary and sufficient conditions for almost sure sample stability of linear Ito equations. SIAM Journal on Applied Mathematics, 21(3):413–424, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kunz:1971:AAT

[769] K. S. Kunz. Application of an algebraic technique to the solution of Laplace’s equation in three dimensions. SIAM Journal on Applied Mathematics, 21 (3):425–441, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohn:1971:WFS

[770] Martin Cohn. Walsh functions, sequency, and Gray codes. SIAM Journal on Applied Mathematics, 21(3): 442–447, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zlobec:1971:EAK

[771] Sanjo Zlobec. Extensions of asymptotic Kuhn–Tucker conditions in mathematical programming. SIAM Journal on Applied Mathematics, 21(3):

448–460, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

English:1971:CIV

[772] William J. English. A computerimplemented vector variational solution of loaded rectangular waveguides. SIAM Journal on Applied Mathematics, 21(3):461–468, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gautesen:1971:SVF

[773] A. K. Gautesen and S. P. Lin. Slow viscous flow past a deforming sphere. SIAM Journal on Applied Mathematics, 21(3):469–477, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Walsh:1971:MPD

[774] T. Walsh. On the maximum of a positive definite quadratic polynomial on a sphere. SIAM Journal on Applied Mathematics, 21(3):478–479, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ward:1971:WPS

[775] J. F. Ward, T. L. Boullion, and T. O. Lewis. Weighted pseudoinverses with singular weights. SIAM Journal on Applied Mathematics, 21(3): 480–482, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grinold:1971:HCL

[776] Richard C. Grinold. The Hirsch conjecture in Leontief substitution systems. SIAM Journal on Applied Mathematics, 21(3):483–485, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burkowski:1971:NOS

[777] Forbes Burkowski. Nonlinear oscillation of a second order sublinear functional differential equation. SIAM Journal on Applied Mathematics, 21 (3):486–490, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schmitt:1971:NPS

[778] Klaus Schmitt. A note on periodic solutions of second order ordinary differential equations. SIAM Journal on Applied Mathematics, 21(3): 491–494, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Martin:1971:ESC

[779] R. H. Martin, Jr. Exponential stability criteria for nonlinear, nonautonomous differential equations. SIAM Journal on Applied Mathematics, 21 (3):495–499, November 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jain:1971:GNB

[780] G. C. Jain and P. C. Consul. A generalized negative binomial distribution. SIAM Journal on Applied Mathematics, 21(4):501–513, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hu:1971:OCS

[781] T. C. Hu and A. C. Tucker. Optimal computer search trees and variablelength alphabetical codes. SIAM Journal on Applied Mathematics, 21(4): 514–532, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1971:PST

[782] A. H. Zemanian. The passivity and semipassivity of time-varying systems under the admittance formulism. SIAM Journal on Applied Mathematics, 21 (4):533–541, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peletier:1971:ABS

[783] L. A. Peletier. Asymptotic behavior of solutions of the porous media equation. SIAM Journal on Applied Mathematics, 21(4):542–551, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eliason:1971:CTR

[784] Stanley B. Eliason. Comparison theorems related to nonlinear eigenvalue problems. SIAM Journal on Applied Mathematics, 21(4):552–564, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lenard:1971:PFF

[785] J. G. Lenard. Plastic flow and fracture of a rotating cylinder. SIAM Journal on Applied Mathematics, 21 (4):565–572, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Krieger:1971:GOI

[786] H. A. Krieger and C. A. Schaffner. The global optimization of incoherent phase signals. SIAM Journal on Applied Mathematics, 21(4):573–589, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gal:1971:SMS

[787] Shmuel Gal. Sequential minimax search for a maximum when prior information is available. SIAM Journal on Applied Mathematics, 21(4): 590–595, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zarghamee:1971:OVV

[788] Mehdi S. Zarghamee. On the optimal value of a variational form. SIAM Journal on Applied Mathematics, 21 (4):596–600, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kayande:1971:TCS

[789] A. A. Kayande. A theorem on contractive stability. SIAM Journal on Applied Mathematics, 21(4):601–604, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lange:1971:PBP

[790] C. G. Lange and A. C. Newell. The post-buckling problem for thin elastic shells. SIAM Journal on Applied Mathematics, 21(4):605–629, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Preparata:1971:RIN

[791] Franco P. Preparata. On the representation of integers in nonadjacent form. SIAM Journal on Applied Mathematics, 21(4):630–635, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stone:1971:EOS

[792] Lawrence D. Stone and Jerome A. Stanshine. Erratum: “Optimal Search Using Uninterrupted Contact Investigation” [SIAM J. Appl. Math. 20 (1971), no. 2, 241–263]. SIAM Journal on Applied Mathematics, 21(4):636, December 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [678].

Wu:1972:WSL

[793] Chien-Heng Wu. Whirling of a string at large angular speeds — A nonlinear eigenvalue problem with moving boundary layers. SIAM Journal on Applied Mathematics, 22(1):1–13, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shamma:1972:AES

[794] S. E. Shamma and S. N. Karp. Asymptotic eigenfunctions of a scattering problem. SIAM Journal on Applied Mathematics, 22(1):14–21, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wall:1972:SIS

[795] J. R. Wall and R. J. Plemmons. Spectral inverses of stochastic matrices. SIAM Journal on Applied Mathematics, 22(1):22–26, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schmidt:1972:FPO

[796] Dieter S. Schmidt. Families of periodic orbits in the restricted problem of three bodies connecting families of direct and retrograde orbits. SIAM Journal on Applied Mathematics, 22 (1):27–37, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Seifert:1972:APS

[797] George Seifert. Almost periodic solutions for limit periodic systems. SIAM Journal on Applied Mathematics, 22 (1):38–44, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Evans:1972:FDR

[798] David S. Evans. Finite-dimensional realizations of discrete-time weighting patterns. SIAM Journal on Applied Mathematics, 22(1):45–67, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Anis:1972:RMM

[799] Abdel-Azim Anis and Emlyn H. Lloyd. Reservoirs with mixed Markovianindependent inflows. SIAM Journal on Applied Mathematics, 22(1):68–76, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Scheid:1972:LSF

[800] Francis Scheid. A least squares family of cubic curves with an application to golf handicapping. SIAM Journal on Applied Mathematics, 22 (1):77–83, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barnett:1972:NBM

[801] S. Barnett. A note on the Bezoutian matrix. SIAM Journal on Applied Mathematics, 22(1):84–86, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pop-Stojanovic:1972:MBS

[802] Zoran R. Pop-Stojanovic. On Mc-Shane’s belated stochastic integral. SIAM Journal on Applied Mathematics, 22(1):87–92, January 1972. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brown:1972:APC

[803] Mark Brown and Sheldon M. Ross. Asymptotic properties of cumulative processes. SIAM Journal on Applied Mathematics, 22(1):93–105, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Beged-Dov:1972:LUB

[804] Aharon Gavriel Beged-Dov. Lower and upper bounds for the number of lattice points in a simplex. SIAM Journal on Applied Mathematics, 22 (1):106–108, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kaminetzky:1972:DCH

[805] Lee Kaminetzky and Joseph B. Keller. Diffraction coefficients for higher order edges and vertices. SIAM Journal on Applied Mathematics, 22(1): 109–134, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kreith:1972:OCC

[806] Kurt Kreith. Oscillation criteria for a class of fourth order differential equations. SIAM Journal on Applied Mathematics, 22(1):135–137, January 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hazony:1972:DRA

[807] D. Hazony and R. Sobecks. A distributed RC active delay line. SIAM Journal on Applied Mathematics, 22 (2):139–141, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bandle:1972:IIS

[808] Catherine Bandle. Isoperimetric inequality for some eigenvalues of an inhomogeneous, free membrane. SIAM Journal on Applied Mathematics, 22 (2):142–147, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1972:IFV

[809] S. H. Smith. On the impulsive flow of a viscous liquid past a semi-infinite flat plate. SIAM Journal on Applied Mathematics, 22(2):148–154, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berman:1972:MGI

[810] Abraham Berman and Robert J. Plemmons. Monotonicity and the generalized inverse. SIAM Journal on Applied Mathematics, 22(2):155–161, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ephremides:1972:MCL

[811] A. Ephremides and J. B. Thomas. Multiplicity considerations in linear least mean-square estimation. SIAM Journal on Applied Mathematics, 22 (2):162–172, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sandberg:1972:EUS

[812] I. W. Sandberg and A. N. Willson, Jr. Existence and uniqueness of solutions for the equations of nonlinear DC networks. SIAM Journal on Applied Mathematics, 22(2):173–186, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Holzmann:1972:TGM

[813] Carlos A. Holzmann and Frank Harary. On the tree graph of a matroid. SIAM Journal on Applied Mathematics, 22 (2):187–193, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Milton:1972:EUR

[814] J. Susan Milton, W. J. Padgett, and Chris P. Tsokos. On the existence and uniqueness of a random solution to a perturbed random integral equation of the Fredholm type. SIAM Journal on Applied Mathematics, 22 (2):194–208, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mortell:1972:PPN

[815] Michael P. Mortell and Brian R. Seymour. Pulse propagation in a nonlinear viscoelastic rod of finite length. SIAM Journal on Applied Mathematics, 22 (2):209–224, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hu:1972:PLB

[816] T. C. Hu and K. C. Tan. Path length of binary search trees. SIAM Journal on Applied Mathematics, 22 (2):225–234, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chikwendu:1972:PMH

[817] S. C. Chikwendu and J. Kevorkian. A perturbation method for hyperbolic equations with small nonlinearities. SIAM Journal on Applied Mathematics, 22(2):235–258, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mason:1972:LTN

[818] J. David Mason. Local theorems for nonidentically distributed lattice random variables. SIAM Journal on Applied Mathematics, 22(2):259–265, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shaw:1972:EAS

[819] Leonard Shaw. Existence and approximation of solutions to an infinite set of linear time-invariant differential equations. SIAM Journal on Applied Mathematics, 22(2):266–279, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Perng:1972:WRP

[820] Daniel Yih Pin Perng. Wave reflection by a periodic medium. SIAM Journal on Applied Mathematics, 22 (2):280–299, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Thompson:1972:CSW

[821] Jon H. Thompson. Closed solutions for wedge diffraction. SIAM Journal on Applied Mathematics, 22(2):300–306, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fujisawa:1972:PLT

[822] T. Fujisawa and E. S. Kuh. Piecewiselinear theory of nonlinear networks. SIAM Journal on Applied Mathematics, 22(2):307–328, March 1972. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hastings:1972:RFS

[823] S. P. Hastings. Reversed flow solutions of the Falkner–Skan equation. SIAM Journal on Applied Mathematics, 22 (2):329–334, March 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pyle:1972:GIL

[824] L. D. Pyle. The generalized inverse in linear programming. Basic structure. SIAM Journal on Applied Mathematics, 22(3):335–355, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sposito:1972:NFL

[825] Vince A. Sposito and H. T. David. A note on Farkas lemmas over cone domains. SIAM Journal on Applied Mathematics, 22(3):356–358, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sweet:1972:PFS

[826] Daniel Sweet. Periodic families of systems possessing a first integral. SIAM Journal on Applied Mathematics, 22 (3):359–372, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Handelsman:1972:ASC

[827] Richard A. Handelsman and W. E. Olmstead. Asymptotic solution to a class of nonlinear Volterra integral equations. SIAM Journal on Applied Mathematics, 22(3):373–384, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schindler:1972:DTW

[828] G. M. Schindler. On N-dimensional traveling waves in a gas with an application to radial flow. SIAM Journal on Applied Mathematics, 22(3):385–405, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sanders:1972:RDS

[829] J. Lyell Sanders, Jr. Random deflections of a string on an elastic foundation. SIAM Journal on Applied Mathematics, 22(3):406–418, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meidan:1972:TVL

[830] Reuven Meidan. Translation varying linear operators. SIAM Journal on Applied Mathematics, 22(3):419–436, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Aronsson:1972:MMS

[831] Gunnar Aronsson. A mathematical model in sand mechanics: Presentation and analysis. SIAM Journal on Applied Mathematics, 22(3):437–458, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matula:1972:CCS

[832] David W. Matula. k-components, clusters, and slicings in graphs. SIAM Journal on Applied Mathematics, 22 (3):459–480, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rahimi:1972:HCM

[833] A. Rahimi and R. W. Brockett. Homotopic classification of minimal periodic realizations of stationary weighting patterns. SIAM Journal on Applied Mathematics, 22(3):481–486, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hendry:1972:NDS

[834] W. L. Hendry. Neutron distribution in a slab reactor with moving boundaries. SIAM Journal on Applied Mathematics, 22(3):487–494, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gautesen:1972:FRT

[835] A. K. Gautesen. A force relationship for two-dimensional viscous flow past a semi-infinite plate and an obstacle. SIAM Journal on Applied Mathematics, 22(3):495–502, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hayes:1972:EPP

[836] John Hayes and Richard Kellner. The eigenvalue problem for a pair of coupled integral equations arising in the numerical solution of Laplace’s equation. SIAM Journal on Applied Mathematics, 22(3):503–513, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ward:1972:WSI

[837] J. F. Ward, T. L. Boullion, and T. O. Lewis. Weak spectral inverses. SIAM Journal on Applied Mathematics, 22(3):514–518, May 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bushard:1972:PSP

[838] Louis B. Bushard. Periodic solutions of perturbed autonomous systems, and locking-in. SIAM Journal on Applied Mathematics, 22(4):519–528, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [988].

Chan:1972:FIB

[839] C. Y. Chan. A first initial boundary value problem for a semilinear heat equation. SIAM Journal on Applied Mathematics, 22(4):529–537, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartwig:1972:RME

[840] Robert E. Hartwig. The resultant and the matrix equation AX = XB. SIAM Journal on Applied Mathematics, 22(4):538–544, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hellman:1972:OSR

[841] Olavi Hellman. On the optimal search for a randomly moving target. SIAM Journal on Applied Mathematics, 22 (4):545–552, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vogt:1972:LFI

[842] Andrew Vogt. On the linearity of form isometries. SIAM Journal on Applied Mathematics, 22(4):553–560, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Komkov:1972:BOD

[843] V. Komkov. On boundedness and oscillation of the differential equation: x′′ + A(t)g(x) = f(t) in R^{n}. SIAM Journal on Applied Mathematics, 22 (4):561–568, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Amin:1972:UBO

[844] A. T. Amin and S. L. Hakimi. Upper bounds on the order of a clique of a graph. SIAM Journal on Applied Mathematics, 22(4):569–573, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lindner:1972:ACR

[845] Charles C. Lindner. An algebraic construction for Room squares. SIAM Journal on Applied Mathematics, 22 (4):574–579, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bowman:1972:PP

[846] V. J. Bowman. Permutation polyhedra. SIAM Journal on Applied Mathematics, 22(4):580–589, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Butterworth:1972:STT

[847] Richard W. Butterworth. A set theoretic treatment of coherent systems. SIAM Journal on Applied Mathematics, 22(4):590–598, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ridenhour:1972:CPF

[848] Jerry R. Ridenhour and Thomas L. Sherman. Conjugate points for fourth order linear differential equations. SIAM Journal on Applied Mathematics, 22(4):599–603, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Katz:1972:TCB

[849] Amnon Katz. Transformations of the compressible boundary layer equations. SIAM Journal on Applied Mathematics, 22(4):604–611, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1972:RCT

[850] A. H. Zemanian. Realizability conditions for time-varying and timeinvariant Hilbert ports. SIAM Journal on Applied Mathematics, 22(4):612– 628, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chow:1972:ASI

[851] Pao-Liu Chow. Asymptotic solutions of inhomogeneous initial boundary value problems for weakly nonlinear partial differential equations. SIAM Journal on Applied Mathematics, 22(4):629– 647, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Utgoff:1972:BSSa

[852] V. A. Utgoff and R. L. Kashyap. On behavior strategy solutions in twoperson zero-sum finite extended games with imperfect information. part I: A method for determination of minimally complex behavior strategy solutions. SIAM Journal on Applied Mathematics, 22(4):648–664, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Utgoff:1972:BSSb

[853] V. A. Utgoff and R. L. Kashyap. On behavior strategy solutions in two-person zero-sum finite extended games with imperfect information. part II: Determination of a minimally complex behavior strategy solution in a medical decision process. SIAM Journal on Applied Mathematics, 22(4):665–675, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bereanu:1972:ERP

[854] B. Bereanu. Errata: “Renewal Processes and Some Stochastic Programming Problems in Economics” [SIAM J. Appl. Math. 19 (1970), no. 2, 308– 322]. SIAM Journal on Applied Mathematics, 22(4):676, June 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [600].

Lientz:1972:PMI

[855] Bennet P. Lientz. Properties of modal intervals. SIAM Journal on Applied Mathematics, 23(1):1–5, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stone:1972:OSP

[856] Lawrence D. Stone, Jerome A. Stanshine, and Carl A. Persinger. Optimal search in the presence of Poissondistributed false targets. SIAM Journal on Applied Mathematics, 23(1):6–27, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Masry:1972:CFU

[857] Elias Masry. On covariance functions of unit processes. SIAM Journal on Applied Mathematics, 23(1):28–33, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kohlberg:1972:NSM

[858] Elon Kohlberg. The nucleolus as a solution of a minimization problem. SIAM Journal on Applied Mathematics, 23(1):34–39, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Saigal:1972:CCP

[859] Romesh Saigal. A characterization of the constant parity property of the number of solutions to the linear complementarity problem. SIAM Journal on Applied Mathematics, 23(1):40–45, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Saigal:1972:CCC

[860] Romesh Saigal. On the class of complementary cones and Lemke’s algorithm. SIAM Journal on Applied Mathematics, 23(1):46–60, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Balas:1972:CCU

[861] Egon Balas and Robert Jeroslow. Canonical cuts on the unit hypercube. SIAM Journal on Applied Mathematics, 23(1):61–69, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

MacCamy:1972:ACF

[862] R. C. MacCamy. Approximations for a class of functional differential equations. SIAM Journal on Applied Mathematics, 23(1):70–83, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nathanson:1972:IBS

[863] Melvyn B. Nathanson. Integrals of binary sequences. SIAM Journal on Applied Mathematics, 23(1):84–86, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Deakin:1972:ASW

[864] A. S. Deakin. Asymptotic solution of the wave equation with variable velocity and boundary conditions. SIAM Journal on Applied Mathematics, 23(1):87–98, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartwig:1972:NJC

[865] Robert E. Hartwig. A note on Jacobson chains. SIAM Journal on Applied Mathematics, 23(1):99–103, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartwig:1972:RS

[866] Robert E. Hartwig. Resultants and the solution of AX − XB = −C. SIAM Journal on Applied Mathematics, 23(1):104–117, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wolfe:1972:DPW

[867] Peter Wolfe. Diffraction of plane waves by a strip; exact and asymptotic solutions. SIAM Journal on Applied Mathematics, 23(1):118–132, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wu:1972:ASD

[868] Chien-Heng Wu and Daniel Y. P. Perng. On the asymptotically spherical deformations of arbitrary membranes of revolution fixed along an edge and inflated by large pressures — A nonlinear boundary layer phenomenon. SIAM Journal on Applied Mathematics, 23(1):133–152, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fink:1972:UEO

[869] James P. Fink. Uniformly elliptic operators and Green’s functions. SIAM Journal on Applied Mathematics, 23 (1):153–155, July 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Halfin:1972:ACC

[870] S. Halfin. Arbitrarily complex corner polyhedra are dense in R^{n}. SIAM Journal on Applied Mathematics, 23 (2):157–163, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Beutler:1972:ASE

[871] Frederick J. Beutler. Almost sure exponential bounds for stochastic operator systems, with applications to randomly sampled control. SIAM Journal on Applied Mathematics, 23(2): 164–170, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Poole:1972:WSI

[872] G. D. Poole and T. L. Boullion. Weak spectral inverses which are partial isometries. SIAM Journal on Applied Mathematics, 23(2):171–172, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garey:1972:OBI

[873] M. R. Garey. Optimal binary identification procedures. SIAM Journal on Applied Mathematics, 23(2): 173–186, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1972:LTP

[874] Kuang ho Chen. A liouville-type problem in partial differential equations. SIAM Journal on Applied Mathematics, 23(2):187–188, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Horn:1972:SMJ

[875] W. A. Horn. Single-machine job sequencing with treelike precedence ordering and linear delay penalties. SIAM Journal on Applied Mathematics, 23(2):189–202, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1972:SPO

[876] Richard H. Smith, Jr. and V. David Vandelinde. A saddle-point optimality criterion for nonconvex programming in normed spaces. SIAM Journal on Applied Mathematics, 23(2): 203–213, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Skala:1972:VFT

[877] H. L. Skala. A variation of the friendship theorem. SIAM Journal on Applied Mathematics, 23(2):214–220, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Flemming:1972:GWE

[878] D. P. Flemming. Generalizations of the wave equation which allow weak and strong discontinuities. SIAM Journal on Applied Mathematics, 23(2): 221–224, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vujicic:1972:CFM

[879] Milan Vujicic, Fedor Herbut, and Gradimir Vujicic. Canonical form for matrices under unitary congruence transformations. I: Conjugate-normal matrices. SIAM Journal on Applied Mathematics, 23(2):225–238, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Consul:1972:ULE

[880] P. C. Consul and L. R. Shenton. Use of Lagrange expansion for generating discrete generalized probability distributions. SIAM Journal on Applied Mathematics, 23(2):239–248, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Althaus:1972:SDC

[881] H. L. Althaus. State diagrams of T - cyclic autonomous nonsingular LSC’s over J_{m} with multiply irreducible elementary divisors. SIAM Journal on Applied Mathematics, 23(2):249–258, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Guderley:1972:EPT

[882] K. G. Guderley. An extremum principle for three-dimensional compressible inviscid flows. SIAM Journal on Applied Mathematics, 23(2):259–275, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Aneja:1972:RWD

[883] K. G. Aneja and Kanwar Sen. Random walk and distributions of rank order statistics. SIAM Journal on Applied Mathematics, 23(2):276–287, September 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Justice:1972:AIP

[884] James H. Justice. An algorithm for inverting positive definite Toeplitz matrices. SIAM Journal on Applied Mathematics, 23(3):289–291, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Helton:1972:CLP

[885] J. W. Helton and A. H. Zemanian. The cascade loading of passive Hilbert ports. SIAM Journal on Applied Mathematics, 23(3):292–306, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Spjotvoll:1972:NTF

[886] Emil Spjotvoll. A note on a theorem of forsythe and golub. SIAM Journal on Applied Mathematics, 23 (3):307–311, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Blake:1972:DPG

[887] Ian F. Blake. Distance properties of group codes for the Gaussian channel. SIAM Journal on Applied Mathematics, 23(3):312–324, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ranger:1972:ASE

[888] K. B. Ranger. The applicability of Stokes expansions to reversed flow.

SIAM Journal on Applied Mathematics, 23(3):325–333, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kreisselmeier:1972:SBM

[889] G. Kreisselmeier. A solution of the bilinear matrix equation AY + Y B = −Q. SIAM Journal on Applied Mathematics, 23(3):334–338, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chinnaswamy:1972:UAT

[890] A. V. Chinnaswamy and R. P. Kanwal. Uniform asymptotic theory of diffraction by a plane screen by the method of boundary layers. SIAM Journal on Applied Mathematics, 23 (3):339–355, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Halfin:1972:SSD

[891] S. Halfin. Steady-state distribution for the buffer content of an M/G/1 queue with varying service rate. SIAM Journal on Applied Mathematics, 23 (3):356–363, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shannon:1972:IFA

[892] A. G. Shannon. Iterative formulas associated with generalized third order recurrence relations. SIAM Journal on Applied Mathematics, 23(3): 364–368, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Halfin:1972:PQM

[893] S. Halfin and M. Segal. A priority queueing model for a mixture of two types of customers. SIAM Journal on Applied Mathematics, 23(3): 369–379, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Curl:1972:AII

[894] M. L. Curl and W. P. Graebel. Application of invariant imbedding techniques to flow instability problems. SIAM Journal on Applied Mathematics, 23 (3):380–395, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Edelstein:1972:PBD

[895] Warren S. Edelstein. A priori bounds for the displacement problem in viscoelasticity. SIAM Journal on Applied Mathematics, 23(3):396–408, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tam:1972:EAS

[896] K. K. Tam. Erratum: “On the Asymptotic Solution of Viscous Incompressible Flow Past a Heated Paraboloid of Revolution” [SIAM J. Appl. Math. 20 (1971), no. 4, 714–721]. SIAM Journal on Applied Mathematics, 23 (3):409, November 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [719].

Chen:1972:IAN

[897] Y. M. Chen and J. S. Chang. An iterative approach to nonlinear dynamic stability problems. SIAM Journal on Applied Mathematics, 23(4):

411–419, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wilson:1972:FMS

[898] G. Tunnicliffe Wilson. The factorization of matricial spectral densities. SIAM Journal on Applied Mathematics, 23(4):420–426, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Davis:1972:ROP

[899] Paul Wm. Davis and Bernard A. Fleishman. Representation of oscillations in piecewise-linear systems by the phase-shift-averaging method. SIAM Journal on Applied Mathematics, 23 (4):427–438, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Longman:1972:ALT

[900] I. M. Longman. Approximate Laplace transform inversion applied to a problem in electrical network theory. SIAM Journal on Applied Mathematics, 23 (4):439–445, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Flaherty:1972:PBB

[901] Joseph E. Flaherty, Joseph B. Keller, and S. I. Rubinow. Post buckling behavior of elastic tubes and rings with opposite sides in contact. SIAM Journal on Applied Mathematics, 23(4): 446–455, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chiou:1972:NTS

[902] Kuo-Liang Chiou. A nonoscillation theorem for the superlinear case of second order differential equations y′′ + yF (y^{2}, x) = 0. SIAM Journal on Applied Mathematics, 23(4):456–459, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bowman:1972:SCG

[903] V. J. Bowman. A structural comparison of Gomory’s fractional cutting planes and Hermitian basic solutions. SIAM Journal on Applied Mathematics, 23(4):460–462, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bergman:1972:TTO

[904] Lawrence G. Bergman and C. Dennis Weiss. A transformation technique for obtaining least cost solutions to covering problems. SIAM Journal on Applied Mathematics, 23 (4):463–476, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Troesch:1972:CIF

[905] B. Andreas Troesch and Patrick D. Weidman. Containers with isochronous fluid oscillations. SIAM Journal on Applied Mathematics, 23(4):477–489, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Corley:1972:DRP

[906] H. W. Corley, Jr. and S. D. Roberts. Duality relationships for a partitioning problem. SIAM Journal on Applied Mathematics, 23(4):490–494, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Alam:1972:AVM

[907] Khursheed Alam. Asymptotic value of the mean of a function of a normal random variable. SIAM Journal on Applied Mathematics, 23(4): 495–498, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Padgett:1972:SID

[908] W. J. Padgett and Chris P. Tsokos. On a stochastic integro-differential equation of Volterra type. SIAM Journal on Applied Mathematics, 23(4): 499–512, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gal:1972:MMS

[909] Shmuel Gal. Multidimensional minimax search for a maximum. SIAM Journal on Applied Mathematics, 23 (4):513–526, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morris:1972:ACG

[910] A. J. Morris. Approximation and complementary geometric programming. SIAM Journal on Applied Mathematics, 23(4):527–531, December 1972. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barnett:1973:MDE

[911] S. Barnett. Matrix differential equations and Kronecker products. SIAM Journal on Applied Mathematics, 24 (1):1–5, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartfiel:1973:BEE

[912] D. J. Hartfiel. Bounds for eigenvalues and eigenvectors of a nonnegative matrix which involve a measure of irreducibility. SIAM Journal on Applied Mathematics, 24(1):6–8, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Foschini:1973:ATT

[913] Gerard J. Foschini. Applications of a theorem of Titchmarsh to bandlimit irreducibility of memoryless transformations. SIAM Journal on Applied Mathematics, 24(1):9–18, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mitchell:1973:OSR

[914] Bill Mitchell. Optimal service-rate selection in an M|G|Î queue. SIAM Journal on Applied Mathematics, 24 (1):19–35, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rodgers:1973:DSPa

[915] L. C. Rodgers and W. H. Warner. Dynamic stability of out-of-plane motion of curved elastic rods. SIAM Journal on Applied Mathematics, 24 (1):36–43, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lovass-Nagy:1973:RBM

[916] V. Lovass-Nagy and D. L. Powers. A relation between the Moore–Penrose and commuting reciprocal inverses. SIAM Journal on Applied Mathematics, 24 (1):44–49, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Thurston:1973:GAS

[917] Louis H. Thurston and James S. W. Wong. On global asymptotic stability of certain second order differential equations with integrable forcing terms. SIAM Journal on Applied Mathematics, 24(1):50–61, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yee:1973:ESI

[918] Kane S. Yee. Explicit solutions for the initial boundary value problem of a system of lossless transmission lines. SIAM Journal on Applied Mathematics, 24(1):62–80, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bassanini:1973:VOF

[919] Piero Bassanini. On a variational operator in fluid mechanics. SIAM Journal on Applied Mathematics, 24 (1):81–87, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tietavainen:1973:NPC

[920] Aimo Tietavainen. On the nonexistence of perfect codes over finite fields. SIAM Journal on Applied Mathematics, 24 (1):88–96, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berger:1973:ESD

[921] Neil Berger. Estimates for stress derivatives and error in interior equations for shells of variable thickness with applied forces. SIAM Journal on Applied Mathematics, 24(1):97–120, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shirakawa:1973:SOO

[922] Isao Shirakawa, Akio Sakamoto, and Hiroshi Ozaki. Sparsity-oriented ordering of pivotal operations on network equations. SIAM Journal on Applied Mathematics, 24(1):121–129, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Genin:1973:PBP

[923] Joseph Genin and Russell Wilhelmsen. Postbuckling behavior of a parametrically excited nonlinear column. SIAM Journal on Applied Mathematics, 24 (1):130–133, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reed:1973:BDC

[924] I. S. Reed. Boolean difference calculus and fault finding. SIAM Journal on Applied Mathematics, 24(1):134– 143, January 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Engelhardt:1973:UBV

[925] Max Engelhardt. On upper bounds for variances in stochastic approximation. SIAM Journal on Applied Mathematics, 24(2):145–151, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bigelow:1973:OPL

[926] J. H. Bigelow and N. Z. Shapiro. Optimization problems with large parameters. SIAM Journal on Applied Mathematics, 24(2):152–163, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mallows:1973:SSF

[927] C. L. Mallows. Sequential sampling of finite populations with and without replacement. SIAM Journal on Applied Mathematics, 24(2):164–168, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kuczura:1973:PMP

[928] A. Kuczura. Piecewise Markov processes. SIAM Journal on Applied Mathematics, 24(2):169–181, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Albert:1973:GMT

[929] Arthur Albert. The Gauss–Markov theorem for regression models with possibly singular covariances. SIAM Journal on Applied Mathematics, 24 (2):182–187, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

DeSantis:1973:TRP

[930] R. M. DeSantis and W. A. Porter. On time-related properties of nonlinear systems. SIAM Journal on Applied Mathematics, 24(2):188–206, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tyagi:1973:SPT

[931] M. S. Tyagi. On some properties of n-tournaments. SIAM Journal on Applied Mathematics, 24(2):207–214, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Flaherty:1973:CPI

[932] Joseph E. Flaherty and Joseph B. Keller. Contact problems involving a buckled elastica. SIAM Journal on Applied Mathematics, 24(2):215–225, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huilgol:1973:UNP

[933] R. R. Huilgol. On uniqueness and nonuniqueness in the plane creeping flows of second order fluids. SIAM Journal on Applied Mathematics, 24 (2):226–233, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Akaike:1973:BTM

[934] Hirotugu Akaike. Block Toeplitz matrix inversion. SIAM Journal on Applied Mathematics, 24(2):234–241, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arnold:1973:SCE

[935] Barry C. Arnold. Some characterizations of the exponential distribution by geometric compounding. SIAM Journal on Applied Mathematics, 24 (2):242–244, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Marshall:1973:LBR

[936] Kneale T. Marshall. Linear bounds on the renewal function. SIAM Journal on Applied Mathematics, 24(2):245–250, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Benson:1973:COT

[937] Donald C. Benson. Comparison and oscillation theory for Lienard’s equation with positive damping. SIAM Journal on Applied Mathematics, 24 (2):251–271, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rao:1973:SMR

[938] J. V. Venugopala Rao. Some more representations for the generalized inverse of a partitioned matrix. SIAM Journal on Applied Mathematics, 24 (2):272–276, March 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Blair:1973:BCS

[939] J. J. Blair. Bounds for the change in the solutions of second order elliptic PDE’s when the boundary is perturbed. SIAM Journal on Applied Mathematics, 24(3):277–285, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosen:1973:IGE

[940] Gerald Rosen and G. W. Ullrich. Invariance group of the equation ∂u/∂t = −u · Δu. SIAM Journal on Applied Mathematics, 24(3):286–288, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wyner:1973:BND

[941] A. D. Wyner. A bound on the number of distinguishable functions which are time-limited and approximately band-limited. SIAM Journal on Applied Mathematics, 24(3):289–297, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cannon:1973:DUC

[942] J. R. Cannon and Paul Duchateau. Determining unknown coefficients in a nonlinear heat conduction problem. SIAM Journal on Applied Mathematics, 24(3):298–314, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1973:GIM

[943] Carl D. Meyer, Jr. Generalized inversion of modified matrices. SIAM Journal on Applied Mathematics, 24 (3):315–323, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nowinski:1973:CSI

[944] J. L. Nowinski. Cauchy–Schwarz inequality and the evaluation of torsional rigidity of anisotropic bars. SIAM Journal on Applied Mathematics, 24 (3):324–331, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Clements:1973:TSA

[945] D. L. Clements. Thermal stress in an anisotropic elastic half-space. SIAM Journal on Applied Mathematics, 24 (3):332–337, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cline:1973:GIL

[946] R. E. Cline and L. D. Pyle. The generalized inverse in linear programming — an intersection projection method and the solution of a class of structured linear programming problems. SIAM Journal on Applied Mathematics, 24(3):338–351, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kogelman:1973:ATN

[947] Stanley Kogelman and Joseph B. Keller. Asymptotic theory of nonlinear wave propagation. SIAM Journal on Applied Mathematics, 24(3):352–361, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Frisch:1973:BOD

[948] R. A. Frisch and B. Ru-Shao Cheo. On a bounded one-dimensional poissonvlasov system. SIAM Journal on Applied Mathematics, 24(3):362–368, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rodgers:1973:DSPb

[949] L. C. Rodgers and W. H. Warner. Dynamic stability of planar rods of noncircular centerline. SIAM Journal on Applied Mathematics, 24(3):369–382, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Licari:1973:DDE

[950] J. P. Licari and W. H. Warner. Domain dependence of eigenvalues of vibrating plates. SIAM Journal on Applied Mathematics, 24(3):383–395, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ra:1973:RTB

[951] J. W. Ra, H. L. Bertoni, and L. B. Felsen. Reflection and transmission of beams at a dielectric interface. SIAM Journal on Applied Mathematics, 24(3):396–413, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Oliveira:1973:MPE

[952] G. N. de Oliveira. Matrices with prescribed entries and eigenvalues. II. SIAM Journal on Applied Mathematics, 24(3):414–417, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eaves:1973:PGJ

[953] B. Curtis Eaves. Polymatrix games with joint constraints. SIAM Journal on Applied Mathematics, 24(3):418– 423, May 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klaasen:1973:LUE

[954] Gene A. Klaasen. Local uniqueness and existence of solutions of a threepoint boundary value problem. SIAM Journal on Applied Mathematics, 24 (4):425–428, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Longman:1973:GRF

[955] I. M. Longman. On the generation of rational function approximations for Laplace transform inversion with an application to viscoelasticity. SIAM Journal on Applied Mathematics, 24(4):429–440, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bhargava:1973:NMW

[956] S. Bhargava and R. J. Duffin. On the nonlinear method of wilkins for cooling fin optimization. SIAM Journal on Applied Mathematics, 24(4):441–448, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Soong:1973:SCR

[957] T. T. Soong and S. N. Chuang. Solutions of a class of random differential equations. SIAM Journal on Applied Mathematics, 24(4):449–459, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burniston:1973:EAS

[958] E. E. Burniston and C. E. Siewert. Exact analytical solutions of the transcendental equation α sin ζ = ζ. SIAM Journal on Applied Mathematics, 24(4):460–466, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dirickx:1973:TTD

[959] Yvo M. I. Dirickx. Turnpike theory in deterministic discrete dynamic programming with discount factor greater than one. SIAM Journal on Applied Mathematics, 24(4):467–472, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rao:1973:TAC

[960] C. Radhakrishna Rao and Sujit Kumar Mitra. Theory and application of constrained inverse of matrices. SIAM Journal on Applied Mathematics, 24 (4):473–488, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bouldin:1973:PIP

[961] Richard Bouldin. The pseudo-inverse of a product. SIAM Journal on Applied Mathematics, 24(4):489–495, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lindstrom:1973:MTK

[962] F. T. Lindstrom and M. N. L. Narasimham. Mathematical theory of a kinetic model for dispersion of previously distributed chemicals in a sorbing porous medium. SIAM Journal on Applied Mathematics, 24(4):496–510, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pyle:1973:GIL

[963] L. D. Pyle and R. E. Cline. The generalized inverse in linear programming — interior gradient projection methods. SIAM Journal on Applied Mathematics, 24(4):511–534, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jackson:1973:USB

[964] Lloyd K. Jackson. Uniqueness of solutions of boundary value problems for ordinary differential equations. SIAM Journal on Applied Mathematics, 24 (4):535–538, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Johnson:1973:IRD

[965] Edward R. Johnson and Raymond Parnes. An integral representation of the Dirac delta function for axisymmetric boundary value problems. SIAM Journal on Applied Mathematics, 24(4):539–544, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ford:1973:CPB

[966] Wayne T. Ford. Convexity in phase behavior. SIAM Journal on Applied Mathematics, 24(4):545–551, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pye:1973:PCM

[967] W. C. Pye, T. L. Boullion, and T. A. Atchison. The pseudoinverse of a composite matrix of circulants. SIAM Journal on Applied Mathematics, 24 (4):552–555, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ranger:1973:NSM

[968] K. B. Ranger. A note on some mixed boundary value problems for the heat equation. SIAM Journal on Applied Mathematics, 24(4):556–561, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Balachandra:1973:PSS

[969] M. Balachandra. Periodic solutions of singularly perturbed equations arising from gyroscopic systems. SIAM Journal on Applied Mathematics, 24 (4):562–574, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fink:1973:PES

[970] James P. Fink, William S. Hall, and Siamak Khalili. Perturbation expansions for some nonlinear wave equations. SIAM Journal on Applied Mathematics, 24(4):575–595, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cannon:1973:OST

[971] J. R. Cannon and R. E. Klein. On the observability and stability of the temperature distribution in a composite heat conductor. SIAM Journal on Applied Mathematics, 24(4):596–602, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Craven:1973:TTC

[972] B. D. Craven and B. Mond. Transportation theorems for cone-convex functions. SIAM Journal on Applied Mathematics, 24(4):603–612, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wolfe:1973:ADP

[973] Peter Wolfe. Addendum to: Diffraction of plane waves by a strip; exact and asymptotic solutions — the solution in the general case. SIAM Journal on Applied Mathematics, 24 (4):613–618, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nunziato:1973:NUL

[974] Jace W. Nunziato. A note on uniqueness in the linear theory of heat conduction with finite wave speeds. SIAM Journal on Applied Mathematics, 25 (1):1–4, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mitra:1973:MSC

[975] Debasis Mitra and Hing C. So. P matrices and solutions to a class of linear inequalities. SIAM Journal on Applied Mathematics, 25(1):5–15, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bebernes:1973:SBV

[976] J. W. Bebernes and W. Kelley. Some boundary value problems for generalized differential equations. SIAM Journal on Applied Mathematics, 25(1):16– 23, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Willett:1973:IS

[977] Michael Willett. The index of an Msequence. SIAM Journal on Applied Mathematics, 25(1):24–27, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Skinner:1973:STS

[978] L. A. Skinner. “shooting technique” for singular perturbations. SIAM Journal on Applied Mathematics, 25(1):28–31, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chang:1973:DFE

[979] Shih-Chih Chang. The distribution of free electrical charge on the surface of a conducting body — the Robin problem. SIAM Journal on Applied Mathematics, 25(1):32–36, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baker:1973:OTS

[980] John W. Baker. Oscillation theorems for a second order damped nonlinear differential equation. SIAM Journal on Applied Mathematics, 25(1):37–40, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Luce:1973:TAS

[981] R. Duncan Luce. Three axiom systems for additive semiordered structures. SIAM Journal on Applied Mathematics, 25(1):41–53, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ting:1973:TRB

[982] Tsuan Wu Ting. Torsional rigidities for bars under fully plastic torsion. SIAM Journal on Applied Mathematics, 25(1):54–68, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sigillito:1973:PIN

[983] V. G. Sigillito and J. C. Pirkle, Jr. A priori inequalities and norm error bounds for solutions of a third order diffusion-like equation. SIAM Journal on Applied Mathematics, 25(1):69–71, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Richards:1973:FMD

[984] J. A. Richards and D. J. Miller. Features of mode diagrams for l th order periodic systems. SIAM Journal on Applied Mathematics, 25(1):72–82, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hu:1973:NPA

[985] T. C. Hu. A new proof of the T -C algorithm. SIAM Journal on Applied Mathematics, 25(1):83–94, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Karlin:1973:FLW

[986] Maurice Karlin and F. J. MacWilliams. On finding low weight vectors in quadratic residue codes for p = 8m−1. SIAM Journal on Applied Mathematics, 25(1):95–104, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kassoy:1973:SBL

[987] D. R. Kassoy. The singularity at boundary layer separation due to mass injection. SIAM Journal on Applied Mathematics, 25(1):105–123, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bushard:1973:EPS

[988] Louis B. Bushard. Erratum: “Periodic Solutions of Perturbed Autonomous Systems, and Locking-In” [SIAM J. Appl. Math. 22 (1972), no. 4, 519– 528]. SIAM Journal on Applied Mathematics, 25(1):124, July 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [838].

Sharma:1973:ITG

[989] Bhu Dev Sharma and Ram Autar. An inversion theorem and generalized entropies for continuous distributions.

SIAM Journal on Applied Mathematics, 25(2):125–132, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [1156].

Shanholt:1973:SSP

[990] Gerald A. Shanholt. Stability and small parameter perturbations of a critical neutral functional differential equation. SIAM Journal on Applied Mathematics, 25(2):133–141, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tosovic:1973:SES

[991] Ljubomir B. Tosovic. Some experiments on sparse sets of linear equations. SIAM Journal on Applied Mathematics, 25(2):142–148, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Richter:1973:CDO

[992] G. Richter. On the convergence of the discrete ordinate method in kinetic theory. SIAM Journal on Applied Mathematics, 25(2):149–157, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1973:GIT

[993] Chi-Tsong Chen. A generalization of the inertia theorem. SIAM Journal on Applied Mathematics, 25(2): 158–161, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kolman:1973:PLA

[994] Bernard Kolman. Preface (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25(2):163–168, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wigner:1973:RDP

[995] Eugene P. Wigner. Reduction of direct products and restriction of representations to subgroups: The everyday tasks of the quantum theorists (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25 (2):169–185, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Judd:1973:LGA

[996] B. R. Judd. Lie groups in atomic and molecular spectroscopy (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25(2):186–192, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moshinsky:1973:CTQ

[997] Marcos Moshinsky. Canonical transformations and quantum mechanics (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25(2): 193–212, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brockett:1973:LTC

[998] R. W. Brockett. Lie theory and control systems defined on spheres (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25(2):213–225, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miller:1973:LTG

[999] Willard Miller, Jr. Lie theory and generalizations of the hypergeometric functions (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25 (2):226–235, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Koornwinder:1973:AFJ

[1000] Tom Koornwinder. The addition formula for Jacobi polynomials and spherical harmonics (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25(2):236–246, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barut:1973:SUA

[1001] A. O. Barut. Some unusual applications of Lie algebra representations in quantum theory (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25(2):247–259, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Belinfante:1973:LAI

[1002] Johan G. F. Belinfante. Lie algebras and inhomogeneous simple materials (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25 (2):260–268, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gruber:1973:PLR

[1003] Bruno Gruber. Properties of linear representations of semisimple Lie algebras (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25 (2):269–286, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McConnell:1973:RRG

[1004] J. McConnell. Reduction of representations of the general linear group using Lie algebras (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25(2):287–299, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kolman:1973:CLAa

[1005] Bernard Kolman and Robert E. Beck. Computers in Lie algebras. I: Calculation of inner multiplicities (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25(2):300–312, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kolman:1973:CLAb

[1006] Bernard Kolman and Robert E. Beck.

Computers in Lie algebras. II: Calculation of outer multiplicities (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25(2):313–323, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Diaz:1973:ISP

[1007] J. B. Diaz. The inequality of schwarz, and pointwise upper and lower bounds for the Dirichlet problem. SIAM Journal on Applied Mathematics, 25(3): 325–334, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Synge:1973:LHO

[1008] John L. Synge. Linked harmonic oscillators. SIAM Journal on Applied Mathematics, 25(3):335–345, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Newell:1973:SLT

[1009] G. F. Newell. Scheduling, location, transportation, and continuum mechanics; some simple approximations to optimization problems. SIAM Journal on Applied Mathematics, 25(3): 346–360, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Plaut:1973:EMW

[1010] R. H. Plaut. Elastic minimum-weight design for specified critical load. SIAM Journal on Applied Mathematics, 25 (3):361–371, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lin:1973:NSL

[1011] C. C. Lin. Note on stretching of line and surface elements in isotropic turbulence. SIAM Journal on Applied Mathematics, 25(3):372–377, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Budiansky:1973:PW

[1012] Bernard Budiansky and George F. Carrier. The pointless wedge. SIAM Journal on Applied Mathematics, 25 (3):378–387, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1973:SCN

[1013] R. E. Meyer and Lynn O. Wilson. Stretching and calculus of negligibility. SIAM Journal on Applied Mathematics, 25(3):388–402, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Adolphson:1973:OLO

[1014] D. Adolphson and T. C. Hu. Optimal linear ordering. SIAM Journal on Applied Mathematics, 25(3): 403–423, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Koiter:1973:PSC

[1015] W. T. Koiter. On the principle of stationary complementary energy in the nonlinear theory of elasticity. SIAM Journal on Applied Mathematics, 25 (3):424–434, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hodge:1973:CSE

[1016] Philip G. Hodge, Jr. Complete solutions for elastic-plastic trusses. SIAM Journal on Applied Mathematics, 25 (3):435–447, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hill:1973:EPS

[1017] R. Hill and J. R. Rice. Elastic potentials and the structure of inelastic constitutive laws. SIAM Journal on Applied Mathematics, 25(3): 448–461, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Symonds:1973:ATI

[1018] P. S. Symonds. Approximation techniques for impulsively loaded structures of rate sensitive plastic behavior. SIAM Journal on Applied Mathematics, 25(3):462–473, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gunnlaugsson:1973:OCF

[1019] Geir A. Gunnlaugsson and John B.

Martin. Optimality conditions for fully stressed designs. SIAM Journal on Applied Mathematics, 25(3): 474–482, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shield:1973:RAL

[1020] R. T. Shield. The rotation associated with large strains. SIAM Journal on Applied Mathematics, 25(3): 483–491, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lee:1973:WCM

[1021] E. H. Lee and Wei H. Yang. On waves in composite materials with periodic structure. SIAM Journal on Applied Mathematics, 25(3):492–499, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gurkok:1973:EGC

[1022] Ayfer Gurkok and H. G. Hopkins. The effect of geometry changes on the load carrying capacity of beams under transverse load. SIAM Journal on Applied Mathematics, 25(3):500–521, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Onat:1973:RCR

[1023] E. T. Onat and F. Fardshisheh. Representation of creep, rate sensitivity and plasticity. SIAM Journal on Applied Mathematics, 25(3):522–538, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nachbar:1973:QEP

[1024] W. Nachbar and J. M. Jong. Quadratic eigenvalue problems from a nonlinear structural analysis. SIAM Journal on Applied Mathematics, 25(3): 539–555, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Germain:1973:MVP

[1025] P. Germain. The method of virtual power in continuum mechanics. part 2: Microstructure. SIAM Journal on Applied Mathematics, 25(3): 556–575, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nagtegaal:1973:BOW

[1026] J. C. Nagtegaal. Bounds on the optimal weight of structures under compliance constraints. SIAM Journal on Applied Mathematics, 25(3):576–581, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keller:1973:UAS

[1027] Joseph B. Keller and Daljit S.

Ahluwalia. Uniform asymptotic solution of eigenvalue problems for convex plane domains. SIAM Journal on Applied Mathematics, 25(4):583–591, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zachmanoglou:1973:CTS

[1028] E. C. Zachmanoglou. Caratheodory’s theorem on the second law of thermodynamics. SIAM Journal on Applied Mathematics, 25(4):592–596, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1973:GIR

[1029] Carl D. Meyer, Jr. Generalized inverses and ranks of block matrices. SIAM Journal on Applied Mathematics, 25 (4):597–602, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jenkins:1973:LHC

[1030] James T. Jenkins. Linearized hydrostatics for cholesteric liquid crystals. SIAM Journal on Applied Mathematics, 25(4):603–612, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dickinson:1973:MIC

[1031] B. W. Dickinson. Matricial indices and controllability of linear differential systems. SIAM Journal on Applied Mathematics, 25(4):613–617, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Holzmann:1973:GRM

[1032] C. A. Holzmann, P. G. Norton, and M. D. Tobey. A graphical representation of matroids. SIAM Journal on Applied Mathematics, 25(4): 618–627, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Masry:1973:RSP

[1033] Elias Masry and Stamatis Cambanis. The representation of stochastic processes without loss of information. SIAM Journal on Applied Mathematics, 25(4):628–633, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bandle:1973:BFI

[1034] Catherine Bandle. Bounds for the frequencies of an inhomogeneous string. SIAM Journal on Applied Mathematics, 25(4):634–639, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1973:MSP

[1035] Donald S. Cohen. Multiple solutions and periodic oscillations in nonlinear diffusion processes. SIAM Journal on Applied Mathematics, 25(4): 640–654, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Messiter:1973:MLB

[1036] A. F. Messiter and R. L. Enlow. A model for laminar boundary-layer flow near a separation point. SIAM Journal on Applied Mathematics, 25(4): 655–670, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mitra:1973:ULS

[1037] Sujit Kumar Mitra. Unified least squares approach to linear estimation in a general Gauss–Markov model. SIAM Journal on Applied Mathematics, 25(4):671–680, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nashed:1973:MGI

[1038] M. Z. Nashed and H. Salehi. Measurability of generalized inverses of random linear operators. SIAM Journal on Applied Mathematics, 25(4): 681–692, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ludford:1973:FDR

[1039] G. S. S. Ludford and R. A. Robertson. Fully diffused regions. SIAM Journal on Applied Mathematics, 25 (4):693–703, December 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barker:1974:NML

[1040] George Phillip Barker. Normal matrices and the Lyapunov equation. SIAM Journal on Applied Mathematics, 26(1):1–4, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1974:BBW

[1041] Ronald Smith. Bouncing ball waves. SIAM Journal on Applied Mathematics, 26(1):5–14, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kucera:1974:ME

[1042] Vladimir Kucera. The matrix equation AX + XB = C. SIAM Journal on Applied Mathematics, 26(1):15–25, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fink:1974:CAP

[1043] A. M. Fink. Convergence and almost periodicity of solutions of forced Lienard equations. SIAM Journal on Applied Mathematics, 26(1):26–34, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gordon:1974:PCE

[1044] Paul Gordon. Paths connecting elementary critical points of dynamical systems. SIAM Journal on Applied Mathematics, 26(1):35–102, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kundu:1974:DRT

[1045] Sukhamay Kundu. Disjoint representation of tree realizable sequences. SIAM Journal on Applied Mathematics, 26 (1):103–107, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ahluwalia:1974:EWP

[1046] Daljit S. Ahluwalia, Joseph B. Keller, and Richard Jarvis. Elastic waves produced by surface displacements. SIAM Journal on Applied Mathematics, 26 (1):108–119, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bowman:1974:GSS

[1047] V. J. Bowman and Claude-Alain Burdet. On the general solution to systems of mixed-integer linear equations. SIAM Journal on Applied Mathematics, 26(1):120–125, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Osteen:1974:CDA

[1048] Robert E. Osteen. Clique detection algorithms based on line addition and line removal. SIAM Journal on Applied Mathematics, 26(1):126–135, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Habets:1974:CTS

[1049] Patrick Habets. A consistency theory of singular perturbations of differential equations. SIAM Journal on Applied Mathematics, 26(1):136–153, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hakimi:1974:EGP

[1050] S. L. Hakimi. On the existence of graphs with prescribed degrees and connectivity. SIAM Journal on Applied Mathematics, 26(1):154–164, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chow:1974:LFS

[1051] Shui-Nee Chow and D. R. Dunninger. Lyapunov functions satisfying V̈ > 0. SIAM Journal on Applied Mathematics, 26(1):165–168, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carlson:1974:GSC

[1052] David Carlson, Emile Haynsworth, and Thomas Markham. A generalization of the Schur complement by means of the Moore–Penrose inverse. SIAM Journal on Applied Mathematics, 26 (1):169–175, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kadota:1974:ISS

[1053] T. T. Kadota. On the information stability of stationary ergodic processes. SIAM Journal on Applied Mathematics, 26(1):176–182, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rogers:1974:NMH

[1054] Colin Rogers and John G. Kingston. Nondissipative magneto-hydrodynamic flows with magnetic and velocity field lines orthogonal geodesics. SIAM Journal on Applied Mathematics, 26 (1):183–195, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Waid:1974:SMP

[1055] Margaret C. Waid. Strong maximum principles for time degenerate parabolic operators. SIAM Journal on Applied Mathematics, 26(1):196–202, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hurd:1974:SPR

[1056] H. L. Hurd. Stationarizing properties of random shifts. SIAM Journal on Applied Mathematics, 26(1):203–212, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pipkin:1974:ESP

[1057] A. C. Pipkin and V. M. Sanchez. Existence of solutions of plane traction problems for ideal composites. SIAM Journal on Applied Mathematics, 26 (1):213–220, January 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosen:1974:ASG

[1058] Gerald Rosen. Approximate solution to the generic initial value problem for nonlinear reaction-diffusion equations. SIAM Journal on Applied Mathematics, 26(2):221–224, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ewbank:1974:MSD

[1059] James B. Ewbank, Bob L. Foote, and Hillel J. Kumin. A method for the solution of the distribution problem of stochastic linear programming. SIAM Journal on Applied Mathematics, 26 (2):225–238, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peleg:1974:CPO

[1060] Bezalel Peleg. On competitive prices for optimal consumption plans. SIAM Journal on Applied Mathematics, 26 (2):239–253, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burns:1974:GIF

[1061] Fennell Burns, David Carlson, Emilie Haynsworth, and Thomas Markham. Generalized inverse formulas using the Schur complement. SIAM Journal on Applied Mathematics, 26(2):254–259, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Roberts:1974:LET

[1062] Fred S. Roberts. Laws of exchange and their applications. SIAM Journal on Applied Mathematics, 26(2):260–284, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1974:BRF

[1063] C. L. Chen, R. T. Chien, and C. K. Liu. On the binary representation form of certain integers. SIAM Journal on Applied Mathematics, 26(2):285–293, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartfiel:1974:CSV

[1064] D. J. Hartfiel and G. L. Curry. Concerning the solution vector for parametric programming problems. SIAM Journal on Applied Mathematics, 26 (2):294–296, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartfiel:1974:IPN

[1065] D. J. Hartfiel. On infinite products of nonnegative matrices. SIAM Journal on Applied Mathematics, 26 (2):297–301, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lambe:1974:BNF

[1066] Thomas A. Lambe. Bounds on the number of feasible solutions to a knapsack problem. SIAM Journal on Applied Mathematics, 26(2):302–305, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Felsen:1974:DPF

[1067] L. B. Felsen. Diffraction of the pulsed field from an arbitrarily oriented electric or magnetic dipole by a perfectly conducting wedge. SIAM Journal on Applied Mathematics, 26(2):306–312, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:1974:NEC

[1068] D. L. Wang and D. J. Kleitman. A note on n-edge-connectivity. SIAM Journal on Applied Mathematics, 26 (2):313–314, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Desoer:1974:NMN

[1069] Charles A. Desoer and Felix F. Wu.

Nonlinear monotone networks. SIAM Journal on Applied Mathematics, 26 (2):315–333, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miloh:1974:UIS

[1070] Touvia Miloh. The ultimate image singularities for external ellipsoidal harmonics. SIAM Journal on Applied Mathematics, 26(2):334–344, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gunderson:1974:TDS

[1071] H. Gunderson, H. Rigas, and F. S. VanVleck. A technique for determining stability regions for the damped Mathieu equation. SIAM Journal on Applied Mathematics, 26(2):345–349, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Datta:1974:RRS

[1072] S. K. Datta and J. D. Sangster. Response of a rigid spheroidal inclusion to an incident plane compressional elastic wave. SIAM Journal on Applied Mathematics, 26(2):350–369, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kaufman:1974:CCF

[1073] Marc T. Kaufman. Counterexample to a conjecture of Fujii, Kasami and Ninomiya. SIAM Journal on Applied Mathematics, 26(2):370–371, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huff:1974:RWC

[1074] Barthel W. Huff. Random walks and the continuity of infinitely divisible distributions. SIAM Journal on Applied Mathematics, 26(2):372–375, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miura:1974:ANL

[1075] Robert M. Miura and Martin D.

Kruskal. Application of a non linear WKB method to the Korteweg– de Vries equation. SIAM Journal on Applied Mathematics, 26(2):376–395, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1974:COP

[1076] Edwin Cohen and Henry Ruston. A class of optimal polynomials related to electric band-pass filters. SIAM Journal on Applied Mathematics, 26 (2):396–412, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bush:1974:AAT

[1077] W. B. Bush and F. E. Fendell. Asymptotic analysis of turbulent channel flow for mixing length theory. SIAM Journal on Applied Mathematics, 26 (2):413–427, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Farooqui:1974:TEI

[1078] A. S. Farooqui and D. L. Sharma. A theorem on electrical induction. SIAM Journal on Applied Mathematics, 26 (2):428–430, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carlitz:1974:AMM

[1079] L. Carlitz. An application of MacMahon’s master theorem. SIAM Journal on Applied Mathematics, 26(2):431– 436, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wu:1974:FTM

[1080] Chien-Heng Wu. Flattening of tubelike membranes of revolution by stretching — asymptotic solution with boundary layer. SIAM Journal on Applied Mathematics, 26(2):437–446, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

DiDonato:1974:MLE

[1081] A. R. DiDonato and M. P. Jarnagin, Jr. Maximum likelihood estimation in quantal response experiments. SIAM Journal on Applied Mathematics, 26 (2):447–454, March 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vaid:1974:IET

[1082] B. K. Vaid and D. L. Jain. An integral equation technique. SIAM Journal on Applied Mathematics, 26(3):455–468, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1974:LIS

[1083] Carl D. Meyer, Jr. Limits and the index of a square matrix. SIAM Journal on Applied Mathematics, 26(3):469–478, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burdet:1974:GAF

[1084] Claude-Alain Burdet. Generating all the faces of a polyhedron. SIAM Journal on Applied Mathematics, 26 (3):479–489, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cheo:1974:SBC

[1085] L. S. Cheo and Edward L. Reiss.

Secondary buckling of circular plates. SIAM Journal on Applied Mathematics, 26(3):490–495, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bazaraa:1974:DSP

[1086] M. S. Bazaraa and R. W. Langley. A dual shortest path algorithm. SIAM Journal on Applied Mathematics, 26(3):496–501, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Salehi:1974:BLP

[1087] H. Salehi. On the bilateral linear predictor for minimal stationary stochastic processes. SIAM Journal on Applied Mathematics, 26(3):502–507, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ponzo:1974:WCH

[1088] Peter J. Ponzo and Nelson Wax.

Weakly coupled harmonic oscillators. SIAM Journal on Applied Mathematics, 26(3):508–527, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gould:1974:NNL

[1089] H. W. Gould. A note on the number of linearly inducible orderings of points in d-space. SIAM Journal on Applied Mathematics, 26(3):528–530, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morgan:1974:MMS

[1090] Richard C. Morgan and Samuel N. Karp. Multi-mode surface wave diffraction by a wedge. SIAM Journal on Applied Mathematics, 26(3):531–538, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Geer:1974:UAS

[1091] James F. Geer. Uniform asymptotic solutions for the two-dimensional potential field about a slender body. SIAM Journal on Applied Mathematics, 26(3):539–553, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chang:1974:ASN

[1092] K. W. Chang. Approximate solutions of nonlinear boundary value problems involving a small parameter. SIAM Journal on Applied Mathematics, 26 (3):554–567, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sarma:1974:FLF

[1093] P. V. B. A. S. Sarma, B. S. Ramachandra Rao, and S. Gopalacharyulu. Flexure of a long flat curved plate under end loadings. SIAM Journal on Applied Mathematics, 26(3):568–577, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lindstrom:1974:SIPa

[1094] F. T. Lindstrom and W. M. Stone. On the start up or initial phase of linear mass transport of chemicals in a water saturated sorbing porous medium. I. SIAM Journal on Applied Mathematics, 26(3):578–591, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lindstrom:1974:SIPb

[1095] F. T. Lindstrom and W. M. Stone. On the start up or initial phase of linear mass transport of chemicals in a water saturated sorbing porous medium. II: Integral equation approach. SIAM Journal on Applied Mathematics, 26 (3):592–606, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Megiddo:1974:NCS

[1096] Nimrod Megiddo. Nucleoluses of compound simple games. SIAM Journal on Applied Mathematics, 26(3):607–621, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hansen:1974:ANE

[1097] Terje Hansen. On the approximation of Nash equilibrium points in an Nperson noncooperative game. SIAM Journal on Applied Mathematics, 26 (3):622–637, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kevorkian:1974:MRR

[1098] J. Kevorkian. On a model for reentry roll resonance. SIAM Journal on Applied Mathematics, 26(3):638–669, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schmitt:1974:BVP

[1099] Klaus Schmitt. Boundary value problems and comparison theorems for ordinary differential equations. SIAM Journal on Applied Mathematics, 26 (3):670–678, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Srivastav:1974:PEA

[1100] R. P. Srivastav and R. P. Khetan. A procedure for elastic analysis of cracked plates. SIAM Journal on Applied Mathematics, 26(3):679–685, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kevorkian:1974:EPT

[1101] J. Kevorkian. Erratum: “Passage through Resonance for a One-Dimensional Oscillator with Slowly Varying Frequency” [SIAM J. Appl. Math. 20 (1971), no. 3, 364–373]. SIAM Journal on Applied Mathematics, 26(3):686, May 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [688].

Parter:1974:SDE

[1102] Seymour V. Parter. Solutions of a differential equation arising in chemical reactor processes. SIAM Journal on Applied Mathematics, 26(4):687–716, June 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1974:AST

[1103] Jye Chen and R. E. O’Malley, Jr. On the asymptotic solution of a twoparameter boundary value problem of chemical reactor theory. SIAM Journal on Applied Mathematics, 26(4):717– 729, June 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kimura:1974:IPS

[1104] Hidenori Kimura. An inverse problem for second order differential systems and waveform design. SIAM Journal on Applied Mathematics, 26 (4):730–744, June 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Clements:1974:ETQ

[1105] John Clements. Existence theorems for a quasilinear evolution equation. SIAM Journal on Applied Mathematics, 26(4):745–752, June 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Crum:1974:SRE

[1106] Larry A. Crum and James A. Heinen.

Simultaneous reduction and expansion of multidimensional Laplace transform kernels. SIAM Journal on Applied Mathematics, 26(4):753–771, June 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Seguin:1974:OOR

[1107] Gerald Seguin. On obtaining orbit representatives for a class of groups with operators. SIAM Journal on Applied Mathematics, 26(4):772–775, June 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cochran:1974:EAC

[1108] James Alan Cochran and Erold W.

Hinds. Eigensystems associated with the complex-symmetric kernels of laser theory. SIAM Journal on Applied Mathematics, 26(4):776–786, June 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sellers:1974:TCE

[1109] Peter H. Sellers. On the theory and computation of evolutionary distances. SIAM Journal on Applied Mathematics, 26(4):787–793, June 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Herbut:1974:CFM

[1110] Fedor Herbut, Philippe Loncke, and Milan Vujicic. Canonical form for matrices under unitary congruence transformations. II: Congruence-normal matrices. SIAM Journal on Applied Mathematics, 26(4):794–805, June 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Saari:1974:RMT

[1111] Donald G. Saari. Restrictions on the motion of the three-body problem. SIAM Journal on Applied Mathematics, 26(4):806–815, June 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Steinmetz:1974:NSP

[1112] William J. Steinmetz. On a nonlinear singular perturbation boundary value problem in gas lubrication theory. SIAM Journal on Applied Mathematics, 26(4):816–827, June 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Greville:1974:SME

[1113] T. N. E. Greville. Solutions of the matrix equation XAX = X, and relations between oblique and orthogonal projectors. SIAM Journal on Applied Mathematics, 26(4):828–832, June 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moore:1974:AGI

[1114] R. H. Moore and M. Z. Nashed. Approximations to generalized inverses of linear operators. SIAM Journal on Applied Mathematics, 27(1):1–16, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gal:1974:MSL

[1115] Shmuel Gal. Minimax solutions for linear search problems. SIAM Journal on Applied Mathematics, 27(1):17–30, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fisher:1974:CDI

[1116] Marshall L. Fisher and Jeremy F.

Shapiro. Constructive duality in integer programming. SIAM Journal on Applied Mathematics, 27(1):31–52, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jensen:1974:RSD

[1117] D. R. Jensen. Robustness of some distribution of products and quotients under ϵ- independence. SIAM Journal on Applied Mathematics, 27(1):53–58, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Duncan:1974:HQT

[1118] George T. Duncan. Heterogeneous questionnaire theory. SIAM Journal on Applied Mathematics, 27(1):59–71, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chan:1974:EUU

[1119] C. Y. Chan. Existence, uniqueness, upper and lower bounds of solutions of nonlinear space-time nuclear reactor kinetics. SIAM Journal on Applied Mathematics, 27(1):72–82, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pao:1974:MSA

[1120] C. V. Pao. Method of successive approximations for nonlinear stationary transport equations. SIAM Journal on Applied Mathematics, 27(1):83–92, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dickey:1974:BPF

[1121] R. W. Dickey. A bifurcation problem for a fourth order nonlinear ordinary differential equation. SIAM Journal on Applied Mathematics, 27(1):93–101, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Simmonds:1974:FNP

[1122] J. G. Simmonds and C. G. Tropf. The fundamental (normal point load) solution for a shallow hyperbolic paraboloidal shell. SIAM Journal on Applied Mathematics, 27(1):102–120, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ojha:1974:LLC

[1123] M. P. Ojha. Location of limit cycles by a comparison method. SIAM Journal on Applied Mathematics, 27(1):121– 130, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mereau:1974:SOC

[1124] Pierre Mereau and Jean-Guy Paquet.

Second order conditions for pseudoconvex functions. SIAM Journal on Applied Mathematics, 27(1):131–137, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1974:RMS

[1125] L. W. Chen and L. B. Felsen. Ray method for scattering by discontinuities in a multiwave wave guide. SIAM Journal on Applied Mathematics, 27 (1):138–158, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baker:1974:SPS

[1126] John W. Baker. Stability properties of a second order damped and forced nonlinear differential equation. SIAM Journal on Applied Mathematics, 27(1):159–166, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kohler:1974:NAS

[1127] Werner E. Kohler and William E.

Boyce. A numerical analysis of some first order stochastic initial value problems. SIAM Journal on Applied Mathematics, 27(1):167–179, July 1974. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eliason:1974:LTI

[1128] Stanley B. Eliason. Lyapunov type inequalities for certain second order functional differential equations. SIAM Journal on Applied Mathematics, 27 (1):180–199, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Perko:1974:POR

[1129] L. M. Perko. Periodic orbits in the restricted three-body problem: Existence and asymptotic approximation. SIAM Journal on Applied Mathematics, 27(1):200–237, July 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stone:1974:STC

[1130] Lawrence D. Stone and Henry R. Richardson. Search for targets with conditionally deterministic motion. SIAM Journal on Applied Mathematics, 27(2):239–255, September 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berger:1974:EDV

[1131] Neil Berger. Estimates for the derivatives of the velocity and pressure in shallow water flow and approximate shallow water equations. SIAM Journal on Applied Mathematics, 27(2): 256–280, September 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartfiel:1974:CSI

[1132] D. J. Hartfiel. Concerning spectral inverses of stochastic matrices. SIAM Journal on Applied Mathematics, 27 (2):281–292, September 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kara:1974:WSM

[1133] H. Kara, V. Mandrekar, and G. L.

Park. Wide-sense martingale approach to linear optimal estimation. SIAM Journal on Applied Mathematics, 27 (2):293–302, September 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rorres:1974:TCE

[1134] Chris Rorres. Transmission coefficients and eigenvalues of a finite onedimensional crystal. SIAM Journal on Applied Mathematics, 27(2):303–321, September 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stallings:1974:SSI

[1135] W. T. Stallings and T. L. Boullion.

A strong spectral inverse for an rcirculant matrix. SIAM Journal on Applied Mathematics, 27(2):322–325, September 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barwick:1974:GRO

[1136] D. T. Barwick and J. D. Gilbert. On generalizations of the reverse order law. SIAM Journal on Applied Mathematics, 27(2):326–330, September 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Warsi:1974:CEF

[1137] Z. U. A. Warsi. A class of external flows exhibiting self-similarity in threedimensional boundary layers. SIAM Journal on Applied Mathematics, 27 (2):331–340, September 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kuzanek:1974:EUS

[1138] Jerry F. Kuzanek. Existence and uniqueness of solutions to a fourth order nonlinear eigenvalue problem. SIAM Journal on Applied Mathematics, 27(2):341–354, September 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Megiddo:1974:NBS

[1139] Nimrod Megiddo. On the nonmonotonicity of the bargaining set, the kernel and the nucleolus of a game. SIAM Journal on Applied Mathematics, 27 (2):355–358, September 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gupta:1974:GHD

[1140] R. P. Gupta and G. C. Jain. A generalized Hermite distribution and its properties. SIAM Journal on Applied Mathematics, 27(2):359–363, September 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Banks:1974:BEC

[1141] D. O. Banks. Bounds for eigenvalues and conditions for existence of conjugate points. SIAM Journal on Applied Mathematics, 27(3):365–375, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chow:1974:EAS

[1142] Pao-Liu Chow. On the exact and approximate solutions of a random parabolic equation. SIAM Journal on Applied Mathematics, 27(3):376–397, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dolan:1974:DB

[1143] J. Michael Dolan and Gene A. Klaasen.

On the disconjugate behavior of y′′′ + py = 0. SIAM Journal on Applied Mathematics, 27(3):398–403, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Idemen:1974:FER

[1144] Mithat Idemen. On the functional equation related to the three-part mixed boundary value problems. SIAM Journal on Applied Mathematics, 27 (3):404–415, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1974:TCR

[1145] Donald S. Cohen and Aubrey B. Poore.

Tubular chemical reactors: The “lumping approximation” and bifurcation of oscillatory states. SIAM Journal on Applied Mathematics, 27(3):416–429, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ludford:1974:SR

[1146] G. S. S. Ludford and Sharon S. Wilson. Subcharacteristic reversal. SIAM Journal on Applied Mathematics, 27 (3):430–440, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lange:1974:SCE

[1147] Charles G. Lange and Alan C. Newell. A stability criterion for envelope equations. SIAM Journal on Applied Mathematics, 27(3):441–456, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosenstock:1974:RWL

[1148] Herbert B. Rosenstock. Random walk on lattices with uncertain traps. SIAM Journal on Applied Mathematics, 27 (3):457–463, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Davis:1974:SSWa

[1149] A. M. J. Davis. Short surface waves in the presence of a submerged sphere. SIAM Journal on Applied Mathematics, 27(3):464–478, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Davis:1974:SSWb

[1150] A. M. J. Davis. Short surface waves in the presence of a submerged circular cylinder. SIAM Journal on Applied Mathematics, 27(3):479–491, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Davis:1974:EDS

[1151] C. W. Davis. Elementary divisors of the Stein transformation X → X − CXC. SIAM Journal on Applied Mathematics, 27(3):492–494, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gibson:1974:UOT

[1152] P. M. Gibson. Unitary and orthogonal transformations on matrices. SIAM Journal on Applied Mathematics, 27 (3):495–499, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stineman:1974:CCM

[1153] Russell W. Stineman. Calculation of a constrained minimal symmetric matrix. SIAM Journal on Applied Mathematics, 27(3):500–502, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1974:LSV

[1154] Ronald Smith. A least squares version of short wave asymptotics. SIAM Journal on Applied Mathematics, 27 (3):503–516, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kymn:1974:SDR

[1155] Kern O. Kymn. A simple derivation of the relation between t- and F -distributions. SIAM Journal on Applied Mathematics, 27(3):517–518, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sharma:1974:EIT

[1156] Bhu Dev Sharma and Ram Autur. Erratum: “An Inversion Theorem and Generalized Entropies for Continuous Distributions” [SIAM J. Appl. Math. 25 (1973), no. 2, 125–132]. SIAM Journal on Applied Mathematics, 27 (3):519, November 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [989].

Novomestky:1974:AEU

[1157] F. Novomestky. Asymptotic expressions for the unit-step and Dirac delta functions. SIAM Journal on Applied Mathematics, 27(4):521–525, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gautesen:1974:RFD

[1158] A. K. Gautesen. Resonance for a forced N-dimensional oscillator. SIAM Journal on Applied Mathematics, 27 (4):526–530, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ridenhour:1974:CCP

[1159] Jerry R. Ridenhour. On the continuity of conjugate point functions. SIAM Journal on Applied Mathematics, 27 (4):531–538, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Buchanan:1974:BGC

[1160] J. Buchanan. Bounds on the growth of a class of oscillatory solutions of y′(x) = my(x−d(x)) with bounded delay. SIAM Journal on Applied Mathematics, 27(4):539–543, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wyner:1974:PBO

[1161] Aaron D. Wyner. On the probability of buffer overflow under an arbitrary bounded input-output distribution. SIAM Journal on Applied Mathematics, 27(4):544–570, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mitchell:1974:SSS

[1162] Richard R. Mitchell and Frank Kozin.

Sample stability of second order linear differential equations with wide band noise coefficients. SIAM Journal on Applied Mathematics, 27(4): 571–605, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [1232].

Achou:1974:NFS

[1163] O. Achou. The number of feasible solutions to a knapsack problem. SIAM Journal on Applied Mathematics, 27 (4):606–610, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sethna:1974:SAR

[1164] P. R. Sethna and M. Balachandra.

Some asymptotic results for systems with multiple time scales. SIAM Journal on Applied Mathematics, 27(4): 611–625, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cook:1974:CGT

[1165] Roger J. Cook. Complementary graphs and total chromatic numbers. SIAM Journal on Applied Mathematics, 27 (4):626–628, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abrams:1974:RRP

[1166] Robert Abrams and Marcus Bunting.

Reducing reversed posynomial programs. SIAM Journal on Applied Mathematics, 27(4):629–640, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gal:1974:DSG

[1167] Shmuel Gal. A discrete search game. SIAM Journal on Applied Mathematics, 27(4):641–648, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ghandour:1974:IVP

[1168] Edmond Ghandour. Initial value problem for boundary values of a Green’s function. SIAM Journal on Applied Mathematics, 27(4):649–655, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abraham-Shrauner:1974:PEP

[1169] Barbara Abraham-Shrauner. Perturbation expansions for the potential of a small radius, charged dielectric sphere in 1-1 electrolytes. SIAM Journal on Applied Mathematics, 27(4): 656–665, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Freedman:1975:PTDa

[1170] H. I. Freedman and P. Waltman. Perturbation of two-dimensional predatorprey equations. SIAM Journal on Applied Mathematics, 28(1):1–10, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rogers:1975:IRP

[1171] C. Rogers. Invariance and reducibility properties under a class of (m, 2)- baecklund transformations. SIAM Journal on Applied Mathematics, 28 (1):11–23, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Komkov:1975:CES

[1172] V. Komkov. Certain estimates for solutions of nonlinear elliptic differential equations applicable to the theory of thin plates. SIAM Journal on Applied Mathematics, 28(1):24–34, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sankoff:1975:MMT

[1173] David Sankoff. Minimal mutation trees of sequences. SIAM Journal on Applied Mathematics, 28(1):35–42, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ewer:1975:ABS

[1174] J. P. G. Ewer and L. A. Peletier. On the asymptotic behavior of solutions of semilinear parabolic equations. SIAM Journal on Applied Mathematics, 28 (1):43–53, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Testa:1975:CDL

[1175] F. J. Testa. Congruence diagonalization and Lie groups. SIAM Journal on Applied Mathematics, 28(1):54– 59, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Anderson:1975:SOI

[1176] W. N. Anderson, Jr. and G. E. Trapp.

Shorted operators. II. SIAM Journal on Applied Mathematics, 28(1):60– 71, January 1975. CODEN SMJMAP.

ISSN 0036-1399 (print), 1095-712X (electronic).

Dobbie:1975:SAT

[1177] James M. Dobbie. Search for an avoiding target. SIAM Journal on Applied Mathematics, 28(1):72–86, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Flynn:1975:SSP

[1178] James Flynn. Steady state policies for a class of deterministic dynamic programming models. SIAM Journal on Applied Mathematics, 28(1):87– 99, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Witsenhausen:1975:SPD

[1179] H. S. Witsenhausen. On sequences of pairs of dependent random variables. SIAM Journal on Applied Mathematics, 28(1):100–113, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Anderson:1975:MMR

[1180] William N. Anderson, Jr. Maximum matching and the rank of a matrix. SIAM Journal on Applied Mathematics, 28(1):114–123, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pitkaranta:1975:SPP

[1181] J. Pitkaranta and P. Silvennoinen. A symmetrized perturbation procedure for the transport operator. SIAM Journal on Applied Mathematics, 28 (1):124–129, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zischka:1975:CBV

[1182] K. A. Zischka and P. S. Chow. On a certain boundary value problem and its relevance in the theory of surface induction hardening. SIAM Journal on Applied Mathematics, 28(1):130–153, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartwig:1975:RGI

[1183] Robert E. Hartwig. AX − XB = C, resultants and generalized inverses. SIAM Journal on Applied Mathematics, 28(1):154–183, January 1975. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nath:1975:UTM

[1184] V. I. Nath. On the uniqueness of transverse MHD flows with same streamline pattern. SIAM Journal on Applied Mathematics, 28(1):184–189, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

deFigueiredo:1975:CMB

[1185] Rui J. P. de Figueiredo and A. N. Netravali. On the construction of minimum bandwidth functions from discrete data. SIAM Journal on Applied Mathematics, 28(1):190–201, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lagerstrom:1975:SNS

[1186] Paco A. Lagerstrom. Solutions of the Navier–Stokes equation at large Reynolds number. SIAM Journal on Applied Mathematics, 28(1):202–214, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sears:1975:BLS

[1187] W. R. Sears and D. P. Telionis. Boundary-layer separation in unsteady flow. SIAM Journal on Applied Mathematics, 28(1):215–235, January 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Downs:1975:SPS

[1188] T. Downs. Some properties of the Souriau-frame algorithm with application to the inversion of rational matrices. SIAM Journal on Applied Mathematics, 28(2):237–251, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosenstark:1975:SSQ

[1189] Jeannette Rosenstark. On a sequence of subcodes of the quadratic residue code for p = 8m + 1. SIAM Journal on Applied Mathematics, 28(2):252–264, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Singh:1975:NFF

[1190] Bhagat Singh. Nonoscillation of forced fourth order retarded equations. SIAM Journal on Applied Mathematics, 28 (2):265–269, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Good:1975:LDB

[1191] I. J. Good. The Lagrange distributions and branching processes. SIAM Journal on Applied Mathematics, 28 (2):270–275, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bagchi:1975:MAC

[1192] Arunabha Bagchi. A martingale approach to continuous-time linear smoothing. SIAM Journal on Applied Mathematics, 28(2):276–281, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Foschini:1975:DAM

[1193] Gerard J. Foschini. Demodulating analytically modulated signals after propagation over certain channels. SIAM Journal on Applied Mathematics, 28 (2):282–289, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kundu:1975:DRT

[1194] Sukhamay Kundu. Disjoint representation of three tree realizable sequences. I. SIAM Journal on Applied Mathematics, 28(2):290–302, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bohm:1975:COP

[1195] Volker Bohm. On the continuity of the optimal policy set for linear programs. SIAM Journal on Applied Mathematics, 28(2):303–306, March 1975. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1975:OPT

[1196] Donald S. Cohen and James P. Keener.

Oscillatory processes in the theory of particulate formation in supersaturated chemical solutions. SIAM Journal on Applied Mathematics, 28 (2):307–318, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Terry:1975:SOC

[1197] Raymond D. Terry. Some oscillation criteria for delay-differential equations of even order. SIAM Journal on Applied Mathematics, 28(2):319–334, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Balas:1975:NQP

[1198] Egon Balas. Nonconvex quadratic programming via generalized polars. SIAM Journal on Applied Mathematics, 28(2):335–349, March 1975. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abramovich:1975:MEU

[1199] Shoshana Abramovich. Monotonicity of eigenvalues under symmetrization. SIAM Journal on Applied Mathematics, 28(2):350–361, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Testa:1975:BSC

[1200] F. J. Testa. Bounds on solutions to certain nonlinear reaction-diffusion equations. SIAM Journal on Applied Mathematics, 28(2):362–366, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pack:1975:COP

[1201] Charles D. Pack. A comparison of the output process of an M/D/1 queue as measured from an arbitrary departure epoch to that measured from an arbitrary instant in time. SIAM Journal on Applied Mathematics, 28 (2):367–375, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gordon:1975:CCT

[1202] M. Gordon and J. W. Kennedy. The counting and coding of trees of fixed diameter. SIAM Journal on Applied Mathematics, 28(2):376–398, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Parihar:1975:SIF

[1203] K. S. Parihar and P. S. Kushwaha.

The stress intensity factors for two symmetrically located Griffith cracks in an elastic strip in which symmetrical body forces are acting. SIAM Journal on Applied Mathematics, 28 (2):399–410, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ungar:1975:EPW

[1204] Abraham Ungar. An electromagnetic progressing wave representation for the field of a uniformly moving point charge. SIAM Journal on Applied Mathematics, 28(2):411–418, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Horgan:1975:KII

[1205] Cornelius O. Horgan. On Korn’s inequality for incompressible media. SIAM Journal on Applied Mathematics, 28(2):419–430, March 1975. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rogler:1975:DBL

[1206] Harold L. Rogler and Eli Reshotko. Disturbances in a boundary layer introduced by a low intensity array of vortices. SIAM Journal on Applied Mathematics, 28(2):431–462, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carrier:1975:ESR

[1207] G. F. Carrier, F. E. Fendell, and F. E. Marble. The effect of strain rate on diffusion flames. SIAM Journal on Applied Mathematics, 28(2):463–500, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stewartson:1975:ATS

[1208] K. Stewartson. On the asymptotic theory of separated and unseparated fluid motions. SIAM Journal on Applied Mathematics, 28(2):501–518, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mair:1975:VWB

[1209] W. A. Mair and P. K. Stansby. Vortex wakes of bluff cylinders in shear flow. SIAM Journal on Applied Mathematics, 28(2):519–540, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cahalon:1975:RBS

[1210] Baruch Cahalon. The relation between the steady state solution and the solution of a minimum problem. SIAM Journal on Applied Mathematics, 28(3):541–546, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tan:1975:SMR

[1211] W. Y. Tan. Some matrix results and extensions of Cochran’s theorem. SIAM Journal on Applied Mathematics, 28(3):547–554, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [1361].

Wright:1975:MHP

[1212] D. J. Wright. Multidimensional Hermite polynomials and Ito equations. SIAM Journal on Applied Mathematics, 28(3):555–558, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neapolitan:1975:PBS

[1213] Richard E. Neapolitan and Warren S.

Edelstein. A priori bounds for the secondary boundary value problem in linear viscoelasticity. SIAM Journal on Applied Mathematics, 28(3):559–564, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gondran:1975:RSS

[1214] M. Gondran and J. F. Maurras. Remarks on the structure of some integer algorithms. SIAM Journal on Applied Mathematics, 28(3):565–567, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ecker:1975:DMQ

[1215] J. G. Ecker and R. D. Niemi. A dual method for quadratic programs with quadratic constraints. SIAM Journal on Applied Mathematics, 28(3):568– 576, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carlson:1975:MDI

[1216] David Carlson. Matrix decompositions involving the Schur complement. SIAM Journal on Applied Mathematics, 28(3):577–587, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Demoulin:1975:IMS

[1217] Y.-M. J. Demoulin and Y. M. Chen. An iteration method for solving nonlinear eigenvalue problems. SIAM Journal on Applied Mathematics, 28(3):588–595, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lemoine:1975:LTG

[1218] Austin J. Lemoine. Limit theorems for generalized single server queues: The exceptional system. SIAM Journal on Applied Mathematics, 28(3):596–606, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartfiel:1975:SMU

[1219] D. J. Hartfiel and C. J. Maxson. A study of minimally unilateral digraphs. SIAM Journal on Applied Mathematics, 28(3):607–610, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Travis:1975:ODT

[1220] S. P. Travis. A one-dimensional two-body problem of classical electrodynamics. SIAM Journal on Applied Mathematics, 28(3):611–632, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bleistein:1975:ASL

[1221] N. Bleistein and J. K. Cohen. An analysis of scaling for linear inviscid internal waves. SIAM Journal on Applied Mathematics, 28(3):633–641, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bush:1975:AAS

[1222] William B. Bush. An asymptotic analysis of strong radiation-resisted normal shock waves. SIAM Journal on Applied Mathematics, 28(3):642–656, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Herbert:1975:NFE

[1223] H. G. Herbert. A note on first emptiness in a discrete storage system with Markovian inputs. SIAM Journal on Applied Mathematics, 28(3):657–661, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Greenberg:1975:ABS

[1224] J. M. Greenberg and F. Hoppensteadt.

Asymptotic behavior of solutions to a population equation. SIAM Journal on Applied Mathematics, 28(3):662–674, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tyagi:1975:SPT

[1225] M. S. Tyagi. On some properties of ntournaments: A note. SIAM Journal on Applied Mathematics, 28(3):675– 677, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hastings:1975:EOS

[1226] S. P. Hastings and J. D. Murray. The existence of oscillatory solutions in the Field–Noyes model for the Belousov– Zhabotinskii reaction. SIAM Journal on Applied Mathematics, 28(3):678– 688, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Liepmann:1975:FDL

[1227] H. W. Liepmann. Fluid dynamics of liquid helium. SIAM Journal on Applied Mathematics, 28(3):689–699, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Witting:1975:HOS

[1228] James Witting. On the highest and other solitary waves. SIAM Journal on Applied Mathematics, 28(3):700–719, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

vanDyke:1975:CEP

[1229] Milton van Dyke. Computer extension of perturbation series in fluid mechanics. SIAM Journal on Applied Mathematics, 28(3):720–734, May 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Landahl:1975:WBT

[1230] Marten T. Landahl. Wave breakdown and turbulence. SIAM Journal on Applied Mathematics, 28(4):735–756, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Trilling:1975:KBL

[1231] Leon Trilling. The Knudsen boundary layer in low density gas flow past a crystal. SIAM Journal on Applied Mathematics, 28(4):757–777, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mitchell:1975:ESS

[1232] Richard R. Mitchell and Frank Kozin.

Erratum: “Sample Stability of Second Order Linear Differential Equations with Wide Band Noise Coefficients” [SIAM J. Appl. Math. 27 (1974), no. 4, 571–605]. SIAM Journal on Applied Mathematics, 28(4):778, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1162].

Eisner:1975:DSP

[1233] Mark J. Eisner and Paul Olsen. Duality for stochastic programming interpreted as L.P. in L_{p}-space. SIAM Journal on Applied Mathematics, 28 (4):779–792, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Masry:1975:CES

[1234] Elias Masry and Ming-Chuan C. Lui.

A consistent estimate of the spectrum by random sampling of the time series. SIAM Journal on Applied Mathematics, 28(4):793–810, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bush:1975:SDM

[1235] William B. Bush. The slip-dominated merged layer regime for Newtonian hypersonic flow past a flat plate. SIAM Journal on Applied Mathematics, 28 (4):811–823, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Young:1975:SCS

[1236] H. P. Young. Social choice scoring functions. SIAM Journal on Applied Mathematics, 28(4):824–838, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kaminetzky:1975:DEV

[1237] Lee Kaminetzky and Joseph B. Keller. Diffraction by edges and vertices of interfaces. SIAM Journal on Applied Mathematics, 28(4):839–856, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shen:1975:URT

[1238] M. C. Shen and J. B. Keller. Uniform ray theory of surface, internal and acoustic wave propagation in a rotating ocean or atmosphere. SIAM Journal on Applied Mathematics, 28(4):857–875, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mallows:1975:TGS

[1239] C. L. Mallows and N. J. A. Sloane. Two-graphs, switching classes and Euler graphs are equal in number. SIAM Journal on Applied Mathematics, 28 (4):876–880, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Looney:1975:LUQ

[1240] Carl G. Looney. Locally uniformly quasi-convex programming. SIAM Journal on Applied Mathematics, 28 (4):881–884, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ghandour:1975:ODR

[1241] Edmond Ghandour. One-dimensional random wave propagation. SIAM Journal on Applied Mathematics, 28 (4):885–898, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yih:1975:VVR

[1242] Chia-Shun Yih. Vortices and vortex rings of stratified fluids. SIAM Journal on Applied Mathematics, 28 (4):899–912, June 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Einaudi:1975:CIM

[1243] F. Einaudi. On the convergence of an iterative method for studying nonlinear oscillations without damping. SIAM Journal on Applied Mathematics, 29(1):1–11, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rappaport:1975:CSI

[1244] David Rappaport and Joshua Dayan.

On convergence for spatially inhomogeneous multiphase transport systems. SIAM Journal on Applied Mathematics, 29(1):12–19, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miller:1975:SRE

[1245] R. K. Miller. A system of renewal equations. SIAM Journal on Applied Mathematics, 29(1):20–34, July 1975.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wilson:1975:BLS

[1246] H. K. Wilson. A boundary layer solution for a flow past a plate with parabolic edge. SIAM Journal on Applied Mathematics, 29(1):35–39, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pao:1975:EDD

[1247] C. V. Pao. The energy-dependent diffusion system in nuclear reactor dynamics. SIAM Journal on Applied Mathematics, 29(1):40–59, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kalai:1975:EFC

[1248] Ehud Kalai. Excess functions for cooperative games without sidepayments. SIAM Journal on Applied Mathematics, 29(1):60–71, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mehri:1975:EPS

[1249] B. Mehri and G. G. Hamedani. On the existence of periodic solutions of nonlinear second order differential equations. SIAM Journal on Applied Mathematics, 29(1):72–76, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jones:1975:CTM

[1250] Robert A. Jones. Comparison theorems for matrix Riccati equations. SIAM Journal on Applied Mathematics, 29(1):77–90, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neufeld:1975:GGC

[1251] G. A. Neufeld and J. Tartar. Generalized graph colorations. SIAM Journal on Applied Mathematics, 29(1):91–98, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klaasen:1975:CDP

[1252] Gene A. Klaasen. Continuous dependence for N-point boundary value problems. SIAM Journal on Applied Mathematics, 29(1):99–102, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lay:1975:SPG

[1253] David C. Lay. Spectral properties of generalized inverses of linear operators. SIAM Journal on Applied Mathematics, 29(1):103–109, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosenblat:1975:ASL

[1254] S. Rosenblat and J. Shepherd. On the asymptotic solution of the Lagerstrom model equation. SIAM Journal on Applied Mathematics, 29(1):110–120, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Adamson:1975:IBS

[1255] T. C. Adamson, Jr. and A. Feo. Interaction between a shock wave and a turbulent boundary layer in transonic flow. SIAM Journal on Applied Mathematics, 29(1):121–145, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosen:1975:SNR

[1256] Gerald Rosen. Solutions to the nonlinear recombination equation for the infinite spatial domain. SIAM Journal on Applied Mathematics, 29(1):146–151, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hall:1975:GIB

[1257] Frank J. Hall. Generalized inverses of a bordered matrix of operators. SIAM Journal on Applied Mathematics, 29(1):152–163, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boersma:1975:ROA

[1258] J. Boersma. Ray-optical analysis of reflection in an open-ended parallel-plane waveguide. I: TM case. SIAM Journal on Applied Mathematics, 29(1):164– 195, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lopes:1975:FON

[1259] Orlando Lopes. Forced oscillations in nonlinear neutral differential equations. SIAM Journal on Applied Mathematics, 29(1):196–207, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Leversen:1975:SSF

[1260] Tore Leversen and Jacqueline Naze Tjotta. Solution of a stationary Fokker–Planck equation. SIAM Journal on Applied Mathematics, 29 (1):208–217, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schmerr:1975:CSW

[1261] Lester W. Schmerr, Jr. A class of shear wave problems in an elastic wedge. SIAM Journal on Applied Mathematics, 29(1):218–224, July 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Roberts:1975:PSI

[1262] Fred S. Roberts. [preface] (in special issue on mathematics and the social and biological sciences). SIAM Journal on Applied Mathematics, 29(2):v, September 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Norman:1975:ASP

[1263] M. Frank Norman. Approximation of stochastic processes by Gaussian diffusions, and applications to Wright– Fisher genetic models (in special issue on mathematics and the social and biological sciences). SIAM Journal on Applied Mathematics, 29(2): 225–242, September 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

May:1975:NAC

[1264] Robert M. May and Warren J.

Leonard. Nonlinear aspects of competition between three species (in special issue on mathematics and the social and biological sciences). SIAM Journal on Applied Mathematics, 29 (2):243–253, September 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bogart:1975:PSI

[1265] Kenneth P. Bogart. Preference structures II: Distances between asymmetric relations (in special issue on mathematics and the social and biological sciences). SIAM Journal on Applied Mathematics, 29(2):254–262, September 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fishburn:1975:FAC

[1266] Peter C. Fishburn. Finite additive choice structures (in special issue on mathematics and the social and biological sciences). SIAM Journal on Applied Mathematics, 29(2):263–272, September 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wolfe:1975:LFD

[1267] Peter Wolfe. Low frequency diffraction by a hard strip. SIAM Journal on Applied Mathematics, 29(2): 273–287, September 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lindstrom:1975:NLT

[1268] F. T. Lindstrom and F. Oberhettinger. A note on a Laplace transform pair associated with mass transport in porous media and heat transport problems. SIAM Journal on Applied Mathematics, 29(2):288–292, September 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bogdonoff:1975:SRS

[1269] S. M. Bogdonoff. Some remarks on shock wave turbulent boundary-layer interaction at high Reynolds number. SIAM Journal on Applied Mathematics, 29(2):293–308, September 1975.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rom:1975:FSI

[1270] J. Rom. Flows with strong interaction between the viscous and inviscid regions. SIAM Journal on Applied Mathematics, 29(2):309–328, September 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abarbanel:1975:DSF

[1271] S. Abarbanel, D. Gottlieb, and E. Turkel. Difference schemes with fourth order accuracy for hyperbolic equations. SIAM Journal on Applied Mathematics, 29(2):329–351, September 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lin:1975:SWG

[1272] C. C. Lin and Y. Y. Lau. On spiral waves in galaxies — A gas dynamic approach. SIAM Journal on Applied Mathematics, 29(2):352–370, September 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fix:1975:FEM

[1273] George J. Fix. Finite element models for ocean circulation problems. SIAM Journal on Applied Mathematics, 29 (3):371–387, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Megiddo:1975:TDC

[1274] Nimrod Megiddo. Tensor decomposition of cooperative games. SIAM Journal on Applied Mathematics, 29 (3):388–405, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gunderson:1975:SCM

[1275] R. W. Gunderson and E. E. Underwood. Stability of a class of matrices similar to nonnegative matrices. SIAM Journal on Applied Mathematics, 29 (3):406–408, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ferland:1975:MQC

[1276] Jacques A. Ferland. On the maximum of a quasi-convex quadratic function on a polyhedral convex set. SIAM Journal on Applied Mathematics, 29 (3):409–415, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Beavers:1975:NSM

[1277] A. N. Beavers, Jr. and E. D. Denman. A new solution method for the Lyapunov matrix equation. SIAM Journal on Applied Mathematics, 29 (3):416–421, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burton:1975:PDD

[1278] T. A. Burton. Perturbations and delays in differential equations. SIAM Journal on Applied Mathematics, 29 (3):422–438, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Levine:1975:OAS

[1279] Lawrence E. Levine and Eric S. Lubot. Optimal approximate solutions and time scales. SIAM Journal on Applied Mathematics, 29(3):439–448, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fletcher:1975:MDU

[1280] John E. Fletcher. A model describing the unsteady transport of substrate to tissue from the microcirculation. SIAM Journal on Applied Mathematics, 29 (3):449–480, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1975:GRS

[1281] R. E. Meyer. Gradual reflection of short waves. SIAM Journal on Applied Mathematics, 29(3):481–492, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tucker:1975:CFC

[1282] Alan Tucker. Coloring a family of circular arcs. SIAM Journal on Applied Mathematics, 29(3):493–502, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bird:1975:CCN

[1283] Charles Bird. A class of convex nuclei solution concepts from differences in coalition excesses. SIAM Journal on Applied Mathematics, 29(3): 503–510, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lavoie:1975:PMB

[1284] J. L. Lavoie. The mth power of an n × n matrix and the Bell polynomials. SIAM Journal on Applied Mathematics, 29(3):511–514, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Witsenhausen:1975:DMP

[1285] H. S. Witsenhausen. A determinant maximization problem occurring in the theory of data communication. SIAM Journal on Applied Mathematics, 29 (3):515–522, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abramovich:1975:SPE

[1286] Shoshana Abramovich. Some properties of eigenfunctions. SIAM Journal on Applied Mathematics, 29(3): 523–529, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Greene:1975:SAP

[1287] Curtis Greene and Thomas L. Magnanti. Some abstract pivot algorithms. SIAM Journal on Applied Mathematics, 29(3):530–539, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Siegmann:1975:NMD

[1288] William L. Siegmann and Lester A. Rubenfeld. A nonlinear model for double-diffusive convection. SIAM Journal on Applied Mathematics, 29 (3):540–557, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yang:1975:ASF

[1289] H. T. Yang and L. C. Chien. Analytic solutions of the Falkner–Skan equation when β = −1 and γ = 0. SIAM Journal on Applied Mathematics, 29 (3):558–569, November 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Soukup:1975:MCN

[1290] J. Soukup. On minimum cost networks with nonlinear costs. SIAM Journal on Applied Mathematics, 29 (4):571–581, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemell:1975:DEW

[1291] S. H. Zemell. Diffraction of elastic waves by a rigid-smooth wedge. SIAM Journal on Applied Mathematics, 29 (4):582–596, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ting:1975:SMS

[1292] Lu Ting. Spherical means of solutions of partial differential equations in a conical region. SIAM Journal on Applied Mathematics, 29(4):597–623, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Coppersmith:1975:GCA

[1293] Don Coppersmith and Edna Grossman. Generators for certain alternating groups with applications to cryptography. SIAM Journal on Applied Mathematics, 29(4):624–627, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gilsinn:1975:MAD

[1294] David E. Gilsinn. The method of averaging and domains of stability for integral manifolds. SIAM Journal on Applied Mathematics, 29(4):628–660, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Corio:1975:OEA

[1295] P. L. Corio. Orthonormal expansions of angular momentum functions. SIAM Journal on Applied Mathematics, 29 (4):661–664, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Drange:1975:LBC

[1296] Hans Birger Drange. The linearized Boltzmann collision operator for cutoff potentials. SIAM Journal on Applied Mathematics, 29(4):665–676, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Alam:1975:NSV

[1297] Khursheed Alam. On norms and singular values of rectangular matrices. SIAM Journal on Applied Mathematics, 29(4):677–679, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rheinboldt:1975:PAM

[1298] Werner C. Rheinboldt and James S. Vandergraft. On piecewise affine mappings in R^{n}. SIAM Journal on Applied Mathematics, 29(4):680–689, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garner:1975:OCG

[1299] J. B. Garner. Oscillatory criteria for a general second order, functional differential equation. SIAM Journal on Applied Mathematics, 29(4):690–698, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ross:1975:IEF

[1300] Dieter K. Ross. The interaction energy, field strength and force acting on a pair of dielectric spheres embedded in a dielectric medium. SIAM Journal on Applied Mathematics, 29 (4):699–707, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gochet:1975:CSG

[1301] Willy Gochet and Yves Smeers. Constraint sets of geometric programs characterized by auxiliary problems. SIAM Journal on Applied Mathematics, 29 (4):708–718, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Freedman:1975:PTDb

[1302] H. I. Freedman and P. Waltman. Perturbation of two-dimensional predatorprey equations with an unperturbed critical point. SIAM Journal on Applied Mathematics, 29(4):719–733, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shapiro:1975:DAR

[1303] Jesse M. Shapiro. Domains of attraction for reciprocals of powers of random variables. SIAM Journal on Applied Mathematics, 29(4):734–739, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pasternack:1975:OGS

[1304] Barry Alan Pasternack. Optimal gambling systems under discounting and disbursement. SIAM Journal on Applied Mathematics, 29(4):740–750, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Crane:1975:BTP

[1305] Michael A. Crane. A bounding technique for polynomial functions. SIAM Journal on Applied Mathematics, 29 (4):751–754, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See note [1615].

Witsenhausen:1975:IDC

[1306] H. S. Witsenhausen. An inequality for discrete convex hulls and applications. SIAM Journal on Applied Mathematics, 29(4):755–762, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cole:1975:MDT

[1307] Julian D. Cole. Modern developments in transonic flow. SIAM Journal on Applied Mathematics, 29(4):763–787, December 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dickey:1976:NBC

[1308] R. W. Dickey. Nonlinear bending of circular plates. SIAM Journal on Applied Mathematics, 30(1):1–9, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cambanis:1976:ZCB

[1309] Stamatis Cambanis and Elias Masry.

Zakai’s class of bandlimited functions and processes: Its characterization and properties. SIAM Journal on Applied Mathematics, 30(1):10–21, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Anderssen:1976:CTS

[1310] R. S. Anderssen, R. P. Brent, D. J.

Daley, and P. A. P. Moran. Concerning ∫_{0}^{1} ⋯ ∫_{0}^{1}(x_{1}^{2}+⋯+x_{k}^{2})^{1/2}dx_{1} ⋯ dx_{k} and a Taylor series method. SIAM Journal on Applied Mathematics, 30 (1):22–30, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Greenberg:1976:PWM

[1311] J. M. Greenberg and S. Hastings. Progressive waves in materials exhibiting long range memory. SIAM Journal on Applied Mathematics, 30(1):31–41, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chang:1976:SPB

[1312] K. W. Chang. Singular perturbations of a boundary problem for a vector second order differential equation. SIAM Journal on Applied Mathematics, 30 (1):42–54, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Christofides:1976:OPG

[1313] Nicos Christofides and P. Brooker. The optimal partitioning of graphs. SIAM Journal on Applied Mathematics, 30 (1):55–69, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kelley:1976:PSG

[1314] Walter G. Kelley. Periodic solutions of generalized differential equations. SIAM Journal on Applied Mathematics, 30(1):70–74, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Davis:1976:HEL

[1315] Paul L. Davis. On the hyperbolicity of the equations of the linear theory of heat conduction for materials with memory. SIAM Journal on Applied Mathematics, 30(1):75–80, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Billera:1976:PSC

[1316] Louis J. Billera and Robert E. Bixby. Pareto surfaces of complexity 1. SIAM Journal on Applied Mathematics, 30 (1):81–89, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Allen:1976:MST

[1317] G. D. Allen. On the multiplicity and spectral type of a class of stochastic processes. SIAM Journal on Applied Mathematics, 30(1):90–97, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stanton:1976:NCC

[1318] R. G. Stanton and R. C. Mullin. On the nonexistence of a class of circulant balanced weighing matrices. SIAM Journal on Applied Mathematics, 30 (1):98–102, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Good:1976:RFC

[1319] I. J. Good. The relationship of a formula of Carlitz to the generalized Lagrange expansion. SIAM Journal on Applied Mathematics, 30(1):103, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hwang:1976:SMT

[1320] F. K. Hwang. On Steiner minimal trees with rectilinear distance. SIAM Journal on Applied Mathematics, 30 (1):104–114, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tjostheim:1976:CRW

[1321] D. Tjostheim. A commutation relation for wide sense stationary processes. SIAM Journal on Applied Mathematics, 30(1):115–122, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boa:1976:BLD

[1322] James A. Boa and Donald S. Cohen.

Bifurcation of localized disturbances in a model biochemical reaction. SIAM Journal on Applied Mathematics, 30 (1):123–135, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1976:ME

[1323] Hans-Bernhard Meyer. The matrix equation AZ + B − ZCZ − ZD = 0. SIAM Journal on Applied Mathematics, 30(1):136–142, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maurer:1976:MGS

[1324] Stephen B. Maurer. Matrix generalizations of some theorems on trees, cycles and cocycles in graphs. SIAM Journal on Applied Mathematics, 30 (1):143–148, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Inselberg:1976:MMC

[1325] Alfred Inselberg and Richard S. Chadwick. Mathematical model of the cochlea. I: Formulation and solution. SIAM Journal on Applied Mathematics, 30(1):149–163, January 1976. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chadwick:1976:MMC

[1326] R. S. Chadwick, A. Inselberg, and K. Johnson. Mathematical model of the cochlea. II: Results and conclusions. SIAM Journal on Applied Mathematics, 30(1):164–179, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Olmstead:1976:ASC

[1327] W. E. Olmstead and Richard A. Handelsman. Asymptotic solution to a class of nonlinear Volterra integral equations. II. SIAM Journal on Applied Mathematics, 30(1):180–189, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rao:1976:ATS

[1328] M. R. Rao. Adjacency of the traveling salesman tours and 0-1 vertices. SIAM Journal on Applied Mathematics, 30 (2):191–198, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Greenberg:1976:PSR

[1329] J. M. Greenberg. Periodic solutions to reaction-diffusion equations. SIAM Journal on Applied Mathematics, 30 (2):199–205, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kleitman:1976:DGR

[1330] D. J. Kleitman and D. L. Wang. Decomposition of a graph realizing a degree sequence into disjoint spanning trees. SIAM Journal on Applied Mathematics, 30(2):206–221, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peskoff:1976:MAE

[1331] A. Peskoff, R. S. Eisenberg, and J. D.

Cole. Matched asymptotic expansions of the Green’s function for the electric potential in an infinite cylindrical cell. SIAM Journal on Applied Mathematics, 30(2):222–239, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Richards:1976:SDA

[1332] J. A. Richards. Stability diagram approximation for the lossy Mathieu equation. SIAM Journal on Applied Mathematics, 30(2):240–247, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dickey:1976:STD

[1333] R. W. Dickey. Stability theory for the damped Sine-Gordon equation. SIAM Journal on Applied Mathematics, 30 (2):248–262, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kohler:1976:PRL

[1334] W. Kohler and G. C. Papanicolaou. Power reflection from a lossy onedimensional random medium. SIAM Journal on Applied Mathematics, 30 (2):263–268, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lee:1976:BLS

[1335] Alan J. Lee. On band limited stochastic processes. SIAM Journal on Applied Mathematics, 30(2):269–277, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1976:CBC

[1336] A. H. Zemanian. The complete behavior of certain infinite networks under Kirchhoff’s node and loop laws. SIAM Journal on Applied Mathematics, 30 (2):278–295, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Howes:1976:ECA

[1337] F. A. Howes. Effective characterization of the asymptotic behavior of solutions of singularly perturbed boundary value problems. SIAM Journal on Applied Mathematics, 30(2):296–306, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kounias:1976:BLB

[1338] Stratis Kounias and Jacqueline Marin.

Best linear Bonferroni bounds. SIAM Journal on Applied Mathematics, 30 (2):307–323, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gal:1976:OEF

[1339] Shmuel Gal and Dan Chazan. On the optimality of the exponential functions for some minimax problems. SIAM Journal on Applied Mathematics, 30 (2):324–348, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [1479].

Foch:1976:SPC

[1340] James D. Foch, Jr. Simple proof of the convergence of Shanks’s transformations for meromorphic functions. SIAM Journal on Applied Mathematics, 30 (2):349–351, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Assmus:1976:NFS

[1341] E. F. Assmus, Jr., H. F. Mattson, Jr., and Howard E. Sachar. A new form of the square root bound. SIAM Journal on Applied Mathematics, 30 (2):352–354, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morgan:1976:HOT

[1342] A. J. A. Morgan. Higher order tensors. SIAM Journal on Applied Mathematics, 30(2):355–380, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosenblat:1976:ETW

[1343] S. Rosenblat. An existence theorem for waves in a channel. SIAM Journal on Applied Mathematics, 30(3):381–390, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gordon:1976:SSL

[1344] Paul Gordon. Scaling of a system of linear differential equations. SIAM Journal on Applied Mathematics, 30 (3):391–401, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grinold:1976:LDL

[1345] Richard C. Grinold and Robert E. Stanford. Limiting distributions in a linear fractional flow model. SIAM Journal on Applied Mathematics, 30 (3):402–406, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tortorella:1976:UEM

[1346] Michael Tortorella and James A. Leise. The unitarity equations for matrices and the problem of determining the scattering amplitude from the differential cross section. SIAM Journal on Applied Mathematics, 30(3):407–420, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Galbiati:1976:PES

[1347] G. Galbiati and F. P. Preparata. On permutation-embedding sequences. SIAM Journal on Applied Mathematics, 30(3):421–423, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hills:1976:QRC

[1348] R. N. Hills and R. J. Knops. Qualitative results for a class of general materials. SIAM Journal on Applied Mathematics, 30(3):424–451, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vaid:1976:SEH

[1349] B. K. Vaid and D. L. Jain. Some electrostatic and hydrodynamic problems for two spherical caps. SIAM Journal on Applied Mathematics, 30 (3):452–462, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eliezer:1976:NTD

[1350] C. J. Eliezer and A. Gray. A note on the time-dependent harmonic oscillator. SIAM Journal on Applied Mathematics, 30(3):463–468, May 1976. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cornwell:1976:BVP

[1351] Peter E. Cornwell and David H. Y. Yen.

Boundary value problems in the improved theory of elastic plates. I: Existence of eigenvibrations for plates of arbitrary shape. SIAM Journal on Applied Mathematics, 30(3):469–482, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nahmias:1976:TPP

[1352] Steven Nahmias and William P.

Pierskalla. A two-product perishable/nonperishable inventory problem. SIAM Journal on Applied Mathematics, 30(3):483–500, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pao:1976:EDN

[1353] C. V. Pao. On the energy-dependent neutron transport problem with arbitrary geometry. SIAM Journal on Applied Mathematics, 30(3):501–514, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carrier:1976:ESR

[1354] G. F. Carrier and F. E. Fendell. The effect of strain rate on diffusion flames. II: Large straining. SIAM Journal on Applied Mathematics, 30(3):515–527, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schal:1976:OPD

[1355] Manfred Schal. On the optimality of (s, S)-policies in dynamic inventory models with finite horizon. SIAM Journal on Applied Mathematics, 30(3):528–537, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ben-Tal:1976:PGP

[1356] Aharon Ben-Tal and Adi Ben-Israel. Primal geometric programs treated by linear programming. SIAM Journal on Applied Mathematics, 30(3):538–556, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

MacCamy:1976:STC

[1357] R. C. MacCamy. Stability theorems for a class of functional differential equations. SIAM Journal on Applied Mathematics, 30(3):557–576, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garg:1976:VOM

[1358] M. R. Garg and O. P. Chandna. Viscous orthogonal MHD flows. SIAM Journal on Applied Mathematics, 30 (3):577–585, May 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mortell:1976:ESF

[1359] Michael P. Mortell and Brian R. Seymour. Exact solutions of a functional equation arising in nonlinear wave propagation. SIAM Journal on Applied Mathematics, 30(4):587–596, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Robinson:1976:FOC

[1360] Stephen M. Robinson. First order conditions for general nonlinear optimization. SIAM Journal on Applied Mathematics, 30(4):597–607, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tan:1976:ESM

[1361] W. Y. Tan. Errata: “Some Matrix Results and Extensions of Cochran’s Theory” [SIAM J. Appl. Math. 28 (1975), no. 3, 547–554]. SIAM Journal on Applied Mathematics, 30(4):608–610, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1211].

Davis:1976:DVH

[1362] A. M. J. Davis. The density variations of a heterogeneous ambient flow. SIAM Journal on Applied Mathematics, 30(4):611–616, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hall:1976:FRG

[1363] Frank J. Hall and Robert E. Hartwig. Further results on generalized inverses of partitioned matrices. SIAM Journal on Applied Mathematics, 30(4):617– 624, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Koren:1976:EDR

[1364] Michael Koren. Edge disjoint realizations of two forest-realizable sequences. SIAM Journal on Applied Mathematics, 30(4):625–629, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kalnins:1976:GSV

[1365] E. G. Kalnins, W. Miller, Jr., and P. Winternitz. The group O(4), separation of variables and the hydrogen atom. SIAM Journal on Applied Mathematics, 30(4):630–664, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cushing:1976:FAP

[1366] J. M. Cushing. Forced asymptotically periodic solutions of predator-prey systems with or without hereditary effects. SIAM Journal on Applied Mathematics, 30(4):665–674, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Poore:1976:DBT

[1367] A. B. Poore. On the dynamical behavior of the two-temperature feedback nuclear reactor model. SIAM Journal on Applied Mathematics, 30(4):675–686, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Loewy:1976:RLT

[1368] Raphael Loewy. On ranges of Lyapunov transformations. III. SIAM Journal on Applied Mathematics, 30 (4):687–702, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Campbell:1976:DDI

[1369] Stephen L. Campbell. Differentiation of the Drazin inverse. SIAM Journal on Applied Mathematics, 30(4):703–707, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bell:1976:CGR

[1370] David E. Bell. Constructive group relaxations for integer programs. SIAM Journal on Applied Mathematics, 30 (4):708–719, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1976:FBC

[1371] B. J. Matkowsky and W. L. Siegmann. The flow between counterrotating disks at high Reynolds number. SIAM Journal on Applied Mathematics, 30(4):720–727, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lew:1976:BNM

[1372] Robert A. Lew. Bounds on negative moments. SIAM Journal on Applied Mathematics, 30(4):728–731, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Larsen:1976:EMW

[1373] E. W. Larsen and P. F. Zweifel. Explicit matrix Wiener–Hopf factorizations. SIAM Journal on Applied Mathematics, 30(4):732–738, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Longo:1976:NCT

[1374] Giuseppe Longo. A noiseless coding theorem for sources having utilities. SIAM Journal on Applied Mathematics, 30(4):739–748, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kohn:1976:ISU

[1375] Samuel Kohn. Investigation of the stability of \frac{d^{2}y}{dt^{2}} + w^{2}y = ε( ∑_{i=1}^{n} b_{i} cos 2a_{i}t)y using Shtokalo’s method. SIAM Journal on Applied Mathematics, 30(4):749–767, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Luk:1976:SRC

[1376] Wo shun Luk. Some results concerning the boundedness of solutions of Lienard equations with delay. SIAM Journal on Applied Mathematics, 30 (4):768–774, June 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemell:1976:HFC

[1377] S. H. Zemell. High frequency cylindrical and spherical elastic waves in a heterogeneous half-space. SIAM Journal on Applied Mathematics, 31(1):1–15, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Davis:1976:SWN

[1378] A. M. J. Davis and M. J. Hood. Surface waves normally incident on a submerged horizontal cylinder. SIAM Journal on Applied Mathematics, 31 (1):16–30, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartwig:1976:SVD

[1379] Robert E. Hartwig. Singular value decomposition and the Moore–Penrose inverse of bordered matrices. SIAM Journal on Applied Mathematics, 31 (1):31–41, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartwig:1976:MSF

[1380] R. E. Hartwig. More on the Souriauframe algorithm and the Drazin inverse. SIAM Journal on Applied Mathematics, 31(1):42–46, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Haberman:1976:NGN

[1381] Richard Haberman. Note on generating nonlinear evolution equations. SIAM Journal on Applied Mathematics, 31(1):47–50, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Harrison:1976:DST

[1382] J. Michael Harrison. Dynamic scheduling of a two-class queue: Small interest rates. SIAM Journal on Applied Mathematics, 31(1):51–61, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lax:1976:MML

[1383] Melvin D. Lax. The method of moments for linear random boundary value problems. SIAM Journal on Applied Mathematics, 31(1):62–83, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lovass-Nagy:1976:LSS

[1384] V. Lovass-Nagy and D. L. Powers. On least squares solutions of an inconsistent singular equation AX + XB = C. SIAM Journal on Applied Mathematics, 31(1):84–88, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mangasarian:1976:ECP

[1385] O. L. Mangasarian. Equivalence of the complementarity problem to a system of nonlinear equations. SIAM Journal on Applied Mathematics, 31(1):89–92, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1976:SUF

[1386] J. A. Morrison. The structure of a utility function under strong risk invariance. SIAM Journal on Applied Mathematics, 31(1):93–98, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lang:1976:SDN

[1387] John Patrick Lang and George Kimeldorf. Silent duels with nondiscrete firing. SIAM Journal on Applied Mathematics, 31(1):99–110, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Coleman:1976:PSC

[1388] Bernard D. Coleman and George H.

Renninger. Periodic solutions of certain nonlinear integral equations with a time lag. SIAM Journal on Applied Mathematics, 31(1):111–120, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kaup:1976:PEZ

[1389] D. J. Kaup. A perturbation expansion for the Zakharov–Shabat inverse scattering transform. SIAM Journal on Applied Mathematics, 31(1):121–133, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Becus:1976:RSSa

[1390] Georges A. Becus and F. A. Cozzarelli. The random steady state diffusion problem. I. Random generalized solutions to Laplace’s equation. SIAM Journal on Applied Mathematics, 31(1):134–147, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Becus:1976:RSSb

[1391] Georges A. Becus and F. A. Cozzarelli. The random steady state diffusion problem. II. Random solutions to nonlinear, inhomogeneous, steady state diffusion problems. SIAM Journal on Applied Mathematics, 31(1):148–158, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Becus:1976:RSSc

[1392] Georges A. Becus and F. A. Cozzarelli. The random steady state diffusion problem. III. Solutions to random diffusion problems by the method of random successive approximations. SIAM Journal on Applied Mathematics, 31(1):159–178, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Clements:1976:PFR

[1393] David J. Clements and Brian D. O. Anderson. Polynomial factorization via the Riccati equation. SIAM Journal on Applied Mathematics, 31(1):179–205, July 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Spielvogel:1976:ESF

[1394] L. Q. Spielvogel and E. R. Spielvogel. Existence of steady free surface gravity flows. SIAM Journal on Applied Mathematics, 31(2):207–232, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1976:CSC

[1395] J. A. Morrison and J. A. Lewis. Charge singularity at the corner of a flat plate. SIAM Journal on Applied Mathematics, 31(2):233–250, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [1574].

Cushing:1976:PSV

[1396] J. M. Cushing. Periodic solutions of Volterra’s population equation with hereditary effects. SIAM Journal on Applied Mathematics, 31(2):251–261, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [1512].

Delsarte:1976:PAR

[1397] Philippe Delsarte. Properties and applications of the recurrence F (i + 1, k + 1, n + 1) = q^{k+1F} (i, k + 1, n) − q^{k}F (i, k, n). SIAM Journal on Applied Mathematics, 31(2):262–270, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Christofides:1976:SAL

[1398] Nicos Christofides and Mary Trypia. A sequential approach to the 0-1 linear programming problem. SIAM Journal on Applied Mathematics, 31(2): 271–285, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chikwendu:1976:ASS

[1399] S. C. Chikwendu. Asymptotic solutions of some weakly nonlinear elliptic equations. SIAM Journal on Applied Mathematics, 31(2):286–303, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Walton:1976:NCA

[1400] J. R. Walton. A note on certain asymptotic expressions for the unitstep and Dirac delta functions. SIAM Journal on Applied Mathematics, 31 (2):304–306, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berman:1976:PSR

[1401] Abraham Berman and Michael Neumann. Proper splittings of rectangular matrices. SIAM Journal on Applied Mathematics, 31(2):307–312, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Witsenhausen:1976:VBC

[1402] H. S. Witsenhausen. Values and bounds for the common information of two discrete random variables. SIAM Journal on Applied Mathematics, 31 (2):313–333, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Obi:1976:ATNa

[1403] Chike Obi. Analytical theory of nonlinear oscillations. II. Small periodic oscillations of equations of the second order. SIAM Journal on Applied Mathematics, 31(2):334–344, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Obi:1976:ATNb

[1404] Chike Obi. Analytical theory of nonlinear oscillations. IV: The periodic oscillations of the equation x − ε(1 − x^{2n+2})x + x^{2n+1} = εa cos ωt, a > 0, ω > 0 independent of ε. SIAM Journal on Applied Mathematics, 31 (2):345–357, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Charnes:1976:CCC

[1405] A. Charnes and Daniel Granot. Coalitional and chance-constrained solutions to n-person games. I. The prior satisficing nucleolus. SIAM Journal on Applied Mathematics, 31(2):358–367, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Glassey:1976:OHT

[1406] C. R. Glassey and R. M. Karp. On the optimality of Huffman trees. SIAM Journal on Applied Mathematics, 31 (2):368–378, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Watson:1976:SPT

[1407] Layne T. Watson. Some perturbation theorems for Q-matrices. SIAM Journal on Applied Mathematics, 31 (2):379–384, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [1585].

Wheeler:1976:IBL

[1408] Lewis Wheeler and Cornelius O. Horgan. Isoperimetric bounds on the lowest nonzero Stekloff eigenvalue for plane strip domains. SIAM Journal on Applied Mathematics, 31(2): 385–391, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hochstadt:1976:IPH

[1409] Harry Hochstadt. An inverse problem for a Hill’s equation. SIAM Journal on Applied Mathematics, 31(2): 392–396, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maybee:1976:SAT

[1410] John S. Maybee. Some aspects of the theory of P N-matrices. SIAM Journal on Applied Mathematics, 31 (2):397–410, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Campbell:1976:ADI

[1411] Stephen L. Campbell, Carl D. Meyer, Jr., and Nicholas J. Rose. Applications of the Drazin inverse to linear systems of differential equations with singular constant coefficients. SIAM Journal on Applied Mathematics, 31 (3):411–425, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mentz:1976:ISC

[1412] Raul Pedro Mentz. On the inverse of some covariance matrices of Toeplitz type. SIAM Journal on Applied Mathematics, 31(3):426–437, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berger:1976:DAL

[1413] Neil Berger. Derivation of approximate long wave equations in a nearly uniform channel of approximately rectangular cross section. SIAM Journal on Applied Mathematics, 31(3): 438–448, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rayner:1976:SAS

[1414] A. A. Rayner and R. M. Pringle. Some aspects of the solution of singular normal equations with the use of linear restrictions. SIAM Journal on Applied Mathematics, 31(3):449–460, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [2450].

Tefteller:1976:OSO

[1415] S. C. Tefteller. Oscillation of second order nonhomogeneous linear differential equations. SIAM Journal on Applied Mathematics, 31(3):461–467, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brown:1976:CUM

[1416] David P. Brown. Circuit and unimodular matrices. SIAM Journal on Applied Mathematics, 31(3):468–473, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bird:1976:ENI

[1417] Charles Bird. Extending the nucleolus to infinite player games. SIAM Journal on Applied Mathematics, 31 (3):474–484, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keller:1976:MRT

[1418] Joseph B. Keller. The minimum ratio of two eigenvalues. SIAM Journal on Applied Mathematics, 31(3): 485–491, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Campbell:1976:DII

[1419] Stephen L. Campbell. The Drazin inverse of an infinite matrix. SIAM Journal on Applied Mathematics, 31 (3):492–503, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Utz:1976:PSS

[1420] W. R. Utz. Periodic solutions of second order differential equations with nonlinear, nondifferentiable damping.

SIAM Journal on Applied Mathematics, 31(3):504–510, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fettis:1976:FSE

[1421] Henry E. Fettis. Fourier series expansions for Pearson type IV distributions and probabilities. SIAM Journal on Applied Mathematics, 31(3): 511–518, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kaper:1976:PNO

[1422] B. Kaper. Perturbed nonlinear oscillations. SIAM Journal on Applied Mathematics, 31(3):519–546, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bushard:1976:IVR

[1423] Louis B. Bushard. Initial value region for a boundary value problem arising in gas glow discharge theory. SIAM Journal on Applied Mathematics, 31 (3):547–557, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cambanis:1976:RBD

[1424] Stamatis Cambanis. The relationship between the detection of sure and stochastic signals in noise. SIAM Journal on Applied Mathematics, 31(3): 558–568, November 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Warren:1976:SMN

[1425] P. Warren, J. Foster, and N. Bleistein. A stochastic model of a nonhomogeneous carrier-borne epidemic. SIAM Journal on Applied Mathematics, 31 (4):569–578, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Khatri:1976:HND

[1426] C. G. Khatri and Sujit Kumar Mitra.

Hermitian and nonnegative definite solutions of linear matrix equations. SIAM Journal on Applied Mathematics, 31(4):579–585, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Frank:1976:HPS

[1427] B. Frank. Hermite polynomial solutions of an infinite set of Green function equations. SIAM Journal on Applied Mathematics, 31(4):586–590, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Belleni-Morante:1976:NTS

[1428] A. Belleni-Morante and R. Farano.

Neutron transport in a slab with moving boundaries. SIAM Journal on Applied Mathematics, 31(4):591–599, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kojima:1976:ELM

[1429] Masakazu Kojima, Hisakazu Nishino, and Tomoharu Sekine. An extension of Lemke’s method to the piecewise linear complementarity problem. SIAM Journal on Applied Mathematics, 31 (4):600–613, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Blair:1976:TRD

[1430] Charles Blair. Two rules for deducing valid inequalities for 0-1 problems. SIAM Journal on Applied Mathematics, 31(4):614–617, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goldman:1976:RTI

[1431] Jay R. Goldman, J. T. Joichi, David L. Reiner, and Dennis E. White. Rook theory. II: Boards of binomial type. SIAM Journal on Applied Mathematics, 31(4):618–633, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sansing:1976:SDE

[1432] R. C. Sansing. The t-statistic for a double exponential distribution. SIAM Journal on Applied Mathematics, 31 (4):634–645, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rothblum:1976:RDI

[1433] Uriel G. Rothblum. A representation of the Drazin inverse and characterizations of the index. SIAM Journal on Applied Mathematics, 31(4): 646–648, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mallows:1976:SDC

[1434] C. L. Mallows, V. Pless, and N. J. A. Sloane. Self-dual codes over GF(3). SIAM Journal on Applied Mathematics, 31(4):649–666, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grasman:1976:ROG

[1435] J. Grasman, E. J. M. Veling, and G. M. Willems. Relaxation oscillations governed by a Van Der Pol equation with periodic forcing term. SIAM Journal on Applied Mathematics, 31 (4):667–676, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ahner:1976:TDE

[1436] John F. Ahner and George C.

Hsiao. On the two-dimensional exterior boundary-value problems of elasticity. SIAM Journal on Applied Mathematics, 31(4):677–685, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chandna:1976:SPF

[1437] O. P. Chandna and P. N. Kaloni.

Steady plane flows of Cosserat fluids. SIAM Journal on Applied Mathematics, 31(4):686–698, December 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Katz:1977:LTE

[1438] H. Katz. A large time expansion for the Stefan problem. SIAM Journal on Applied Mathematics, 32(1):1–20, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rubenfeld:1977:UAS

[1439] L. A. Rubenfeld and B. Willner. Uniform asymptotic solutions for linear second order ordinary differential equations with turning points: Formal theory. SIAM Journal on Applied Mathematics, 32(1):21–38, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Thompson:1977:SVD

[1440] R. C. Thompson. Singular values, diagonal elements, and convexity. SIAM Journal on Applied Mathematics, 32 (1):39–63, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Benson:1977:VMP

[1441] H. P. Benson and T. L. Morin. The vector maximization problem: Proper efficiency and stability. SIAM Journal on Applied Mathematics, 32(1):64– 72, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hoppensteadt:1977:PSL

[1442] F. C. Hoppensteadt and J. M. Hyman. Periodic solutions of a logistic difference equation. SIAM Journal on Applied Mathematics, 32(1):73–81, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cushing:1977:PTD

[1443] J. M. Cushing. Periodic timedependent predator-prey systems. SIAM Journal on Applied Mathematics, 32(1):82–95, January 1977. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Macken:1977:DMS

[1444] Catherine A. Macken and Howard M. Taylor. On deviations from the maximum in a stochastic process. SIAM Journal on Applied Mathematics, 32 (1):96–104, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pursiheimo:1977:OST

[1445] Ulla Pursiheimo. On the optimal search for a target whose motion is conditionally deterministic with stochastic initial conditions on location and parameters. SIAM Journal on Applied Mathematics, 32(1):105–114, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rath:1977:OPC

[1446] John H. Rath. The optimal policy for a controlled Brownian motion process. SIAM Journal on Applied Mathematics, 32(1):115–125, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lukka:1977:OST

[1447] Markku Lukka. On the optimal searching tracks for a moving target. SIAM Journal on Applied Mathematics, 32 (1):126–132, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Henrard:1977:ABS

[1448] Jacques Henrard and Kenneth R.

Meyer. Averaging and bifurcations in symmetric systems. SIAM Journal on Applied Mathematics, 32(1): 133–145, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Troy:1977:OTO

[1449] William C. Troy. Oscillations in a third order differential equation modeling a nuclear reactor. SIAM Journal on Applied Mathematics, 32(1):146–153, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [1677].

Haberman:1977:SBS

[1450] Richard Haberman. On the singular behavior of spatially dependent nonlinear wave envelopes. SIAM Journal on Applied Mathematics, 32(1):154– 163, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pao:1977:PSC

[1451] C. V. Pao. Positive solutions and criticality of the linear and some nonlinear neutron transport problems. SIAM Journal on Applied Mathematics, 32 (1):164–176, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yuen:1977:ACC

[1452] C. K. Yuen. The auto- and crosscorrelations of Walsh functions. SIAM Journal on Applied Mathematics, 32 (1):177–190, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kaper:1977:ABS

[1453] Hans G. Kaper and R. Bruce Kellogg. Asymptotic behavior of the solution of the integral transport equation in slab geometry. SIAM Journal on Applied Mathematics, 32(1):191–200, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kaper:1977:CDP

[1454] Hans G. Kaper and R. Bruce Kellogg. Continuity and differentiability properties of the solution of the linear transport equation. SIAM Journal on Applied Mathematics, 32(1):201–214, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lambe:1977:UBN

[1455] Thomas A. Lambe. Upper bound on the number of nonnegative integer solutions to a linear equation. SIAM Journal on Applied Mathematics, 32 (1):215–219, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jones:1977:NTS

[1456] Lee Jones. A note on the traveling salesman problem. SIAM Journal on Applied Mathematics, 32(1): 220–222, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boudjelkha:1977:UIR

[1457] M. T. Boudjelkha and J. B. Diaz.

Uniform integral representation of the Macdonald function of positive integer order. SIAM Journal on Applied Mathematics, 32(1):223–231, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gribik:1977:ETC

[1458] P. R. Gribik and K. O. Kortanek.

Equivalence theorems and cutting plane algorithms for a class of experimental design problems. SIAM Journal on Applied Mathematics, 32 (1):232–259, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lin:1977:LCR

[1459] Juan Lin and Peter B. Kahn. Limit cycles in random environments. SIAM Journal on Applied Mathematics, 32 (1):260–291, January 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Whittemore:1977:OIU

[1460] A. S. Whittemore and S. C. Saunders. Optimal inventory under stochastic demand with two supply options. SIAM Journal on Applied Mathematics, 32 (2):293–305, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Troy:1977:ABD

[1461] William C. Troy and Richard J. Field. The amplification before decay of perturbations around stable states in a model of the zhabotinskii reaction. SIAM Journal on Applied Mathematics, 32(2):306–322, March 1977. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gaiha:1977:AVP

[1462] P. Gaiha and S. K. Gupta. Adjacent vertices on a permutohedron. SIAM Journal on Applied Mathematics, 32 (2):323–327, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wu:1977:DTC

[1463] Lilian Shiao-Yen Wu. A dynamic theory for the class of games with nonempty cores. SIAM Journal on Applied Mathematics, 32(2):328–338, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Crain:1977:ITA

[1464] Bradford R. Crain. An information theoretic approach to approximating a probability distribution. SIAM Journal on Applied Mathematics, 32 (2):339–346, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abbott:1977:SPC

[1465] H. L. Abbott and P. F. Dierker. Snakes in powers of complete graphs. SIAM Journal on Applied Mathematics, 32 (2):347–355, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Golomb:1977:ICF

[1466] S. W. Golomb, I. S. Reed, and T. K.

Truong. Integer convolutions over the finite field GF (3 · 2^{n} + 1). SIAM Journal on Applied Mathematics, 32 (2):356–365, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hsu:1977:MTS

[1467] S. B. Hsu, S. Hubbell, and P. Waltman. A mathematical theory for single-nutrient competition in continuous cultures of micro-organisms. SIAM Journal on Applied Mathematics, 32 (2):366–383, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosenthal:1977:CRC

[1468] Arnie Rosenthal. Computing the reliability of complex networks. SIAM Journal on Applied Mathematics, 32 (2):384–393, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mahar:1977:MBR

[1469] T. J. Mahar and B. J. Matkowsky. A model biochemical reaction exhibiting secondary bifurcation. SIAM Journal on Applied Mathematics, 32(2):394– 404, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sharpe:1977:IPA

[1470] Charles B. Sharpe. An inverse problem associated with the discrete Schrödinger equation. SIAM Journal on Applied Mathematics, 32(2):405– 417, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goodman:1977:DRH

[1471] S. Goodman and D. Shier. On designing a reliable hierarchical structure. SIAM Journal on Applied Mathematics, 32(2):418–430, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Danaraj:1977:CGC

[1472] Gopal Danaraj and Victor Klee. The connectedness game and the ccomplexity of certain graphs. SIAM Journal on Applied Mathematics, 32 (2):431–442, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brylawski:1977:DIR

[1473] Tom Brylawski. A determinantal identity for resistive networks. SIAM Journal on Applied Mathematics, 32 (2):443–449, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Truemper:1977:MFG

[1474] K. Truemper. On max flows with gains and pure min-cost flows. SIAM Journal on Applied Mathematics, 32 (2):450–456, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kreps:1977:OSP

[1475] David M. Kreps and Evan L. Porteus.

On the optimality of structured policies in countable stage decision processes. II: Positive and negative problems. SIAM Journal on Applied Mathematics, 32(2):457–466, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Padgett:1977:RDE

[1476] W. J. Padgett, G. Schultz, and Chris P.

Tsokos. A random differential equation approach to the probability distribution of BOD and DO in streams. SIAM Journal on Applied Mathematics, 32 (2):467–483, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Swick:1977:NAD

[1477] K. E. Swick. A nonlinear agedependent model of single species population dynamics. SIAM Journal on Applied Mathematics, 32(2):484–498, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grenander:1977:SAN

[1478] Ulf Grenander and Jack W. Silverstein.

Spectral analysis of networks with random topologies. SIAM Journal on Applied Mathematics, 32(2):499–519, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gal:1977:EOE

[1479] Shmuel Gal and Dan Chazan. Errata: “On the Optimality of the Exponential Functions for Some Minimax Problems” [SIAM J. Appl. Math. 30 (1976), no. 2, 324–348]. SIAM Journal on Applied Mathematics, 32(2):520, March 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1339].

Kohler:1977:PRI

[1480] Werner Kohler. Power reflection at the input of a randomly perturbed rectangular waveguide. SIAM Journal on Applied Mathematics, 32(3):521–533, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brown:1977:MSE

[1481] Harold Brown. Molecular structure elucidation. III. SIAM Journal on Applied Mathematics, 32(3):534–551, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

VanSchuppen:1977:FPS

[1482] J. H. Van Schuppen. Filtering, prediction and smoothing for counting process observations, a martingale approach. SIAM Journal on Applied Mathematics, 32(3):552–570, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pack:1977:ODM

[1483] Charles D. Pack. The output of a D/M/1 queue. SIAM Journal on Applied Mathematics, 32(3):571–587, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grasman:1977:VAS

[1484] J. Grasman and B. J. Matkowsky. A variational approach to singularly perturbed boundary value problems for ordinary and partial differential equations with turning points. SIAM Journal on Applied Mathematics, 32 (3):588–597, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schmeichel:1977:PGD

[1485] E. F. Schmeichel and S. L. Hakimi.

On planar graphical degree sequences. SIAM Journal on Applied Mathematics, 32(3):598–609, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lardy:1977:SIM

[1486] L. J. Lardy. Some iterative methods for linear operator equations with applications to generalized inverses. SIAM Journal on Applied Mathematics, 32(3):610–618, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kleitman:1977:ESE

[1487] D. J. Kleitman, M. Koren, and S.- Y. R. Li. On the existence of simultaneous edge disjoint realizations of degree sequences with “few” edges. SIAM Journal on Applied Mathematics, 32(3):619–626, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peleg:1977:CPO

[1488] Bezalel Peleg and Itzhak Zilcha. On competitive prices for optimal consumptions plans, II. SIAM Journal on Applied Mathematics, 32(3):627–630, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Freedman:1977:MMP

[1489] H. I. Freedman and Paul Waltman. Mathematical models of population interactions with dispersal. I: Stability of two habitats with an without a predator. SIAM Journal on Applied Mathematics, 32(3):631–648, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kassam:1977:NWS

[1490] Saleem A. Kassam and John B. Thomas. Nonparametric weightedsigns tests for location. SIAM Journal on Applied Mathematics, 32(3):649– 652, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Segel:1977:TSR

[1491] Lee A. Segel. A theoretical study of receptor mechanisms in bacterial chemotaxis. SIAM Journal on Applied Mathematics, 32(3):653–665, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

White:1977:ERF

[1492] Benjamin S. White. The effects of a rapidly-fluctuating random environment on systems of interacting species. SIAM Journal on Applied Mathematics, 32(3):666–693, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lakin:1977:TSP

[1493] W. D. Lakin and P. Van Den Driessche. Time scales in population biology. SIAM Journal on Applied Mathematics, 32(3):694–705, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moon:1977:EDC

[1494] J. W. Moon. Erratum: “On the Distribution of Crossings in Random Complete Graphs” [J. SIAM 13 (1965), no. 2, 506–510]. SIAM Journal on Applied Mathematics, 32(3):706, May 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [3].

Flanders:1977:ME

[1495] Harley Flanders and Harald K. Wimmer. On the matrix equations AX − XB = C and AX − Y B = C. SIAM Journal on Applied Mathematics, 32(4):707–710, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hammer:1977:RMP

[1496] Peter L. Hammer, Ellis L. Johnson, and Uri N. Peled. The role of master polytopes in the unit cube. SIAM Journal on Applied Mathematics, 32 (4):711–716, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stolarsky:1977:PES

[1497] Kenneth B. Stolarsky. Power and exponential sums of digital sums related to binomial coefficient parity.

SIAM Journal on Applied Mathematics, 32(4):717–730, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lee:1977:AIS

[1498] Alan J. Lee. Approximate interpolation and the sampling theorem. SIAM Journal on Applied Mathematics, 32 (4):731–744, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pakes:1977:LDT

[1499] A. G. Pakes and T. P. Speed. Lagrange distributions and their limit theorems. SIAM Journal on Applied Mathematics, 32(4):745–754, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jenkins:1977:EME

[1500] James T. Jenkins. The equations of mechanical equilibrium of a model membrane. SIAM Journal on Applied Mathematics, 32(4):755–764, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bowden:1977:CMT

[1501] Robert L. Bowden, William Greenberg, and P. F. Zweifel. Critical multigroup transport. SIAM Journal on Applied Mathematics, 32(4):765–777, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gladstien:1977:HPS

[1502] Keith Gladstien. Haploid populations subject to varying environment: The characteristic values and the rate of loss of alleles. SIAM Journal on Applied Mathematics, 32(4):778–783, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1977:IMD

[1503] Jack K. Cohen and Norman Bleistein.

An inverse method for determining small variations in propagation speed. SIAM Journal on Applied Mathematics, 32(4):784–799, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hinich:1977:ADM

[1504] Melvin Hinich. Array design for measuring source depth in a horizontal waveguide. SIAM Journal on Applied Mathematics, 32(4):800–806, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Griggs:1977:SCS

[1505] Jerrold R. Griggs. Sufficient conditions for a symmetric chain order. SIAM Journal on Applied Mathematics, 32(4):807–809, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Conolly:1977:TDF

[1506] Brian Conolly and I. J. Good. A table of discrete Fourier transform pairs. SIAM Journal on Applied Mathematics, 32(4):810–822, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [1546].

Mitra:1977:ME

[1507] Sujit Kumar Mitra. The matrix equation AXB + CXD = E. SIAM Journal on Applied Mathematics, 32 (4):823–825, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garey:1977:RST

[1508] M. R. Garey and D. S. Johnson. The rectilinear Steiner tree problem is NP -complete. SIAM Journal on Applied Mathematics, 32(4):826–834, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garey:1977:CCS

[1509] M. R. Garey, R. L. Graham, and D. S. Johnson. The complexity of computing Steiner minimal trees. SIAM Journal on Applied Mathematics, 32 (4):835–859, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:1977:DIP

[1510] Da-Lun Wang and Ping Wang. Discrete isoperimetric problems. SIAM Journal on Applied Mathematics, 32 (4):860–870, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rubenfeld:1977:NDT

[1511] L. A. Rubenfeld and W. L. Siegmann. Nonlinear dynamic theory for a double-diffusive convection model. SIAM Journal on Applied Mathematics, 32(4):871–894, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cushing:1977:EPS

[1512] J. M. Cushing. Errata: “Periodic Solutions of Volterra’s Population Equation with Hereditary Effects” [SIAM J. Appl. Math. 31 (1976), no. 2, 251– 261]. SIAM Journal on Applied Mathematics, 32(4):895, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1396].

Meyer:1977:IDI

[1513] Carl D. Meyer, Jr. and Nicholas J.

Rose. The index and the Drazin inverse of block triangular matrices. SIAM Journal on Applied Mathematics, 33(1):1–7, July 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pechuzal:1977:STF

[1514] G. Pechuzal and J. Kevorkian.

Supersonic-transonic flow generated by a thin airfoil in a stratified atmosphere. SIAM Journal on Applied Mathematics, 33(1):8–33, July 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ward:1977:LFW

[1515] James F. Ward, Jr. On a limit formula for weighted pseudoinverses. SIAM Journal on Applied Mathematics, 33(1):34–38, July 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1977:PSV

[1516] Ronald Smith. Propagation in slowlyvarying wave-guides. SIAM Journal on Applied Mathematics, 33(1):39–50, July 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tourlakis:1977:HMC

[1517] George Tourlakis. Homological methods for the classification of discrete Euclidean structures. SIAM Journal on Applied Mathematics, 33(1):51–54, July 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:1977:ECD

[1518] Da-Lun Wang and Ping Wang. Extremal configurations on a discrete torus and a generalization of the generalized Macaulay theorem. SIAM Journal on Applied Mathematics, 33(1):55– 59, July 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bazer:1977:MSO

[1519] J. Bazer. Multiple scattering in one dimension. II. SIAM Journal on Applied Mathematics, 33(1):60–94, July 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Amari:1977:MFS

[1520] Shun-Ichi Amari, Kiyonori Yoshida, and Ken-Ichi Kanatani. A mathematical foundation for statistical neurodynamics. SIAM Journal on Applied Mathematics, 33(1):95–126, July 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1977:CSV

[1521] J. A. Morrison. Charge singularity at the vertex of a slender cone of general cross-section. SIAM Journal on Applied Mathematics, 33(1):127–132, July 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Newell:1977:FAI

[1522] Alan C. Newell. Finite amplitude instabilities of partial difference equations. SIAM Journal on Applied Mathematics, 33(1):133–160, July 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Melvin:1977:CPL

[1523] Peter J. Melvin. On the construction of Poincaré–Lindstedt solutions: the nonlinear oscillator equation. SIAM Journal on Applied Mathematics, 33 (1):161–194, July 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Palm:1977:VRW

[1524] G. Palm and T. Poggio. The Volterra representation and the Wiener expansion: Validity and pitfalls. SIAM Journal on Applied Mathematics, 33 (2):195–216, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1977:SMO

[1525] Donald S. Cohen, Frank C. Hoppensteadt, and Robert M. Miura. Slowly modulated oscillations in nonlinear diffusion processes. SIAM Journal on Applied Mathematics, 33(2):217–229, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1977:SPB

[1526] Bernard J. Matkowsky and Edward L.

Reiss. Singular perturbations of bifurcations. SIAM Journal on Applied Mathematics, 33(2):230–255, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mangel:1977:PES

[1527] Marc Mangel and Donald Ludwig. Probability of extinction in a stochastic competition. SIAM Journal on Applied Mathematics, 33(2):256–266, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Franklin:1977:SSP

[1528] Joel N. Franklin. Some stereological principles in morphometric cytology. SIAM Journal on Applied Mathematics, 33(2):267–278, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Charalambides:1977:NKN

[1529] Ch. A. Charalambides. A new kind of numbers appearing in the n-fold convolution of truncated binomial and negative binomial distributions. SIAM Journal on Applied Mathematics, 33 (2):279–288, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fowler:1977:CDE

[1530] A. C. Fowler. Convective diffusion of an enzyme reaction. SIAM Journal on Applied Mathematics, 33(2): 289–297, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pimbley:1977:SHO

[1531] George H. Pimbley, Jr. Superlinear Hammerstein operators and the nonoccurrence of secondary bifurcation. SIAM Journal on Applied Mathematics, 33(2):298–322, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bauer:1977:FAC

[1532] W. R. Bauer and W. A. Blankinship.

A fast algorithm for certain dynamic programming minimizations. SIAM Journal on Applied Mathematics, 33 (2):323–336, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Joseph:1977:CBS

[1533] Daniel D. Joseph. The convergence of biorthogonal series for biharmonic and Stokes flow edge problems. part I. SIAM Journal on Applied Mathematics, 33(2):337–347, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bigelow:1977:STA

[1534] James H. Bigelow and Norman Z.

Shapiro. A scaling theorem and algorithms. SIAM Journal on Applied Mathematics, 33(2):348–352, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Casten:1977:SPS

[1535] Richard G. Casten and Charles J. Holland. Stability properties of solutions to systems of reaction-diffusion equations. SIAM Journal on Applied Mathematics, 33(2):353–364, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1977:EPR

[1536] B. J. Matkowsky and Z. Schuss. The exit problem for randomly perturbed dynamical systems. SIAM Journal on Applied Mathematics, 33(2):365–382, September 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Picci:1977:SCB

[1537] Giorgio Picci. Some connections between the theory of sufficient statistics and the identifiability problem. SIAM Journal on Applied Mathematics, 33 (3):383–398, November 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carasso:1977:VFF

[1538] Alfred Carasso and Mei-Chang Shen. On viscous fluid flow down an inclined plane and the development of roll waves. SIAM Journal on Applied Mathematics, 33(3):399–426, November 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Larsen:1977:ATL

[1539] Edward W. Larsen. Asymptotic theory of the linear transport equation for small mean free paths. II. SIAM Journal on Applied Mathematics, 33 (3):427–445, November 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Payne:1977:SAM

[1540] L. E. Payne and G. A. Philippin. Some applications of the maximum principle in the problem of torsional creep. SIAM Journal on Applied Mathematics, 33 (3):446–455, November 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stone:1977:STG

[1541] Lawrence D. Stone. Search for targets with generalized conditionally deterministic motion. SIAM Journal on Applied Mathematics, 33(3):456–468, November 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fishburn:1977:CSC

[1542] Peter C. Fishburn. Condorcet social choice functions. SIAM Journal on Applied Mathematics, 33(3): 469–489, November 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Swick:1977:HRA

[1543] D. A. Swick. Harmonic relationships among random variables. SIAM Journal on Applied Mathematics, 33(3): 490–498, November 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zeppetella:1977:AAB

[1544] Anthony Zeppetella. Analysis of the adapting bump model of the horseshoe crab eye. SIAM Journal on Applied Mathematics, 33(3):499–510, November 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Herman:1977:FIR

[1545] Gabor T. Herman and Abraham Naparstek. Fast image reconstruction based on a Radon inversion formula appropriate for rapidly collected data.

SIAM Journal on Applied Mathematics, 33(3):511–533, November 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Conolly:1977:ETD

[1546] Brian Conolly and I. J. Good. Errata: “A Table of Discrete Fourier Transform Pairs” [SIAM J. Appl. Math. 32 (1977), no. 4, 810–822]. SIAM Journal on Applied Mathematics, 33 (3):534, November 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1506].

Griffel:1977:IVP

[1547] D. H. Griffel and C. A. N. Morris. An initial-value problem for waves over a sloping beach. SIAM Journal on Applied Mathematics, 33(4):535–541, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carter:1977:DPQ

[1548] Bradley D. Carter and Melvin D. Springer. The distribution of products, quotients and powers of independent H-function variates. SIAM Journal on Applied Mathematics, 33 (4):542–558, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gray:1977:SOM

[1549] Robert M. Gray and Lee D. Davisson. Second order moments and prediction for doubly-reflected symmetric independent increment processes. SIAM Journal on Applied Mathematics, 33 (4):559–567, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Petty:1977:CAP

[1550] J. V. Petty. A containment algorithm for point sets determined by polygonal paths. SIAM Journal on Applied Mathematics, 33(4):568–574, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kapadia:1977:RPR

[1551] Asha S. Kapadia, Bartholomew P. Hsi, and Harvey Diamond. Renewal processes with restriction on success runs. SIAM Journal on Applied Mathematics, 33(4):575–580, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Levine:1977:MEU

[1552] Lawrence E. Levine and Wilson C.

Obi. Multivariable expansions using time scales generated from differential equations. SIAM Journal on Applied Mathematics, 33(4):581–586, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Golomb:1977:SOP

[1553] S. W. Golomb and A. Lempel. Second order polynomial recursions. SIAM Journal on Applied Mathematics, 33 (4):587–592, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ziv:1977:NSS

[1554] Abraham Ziv and Vardy Amdursky.

On the numerical solution of stiff linear systems of the oscillatory type. SIAM Journal on Applied Mathematics, 33 (4):593–606, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Balinski:1977:HMA

[1555] M. L. Balinski and H. P. Young. On huntington methods of apportionment. SIAM Journal on Applied Mathematics, 33(4):607–618, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1977:PNM

[1556] A. H. Zemanian. Proportioning networks: A model for a two-level marketing system in an underdeveloped economy. SIAM Journal on Applied Mathematics, 33(4):619–639, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cushing:1977:BPS

[1557] J. M. Cushing. Bifurcation of periodic solutions of integrodifferential systems with applications to time delay models in population dynamics. SIAM Journal on Applied Mathematics, 33 (4):640–654, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bitmead:1977:MCI

[1558] Robert R. Bitmead and Brian D. O. Anderson. The matrix Cauchy index: Properties and applications. SIAM Journal on Applied Mathematics, 33 (4):655–672, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Conway:1977:DPP

[1559] Edward D. Conway and Joel A. Smoller. Diffusion and the predatorprey interaction. SIAM Journal on Applied Mathematics, 33(4):673–686, December 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jones:1978:DSV

[1560] Lee Jones. On the distribution of sums of vectors. SIAM Journal on Applied Mathematics, 34(1):1–6, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Joseph:1978:CBS

[1561] Daniel D. Joseph and Leroy Sturges.

The convergence of biorthogonal series for biharmonic and Stokes flow edge problems: Part II. SIAM Journal on Applied Mathematics, 34(1):7–26, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chung:1978:LBS

[1562] F. R. K. Chung and F. K. Hwang. A lower bound for the Steiner tree problem. SIAM Journal on Applied Mathematics, 34(1):27–36, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kaup:1978:GCP

[1563] David J. Kaup and Alan C. Newell.

The Goursat and Cauchy problems for the Sine-Gordon equation. SIAM Journal on Applied Mathematics, 34 (1):37–54, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Janowitz:1978:OTM

[1564] M. F. Janowitz. An order theoretic model for cluster analysis. SIAM Journal on Applied Mathematics, 34 (1):55–72, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mustafi:1978:ADE

[1565] Chandan K. Mustafi. On the asymptotic distribution of the estimates of the derivatives of a distribution function. SIAM Journal on Applied Mathematics, 34(1):73–77, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bazer:1978:EPR

[1566] J. Bazer and R. Burridge. Energy partition in the reflection and refraction of plane waves. SIAM Journal on Applied Mathematics, 34(1):78–92, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Larson:1978:TBT

[1567] D. A. Larson. Transient bounds and time-asymptotic behavior of solutions to nonlinear equations of Fisher type. SIAM Journal on Applied Mathematics, 34(1):93–103, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Udwadia:1978:SUR

[1568] F. E. Udwadia and D. K. Sharma. Some uniqueness results related to building structural identification. SIAM Journal on Applied Mathematics, 34(1):104–118, January 1978. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Balas:1978:FKP

[1569] Egon Balas and Eitan Zemel. Facets of the knapsack polytope from minimal covers. SIAM Journal on Applied Mathematics, 34(1):119–148, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kerr:1978:NIS

[1570] Arnold D. Kerr. A new iterative scheme for the solution of partial differential equations based on the principle of least squares. SIAM Journal on Applied Mathematics, 34(1):149–156, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Delsarte:1978:HPD

[1571] Ph. Delsarte. Hahn polynomials, discrete harmonics, and t-designs. SIAM Journal on Applied Mathematics, 34 (1):157–166, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wan:1978:CSC

[1572] Yieh-Hei Wan. Computation of the stability condition for the Hopf bifurcation of difeomorphisms on R^{2}. SIAM Journal on Applied Mathematics, 34 (1):167–175, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rose:1978:AAV

[1573] Donald J. Rose and Robert Endre Tarjan. Algorithmic aspects of vertex elimination on directed graphs. SIAM Journal on Applied Mathematics, 34 (1):176–197, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Taylor:1978:NPJ

[1574] R. S. Taylor. Note on a paper by J. A. Morrison and J. A. Lewis: “Charge Singularity at the Corner of a Flat Plate” [SIAM J. Appl. Math. 31 (1976), no. 2, 233–250, MR 54 #4348]. SIAM Journal on Applied Mathematics, 34 (1):198–199, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1395].

Kriegsmann:1978:PAR

[1575] Gregory A. Kriegsmann and Edward W. Larsen. On the parabolic approximation to the reduced wave equation. SIAM Journal on Applied Mathematics, 34(1):200–204, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gal:1978:SSG

[1576] Shmuel Gal. A stochastic search game. SIAM Journal on Applied Mathematics, 34(1):205–210, January 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartley:1978:CEC

[1577] Roger Hartley. On cone-efficiency, cone-convexity and cone-compactness. SIAM Journal on Applied Mathematics, 34(2):211–222, March 1978. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ali-Amidi:1978:NMG

[1578] Ali-Amidi. A note on the moments of the generalized negative binomial distribution and on certain properties of this distribution. SIAM Journal on Applied Mathematics, 34(2):223–224, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cheung:1978:NMC

[1579] To-Yat Cheung and Jiri Soukup. A note on minimum cost networks with nonlinear costs. SIAM Journal on Applied Mathematics, 34(2):225–226, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Phoenix:1978:ATF

[1580] S. Leigh Phoenix. The asymptotic time to failure of a mechanical system of parallel members. SIAM Journal on Applied Mathematics, 34(2):227–246, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yoo:1978:SFT

[1581] Jung Y. Yoo and Daniel D. Joseph.

Stokes flow in a trench between concentric cylinders. SIAM Journal on Applied Mathematics, 34(2):247–285, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Liu:1978:SFC

[1582] C. H. Liu and Daniel D. Joseph.

Stokes flow in conical trenches. SIAM Journal on Applied Mathematics, 34 (2):286–296, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keer:1978:NSC

[1583] L. M. Keer and K. S. Parihar. A note on the singularity at the corner of a wedge-shaped punch or crack. SIAM Journal on Applied Mathematics, 34 (2):297–302, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chapman:1978:RWS

[1584] P. B. Chapman and J. J. Mahony. Reflection of waves in a slowly varying medium. SIAM Journal on Applied Mathematics, 34(2):303–319, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kelly:1978:ESP

[1585] L. M. Kelly and L. T. Watson. Erratum: “Some Perturbation Theorems for Q-Matrices” [SIAM J. Appl. Math. 31 (1976), no. 2, 379–384]. SIAM Journal on Applied Mathematics, 34 (2):320–321, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1407].

Tavantzis:1978:STC

[1586] John Tavantzis, Edward L. Reiss, and Bernard J. Matkowsky. On the smooth transition to convection. SIAM Journal on Applied Mathematics, 34(2):322– 337, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pao:1978:BCL

[1587] C. V. Pao. Bounds for the critical length of a slab with finite reflectors in neutron transport. SIAM Journal on Applied Mathematics, 34 (2):338–347, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Geer:1978:SSW

[1588] James Geer. The scattering of a scalar wave by a slender body of revolution. SIAM Journal on Applied Mathematics, 34(2):348–370, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hwang:1978:NHG

[1589] F. K. Hwang. A note on hypergeometric group testing procedures. SIAM Journal on Applied Mathematics, 34 (2):371–375, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Merrill:1978:CVP

[1590] Samuel Merrill, III. Citizen voting power under the electoral college: A stochastic model based on state voting patterns. SIAM Journal on Applied Mathematics, 34(2):376–390, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Greenberg:1978:AST

[1591] J. M. Greenberg. Axi-symmetric, timeperiodic solutions of reaction-diffusion equations. SIAM Journal on Applied Mathematics, 34(2):391–397, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Albertsen:1978:HDC

[1592] N. C. Albertsen and P. L. Christiansen. Hybrid diffraction coefficients for first and second order discontinuities of two-dimensional scatterers. SIAM Journal on Applied Mathematics, 34(2):398–414, March 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Feintuch:1978:SCL

[1593] Avraham Feintuch. Strictly cyclic linear systems. SIAM Journal on Applied Mathematics, 34(3):415–422, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Besieris:1978:USW

[1594] Ioannis M. Besieris and Werner E. Kohler. Underwater sound wave propagation in the presence of a randomly perturbed sound speed profile. part I. SIAM Journal on Applied Mathematics, 34(3):423–436, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Blankenship:1978:SCS

[1595] G. Blankenship and G. C. Papanicolaou. Stability and control of stochastic systems with wide-band noise disturbances. I. SIAM Journal on Applied Mathematics, 34(3):437–476, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garey:1978:CRB

[1596] M. R. Garey, R. L. Graham, D. S. Johnson, and D. E. Knuth. Complexity results for bandwidth minimization. SIAM Journal on Applied Mathematics, 34(3):477–495, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Leach:1978:GLI

[1597] P. G. L. Leach. On a generalization of the Lewis invariant for the timedependent harmonic oscillator. SIAM Journal on Applied Mathematics, 34 (3):496–503, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ting:1978:PSN

[1598] Lu Ting. Periodic solutions of nonlinear wave equations in n-dimensional space. SIAM Journal on Applied Mathematics, 34(3):504–514, May 1978. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Greenberg:1978:SPD

[1599] J. M. Greenberg and S. P. Hastings. Spatial patterns for discrete models of diffusion in excitable media. SIAM Journal on Applied Mathematics, 34(3):515–523, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Palm:1978:SIM

[1600] G. Palm and T. Poggio. Stochastic identification methods for nonlinear systems: An extension of the Wiener theory. SIAM Journal on Applied Mathematics, 34(3):524–534, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Skalak:1978:NSL

[1601] Francis M. Skalak and Chang-Yi Wang.

On the nonunique solutions of laminar flow through a porous tube or channel. SIAM Journal on Applied Mathematics, 34(3):535–544, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Singh:1978:CPP

[1602] Jagbir Singh. A characterization of positive Poisson distribution and its statistical application. SIAM Journal on Applied Mathematics, 34(3):545– 548, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tier:1978:AAD

[1603] Charles Tier and Joseph B. Keller. Asymptotic analysis of diffusion equations in population genetics. SIAM Journal on Applied Mathematics, 34 (3):549–576, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Drange:1978:BEE

[1604] Hans Birger Drange. On the Boltzmann equation with external forces. SIAM Journal on Applied Mathematics, 34(3):577–592, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Druffel:1978:GSR

[1605] Larry E. Druffel, Douglas C. Schmidt, and Da-Lun Wang. A generator set for representing all automorphisms of a graph. SIAM Journal on Applied Mathematics, 34(3):593–596, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1978:BSP

[1606] A. H. Zemanian. The balanced states of a proportioning network. SIAM Journal on Applied Mathematics, 34 (3):597–610, May 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McEliece:1978:PEA

[1607] Robert J. McEliece and James B. Shearer. A property of Euclid’s algorithm and an application to Padé approximation. SIAM Journal on Applied Mathematics, 34(4):611–615, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Delsarte:1978:AGC

[1608] Ph. Delsarte and Ph. Piret. Automorphism groups of convolutional codes. SIAM Journal on Applied Mathematics, 34(4):616–629, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gladstien:1978:CVV

[1609] Keith Gladstien. The characteristic values and vectors for a class of stochastic matrices arising in genetics. SIAM Journal on Applied Mathematics, 34(4):630–642, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Parthasarathy:1978:ASS

[1610] A. Parthasarathy and R. M. Evan-Iwanowski. On the almost sure stability of linear stochastic systems. SIAM Journal on Applied Mathematics, 34 (4):643–656, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boesch:1978:GLC

[1611] F. T. Boesch and S. Chen. A generalization of line connectivity and optimally invulnerable graphs. SIAM Journal on Applied Mathematics, 34 (4):657–665, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Profant:1978:SHE

[1612] M. Profant, K. Vyska, and U. Eckhardt. The Stewart–Hamilton equations and the indicator dilution method. SIAM Journal on Applied Mathematics, 34(4):666–675, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hochstadt:1978:ISL

[1613] Harry Hochstadt and Burton Lieberman. An inverse Sturm–Liouville problem with mixed given data. SIAM Journal on Applied Mathematics, 34 (4):676–680, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gideon:1978:PTA

[1614] Rudy A. Gideon and John Gurland. A polynomial type approximation for bivariate normal variates. SIAM Journal on Applied Mathematics, 34(4):681– 684, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Herzberger:1978:NBT

[1615] Jürgen Herzberger. Note on a bounding technique for polynomial functions. SIAM Journal on Applied Mathematics, 34(4):685–686, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1305].

Allwright:1978:HFB

[1616] D. J. Allwright. Hypergraphic functions and bifurcations in recurrence relations. SIAM Journal on Applied Mathematics, 34(4):687–691, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Padgett:1978:BER

[1617] W. J. Padgett and C. P. Tsokos. On Bayes estimation of reliability for mixtures of life distributions. SIAM Journal on Applied Mathematics, 34 (4):692–703, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Thakur:1978:EAC

[1618] Lakshman S. Thakur. Error analysis for convex separable programs: The piecewise linear approximation and the bounds on the optimal objective value. SIAM Journal on Applied Mathematics, 34(4):704–714, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Karlof:1978:OGC

[1619] John Karlof and Charles Downey. Odd group codes for the Gaussian channel. SIAM Journal on Applied Mathematics, 34(4):715–716, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chernoff:1978:OCB

[1620] Herman Chernoff and Albert John Petkau. Optimal control of a Brownian motion. SIAM Journal on Applied Mathematics, 34(4):717–731, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Clasen:1978:MSC

[1621] Richard J. Clasen, David Garfinkel, Norman Z. Shapiro, and Gruia-Catalin Roman. A method for solving certain stiff differential equations. SIAM Journal on Applied Mathematics, 34 (4):732–742, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Landman:1978:BME

[1622] K. A. Landman and S. Rosenblat. Bifurcation from a multiple eigenvalue and stability of solutions. SIAM Journal on Applied Mathematics, 34 (4):743–759, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hsu:1978:LBC

[1623] S. B. Hsu. Limiting behavior for competing species. SIAM Journal on Applied Mathematics, 34(4):760–763, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Clements:1978:BOM

[1624] D. L. Clements and C. Rogers. On the Bergman operator method and anti-plane contact problems involving an inhomogeneous half-space. SIAM Journal on Applied Mathematics, 34 (4):764–773, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shepherd:1978:ASR

[1625] J. J. Shepherd. On the asymptotic solution of the Reynolds equation. SIAM Journal on Applied Mathematics, 34 (4):774–791, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Todd:1978:ODS

[1626] Michael J. Todd. Optimal dissection of simplices. SIAM Journal on Applied Mathematics, 34(4):792–803, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

MacGillivray:1978:MEL

[1627] A. D. MacGillivray. On a model equation of Lagerstrom. SIAM Journal on Applied Mathematics, 34(4):804–812, June 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Conway:1978:LTB

[1628] Edward Conway, David Hoff, and Joel Smoller. Large time behavior of solutions of systems of nonlinear reactiondiffusion equations. SIAM Journal on Applied Mathematics, 35(1):1–16, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huff:1978:SCM

[1629] Barthel W. Huff. Some comments on mixtures and infinite divisibility. SIAM Journal on Applied Mathematics, 35(1):17–20, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Duffin:1978:ADO

[1630] R. J. Duffin and T. D. Morley. Almost definite operators and electromechanical systems. SIAM Journal on Applied Mathematics, 35(1):21–30, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kadota:1978:SOC

[1631] T. T. Kadota. Sequential operation of communication channels and the capacity using variable-length codes. SIAM Journal on Applied Mathematics, 35(1):31–47, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Taylor:1978:SRS

[1632] G. C. Taylor. Stability in recurrence systems whose recurrence relations possess a certain positivity property. SIAM Journal on Applied Mathematics, 35(1):48–67, July 1978. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nussinov:1978:ALM

[1633] Ruth Nussinov, George Pieczenik, Jerrold R. Griggs, and Daniel J. Kleitman. Algorithms for loop matchings. SIAM Journal on Applied Mathematics, 35(1):68–82, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berkey:1978:BAS

[1634] Dennis D. Berkey, Marvin I. Freedman, and Barry Granoff. Bifurcation analysis in systems with switch-type coefficients. SIAM Journal on Applied Mathematics, 35(1):83–96, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Horgan:1978:SDE

[1635] C. O. Horgan and L. T. Wheeler. Spatial decay estimates for the Navier– Stokes equations with application to the problem of entry flow. SIAM Journal on Applied Mathematics, 35(1):97– 116, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arnold:1978:SEV

[1636] Barry C. Arnold. Some elementary variations of the Lyapunov inequality. SIAM Journal on Applied Mathematics, 35(1):117–118, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hartfiel:1978:EME

[1637] D. J. Hartfiel. The effect of D on the maximum eigenvalue of DA where A is an ML-matrix. SIAM Journal on Applied Mathematics, 35(1):119–122, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Worz-Busekros:1978:GSE

[1638] Angelika Worz-Busekros. Global stability in ecological systems with continuous time delay. SIAM Journal on Applied Mathematics, 35(1):123–134, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wilson:1978:EUS

[1639] D. G. Wilson. Existence and uniqueness for similarity solutions of one dimensional multi-phase Stefan problems. SIAM Journal on Applied Mathematics, 35(1):135–147, July 1978. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Callegari:1978:MCV

[1640] A. J. Callegari and L. Ting. Motion of a curved vortex filament with decaying vortical core and axial velocity. SIAM Journal on Applied Mathematics, 35(1):148–175, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shepherd:1978:ASN

[1641] J. J. Shepherd. Asymptotic solution of a nonlinear singular perturbation problem. SIAM Journal on Applied Mathematics, 35(1):176–186, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1978:PSA

[1642] D. S. Cohen, A. Fokas, and P. A.

Lagerstrom. Proof of some asymptotic results for a model equation for low Reynolds number flow. SIAM Journal on Applied Mathematics, 35(1):187– 207, July 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cook:1978:LLT

[1643] L. Pamela Cook and Julian D. Cole. Lifting line theory for transonic flow. SIAM Journal on Applied Mathematics, 35(2):209–228, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barnett:1978:TTC

[1644] Arnold Barnett and Daniel J. Kleitman. On two-terminal control of a shuttle service. SIAM Journal on Applied Mathematics, 35(2):229–234, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chow:1978:PMR

[1645] Pao-Liu Chow. Probabilistic model for the reflection of waves from a rough surface. SIAM Journal on Applied Mathematics, 35(2):235–248, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lugannani:1978:SFR

[1646] Robert Lugannani. Sample function regularity of shot processes. SIAM Journal on Applied Mathematics, 35 (2):249–259, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Singer:1978:SOB

[1647] David Singer. Stable orbits and bifurcation of maps of the interval. SIAM Journal on Applied Mathematics, 35 (2):260–267, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grunbaum:1978:IEP

[1648] F. Alberto Grünbaum. An inverse eigenvalue problem for random matrices. SIAM Journal on Applied Mathematics, 35(2):268–273, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Capasso:1978:GSD

[1649] Vincenzo Capasso. Global solution for a diffusive nonlinear deterministic epidemic model. SIAM Journal on Applied Mathematics, 35(2): 274–284, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [1721].

Young:1978:CEC

[1650] H. P. Young and A. Levenglick. A consistent extension of Condorcet’s election principle. SIAM Journal on Applied Mathematics, 35(2):285–300, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gautesen:1978:ESW

[1651] A. K. Gautesen and J. D. Achenbach. On the existence of surface waves guided by a slit. SIAM Journal on Applied Mathematics, 35(2): 301–306, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1978:IRC

[1652] Donald S. Cohen and Bernard J. Matkowsky. On inhibiting runaway in catalytic reactors. SIAM Journal on Applied Mathematics, 35(2): 307–314, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Owen:1978:CBC

[1653] Guillermo Owen. Characterization of the Banzhaf–Coleman index. SIAM Journal on Applied Mathematics, 35 (2):315–327, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Truemper:1978:ACU

[1654] Klaus Truemper. Algebraic characterizations of unimodular matrices. SIAM Journal on Applied Mathematics, 35 (2):328–332, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Willett:1978:FPF

[1655] Michael Willett. Factoring polynomials over a finite field. SIAM Journal on Applied Mathematics, 35(2): 333–337, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rulf:1978:BLM

[1656] Benjamin Rulf. Bounds for the lowest mode in general open waveguides. SIAM Journal on Applied Mathematics, 35(2):338–342, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gustafsson:1978:IPP

[1657] Bertil Gustafsson and Arne Sundstrom. Incompletely parabolic problems in fluid dynamics. SIAM Journal on Applied Mathematics, 35(2): 343–357, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mohanty:1978:GIF

[1658] N. C. Mohanty. On generation of an isotropic field. SIAM Journal on Applied Mathematics, 35(2):358–361, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peters:1978:IPO

[1659] B. Charles Peters, Jr. and Homer F.

Walker. An iterative procedure for obtaining maximum-likelihood estimates of the parameters for a mixture of normal distributions. SIAM Journal on Applied Mathematics, 35(2): 362–378, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bazer:1978:MSO

[1660] J. Bazer and F. Karal. Multiple scattering in one dimension. III. SIAM Journal on Applied Mathematics, 35 (2):379–400, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Guibas:1978:MPS

[1661] L. J. Guibas and A. M. Odlyzko. Maximal prefix-synchronized codes. SIAM Journal on Applied Mathematics, 35 (2):401–418, September 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Oliger:1978:TPA

[1662] Joseph Oliger and Arne Sundstrom. Theoretical and practical aspects of some initial boundary value problems in fluid dynamics. SIAM Journal on Applied Mathematics, 35(3):419–446, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peters:1978:NEM

[1663] B. Charles Peters, Jr. and Homer F. Walker. The numerical evaluation of the maximum-likelihood estimate of a subset of mixture proportions. SIAM Journal on Applied Mathematics, 35 (3):447–452, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Krishnamurthy:1978:EIR

[1664] E. V. Krishnamurthy. Exact inversion of a rational polynomial matrix using finite field transforms. SIAM Journal on Applied Mathematics, 35 (3):453–464, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1978:PPR

[1665] B. J. Matkowsky and G. I. Sivashinsky. Propagation of a pulsating reaction front in solid fuel combustion. SIAM Journal on Applied Mathematics, 35 (3):465–478, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Milstein:1978:EER

[1666] Jaime Milstein. Error estimates for rate constants of inverse problems. SIAM Journal on Applied Mathematics, 35 (3):479–487, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fishburn:1978:SAN

[1667] Peter C. Fishburn. A strategic analysis of nonranked voting systems. SIAM Journal on Applied Mathematics, 35 (3):488–495, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1978:SOS

[1668] A. H. Zemanian. The steady and oscillating states of a proportioning network. SIAM Journal on Applied Mathematics, 35(3):496–507, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bauer:1978:CWM

[1669] Louis Bauer and Edward L. Reiss. Cutoff wavenumbers and modes of hexagonal waveguides. SIAM Journal on Applied Mathematics, 35(3):508–514, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baksalary:1978:RBS

[1670] J. K. Baksalary and R. Kala. Relationships between some representations of the best linear unbiased estimator in the general Gauss-Markoff model. SIAM Journal on Applied Mathematics, 35(3):515–520, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tier:1978:TAD

[1671] Charles Tier and Joseph B. Keller. A tri-allelic diffusion model with selection. SIAM Journal on Applied Mathematics, 35(3):521–535, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1978:RSW

[1672] Donald S. Cohen, John C. Neu, and Rodolfo R. Rosales. Rotating spiral wave solutions of reaction-diffusion equations. SIAM Journal on Applied Mathematics, 35(3):536–547, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Olmstead:1978:OB

[1673] W. E. Olmstead and W. E. Schmitendorf. Optimal blowing. SIAM Journal on Applied Mathematics, 35 (3):548–563, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stokes:1978:CTI

[1674] A. Stokes. Comparison theorems and instability with applications to numerical integration, satellite orbits and averaging. SIAM Journal on Applied Mathematics, 35(3):564–574, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zauderer:1978:ADP

[1675] Erich Zauderer. Asymptotic decomposition of partial differential equations. SIAM Journal on Applied Mathematics, 35(3):575–592, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

DiBlasio:1978:IBV

[1676] Gabriella Di Blasio and Lamberto Lamberti. An initial-boundary value problem for age-dependent population diffusion. SIAM Journal on Applied Mathematics, 35(3):593–615, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Troy:1978:EOT

[1677] William C. Troy. Errata: “Oscillations in a Third-Order Differential Equation Modeling a Nuclear Reactor” [SIAM J. Appl. Math. 32 (1977), no. 1, 146– 153]. SIAM Journal on Applied Mathematics, 35(3):616, November 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1449].

Hsu:1978:CP

[1678] S. B. Hsu, S. P. Hubbell, and Paul Waltman. Competing predators. SIAM Journal on Applied Mathematics, 35 (4):617–625, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nott:1978:DIG

[1679] David Nott. Definite integrals of generalized functions. SIAM Journal on Applied Mathematics, 35(4):626–636, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Singh:1978:NED

[1680] R. S. Singh. Nonparametric estimation of derivatives of average of μdensities with convergence rates and applications. SIAM Journal on Applied Mathematics, 35(4):637–649, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Newell:1978:LWS

[1681] Alan C. Newell. Long waves — short waves; A solvable model. SIAM Journal on Applied Mathematics, 35(4): 650–664, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chow:1978:SIP

[1682] P. L. Chow and L. Maestrello. Stochastic inverse problem in the radiation of noise. SIAM Journal on Applied Mathematics, 35(4):665–677, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bell:1978:SNC

[1683] Jonathan Bell and L. Pamela Cook. On the solutions of a nerve conduction equation. SIAM Journal on Applied Mathematics, 35(4):678–688, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Metropolis:1978:CSF

[1684] N. Metropolis and Gian-Carlo Rota. Combinatorial structure of the faces of the n-cube. SIAM Journal on Applied Mathematics, 35(4):689–694, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Heuch:1978:GAC

[1685] Ivar Heuch. Genetic algebras considered as elements in a vector space. SIAM Journal on Applied Mathematics, 35(4):695–703, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Busenberg:1978:PSP

[1686] Stavros Busenberg and Kenneth L.

Cooke. Periodic solutions of a periodic nonlinear delay differential equation. SIAM Journal on Applied Mathematics, 35(4):704–721, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Balasko:1978:CSS

[1687] Yves Balasko. Connectedness of the set of stable equilibria. SIAM Journal on Applied Mathematics, 35(4): 722–728, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Seymour:1978:SOH

[1688] Brian R. Seymour and Michael P.

Mortell. Second order hyperbolic equations with small nonlinearities. SIAM Journal on Applied Mathematics, 35 (4):729–737, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lewin:1978:PSR

[1689] L. Lewin and J. Kevorkian. On the problem of sustained resonance. SIAM Journal on Applied Mathematics, 35 (4):738–754, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Clarke:1978:AIR

[1690] Bruce L. Clarke. Asymptotes and intercepts of real-power polynomial surfaces from the geometry of the exponent polytope. SIAM Journal on Applied Mathematics, 35(4):755–786, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lehnigk:1978:CSO

[1691] Siegfried H. Lehnigk. Conservative solutions of one-dimensional nonlinear distribution processes with drift. SIAM Journal on Applied Mathematics, 35 (4):787–802, December 1978. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Weeks:1979:CSD

[1692] Jeffrey R. Weeks and Kenneth P. Bogart. Consensus signed digraphs. SIAM Journal on Applied Mathematics, 36 (1):1–14, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yohe:1979:SCR

[1693] J. M. Yohe. A sufficient condition for realizability of constrained graphs. SIAM Journal on Applied Mathematics, 36(1):15–25, February 1979. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Larsen:1979:TTH

[1694] Edward W. Larsen. Two types of homogenization. SIAM Journal on Applied Mathematics, 36(1):26–33, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Delsarte:1979:SPP

[1695] Ph. Delsarte, Y. Genin, and Y. Kamp.

Schur parametrization of positive definite block-toeplitz systems. SIAM Journal on Applied Mathematics, 36 (1):34–46, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Delsarte:1979:NPP

[1696] Ph. Delsarte, Y. Genin, and Y. Kamp.

The Nevanlinna–Pick problem for matrix-valued functions. SIAM Journal on Applied Mathematics, 36(1):47– 61, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

vanLint:1979:PRA

[1697] J. H. van Lint, F. J. MacWilliams, and N. J. A. Sloane. On pseudorandom arrays. SIAM Journal on Applied Mathematics, 36(1):62–72, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://www.jstor.org/
stable/2100768.

Lo:1979:EFD

[1698] James Ting Ho Lo and Linda R. Eshleman. Exponential Fourier densities on SO(3) and optimal estimation and detection for rotational processes. SIAM Journal on Applied Mathematics, 36 (1):73–82, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Raghavan:1979:SRM

[1699] T. E. S. Raghavan. Some remarks on matrix games and nonnegative matrices. SIAM Journal on Applied Mathematics, 36(1):83–85, February 1979.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Geman:1979:SAS

[1700] Stuart Geman. Some averaging and stability results for random differential equations. SIAM Journal on Applied Mathematics, 36(1):86–105, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

MacGillivray:1979:EOD

[1701] A. D. MacGillivray. The existence of an overlap domain for a singular perturbation problem. SIAM Journal on Applied Mathematics, 36(1): 106–114, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dubi:1979:ICM

[1702] A. Dubi and Y. S. Horowitz. The interpretation of conditional Monte Carlo as a form of importance sampling. SIAM Journal on Applied Mathematics, 36 (1):115–122, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Urenko:1979:ICN

[1703] John B. Urenko. Improbability of collisions in Newtonian gravitational systems of specified angular momentum. SIAM Journal on Applied Mathematics, 36(1):123–136, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dellwo:1979:UAS

[1704] David R. Dellwo and Morton B. Friedman. Uniform asymptotic solutions for a differential equation with an almost periodic coefficient. SIAM Journal on Applied Mathematics, 36(1): 137–147, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fleming:1979:EDC

[1705] Wendell H. Fleming. Equilibrium distributions of continuous polygenic traits. SIAM Journal on Applied Mathematics, 36(1):148–168, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1979:CDO

[1706] Joel E. Cohen. The cumulative distance from an observed to a stable age structure. SIAM Journal on Applied Mathematics, 36(1):169–175, February 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lipton:1979:STP

[1707] Richard J. Lipton and Robert Endre Tarjan. A separator theorem for planar graphs. SIAM Journal on Applied Mathematics, 36(2):177–189, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cebeci:1979:RSB

[1708] Tuncer Cebeci, Keith Stewartson, and Peter G. Williams. On the response of a stagnation boundary layer to a change in the external velocity. SIAM Journal on Applied Mathematics, 36 (2):190–199, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pitkaranta:1979:ABS

[1709] Juhani Pitkäranta. Asymptotic behavior of the solution of the integral transport equation in the vicinity of a curved material interface. SIAM Journal on Applied Mathematics, 36 (2):200–218, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Selgrade:1979:MAC

[1710] James F. Selgrade. Mathematical analysis of a cellular control process with positive feedback. SIAM Journal on Applied Mathematics, 36(2):219–229, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wilson:1979:NTM

[1711] H. L. Wilson, W. H. Scott, Jr., and G. C. Pomraning. Neutron transport in moving media. SIAM Journal on Applied Mathematics, 36(2):230–262, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kriegsmann:1979:TLR

[1712] Gregory A. Kriegsmann and Michael Williams. On the tracing of light rays through deformed glass rods with radially graded refractive indices. SIAM Journal on Applied Mathematics, 36 (2):263–272, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Charalambides:1979:SPA

[1713] Ch. A. Charalambides. Some properties and applications of the differences of the generalized factorials. SIAM Journal on Applied Mathematics, 36 (2):273–280, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Leung:1979:CGS

[1714] Anthony Leung. Conditions for global stability concerning a prey-predator model with delay effects. SIAM Journal on Applied Mathematics, 36(2):281– 286, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eisenberg:1979:TQA

[1715] Martin Eisenberg. Two queues with alternating service. SIAM Journal on Applied Mathematics, 36(2):287–303, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lee:1979:DRC

[1716] Ru Ying Lee, Burt S. Holland, and John A. Flueck. Distribution of a ratio of correlated gamma random variables. SIAM Journal on Applied Mathematics, 36(2):304–320, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Getz:1979:ISS

[1717] Wayne M. Getz. Invariant structures in a system of competitive logistic ordinary differential equations. SIAM Journal on Applied Mathematics, 36 (2):321–333, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carpenter:1979:BPE

[1718] Gail A. Carpenter. Bursting phenomena in excitable membranes. SIAM Journal on Applied Mathematics, 36 (2):334–372, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kapila:1979:RDS

[1719] A. K. Kapila and B. J. Matkowsky. Reactive-diffuse systems with Arrhenius kinetics: Multiple solutions, ignition and extinction. SIAM Journal on Applied Mathematics, 36(2):373–389, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kohler:1979:RRT

[1720] Werner E. Kohler. The rapid reaction of two chemical species in the presence of weak random advection and molecular diffusion. SIAM Journal on Applied Mathematics, 36(2):390–405, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Capasso:1979:EGS

[1721] Vincenzo Capasso. Erratum: “Global Solution for a Diffusive Nonlinear Deterministic Epidemic Model” [SIAM J. Appl. Math. 35 (1978), no. 2, 274–284]. SIAM Journal on Applied Mathematics, 36(2):406, April 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1649].

Bertin:1979:LDP

[1722] Giuseppe Bertin and James W.-K. Mark. Local density-potential relation for spiral density waves in galaxies. SIAM Journal on Applied Mathematics, 36(3):407–420, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nakajima:1979:PTD

[1723] F. Nakajima. Periodic time dependent gross-substitute systems. SIAM Journal on Applied Mathematics, 36 (3):421–427, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ablowitz:1979:SWC

[1724] M. J. Ablowitz, M. D. Kruskal, and J. F. Ladik. Solitary wave collisions. SIAM Journal on Applied Mathematics, 36(3):428–437, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Taylor:1979:PZP

[1725] James G. Taylor. Prediction of zero points of solutions to Lanchestertype differential combat equations for modern warfare. SIAM Journal on Applied Mathematics, 36(3):438–456, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Robbins:1979:PSB

[1726] K. A. Robbins. Periodic solutions and bifurcation structure at high R in the Lorenz model. SIAM Journal on Applied Mathematics, 36(3):457–472, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bolker:1979:BRFa

[1727] Ethan D. Bolker and Henry Crapo.

Bracing rectangular frameworks. I. SIAM Journal on Applied Mathematics, 36(3):473–490, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bolker:1979:BRFb

[1728] Ethan D. Bolker. Bracing rectangular frameworks. II. SIAM Journal on Applied Mathematics, 36(3):491–508, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neu:1979:CWD

[1729] John C. Neu. Chemical waves and the diffussive [sic] coupling of limit cycle oscillators. SIAM Journal on Applied Mathematics, 36(3):509–515, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Coste:1979:ABD

[1730] J. Coste, J. Peyraud, and P. Coullet. Asymptotic behaviors in the dynamics of competing species. SIAM Journal on Applied Mathematics, 36 (3):516–543, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mangel:1979:SFS

[1731] Marc Mangel. Small fluctuations in systems with multiple steady states. SIAM Journal on Applied Mathematics, 36(3):544–572, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brown:1979:FAB

[1732] S. N. Brown and K. Stewartson. On finite amplitude Bénard convection in a cylindrical container. part II. SIAM Journal on Applied Mathematics, 36 (3):573–586, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Linan:1979:IRS

[1733] A. Liñán and F. A. Williams. Ignition of a reactive solid exposed to a step in surface temperature. SIAM Journal on Applied Mathematics, 36 (3):587–603, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schuss:1979:EPN

[1734] Zeev Schuss and Bernard J. Matkowsky.

The exit problem: A new approach to diffusion across potential barriers. SIAM Journal on Applied Mathematics, 36(3):604–623, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kassoy:1979:RCG

[1735] D. R. Kassoy. The response of a confined gas to a thermal disturbance. I. Slow transients. SIAM Journal on Applied Mathematics, 36(3):624–634, June 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosenblat:1979:BI

[1736] S. Rosenblat and Stephen H. Davis. Bifurcation from infinity. SIAM Journal on Applied Mathematics, 37(1):1–19, August 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baksalary:1979:CAW

[1737] J. K. Baksalary and R. Kala. Covariance adjustment when a vector of parameters is restricted to a given subspace. SIAM Journal on Applied Mathematics, 37(1):20–21, August 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Langford:1979:PSS

[1738] W. F. Langford. Periodic and steadystate mode interactions lead to tori. SIAM Journal on Applied Mathematics, 37(1):22–48, August 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schuster:1979:LCB

[1739] P. Schuster, K. Sigmund, and R. Wolff. On ω-limits for competition between three species. SIAM Journal on Applied Mathematics, 37(1):49–54, August 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berkey:1979:BEC

[1740] Dennis D. Berkey and Marvin I. Freedman. Buckled elastica at contact: A perturbation analysis. SIAM Journal on Applied Mathematics, 37(1):55–68, August 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Haberman:1979:SVJ

[1741] Richard Haberman. Slowly varying jump and transition phenomena associated with algebraic bifurcation problems. SIAM Journal on Applied Mathematics, 37(1):69–106, August 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matalon:1979:ADI

[1742] Moshe Matalon and G. S. S. Ludford. The asymptotic derivation of Isola and S responses for chambered diffusion flames. SIAM Journal on Applied Mathematics, 37(1):107–123, August 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hwang:1979:GHT

[1743] F. K. Hwang. Generalized Huffman trees. SIAM Journal on Applied Mathematics, 37(1):124–127, August 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Flynn:1979:SSP

[1744] James Flynn. Steady state policies for deterministic dynamic programs. SIAM Journal on Applied Mathematics, 37(1):128–147, August 1979. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Janowitz:1979:MEC

[1745] M. F. Janowitz. Monotone equivariant cluster methods. SIAM Journal on Applied Mathematics, 37(1):148–165, August 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ludwig:1979:OHRa

[1746] Donald Ludwig. Optimal harvesting of a randomly fluctuating resource. I. Application of perturbation methods. SIAM Journal on Applied Mathematics, 37(1):166–184, August 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ludwig:1979:OHRb

[1747] Donald Ludwig and James M. Varah.

Optimal harvesting of a randomly fluctuating resource. II. Numerical methods and results. SIAM Journal on Applied Mathematics, 37(1):185–205, August 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

White:1979:SMN

[1748] Benjamin S. White and Samuel Ellias. A stochastic model for neuronal spike generation. SIAM Journal on Applied Mathematics, 37(1):206–233, August 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Silverstein:1979:REG

[1749] Jack W. Silverstein. On the randomness of eigenvectors generated from networks with random topologies. SIAM Journal on Applied Mathematics, 37 (2):235–245, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hu:1979:BTO

[1750] T. C. Hu, Daniel J. Kleitman, and Jeanne K. Tamaki. Binary trees optimum under various criteria. SIAM Journal on Applied Mathematics, 37 (2):246–256, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morley:1979:GMT

[1751] T. D. Morley. A Gauss–Markov theorem for infinite dimensional regression models with possibly singular covariance. SIAM Journal on Applied Mathematics, 37(2):257–260, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carr:1979:BP

[1752] J. Carr and M. Z. M. Malhardeen. Beck’s problem. SIAM Journal on Applied Mathematics, 37(2):261–262, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Conolly:1979:WTP

[1753] B. W. Conolly and Q. H. Choo.

The waiting time process for a generalized correlated queue with exponential demand and service. SIAM Journal on Applied Mathematics, 37 (2):263–275, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zak:1979:DLF

[1754] Michail Zak. Dynamics of liquid films and thin jets. SIAM Journal on Applied Mathematics, 37(2):276–289, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sapir:1979:DBN

[1755] Marilyn Hirschhorn Sapir and Edward L. Reiss. Dynamic buckling of a nonlinear Timoshenko beam. SIAM Journal on Applied Mathematics, 37 (2):290–301, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [1788].

Rubenfeld:1979:MBP

[1756] L. A. Rubenfeld. A model bifurcation problem exhibiting the effects of slow passage through critical. SIAM Journal on Applied Mathematics, 37 (2):302–306, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neu:1979:CCO

[1757] John C. Neu. Coupled chemical oscillators. SIAM Journal on Applied Mathematics, 37(2):307–315, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cope:1979:RDS

[1758] Davis Cope. Reaction-diffusion systems with explicit traveling wave and transient solutions. SIAM Journal on Applied Mathematics, 37(2):316–324, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gualtierotti:1979:CCO

[1759] A. F. Gualtierotti. On cross-covariance operators. SIAM Journal on Applied Mathematics, 37(2):325–329, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keener:1979:SBM

[1760] James P. Keener. Secondary bifurcation and multiple eigenvalues. SIAM Journal on Applied Mathematics, 37 (2):330–349, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Collinge:1979:TTR

[1761] I. R. Collinge and J. R. Ockendon. Transition through resonance of a Duffing oscillator. SIAM Journal on Applied Mathematics, 37(2):350–357, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Solomon:1979:CBB

[1762] G. Solomon and H. C. A. van Tilborg. A connection between block and convolutional codes. SIAM Journal on Applied Mathematics, 37(2):358–369, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zarrop:1979:EDT

[1763] M. B. Zarrop, R. L. Payne, and G. C. Goodwin. Experiment design for time series analysis: The multivariate case. SIAM Journal on Applied Mathematics, 37(2):370–381, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kurss:1979:RRR

[1764] H. Kurss and W. Vojir. A resonant response to random excitation. SIAM Journal on Applied Mathematics, 37 (2):382–395, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hwang:1979:SRT

[1765] F. K. Hwang, S. Lin, and C. L. Mallows. Some realizability theorems in group testing. SIAM Journal on Applied Mathematics, 37(2):396–400, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Still:1979:CNM

[1766] Jonathan W. Still. A class of new methods for Congressional apportionment. SIAM Journal on Applied Mathematics, 37(2):401–418, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McLoughlin:1979:CRO

[1767] Aileen M. McLoughlin. The covering radius of the (m − 3) rd order Reed Muller codes and a lower bound on the (m − 4) th order Reed Muller codes. SIAM Journal on Applied Mathematics, 37(2):419–422, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rorres:1979:NMP

[1768] Chris Rorres. A nonlinear model of population growth in which fertility is dependent on birth rate. SIAM Journal on Applied Mathematics, 37 (2):423–432, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abrahams:1979:NCB

[1769] Julia Abrahams and John B. Thomas. A note on the characterization of bivariate densities with Gaussian marginal densities. SIAM Journal on Applied Mathematics, 37(2):433–435, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [1789].

Kopell:1979:GAB

[1770] Nancy Kopell. A geometric approach to boundary layer problems exhibiting resonance. SIAM Journal on Applied Mathematics, 37(2):436–458, October 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ufimtsev:1979:UAT

[1771] Petr Ya. Ufimtsev. Uniform asymptotic theory of diffraction by a finite cylinder. SIAM Journal on Applied Mathematics, 37(3):459–466, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kailath:1979:FRO

[1772] T. Kailath, B. Lévy, L. Ljung, and M. Morf. The factorization and representation of operators in the algebra generated by Toeplitz operators. SIAM Journal on Applied Mathematics, 37 (3):467–484, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bhattacharya:1979:STS

[1773] R. N. Bhattacharya and V. K. Gupta.

On a statistical theory of solute transport in porous media. SIAM Journal on Applied Mathematics, 37(3): 485–498, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mimura:1979:ABP

[1774] Masayasu Mimura. Asymptotic behaviors of a parabolic system related to a planktonic prey and predator model. SIAM Journal on Applied Mathematics, 37(3):499–512, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kariv:1979:AANa

[1775] O. Kariv and S. L. Hakimi. An algorithmic approach to network location problems. I. The p-centers. SIAM Journal on Applied Mathematics, 37 (3):513–538, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kariv:1979:AANb

[1776] O. Kariv and S. L. Hakimi. An algorithmic approach to network location problems. II. The p-medians. SIAM Journal on Applied Mathematics, 37 (3):539–560, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Field:1979:EST

[1777] Richard J. Field and William C. Troy.

The existence of solitary traveling wave solutions of a model of the Belousov– Zhabotinskiĭ reaction. SIAM Journal on Applied Mathematics, 37(3): 561–587, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Antman:1979:MES

[1778] Stuart S. Antman. Multiple equilibrium states of nonlinearly elastic strings. SIAM Journal on Applied Mathematics, 37(3):588–604, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barcilon:1979:MSI

[1779] Victor Barcilon. On the multiplicity of solutions of the inverse problem for a vibrating beam. SIAM Journal on Applied Mathematics, 37 (3):605–613, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berger:1979:PKS

[1780] Marc A. Berger. Positive kernels and stochastic integrals. SIAM Journal on Applied Mathematics, 37(3): 614–618, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Guenther:1979:RPI

[1781] R. B. Guenther and R. Schmidt. Remarks on parameter identification. SIAM Journal on Applied Mathematics, 37(3):619–623, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chossat:1979:BSC

[1782] Pascal Chossat. Bifurcation and stability of convective flows in a rotating or not rotating spherical shell. SIAM Journal on Applied Mathematics, 37 (3):624–647, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

deMottoni:1979:CHE

[1783] P. de Mottoni and F. Rothe. Convergence to homogeneous equilibrium state for generalized Volterra–Lotka systems with diffusion. SIAM Journal on Applied Mathematics, 37(3): 648–663, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Braun:1979:SVA

[1784] Martin Braun. The stability of the Van Allen radiation belt. SIAM Journal on Applied Mathematics, 37(3): 664–668, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1979:AES

[1785] B. J. Matkowsky and G. I. Sivashinsky. Acceleration effects on the stability of flame propagation. SIAM Journal on Applied Mathematics, 37 (3):669–685, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1979:ADT

[1786] B. J. Matkowsky and G. I. Sivashinsky.

An asymptotic derivation of two models in flame theory associated with the constant density approximation. SIAM Journal on Applied Mathematics, 37 (3):686–699, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

JaJa:1979:AKT

[1787] Joseph Ja’Ja’. An addendum to Kronecker’s theory of pencils. SIAM Journal on Applied Mathematics, 37(3): 700–712, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sapir:1979:EDB

[1788] Marilyn Hirschhorn Sapir and Edward L. Reiss. Erratum: “Dynamic Buckling of a Nonlinear Timoshenko Beam” [SIAM J. Appl. Math. 37 (1979), no. 2, 290–301]. SIAM Journal on Applied Mathematics, 37(3):713, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1755].

Abrahams:1979:ENC

[1789] Julia Abrahams and John B. Thomas. Erratum: “A Note on the Characterization of Bivariate Densities with Gaussian Marginal Densities” [SIAM J. Appl. Math. 37 (1979), no. 2, 433– 435]. SIAM Journal on Applied Mathematics, 37(3):714, December 1979. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1769].

Ermentrout:1980:LSS

[1790] G. B. Ermentrout and J. D. Cowan. Large scale spatially organized activity in neural nets. SIAM Journal on Applied Mathematics, 38(1):1–21, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brown:1980:DUS

[1791] Peter N. Brown. Decay to uniform states in ecological interactions. SIAM Journal on Applied Mathematics, 38 (1):22–37, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1980:SSR

[1792] Bernard J. Matkowsky, Leonard J. Putnick, and Edward L. Reiss. Secondary states of rectangular plates. SIAM Journal on Applied Mathematics, 38 (1):38–51, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lentini:1980:KSF

[1793] Marianela Lentini and Herbert B.

Keller. The von Kármán swirling flows. SIAM Journal on Applied Mathematics, 38(1):52–64, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Holmes:1980:ACM

[1794] Philip J. Holmes. Averaging and chaotic motions in forced oscillations. SIAM Journal on Applied Mathematics, 38(1):65–80, February 1980. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata and addenda [1885].

Bernard:1980:MCT

[1795] Peter S. Bernard. A method for computing two-dimensional turbulent flows. SIAM Journal on Applied Mathematics, 38(1):81–92, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Geer:1980:ESS

[1796] James Geer. Electromagnetic scattering by a slender body of revolution: Axially incident plane wave. SIAM Journal on Applied Mathematics, 38 (1):93–102, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kubicek:1980:AEC

[1797] Milan Kubíček. Algorithm for evaluation of complex bifurcation points in ordinary differential equations. SIAM Journal on Applied Mathematics, 38 (1):103–107, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rubinstein:1980:RPT

[1798] Ariel Rubinstein. Ranking the participants in a tournament. SIAM Journal on Applied Mathematics, 38 (1):108–111, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Corwin:1980:MFD

[1799] Thomas L. Corwin. On multiplicative functionals on diffusion processes. SIAM Journal on Applied Mathematics, 38(1):112–119, February 1980. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mangel:1980:SFS

[1800] Marc Mangel. Small fluctuations in systems with multiple limit cycles. SIAM Journal on Applied Mathematics, 38 (1):120–138, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Davis:1980:QMH

[1801] Stephen H. Davis and S. Rosenblat.

A quasiperiodic Mathieu–Hill equation. SIAM Journal on Applied Mathematics, 38(1):139–155, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mahoney:1980:PSE

[1802] J. F. Mahoney and B. D. Sivazlian.

Probability of shortage for Erlang distributed demand in a (σ, S) inventory problem. SIAM Journal on Applied Mathematics, 38(1):156–162, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arnold:1980:DFB

[1803] Barry C. Arnold. Distribution-free bounds on the mean of the maximum of a dependent sample. SIAM Journal on Applied Mathematics, 38 (1):163–167, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Antoun:1980:LER

[1804] John T. Antoun and F. Alberto Grünbaum. The lowest eigenvalue of a random matrix. SIAM Journal on Applied Mathematics, 38(1): 168–174, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yasuda:1980:IOS

[1805] Yutaka Yasuda, Hisataka Ito, and Hisatsugu Ishizu. On the input-output stability of linear time-varying systems. SIAM Journal on Applied Mathematics, 38(1):175–188, February 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neu:1980:MNI

[1806] John C. Neu. The method of nearidentity transformations and its applications. SIAM Journal on Applied Mathematics, 38(2):189–208, April 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maiti:1980:EGC

[1807] M. Maiti. On the equilibrium of a Griffith crack opened by a rigid inclusion. SIAM Journal on Applied Mathematics, 38(2):209–214, April 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Williams:1980:BLD

[1808] J. C. Williams, III and T. B. Rhyne. Boundary layer development on a wedge impulsively set into motion. SIAM Journal on Applied Mathematics, 38(2):215–224, April 1980. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

OMalley:1980:ANM

[1809] R. E. O’Malley, Jr. and J. E. Flaherty. Analytical and numerical methods for nonlinear singular singularlyperturbed initial value problems. SIAM Journal on Applied Mathematics, 38 (2):225–248, April 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reiss:1980:MTT

[1810] Edward L. Reiss. A modified twotime method for the dynamic transitions of bifurcation. SIAM Journal on Applied Mathematics, 38(2):249–260, April 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Artstein:1980:CDP

[1811] Zvi Artstein and George Karakostas.

Convergence in the delay population equation. SIAM Journal on Applied Mathematics, 38(2):261–272, April 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lardner:1980:CPS

[1812] R. W. Lardner and G. N. Bojadziev.

Comments on a paper by Seymour and Mortell. SIAM Journal on Applied Mathematics, 38(2):273–274, April 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nachman:1980:NSB

[1813] A. Nachman and A. Callegari. A nonlinear singular boundary value problem in the theory of pseudoplastic fluids. SIAM Journal on Applied Mathematics, 38(2):275–281, April 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hofbauer:1980:DSU

[1814] J. Hofbauer, P. Schuster, K. Sigmund, and R. Wolff. Dynamical systems under constant organization. II. Homogeneous growth functions of degree p = 2. SIAM Journal on Applied Mathematics, 38(2):282–304, April 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neu:1980:LPC

[1815] John C. Neu. Large populations of coupled chemical oscillators. SIAM Journal on Applied Mathematics, 38 (2):305–316, April 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Johnson:1980:GMA

[1816] Harold H. Johnson and Andrew Vogt. A geometric method for approximating convex arcs. SIAM Journal on Applied Mathematics, 38(2):317–325, April 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Normandia:1980:SED

[1817] M. J. Normandia and G. S. S. Ludford. Surface equilibrium in drop combustion. SIAM Journal on Applied Mathematics, 38(2):326–339, April 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Simpson:1980:SPS

[1818] Henry C. Simpson. Stability of periodic solutions of nonlinear integrodifferential systems. SIAM Journal on Applied Mathematics, 38(3):341–363, June 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yannakakis:1980:EDS

[1819] M. Yannakakis and F. Gavril. Edge dominating sets in graphs. SIAM Journal on Applied Mathematics, 38 (3):364–372, June 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dassios:1980:SOL

[1820] George Dassios. Second order lowfrequency scattering by the soft ellipsoid. SIAM Journal on Applied Mathematics, 38(3):373–381, June 1980.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kapila:1980:RDSa

[1821] A. K. Kapila, B. J. Matkowsky, and J. Vega. Reactive-diffusive system with Arrhenius kinetics: Peculiarities of the spherical geometry. SIAM Journal on Applied Mathematics, 38(3):382–401, June 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Harary:1980:GCS

[1822] Frank Harary and Andrew Vince.

Graphical completions of a sequence. SIAM Journal on Applied Mathematics, 38(3):402, June 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Marzec:1980:ODE

[1823] C. J. Marzec and E. A. Spiegel.

Ordinary differential equations with strange attractors. SIAM Journal on Applied Mathematics, 38(3):403–421, June 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gibbs:1980:TWB

[1824] R. G. Gibbs. Traveling waves in the Belousov–Zhabotinskiĭ reaction. SIAM Journal on Applied Mathematics, 38 (3):422–444, June 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Holmes:1980:LFA

[1825] Mark H. Holmes. Low frequency asymptotics for a hydroelastic model of the cochlea. SIAM Journal on Applied Mathematics, 38(3):445–456, June 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cope:1980:SLC

[1826] Davis Cope. Stability of limit cycle solutions of reaction-diffusion equations. SIAM Journal on Applied Mathematics, 38(3):457–479, June 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Devroye:1980:DAB

[1827] Luc Devroye and Gary L. Wise. Detection of abnormal behavior via nonparametric estimation of the support. SIAM Journal on Applied Mathematics, 38(3):480–488, June 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1980:NTC

[1828] B. J. Matkowsky, L. J. Putnick, and G. I. Sivashinsky. A nonlinear theory of cellular flames. SIAM Journal on Applied Mathematics, 38(3):489–504, June 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wilson:1980:SIE

[1829] J. H. Wilson. Solution of integral equation important in signal processing of radar and sonar hydrophone array outputs. SIAM Journal on Applied Mathematics, 39(1):1–7, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Duffy:1980:SWS

[1830] M. R. Duffy, N. F. Britton, and J. D. Murray. Spiral wave solutions of practical reaction-diffusion systems.

SIAM Journal on Applied Mathematics, 39(1):8–13, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Friedman:1980:MSN

[1831] Mark J. Friedman. Mathematical study of the nonlinear singular integral magnetic field equation. I. SIAM Journal on Applied Mathematics, 39(1):14–20, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kapila:1980:RDSb

[1832] A. K. Kapila. Reactive-diffusive system with Arrhenius kinetics: Dynamics of ignition. SIAM Journal on Applied Mathematics, 39(1):21–36, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Capasso:1980:SRS

[1833] Vincenzo Capasso and Donato Fortunato. Stability results for semilinear evolution equations and their application to some reaction-diffusion problems. SIAM Journal on Applied Mathematics, 39(1):37–47, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Diekmann:1980:SBV

[1834] O. Diekmann, D. Hilhorst, and L. A.

Peletier. A singular boundary value problem arising in a pre-breakdown gas discharge. SIAM Journal on Applied Mathematics, 39(1):48–66, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sivashinsky:1980:FPU

[1835] G. I. Sivashinsky. On flame propagation under conditions of stoichiometry. SIAM Journal on Applied Mathematics, 39(1):67–82, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chang:1980:SSF

[1836] Tao Chang and Gabor T. Herman. A scientific study of filter selection for a fan-beam convolution reconstruction algorithm. SIAM Journal on Applied Mathematics, 39(1):83–105, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bultheel:1980:RAN

[1837] A. Bultheel. Recursive algorithms for nonnormal Padé tables. SIAM Journal on Applied Mathematics, 39 (1):106–118, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Howell:1980:CGF

[1838] J. A. Howell, T. F. Smith, and M. S. Waterman. Computation of generating functions for biological molecules. SIAM Journal on Applied Mathematics, 39(1):119–133, August 1980. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sentis:1980:AHT

[1839] Rémi Sentis. Approximation and homogenization of a transport process. SIAM Journal on Applied Mathematics, 39(1):134–141, August 1980. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sirovich:1980:BEN

[1840] Lawrence Sirovich. Boundary effects in neural networks. SIAM Journal on Applied Mathematics, 39(1):142–160, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McKelvey:1980:CVE

[1841] Richard D. McKelvey, Peter C. Ordeshook, and Peter Ungar. Conditions for voting equilibria in continuous voter distributions. SIAM Journal on Applied Mathematics, 39(1):161–168, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hryniewicz:1980:NGM

[1842] Z. Hryniewicz. Numerical generation methods of homogeneous and nonhomogeneous two-dimensional Gaussian random fields. SIAM Journal on Applied Mathematics, 39(1):169–172, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tydeman:1980:NKM

[1843] John Tydeman and Robert Mitchell. A note on the Kounias and Marin method of best linear Bonferroni bounds. SIAM Journal on Applied Mathematics, 39 (1):173–177, August 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mei:1980:FSL

[1844] Chiang C. Mei and E. O. Tuck. Forward scattering by long thin bodies. SIAM Journal on Applied Mathematics, 39(1):178–191, August 1980.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nowinski:1980:BBS

[1845] J. L. Nowinski. Bilateral bounds for the solution of a generalized biharmonic boundary value problem. SIAM Journal on Applied Mathematics, 39 (2):193–200, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Larsen:1980:AAN

[1846] E. W. Larsen and G. C. Pomraning. Asymptotic analysis of nonlinear Marshak waves. SIAM Journal on Applied Mathematics, 39(2):201–212, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bajaj:1980:HBP

[1847] A. K. Bajaj, P. R. Sethna, and T. S. Lundgren. Hopf bifurcation phenomena in tubes carrying a fluid. SIAM Journal on Applied Mathematics, 39 (2):213–230, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Consul:1980:GNB

[1848] P. C. Consul and H. C. Gupta. The generalized negative binomial distribution and its characterization by zero regression. SIAM Journal on Applied Mathematics, 39(2):231–237, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bluman:1980:TDP

[1849] George W. Bluman. On the transformation of diffusion processes into the Wiener process. SIAM Journal on Applied Mathematics, 39(2):238–247, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hunter:1980:DPS

[1850] C. Hunter and B. Guerrieri. Deducing the properties of singularities of functions from their Taylor series coefficients. SIAM Journal on Applied Mathematics, 39(2):248–263, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [1932].

Lugt:1980:MDS

[1851] Hans J. Lugt. Multipole decomposition of solutions of the vector diffusion equation. SIAM Journal on Applied Mathematics, 39(2):264–271, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cannon:1980:IPN

[1852] J. R. Cannon and Paul DuChateau. An inverse problem for a nonlinear diffusion equation. SIAM Journal on Applied Mathematics, 39(2):272–289, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1980:PPF

[1853] Bernard J. Matkowsky and David O.

Olagunju. Propagation of a pulsating flame front in a gaseous combustible mixture. SIAM Journal on Applied Mathematics, 39(2):290–300, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Greenberg:1980:SWS

[1854] J. M. Greenberg. Spiral waves for λ−ω systems. SIAM Journal on Applied Mathematics, 39(2):301–309, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tuckwell:1980:ESL

[1855] Henry C. Tuckwell. Evidence of solitonlike behavior of solitary waves in a nonlinear reaction-diffusion system. SIAM Journal on Applied Mathematics, 39 (2):310–322, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ermentrout:1980:SBN

[1856] G. B. Ermentrout and J. D. Cowan. Secondary bifurcation in neuronal nets. SIAM Journal on Applied Mathematics, 39(2):323–340, October 1980. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tretiak:1980:ERT

[1857] Oleh Tretiak and Charles Metz. The exponential Radon transform. SIAM Journal on Applied Mathematics, 39 (2):341–354, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fung:1980:SWF

[1858] K. Y. Fung. Shock wave formation at a caustic. SIAM Journal on Applied Mathematics, 39(2):355–371, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Beckett:1980:CNF

[1859] P. M. Beckett. Combined natural- and forced-convection between parallel vertical walls. SIAM Journal on Applied Mathematics, 39(2):372–384, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Farley:1980:BTC

[1860] Arthur M. Farley. Broadcast time in communication networks. SIAM Journal on Applied Mathematics, 39 (2):385–390, October 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kapila:1980:RDSc

[1861] A. K. Kapila and B. J. Matkowsky.

Reactive-diffusive system with Arrhenius kinetics: The Robin problem. SIAM Journal on Applied Mathematics, 39(3):391–401, December 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Natterer:1980:SSA

[1862] Frank Natterer. A Sobolev space analysis of picture reconstruction. SIAM Journal on Applied Mathematics, 39 (3):402–411, December 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kassoy:1980:TEC

[1863] D. R. Kassoy and Justin Poland. The thermal explosion confined by a constant temperature boundary: I — the induction period solution. SIAM Journal on Applied Mathematics, 39(3): 412–430, December 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Singh:1980:ANN

[1864] R. S. Singh. Asymptotic normality of nonparametric estimators of derivatives of average of μ-densities. SIAM Journal on Applied Mathematics, 39 (3):431–439, December 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reiss:1980:NAM

[1865] Edward L. Reiss. A new asymptotic method for jump phenomena. SIAM Journal on Applied Mathematics, 39 (3):440–455, December 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gray:1980:SOP

[1866] Samuel H. Gray. A second order procedure for one-dimensional velocity inversion. SIAM Journal on Applied Mathematics, 39(3):456–462, December 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1980:FTF

[1867] Donald S. Cohen and S. Rosenblat. Flickering and thermal flicker waves on catalytic wires and gauzes and in chemical reactors. SIAM Journal on Applied Mathematics, 39(3):463–474, December 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Skar:1980:SMM

[1868] Sherwin J. Skar. Stability of multimachine power systems with nontrivial transfer conductances. SIAM Journal on Applied Mathematics, 39(3): 475–491, December 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Harper:1980:MIS

[1869] E. Y. Harper. Motion-induced singularities in power spectra associated with ocean gravity-wave fluctuations. SIAM Journal on Applied Mathematics, 39 (3):492–511, December 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Logan:1980:SSR

[1870] J. David Logan and José de Jesús Pérez.

Similarity solutions for reactive shock hydrodynamics. SIAM Journal on Applied Mathematics, 39(3):512–527, December 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keener:1980:WEM

[1871] J. P. Keener. Waves in excitable media. SIAM Journal on Applied Mathematics, 39(3):528–548, December 1980. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hethcote:1981:NOE

[1872] Herbert W. Hethcote, Harlan W.

Stech, and P. van den Driessche. Nonlinear oscillations in epidemic models. SIAM Journal on Applied Mathematics, 40(1):1–9, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grimshaw:1981:MWN

[1873] R. Grimshaw. The modulation of a weakly nonlinear wave packet in an inhomogeneous medium. SIAM Journal on Applied Mathematics, 40 (1):10–34, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lange:1981:BCS

[1874] Charles G. Lange. Branching from closely spaced eigenvalues with application to a model biochemical reaction. SIAM Journal on Applied Mathematics, 40(1):35–51, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lange:1981:BLE

[1875] Charles G. Lange. Branching from large eigenvalues of nonlinear Sturm– Liouville systems. SIAM Journal on Applied Mathematics, 40(1):52–70, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bardsley:1981:ATP

[1876] J. N. Bardsley and A. Dubi. The average transport path length in scattering media. SIAM Journal on Applied Mathematics, 40(1):71–77, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1981:DVL

[1877] S. H. Smith. The development of vertical layers in a rotating fluid. SIAM Journal on Applied Mathematics, 40 (1):78–89, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ikeuchi:1981:ASH

[1878] Masatoshi Ikeuchi, Kazuo Inoue, Hideo Sawami, and Hiroshi Niki. Arbitrarily shaped hollow waveguide analysis by the α-interpolation method. SIAM Journal on Applied Mathematics, 40 (1):90–98, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Homentcovschi:1981:ASO

[1879] Dorel Homentcovschi. Axially symmetric Oseen flow past a slender body of revolution. SIAM Journal on Applied Mathematics, 40(1):99–112, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hanson:1981:ASF

[1880] Floyd B. Hanson and Charles Tier.

An asymptotic solution of the first passage problem for singular diffusion in population biology. SIAM Journal on Applied Mathematics, 40(1): 113–132, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pfeifer:1981:VSS

[1881] Phillip E. Pfeifer and Stuart Jay Deutsch. Variance of the sample spacetime correlation function of contemporaneously correlated variables. SIAM Journal on Applied Mathematics, 40 (1):133–136, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Samson:1981:GSV

[1882] J. C. Samson and J. V. Olson. Generalized Stokes vectors and generalized power spectra for second-order stationary vector-processes. SIAM Journal on Applied Mathematics, 40(1): 137–149, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Plant:1981:FDD

[1883] Richard E. Plant. A Fitzhugh differential-difference equation modeling recurrent neural feedback. SIAM Journal on Applied Mathematics, 40 (1):150–162, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wenocur:1981:RMR

[1884] Roberta S. Wenocur and Stephen W.

Salant. Recurrence of a modified random walk and its application to an economic model. SIAM Journal on Applied Mathematics, 40(1):163–166, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Holmes:1981:EAA

[1885] Philip J. Holmes. Errata and addenda: “Averaging and Chaotic Motions in Forced Oscillations” [SIAM J. Appl. Math. 38 (1980), no. 1, 65–80, MR 80m:34044]. SIAM Journal on Applied Mathematics, 40(1):167–168, February 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1794].

Thrall:1981:SPL

[1886] Anthony D. Thrall. Some properties of least-squares digital filters. SIAM Journal on Applied Mathematics, 40 (2):169–178, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Franklin:1981:CIS

[1887] Joel N. Franklin. Confidence intervals for stereological estimators with infinite variance. SIAM Journal on Applied Mathematics, 40(2):179–190, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Munier:1981:GTN

[1888] A. Munier, J. R. Burgan, J. Gutierrez, E. Fijalkow, and M. R. Feix. Group transformations and the nonlinear heat diffusion equation. SIAM Journal on Applied Mathematics, 40 (2):191–207, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Williams:1981:PSE

[1889] Randall G. Williams. The problem of stochastic exit. SIAM Journal on Applied Mathematics, 40(2):208–223, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bissett:1981:ISP

[1890] Edward J. Bissett. Interaction of steady and periodic bifurcating modes with imperfection effects in reactiondiffusion systems. SIAM Journal on Applied Mathematics, 40(2):224–241, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1981:EFP

[1891] B. J. Matkowsky and Z. Schuss. Eigenvalues of the Fokker–Planck operator and the approach to equilibrium for diffusions in potential fields. SIAM Journal on Applied Mathematics, 40 (2):242–254, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sivashinsky:1981:SNF

[1892] G. I. Sivashinsky and B. J. Matkowsky.

On the stability of nonadiabatic flames. SIAM Journal on Applied Mathematics, 40(2):255–260, April 1981. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Linan:1981:NIH

[1893] A. Linán and F. A. Williams. Note on ignition by a hot catalytic surface. SIAM Journal on Applied Mathematics, 40(2):261–265, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Swick:1981:NMH

[1894] K. E. Swick. A nonlinear model for human population dynamics. SIAM Journal on Applied Mathematics, 40 (2):266–278, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Denman:1981:VTE

[1895] Harry H. Denman and Lawrence H. Buch. Variational treatment for equations of the Langmuir type. SIAM Journal on Applied Mathematics, 40 (2):279–282, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McIlroy:1981:EFDa

[1896] D. K. McIlroy and D. P. Mason. Electric field dissociation of charged macromolecular sites in polarized biomembranes. I. The mathematical model. SIAM Journal on Applied Mathematics, 40(2):283–297, April 1981. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mason:1981:EFD

[1897] D. P. Mason and D. K. McIlroy. Electric field dissociation of charged macromolecular sites in polarized biomembranes. II. A dual perturbation solution. SIAM Journal on Applied Mathematics, 40(2):298–314, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chow:1981:ASR

[1898] P. L. Chow and K. L. Chiou. Asymptotic stability of randomly perturbed linear periodic systems. SIAM Journal on Applied Mathematics, 40(2):315– 326, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mangel:1981:SRM

[1899] Marc Mangel. Search for a randomly moving object. SIAM Journal on Applied Mathematics, 40(2):327–338, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bezdek:1981:DCCa

[1900] James C. Bezdek, Chris Coray, Robert Gunderson, and James Watson. Detection and characterization of cluster substructure. I. Linear structure: fuzzy c-lines. SIAM Journal on Applied Mathematics, 40(2):339–357, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bezdek:1981:DCCb

[1901] James C. Bezdek, Chris Coray, Robert Gunderson, and James Watson. Detection and characterization of cluster substructure. II. Fuzzy c-varieties and convex combinations thereof. SIAM Journal on Applied Mathematics, 40 (2):358–372, April 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Payton:1981:EEA

[1902] Robert G. Payton. Epicenter and epicentral-axis motion of a transversely isotropic elastic half-space. SIAM Journal on Applied Mathematics, 40 (3):373–389, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Guy:1981:UGT

[1903] J. Guy and B. Mangeot. Use of group theory in various integral equations. SIAM Journal on Applied Mathematics, 40(3):390–399, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Phatarfod:1981:IDD

[1904] Ravindra M. Phatarfod. The infinitely deep dam with seasonal Markov inflows. SIAM Journal on Applied Mathematics, 40(3):400–408, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ramaswamy:1981:NEE

[1905] Mythily Ramaswamy. On a nonlinear elliptic equation with periodic coefficients. SIAM Journal on Applied Mathematics, 40(3):409–418, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wolf:1981:TDQ

[1906] Kurt Bernardo Wolf. On timedependent quadratic quantum Hamiltonians. SIAM Journal on Applied Mathematics, 40(3):419–431, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sivashinsky:1981:SPC

[1907] G. Sivashinsky. On spinning propagation of combustion waves. SIAM Journal on Applied Mathematics, 40 (3):432–438, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hagin:1981:SAS

[1908] Frank Hagin. A stable approach to solving one-dimensional inverse problems. SIAM Journal on Applied Mathematics, 40(3):439–453, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gaver:1981:CRL

[1909] D. P. Gaver and P. A. Jacobs. On combinations of random loads. SIAM Journal on Applied Mathematics, 40 (3):454–466, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hadidi:1981:QPC

[1910] Nasser Hadidi. Queues with partial correlation. SIAM Journal on Applied Mathematics, 40(3):467–475, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bebernes:1981:MAB

[1911] J. W. Bebernes and D. R. Kassoy. A mathematical analysis of blowup for thermal reactions — the spatially nonhomogeneous case. SIAM Journal on Applied Mathematics, 40(3):476–484, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Percus:1981:ODR

[1912] O. E. Percus and J. K. Percus.

One-dimensional random walk with phase transition. SIAM Journal on Applied Mathematics, 40(3):485–497, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1981:CCO

[1913] Hal L. Smith. Competitive coexistence in an oscillating chemostat. SIAM Journal on Applied Mathematics, 40(3):498–522, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Case:1981:RME

[1914] J. Case. On the relative merits of the English and American rules for allocating court costs. SIAM Journal on Applied Mathematics, 40(3):523–527, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kushner:1981:WCS

[1915] H. J. Kushner and Hai Huang. On the weak convergence of a sequence of general stochastic difference equations to a diffusion. SIAM Journal on Applied Mathematics, 40(3):528–541, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shaked:1981:EII

[1916] Moshe Shaked. Extensions of the IHR and IHRA aging notions. SIAM Journal on Applied Mathematics, 40 (3):542–550, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1981:PBS

[1917] Bernard J. Matkowsky and David O. Olagunju. Pulsations in a burnerstabilized premixed plane flame. SIAM Journal on Applied Mathematics, 40 (3):551–562, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dean:1981:FAM

[1918] Donald L. Dean and Robert T. Ramsay. Field analysis of models with multiple node sets along one coordinate. SIAM Journal on Applied Mathematics, 40(3):563–568, June 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hadeler:1981:SPS

[1919] K. P. Hadeler. Stable polymorphisms in a selection model with mutation. SIAM Journal on Applied Mathematics, 41(1):1–7, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Colton:1981:DSI

[1920] David Colton and Andreas Kirsch. The determination of the surface impedance of an obstacle from measurements of the far field pattern. SIAM Journal on Applied Mathematics, 41(1):8–15, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Persek:1981:ASS

[1921] Stephen C. Persek. Asymptotic stability on slow time scales for periodic systems. SIAM Journal on Applied Mathematics, 41(1):16–28, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kapila:1981:ASD

[1922] A. K. Kapila. Arrhenius systems: Dynamics of jump due to slow passage through criticality. SIAM Journal on Applied Mathematics, 41(1):29–42, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cambanis:1981:LPO

[1923] Stamatis Cambanis and Grady Miller. Linear problems in p th order and stable processes. SIAM Journal on Applied Mathematics, 41(1):43–69, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Majda:1981:QMD

[1924] Andrew Majda. A qualitative model for dynamic combustion. SIAM Journal on Applied Mathematics, 41(1):70–93, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Seymour:1981:DSM

[1925] B. R. Seymour and M. P. Mortell. Discontinuous solutions of a measurepreserving mapping. SIAM Journal on Applied Mathematics, 41(1):94–106, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jones:1981:CMS

[1926] A. S. Jones. On a conjecture of Mortell and Seymour. SIAM Journal on Applied Mathematics, 41(1):107–111, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barshinger:1981:EPF

[1927] Richard N. Barshinger and James F. Geer. The electrostatic potential field about a thin oblate body of revolution. SIAM Journal on Applied Mathematics, 41(1):112–126, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keener:1981:IPB

[1928] James P. Keener. Infinite period bifurcation and global bifurcation branches. SIAM Journal on Applied Mathematics, 41(1):127–144, August 1981. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klaasen:1981:STW

[1929] Gene A. Klaasen and William C. Troy.

The stability of traveling wave front solutions of a reaction-diffusion system. SIAM Journal on Applied Mathematics, 41(1):145–167, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bell:1981:MPU

[1930] Jonathan Bell. Modeling parallel, unmyelinated axons: Pulse trapping and ephaptic transmission. SIAM Journal on Applied Mathematics, 41 (1):168–180, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Perko:1981:POR

[1931] L. M. Perko. Periodic orbits in the restricted problem: an analysis in the neighborhood of a 1st species–2nd species bifurcation. SIAM Journal on Applied Mathematics, 41(1):181–202, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hunter:1981:EDP

[1932] C. Hunter and B. Guerrieri. Erratum: “Deducing the Properties of Singularities of Functions from their Taylor Series Coefficients” [SIAM J. Appl. Math. 39 (1980), no. 2, 248–263].

SIAM Journal on Applied Mathematics, 41(1):203, August 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1850].

Sivazlian:1981:MEG

[1933] B. D. Sivazlian. On a multivariate extension of the gamma and beta distributions. SIAM Journal on Applied Mathematics, 41(2):205–209, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Richter:1981:IPS

[1934] Gerard R. Richter. An inverse problem for the steady state diffusion equation. SIAM Journal on Applied Mathematics, 41(2):210–221, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Britten:1981:STR

[1935] D. J. Britten and F. W. Lemire. A structure theorem for rings supporting a discrete Fourier transform. SIAM Journal on Applied Mathematics, 41 (2):222–226, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miles:1981:ESP

[1936] John W. Miles. An envelope soliton problem. SIAM Journal on Applied Mathematics, 41(2):227–230, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kassoy:1981:TEC

[1937] D. R. Kassoy and J. Poland. The thermal explosion confined by a constant temperature boundary. II. The extremely rapid transient. SIAM Journal on Applied Mathematics, 41(2):

231–246, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Duck:1981:TOF

[1938] P. W. Duck. Torsional oscillations of a finite disk. SIAM Journal on Applied Mathematics, 41(2):247–264, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kates:1981:UTM

[1939] Ronald E. Kates. Uniformity test for the method of strained coordinates. SIAM Journal on Applied Mathematics, 41(2):265–270, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cook:1981:WDR

[1940] L. Pamela Cook and Mark Holmes.

Waves and dispersion relations for hydroelastic systems. SIAM Journal on Applied Mathematics, 41(2):271–287, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Williams:1981:ALA

[1941] Michael Williams. Another look at Ackerberg–O’Malley resonance. SIAM Journal on Applied Mathematics, 41 (2):288–293, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keller:1981:ODB

[1942] Joseph B. Keller. Oblique derivative boundary conditions and the image method. SIAM Journal on Applied Mathematics, 41(2):294–300, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Levine:1981:CIC

[1943] Lawrence E. Levine and Sylvester L. Tuohy. Correcting for initial conditions in multitime problems. SIAM Journal on Applied Mathematics, 41 (2):301–305, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vanden-Broeck:1981:DLS

[1944] Jean Marc Vanden-Broeck. Deformation of a liquid surface by an impinging gas jet. SIAM Journal on Applied Mathematics, 41(2):306–309, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

DuChateau:1981:MUR

[1945] Paul DuChateau. Monotonicity and uniqueness results in identifying an unknown coefficient in a nonlinear diffusion equation. SIAM Journal on Applied Mathematics, 41(2):310–323, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sverdlove:1981:LMI

[1946] Ronald Sverdlove. Leslie models with input. SIAM Journal on Applied Mathematics, 41(2):324–330, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kreiss:1981:RSP

[1947] Heinz-Otto Kreiss. Resonance for singular perturbation problems. SIAM Journal on Applied Mathematics, 41 (2):331–344, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Harrison:1981:DMR

[1948] J. Michael Harrison and Martin I.

Reiman. On the distribution of multidimensional reflected Brownian motion. SIAM Journal on Applied Mathematics, 41(2):345–361, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Skinner:1981:NLS

[1949] L. A. Skinner. Note on the Lagerstrom singular perturbation models. SIAM Journal on Applied Mathematics, 41 (2):362–364, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neu:1981:SPR

[1950] John C. Neu. Stochastically perturbed resonance. SIAM Journal on Applied Mathematics, 41(2):365–369, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bouc:1981:MSR

[1951] R. Bouc and E. Pardoux. Moments of semilinear random evolutions. SIAM Journal on Applied Mathematics, 41 (2):370–399, October 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knowles:1981:NER

[1952] James K. Knowles. A note on the energy release rate in quasi-static elastic crack propagation. SIAM Journal on Applied Mathematics, 41(3): 401–412, December 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ahmad:1981:APS

[1953] Ibrahim A. Ahmad. Asymptotic properties of system reliability estimators. SIAM Journal on Applied Mathematics, 41(3):413–421, December 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hsu:1981:ECM

[1954] Sze Bi Hsu, Kuo Shung Cheng, and S. P. Hubbell. Exploitative competition of microorganisms for two complementary nutrients in continuous cultures. SIAM Journal on Applied Mathematics, 41(3):422–444, December 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rahman:1981:HLC

[1955] Matiur Rahman and I. C. Chakravartty. Hydrodynamic loading calculations for offshore structures. SIAM Journal on Applied Mathematics, 41(3):445–458, December 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Glaz:1981:SBC

[1956] Harland M. Glaz. Statistical behavior and coherent structures in twodimensional inviscid turbulence. SIAM Journal on Applied Mathematics, 41 (3):459–479, December 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lardner:1981:SSE

[1957] R. W. Lardner and G. Nicklason. The stable self-excitations of the nonlinear wave equation of Van Der Pol type. SIAM Journal on Applied Mathematics, 41(3):480–492, December 1981.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pitts:1981:ORP

[1958] E. Pitts. The overlap of random particles and similar problems: Expressions for variance of coverage and its analogue. SIAM Journal on Applied Mathematics, 41(3):493–498, December 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fishburn:1981:ASV

[1959] Peter C. Fishburn. An analysis of simple voting systems for electing committees. SIAM Journal on Applied Mathematics, 41(3):499–502, December 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keener:1981:IFM

[1960] J. P. Keener, F. C. Hoppensteadt, and J. Rinzel. Integrate-and-fire models of nerve membrane response to oscillatory input. SIAM Journal on Applied Mathematics, 41(3):503–517, December 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Blankenship:1981:AES

[1961] G. L. Blankenship and J. S. Baras. Accurate evaluation of stochastic Wiener integrals with applications to scattering in random media and to nonlinear filtering. SIAM Journal on Applied Mathematics, 41(3):518–552, December 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lee:1981:NCS

[1962] Alan J. Lee. A note on the Campbell sampling theorem. SIAM Journal on Applied Mathematics, 41(3): 553–557, December 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maine:1981:FIP

[1963] Richard E. Maine and Kenneth W. Iliff. Formulation and implementation of a practical algorithm for parameter estimation with process and measurement noise. SIAM Journal on Applied Mathematics, 41(3):558–579, December 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McIlroy:1981:EFDb

[1964] D. K. McIlroy and D. P. Mason. Electric field dissociation of charged macromolecular sites in polarized biomembranes. III. Numerical results and application to nervous conduction. SIAM Journal on Applied Mathematics, 41 (3):580–591, December 1981. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shen:1982:UAE

[1965] M. C. Shen. Uniform asymptotic expansions related to unsteady flow in unconfined aquifers. SIAM Journal on Applied Mathematics, 42(1):1– 9, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schindler:1982:IGP

[1966] G. M. Schindler. The intrinsic geometrical and physical properties of plane steady gas flow. SIAM Journal on Applied Mathematics, 42(1):10–26, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1982:ISS

[1967] Hal L. Smith. The interaction of steady state and Hopf bifurcations in a twopredator–one-prey competition model. SIAM Journal on Applied Mathematics, 42(1):27–43, February 1982. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Homentcovschi:1982:UAS

[1968] Dorel Homentcovschi. Uniform asymptotic solutions for the potential field around a thin oblate body of revolution. SIAM Journal on Applied Mathematics, 42(1):44–65, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bertsch:1982:ABS

[1969] M. Bertsch. Asymptotic behavior of solutions of a nonlinear diffusion equation. SIAM Journal on Applied Mathematics, 42(1):66–76, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lakin:1982:URV

[1970] W. D. Lakin and A. Nachman. Uniform real-variable asymptotic approximations for vibrations of a rotating flexible rod. SIAM Journal on Applied Mathematics, 42(1):77–85, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rusinek:1982:REL

[1971] Roza Rusinek. Rate of extinction and limiting distribution for a geographically structured population. SIAM Journal on Applied Mathematics, 42 (1):86–93, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gurtin:1982:PCT

[1972] Morton E. Gurtin and Daniel S. Levine. On populations that cannibalize their young. SIAM Journal on Applied Mathematics, 42(1):94–108, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miles:1982:NBE

[1973] John W. Miles. On a nonlinear Bessel equation. SIAM Journal on Applied Mathematics, 42(1):109–112, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gilsinn:1982:HOG

[1974] David E. Gilsinn. A high order generalized method of averaging. SIAM Journal on Applied Mathematics, 42 (1):113–134, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Watson:1982:STI

[1975] John G. Watson and Edward L. Reiss. A statistical theory for imperfect bifurcation. SIAM Journal on Applied Mathematics, 42(1):135–148, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Williams:1982:AET

[1976] Michael Williams. Asymptotic exit time distributions. SIAM Journal on Applied Mathematics, 42(1):149–154, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Spence:1982:NEE

[1977] D. A. Spence. A note on the eigenfunction expansion for the elastic strip. SIAM Journal on Applied Mathematics, 42(1):155–173, February 1982. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bobrovsky:1982:SPM

[1978] B. Z. Bobrovsky and Z. Schuss. A singular perturbation method for the computation of the mean first passage time in a nonlinear filter. SIAM Journal on Applied Mathematics, 42 (1):174–187, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Britton:1982:TPS

[1979] N. F. Britton. Threshold phenomena and solitary traveling waves in a class of reaction-diffusion systems. SIAM Journal on Applied Mathematics, 42 (1):188–217, February 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Evans:1982:DIS

[1980] John W. Evans, Neil Fenichel, and John A. Feroes. Double impulse solutions in nerve axon equations. SIAM Journal on Applied Mathematics, 42 (2):219–234, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Feroe:1982:ESM

[1981] John A. Feroe. Existence and stability of multiple impulse solutions of a nerve equation. SIAM Journal on Applied Mathematics, 42(2):235–246, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hastings:1982:SMP

[1982] Stuart P. Hastings. Single and multiple pulse waves for the Fitzhugh– Nagumo equations. SIAM Journal on Applied Mathematics, 42(2):247–260, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Srivastava:1982:NRR

[1983] R. C. Srivastava. A note on the Rao– Rubin characterization of the Poisson distribution. SIAM Journal on Applied Mathematics, 42(2):261–265, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Buhite:1982:TEP

[1984] Jack L. Buhite. Two examples of pathological elastic-plastic behavior. SIAM Journal on Applied Mathematics, 42(2):266–271, April 1982. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dassios:1982:LFS

[1985] George Dassios. Low frequency scattering theory for a penetrable body with an impenetrable core. SIAM Journal on Applied Mathematics, 42 (2):272–280, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lotstedt:1982:MSR

[1986] Per Lötstedt. Mechanical systems of rigid bodies subject to unilateral constraints. SIAM Journal on Applied Mathematics, 42(2):281–296, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jensen:1982:IJD

[1987] D. R. Jensen. Inequalities for joint distributions of quadratic forms. SIAM Journal on Applied Mathematics, 42 (2):297–301, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chew:1982:MCC

[1988] W. C. Chew and J. A. Kong. Microstrip capacitance for a circular disk through matched asymptotic expansions. SIAM Journal on Applied Mathematics, 42 (2):302–317, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fokas:1982:ESE

[1989] A. S. Fokas and Y. C. Yortsos. On the exactly solvable equation S_{t} = [(βS + γ)^{−2}S_{x}]_{x} + α(βS + γ)_{x}^{−2S} occurring in two-phase flow in porous media. SIAM Journal on Applied Mathematics, 42(2):318–332, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Luke:1982:MCV

[1990] Jon C. Luke. A method for the calculation of vesicle shapes. SIAM Journal on Applied Mathematics, 42 (2):333–345, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [2044].

Gray:1982:TPS

[1991] Samuel H. Gray and Frank Hagin. Toward precise solution of onedimensional velocity inverse problems. SIAM Journal on Applied Mathematics, 42(2):346–355, April 1982. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sirovich:1982:CEP

[1992] L. Sirovich and B. W. Knight. Contributions to the eigenvalue problem for slowly varying operators. SIAM Journal on Applied Mathematics, 42 (2):356–377, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knight:1982:WTS

[1993] B. W. Knight and L. Sirovich. The Wigner transform and some exact properties of linear operators. SIAM Journal on Applied Mathematics, 42 (2):378–389, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hall:1982:IRE

[1994] Peter Hall. The influence of rounding errors on some nonparametric estimators of a density and its derivatives. SIAM Journal on Applied Mathematics, 42(2):390–399, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Glowinski:1982:NSP

[1995] R. Glowinski and P. Le Tallec. Numerical solution of problems in incompressible finite elasticity by augmented Lagrangian methods. I. Two-dimensional and axisymmetric problems. SIAM Journal on Applied Mathematics, 42 (2):400–429, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bayliss:1982:BCN

[1996] Alvin Bayliss, Max Gunzburger, and Eli Turkel. Boundary conditions for the numerical solution of elliptic equations in exterior regions. SIAM Journal on Applied Mathematics, 42(2):430–451, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bernard:1982:MCT

[1997] Peter S. Bernard and Bruce S.

Berger. A method for computing threedimensional turbulent flows. SIAM Journal on Applied Mathematics, 42 (3):453–470, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Videc:1982:GOF

[1998] M. F. Videc and C. Kooy. On the geometric optic formulation of radiation in a rotating medium. SIAM Journal on Applied Mathematics, 42(3):471–485, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1982:TWA

[1999] B. J. Matkowsky and D. O. Olagunju.

Traveling waves along the front of a pulsating flame. SIAM Journal on Applied Mathematics, 42(3):486–501, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lange:1982:SPA

[2000] Charles G. Lange and Robert M. Miura. Singular perturbation analysis of boundary-value problems for differential-difference equations. SIAM Journal on Applied Mathematics, 42 (3):502–531, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sarkar:1982:AAW

[2001] Purnā Sarkār. Asymptotic analysis of waves due to oscillatory surface pressure over a sloping beach. SIAM Journal on Applied Mathematics, 42(3):532–541, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Falkovitz:1982:SAR

[2002] Meira S. Falkovitz and Lee A. Segel. Some analytic results concerning the accuracy of the continuous approximation in a polymerization problem. SIAM Journal on Applied Mathematics, 42(3):542–548, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Markowich:1982:AAK

[2003] Peter A. Markowich. Asymptotic analysis of von Kármán flows. SIAM Journal on Applied Mathematics, 42 (3):549–557, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carasso:1982:DST

[2004] Alfred Carasso. Determining surface temperatures from interior observations. SIAM Journal on Applied Mathematics, 42(3):558–574, June 1982.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Manoukian:1982:BAB

[2005] E. B. Manoukian. Bound on the accuracy of Bartlett’s chi-square approximation for testing the homogeneity of variances. SIAM Journal on Applied Mathematics, 42(3):575–587, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1982:SIE

[2006] J. A. Morrison. The solution of an integral equation related to an optimal linear estimation problem. SIAM Journal on Applied Mathematics, 42 (3):588–607, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1982:DLE

[2007] Donald S. Cohen and S. Rosenblat.

A delay logistic equation with variable growth rate. SIAM Journal on Applied Mathematics, 42(3):608–624, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lu:1982:AAP

[2008] G. C. Lu and G. S. S. Ludford. Asymptotic analysis of plane steady detonations. SIAM Journal on Applied Mathematics, 42(3):625–635, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Idemen:1982:ITK

[2009] Mithat İdemen. On an integral transform with kernel Q_{ν−1/2}^{−μ}(cos θ) and its applications to second order canonical problems of GTD. SIAM Journal on Applied Mathematics, 42(3):636–652, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dixit:1982:SSF

[2010] P. Dixit and D. D. Joseph. The shape of stress-free surfaces on a sheared block. SIAM Journal on Applied Mathematics, 42(3):653–677, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Andersen:1982:PSE

[2011] C. M. Andersen and James F. Geer. Power series expansions for the frequency and period of the limit cycle of the Van Der Pol equation. SIAM Journal on Applied Mathematics, 42 (3):678–693, June 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Geman:1982:ASS

[2012] Stuart Geman. Almost sure stable oscillations in a large system of randomly coupled equations. SIAM Journal on Applied Mathematics, 42 (4):695–703, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Browning:1982:PDT

[2013] G. Browning and H. O. Kreiss. Problems with different time scales for nonlinear partial differential equations. SIAM Journal on Applied Mathematics, 42(4):704–718, August 1982. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tan:1982:DTN

[2014] W. Y. Tan. On distribution theories for the number of mutants in cell populations. SIAM Journal on Applied Mathematics, 42(4):719–730, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Steele:1982:LCS

[2015] J. Michael Steele. Long common subsequences and the proximity of two random strings. SIAM Journal on Applied Mathematics, 42(4):731–737, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Benson:1982:MCF

[2016] Donald C. Benson. Motion in a central force field with drag or tangential propulsion. SIAM Journal on Applied Mathematics, 42(4):738–750, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hoffman:1982:MSV

[2017] Dale T. Hoffman. A model for strategic voting. SIAM Journal on Applied Mathematics, 42(4):751–761, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hagan:1982:SWR

[2018] Patrick S. Hagan. Spiral waves in reaction-diffusion equations. SIAM Journal on Applied Mathematics, 42 (4):762–786, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matalon:1982:SBS

[2019] Moshe Matalon. The steady burning of a solid particle. SIAM Journal on Applied Mathematics, 42(4):787–803, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Seymour:1982:ESD

[2020] Brian R. Seymour and Eric Varley. Exact solutions describing soliton-like interactions in a nondispersive medium. SIAM Journal on Applied Mathematics, 42(4):804–821, August 1982. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1982:DAC

[2021] B. J. Matkowsky and Z. Schuss. Diffusion across characteristic boundaries. SIAM Journal on Applied Mathematics, 42(4):822–834, August 1982. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1982:SPA

[2022] B. J. Matkowsky, Z. Schuss, and E. Ben-Jacobs. A singular perturbation approach to Kramers’ diffusion problem. SIAM Journal on Applied Mathematics, 42(4):835–849, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

vanHarten:1982:NMF

[2023] A. van Harten and B. J. Matkowsky. A new model in flame theory. SIAM Journal on Applied Mathematics, 42 (4):850–867, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schmitt:1982:SLD

[2024] Bruno V. Schmitt. Sur la structure de L’Équation de Duffing sans dissipation. (French) [on the structure of the Duffing equation without dissipation]. SIAM Journal on Applied Mathematics, 42(4):868–894, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wilken:1982:SRC

[2025] D. R. Wilken. Some remarks on a competing predators problem. SIAM Journal on Applied Mathematics, 42 (4):895–902, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mark:1982:PPAa

[2026] James W. K. Mark and Lloyd Smith.

Plasmas in particle accelerators: The warm-beam equilibria. SIAM Journal on Applied Mathematics, 42(4):903– 913, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mark:1982:PPAb

[2027] James W. K. Mark. Plasmas in particle accelerators: Adiabatic theories for bunched beams. SIAM Journal on Applied Mathematics, 42(4):914–922, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shantaram:1982:CN

[2028] R. Shantaram. On a conjecture of Norton’s. SIAM Journal on Applied Mathematics, 42(4):923–925, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kassoy:1982:NAM

[2029] D. R. Kassoy. A note on asymptotic methods for jump phenomena. SIAM Journal on Applied Mathematics, 42 (4):926–932, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Byrne:1982:RPI

[2030] Charles L. Byrne and Raymond M. Fitzgerald. Reconstruction from partial information with applications to tomography. SIAM Journal on Applied Mathematics, 42(4):933–940, August 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grunbaum:1982:DOC

[2031] F. A. Grünbaum, L. Longhi, and M. Perlstadt. Differential operators commuting with finite convolution integral operators: Some non-Abelian examples. SIAM Journal on Applied Mathematics, 42(5):941–955, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dellwo:1982:BI

[2032] David Dellwo, Herbert B. Keller, Bernard J. Matkowsky, and Edward L. Reiss. On the birth of isolas. SIAM Journal on Applied Mathematics, 42 (5):956–963, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Teng:1982:RPR

[2033] Zhen Huan Teng, Alexandre Joel Chorin, and Tai Ping Liu. Riemann problems for reacting gas, with applications to transition. SIAM Journal on Applied Mathematics, 42(5):964– 981, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Margolis:1982:FPM

[2034] Stephen B. Margolis and Bernard J.

Matkowsky. Flame propagation with multiple fuels. SIAM Journal on Applied Mathematics, 42(5):982–1003, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klamer:1982:PIR

[2035] Dale M. Klamer and Elias Masry. Polynomial interpolation of randomly sampled bandlimited functions and processes. SIAM Journal on Applied Mathematics, 42(5):1004–1019, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Roehner:1982:SBD

[2036] Bertrand Roehner and Galliano Valent. Solving the birth and death processes with quadratic asymptotically symmetric transition rates. SIAM Journal on Applied Mathematics, 42 (5):1020–1046, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vega:1982:SLH

[2037] José M. Vega and Amable Li nán. Singular Langmuir–Hinshelwood reactiondiffusion problems: Strong adsorption under quasi-isothermal conditions. SIAM Journal on Applied Mathematics, 42(5):1047–1068, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Figari:1982:MFG

[2038] R. Figari, E. Orlandi, and G. Papanicolaou. Mean field and Gaussian approximation for partial differential equations with random coefficients. SIAM Journal on Applied Mathematics, 42 (5):1069–1077, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bajaj:1982:BPS

[2039] A. K. Bajaj. Bifurcating periodic solutions in rotationally symmetric systems. SIAM Journal on Applied Mathematics, 42(5):1078–1098, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Persek:1982:CTW

[2040] Stephen C. Persek. Chaotic traveling waves in general autonomous systems of nonlinear partial differential equations. SIAM Journal on Applied Mathematics, 42(5):1099–1110, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rinzel:1982:PPB

[2041] John Rinzel and David Terman. Propagation phenomena in a bistable reaction-diffusion system. SIAM Journal on Applied Mathematics, 42(5): 1111–1137, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1982:SWG

[2042] B. J. Matkowsky and D. O. Olagunju. Spinning waves in gaseous combustion. SIAM Journal on Applied Mathematics, 42(5):1138–1156, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kumei:1982:WND

[2043] Sukeyuki Kumei and George W.

Bluman. When nonlinear differential equations are equivalent to linear differential equations. SIAM Journal on Applied Mathematics, 42(5): 1157–1173, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Luke:1982:EMC

[2044] Jon C. Luke. Erratum: “A Method for the Calculation of Vesicle Shapes” [SIAM J. Appl. Math. 42 (1982), no. 2, 333–345]. SIAM Journal on Applied Mathematics, 42(5):1174, October 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1990].

Margolis:1982:FPS

[2045] Stephen B. Margolis and Bernard J.

Matkowsky. Flame propagation with a sequential reaction mechanism. SIAM Journal on Applied Mathematics, 42 (6):1175–1188, December 1982. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huilgol:1982:MPB

[2046] R. R. Huilgol. A multi-parameter bifurcation theorem. SIAM Journal on Applied Mathematics, 42(6):1189–1194, December 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wilson:1982:LCM

[2047] D. G. Wilson. Lagrangian coordinates for moving boundary problems. SIAM Journal on Applied Mathematics, 42(6):1195–1201, December 1982.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pedley:1982:NCD

[2048] T. J. Pedley. Natural convection driven by osmosis. SIAM Journal on Applied Mathematics, 42(6):1202–1216, December 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Banks:1982:UCF

[2049] W. H. H. Banks and M. Campbell. The unsteady collision of free-convective boundary layers. SIAM Journal on Applied Mathematics, 42(6):1217–1230, December 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burridge:1982:SSV

[2050] Robert Burridge, Jay Kappraff, and Christine Morshedi. The sitar string, a vibrating string with a one-sided inelastic constraint. SIAM Journal on Applied Mathematics, 42(6):1231–1251, December 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lacey:1982:WTS

[2051] A. A. Lacey, J. R. Ockendon, and A. B. Tayler. “waiting-time” solutions of a nonlinear diffusion equation. SIAM Journal on Applied Mathematics, 42(6):1252–1264, December 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dhariyal:1982:ELM

[2052] Ishwari D. Dhariyal, Edward J. Dudewicz, and Saul Blumenthal. Estimation of the larger mean. SIAM Journal on Applied Mathematics, 42(6): 1265–1275, December 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wollkind:1982:KHI

[2053] David J. Wollkind and J. Iwan D.

Alexander. Kelvin–Helmholtz instability in a layered Newtonian fluid model of the geological phenomenon of rock folding. SIAM Journal on Applied Mathematics, 42(6):1276–1295, December 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [2100].

Louton:1982:GMT

[2054] Tom Louton. The Gauss–Markov theorem for nonlinear models. SIAM Journal on Applied Mathematics, 42(6): 1296–1301, December 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vogelius:1982:PMA

[2055] Michael Vogelius and George Papanicolaou. A projection method applied to diffusion in a periodic structure. SIAM Journal on Applied Mathematics, 42(6):1302–1322, December 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chandna:1982:RPF

[2056] O. P. Chandna, R. M. Barron, and A. C. Smith. Rotational plane flows of a viscous fluid. SIAM Journal on Applied Mathematics, 42(6):1323–1336, December 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Christov:1982:COS

[2057] C. I. Christov. A complete orthonormal system of functions in L^{2}(−∞, ∞) space. SIAM Journal on Applied Mathematics, 42(6):1337–1344, December 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Steen:1982:QBN

[2058] Paul H. Steen and Stephen H. Davis. Quasiperiodic bifurcation in nonlinearly-coupled oscillators near a point of strong resonance. SIAM Journal on Applied Mathematics, 42(6): 1345–1368, December 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Olmstead:1983:ICH

[2059] W. E. Olmstead. Ignition of a combustible half space. SIAM Journal on Applied Mathematics, 43(1):1–15, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Magnan:1983:PBS

[2060] J. F. Magnan and R. A. Goldstein. Perturbed bifurcation of stationary striations in a contaminated, nonuniform plasma. SIAM Journal on Applied Mathematics, 43(1):16–30, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Karlin:1983:CDP

[2061] Samuel Karlin and Simon Tavaré. A class of diffusion processes with killing arising in population genetics. SIAM Journal on Applied Mathematics, 43 (1):31–41, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wen:1983:EAC

[2062] Shih Liang Wen and Harold Grad.

Equilibrium and adiabatic compression of a free boundary doublet. SIAM Journal on Applied Mathematics, 43 (1):42–56, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reiss:1983:CB

[2063] Edward L. Reiss. Cascading bifurcations. SIAM Journal on Applied Mathematics, 43(1):57–65, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Murray:1983:NCD

[2064] J. D. Murray and J. E. R. Cohen. On nonlinear convective dispersal effects in an interacting population model. SIAM Journal on Applied Mathematics, 43 (1):66–78, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gorell:1983:TOP

[2065] Sheldon B. Gorell and G. M. Homsy. A theory of the optimal policy of oil recovery by secondary displacement processes. SIAM Journal on Applied Mathematics, 43(1):79–98, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mangel:1983:OSM

[2066] Marc Mangel. Optimal search for and mining of underwater mineral resources. SIAM Journal on Applied Mathematics, 43(1):99–106, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Finden:1983:AAS

[2067] W. F. Finden. An asymptotic approximation for singular perturbations. SIAM Journal on Applied Mathematics, 43(1):107–119, February 1983. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wingate:1983:NSB

[2068] Charles A. Wingate. Numerical search for a ϕ^{4} breather mode. SIAM Journal on Applied Mathematics, 43(1): 120–140, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neu:1983:RIW

[2069] John C. Neu. Resonantly interacting waves. SIAM Journal on Applied Mathematics, 43(1):141–156, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Madych:1983:PBA

[2070] W. R. Madych and S. A. Nelson. Polynomial based algorithms for computed tomography. SIAM Journal on Applied Mathematics, 43(1):157–185, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berntsen:1983:DEP

[2071] Svend Berntsen. Diffraction of an electric polarized wave by a dielectric wedge. SIAM Journal on Applied Mathematics, 43(1):186–211, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barshinger:1983:PFA

[2072] Richard N. Barshinger and James F. Geer. Potential flow around a thin oblate body of revolution. SIAM Journal on Applied Mathematics, 43 (1):212–224, February 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maginu:1983:SHI

[2073] Kenjiro Maginu. Spatially homogeneous and inhomogeneous oscillations and chaotic motion in the active Josephson junction line. SIAM Journal on Applied Mathematics, 43 (2):225–243, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Haberman:1983:EBS

[2074] Richard Haberman. Energy bounds for the slow capture by a center in sustained resonance. SIAM Journal on Applied Mathematics, 43(2):244–256, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Whittaker:1983:SID

[2075] James V. Whittaker. On scaleinvariant distributions. SIAM Journal on Applied Mathematics, 43(2):257– 267, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keller:1983:STD

[2076] Joseph B. Keller and Michael J. Miksis. Surface tension driven flows. SIAM Journal on Applied Mathematics, 43 (2):268–277, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hummel:1983:SSR

[2077] Robert Hummel. Sampling for spline reconstruction. SIAM Journal on Applied Mathematics, 43(2):278–288, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Davis:1983:SFO

[2078] J. H. Davis and R. G. Dickinson. Spectral factorization by optimal gain iteration. SIAM Journal on Applied Mathematics, 43(2):289–301, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Janardan:1983:LDS

[2079] Konanur G. Janardan and B. Raja Rao. Lagrange distributions of the second kind and weighted distributions. SIAM Journal on Applied Mathematics, 43(2):302–313, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kath:1983:NCS

[2080] William L. Kath. Necessary conditions for sustained reentry roll resonance. SIAM Journal on Applied Mathematics, 43(2):314–324, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Delebecque:1983:RPP

[2081] F. Delebecque. A reduction process for perturbed Markov chains. SIAM Journal on Applied Mathematics, 43 (2):325–350, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Margolis:1983:ATC

[2082] Stephen B. Margolis. An asymptotic theory of condensed two-phase flame propagation. SIAM Journal on Applied Mathematics, 43(2):351–369, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1983:HBS

[2083] Hal Smith. Hopf bifurcation in a system of functional equations modeling the spread of an infectious disease. SIAM Journal on Applied Mathematics, 43(2):370–385, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Falkovitz:1983:SIP

[2084] Meira S. Falkovitz and Lee A. Segel.

Spatially inhomogeneous polymerization in unstirred bulk. SIAM Journal on Applied Mathematics, 43(2):386– 416, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Capasso:1983:PSR

[2085] V. Capasso and L. Maddalena. Periodic solutions for a reaction-diffusion system modelling the spread of a class of epidemics. SIAM Journal on Applied Mathematics, 43(2):417–427, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Davison:1983:ICN

[2086] Mark E. Davison. The ill-conditioned nature of the limited angle tomography problem. SIAM Journal on Applied Mathematics, 43(2):428–448, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nagai:1983:ABN

[2087] Toshitaka Nagai and Masayasu Mimura. Asymptotic behavior for a nonlinear degenerate diffusion equation in population dynamics. SIAM Journal on Applied Mathematics, 43(3):449–464, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lenhart:1983:SNP

[2088] Suzanne M. Lenhart and Stavros A. Belbas. A system of nonlinear partial differential equations arising in the optimal control of stochastic systems with switching costs. SIAM Journal on Applied Mathematics, 43(3):465–475, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Karr:1983:EBR

[2089] Alan F. Karr and Robert J. Serfling. Error bounds for reconstruction of a function f from a finite sequence {sgn(f(t_{i}) + x_{i})}. SIAM Journal on Applied Mathematics, 43(3):476–490, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kapila:1983:ATD

[2090] A. K. Kapila, B. J. Matkowsky, and A. Van Harten. An asymptotic theory of deflagrations and detonations. I. The steady solutions. SIAM Journal on Applied Mathematics, 43(3):491–519, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Skjolding:1983:BDD

[2091] Henrik Skjolding, Bodil Branner-Jørgensen, Peter L. Christiansen, and Helge E. Jensen. Bifurcations in discrete dynamical systems with cubic maps. SIAM Journal on Applied Mathematics, 43(3):520–534, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Glasser:1983:EEE

[2092] M. L. Glasser and J. Boersma. Exchange energy of an electron gas of arbitrary dimensionality. SIAM Journal on Applied Mathematics, 43(3):535– 545, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tuy:1983:IFC

[2093] Heang K. Tuy. An inversion formula for cone-beam reconstruction. SIAM Journal on Applied Mathematics, 43 (3):546–552, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cannon:1983:IPU

[2094] J. R. Cannon and P. Duchateau. An inverse problem for an unknown source term in a wave equation. SIAM Journal on Applied Mathematics, 43(3):553– 564, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Benzi:1983:TSR

[2095] Roberto Benzi, Giorgio Parisi, and Angelo Vulpiani. A theory of stochastic resonance in climatic change. SIAM Journal on Applied Mathematics, 43 (3):565–578, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kath:1983:CSR

[2096] William L. Kath. Conditions for sustained resonance. II. SIAM Journal on Applied Mathematics, 43(3):579–583, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cho:1983:CPS

[2097] S. K. Cho. On complex poles in scattering theory. SIAM Journal on Applied Mathematics, 43(3):584–593, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

MacGillivray:1983:LTI

[2098] A. D. MacGillivray. On the leading term of the inner asymptotic expansion of Van Der Pol’s equation. SIAM Journal on Applied Mathematics, 43(3):594–612, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Erneux:1983:SSM

[2099] Thomas Erneux and Edward L. Reiss. Splitting of steady multiple eigenvalues may lead to periodic cascading bifurcation. SIAM Journal on Applied Mathematics, 43(3):613–624, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wollkind:1983:EKH

[2100] David J. Wollkind and J. Iwan D. Alexander. Erratum: “Kelvin– Helmholtz Instability in a Layered Newtonian Fluid Model of the Geological Phenomenon of Rock Folding” [SIAM J. Appl. Math. 42 (1982), no. 6, 1276–1295, MR 83j:86004]. SIAM Journal on Applied Mathematics, 43 (3):625, June 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [2053].

Hansen:1983:E

[2101] Erik B. Hansen and Joseph A. Mc-Geough. On electropainting. SIAM Journal on Applied Mathematics, 43 (4):627–638, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1983:VTD

[2102] R. E. Meyer. A view of the triple deck. SIAM Journal on Applied Mathematics, 43(4):639–663, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ludwig:1983:PA

[2103] Donald Ludwig. Parsimonious asymptotics. SIAM Journal on Applied Mathematics, 43(4):664–672, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matkowsky:1983:DAC

[2104] B. J. Matkowsky, Z. Schuss, and C. Tier. Diffusion across characteristic boundaries with critical points. SIAM Journal on Applied Mathematics, 43 (4):673–695, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

VanDyke:1983:LFM

[2105] Milton Van Dyke. Laminar flow in a meandering channel. SIAM Journal on Applied Mathematics, 43(4):696–702, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Whittemore:1983:TLB

[2106] Alice S. Whittemore. Transformations to linearity in binary regression. SIAM Journal on Applied Mathematics, 43 (4):703–710, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Twersky:1983:SNO

[2107] Victor Twersky. Scattering and nonscattering obstacles. SIAM Journal on Applied Mathematics, 43(4):711–725, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lahlou:1983:HAI

[2108] Mourad Lahlou, Jack K. Cohen, and Norman Bleistein. Highly accurate inversion methods for three-dimensional stratified media. SIAM Journal on Applied Mathematics, 43(4):726–758, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ting:1983:SCM

[2109] Lu Ting, Daljit S. Ahluwalia, and Michael J. Miksis. Solutions of a class of mixed free boundary problems. SIAM Journal on Applied Mathematics, 43(4):759–775, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Coullet:1983:AES

[2110] P. H. Coullet and E. A. Spiegel. Amplitude equations for systems with competing instabilities. SIAM Journal on Applied Mathematics, 43(4):776–821, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shen:1983:RMF

[2111] M. C. Shen. Ray method for flow of a compressible viscous fluid. SIAM Journal on Applied Mathematics, 43 (4):822–833, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hoppensteadt:1983:AAS

[2112] F. C. Hoppensteadt. An algorithm for approximate solutions to weakly filtered synchronous control systems and nonlinear renewal processes. SIAM Journal on Applied Mathematics, 43 (4):834–843, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morawetz:1983:CIP

[2113] Cathleen S. Morawetz and Gregory A.

Kriegsmann. The calculations of an inverse potential problem. SIAM Journal on Applied Mathematics, 43(4):844– 854, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Heyman:1983:EWC

[2114] E. Heyman and L. B. Felsen. Evanescent waves and complex rays for modal propagation in curved open waveguides. SIAM Journal on Applied Mathematics, 43(4):855–884, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Caflisch:1983:DTS

[2115] R. Caflisch and G. C. Papanicolaou. Dynamic theory of suspensions with Brownian effects. SIAM Journal on Applied Mathematics, 43(4):885–906, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rinzel:1983:HBR

[2116] John Rinzel and James P. Keener. Hopf bifurcation to repetitive activity in nerve. SIAM Journal on Applied Mathematics, 43(4):907–922, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kriegsmann:1983:LFS

[2117] Gregory A. Kriegsmann and Edward L. Reiss. Low frequency scattering by local inhomogeneities. SIAM Journal on Applied Mathematics, 43(4):923–934, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

OMalley:1983:STL

[2118] Robert E. O’Malley, Jr. Shock and transition layers for singularly perturbed second-order vector systems. SIAM Journal on Applied Mathematics, 43(4):935–943, August 1983. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cole:1983:NAS

[2119] Julian D. Cole. Note on the axisymmetric sonic jet. SIAM Journal on Applied Mathematics, 43(4):944–948, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1983:GDG

[2120] Donald S. Cohen and E. Ann Stanley. Gaseous diffusion in glassy polymers. SIAM Journal on Applied Mathematics, 43(4):949–970, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gutkin:1983:PCB

[2121] E. Gutkin and M. Kac. Propagation of chaos and the Burgers equation. SIAM Journal on Applied Mathematics, 43 (4):971–980, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keller:1983:FSR

[2122] H. B. Keller and D. J. Perozzi. Fast seismic ray tracing. SIAM Journal on Applied Mathematics, 43(4):981–992, August 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Howes:1983:ASC

[2123] F. A. Howes. The asymptotic solution of a class of third-order boundary value problems arising in the theory of thin film flows. SIAM Journal on Applied Mathematics, 43(5):993–1004, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keener:1983:OCC

[2124] James P. Keener. Oscillatory coexistence in the chemostat: A codimension two unfolding. SIAM Journal on Applied Mathematics, 43(5): 1005–1018, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ghil:1983:GHB

[2125] Michael Ghil and John Tavantzis. Global Hopf bifurcation in a simple climate model. SIAM Journal on Applied Mathematics, 43(5):1019–1041, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Erneux:1983:IBN

[2126] T. Erneux and D. S. Cohen. Imperfect bifurcation near a double eigenvalue: Transitions between nonsymmetric and symmetric patterns. SIAM Journal on Applied Mathematics, 43 (5):1042–1060, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1983:TBN

[2127] D. S. Cohen and T. Erneux. Transitions between nonsymmetric and symmetric steady states near a triple eigenvalue. SIAM Journal on Applied Mathematics, 43(5):1061–1074, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Roose:1983:NCH

[2128] D. Roose and V. Hlavacek. Numerical computation of Hopf bifurcation points for parabolic diffusion-reaction differential equations. SIAM Journal on Applied Mathematics, 43(5): 1075–1085, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosales:1983:WND

[2129] Rodolfo R. Rosales and Andrew Majda. Weakly nonlinear detonation waves.

SIAM Journal on Applied Mathematics, 43(5):1086–1118, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brown:1983:IEM

[2130] S. N. Brown and K. Stewartson. On an integral equation of marginal separation. SIAM Journal on Applied Mathematics, 43(5):1119–1126, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dubbelday:1983:ANC

[2131] Pieter S. Dubbelday. Application of a new complex root-finding technique to the dispersion relations for elastic waves in a fluid-loaded plate. SIAM Journal on Applied Mathematics, 43 (5):1127–1139, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miksis:1983:EDC

[2132] Michael J. Miksis. Effective dielectric constant of a nonlinear composite material. SIAM Journal on Applied Mathematics, 43(5):1140–1155, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maginu:1983:SWO

[2133] Kenjiro Maginu. Standing wave oscillations with periodical spatial structures in an active Josephson junction line and in reaction-diffusion systems. SIAM Journal on Applied Mathematics, 43 (5):1156–1173, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chandra:1983:SAC

[2134] Jagdish Chandra, G. S. Ladde, and V. Lakshmikantham. Stochastic analysis of a compressible gas lubrication slider bearing problem. SIAM Journal on Applied Mathematics, 43(5): 1174–1186, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hall:1983:UAC

[2135] Peter Hall. A unified approach to the correction of normal approximations. SIAM Journal on Applied Mathematics, 43(5):1187–1193, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jensen:1983:RED

[2136] D. R. Jensen and I. J. Good. A representation for ellipsoidal distributions. SIAM Journal on Applied Mathematics, 43(5):1194–1200, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Natterer:1983:CTU

[2137] Frank Natterer. Computerized tomography with unknown sources. SIAM Journal on Applied Mathematics, 43 (5):1201–1212, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hoffman:1983:REV

[2138] Dale T. Hoffman. Relative efficiency of voting systems. SIAM Journal on Applied Mathematics, 43(5):1213–1219, October 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

MacGillivray:1983:LTO

[2139] A. D. MacGillivray. On the leading term of the outer asymptotic expansion of Van Der Pol’s equation. SIAM Journal on Applied Mathematics, 43(6):1221–1239, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Erneux:1983:BI

[2140] Thomas Erneux and Edward L. Reiss.

Brussellator isolas. SIAM Journal on Applied Mathematics, 43(6):1240–1246, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jing:1983:AQM

[2141] Zhu Jun Jing. Application of qualitative methods of differential equations to study phase-locked loops. SIAM Journal on Applied Mathematics, 43(6): 1247–1258, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bluman:1983:MLP

[2142] George W. Bluman. On mapping linear partial differential equations to constant coefficient equations. SIAM Journal on Applied Mathematics, 43(6): 1259–1273, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kersner:1983:NHC

[2143] Robert Kersner. Nonlinear heat conduction with absorption: Space localization and extinction in finite time. SIAM Journal on Applied Mathematics, 43(6):1274–1285, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nakamura:1983:ESE

[2144] Mitsunobu Nakamura and Masahide Hirasawa. Eigenvalues of the Schrödinger equation by the αinterpolation method. SIAM Journal on Applied Mathematics, 43(6): 1286–1293, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Joly:1983:TIP

[2145] G. Joly, J.-P. Kernévez, and M. Llory. Thermal instability in pool boiling on wires at constant pressure. SIAM Journal on Applied Mathematics, 43(6): 1294–1309, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Majda:1983:TSM

[2146] Andrew Majda and Rodolfo Rosales. A theory for spontaneous Mach stem formation in reacting shock fronts. I. The basic perturbation analysis. SIAM Journal on Applied Mathematics, 43(6):1310–1334, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Buckmaster:1983:SPP

[2147] J. Buckmaster. Stability of the porous plug burner flame. SIAM Journal on Applied Mathematics, 43(6):1335–1349, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lacey:1983:MAT

[2148] A. A. Lacey. Mathematical analysis of thermal runaway for spatially inhomogeneous reactions. SIAM Journal on Applied Mathematics, 43(6):

1350–1366, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vega:1983:SLH

[2149] José M. Vega. Singular Langmuir– Hinshelwood reaction-diffusion problems: Strongly nonisothermal conditions. SIAM Journal on Applied Mathematics, 43(6):1367–1389, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Krafft:1983:PPA

[2150] Geoffrey A. Krafft, James W.-K. Mark, and Tai Sen F. Wang. Plasmas in particle accelerators: A hydrodynamic model of three-dimensional electrostatic instabilities. SIAM Journal on Applied Mathematics, 43(6): 1390–1399, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Censor:1983:SWN

[2151] Dan Censor. Scattering by weakly nonlinear objects. SIAM Journal on Applied Mathematics, 43(6):1400–1417, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gyllenberg:1983:SNA

[2152] Mats Gyllenberg. Stability of a nonlinear age-dependent population model containing a control variable. SIAM Journal on Applied Mathematics, 43 (6):1418–1438, December 1983. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chow:1983:RMC

[2153] P. L. Chow. Reconstruction of the mutual coherence function for a moving source. SIAM Journal on Applied Mathematics, 43(6):1439–1445, December 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Homentcovschi:1984:UAS

[2154] Dorel Homentcovschi. Uniform asymptotic solutions of two-dimensional problems of elasticity for the domain exterior to a thin region. SIAM Journal on Applied Mathematics, 44(1):1– 10, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burgan:1984:HNH

[2155] J. R. Burgan, A. Munier, M. R. Feix, and E. Fijalkow. Homology and the nonlinear heat diffusion equation. SIAM Journal on Applied Mathematics, 44(1):11–18, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barshinger:1984:SFP

[2156] Richard N. Barshinger and James F. Geer. Stokes flow past a thin oblate body of revolution: Axially incident uniform flow. SIAM Journal on Applied Mathematics, 44(1):19–32, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bhattacharya:1984:TAT

[2157] R. N. Bhattacharya and Vijay K. Gupta. On the Taylor–Aris theory of solute transport in a capillary. SIAM Journal on Applied Mathematics, 44 (1):33–39, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Buckmaster:1984:PFE

[2158] J. Buckmaster. Polyhedral flames — an exercise in bimodal bifurcation analysis. SIAM Journal on Applied Mathematics, 44(1):40–55, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gardner:1984:ETW

[2159] Robert A. Gardner. Existence of travelling wave solutions of predatorprey systems via the connection index. SIAM Journal on Applied Mathematics, 44(1):56–79, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ermentrout:1984:PDP

[2160] G. Bard Ermentrout. Period doublings and possible chaos in neural models. SIAM Journal on Applied Mathematics, 44(1):80–95, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klaasen:1984:SWS

[2161] Gene A. Klaasen and William C. Troy.

Stationary wave solutions of a system of reaction-diffusion equations derived from the Fitzhugh–Nagumo equations. SIAM Journal on Applied Mathematics, 44(1):96–110, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dee:1984:BDE

[2162] D. Dee and M. Ghil. Boolean difference equations, I: Formulation and dynamic behavior. SIAM Journal on Applied Mathematics, 44(1):111–126, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

White:1984:SC

[2163] B. S. White. The stochastic caustic. SIAM Journal on Applied Mathematics, 44(1):127–149, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Elmroth:1984:FCT

[2164] Tony Elmroth. On the H-function and convergence towards equilibrium for a space-homogeneous molecular density. SIAM Journal on Applied Mathematics, 44(1):150–159, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kushner:1984:RAE

[2165] Harold J. Kushner. Robustness and approximation of escape times and large deviations estimates for systems with small noise effects. SIAM Journal on Applied Mathematics, 44(1): 160–182, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abaya:1984:CVQ

[2166] Efren F. Abaya and Gary L. Wise. Convergence of vector quantizers with applications to optimal quantization. SIAM Journal on Applied Mathematics, 44(1):183–189, February 1984. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abrahams:1984:MSO

[2167] Julia Abrahams. On Miroshin’s secondorder reciprocal processes. SIAM Journal on Applied Mathematics, 44(1): 190–192, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Madych:1984:PBA

[2168] W. R. Madych and S. A. Nelson.

Polynomial based algorithms for computed tomography. II. SIAM Journal on Applied Mathematics, 44(1): 193–208, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cook:1984:RBM

[2169] Wade D. Cook and Moshe Kress. Relationships between l^{1} metrics on linear ranking spaces. SIAM Journal on Applied Mathematics, 44(1):209–220, February 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chikwendu:1984:MSS

[2170] S. C. Chikwendu. Multiple-scale solution of weakly nonlinear elliptic equations on the upper half-plane. SIAM Journal on Applied Mathematics, 44 (2):221–230, April 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Markowich:1984:SPB

[2171] Peter A. Markowich and C. A.

Ringhofer. A singularly perturbed boundary value problem modelling a semiconductor device. SIAM Journal on Applied Mathematics, 44(2):231– 256, April 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kolkka:1984:SPB

[2172] Robert W. Kolkka. Singular perturbations of bifurcations with multiple independent bifurcation parameters. SIAM Journal on Applied Mathematics, 44 (2):257–269, April 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bajaj:1984:FIB

[2173] A. K. Bajaj and P. R. Sethna. Flow induced bifurcations to threedimensional oscillatory motions in continuous tubes. SIAM Journal on Applied Mathematics, 44(2):270–286, April 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Forest:1984:MPK

[2174] M. Gregory Forest and David W. McLaughlin. Modulations of perturbed KdV wavetrains. SIAM Journal on Applied Mathematics, 44(2):287–300, April 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Parker:1984:FPD

[2175] David F. Parker and Frederic Y. M. Wan. Finite polar dimpling of shallow caps under sub-buckling axisymmetric pressure distributions. SIAM Journal on Applied Mathematics, 44 (2):301–326, April 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matalon:1984:SPC

[2176] M. Matalon and B. J. Matkowsky. On the stability of plane and curved flames.

SIAM Journal on Applied Mathematics, 44(2):327–343, April 1984. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Margolis:1984:FPV

[2177] Stephen B. Margolis and Gregory I.

Sivashinsky. Flame propagation in vertical channels: Bifurcation to bimodal cellular flames. SIAM Journal on Applied Mathematics, 44(2):344–368, April 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bonilla:1984:ROP

[2178] Luis L. Bonilla and Amable Liñán. Relaxation oscillations, pulses, and travelling waves in the diffusive Volterra delay-differential equation. SIAM Journal on Applied Mathematics, 44 (2):369–391, April 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cushing:1984:PTP

[2179] J. M. Cushing. Periodic two-predator, one-prey interactions and the time sharing of a resource niche. SIAM Journal on Applied Mathematics, 44 (2):392–410, April 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Winter:1984:PED

[2180] C. L. Winter, C. M. Newman, and S. P. Neuman. A perturbation expansion for diffusion in a random velocity field. SIAM Journal on Applied Mathematics, 44(2):411–424, April 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Byrne:1984:SEE

[2181] Charles L. Byrne and Raymond M. Fitzgerald. Spectral estimators that extend the maximum entropy and maximum likelihood methods. SIAM Journal on Applied Mathematics, 44 (2):425–442, April 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Heidtmann:1984:PEE

[2182] Klaus D. Heidtmann. A priori error estimates for the method of inclusionexclusion with applications. SIAM Journal on Applied Mathematics, 44 (2):443–450, April 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lagerstrom:1984:NLS

[2183] P. A. Lagerstrom and D. A. Reinelt. Note on logarithmic switchback terms in regular and singular perturbation expansions. SIAM Journal on Applied Mathematics, 44(3):451–462, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Erneux:1984:SSB

[2184] Thomas Erneux and Edward L. Reiss. Singular secondary bifurcation. SIAM Journal on Applied Mathematics, 44 (3):463–478, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Parry:1984:FKT

[2185] Gareth P. Parry. First kind transitions in ideal crystals. SIAM Journal on Applied Mathematics, 44(3):479–492, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Liron:1984:CFS

[2186] Nadav Liron and Jacob Rubinstein.

Calculating the fundamental solution to linear convection-diffusion problems. SIAM Journal on Applied Mathematics, 44(3):493–511, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neil:1984:EUB

[2187] Charles H. Neil. Existence, uniqueness and behavior of solutions of a singular nonlinear system from fluid dynamics. SIAM Journal on Applied Mathematics, 44(3):512–523, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kinoshita:1984:EPF

[2188] N. Kinoshita and T. Mura. Eigenstrain problems in a finite elastic body. SIAM Journal on Applied Mathematics, 44(3):524–535, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Erneux:1984:QPW

[2189] T. Erneux and B. J. Matkowsky. Quasiperiodic waves along a pulsating propagating front in a reaction-diffusion system. SIAM Journal on Applied Mathematics, 44(3):536–544, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Holmes:1984:CTS

[2190] Mark H. Holmes. Comparison theorems and similarity solution approximations for a nonlinear diffusion equation arising in the study of soft tissue.

SIAM Journal on Applied Mathematics, 44(3):545–556, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wilbur:1984:CDC

[2191] W. John Wilbur and David J. Lipman. The context dependent comparison of biological sequences. SIAM Journal on Applied Mathematics, 44(3):557–567, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bigg:1984:SPA

[2192] G. R. Bigg. Scattering and the parabolic approximation for slender bodies. SIAM Journal on Applied Mathematics, 44(3):568–586, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Proschan:1984:RAV

[2193] Frank Proschan and Moshe Shaked. Random averaging of vector elements. SIAM Journal on Applied Mathematics, 44(3):587–590, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Katzur:1984:AAOa

[2194] R. Katzur, B. Z. Bobrovsky, and Z. Schuss. Asymptotic analysis of the optimal filtering problem for onedimensional diffusions measured in a low noise channel, part I. SIAM Journal on Applied Mathematics, 44 (3):591–604, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hoskins:1984:SEF

[2195] R. F. Hoskins and J. De Sousa Pinto. Sampling expansions for functions band-limited in the distributional sense. SIAM Journal on Applied Mathematics, 44(3):605–610, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hoskins:1984:GSE

[2196] R. F. Hoskins and J. De Sousa Pinto.

Generalized sampling expansions in the sense of Papoulis. SIAM Journal on Applied Mathematics, 44(3):611–617, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Borden:1984:VIA

[2197] Brett Borden. A vector inverse algorithm for electromagnetic scattering. SIAM Journal on Applied Mathematics, 44(3):618–626, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Alliney:1984:ICV

[2198] S. Alliney and F. Sgallari. An “illconditioned” Volterra integral equation related to the reconstruction of images from projections. SIAM Journal on Applied Mathematics, 44(3):627–645, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chang:1984:ART

[2199] Jin Fu Chang and Rong Feng Chang.

The application of the residue theorem to the study of a finite queue with batch Poisson arrivals and synchronous servers. SIAM Journal on Applied Mathematics, 44(3):646–656, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Waldmann:1984:ICR

[2200] K.-H. Waldmann. Inventory control in randomly varying environments. SIAM Journal on Applied Mathematics, 44 (3):657–666, June 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Daniele:1984:STC

[2201] V. G. Daniele. On the solution of two coupled Wiener–Hopf equations. SIAM Journal on Applied Mathematics, 44 (4):667–680, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fowler:1984:NTC

[2202] A. C. Fowler and M. J. McGuinness. On the nature of the torus in the complex Lorenz equations. SIAM Journal on Applied Mathematics, 44 (4):681–700, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mosher:1984:EMF

[2203] M. Calvin Mosher. Ergodicity and mixing in four-mode spectral models of two-dimensional turbulence. SIAM Journal on Applied Mathematics, 44 (4):701–709, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Glowinski:1984:NSP

[2204] R. Glowinski and P. Le Tallec. Numerical solution of problems in incompressible finite elasticity by augmented Lagrangian methods. II. Threedimensional problems. SIAM Journal on Applied Mathematics, 44(4):710– 733, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Matalon:1984:EFM

[2205] Moshe Matalon and Thomas Erneux.

Expanding flames may delay the transition to cellular structures. SIAM Journal on Applied Mathematics, 44 (4):734–744, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mooney:1984:SSR

[2206] J. R. Mooney. Steady states of a reaction-diffusion system on the offcentre annulus. SIAM Journal on Applied Mathematics, 44(4):745–761, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hofbauer:1984:DEM

[2207] Josef Hofbauer. A difference equation model for the hypercycle. SIAM Journal on Applied Mathematics, 44 (4):762–772, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Letessier:1984:GFM

[2208] J. Letessier and G. Valent. The generating function method for quadratic asymptotically symmetric birth and death processes. SIAM Journal on Applied Mathematics, 44(4):773–783, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sellers:1984:CCC

[2209] Peter H. Sellers. Combinatorial classification of chemical mechanisms. SIAM Journal on Applied Mathematics, 44 (4):784–792, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arnold:1984:FCS

[2210] Ludwig Arnold. A formula connecting sample and moment stability of linear stochastic systems. SIAM Journal on Applied Mathematics, 44(4):793–802, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hamada:1984:UCC

[2211] N. Hamada and B. D. O. Anderson. The unit circle Cauchy index: Definition, characterization and polynomial zero distribution. SIAM Journal on Applied Mathematics, 44(4):803–818, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sanz:1984:CTC

[2212] Jorge L. C. Sanz and Thomas S. Huang. Continuation techniques for a certain class of analytic functions. SIAM Journal on Applied Mathematics, 44(4):819–838, August 1984. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cowan:1984:MRP

[2213] Richard Cowan. A model for random packing of disks in the neighbourhood of one disk. SIAM Journal on Applied Mathematics, 44(4):839–853, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Janardan:1984:MCS

[2214] Konanur G. Janardan. Moments of certain series distributions and their applications. SIAM Journal on Applied Mathematics, 44(4):854–868, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pinedo:1984:IBS

[2215] Michael Pinedo and Richard Weber.

Inequalities and bounds in stochastic shop scheduling. SIAM Journal on Applied Mathematics, 44(4):869–879, August 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dadfar:1984:PAL

[2216] Mohammad B. Dadfar, James Geer, and Carl M. Andersen. Perturbation analysis of the limit cycle of the free Van Der Pol equation. SIAM Journal on Applied Mathematics, 44 (5):881–895, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Markowich:1984:SPA

[2217] Peter A. Markowich. A singular perturbation analysis of the fundamental semiconductor device equations. SIAM Journal on Applied Mathematics, 44 (5):896–928, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kida:1984:NPA

[2218] Teruhiko Kida. A new perturbation approach to the laminar fluid flow behind a two-dimensional solid body. SIAM Journal on Applied Mathematics, 44 (5):929–951, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boyd:1984:CWE

[2219] John P. Boyd. Cnoidal waves as exact sums of repeated solitary waves: New series for elliptic functions. SIAM Journal on Applied Mathematics, 44 (5):952–955, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tseng:1984:THS

[2220] W. F. Tseng and S. P. Lin. Theory of a heat sensing velocimeter. SIAM Journal on Applied Mathematics, 44 (5):956–968, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Clements:1984:SBV

[2221] David L. Clements, D. L. Hill, C. Rogers, and J. Mazumdar. On some boundary value problems in anisotropic inhomogeneous thermostatics and elastostatics. SIAM Journal on Applied Mathematics, 44(5):969–981, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

vanHarten:1984:ACF

[2222] A. van Harten, A. K. Kapila, and B. J. Matkowsky. Acoustic coupling of flames. SIAM Journal on Applied Mathematics, 44(5):982–995, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Colton:1984:NSI

[2223] David Colton and Rembert Reemtsen. The numerical solution of the inverse Stefan problem in two space variables. SIAM Journal on Applied Mathematics, 44(5):996–1013, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vassiliou:1984:MSN

[2224] P.-C. G. Vassiliou and N. Tsantas.

Maintainability of structures in nonhomogeneous Markov systems under cyclic behavior and input control. SIAM Journal on Applied Mathematics, 44(5):1014–1022, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pourahmadi:1984:MIH

[2225] Mohsen Pourahmadi. On minimality and interpolation of harmonizable stable processes. SIAM Journal on Applied Mathematics, 44(5):1023–1030, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lloyd:1984:ODO

[2226] Stuart P. Lloyd. Optimal dual output extended hydrophone. SIAM Journal on Applied Mathematics, 44(5): 1031–1040, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Flatto:1984:TPQ

[2227] L. Flatto and S. Hahn. Two parallel queues created by arrivals with two demands. I. SIAM Journal on Applied Mathematics, 44(5):1041–1053, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [2252].

Hendy:1984:CTP

[2228] M. D. Hendy, C. H. C. Little, and David Penny. Comparing trees with pendant vertices labelled. SIAM Journal on Applied Mathematics, 44(5): 1054–1065, October 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Illner:1984:DEC

[2229] Reinhard Illner and Michael C. Reed. Decay to equilibrium for the Carleman model in a box. SIAM Journal on Applied Mathematics, 44(6): 1067–1075, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mills:1984:SMB

[2230] P. L. Mills and M. P. Duduković. Solution of mixed boundary value problems by integral equations and methods of weighted residuals with application to heat conduction and diffusion-reaction systems. SIAM Journal on Applied Mathematics, 44(6):1076–1091, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eckhoff:1984:DLW

[2231] Knut S. Eckhoff. On dispersion for linear waves in nonuniform media. SIAM Journal on Applied Mathematics, 44(6):1092–1105, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ioakimidis:1984:RSI

[2232] N. I. Ioakimidis. A remark on singular integral equations with generalized kernels. SIAM Journal on Applied Mathematics, 44(6):1106–1111, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cosner:1984:SCS

[2233] Chris Cosner and A. C. Lazer. Stable coexistence states in the Volterra– Lotka competition model with diffusion. SIAM Journal on Applied Mathematics, 44(6):1112–1132, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ermentrout:1984:LAS

[2234] G. B. Ermentrout, S. P. Hastings, and W. C. Troy. Large amplitude stationary waves in an excitable lateralinhibitory medium. SIAM Journal on Applied Mathematics, 44(6):1133–1149, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ben-Haim:1984:MSR

[2235] Yakov Ben-Haim and Natan Shenhav.

The measurement of spatially random phenomena. SIAM Journal on Applied Mathematics, 44(6):1150–1163, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Percus:1984:BKS

[2236] O. E. Percus and J. K. Percus. On the Bechhofer–Kulkarni stopping rule for sequential clinical trials. SIAM Journal on Applied Mathematics, 44(6): 1164–1175, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Katzur:1984:AAOb

[2237] R. Katzur, B. Z. Bobrovsky, and Z. Schuss. Asymptotic analysis of the optimal filtering problem for onedimensional diffusions measured in a low noise channel, part II. SIAM Journal on Applied Mathematics, 44(6): 1176–1191, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Xia:1984:DFS

[2238] Ning Mao Xia and William E. Boyce. The density function of the solution of a random initial value problem containing small stochastic processes. SIAM Journal on Applied Mathematics, 44(6):1192–1209, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Xu:1984:PDP

[2239] Wen Yuan Xu and Christodoulos Chamzas. On the periodic discrete prolate spheroidal sequences. SIAM Journal on Applied Mathematics, 44(6): 1210–1217, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stickler:1984:ATM

[2240] D. C. Stickler. Application of the trace method for inverse scattering to a medium which supports N types of waves. SIAM Journal on Applied Mathematics, 44(6):1218–1225, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jensen:1984:OEM

[2241] D. R. Jensen. Ordering ellipsoidal measures: Scale and peakedness orderings. SIAM Journal on Applied Mathematics, 44(6):1226–1231, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Landau:1984:DCM

[2242] H. J. Landau, B. F. Logan, L. A.

Shepp, and N. Bauman. Diffusion, cell mobility, and bandlimited functions. SIAM Journal on Applied Mathematics, 44(6):1232–1245, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Angell:1984:GPO

[2243] Thomas S. Angell and Ralph E. Kleinman. A Galerkin procedure for optimization in radiation problems. SIAM Journal on Applied Mathematics, 44 (6):1246–1257, December 1984. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bhattacharya:1984:APT

[2244] R. N. Bhattacharya and Vijay K.

Gupta. Acknowledgment of priority: “On the Taylor–Aris theory of solute transport in a capillary” [SIAM J. Appl. Math. 44 (1984), no. 1, 33–39, MR 85d:82114]. SIAM Journal on Applied Mathematics, 44(6):1258, December 1984. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eckhaus:1985:SPF

[2245] Wiktor Eckhaus, Aart Van Harten, and Zbigniew Peradzynski. A singularly perturbed free boundary problem describing a laser sustained plasma. SIAM Journal on Applied Mathematics, 45(1):1–31, February 1985. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Prufer:1985:TMM

[2246] Michael Prüfer. Turbulence in multistep methods for initial value problems. SIAM Journal on Applied Mathematics, 45(1):32–69, February 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McTaggart:1985:NEB

[2247] Carol L. McTaggart and K. A. Lindsay. Nonclassical effects in the Bénard problem. SIAM Journal on Applied Mathematics, 45(1):70–92, February 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Margolis:1985:FPC

[2248] Stephen B. Margolis and Bernard J. Matkowsky. Flame propagation in channels: Secondary bifurcation to quasi-periodic pulsations. SIAM Journal on Applied Mathematics, 45(1): 93–129, February 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jackson:1985:SIT

[2249] T. L. Jackson and A. K. Kapila. Shockinduced thermal runaway. SIAM Journal on Applied Mathematics, 45(1): 130–137, February 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Butler:1985:MMCa

[2250] G. J. Butler and G. S. K. Wolkowicz. A mathematical model of the chemostat with a general class of functions describing nutrient uptake. SIAM Journal on Applied Mathematics, 45 (1):138–151, February 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1985:RTD

[2251] J. A. Morrison. Response-time distribution for a processor-sharing system. SIAM Journal on Applied Mathematics, 45(1):152–167, February 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Flatto:1985:ETP

[2252] L. Flatto and S. Hahn. Erratum:

“Two parallel queues created by arrivals with two demands. I” [SIAM J. Appl. Math. 44 (1984), no. 5, 1041–1053]. SIAM Journal on Applied Mathematics, 45(1):168, February 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [2227].

Grotberg:1985:SBQ

[2253] James B. Grotberg and Edward L.

Reiss. Secondary bifurcation of quasiperiodic solutions can lead to period multiplication. SIAM Journal on Applied Mathematics, 45(2):169–174, April 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lange:1985:BAN

[2254] Charles G. Lange and Gregory A.

Kriegsmann. Bifurcation analysis of nonlinear turning point problems. SIAM Journal on Applied Mathematics, 45(2):175–199, April 1985. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Weinstein:1985:HEL

[2255] Michael I. Weinstein and Joseph B.

Keller. Hill’s equation with a large potential. SIAM Journal on Applied Mathematics, 45(2):200–214, April 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lim:1985:CQS

[2256] Chee-Seng Lim. Coupled quasiisotropic systems. SIAM Journal on Applied Mathematics, 45(2):215–255, April 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Swick:1985:SRM

[2257] K. E. Swick. Some reducible models of age dependent dynamics. SIAM Journal on Applied Mathematics, 45 (2):256–267, April 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1985:UOS

[2258] Michael A. Cohen and Daniel S. Levine. Unbounded oscillatory solutions for a system of interacting populations. SIAM Journal on Applied Mathematics, 45(2):268–279, April 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Richards:1985:AID

[2259] Donald St. P. Richards. Applications of invariant differential operators to multivariate distribution theory. SIAM Journal on Applied Mathematics, 45 (2):280–288, April 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Christov:1985:SFE

[2260] C. I. Christov and K. Z. Markov. Stochastic functional expansion for random media with perfectly disordered constitution. SIAM Journal on Applied Mathematics, 45(2):289–311, April 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bruckstein:1985:DMI

[2261] Alfred M. Bruckstein, Bernard C. Levy, and Thomas Kailath. Differential methods in inverse scattering. SIAM Journal on Applied Mathematics, 45 (2):312–335, April 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fawcett:1985:IDS

[2262] John A. Fawcett. Inversion of Ndimensional spherical averages. SIAM Journal on Applied Mathematics, 45 (2):336–341, April 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sethna:1985:MAI

[2263] P. R. Sethna, K. R. Meyer, and A. K.

Bajaj. On the method of averaging, integral manifolds and systems with symmetry. SIAM Journal on Applied Mathematics, 45(3):343–359, June 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lenhoff:1985:CEC

[2264] A. M. Lenhoff. Computation of the eigenvalues of a class of non-self-adjoint operators. SIAM Journal on Applied Mathematics, 45(3):360–368, June 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carasso:1985:PWD

[2265] Alfred S. Carasso and Nelson N. Hsu.

Probe waveforms and deconvolution in the experimental determination of elastic Green’s functions. SIAM Journal on Applied Mathematics, 45(3):369–382, June 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Katz:1985:RES

[2266] A. Katz and Z. Schuss. Reliability of elastic structures driven by random loads. SIAM Journal on Applied Mathematics, 45(3):383–402, June 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [2309].

Rubinstein:1985:APF

[2267] I. Rubinstein. Asymptotics of propagating front formation in diffusion kinetics. SIAM Journal on Applied Mathematics, 45(3):403–419, June 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Joulin:1985:FCB

[2268] G. Joulin, A. Liñán, G. S. S. Ludford, N. Peters, and C. Schmidt-Lainé. Flames with chain-branching/ chain-breaking kinetics. SIAM Journal on Applied Mathematics, 45(3):420– 434, June 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Butler:1985:MMCb

[2269] G. J. Butler, S. B. Hsu, and Paul Waltman. A mathematical model of the chemostat with periodic washout rate. SIAM Journal on Applied Mathematics, 45(3):435–449, June 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

LeVeque:1985:STD

[2270] Randall J. LeVeque, Charles S. Peskin, and Peter D. Lax. Solution of a twodimensional cochlea model using transform techniques. SIAM Journal on Applied Mathematics, 45(3):450–464, June 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brown:1985:DUS

[2271] Peter N. Brown. Decay to uniform states in predator-prey chains. SIAM Journal on Applied Mathematics, 45 (3):465–478, June 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goovaerts:1985:ATP

[2272] M. J. Goovaerts and F. Broeckx. Analytical treatment of a periodic δfunction potential in the path-integral formalism. SIAM Journal on Applied Mathematics, 45(3):479–490, June 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fawcett:1985:TDR

[2273] John Fawcett and H. B. Keller. Threedimensional ray tracing and geophysical inversion in layered media. SIAM Journal on Applied Mathematics, 45 (3):491–501, June 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pelaez:1985:SSF

[2274] J. Peláez and A. Li nán. Structure and stability of flames with two sequential reactions. SIAM Journal on Applied Mathematics, 45(4):503–522, August 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hannsgen:1985:SSS

[2275] Kenneth B. Hannsgen, John J. Tyson, and Layne T. Watson. Steady-state size distributions in probabilistic models of the cell division cycle. SIAM Journal on Applied Mathematics, 45 (4):523–540, August 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Phillipson:1985:ASM

[2276] Paul E. Phillipson, Peter Schuster, and Roger G. Johnston. An analytic study of the May–Leonard equations. SIAM Journal on Applied Mathematics, 45 (4):541–554, August 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Coste:1985:DRM

[2277] Jean Coste. Dynamical regression in many species ecosystems: The case of many predators competing for several preys. SIAM Journal on Applied Mathematics, 45(4):555–564, August 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jerome:1985:CSM

[2278] Joseph W. Jerome. Consistency of semiconductor modeling: An existence/stability analysis for the stationary van Roosbroeck system. SIAM Journal on Applied Mathematics, 45 (4):565–590, August 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1985:OVT

[2279] A. H. Zemanian. Operator-valued transmission lines in the analysis of two-dimensional anomalies imbedded in a horizontally layered earth under transient polarized electromagnetic excitation. SIAM Journal on Applied Mathematics, 45(4):591–620, August 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Percus:1985:PBS

[2280] Ora E. Percus and Jerome K. Percus. Probability bounds on the sum of independent nonidentically distributed binomial random variables. SIAM Journal on Applied Mathematics, 45 (4):621–640, August 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Delsarte:1985:DPS

[2281] Ph. Delsarte, A. J. E. M. Janssen, and L. B. Vries. Discrete prolate spheroidal wave functions and interpolation. SIAM Journal on Applied Mathematics, 45(4):641–650, August 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sanz:1985:MCP

[2282] Jorge L. C. Sanz. Mathematical considerations for the problem of Fourier transform phase retrieval from magnitude. SIAM Journal on Applied Mathematics, 45(4):651–664, August 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Finch:1985:CBR

[2283] David V. Finch. Cone beam reconstruction with sources on a curve. SIAM Journal on Applied Mathematics, 45 (4):665–673, August 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Udwadia:1985:SUR

[2284] F. E. Udwadia. Some uniqueness results related to soil and building structural identification. SIAM Journal on Applied Mathematics, 45(4):674–685, August 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lange:1985:SPAa

[2285] Charles G. Lange and Robert M. Miura. Singular perturbation analysis of boundary value problems for differential-difference equations. II. Rapid oscillations and resonances. SIAM Journal on Applied Mathematics, 45(5):687–707, October 1985. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lange:1985:SPAb

[2286] Charles G. Lange and Robert M. Miura. Singular perturbation analysis of boundary value problems for differential-difference equations. III. Turning point problems. SIAM Journal on Applied Mathematics, 45(5): 708–734, October 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hunter:1985:RMS

[2287] John K. Hunter. A ray method for slowly modulated nonlinear waves. SIAM Journal on Applied Mathematics, 45(5):735–749, October 1985. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maginu:1985:GCA

[2288] Kenjiro Maginu. Geometrical characteristics associated with stability and bifurcations of periodic travelling waves in reaction-diffusion systems. SIAM Journal on Applied Mathematics, 45 (5):750–774, October 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Herron:1985:ESG

[2289] Isom H. Herron. Exchange of stabilities for Görtler flow. SIAM Journal on Applied Mathematics, 45(5):775–779, October 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McLaughlin:1985:CMR

[2290] D. W. McLaughlin, G. C. Papanicolaou, and O. R. Pironneau. Convection of microstructure and related problems. SIAM Journal on Applied Mathematics, 45(5):780–797, October 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cannon:1985:FBV

[2291] John R. Cannon, James C. Cavendish, and Antonio Fasano. A free boundaryvalue problem related to the combustion of a solid. SIAM Journal on Applied Mathematics, 45(5):798–809, October 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sankoff:1985:SSR

[2292] David Sankoff. Simultaneous solution of the RNA folding, alignment and protosequence problems. SIAM Journal on Applied Mathematics, 45 (5):810–825, October 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hagan:1985:ESP

[2293] P. S. Hagan, R. S. Polizzotti, and G. Luckman. The effects of surface poisons on the oxidation of binary alloys. SIAM Journal on Applied Mathematics, 45(5):826–842, October 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schuss:1985:FFN

[2294] Z. Schuss, C. Tier, and B. J. Matkowsky. Frequency fluctuations in noisy oscillators. SIAM Journal on Applied Mathematics, 45(5):843–854, October 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Trietsch:1985:AEN

[2295] Dan Trietsch. Augmenting Euclidean networks — the Steiner case. SIAM Journal on Applied Mathematics, 45 (5):855–860, October 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Flatto:1985:TPQ

[2296] Leopold Flatto. Two parallel queues created by arrivals with two demands. II. SIAM Journal on Applied Mathematics, 45(5):861–878, October 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Roose:1985:DMC

[2297] D. Roose and V. Hlavacek. A direct method for the computation of Hopf bifurcation points. SIAM Journal on Applied Mathematics, 45(6): 879–894, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dangelmayr:1985:SBC

[2298] Gerhard Dangelmayr and Ian Stewart. Sequential bifurcations in continuous stirred tank chemical reactors coupled in series. SIAM Journal on Applied Mathematics, 45(6):895–918, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Haberman:1985:SVS

[2299] Richard Haberman and Darrell Allgaier. Slowly varying solitary wave tails: Focusing, cusped caustics, wave number shocks and birth of tails. SIAM Journal on Applied Mathematics, 45 (6):919–927, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pramanik:1985:TDG

[2300] A. K. Pramanik and S. R. Majumdar. Transient development of gravity waves on a sloping beach. SIAM Journal on Applied Mathematics, 45 (6):928–942, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kath:1985:AMB

[2301] W. L. Kath and J. D. Murray. Analysis of a model biological switch. SIAM Journal on Applied Mathematics, 45 (6):943–955, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hagan:1985:IOB

[2302] P. S. Hagan, R. S. Polizzotti, and G. Luckman. Internal oxidation of binary alloys. SIAM Journal on Applied Mathematics, 45(6):956–971, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kipnis:1985:MBI

[2303] C. Kipnis and C. M. Newman. The metastable behavior of infrequently observed, weakly random, onedimensional diffusion processes. SIAM Journal on Applied Mathematics, 45 (6):972–982, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Iyengar:1985:HLT

[2304] Satish Iyengar. Hitting lines with twodimensional Brownian motion. SIAM Journal on Applied Mathematics, 45 (6):983–989, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Spigler:1985:SMN

[2305] Renato Spigler. A stochastic model for nonlinear oscillators of Duffing type. SIAM Journal on Applied Mathematics, 45(6):990–1005, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:1985:ATL

[2306] C. Knessl, B. J. Matkowsky, Z. Schuss, and C. Tier. An asymptotic theory of large deviations for Markov jump processes. SIAM Journal on Applied Mathematics, 45(6):1006–1028, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carasso:1985:EBP

[2307] Alfred S. Carasso and Nelson N. Hsu. L^{∞} error bounds in partial deconvolution of the inverse Gaussian pulse.

SIAM Journal on Applied Mathematics, 45(6):1029–1038, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Colton:1985:NMS

[2308] David Colton and Peter Monk. A novel method for solving the inverse scattering problem for time-harmonic acoustic waves in the resonance region. SIAM Journal on Applied Mathematics, 45(6):1039–1053, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Katz:1985:ERE

[2309] A. Katz and Z. Schuss. Erratum: “Reliability of Elastic Structures Driven by Random Loads” [SIAM J. Appl. Math. 45 (1985), no. 3, 383–402, MR 86j:73076]. SIAM Journal on Applied Mathematics, 45(6):1054, December 1985. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [2266].

Erneux:1986:IBS

[2310] Thomas Erneux and Paul Mandel.

Imperfect bifurcation with a slowlyvarying control parameter. SIAM Journal on Applied Mathematics, 46 (1):1–15, February 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Haberman:1986:NIF

[2311] Richard Haberman. Note on the initial formation of shocks. SIAM Journal on Applied Mathematics, 46 (1):16–19, February 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Howison:1986:CDH

[2312] S. D. Howison. Cusp development in Hele–Shaw flow with a free surface. SIAM Journal on Applied Mathematics, 46(1):20–26, February 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cook:1986:SLT

[2313] L. Pamela Cook and F. J. Zeigler. The stabilization law for transonic flow. SIAM Journal on Applied Mathematics, 46(1):27–48, February 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mauri:1986:AWP

[2314] R. Mauri and S. Haber. Applications of Wiener’s path integral for the diffusion of Brownian particles in shear flows. SIAM Journal on Applied Mathematics, 46(1):49–55, February 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Storti:1986:DTS

[2315] D. W. Storti and R. H. Rand. Dynamics of two strongly coupled relaxation oscillators. SIAM Journal on Applied Mathematics, 46(1):56–67, February 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kopell:1986:BCR

[2316] Nancy Kopell and David Ruelle. Bounds on complexity in reactiondiffusion systems. SIAM Journal on Applied Mathematics, 46(1):68–80, February 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Leung:1986:SCS

[2317] Anthony W. Leung and Boualem Bendjilali. N species competition for a spatially heterogeneous prey with Neumann boundary conditions: Steady states and stability. SIAM Journal on Applied Mathematics, 46(1):81–98, February 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Idemen:1986:HFS

[2318] M. İdemen and T. Şengör. Highfrequency surface currents induced on a spherical cap. SIAM Journal on Applied Mathematics, 46(1):99–117, February 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Darling:1986:EVD

[2319] R. W. R. Darling and Michael S. Waterman. Extreme value distribution for the largest cube in a random lattice. SIAM Journal on Applied Mathematics, 46(1):118–132, February 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rawlins:1986:MWH

[2320] A. D. Rawlins. Matrix Wiener–Hopf– Hilbert factorisation. SIAM Journal on Applied Mathematics, 46(1): 133–139, February 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1986:AAW

[2321] J. A. Morrison. Asymptotic analysis of the waiting-time distribution for a large closed processor-sharing system. SIAM Journal on Applied Mathematics, 46 (1):140–170, February 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Olmstead:1986:BM

[2322] W. E. Olmstead, S. H. Davis, S. Rosenblat, and W. L. Kath. Bifurcation with memory. SIAM Journal on Applied Mathematics, 46(2):171–188, April 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tu:1986:MTV

[2323] Son Tu and Edward L. Reiss. Minimum transition values and the dynamics of subcritical bifurcation. SIAM Journal on Applied Mathematics, 46 (2):189–198, April 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Alexander:1986:PPH

[2324] J. C. Alexander. Patterns at primary Hopf bifurcations of a plexus of identical oscillators. SIAM Journal on Applied Mathematics, 46(2):199–221, April 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Freidlin:1986:GOA

[2325] M. I. Freĭdlin. Geometric optics approach to reaction-diffusion equations. SIAM Journal on Applied Mathematics, 46(2):222–232, April 1986. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ermentrout:1986:PBE

[2326] G. B. Ermentrout and N. Kopell. Parabolic bursting in an excitable system coupled with a slow oscillation.

SIAM Journal on Applied Mathematics, 46(2):233–253, April 1986. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tao:1986:MSM

[2327] L. N. Tao. A method for solving moving boundary problems. SIAM Journal on Applied Mathematics, 46 (2):254–264, April 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dygas:1986:SPA

[2328] M. M. Dygas, B. J. Matkowsky, and Z. Schuss. A singular perturbation approach to non-Markovian escape rate problems. SIAM Journal on Applied Mathematics, 46(2):265–298, April 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Butzer:1986:SSS

[2329] P. L. Butzer, W. Engels, S. Ries, and R. L. Stens. The Shannon sampling series and the reconstruction of signals in terms of linear, quadratic and cubic splines. SIAM Journal on Applied Mathematics, 46(2):299–323, April 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Strakhov:1986:UIL

[2330] V. N. Strakhov and M. A. Brodsky.

On the uniqueness of the inverse logarithmic potential problem. SIAM Journal on Applied Mathematics, 46 (2):324–344, April 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brodsky:1986:UIP

[2331] M. A. Brodsky. On the uniqueness of the inverse potential problem for homogeneous polyhedrons. SIAM Journal on Applied Mathematics, 46(2):345–350, April 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reid:1986:AAI

[2332] W. H. Reid and S. J. Skates. On the asymptotic approximation of integrals. SIAM Journal on Applied Mathematics, 46(3):351–358, June 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ermentrout:1986:PRD

[2333] G. B. Ermentrout and W. C. Troy. Phaselocking in a reaction-diffusion system with a linear frequency gradient. SIAM Journal on Applied Mathematics, 46(3):359–367, June 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1986:ABC

[2334] Hal L. Smith. On the asymptotic behavior of a class of deterministic models of cooperating species. SIAM Journal on Applied Mathematics, 46 (3):368–375, June 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brown:1986:DUS

[2335] Peter N. Brown. Decay to uniform states in food webs. SIAM Journal on Applied Mathematics, 46(3):376–392, June 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Letessier:1986:DBD

[2336] J. Letessier and G. Valent. Dual birth and death processes and orthogonal polynomials. SIAM Journal on Applied Mathematics, 46(3):393–405, June 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dudewicz:1986:ETS

[2337] Edward J. Dudewicz and Ishwari D.

Dhariyal. Estimation theory of selective reporting: the best-1-of-k report. SIAM Journal on Applied Mathematics, 46(3):406–416, June 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goodrich:1986:SE

[2338] Robert K. Goodrich and Allan Steinhardt. L_{2} spectral estimation. SIAM Journal on Applied Mathematics, 46 (3):417–426, June 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arnold:1986:AAL

[2339] Ludwig Arnold, George Papanicolaou, and Volker Wihstutz. Asymptotic analysis of the Lyapunov exponent and rotation number of the random oscillator and applications. SIAM Journal on Applied Mathematics, 46(3):427–450, June 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pinsky:1986:IHO

[2340] Mark A. Pinsky. Instability of the harmonic oscillator with small noise. SIAM Journal on Applied Mathematics, 46(3):451–463, June 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kohler:1986:ROD

[2341] Werner E. Kohler. Reflection from a one-dimensional, totally refracting random multilayer. SIAM Journal on Applied Mathematics, 46(3):464–482, June 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:1986:AAS

[2342] C. Knessl, B. J. Matkowsky, Z. Schuss, and C. Tier. Asymptotic analysis of a state-dependent M/G/1 queueing system. SIAM Journal on Applied Mathematics, 46(3):483–505, June 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Colton:1986:NMS

[2343] David Colton and Peter Monk. A novel method for solving the inverse scattering problem for time-harmonic acoustic waves in the resonance region. II. SIAM Journal on Applied Mathematics, 46(3):506–523, June 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McClendon:1986:VMA

[2344] Marvin McClendon and Herschel Rabitz. Variational methods for approximating solutions of ∇^{2}u(x) = f(x) + k(x)u(x) and generalizations. SIAM Journal on Applied Mathematics, 46 (4):525–544, August 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miles:1986:NTP

[2345] John Miles. A note on the threeparticle lattice and the Hénon-Heiles problem. SIAM Journal on Applied Mathematics, 46(4):545–551, August 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Freedman:1986:SCS

[2346] H. I. Freedman and V. Sree Hari Rao.

Stability criteria for a system involving two time delays. SIAM Journal on Applied Mathematics, 46(4):552–560, August 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Deshaies:1986:RTP

[2347] B. Deshaies and G. Joulin. Radiative transfer as a propagation mechanism for rich flames of reactive suspensions. SIAM Journal on Applied Mathematics, 46(4):561–581, August 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stewart:1986:SLP

[2348] D. S. Stewart and J. Buckmaster. The stability of Liñán’s “premixed flame regime” revisited. SIAM Journal on Applied Mathematics, 46(4):582–587, August 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Auchmuty:1986:SSP

[2349] G. Auchmuty, J. Chadam, E. Merino, P. Ortoleva, and E. Ripley. The structure and stability of propagating redox fronts. SIAM Journal on Applied Mathematics, 46(4):588–604, August 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hagan:1986:UCF

[2350] Patrick S. Hagan. Uptake curves for Fickian diffusion. SIAM Journal on Applied Mathematics, 46(4):605–613, August 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Frenzen:1986:CKJ

[2351] C. L. Frenzen and J. D. Murray. A cell kinetics justification for Gompertz’ equation. SIAM Journal on Applied Mathematics, 46(4):614–629, August 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Conway:1986:GAS

[2352] E. D. Conway and J. A. Smoller. Global analysis of a system of predator-prey equations. SIAM Journal on Applied Mathematics, 46(4):630–642, August 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jacobs:1986:FPT

[2353] P. A. Jacobs. First passage times for combinations of random loads. SIAM Journal on Applied Mathematics, 46 (4):643–656, August 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:1986:PSD

[2354] C. Knessl, B. J. Matkowsky, Z. Schuss, and C. Tier. On the performance of state-dependent single server queues. SIAM Journal on Applied Mathematics, 46(4):657–697, August 1986. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Devroye:1986:AMG

[2355] Luc Devroye. An automatic method for generating random variates with a given characteristic function. SIAM Journal on Applied Mathematics, 46 (4):698–719, August 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baer:1986:SHB

[2356] S. M. Baer and T. Erneux. Singular Hopf bifurcation to relaxation oscillations. SIAM Journal on Applied Mathematics, 46(5):721–739, October 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dellwo:1986:CTI

[2357] David R. Dellwo. A constructive theory of isolas supported by parabolic cusps, centers and bifurcation points. SIAM Journal on Applied Mathematics, 46 (5):740–764, October 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Overman:1986:SSS

[2358] Edward A. Overman, II. Steady-state solutions of the Euler equations in two dimensions. II. Local analysis of limiting V -states. SIAM Journal on Applied Mathematics, 46(5):765–800, October 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Booty:1986:IPS

[2359] Michael R. Booty, Stephen B. Margolis, and Bernard J. Matkowsky. Interaction of pulsating and spinning waves in condensed phase combustion. SIAM Journal on Applied Mathematics, 46 (5):801–843, October 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bell:1986:TCT

[2360] Jonathan Bell and Chris Cosner. Threshold conditions for two diffusion models suggested by nerve impulse conduction. SIAM Journal on Applied Mathematics, 46(5):844–855, October 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1986:CSM

[2361] Hal L. Smith. Competing subcommunities of mutualists and a generalized Kamke theorem. SIAM Journal on Applied Mathematics, 46(5):856–874, October 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Li:1986:ASS

[2362] C. W. Li and G. L. Blankenship. Almost sure stability of linear stochastic systems with Poisson process coefficients. SIAM Journal on Applied Mathematics, 46(5):875–911, October 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hordijk:1986:OPC

[2363] Arie Hordijk and Frank A. van der Duyn Schouten. On the optimality of (s, S)-policies in continuous review inventory models. SIAM Journal on Applied Mathematics, 46(5):912–929, October 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mugler:1986:DMP

[2364] Dale H. Mugler and W. Splettstösser. Difference methods for the prediction of band-limited signals. SIAM Journal on Applied Mathematics, 46 (5):930–941, October 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wazwaz:1986:MUA

[2365] Abdul-Majid Wazwaz and Floyd B.

Hanson. Matched uniform approximations for a singular boundary point and an interior turning point. SIAM Journal on Applied Mathematics, 46 (6):943–961, December 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wazwaz:1986:SBR

[2366] Abdul-Majid Wazwaz and Floyd B.

Hanson. Singular boundary resonance with turning point resonance. SIAM Journal on Applied Mathematics, 46 (6):962–977, December 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hunter:1986:ASO

[2367] C. Hunter and Sang Myung Lee. The analytic structure of Oseen flow past a sphere as a function of Reynolds number. SIAM Journal on Applied Mathematics, 46(6):978–999, December 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bell:1986:CLM

[2368] John B. Bell, John A. Trangenstein, and Gregory R. Shubin. Conservation laws of mixed type describing three-phase flow in porous media. SIAM Journal on Applied Mathematics, 46(6):1000–1017, December 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rubinstein:1986:DCP

[2369] Jacob Rubinstein and Roberto Mauri. Dispersion and convection in periodic porous media. SIAM Journal on Applied Mathematics, 46(6):1018–1023, December 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rubinstein:1986:CCE

[2370] I. Rubinstein. Counterion condensation as an exact limiting property of solutions of the Poisson–Boltzmann equation. SIAM Journal on Applied Mathematics, 46(6):1024–1038, December 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keener:1986:GTS

[2371] James P. Keener. A geometrical theory for spiral waves in excitable media. SIAM Journal on Applied Mathematics, 46(6):1039–1056, December 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dunbar:1986:TWD

[2372] Steven R. Dunbar. Traveling waves in diffusive predator-prey equations: Periodic orbits and point-to-periodic heteroclinic orbits. SIAM Journal on Applied Mathematics, 46(6):1057–1078, December 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Feroe:1986:ETW

[2373] John A. Feroe. Existence of traveling wave trains in nerve axon equations. SIAM Journal on Applied Mathematics, 46(6):1079–1097, December 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Picard:1986:NFO

[2374] Jean Picard. Nonlinear filtering of onedimensional diffusions in the case of a high signal-to-noise ratio. SIAM Journal on Applied Mathematics, 46(6): 1098–1125, December 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemanian:1986:DCA

[2375] A. H. Zemanian and T. S. Zemanian. Domain contractions around three-dimensional anomalies in spherical finite-difference computations of Poisson’s equation. SIAM Journal on Applied Mathematics, 46(6):1126–1149, December 1986. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Angell:1987:SPVa

[2376] J. S. Angell and W. E. Olmstead.

Singularly perturbed Volterra integral equations. SIAM Journal on Applied Mathematics, 47(1):1–14, February 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Namachchivaya:1987:PPH

[2377] N. Sri Namachchivaya and S. T. Ariaratnam. Periodically perturbed Hopf bifurcation. SIAM Journal on Applied Mathematics, 47(1):15–39, February 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bridges:1987:SBT

[2378] Thomas J. Bridges. On the secondary bifurcation of three-dimensional standing waves. SIAM Journal on Applied Mathematics, 47(1):40–59, February 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Byatt-Smith:1987:PSD

[2379] J. G. Byatt-Smith. 2π periodic solutions of Duffing’s equation with negative stiffness. SIAM Journal on Applied Mathematics, 47(1):60–91, February 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sanz-Serna:1987:SNN

[2380] J. M. Sanz-Serna and F. Vadillo. Studies in numerical nonlinear instability. III. Augmented Hamiltonian systems. SIAM Journal on Applied Mathematics, 47(1):92–108, February 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gautesen:1987:SCP

[2381] A. K. Gautesen and J. Dundurs. On the solution to a Cauchy principal value integral equation which arises in fracture mechanics. SIAM Journal on Applied Mathematics, 47(1):109–116, February 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Please:1987:MMC

[2382] C. P. Please, A. A. Wheeler, and P. Wilmott. A mathematical model of cliff blasting. SIAM Journal on Applied Mathematics, 47(1):117–127, February 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arino:1987:AAC

[2383] Ovide Arino and Marek Kimmel. Asymptotic analysis of a cell cycle model based on unequal division.

SIAM Journal on Applied Mathematics, 47(1):128–145, February 1987. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burridge:1987:SPR

[2384] R. Burridge, G. Papanicolaou, and B. White. Statistics for pulse reflection from a randomly layered medium. SIAM Journal on Applied Mathematics, 47(1):146–168, February 1987. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kushner:1987:AGB

[2385] H. J. Kushner. Asymptotic global behavior for stochastic approximation and diffusions with slowly decreasing noise effects: global minimization via Monte Carlo. SIAM Journal on Applied Mathematics, 47(1):169–185, February 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fox:1987:ETA

[2386] Bennett L. Fox and Peter W.

Glynn. Estimating time averages via randomly-spaced observations. SIAM Journal on Applied Mathematics, 47 (1):186–200, February 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Friedman:1987:DME

[2387] Avner Friedman. Detection of mines by electric measurements. SIAM Journal on Applied Mathematics, 47(1): 201–212, February 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pittel:1987:SR

[2388] Boris Pittel. On spreading a rumor. SIAM Journal on Applied Mathematics, 47(1):213–223, February 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Skinner:1987:USB

[2389] L. A. Skinner. Uniform solution of boundary layer problems exhibiting resonance. SIAM Journal on Applied Mathematics, 47(2):225–231, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Salam:1987:MTH

[2390] Fathi M. A. Salam. The Mel′nikov technique for highly dissipative systems. SIAM Journal on Applied Mathematics, 47(2):232–243, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chan:1987:SSS

[2391] Whei-Ching C. Chan. Stability of subharmonic solutions. SIAM Journal on Applied Mathematics, 47(2):244–253, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

LeVeque:1987:INE

[2392] Randall J. LeVeque. On the interaction of nearly equal solitons in the KdV equation. SIAM Journal on Applied Mathematics, 47(2):254–262, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Olver:1987:GIS

[2393] Peter J. Olver and Philip Rosenau.

Group-invariant solutions of differential equations. SIAM Journal on Applied Mathematics, 47(2):263–278, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kwok:1987:SAT

[2394] Yue-Kuen Kwok and Lawrence Sirovich.

On some aspects of the transonic controversy. SIAM Journal on Applied Mathematics, 47(2):279–295, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Clavin:1987:MRP

[2395] P. Clavin, P. Fife, and B. Nicolaenko.

Multiplicity and related phenomena in competing reaction flames. SIAM Journal on Applied Mathematics, 47 (2):296–331, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lasseigne:1987:ICS

[2396] D. Glenn Lasseigne and W. E. Olmstead. Ignition of a combustible solid with reactant consumption. SIAM Journal on Applied Mathematics, 47 (2):332–342, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dunkl:1987:OPH

[2397] Charles F. Dunkl. Orthogonal polynomials on the hexagon. SIAM Journal on Applied Mathematics, 47(2):343– 351, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chang:1987:ERB

[2398] Chung Ju Chang and Jin Fu Chang. The effect of routing and buffer sharing on the behavior of a finite queue with batch Poisson inputs and synchronous servers. SIAM Journal on Applied Mathematics, 47(2):352–366, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:1987:TRP

[2399] C. Knessl, B. J. Matkowsky, Z. Schuss, and C. Tier. The two repairmen problem: A finite source M/G/2 queue. SIAM Journal on Applied Mathematics, 47(2):367–397, April 1987. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1987:MGG

[2400] J. A. Morrison. The minimum of gaps generated by random packing of unit intervals into a large interval. SIAM Journal on Applied Mathematics, 47 (2):398–410, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Janardan:1987:WLD

[2401] K. G. Janardan. Weighted Lagrange distributions and their characterizations. SIAM Journal on Applied Mathematics, 47(2):411–415, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Springer:1987:EFI

[2402] Melvin D. Springer. Evaluation of the H-function inversion integral for real variables using Jordan’s lemma and residues. SIAM Journal on Applied Mathematics, 47(2):416–424, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dougherty:1987:EBM

[2403] Edward R. Dougherty and Charles R.

Giardina. Error bounds for morphologically derived measurements. SIAM Journal on Applied Mathematics, 47 (2):425–440, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hall:1987:PMB

[2404] Peter Hall. On the processing of a motion-blurred image. SIAM Journal on Applied Mathematics, 47(2):441– 453, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [2539].

Kogelman:1987:AAT

[2405] Stanley Kogelman and Joseph B.

Keller. Addendum: “Asymptotic Theory of Nonlinear Wave Propagation” [SIAM J. Appl. Math. 24 (1973), no. 3, 352–361, MR 47 #5412]. SIAM Journal on Applied Mathematics, 47(2):454, April 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gardini:1987:BTC

[2406] L. Gardini, R. Lupini, C. Mammana, and M. G. Messia. Bifurcations and transitions to chaos in the three-dimensional Lotka–Volterra map. SIAM Journal on Applied Mathematics, 47(3):455–482, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hunter:1987:OCP

[2407] C. Hunter. Oscillations in the coefficients of power series. SIAM Journal on Applied Mathematics, 47(3):483–497, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nusse:1987:APB

[2408] H. E. Nusse. Asymptotically periodic behaviour in the dynamics of chaotic mappings. SIAM Journal on Applied Mathematics, 47(3):498–515, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vogel:1987:SLD

[2409] Thomas I. Vogel. Stability of a liquid drop trapped between two parallel planes. SIAM Journal on Applied Mathematics, 47(3):516–525, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shih:1987:ESE

[2410] Ke-Gang Shih. On the existence of solutions of an equation arising in the theory of laminar flow in a uniformly porous channel with injection. SIAM Journal on Applied Mathematics, 47(3):526–533, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bell:1987:WLS

[2411] Jonathan Bell and Chris Cosner. Wavelike solutions to reaction-diffusion equations on a cylinder: Dependence on cylinder width. SIAM Journal on Applied Mathematics, 47(3):534–543, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kaper:1987:DNP

[2412] Hans G. Kaper, G. K. Leaf, B. J.

Matkowsky, and W. E. Olmstead. Dynamics of nonadiabatic premixed flames in a gravitational field. SIAM Journal on Applied Mathematics, 47 (3):544–555, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keener:1987:PFC

[2413] James P. Keener. Propagation and its failure in coupled systems of discrete excitable cells. SIAM Journal on Applied Mathematics, 47(3):556–572, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Castelfranco:1987:PSM

[2414] Ann M. Castelfranco and Harlan W.

Stech. Periodic solutions in a model of recurrent neural feedback. SIAM Journal on Applied Mathematics, 47 (3):573–588, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Weinstock:1987:LSA

[2415] Evelyn Weinstock and Chris Rorres.

Local stability of an age-structured population with density-dependent fertility and mortality. SIAM Journal on Applied Mathematics, 47(3):589–604, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barshinger:1987:EFA

[2416] Richard Barshinger and James Geer. The electrostatic field about a slender dielectric body. SIAM Journal on Applied Mathematics, 47(3):605–623, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ringhofer:1987:AAT

[2417] Christian Ringhofer. An asymptotic analysis of a transient p-n-junction model. SIAM Journal on Applied Mathematics, 47(3):624–642, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dupuis:1987:SSS

[2418] P. Dupuis and H. J. Kushner. Stochastic systems with small noise, analysis and simulation; A phase locked loop example. SIAM Journal on Applied Mathematics, 47(3):643–661, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chihara:1987:SCB

[2419] T. S. Chihara. On the spectra of certain birth and death processes. SIAM Journal on Applied Mathematics, 47 (3):662–669, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Butler:1987:SDA

[2420] James P. Butler and James A. Reeds. Stereology of dihedral angles I: First two moments. SIAM Journal on Applied Mathematics, 47(3):670–677, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reeds:1987:SDA

[2421] James A. Reeds and James P. Butler.

Stereology of dihedral angles. II. Distribution function. SIAM Journal on Applied Mathematics, 47(3):678–687, June 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Power:1987:SKI

[2422] Henry Power and Guillermo Miranda.

Second kind integral equation formulation of Stokes’ flows past a particle of arbitrary shape. SIAM Journal on Applied Mathematics, 47(4):689–698, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hussaini:1987:SSB

[2423] M. Y. Hussaini, W. D. Lakin, and A. Nachman. On similarity solutions of a boundary layer problem with an upstream moving wall. SIAM Journal on Applied Mathematics, 47 (4):699–709, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Birnir:1987:EBC

[2424] Björn Birnir. An example of blowup, for the complex KdV equation and existence beyond the blow-up. SIAM Journal on Applied Mathematics, 47 (4):710–725, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ganser:1987:NPW

[2425] Gary H. Ganser and Donald A. Drew. Nonlinear periodic waves in a twophase flow model. SIAM Journal on Applied Mathematics, 47(4):726–736, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dadfar:1987:PSS

[2426] Mohammad B. Dadfar and James F. Geer. Power series solution to a simple pendulum with oscillating support. SIAM Journal on Applied Mathematics, 47(4):737–750, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1987:PSM

[2427] Goong Chen, Matthew Coleman, and Harry H. West. Pointwise stabilization in the middle of the span for second order systems, nonuniform and uniform exponential decay of solutions. SIAM Journal on Applied Mathematics, 47 (4):751–780, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pelaez:1987:SPF

[2428] J. Peláez. Stability of premixed flames with two thin reaction layers. SIAM Journal on Applied Mathematics, 47 (4):781–799, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arditi:1987:OFN

[2429] Roger Arditi and Bernard Dacorogna. Optimal foraging in nonpatchy habitats. 2: Unbounded one-dimensional habitat. SIAM Journal on Applied Mathematics, 47(4):800–821, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Percus:1987:PHR

[2430] Ora E. Percus and Jerome K. Percus. Piecewise homogeneous random walk with a moving boundary. SIAM Journal on Applied Mathematics, 47 (4):822–830, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Markov:1987:AVWa

[2431] Konstantin Z. Markov. Application of Volterra–Wiener series for bounding the overall conductivity of heterogeneous media. I. General procedure. SIAM Journal on Applied Mathematics, 47(4):831–849, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Markov:1987:AVWb

[2432] Konstantin Z. Markov. Application of Volterra–Wiener series for bounding the overall conductivity of heterogeneous media. II. Suspensions of equisized spheres. SIAM Journal on Applied Mathematics, 47(4):850–870, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhang:1987:QCI

[2433] Jian Zhong Zhang and Patrick L.

Brockett. Quadratically constrained information theoretic analysis. SIAM Journal on Applied Mathematics, 47 (4):871–885, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cabana:1987:APT

[2434] E. M. Cabaña. Affine processes: A test of isotropy based on level sets. SIAM Journal on Applied Mathematics, 47 (4):886–891, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carasso:1987:IDP

[2435] Alfred S. Carasso. Infinitely divisible pulses, continuous deconvolution, and the characterization of linear time invariant systems. SIAM Journal on Applied Mathematics, 47(4):892–927, August 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gilsinn:1987:AAI

[2436] David E. Gilsinn. Asymptotic approximations of integral manifolds. SIAM Journal on Applied Mathematics, 47 (5):929–940, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Weinstein:1987:ABS

[2437] Michael I. Weinstein and Joseph B. Keller. Asymptotic behavior of stability regions for Hill’s equation. SIAM Journal on Applied Mathematics, 47 (5):941–958, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Newton:1987:SPP

[2438] Paul K. Newton and Joseph B. Keller. Stability of periodic plane waves. SIAM Journal on Applied Mathematics, 47 (5):959–964, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ehrenmark:1987:FFA

[2439] Ulf Torsten Ehrenmark. Far field asymptotics of the two-dimensional linearised sloping beach problem. SIAM Journal on Applied Mathematics, 47 (5):965–981, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vijayakumar:1987:GFB

[2440] Sinnathurai Vijayakumar and Donald E. Cormack. Green’s functions for the biharmonic equation: Bonded elastic media. SIAM Journal on Applied Mathematics, 47(5):982–997, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Joulin:1987:PDI

[2441] Guy Joulin. Preferential diffusion and the initiation of lean flames of light fuels. SIAM Journal on Applied Mathematics, 47(5):998–1016, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Majda:1987:NMF

[2442] Andrew Majda and Rodolfo R. Rosales.

Nonlinear mean field-high frequency wave interactions in the induction zone. SIAM Journal on Applied Mathematics, 47(5):1017–1039, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Merkin:1987:CGE

[2443] J. H. Merkin, D. J. Needham, and S. K. Scott. On the creation, growth and extinction of oscillatory solutions for a simple pooled chemical reaction scheme. SIAM Journal on Applied Mathematics, 47(5):1040–1060, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shapiro:1987:CDA

[2444] Michael Shapiro and Howard Brenner. Convection and diffusion accompanied by bulk and surface chemical reactions in time-periodic one-dimensional flows. SIAM Journal on Applied Mathematics, 47(5):1061–1075, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rubinstein:1987:MSS

[2445] I. Rubinstein. Multiple steady states in one-dimensional electrodiffusion with local electroneutrality. SIAM Journal on Applied Mathematics, 47(5): 1076–1093, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Leong:1987:SAM

[2446] Y. K. Leong. Stochastic approach to matrices. SIAM Journal on Applied Mathematics, 47(5):1094–1102, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Percus:1987:ODR

[2447] Ora E. Percus and Jerome K. Percus. One-dimensional random walk in alternating homogeneous domains. SIAM Journal on Applied Mathematics, 47 (5):1103–1111, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Seip:1987:IST

[2448] Kristian Seip. An irregular sampling theorem for functions bandlimited in a generalized sense. SIAM Journal on Applied Mathematics, 47 (5):1112–1116, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1987:CRT

[2449] John A. Morrison. Conditioned response-time distribution for a large closed processor-sharing system in very heavy usage. SIAM Journal on Applied Mathematics, 47(5):1117–1129, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rayner:1987:ESA

[2450] A. A. Rayner and R. M. Pringle. Errata: “Some Aspects of the Solution of Singular Normal Equations with the Use of Linear Restrictions” [SIAM J. Appl. Math. 31 (1976), no. 3, 449–460, MR 55 #6819]. SIAM Journal on Applied Mathematics, 47(5):1130, October 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1414].

Steinruck:1987:SPE

[2451] Herbert Steinrück. Singularly perturbed eigenvalue problems. SIAM Journal on Applied Mathematics, 47 (6):1131–1149, December 1987. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Angell:1987:SPVb

[2452] J. S. Angell and W. E. Olmstead.

Singularly perturbed Volterra integral equations. II. SIAM Journal on Applied Mathematics, 47(6):1150–1162, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Erneux:1987:NSC

[2453] Thomas Erneux, Edward L. Reiss, J. F. Magnan, and P. K. Jayakumar. Nonlinear stability control and λ-bifurcation. SIAM Journal on Applied Mathematics, 47(6):1163–1176, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stensby:1987:FLB

[2454] John Stensby. False lock and bifurcation in the phase locked loop. SIAM Journal on Applied Mathematics, 47(6):1177–1184, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maddocks:1987:RE

[2455] John H. Maddocks and Joseph B. Keller. Ropes in equilibrium. SIAM Journal on Applied Mathematics, 47 (6):1185–1200, December 1987. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sedov:1987:PDE

[2456] Alexander Sedov and Lester W. Schmerr, Jr. Pulse distortion and the elastodynamic Kirchhoff approximation for cracks. The direct and inverse problems. SIAM Journal on Applied Mathematics, 47(6):1201–1215, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Avellaneda:1987:OBM

[2457] M. Avellaneda. Optimal bounds and microgeometries for elastic two-phase composites. SIAM Journal on Applied Mathematics, 47(6):1216–1228, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hastings:1987:ETF

[2458] S. P. Hastings, C. Lu, and Y. H. Wan.

Existence of a traveling flame front in a model with no cold boundary difficulty. SIAM Journal on Applied Mathematics, 47(6):1229–1240, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Booty:1987:IPS

[2459] Michael R. Booty, Stephen B. Margolis, and Bernard J. Matkowsky. Interaction of pulsating and spinning waves in nonadiabatic flame propagation. SIAM Journal on Applied Mathematics, 47(6):1241–1286, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hagan:1987:ETP

[2460] P. S. Hagan, M. Herskowitz, and C. Pirkle. Equilibrium temperature profiles in highly sensitive tubular reactors. SIAM Journal on Applied Mathematics, 47(6):1287–1305, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brendel:1987:KAA

[2461] Volker Brendel and Alan S. Perelson. Kinetic analysis of adsorption processes. SIAM Journal on Applied Mathematics, 47(6):1306–1319, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ochs:1987:LAP

[2462] Robert L. Ochs, Jr. The limited aperture problem of inverse acoustic scattering: Dirichlet boundary conditions. SIAM Journal on Applied Mathematics, 47(6):1320–1341, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Janeczko:1987:ESL

[2463] Stanisław Janeczko and Adam Kowalczyk. Equivariant singularities of Lagrangian manifolds and uniaxial ferromagnet. SIAM Journal on Applied Mathematics, 47(6):1342–1360, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gill:1987:CSD

[2464] Jelena B. Gill and Ljiljana Petrović. Causality and stochastic dynamic systems. SIAM Journal on Applied Mathematics, 47(6):1361–1366, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baccelli:1987:AMR

[2465] François Baccelli and Guy Fayolle. Analysis of models reducible to a class of diffusion processes in the positive quarter plane. SIAM Journal on Applied Mathematics, 47(6):1367–1385, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

MacBain:1987:ITP

[2466] John A. MacBain. Inversion theory for a parameterized diffusion problem.

SIAM Journal on Applied Mathematics, 47(6):1386–1391, December 1987. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hunter:1988:TDN

[2467] John K. Hunter. Transverse diffraction of nonlinear waves and singular rays. SIAM Journal on Applied Mathematics, 48(1):1–37, February 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Alexander:1988:MV

[2468] James C. Alexander and J. H. Maddocks. On the maneuvering of vehicles. SIAM Journal on Applied Mathematics, 48(1):38–51, February 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wollkind:1988:NSA

[2469] David J. Wollkind and Shingmin Wang.

A nonlinear stability analysis of a model equation for liquid phase electroepitaxial growth of a dilute binary substance. SIAM Journal on Applied Mathematics, 48(1):52–78, February 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [2562].

Dassios:1988:SAC

[2470] George Dassios and Vassilis Kostopoulos. The scattering amplitudes and cross sections in the theory of thermoelasticity. SIAM Journal on Applied Mathematics, 48(1):79–98, February 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [2639].

Bamberger:1988:PWE

[2471] A. Bamberger, B. Engquist, L. Halpern, and P. Joly. Parabolic wave equation approximations in heterogeneous media. SIAM Journal on Applied Mathematics, 48(1):99–128, February 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bamberger:1988:HOP

[2472] A. Bamberger, B. Engquist, L. Halpern, and P. Joly. Higher order paraxial wave equation approximations in heterogeneous media. SIAM Journal on Applied Mathematics, 48(1):129–154, February 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Norbury:1988:TCWa

[2473] J. Norbury and A. M. Stuart. Travelling combustion waves in a porous medium. part I — existence. SIAM Journal on Applied Mathematics, 48 (1):155–169, February 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hainzl:1988:SHB

[2474] J. Hainzl. Stability and Hopf bifurcation in a predator-prey system with several parameters. SIAM Journal on Applied Mathematics, 48(1): 170–190, February 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Leveque:1988:STD

[2475] Randall J. Leveque, Charles S. Peskin, and Peter D. Lax. Solution of a twodimensional cochlea model with fluid viscosity. SIAM Journal on Applied Mathematics, 48(1):191–213, February 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Letessier:1988:SIR

[2476] J. Letessier and G. Valent. Some integral relations involving hypergeometric functions. SIAM Journal on Applied Mathematics, 48(1):214–221, February 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lo:1988:CFS

[2477] James Ting Ho Lo and Sze Kui Ng.

Characterizing Fourier series representation of probability distributions on compact Lie groups. SIAM Journal on Applied Mathematics, 48(1): 222–228, February 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yu:1988:BAD

[2478] Pei Yu and K. Huseyin. Bifurcations associated with a double zero and a pair of pure imaginary eigenvalues. SIAM Journal on Applied Mathematics, 48 (2):229–261, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wiggins:1988:DDC

[2479] Stephen Wiggins. On the detection and dynamical consequences of orbits homoclinic to hyperbolic periodic orbits and normally hyperbolic invariant tori in a class of ordinary differential equations. SIAM Journal on Applied Mathematics, 48(2):262–285, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hillion:1988:RST

[2480] P. Hillion. Remark on a system of two nonlinear differential equations. SIAM Journal on Applied Mathematics, 48 (2):286–293, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gautesen:1988:ASC

[2481] A. K. Gautesen. On the asymptotic solution to a class of linear integral equations. SIAM Journal on Applied Mathematics, 48(2):294–306, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kleinman:1988:SIE

[2482] R. E. Kleinman and P. A. Martin. On single integral equations for the transmission problem of acoustics. SIAM Journal on Applied Mathematics, 48 (2):307–325, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chee-Seng:1988:CPP

[2483] Lim Chee-Seng. A class of problems of pulsatory radiation from a travelling source distribution. SIAM Journal on Applied Mathematics, 48(2):326–352, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Daripa:1988:PFC

[2484] Prabir Daripa, James Glimm, Brent Lindquist, and Oliver McBryan. Polymer floods: A case study of nonlinear wave analysis and of instability control in tertiary oil recovery. SIAM Journal on Applied Mathematics, 48 (2):353–373, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Norbury:1988:TCWb

[2485] J. Norbury and A. M. Stuart. Travelling combustion waves in a porous medium. part II — stability. SIAM Journal on Applied Mathematics, 48 (2):374–392, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gautesen:1988:ECC

[2486] A. K. Gautesen. The effective conductivity of a composite material with a periodic rectangular geometry. SIAM Journal on Applied Mathematics, 48 (2):393–404, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

King:1988:IDD

[2487] J. R. King. Interacting dopant diffusions in crystalline silicon. SIAM Journal on Applied Mathematics, 48 (2):405–415, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kogan:1988:PMG

[2488] V. I. Kogan. Polynomial models of generalized bathtub curves and related moments of the order statistics. SIAM Journal on Applied Mathematics, 48 (2):416–424, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klosek-Dygas:1988:CND

[2489] M. M. Kłosek-Dygas, B. J. Matkowsky, and Z. Schuss. Colored noise in dynamical systems. SIAM Journal on Applied Mathematics, 48(2):425–441, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pardoux:1988:LER

[2490] É. Pardoux and V. Wihstutz. Lyapunov exponent and rotation number of two-dimensional linear stochastic systems with small diffusion. SIAM Journal on Applied Mathematics, 48 (2):442–457, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goodrich:1988:CWL

[2491] Robert K. Goodrich and Ranjit M. Passi. Characterization of the weights of least squares adaptive polynomials. SIAM Journal on Applied Mathematics, 48(2):458–464, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Malhame:1988:SPC

[2492] Roland Malhamé and Chee-Yee Chong. On the statistical properties of a cyclic diffusion process arising in the modeling of thermostat-controlled electric power system loads. SIAM Journal on Applied Mathematics, 48(2):465–480, April 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Joyal:1988:GHB

[2493] Pierre Joyal. Generalized Hopf bifurcation and its dual generalized homoclinic bifurcation. SIAM Journal on Applied Mathematics, 48(3):481–496, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baxley:1988:SNB

[2494] John V. Baxley. A singular nonlinear boundary value problem: Membrane response of a spherical cap. SIAM Journal on Applied Mathematics, 48(3):497–505, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Caginalp:1988:DLI

[2495] G. Caginalp and P. C. Fife. Dynamics of layered interfaces arising from phase boundaries. SIAM Journal on Applied Mathematics, 48(3):506–518, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Booty:1988:NWE

[2496] M. R. Booty, M. Matalon, and B. J.

Matkowsky. A nonlinear wave equation in nonadiabatic flame propagation. SIAM Journal on Applied Mathematics, 48(3):519–535, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meinkohn:1988:DCNa

[2497] Dirk Meinköhn. Disappearance of criticality in nonuniform systems I — an analytically closed representation of the phenomenon. SIAM Journal on Applied Mathematics, 48(3):536–548, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tucker:1988:NMP

[2498] Susan L. Tucker and Stuart O. Zimmerman. A nonlinear model of population dynamics containing an arbitrary number of continuous structure variables.

SIAM Journal on Applied Mathematics, 48(3):549–591, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wolkowicz:1988:BAP

[2499] G. S. K. Wolkowicz. Bifurcation analysis of a predator-prey system involving group defence. SIAM Journal on Applied Mathematics, 48(3):592–606, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Beretta:1988:GVM

[2500] E. Beretta and F. Solimano. A generalization of Volterra models with continuous time delay in population dynamics: Boundedness and global asymptotic stability. SIAM Journal on Applied Mathematics, 48(3):607–626, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Beretta:1988:GAS

[2501] E. Beretta and Y. Takeuchi. Global asymptotic stability of Lotka–Volterra diffusion models with continuous time delay. SIAM Journal on Applied Mathematics, 48(3):627–651, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kohler:1988:ROD

[2502] Werner E. Kohler. Reflection from a one-dimensional, totally refracting random multilayer II — internal field statistics. SIAM Journal on Applied Mathematics, 48(3):652–661, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rychlik:1988:RFC

[2503] Igor Rychlik. Rain-flow-cycle distribution for ergodic load processes. SIAM Journal on Applied Mathematics, 48(3):662–679, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Walter:1988:FCG

[2504] Gilbert G. Walter. A finite continuous Gegenbauer transform and its inverse. II. SIAM Journal on Applied Mathematics, 48(3):680–688, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bava:1988:FEI

[2505] G. P. Bava, G. Ghione, and I. Maio.

Fast exact inversion of the generalized Zakharov–Shabat problem for rational scattering data: Application to the synthesis of optical couplers. SIAM Journal on Applied Mathematics, 48(3):689–702, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Idemen:1988:SGI

[2506] Mithat İdemen and Ibrahim Akduman.

Some geometrical inverse problems connected with two-dimensional static fields. SIAM Journal on Applied Mathematics, 48(3):703–718, June 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

vanHorssen:1988:IBV

[2507] W. T. van Horssen and A. H. P. van der Burgh. On initial boundary value problems for weakly semilinear telegraph equations. Asymptotic theory and application. SIAM Journal on Applied Mathematics, 48(4):719–736, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bourland:1988:MPS

[2508] F. Jay Bourland and Richard Haberman. The modulated phase shift for strongly nonlinear, slowly varying, and weakly damped oscillators. SIAM Journal on Applied Mathematics, 48 (4):737–748, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Metzener:1988:BCS

[2509] P. Metzener and S. H. Davis. Bifurcations of convection states in toroidal shells. SIAM Journal on Applied Mathematics, 48(4):749–769, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goldberg:1988:SSN

[2510] Irwin S. Goldberg and Robert T. Folk. Solutions for steady and nonsteady entrance flow in a semi-infinite circular tube at very low Reynolds numbers. SIAM Journal on Applied Mathematics, 48(4):770–791, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meinkohn:1988:DCNb

[2511] Dirk Meinköhn. Disappearance of criticality in nonuniform systems. II. A general investigation by upper and lower bounds. SIAM Journal on Applied Mathematics, 48(4):792–807, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Magnan:1988:RTC

[2512] Jerry F. Magnan and Edward L.

Reiss. Rotating thermal convection: Neosteady and neoperiodic solutions. SIAM Journal on Applied Mathematics, 48(4):808–827, August 1988. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Margolis:1988:NMQ

[2513] Stephen B. Margolis and Bernard J.

Matkowsky. New modes of quasiperiodic combustion near a degenerate Hopf bifurcation point. SIAM Journal on Applied Mathematics, 48 (4):828–853, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Parra:1988:LNS

[2514] Ignacio E. Parra and José M. Vega.

Local nonlinear stability of the steady state in an isothermal catalyst. SIAM Journal on Applied Mathematics, 48 (4):854–881, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Busenberg:1988:EDS

[2515] S. N. Busenberg and J. M. Mahaffy.

The effects of dimension and size for a compartmental model of repression. SIAM Journal on Applied Mathematics, 48(4):882–903, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Belardinelli:1988:IPE

[2516] Enzo Belardinelli. On infinite product expansion of the arterial longitudinal impedance. SIAM Journal on Applied Mathematics, 48(4):904–916, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hagan:1988:SLI

[2517] P. S. Hagan and D. Brenner. Sorption limits with increasing diffusion coefficients. SIAM Journal on Applied Mathematics, 48(4):917–920, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kuiken:1988:NWS

[2518] H. K. Kuiken. A note on the wall singularity of a solid-liquid interface caused by a difference between the thermal conductivities of the solid and the liquid phases. SIAM Journal on Applied Mathematics, 48(4):921–924, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Beumee:1988:UFT

[2519] J. G. B. Beumee and H. Rabitz. On the use of filtering theory for the inversion of temporal chemical systems: Mathematical formulation. SIAM Journal on Applied Mathematics, 48(4):925–935, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kogan:1988:MDC

[2520] V. I. Kogan. On monotonicity of distributions of certain functions of order statistics. SIAM Journal on Applied Mathematics, 48(4):936–947, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gottlieb:1988:DDI

[2521] H. P. W. Gottlieb. Density distribution for isospectral circular membranes. SIAM Journal on Applied Mathematics, 48(4):948–951, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Levy:1988:VBB

[2522] Bernard C. Levy and Cengiz Esmersoy.

Variable background Born inversion by wavefield backpropagation. SIAM Journal on Applied Mathematics, 48 (4):952–972, August 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berger:1988:IVS

[2523] S. A. Berger. Initial-value stability analysis of a liquid jet. SIAM Journal on Applied Mathematics, 48 (5):973–991, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barta:1988:SBIa

[2524] E. Barta and N. Liron. Slender body interactions for low Reynolds numbers — part I: Body-wall interactions. SIAM Journal on Applied Mathematics, 48 (5):992–1008, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Isaacson:1988:RPNa

[2525] E. Isaacson, D. Marchesin, B. Plohr, and B. Temple. The Riemann problem near a hyperbolic singularity: the classification of solutions of quadratic Riemann problems. I. SIAM Journal on Applied Mathematics, 48(5): 1009–1032, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pitman:1988:SGF

[2526] E. Bruce Pitman. The stability of granular flow in converging hoppers. SIAM Journal on Applied Mathematics, 48 (5):1033–1053, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kaper:1988:DNE

[2527] Hans G. Kaper, G. K. Leaf, B. J. Matkowsky, and W. E. Olmstead. Dynamics of nearly extinguished nonadiabatic premixed flames in a gravitational field. SIAM Journal on Applied Mathematics, 48(5):1054–1063, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maini:1988:NAM

[2528] P. K. Maini and J. D. Murray. A nonlinear analysis of a mechanical model for biological pattern formation. SIAM Journal on Applied Mathematics, 48 (5):1064–1072, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carrillo:1988:MSA

[2529] Humberto Carrillo and David Lipman. The multiple sequence alignment problem in biology. SIAM Journal on Applied Mathematics, 48(5): 1073–1082, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hagan:1988:SAH

[2530] P. S. Hagan, M. Herskowitz, and C. Pirkle. A simple approach to highly sensitive tubular reactors. SIAM Journal on Applied Mathematics, 48(5):

1083–1101, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sturgeon:1988:RES

[2531] M. L. Sturgeon. Range error simulation in optical classifier design. SIAM Journal on Applied Mathematics, 48 (5):1102–1114, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klosek-Dygas:1988:SSN

[2532] M. M. Kłosek-Dygas, B. J. Matkowsky, and Z. Schuss. Stochastic stability of nonlinear oscillators. SIAM Journal on Applied Mathematics, 48(5): 1115–1127, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rebollo:1988:ODT

[2533] T. Chácon Rebollo. Oscillations due to the transport of microstructures. SIAM Journal on Applied Mathematics, 48 (5):1128–1146, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baras:1988:DOA

[2534] J. S. Baras, A. Bensoussan, and M. R.

James. Dynamic observers as asymptotic limits of recursive filters: Special cases. SIAM Journal on Applied Mathematics, 48(5):1147–1158, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hooghiemstra:1988:PSS

[2535] Gerard Hooghiemstra, Michael Keane, and Simon van de Ree. Power series for stationary distributions of coupled processor models. SIAM Journal on Applied Mathematics, 48(5): 1159–1166, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sacks:1988:IPW

[2536] Paul E. Sacks. The inverse problem for a weakly inhomogeneous twodimensional acoustic medium. SIAM Journal on Applied Mathematics, 48 (5):1167–1193, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carroll:1988:SIP

[2537] Robert Carroll and Louise Raphael. Some inverse problems with spectral limitations on data. SIAM Journal on Applied Mathematics, 48(5): 1194–1205, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tongue:1988:TBI

[2538] Benson H. Tongue and Keqin Gu. A theoretical basis for interpolated cell mapping. SIAM Journal on Applied Mathematics, 48(5):1206–1214, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hall:1988:EPM

[2539] Peter Hall. Errata: “On the Processing of a Motion-Blurred Image” [SIAM J. Appl. Math. 47 (1987), no. 2, 441–453, MR 88d:68103]. SIAM Journal on Applied Mathematics, 48(5):1215, October 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [2404].

Ciarkowski:1988:UQU

[2540] Adam Ciarkowski. Uniform and quasiuniform asymptotic expansions of incomplete diffraction integrals. SIAM Journal on Applied Mathematics, 48 (6):1217–1226, December 1988. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

vanHorssen:1988:ATC

[2541] W. T. van Horssen. An asymptotic theory for a class of initial-boundary value problems for weakly nonlinear wave equations with an application to a model of the galloping oscillations of overhead transmission lines. SIAM Journal on Applied Mathematics, 48(6):1227–1243, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hafner:1988:SCS

[2542] Peter Hafner. Strong convergent shock waves near the center of convergence: A power series solution. SIAM Journal on Applied Mathematics, 48(6): 1244–1261, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barta:1988:SBIb

[2543] E. Barta and N. Liron. Slender body interactions for low Reynolds numbers. part II: Body-body interactions. SIAM Journal on Applied Mathematics, 48(6):1262–1280, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Magen:1988:DNE

[2544] Michael Magen. Discussion of a “new”

elliptic linear boundary condition with applications to hydrodynamics. SIAM Journal on Applied Mathematics, 48 (6):1281–1286, December 1988. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Isaacson:1988:RPNb

[2545] E. Isaacson and B. Temple. The Riemann problem near a hyperbolic singularity II. SIAM Journal on Applied Mathematics, 48(6):1287–1301, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Isaacson:1988:RPNc

[2546] E. Isaacson and B. Temple. The Riemann problem near a hyperbolic singularity III. SIAM Journal on Applied Mathematics, 48(6):1302–1318, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Antman:1988:PAS

[2547] Stuart S. Antman. The paradoxical asymptotic status of massless springs. SIAM Journal on Applied Mathematics, 48(6):1319–1334, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vijayakumar:1988:IIM

[2548] Sinnathurai Vijayakumar and Donald E. Cormack. An invariant imbedding method for singular integral evaluation on finite domains. SIAM Journal on Applied Mathematics, 48(6): 1335–1349, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Udwadia:1988:UIB

[2549] F. E. Udwadia. Uniqueness in the identification of a bending beam modeled by finite elements. SIAM Journal on Applied Mathematics, 48(6): 1350–1361, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chadam:1988:WNS

[2550] J. Chadam, P. Ortoleva, and A. Sen.

A weakly nonlinear stability analysis of the reactive infiltration interface. SIAM Journal on Applied Mathematics, 48(6):1362–1378, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Busenberg:1988:ETS

[2551] Stavros Busenberg, Kenneth Cooke, and Mimmo Iannelli. Endemic thresholds and stability in a class of agestructured epidemics. SIAM Journal on Applied Mathematics, 48(6): 1379–1395, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hardin:1988:CDS

[2552] Douglas P. Hardin, Peter Takáč, and G. F. Webb. A comparison of dispersal strategies for survival of spatially heterogeneous populations. SIAM Journal on Applied Mathematics, 48(6): 1396–1423, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gerardi:1988:MAM

[2553] A. Gerardi and G. Nappo. Martingale approach for modeling DNA synthesis.

SIAM Journal on Applied Mathematics, 48(6):1424–1436, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hagan:1988:RHS

[2554] P. S. Hagan, M. Herskowitz, and C. Pirkle. Runaway in highly sensitive tubular reactors. SIAM Journal on Applied Mathematics, 48(6): 1437–1450, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1988:FBPa

[2555] Donald S. Cohen and Thomas Erneux. Free boundary problems in controlled release pharmaceuticals. I. Diffusion in glassy polymers. SIAM Journal on Applied Mathematics, 48(6):1451–1465, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1988:FBPb

[2556] Donald S. Cohen and Thomas Erneux. Free boundary problems in controlled release pharmaceuticals. II. Swellingcontrolled release. SIAM Journal on Applied Mathematics, 48(6):1466–1474, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rubinstein:1988:FFS

[2557] I. Rubinstein. Field and force saturation in ionic equilibrium. SIAM Journal on Applied Mathematics, 48(6): 1475–1486, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gautesen:1988:ODB

[2558] A. K. Gautesen. Oblique derivative boundary conditions and the image method for wedges. SIAM Journal on Applied Mathematics, 48(6): 1487–1492, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1988:CRT

[2559] J. A. Morrison. Conditioned responsetime distribution for a large closed processor-sharing system with multiple classes in very heavy usage. SIAM Journal on Applied Mathematics, 48 (6):1493–1509, December 1988. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dunbar:1988:BRE

[2560] Steven R. Dunbar. A branching random evolution and a nonlinear hyperbolic equation. SIAM Journal on Applied Mathematics, 48(6):1510–1526, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ogawa:1988:OPI

[2561] Hidemitsu Ogawa. An operator pseudo-inversion lemma. SIAM Journal on Applied Mathematics, 48(6): 1527–1531, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wollkind:1988:ENS

[2562] David J. Wollkind and Shingmin Wang.

Erratum: “A Nonlinear Stability Analysis of a Model Equation for Liquid Phase Electro-Epitaxial Growth of a Dilute Binary Substance” [SIAM J.

Appl. Math. 48 (1988), no. 1, 52–78, MR 89a:80001]. SIAM Journal on Applied Mathematics, 48(6):1532, December 1988. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [2469].

Gu:1989:SSP

[2563] Zhong-Mei Gu, N. N. Nefëdov, and R. E. O’Malley, Jr. On singular singularly perturbed initial value problems. SIAM Journal on Applied Mathematics, 49(1):1–25, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lin:1989:SLS

[2564] Xiao-Biao Lin. Shadowing lemma and singularly perturbed boundary value problems. SIAM Journal on Applied Mathematics, 49(1):26–54, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baer:1989:SPT

[2565] S. M. Baer, T. Erneux, and J. Rinzel. The slow passage through a Hopf bifurcation: Delay, memory effects, and resonance. SIAM Journal on Applied Mathematics, 49(1):55–71, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stuart:1989:SFB

[2566] A. M. Stuart. Singular free boundary problems and local bifurcation theory. SIAM Journal on Applied Mathematics, 49(1):72–85, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bhattacharya:1989:ASD

[2567] R. N. Bhattacharya, V. K. Gupta, and H. F. Walker. Asymptotics of solute dispersion in periodic porous media. SIAM Journal on Applied Mathematics, 49(1):86–98, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Guetter:1989:FBP

[2568] Arthur A. Guetter. A free boundary problem in plasma containment. SIAM Journal on Applied Mathematics, 49 (1):99–115, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rubinstein:1989:FRS

[2569] Jacob Rubinstein, Peter Sternberg, and Joseph B. Keller. Fast reaction, slow diffusion, and curve shortening. SIAM Journal on Applied Mathematics, 49 (1):116–133, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Conrad:1989:MSR

[2570] F. Conrad and N. Yebari. Multiple solutions in reaction-diffusion systems with a free boundary. SIAM Journal on Applied Mathematics, 49 (1):134–151, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Young:1989:MID

[2571] Gerald W. Young and Stephen H.

Davis. Morphological instabilities in directional solidification of a binary alloy: End effects. SIAM Journal on Applied Mathematics, 49(1):152–164, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ascher:1989:CSS

[2572] U. Ascher, P. A. Markowich, C. Schmeiser, H. Steinrück, and R. Weiss. Conditioning of the steady state semiconductor device problem. SIAM Journal on Applied Mathematics, 49(1):165–185, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tan:1989:DMC

[2573] W. Y. Tan. On the distribution of mutants in cell populations with both forward and backward mutation. SIAM Journal on Applied Mathematics, 49 (1):186–196, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Altschul:1989:TSM

[2574] Stephen F. Altschul and David J. Lipman. Trees, stars, and multiple biological sequence alignment. SIAM Journal on Applied Mathematics, 49 (1):197–209, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keener:1989:FDD

[2575] James P. Keener. Frequency dependent decoupling of parallel excitable fibers. SIAM Journal on Applied Mathematics, 49(1):210–230, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fowler:1989:GCM

[2576] A. C. Fowler. Generation and creep of magma in the Earth. SIAM Journal on Applied Mathematics, 49(1): 231–245, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fowler:1989:MAG

[2577] A. C. Fowler. A mathematical analysis of glacier surges. SIAM Journal on Applied Mathematics, 49(1): 246–263, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

King:1989:IOSa

[2578] J. R. King. The isolation oxidation of silicon. SIAM Journal on Applied Mathematics, 49(1):264–280, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sanz:1989:MSR

[2579] Jorge L. Sanz. Multidimensional signal representation by zero crossings: An algebraic study. SIAM Journal on Applied Mathematics, 49(1): 281–295, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kotulski:1989:ECF

[2580] Zbigniew Kotulski. Equations for the characteristic functional and moments of the stochastic evolutions with an application. SIAM Journal on Applied Mathematics, 49(1):296–313, February 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burns:1989:SBU

[2581] Timothy J. Burns. Similarity and bifurcation in unstable viscoplastic shear. SIAM Journal on Applied Mathematics, 49(1):314–329, February 1989. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tu:1989:PPA

[2582] Son T. Tu. A phase-plane analysis of bursting in the three-dimensional Bonhoeffer–Van Der Pol equations. SIAM Journal on Applied Mathematics, 49(2):331–343, April 1989. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Geer:1989:HPG

[2583] James F. Geer and Carl M. Andersen. A hybrid perturbation Galerkin technique with applications to slender body theory. SIAM Journal on Applied Mathematics, 49(2):344–361, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reynolds:1989:BHS

[2584] D. W. Reynolds. Bifurcation of harmonic solutions of an integrodifferential equation modelling resonant sloshing. SIAM Journal on Applied Mathematics, 49(2):362–372, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kaklis:1989:LFH

[2585] P. D. Kaklis and G. A. Athanassoulis. The low-frequency heaving motion of a cylinder on the free surface of a liquid. SIAM Journal on Applied Mathematics, 49(2):373–389, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kriegsmann:1989:BCB

[2586] Gregory A. Kriegsmann. Bifurcation in classical bipolar transistor oscillator circuits. SIAM Journal on Applied Mathematics, 49(2):390–403, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Leizarowitz:1989:LEH

[2587] Arie Leizarowitz. On the Lyapunov exponent of a harmonic oscillator driven by a finite-state Markov process. SIAM Journal on Applied Mathematics, 49 (2):404–419, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stensby:1989:FLB

[2588] John Stensby. False lock and bifurcation in Costas loops. SIAM Journal on Applied Mathematics, 49(2):420–431, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jackson:1989:ERC

[2589] T. L. Jackson, A. K. Kapila, and D. S.

Stewart. Evolution of a reaction center in an explosive material. SIAM Journal on Applied Mathematics, 49 (2):432–458, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dold:1989:ATR

[2590] J. W. Dold. Analysis of thermal runaway in the ignition process. SIAM Journal on Applied Mathematics, 49 (2):459–480, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nishiura:1989:LOR

[2591] Yasumasa Nishiura and Masayasu Mimura. Layer oscillations in reactiondiffusion systems. SIAM Journal on Applied Mathematics, 49(2):481–514, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ikeda:1989:WBP

[2592] H. Ikeda and M. Mimura. Waveblocking phenomena in bistable reactiondiffusion systems. SIAM Journal on Applied Mathematics, 49(2):515–538, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dockery:1989:DED

[2593] J. D. Dockery and J. P. Keener. Diffusive effects on dispersion in excitable media. SIAM Journal on Applied Mathematics, 49(2):539–566, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lohofer:1989:TEL

[2594] G. Lohöfer. Theory of an electromagnetically levitated metal sphere. I. Absorbed power. SIAM Journal on Applied Mathematics, 49(2):567–581, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burridge:1989:PRM

[2595] R. Burridge, G. Papanicolaou, P. Sheng, and B. White. Probing a random medium with a pulse. SIAM Journal on Applied Mathematics, 49(2):582–607, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berman:1989:STC

[2596] Simeon M. Berman. Sojourn times in a cone for a class of vector Gaussian processes. SIAM Journal on Applied Mathematics, 49(2):608–616, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1989:AAD

[2597] J. A. Morrison. Asymptotic analysis of a data-handling system with many sources. SIAM Journal on Applied Mathematics, 49(2):617–637, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rawn:1989:GST

[2598] Michael David Rawn. Generalized sampling theorems for Bessel-type transforms of bandlimited functions and distributions. SIAM Journal on Applied Mathematics, 49(2):638–649, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jones:1989:ATD

[2599] Lee K. Jones. Approximation-theoretic derivation of logarithmic entropy principles for inverse problems and unique extension of the maximum entropy method to incorporate prior knowledge. SIAM Journal on Applied Mathematics, 49(2):650–661, April 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gautesen:1989:CFI

[2600] Arthur K. Gautesen. On a class of Fredholm integral operators — asymptotic expansions of solutions and eigenfunctions. SIAM Journal on Applied Mathematics, 49(3):663–675, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Armbruster:1989:KSD

[2601] Dieter Armbruster, John Guckenheimer, and Philip Holmes. Kuramoto– Sivashinsky dynamics on the centerunstable manifold. SIAM Journal on Applied Mathematics, 49(3):676–691, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Glasser:1989:MFT

[2602] M. L. Glasser, V. G. Papageorgiou, and T. C. Bountis. Mel′nikov’s function for two-dimensional mappings. SIAM Journal on Applied Mathematics, 49 (3):692–703, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Forbes:1989:SAF

[2603] Lawrence K. Forbes. A series analysis of forced transverse oscillations in a spring-mass system. SIAM Journal on Applied Mathematics, 49(3):704–719, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cannon:1989:RIT

[2604] John R. Cannon, Ronald B. Guenther, and Fouad A. Mohamed. On the rainfall infiltration through a soil medium. SIAM Journal on Applied Mathematics, 49(3):720–729, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yortsos:1989:LSI

[2605] Y. C. Yortsos and F. J. Hickernell. Linear stability of immiscible displacement in porous media. SIAM Journal on Applied Mathematics, 49(3):730–748, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Trangenstein:1989:MSB

[2606] John A. Trangenstein and John B. Bell.

Mathematical structure of the black-oil model for petroleum reservoir simulation. SIAM Journal on Applied Mathematics, 49(3):749–783, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barkve:1989:RPN

[2607] Tor Barkve. The Riemann problem for a nonstrictly hyperbolic system modeling nonisothermal, two-phase flow in a porous medium. SIAM Journal on Applied Mathematics, 49(3):784–798, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ehrenmark:1989:ONF

[2608] Ulf Torsten Ehrenmark. Overconvergence of the near-field expansion for linearized waves normally incident on a sloping beach. SIAM Journal on Applied Mathematics, 49(3):799–815, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arnold:1989:CID

[2609] E. M. Arnold. Coefficient identification for a divergence form partial differential equation. SIAM Journal on Applied Mathematics, 49(3):816–823, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Avellaneda:1989:OBE

[2610] Marco Avellaneda and Graeme W. Milton. Optimal bounds on the effective bulk modulus of polycrystals. SIAM Journal on Applied Mathematics, 49(3):824–837, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cushing:1989:CMS

[2611] J. M. Cushing. A competition model for size-structured species. SIAM Journal on Applied Mathematics, 49 (3):838–858, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Freedman:1989:CMC

[2612] H. I. Freedman, Joseph W.-H. So, and Paul Waltman. Coexistence in a model of competition in the chemostat incorporating discrete delays. SIAM Journal on Applied Mathematics, 49 (3):859–870, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1989:PSR

[2613] Gen Shun Chen and Anthony W. Leung. Positive solutions for reactor multigroup neutron transport systems: Criticality problem. SIAM Journal on Applied Mathematics, 49(3):871–887, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Narita:1989:SLE

[2614] K. Narita. Stochastic Liénard equation with mean-field interaction. SIAM Journal on Applied Mathematics, 49 (3):888–905, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Donoho:1989:UPS

[2615] David L. Donoho and Philip B.

Stark. Uncertainty principles and signal recovery. SIAM Journal on Applied Mathematics, 49(3):906–931, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kreider:1989:WSA

[2616] K. L. Kreider. A wave splitting approach to time dependent inverse scattering for the stratified cylinder. SIAM Journal on Applied Mathematics, 49(3):932–943, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Somersalo:1989:ODE

[2617] Erkki Somersalo. One-dimensional electromagnetic inverse reflection problem: Formulation as a Riemann– Hilbert problem and imaging of discontinuities. SIAM Journal on Applied Mathematics, 49(3):944–951, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Polewczak:1989:GEA

[2618] Jacek Polewczak. Global existence and asymptotic behavior for the nonlinear Enskog equation. SIAM Journal on Applied Mathematics, 49(3):952–959, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Friedel:1989:SPS

[2619] J. Friedel and P. Widmayer. A simple proof of the Steiner ratio conjecture for five points. SIAM Journal on Applied Mathematics, 49(3):960–967, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miles:1989:RSB

[2620] John Miles and Peter J. Bryant. Resonance and symmetry breaking for a Duffing oscillator. SIAM Journal on Applied Mathematics, 49(3):968–981, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). With an appendix by Peter J. Bryant.

Dewynne:1989:MMF

[2621] J. Dewynne, J. R. Ockendon, and P. Wilmott. On A mathematical model for fiber tapering. SIAM Journal on Applied Mathematics, 49(4):983–990, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fowler:1989:SFH

[2622] A. C. Fowler. Secondary frost heave in freezing soils. SIAM Journal on Applied Mathematics, 49(4):991–1008, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vogel:1989:SLD

[2623] Thomas I. Vogel. Stability of a liquid drop trapped between two parallel planes. II. General contact angles. SIAM Journal on Applied Mathematics, 49(4):1009–1028, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ronney:1989:TSP

[2624] P. D. Ronney and G. I. Sivashinsky. A theoretical study of propagation and extinction of nonsteady spherical flame fronts. SIAM Journal on Applied Mathematics, 49(4):1029–1046, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bayliss:1989:PDG

[2625] A. Bayliss, B. J. Matkowsky, and M. Minkoff. Period doubling gained, period doubling lost. SIAM Journal on Applied Mathematics, 49(4): 1047–1063, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

King:1989:IOSb

[2626] J. R. King. The isolation oxidation of silicon: The reaction-controlled case. SIAM Journal on Applied Mathematics, 49(4):1064–1080, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

King:1989:DPD

[2627] J. R. King. On the diffusion of point defects in silicon. SIAM Journal on Applied Mathematics, 49(4):1081–1101, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Steinruck:1989:BAO

[2628] Herbert Steinrück. A bifurcation analysis of the one-dimensional steady-state semiconductor device equations. SIAM Journal on Applied Mathematics, 49 (4):1102–1121, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Szmolyan:1989:SPA

[2629] Peter Szmolyan. A singular perturbation analysis of the transient semiconductor device equations. SIAM Journal on Applied Mathematics, 49 (4):1122–1135, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Edelstein-Keshet:1989:MBN

[2630] Leah Edelstein-Keshet and Bard Ermentrout. Models for branching networks in two dimensions. SIAM Journal on Applied Mathematics, 49(4): 1136–1157, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yanagida:1989:SDP

[2631] Eiji Yanagida and Kenjiro Maginu. Stability of double-pulse solutions in nerve axon equations. SIAM Journal on Applied Mathematics, 49(4): 1158–1173, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hodnett:1989:SDI

[2632] P. F. Hodnett and T. P. Moloney. On the structure during interaction of the two-soliton solution of the Korteweg– de Vries equation. SIAM Journal on Applied Mathematics, 49(4):1174–1187, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Clarkson:1989:HTL

[2633] P. A. Clarkson, A. S. Fokas, and M. J.

Ablowitz. Hodograph transformations of linearizable partial differential equations. SIAM Journal on Applied Mathematics, 49(4):1188–1209, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Newton:1989:BNN

[2634] Paul K. Newton. Branching near nonlinear plane waves in dispersive systems. SIAM Journal on Applied Mathematics, 49(4):1210–1222, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Turner:1989:NMS

[2635] R. E. L. Turner and P. Villaggio. A note on the motion of a string on a unilaterally reacting foundation. SIAM Journal on Applied Mathematics, 49 (4):1223–1230, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lachowicz:1989:EQA

[2636] M. Lachowicz and R. Monaco. Existence and quantitative analysis of the solutions to the initial value problem for the discrete Boltzmann equation in all space. SIAM Journal on Applied Mathematics, 49(4):1231–1241, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arnold:1989:LER

[2637] Ludwig Arnold and Peter Kloeden.

Lyapunov exponents and rotation number of two-dimensional systems with telegraphic noise. SIAM Journal on Applied Mathematics, 49(4):1242–1274, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Krishna:1989:CER

[2638] Bal Krishna and Hari Krishna. Computationally efficient reduced polynomial based algorithms for Hermitian Toeplitz matrices. SIAM Journal on Applied Mathematics, 49(4):1275–1282, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dassios:1989:ESA

[2639] George Dassios and Vassilis Kostopoulos. Errata: “The Scattering Amplitudes and Cross Sections in the Theory of Thermoelasticity” [SIAM J. Appl. Math. 48 (1988), no. 1, 79–98, MR 89c:73090]. SIAM Journal on Applied Mathematics, 49(4):1283–1284, August 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [2470].

Vijayakumar:1989:NCN

[2640] Sinnathurai Vijayakumar and Donald E. Cormack. A new concept in near-singular integral evaluation: The continuation approach. SIAM Journal on Applied Mathematics, 49(5): 1285–1295, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ikegami:1989:RNE

[2641] Gikō Ikegami. On regularizations of nonlinear electrical circuits. SIAM Journal on Applied Mathematics, 49 (5):1296–1309, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Artola:1989:NDI

[2642] Miguel Artola and Andrew J. Majda.

Nonlinear development of instabilities in supersonic vortex sheets. II. Resonant interaction among kink modes. SIAM Journal on Applied Mathematics, 49(5):1310–1349, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Riley:1989:MSF

[2643] N. Riley and P. D. Weidman. Multiple solutions of the Falkner–Skan equation for flow past a stretching boundary. SIAM Journal on Applied Mathematics, 49(5):1350–1358, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maas:1989:CFG

[2644] Leo R. M. Maas. A closed form Green function describing diffusion in a strained flow field. SIAM Journal on Applied Mathematics, 49(5): 1359–1373, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Panayotounakos:1989:CFS

[2645] Dimitrios E. Panayotounakos and Basilios Zisis. Closed-form solutions of the nonlinear partial differential equations governing the stress-field in a rigid, perfectly-plastic body. SIAM Journal on Applied Mathematics, 49 (5):1374–1389, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Margolis:1989:DTI

[2646] Stephen B. Margolis and Forman A.

Williams. Diffusional/thermal instability of a solid propellant flame. SIAM Journal on Applied Mathematics, 49 (5):1390–1420, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bayliss:1989:CCF

[2647] A. Bayliss, B. J. Matkowsky, and M. Minkoff. Cascading cellular flames. SIAM Journal on Applied Mathematics, 49(5):1421–1432, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bukiet:1989:ECD

[2648] Bruce Bukiet. The effect of curvature on detonation speed. SIAM Journal on Applied Mathematics, 49(5): 1433–1446, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lui:1989:ABS

[2649] Roger Lui. Asymptotic behavior of solutions to an evolutionary ecology model with diffusion. SIAM Journal on Applied Mathematics, 49(5): 1447–1461, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Muratori:1989:RCC

[2650] Simona Muratori and Sergio Rinaldi. Remarks on competitive coexistence. SIAM Journal on Applied Mathematics, 49(5):1462–1472, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cho:1989:OPB

[2651] S. K. Cho and C. M. Chu. Optimal polarization in bistatic scattering. SIAM Journal on Applied Mathematics, 49 (5):1473–1479, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hagan:1989:MET

[2652] Patrick S. Hagan, Charles R. Doering, and C. David Levermore. Mean exit times for particles driven by weakly colored noise. SIAM Journal on Applied Mathematics, 49(5):1480–1513, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lefebvre:1989:FPD

[2653] Mario Lefebvre. First-passage densities of a two-dimensional process. SIAM Journal on Applied Mathematics, 49 (5):1514–1523, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nadim:1989:DSP

[2654] Ali Nadim. Dispersion in spatially periodic media: The moment-difference expansion method. SIAM Journal on Applied Mathematics, 49(5):1524–1537, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stadje:1989:CPR

[2655] Wolfgang Stadje. Coverage problems for random intervals. SIAM Journal on Applied Mathematics, 49(5): 1538–1551, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Spouge:1989:SDP

[2656] John L. Spouge. Speeding up dynamic programming algorithms for finding optimal lattice paths. SIAM Journal on Applied Mathematics, 49(5):

1552–1566, October 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Olmstead:1989:SPI

[2657] W. E. Olmstead and J. S. Angell. Singularly perturbed integral equations with endpoint boundary layers. SIAM Journal on Applied Mathematics, 49(6):1567–1584, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wolansky:1989:BVE

[2658] G. Wolansky. The barotropic vorticity equation under forcing and dissipation: Bifurcations of nonsymmetric responses and multiplicity of solutions. SIAM Journal on Applied Mathematics, 49(6):1585–1607, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wilmott:1989:SSA

[2659] P. Wilmott. The stretching of a slender, axisymmetric, viscous inclusion — part I: Asymptotic analysis. SIAM Journal on Applied Mathematics, 49(6): 1608–1616, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fitt:1989:SSA

[2660] A. D. Fitt and P. Wilmott. The stretching of a slender, axisymmetric, viscous inclusion — part II: Numerical solution and results. SIAM Journal on Applied Mathematics, 49(6): 1617–1634, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Luke:1989:CMR

[2661] Jonathan H. C. Luke. Convergence of a multiple reflection method for calculating Stokes flow in a suspension. SIAM Journal on Applied Mathematics, 49(6):1635–1651, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abrahams:1989:DCS

[2662] I. David Abrahams, Gregory A. Kriegsmann, and Edward L. Reiss. On the development of caustics in shear flows over rigid walls. SIAM Journal on Applied Mathematics, 49(6): 1652–1664, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1989:ADB

[2663] G. Chen, S. G. Krantz, D. L. Russell, C. E. Wayne, H. H. West, and M. P. Coleman. Analysis, designs, and behavior of dissipative joints for coupled beams. SIAM Journal on Applied Mathematics, 49(6):1665–1693, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Liu:1989:ESA

[2664] Kang Sheng Liu, Fa Lun Huang, and Goong Chen. Exponential stability analysis of a long chain of coupled vibrating strings with dissipative linkage. SIAM Journal on Applied Mathematics, 49(6):1694–1707, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sarason:1989:COT

[2665] Leonard Sarason. Comparison of one- and three-dimensional models for tracer evolution in tissue cylinders. SIAM Journal on Applied Mathematics, 49(6):1708–1721, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rubinstein:1989:RDP

[2666] Jacob Rubinstein, Peter Sternberg, and Joseph B. Keller. Reaction-diffusion processes and evolution to harmonic maps. SIAM Journal on Applied Mathematics, 49(6):1722–1733, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schmeiser:1989:SRB

[2667] Christian Schmeiser. On strongly reverse biased semiconductor diodes. SIAM Journal on Applied Mathematics, 49(6):1734–1748, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jing:1989:CBJ

[2668] Zhu Jun Jing. Chaotic behavior in the Josephson equations with periodic force. SIAM Journal on Applied Mathematics, 49(6):1749–1758, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kuang:1989:LCC

[2669] Yang Kuang. Limit cycles in a chemostat-related model. SIAM Journal on Applied Mathematics, 49(6): 1759–1767, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Munoz:1989:ABC

[2670] Eduardo M. Muñoz and James F. Selgrade. Asymptotic behavior for a competitive model with genetic variation.

SIAM Journal on Applied Mathematics, 49(6):1768–1778, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Noonburg:1989:NNM

[2671] V. W. Noonburg. A neural network modeled by an adaptive Lotka–Volterra system. SIAM Journal on Applied Mathematics, 49(6):1779–1792, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Roozen:1989:AST

[2672] H. Roozen. An asymptotic solution to a two-dimensional exit problem arising in population dynamics. SIAM Journal on Applied Mathematics, 49(6): 1793–1810, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klosek:1989:FOD

[2673] M. M. Kłosek, B. J. Matkowsky, and Z. Schuss. First-order dynamics driven by rapid Markovian jumps. SIAM Journal on Applied Mathematics, 49 (6):1811–1833, December 1989. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jensen:1989:SAS

[2674] D. R. Jensen and R. V. Foutz. The structure and analysis of spherical time-dependent processes. SIAM Journal on Applied Mathematics, 49(6): 1834–1844, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Teboulle:1989:SDP

[2675] Marc Teboulle. A simple duality proof for quadratically constrained entropy functionals and extension to convex constraints. SIAM Journal on Applied Mathematics, 49(6):1845–1850, December 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Strain:1990:VEU

[2676] John Strain. Velocity effects in unstable solidification. SIAM Journal on Applied Mathematics, 50(1):1–15, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Villamizar:1990:ESV

[2677] Vianey Villamizar. Elastic scattering from a viscous incompressible fluid sphere at low Reynolds number. SIAM Journal on Applied Mathematics, 50 (1):16–32, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schaeffer:1990:ICS

[2678] David G. Schaeffer, Michael Shearer, and E. Bruce Pitman. Instability in critical state theories of granular flow. SIAM Journal on Applied Mathematics, 50(1):33–47, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hunter:1990:LMS

[2679] C. Hunter, M. Tajdari, and S. D. Boyer. On Lagerstrom’s model of slow incompressible viscous flow. SIAM Journal on Applied Mathematics, 50 (1):48–63, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sangani:1990:CDC

[2680] A. S. Sangani. Conductivity of Ndimensional composites containing hyperspherical inclusion. SIAM Journal on Applied Mathematics, 50(1):64–73, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carasso:1990:IRA

[2681] Alfred S. Carasso. Impulse response acquisition as an inverse heat conduction problem. SIAM Journal on Applied Mathematics, 50(1):74–90, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See corrigendum [2731].

Segel:1990:PIC

[2682] Lee A. Segel and Alan S. Perelson. A paradoxical instability caused by relatively short-range inhibition. SIAM Journal on Applied Mathematics, 50 (1):91–107, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mirollo:1990:JBH

[2683] Renato E. Mirollo and Steven H. Strogatz. Jump bifurcation and hysteresis in an infinite-dimensional dynamical system of coupled spins. SIAM Journal on Applied Mathematics, 50 (1):108–124, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ermentrout:1990:ODS

[2684] G. B. Ermentrout and N. Kopell. Oscillator death in systems of coupled neural oscillators. SIAM Journal on Applied Mathematics, 50(1):125–146, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Holmes:1990:MSM

[2685] Mark H. Holmes and Jonathan Bell. A model of a sensory mechanoreceptor derived from homogenization. SIAM Journal on Applied Mathematics, 50 (1):147–166, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reed:1990:ATW

[2686] Michael C. Reed, Stephanos Venakides, and Jacob Joseph Blum. Approximate traveling waves in linear reactionhyperbolic equations. SIAM Journal on Applied Mathematics, 50(1): 167–180, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Budd:1990:SCP

[2687] Chris J. Budd, Avner Friedman, Bryce Mcleod, and Adam A. Wheeler. The space charge problem. SIAM Journal on Applied Mathematics, 50(1): 181–198, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stevenson:1990:GFH

[2688] Randy C. Stevenson. Green’s function for the Helmholtz equation of a scalar wave in an anisotropic halfspace. SIAM Journal on Applied Mathematics, 50 (1):199–215, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Santosa:1990:BAE

[2689] Fadil Santosa and Michael Vogelius.

A backprojection algorithm for electrical impedance imaging. SIAM Journal on Applied Mathematics, 50(1): 216–243, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Trietsch:1990:IAS

[2690] Dan Trietsch and Frank Hwang. An improved algorithm for Steiner trees. SIAM Journal on Applied Mathematics, 50(1):244–263, February 1990. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:1990:FCP

[2691] Charles Knessl. On finite capacity processor-shared queues. SIAM Journal on Applied Mathematics, 50(1): 264–287, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fishburn:1990:SUD

[2692] P. C. Fishburn, J. C. Lagarias, J. A.

Reeds, and L. A. Shepp. Sets uniquely determined by projections on axes I. continuous case. SIAM Journal on Applied Mathematics, 50(1):288–306, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hunt:1990:EAC

[2693] Fern Hunt. Error analysis and convergence of capacity dimension algorithms. SIAM Journal on Applied Mathematics, 50(1):307–321, February 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:1990:PAR

[2694] Charles Knessl and Joseph B. Keller. Partition asymptotics from recursion equations. SIAM Journal on Applied Mathematics, 50(2):323–338, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ablowitz:1990:HSN

[2695] Mark J. Ablowitz and B. M. Herbst. On homoclinic structure and numerically induced chaos for the nonlinear Schrödinger equation. SIAM Journal on Applied Mathematics, 50(2):339– 351, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lasseigne:1990:SVB

[2696] D. Glenn Lasseigne and W. E. Olmstead. Stability of a viscoelastic Burgers flow. SIAM Journal on Applied Mathematics, 50(2):352–360, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rotenberry:1990:ECB

[2697] James M. Rotenberry and Philip G. Saffman. Effect of compliant boundaries on weakly nonlinear shear waves in channel flow. SIAM Journal on Applied Mathematics, 50(2):361–394, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Antman:1990:IGS

[2698] Stuart S. Antman and John F. Pierce. The intricate global structure of buckled states of compressible columns.

SIAM Journal on Applied Mathematics, 50(2):395–419, April 1990. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Riley:1990:LWM

[2699] D. S. Riley and S. H. Davis. Longwave morphological instabilities in the directional solidification of a dilute binary mixture. SIAM Journal on Applied Mathematics, 50(2):420–436, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bayliss:1990:TRC

[2700] A. Bayliss and B. J. Matkowsky.

Two routes to chaos in condensed phase combustion. SIAM Journal on Applied Mathematics, 50(2):437–459, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Henderson:1990:CBR

[2701] M. E. Henderson and H. B. Keller.

Complex bifurcation from real paths. SIAM Journal on Applied Mathematics, 50(2):460–482, April 1990. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1990:CTH

[2702] Donald S. Cohen and Thomas Erneux.

Changing time history in moving boundary problems. SIAM Journal on Applied Mathematics, 50(2):483–489, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Elphick:1990:PPP

[2703] C. Elphick, E. Meron, and E. A. Spiegel. Patterns of propagating pulses. SIAM Journal on Applied Mathematics, 50(2):490–503, April 1990. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

OMalley:1990:ASS

[2704] R. E. O’Malley, Jr. and C. Schmeiser. The asymptotic solution of a semiconductor device problem involving reverse bias. SIAM Journal on Applied Mathematics, 50(2):504–520, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ting:1990:VFS

[2705] Lu Ting and Michael J. Miksis. On vortical flow and sound generation. SIAM Journal on Applied Mathematics, 50 (2):521–536, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kath:1990:BLA

[2706] W. L. Kath, J. Jiao, and M. E. Marhic. Boundary layer analysis of infrared whispering-gallery waveguides. SIAM Journal on Applied Mathematics, 50 (2):537–546, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jones:1990:TWA

[2707] D. S. Jones. A thin wire approximation. SIAM Journal on Applied Mathematics, 50(2):547–558, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jones:1990:NSR

[2708] D. S. Jones and G. A. Kriegsmann.

Note on surface radiation conditions. SIAM Journal on Applied Mathematics, 50(2):559–568, April 1990. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bleistein:1990:SNA

[2709] Norman Bleistein and Jack K. Cohen.

Stacking of narrow aperture common shot inversions. SIAM Journal on Applied Mathematics, 50(2):569–594, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Naeh:1990:DAE

[2710] T. Naeh, M. M. Kłosek, B. J.

Matkowsky, and Z. Schuss. A direct approach to the exit problem. SIAM Journal on Applied Mathematics, 50 (2):595–627, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shaw:1990:AMC

[2711] L. J. Shaw and J. D. Murray. Analysis of a model for complex skin patterns. SIAM Journal on Applied Mathematics, 50(2):628–648, April 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Drew:1990:EGS

[2712] Donald A. Drew. Evolution of geometric statistics. SIAM Journal on Applied Mathematics, 50(3):649–666, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Holden:1990:SHB

[2713] Lars Holden. On the strict hyperbolicity of the Buckley–Leverett equations for three-phase flow in a porous medium. SIAM Journal on Applied Mathematics, 50(3):667–682, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

James:1990:LEC

[2714] Richard James, Robert Lipton, and Adam Lutoborski. Laminar elastic composites with crystallographic symmetry. SIAM Journal on Applied Mathematics, 50(3):683–702, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McKenna:1990:TWS

[2715] P. J. McKenna and W. Walter. Travelling waves in a suspension bridge. SIAM Journal on Applied Mathematics, 50(3):703–715, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hori:1990:ASC

[2716] Muneo Hori and Siavouche Nemat-Nasser. Asymptotic solution of a class of strongly singular integral equations. SIAM Journal on Applied Mathematics, 50(3):716–725, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stone:1990:RPH

[2717] Emily Stone and Philip Holmes. Random perturbations of heteroclinic attractors. SIAM Journal on Applied Mathematics, 50(3):726–743, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Feng:1990:ASI

[2718] Xiangbo Feng and Kenneth A. Loparo.

Almost sure instability of the random harmonic oscillator. SIAM Journal on Applied Mathematics, 50(3):744–759, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kevrekidis:1990:BSA

[2719] Ioannis G. Kevrekidis, Basil Nicolaenko, and James C. Scovel. Back in the saddle again: A computer assisted study of the Kuramoto–Sivashinsky equation. SIAM Journal on Applied Mathematics, 50(3):760–790, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Terrones:1990:SBS

[2720] Guillermo Terrones, David W. McLaughlin, Edward A. Overman, and Arne J. Pearlstein. Stability and bifurcation of spatially coherent solutions of the damped-driven NLS equation. SIAM Journal on Applied Mathematics, 50 (3):791–818, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abrahams:1990:GWH

[2721] I. D. Abrahams and G. R. Wickham. General Wiener–Hopf factorization of matrix kernels with exponential phase factors. SIAM Journal on Applied Mathematics, 50(3):819–838, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Holmes:1990:NID

[2722] Mark H. Holmes. Nonlinear ionic diffusion through charged polymeric gels.

SIAM Journal on Applied Mathematics, 50(3):839–852, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rao:1990:CGC

[2723] Nagaraj S. Rao and Emilio O. Roxin. Controlled growth of competing species. SIAM Journal on Applied Mathematics, 50(3):853–864, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kasonga:1990:MLT

[2724] Raphael A. Kasonga. Maximum likelihood theory for large interacting systems. SIAM Journal on Applied Mathematics, 50(3):865–875, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1990:UPG

[2725] Kennan T. Smith. The uncertainty principle on groups. SIAM Journal on Applied Mathematics, 50(3):876–882, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wazwaz:1990:TTP

[2726] Abdul-Majid Wazwaz. Two turning points of second order. SIAM Journal on Applied Mathematics, 50(3):883– 892, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zayed:1990:LIK

[2727] Ahmed I. Zayed, Guido Hinsen, and Paul L. Butzer. On Lagrange interpolation and Kramer-type sampling theorems associated with Sturm– Liouville problems. SIAM Journal on Applied Mathematics, 50(3):893–909, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kobayashi:1990:EED

[2728] Mei Kobayashi. Eigenfunction expansions: A discontinuous version. SIAM Journal on Applied Mathematics, 50 (3):910–917, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Curtis:1990:DRN

[2729] Edward B. Curtis and James A. Morrow. Determining the resistors in a network. SIAM Journal on Applied Mathematics, 50(3):918–930, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sacks:1990:VIP

[2730] Paul E. Sacks. A velocity inversion problem involving an unknown source. SIAM Journal on Applied Mathematics, 50(3):931–941, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carasso:1990:CIR

[2731] Alfred S. Carasso. Corrigendum:

“Impulse Response Acquisition as an Inverse Heat Conduction Problem” [SIAM J. Appl. Math. 50 (1990), no. 1, 74–90]. SIAM Journal on Applied Mathematics, 50(3):942, June 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See

Levi:1990:PAP

[2732] Mark Levi. A period-adding phenomenon. SIAM Journal on Applied Mathematics, 50(4):943–955, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ingraham:1990:SGA

[2733] R. L. Ingraham. A sequence of Galerkin approximations to the center manifold for a certain class of nonlinear partial differential equational dynamical systems. SIAM Journal on Applied Mathematics, 50(4):956–971, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Murphy:1990:ETS

[2734] K. A. Murphy. Estimation of time- and state-dependent delays and other parameters in functional differential equations. SIAM Journal on Applied Mathematics, 50(4):972–1000, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cannon:1990:SIN

[2735] John R. Cannon and David M. Elwood. A system involving a nonlinear firstorder partial differential equation coupled with a nonlinear Volterra integral equation. SIAM Journal on Applied Mathematics, 50(4):1001–1013, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kopell:1990:PTO

[2736] N. Kopell and G. B. Ermentrout. Phase transitions and other phenomena in chains of coupled oscillators. SIAM Journal on Applied Mathematics, 50 (4):1014–1052, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Woodward:1990:SPL

[2737] D. E. Woodward. Synthesis of phaselocking patterns in networks with cyclic group symmetries. SIAM Journal on Applied Mathematics, 50(4):1053–1072, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Taylor:1990:NNN

[2738] J. G. Taylor. Noisy neural net states and their time evolution. SIAM Journal on Applied Mathematics, 50(4): 1073–1087, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kriegsmann:1990:MHC

[2739] G. A. Kriegsmann, M. E. Brodwin, and D. G. Watters. Microwave heating of a ceramic halfspace. SIAM Journal on Applied Mathematics, 50 (4):1088–1098, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ward:1990:AMM

[2740] M. J. Ward, F. M. Odeh, and D. S. Cohen. Asymptotic methods for metal oxide semiconductor field effect transistor modeling. SIAM Journal on Applied Mathematics, 50(4):1099–1125, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grayhack:1990:BRS

[2741] William T. Grayhack and Thomas J.

Mahar. Buckling of rib-stiffened plates: An asymptotic approach. SIAM Journal on Applied Mathematics, 50(4):

1126–1133, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yaesh:1990:AAO

[2742] I. Yaesh, B. Z. Bobrovsky, and Z. Schuss. Asymptotic analysis of the optimal filtering problem for twodimensional diffusions measured in a low noise channel. SIAM Journal on Applied Mathematics, 50(4):1134–1155, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Leizarowitz:1990:ERL

[2743] Arie Leizarowitz. Exact results for the Lyapunov exponents of certain linear Ito systems. SIAM Journal on Applied Mathematics, 50(4):1156–1165, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Percus:1990:EPC

[2744] Ora E. Percus and J. K. Percus. Elementary properties of clock-regulated queues. SIAM Journal on Applied Mathematics, 50(4):1166–1175, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhang:1990:ARW

[2745] Zhensheng Zhang. Analytical results for waiting time and system size distributions in two parallel queueing systems. SIAM Journal on Applied Mathematics, 50(4):1176–1193, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Miamee:1990:PCP

[2746] A. G. Miamee. Periodically correlated processes and their stationary dilations. SIAM Journal on Applied Mathematics, 50(4):1194–1199, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pittel:1990:GVM

[2747] B. Pittel and W. A. Woyczyński. A graph-valued Markov process as ringsallowed polymerization model: Subcritical behavior. SIAM Journal on Applied Mathematics, 50(4):1200–1220, August 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mandal:1990:SWW

[2748] B. N. Mandal and P. K. Kundu. Scattering of water waves by a submerged nearly vertical plate. SIAM Journal on Applied Mathematics, 50(5): 1221–1231, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McCaughan:1990:BAA

[2749] F. E. McCaughan. Bifurcation analysis of axial flow compressor stability. SIAM Journal on Applied Mathematics, 50(5):1232–1253, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1990:WPM

[2750] Goong Chen and Jian Xin Zhou. The wave propagation method for the analysis of boundary stabilization in vibrating structures. SIAM Journal on Applied Mathematics, 50(5):1254–1283, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cox:1990:SMS

[2751] Robert W. Cox. Shocks in a model for stress-driven diffusion. SIAM Journal on Applied Mathematics, 50(5): 1284–1299, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dassios:1990:RET

[2752] George Dassios and Vassilis Kostopoulos. On Rayleigh expansions in thermoelastic scattering. SIAM Journal on Applied Mathematics, 50(5): 1300–1324, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moses:1990:AEB

[2753] Harry E. Moses and Reese T. Prosser. Acoustic and electromagnetic bullets: Derivation of new exact solutions of the acoustic and Maxwell’s equations. SIAM Journal on Applied Mathematics, 50(5):1325–1340, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1990:AAD

[2754] G. Chen, S. A. Fulling, F. J. Narcowich, and C. Qi. An asymptotic average decay rate for the wave equation with variable coefficient viscous damping. SIAM Journal on Applied Mathematics, 50(5):1341–1347, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Collings:1990:GAT

[2755] John B. Collings and David J. Wollkind. A global analysis of a temperature-dependent model system for a mite predator-prey interaction.

SIAM Journal on Applied Mathematics, 50(5):1348–1372, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Alexander:1990:RSF

[2756] James C. Alexander, Eusebius J.

Doedel, and Hans G. Othmer. On the resonance structure in a forced excitable system. SIAM Journal on Applied Mathematics, 50(5):1373–1418, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moll:1990:CTS

[2757] V. Moll and S. I. Rosencrans. Calculation of the threshold surface for nerve equations. SIAM Journal on Applied Mathematics, 50(5):1419–1441, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peters:1990:NMP

[2758] Craig Steven Peters and Marc Mangel.

New methods for the problem of collective ruin. SIAM Journal on Applied Mathematics, 50(5):1442–1456, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fox:1990:DTC

[2759] Bennett L. Fox and Peter W. Glynn.

Discrete-time conversion for simulating finite-horizon Markov processes. SIAM Journal on Applied Mathematics, 50 (5):1457–1473, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Geer:1990:HPG

[2760] James F. Geer and Carl M. Andersen. A hybrid perturbation-Galerkin technique that combines multiple expansions. SIAM Journal on Applied Mathematics, 50(5):1474–1495, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dadfar:1990:RPS

[2761] Mohammad B. Dadfar and James F. Geer. Resonances and power series solutions of the forced Van Der Pol oscillator. SIAM Journal on Applied Mathematics, 50(5):1496–1506, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gautesen:1990:ASS

[2762] A. K. Gautesen. Asymptotic solutions to some two-dimensional singularly perturbed Fredholm integral equations. SIAM Journal on Applied Mathematics, 50(5):1507–1515, October 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Caflisch:1990:LAM

[2763] R. E. Caflisch, O. F. Orellana, and M. Siegel. A localized approximation method for vortical flows. SIAM Journal on Applied Mathematics, 50(6): 1517–1532, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ting:1990:SJT

[2764] Lu Ting and Joseph B. Keller. Slender jets and thin sheets with surface tension. SIAM Journal on Applied Mathematics, 50(6):1533–1546, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mercer:1990:CMD

[2765] G. N. Mercer and A. J. Roberts. A centre manifold description of contaminant dispersion in channels with varying flow properties. SIAM Journal on Applied Mathematics, 50(6): 1547–1565, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kida:1990:TRA

[2766] T. Kida, R. Hayashi, and Z. Yasutomi. A theoretical remark about waves on a static water surface beneath a layer of moving air. SIAM Journal on Applied Mathematics, 50(6): 1566–1579, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Coffey:1990:SEG

[2767] Mark W. Coffey. On series expansions giving closed-form solutions of Korteweg–de Vries-like equations. SIAM Journal on Applied Mathematics, 50(6):1580–1592, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Poupaud:1990:SNB

[2768] F. Poupaud. On a system of nonlinear Boltzmann equations of semiconductor physics. SIAM Journal on Applied Mathematics, 50(6):1593–1606, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Angell:1990:RCP

[2769] T. S. Angell, R. E. Kleinman, and F. Hettlich. The resistive and conductive problems for the exterior Helmholtz equation. SIAM Journal on Applied Mathematics, 50(6):1607–1622, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gisser:1990:ECC

[2770] D. G. Gisser, D. Isaacson, and J. C. Newell. Electric current computed tomography and eigenvalues. SIAM Journal on Applied Mathematics, 50 (6):1623–1634, December 1990. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ikehata:1990:IFL

[2771] Masaru Ikehata. Inversion formulas for the linearized problem for an inverse boundary value problem in elastic prospection. SIAM Journal on Applied Mathematics, 50(6):1635–1644, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mirollo:1990:SPC

[2772] Renato E. Mirollo and Steven H. Strogatz. Synchronization of pulse-coupled biological oscillators. SIAM Journal on Applied Mathematics, 50(6): 1645–1662, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Britton:1990:SSP

[2773] N. F. Britton. Spatial structures and periodic travelling waves in an integrodifferential reaction-diffusion population model. SIAM Journal on Applied Mathematics, 50(6):1663–1688, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bluman:1990:RAO

[2774] George Bluman. A reduction algorithm for an ordinary differential equation admitting a solvable Lie group. SIAM Journal on Applied Mathematics, 50(6):1689–1705, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bluman:1990:ISO

[2775] George Bluman. Invariant solutions for ordinary differential equations. SIAM Journal on Applied Mathematics, 50 (6):1706–1715, December 1990. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bourland:1990:SCT

[2776] F. Jay Bourland and Richard Haberman. Separatrix crossing: Timeinvariant potentials with dissipation. SIAM Journal on Applied Mathematics, 50(6):1716–1744, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

delrioCastillo:1990:BCI

[2777] Rafael del rio Castillo. On boundary conditions of an inverse Sturm– Liouville problem. SIAM Journal on Applied Mathematics, 50(6):1745–1751, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Littlefield:1990:FAH

[2778] David L. Littlefield and Prateen V. Desai. Frobenius analysis of higher order equations: Incipient buoyant thermal convection. SIAM Journal on Applied Mathematics, 50(6):1752–1763, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hunter:1990:SCP

[2779] C. Hunter and M. Tajdari. Singular complex periodic solutions of Van Der Pol’s equation. SIAM Journal on Applied Mathematics, 50(6):1764–1779, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yagle:1990:ASL

[2780] Andrew E. Yagle. Analogues of split Levinson, Schur, and lattice algorithms for three-dimensional random field estimation problems. SIAM Journal on Applied Mathematics, 50(6):1780–1799, December 1990. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ehrenmark:1991:VWM

[2781] Ulf Torsten Ehrenmark. On viscous wave motion over a plane beach. SIAM Journal on Applied Mathematics, 51 (1):1–19, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cullen:1991:GLS

[2782] M. J. P. Cullen, J. Norbury, and R. J. Purser. Generalised Lagrangian solutions for atmospheric and oceanic flows. SIAM Journal on Applied Mathematics, 51(1):20–31, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boules:1991:EUF

[2783] Adel N. Boules. On the existence and uniqueness of the flow in a torus. SIAM Journal on Applied Mathematics, 51 (1):32–39, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Macdonald:1991:ZAS

[2784] D. A. Macdonald. The zeros of J_{1}^{2}(ζ)− J_{0}(ζ)J_{2}(ζ) = 0 with an application to swirling flow in a tube. SIAM Journal on Applied Mathematics, 51 (1):40–48, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Golubitsky:1991:TRP

[2785] Martin Golubitsky, Martin Krupa, and Chjan Lim. Time reversibility and particle sedimentation. SIAM Journal on Applied Mathematics, 51(1):49– 72, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Olagunju:1991:PF

[2786] David O. Olagunju and Bernard J.

Matkowsky. Polyhedral flames. SIAM Journal on Applied Mathematics, 51 (1):73–89, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ward:1991:MSS

[2787] M. J. Ward, L. G. Reyna, and F. M.

Odeh. Multiple steady-state solutions in a multijunction semiconductor device. SIAM Journal on Applied Mathematics, 51(1):90–123, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McNabb:1991:DRM

[2788] A. McNabb and L. Bass. A diffusionreaction model for the cellular uptake of protein-bound ligands. SIAM Journal on Applied Mathematics, 51 (1):124–149, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Calzolari:1991:LLN

[2789] A. Calzolari, C. Costantini, and A. Gerardi. Law of large numbers for DNA distribution in an interacting cell population. SIAM Journal on Applied Mathematics, 51(1):150–159, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:1991:DAF

[2790] Charles Knessl. On the diffusion approximation to a fork and join queueing model. SIAM Journal on Applied Mathematics, 51(1):160–171, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Markov:1991:FFS

[2791] Konstantin Z. Markov. On the factorial functional series and their application to random media. SIAM Journal on Applied Mathematics, 51 (1):172–186, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:1991:HTAa

[2792] Charles Knessl and J. A. Morrison. Heavy-traffic analysis of a datahandling system with many sources. SIAM Journal on Applied Mathematics, 51(1):187–213, February 1991. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:1991:APE

[2793] Charles Knessl and Joseph B. Keller.

Asymptotic properties of eigenvalues of integral equations. SIAM Journal on Applied Mathematics, 51(1): 214–232, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wazwaz:1991:MES

[2794] A.-M. Wazwaz and F. B. Hanson. Moments of extinction for a singular perturbation problem in the diffusion process. SIAM Journal on Applied Mathematics, 51(1):233–265, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1991:EDE

[2795] G. Chen, S. A. Fulling, F. J. Narcowich, and S. Sun. Exponential decay of energy of evolution equations with locally distributed damping. SIAM Journal on Applied Mathematics, 51 (1):266–301, February 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bourlioux:1991:TNS

[2796] Anne Bourlioux, Andrew J. Majda, and Victor Roytburd. Theoretical and numerical structure for unstable one- dimensional detonations. SIAM Journal on Applied Mathematics, 51 (2):303–343, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Berestycki:1991:FEP

[2797] Henri Berestycki and Gregory I.

Sivashinsky. Flame extinction by periodic flow field. SIAM Journal on Applied Mathematics, 51(2):344–350, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Almgren:1991:HFA

[2798] Robert F. Almgren. High-frequency acoustic waves in a reacting gas. SIAM Journal on Applied Mathematics, 51 (2):351–373, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nair:1991:HFW

[2799] Balan Nair and Benjamin S. White. High-frequency wave propagation in random media — a unified approach. SIAM Journal on Applied Mathematics, 51(2):374–411, April 1991. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hesaaraki:1991:SSW

[2800] Mahmoud Hesaaraki. The structure of shock waves in magnetohydrodynamics for purely transverse magnetic fields. SIAM Journal on Applied Mathematics, 51(2):412–428, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cox:1991:ASF

[2801] Stephen M. Cox. Analysis of steady flow in a channel with one porous wall, or with accelerating walls. SIAM Journal on Applied Mathematics, 51 (2):429–438, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bosley:1991:SRV

[2802] D. L. Bosley and J. Kevorkian. Sustained resonance in very slowly varying oscillatory Hamiltonian systems. SIAM Journal on Applied Mathematics, 51 (2):439–471, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:1991:SFD

[2803] Donald S. Cohen and Andrew B.

White, Jr. Sharp fronts due to diffusion and viscoelastic relaxation in polymers. SIAM Journal on Applied Mathematics, 51(2):472–483, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bell:1991:SQL

[2804] Jonathan Bell, Avner Friedman, and Andrew A. Lacey. On solutions to a quasi-linear diffusion problem from the study of soft tissue. SIAM Journal on Applied Mathematics, 51(2):484–493, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pilant:1991:DCF

[2805] Michael Pilant and William Rundell.

Determining a coefficient in a firstorder hyperbolic equation. SIAM Journal on Applied Mathematics, 51 (2):494–506, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Szulga:1991:PPF

[2806] J. Szulga, W. A. Woyczyński, J. A.

Mann, Jr., and G. J. Tjatjopoulos. Poissonized partition function and large deviation techniques in analysis of one-dimensional monomolecular layers. SIAM Journal on Applied Mathematics, 51(2):507–524, April 1991.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bersani:1991:AAC

[2807] Alberto Bersani and Cecilia Sciarretta. Asymptotic analysis for a closed processor-sharing system with switching times: Normal usage. SIAM Journal on Applied Mathematics, 51 (2):525–541, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Newton:1991:AER

[2808] Nigel J. Newton. Asymptotically efficient Runge–Kutta methods for a class of Ito and Stratonovich equations. SIAM Journal on Applied Mathematics, 51(2):542–567, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dankel:1991:DIS

[2809] Thad Dankel, Jr. On the distribution of the integrated square of the Ornstein–Uhlenbeck process. SIAM Journal on Applied Mathematics, 51 (2):568–574, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zayed:1991:KST

[2810] Ahmed I. Zayed. On Kramer’s sampling theorem associated with general Sturm–Liouville problems and Lagrange interpolation. SIAM Journal on Applied Mathematics, 51(2):575–604, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Affouf:1991:NSR

[2811] Mahmoud Affouf and Russel E.

Caflisch. A numerical study of Riemann problem solutions and stability for a system of viscous conservation laws of mixed type. SIAM Journal on Applied Mathematics, 51(3):605–634, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dodd:1991:RPP

[2812] Nicholas Dodd. Reflective properties of parabolic approximations in shallow water wave propagation. SIAM Journal on Applied Mathematics, 51 (3):635–657, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sutrisno:1991:WSI

[2813] Sutrisno and D. R. Kassoy. Weak shocks initiated by power deposition on a spherical source boundary. SIAM Journal on Applied Mathematics, 51 (3):658–672, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Frankel:1991:QAO

[2814] Michael L. Frankel. Qualitative approximation of oscillatory flames in premixed gas combustion by a local equation of front dynamics. SIAM Journal on Applied Mathematics, 51 (3):673–683, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chadam:1991:SRF

[2815] J. Chadam, A. Peirce, and P. Ortoleva. Stability of reactive flows in porous media: Coupled porosity and viscosity changes. SIAM Journal on Applied Mathematics, 51(3):684–692, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Margolis:1991:NRQ

[2816] Stephen B. Margolis. New routes to quasi-periodic combustion of solids and high-density fluids near resonant Hopf bifurcation points. SIAM Journal on Applied Mathematics, 51(3):693–726, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [2903].

Bonilla:1991:SWS

[2817] Luis L. Bonilla. Solitary waves in semiconductors with finite geometry and the gunn effect. SIAM Journal on Applied Mathematics, 51(3):727–747, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kerkhoven:1991:ODC

[2818] Thomas Kerkhoven. On the onedimensional current driven semiconductor equations. SIAM Journal on Applied Mathematics, 51(3):748–774, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goodall:1991:EAC

[2819] Colin Goodall. Eigenshape analysis of a cut-grow mapping for triangles, and its application to phyllotaxis in plants. SIAM Journal on Applied Mathematics, 51(3):775–798, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Macken:1991:EWR

[2820] Catherine A. Macken, Patrick S. Hagan, and Alan S. Perelson. Evolutionary walks on rugged landscapes. SIAM Journal on Applied Mathematics, 51(3):799–827, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Simonian:1991:SAF

[2821] Alain Simonian. Stationary analysis of a fluid queue with input rate varying as an Ornstein–Uhlenbeck process. SIAM Journal on Applied Mathematics, 51(3):828–842, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Picard:1991:EEK

[2822] Jean Picard. Efficiency of the extended Kalman filter for nonlinear systems with small noise. SIAM Journal on Applied Mathematics, 51(3):843–885, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hoff:1991:FCD

[2823] David Hoff and Denis Serre. The failure of continuous dependence on initial data for the Navier–Stokes equations of compressible flow. SIAM Journal on Applied Mathematics, 51 (4):887–898, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Malkus:1991:ANP

[2824] David S. Malkus, John A. Nohel, and Bradley J. Plohr. Analysis of new phenomena in shear flow of non-Newtonian fluids. SIAM Journal on Applied Mathematics, 51(4):899–929, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mandal:1991:NWD

[2825] B. N. Mandal and S. Banerjea. A note on waves due to rolling of a partially immersed nearly vertical plate. SIAM Journal on Applied Mathematics, 51 (4):930–939, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moloney:1991:NST

[2826] T. P. Moloney and P. F. Hodnett. Note on the structure of a three-soliton solution of the Korteweg–de Vries equation. SIAM Journal on Applied Mathematics, 51(4):940–947, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jacovkis:1991:ODH

[2827] Pablo M. Jacovkis. One-dimensional hydrodynamic flow in complex networks and some generalizations. SIAM Journal on Applied Mathematics, 51 (4):948–966, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1991:AVE

[2828] Goong Chen, Matthew P. Coleman, and Jian-Xin Zhou. Analysis of vibration eigenfrequencies of a thin plate by the Keller–Rubinow wave method. I. Clamped boundary conditions with rectangular or circular geometry. SIAM Journal on Applied Mathematics, 51 (4):967–983, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Santosa:1991:DEM

[2829] Fadil Santosa and William W. Symes.

A dispersive effective medium for wave propagation in periodic composites. SIAM Journal on Applied Mathematics, 51(4):984–1005, August 1991. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maccluer:1991:DAM

[2830] C. R. Maccluer, C. J. Radcliffe, and Andrew J. Hull. Diagonalizing acoustic models. SIAM Journal on Applied Mathematics, 51(4):1006–1010, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Curtis:1991:DNM

[2831] Edward B. Curtis and James A. Morrow. The Dirichlet to Neumann map for a resistor network. SIAM Journal on Applied Mathematics, 51(4): 1011–1029, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Busenberg:1991:DCP

[2832] Stavros Busenberg, Kenneth Cooke, and Horst Thieme. Demographic change and persistence of HIV/AIDS in a heterogeneous population. SIAM Journal on Applied Mathematics, 51 (4):1030–1052, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klebanoff:1991:ADP

[2833] Aaron Klebanoff, Steven Minta, Alan Hastings, and Tim W. Clark. Agedependent predation model of blackfooted ferrets and prairie dogs. SIAM Journal on Applied Mathematics, 51 (4):1053–1073, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Volpert:1991:PVE

[2834] Vit. A. Vol′pert and Vl. A. Vol′pert. Propagation velocity estimation for condensed phase combustion. SIAM Journal on Applied Mathematics, 51 (4):1074–1089, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kirkkopru:1991:HSR

[2835] Kadir Kirkköprü and David R. Kassoy. High-speed reaction zone structure for variable mole chemistry. SIAM Journal on Applied Mathematics, 51 (4):1090–1118, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:1991:WAG

[2836] Charles Knessl. The WKB approximation to the G/M/m queue. SIAM Journal on Applied Mathematics, 51 (4):1119–1133, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gaver:1991:HTA

[2837] D. P. Gaver and J. A. Morrison. Heavy-traffic analysis of multitype queueing under probabilistically loadpreferential service order. SIAM Journal on Applied Mathematics, 51(4): 1134–1149, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bourland:1991:AMP

[2838] F. Jay Bourland, Richard Haberman, and William L. Kath. Averaging methods for the phase shift of arbitrarily perturbed strongly nonlinear oscillators with an application to capture. SIAM Journal on Applied Mathematics, 51(4):1150–1167, August 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lim:1991:SCB

[2839] Chee-Seng Lim. Semi-Cherenkov and bi-Cherenkov problems. SIAM Journal on Applied Mathematics, 51(5): 1169–1193, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Luke:1991:NPL

[2840] Jon C. Luke. A note on particlelike solutions for a nonlinear complexvalued Klein–Gordon equation. SIAM Journal on Applied Mathematics, 51 (5):1194–1204, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abeyaratne:1991:IVS

[2841] Rohan Abeyaratne and James K.

Knowles. Implications of viscosity and strain-gradient effects for the kinetics of propagating phase boundaries in solids. SIAM Journal on Applied Mathematics, 51(5):1205–1221, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Murota:1991:CIS

[2842] Kazuo Murota and Kiyohiro Ikeda.

Critical imperfection of symmetric structures. SIAM Journal on Applied Mathematics, 51(5):1222–1254, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brewster:1991:TPC

[2843] M. E. Brewster. Thin plates and compressive membrane solutions I: Global breakdown. SIAM Journal on Applied Mathematics, 51(5):1255–1283, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Silverman:1991:AAP

[2844] I. Silverman, J. B. Greenberg, and Y. Tambour. Asymptotic analysis of a premixed polydisperse spray flame. SIAM Journal on Applied Mathematics, 51(5):1284–1303, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Henderson:1991:TTR

[2845] K. L. Henderson and J. W. Dold. Transition from thermal runaway to propagating flames. SIAM Journal on Applied Mathematics, 51(5):1304–1314, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lee:1991:SSF

[2846] C. J. Lee and J. Buckmaster. The structure and stability of flame balls: A near- equidiffusional flame analysis. SIAM Journal on Applied Mathematics, 51(5):1315–1326, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fang:1991:MMO

[2847] W. Fang, M. Shillor, E. Stahel, E. Epstein, C. Ly, J. McNiel, and E. Zaron. A mathematical model for outgassing and contamination. SIAM Journal on Applied Mathematics, 51(5):1327–1355, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bechtold:1991:NHS

[2848] John K. Bechtold and Stephen B. Margolis. Nonlinear hydrodynamic stability and spinning deflagration of liquid propellants. SIAM Journal on Applied Mathematics, 51(5):1356–1379, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Forbes:1991:FTO

[2849] Lawrence K. Forbes. Forced transverse oscillations in a simple springmass system. SIAM Journal on Applied Mathematics, 51(5):1380–1396, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kopell:1991:COF

[2850] N. Kopell, G. B. Ermentrout, and T. L. Williams. On chains of oscillators forced at one end. SIAM Journal on Applied Mathematics, 51 (5):1397–1417, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Terman:1991:CSA

[2851] David Terman. Chaotic spikes arising from a model of bursting in excitable membranes. SIAM Journal on Applied Mathematics, 51(5):1418–1450, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:1991:CVG

[2852] H. L. Smith, Betty Tang, and Paul Waltman. Competition in an n-vessel gradostat. SIAM Journal on Applied Mathematics, 51(5):1451–1471, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Holden:1991:SPS

[2853] Helge Holden and Nils Henrik Risebro. Stochastic properties of the scalar Buckley–Leverett equation. SIAM Journal on Applied Mathematics, 51 (5):1472–1488, October 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cole:1991:GPD

[2854] S. L. Cole and J. W. Wilder. Gas phase decomposition by the Lindemann mechanism. SIAM Journal on Applied Mathematics, 51(6):1489–1497, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hunter:1991:DDF

[2855] John K. Hunter and Ralph Saxton. Dynamics of director fields. SIAM Journal on Applied Mathematics, 51(6): 1498–1521, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cox:1991:SAD

[2856] Robert W. Cox. Stress-assisted diffusion: A free boundary problem. SIAM Journal on Applied Mathematics, 51(6):1522–1537, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mauri:1991:TDD

[2857] Roberto Mauri and Shimon Haber.

Time-dependent dispersion of small particles in rectangular conduits. SIAM Journal on Applied Mathematics, 51 (6):1538–1555, December 1991. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gautesen:1991:SSE

[2858] A. K. Gautesen, R. W. Ziolkowski, and R. R. McLeod. A solution to the scattering of electromagnetic waves from a dielectric semi-cylinder. SIAM Journal on Applied Mathematics, 51(6): 1556–1567, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Budreck:1991:ECR

[2859] David E. Budreck and James H. Rose.

Elastodynamic completeness relations for scattered wavefields. SIAM Journal on Applied Mathematics, 51(6): 1568–1584, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1991:EANa

[2860] Richard E. Meyer. On exponential asymptotics for nonseparable wave equations I: Complex geometrical optics and connection. SIAM Journal on Applied Mathematics, 51(6): 1585–1601, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1991:EANb

[2861] Richard E. Meyer. On exponential asymptotics for nonseparable wave equations. II. EBK quantization. SIAM Journal on Applied Mathematics, 51 (6):1602–1615, December 1991. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pistiner:1991:NSN

[2862] Arieh Pistiner, Michael Shapiro, and Hillel Rubin. A new solution for the nonlinear diffusion-convection equation. SIAM Journal on Applied Mathematics, 51(6):1616–1621, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Krol:1991:AMN

[2863] M. S. Krol. On the averaging method in nearly time-periodic advectiondiffusion problems. SIAM Journal on Applied Mathematics, 51(6):1622–1637, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Haberman:1991:SFM

[2864] Richard Haberman. Standard form and a method of averaging for strongly nonlinear oscillatory dispersive traveling waves. SIAM Journal on Applied Mathematics, 51(6):1638–1652, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klibanov:1991:CQR

[2865] Michael V. Klibanov and Fadil Santosa. A computational quasi-reversibility method for Cauchy problems for Laplace’s equation. SIAM Journal on Applied Mathematics, 51(6):1653–1675, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Babovsky:1991:ISM

[2866] Hans Babovsky. Identification of scattering media from reflected flows. SIAM Journal on Applied Mathematics, 51(6):1676–1704, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Isaacson:1991:EMP

[2867] David Isaacson and Margaret Cheney.

Effects of measurement precision and finite numbers of electrodes on linear impedance imaging algorithms. SIAM Journal on Applied Mathematics, 51 (6):1705–1731, December 1991. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Simonian:1991:TSD

[2868] Alain Simonian and Jorma Virtamo.

Transient and stationary distributions for fluid queues and input processes with a density. SIAM Journal on Applied Mathematics, 51(6):1732–1739, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:1991:HTAb

[2869] C. Knessl and J. A. Morrison. Heavy traffic analysis of the sojourn time in tandem queues with overtaking. SIAM Journal on Applied Mathematics, 51(6):1740–1763, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dougherty:1991:SAR

[2870] Edward R. Dougherty and Eugene J.

Kraus. Shape analysis and reduction of the morphological basis for digital moving-average filters. SIAM Journal on Applied Mathematics, 51(6): 1764–1781, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Banon:1991:MRT

[2871] Gerald Jean Francis Banon and Junior Barrera. Minimal representations for translation-invariant set mappings by mathematical morphology. SIAM Journal on Applied Mathematics, 51(6):1782–1798, December 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Feng:1992:OBM

[2872] James Q. Feng. The oscillations of a bubble moving in an inviscid fluid. SIAM Journal on Applied Mathematics, 52(1):1–14, February 1992. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fowler:1992:FFR

[2873] A. C. Fowler and P. E. Lisseter. Flooding and flow reversal in annular twophase flows. SIAM Journal on Applied Mathematics, 52(1):15–33, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hou:1992:HLT

[2874] Thomas Y. Hou and Xue Xin. Homogenization of linear transport equations with oscillatory vector fields. SIAM Journal on Applied Mathematics, 52 (1):34–45, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Friedman:1992:HAL

[2875] Avner Friedman and Bei Hu. Homogenization approach to light scattering from polymer-dispersed liquid crystal films. SIAM Journal on Applied Mathematics, 52(1):46–64, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dahl:1992:MCT

[2876] Olav Dahl, Thormod Johansen, Aslak Tveito, and Ragnar Winther. Multicomponent chromatography in a two phase environment. SIAM Journal on Applied Mathematics, 52(1):65–104, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Spiga:1992:SWS

[2877] G. Spiga and S. Oggioni. On shock wave solutions in extended discrete kinetic theory. SIAM Journal on Applied Mathematics, 52(1):105–111, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ward:1992:SPF

[2878] Michael J. Ward. Singular perturbations and a free boundary problem in the modeling of field effect transistors. SIAM Journal on Applied Mathematics, 52(1):112–139, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ram:1992:PPR

[2879] Yitshak M. Ram and James Caldwell.

Physical parameters reconstruction of a free–free mass-spring system from its spectra. SIAM Journal on Applied Mathematics, 52(1):140–152, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Adler:1992:DCS

[2880] J. Adler and D. M. Herbert. Diffusion controlled smoulder propagation parallel to a plane surface. SIAM Journal on Applied Mathematics, 52 (1):153–162, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dawson:1992:DRC

[2881] Donald A. Dawson and Klaus Fleischmann. Diffusion and reaction caused by point catalysts. SIAM Journal on Applied Mathematics, 52(1):163–180, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Farr:1992:RCW

[2882] W. W. Farr and M. Golubitsky. Rotating chemical waves in the Gray– Scott model. SIAM Journal on Applied Mathematics, 52(1):181–221, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wolkowicz:1992:GDM

[2883] Gail S. K. Wolkowicz and Zhiqi Lu. Global dynamics of a mathematical model of competition in the chemostat: General response functions and differential death rates. SIAM Journal on Applied Mathematics, 52(1): 222–233, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mankus:1992:AAI

[2884] Margo L. Mankus and Charles Tier. Asymptotic analysis of an integrated digital network. SIAM Journal on Applied Mathematics, 52(1):234–269, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Solo:1992:IRF

[2885] Victor Solo. Intrinsic random functions and the paradox of 1/f noise. SIAM Journal on Applied Mathematics, 52 (1):270–291, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hall:1992:FCV

[2886] Peter Hall and Inge Koch. On the feasibility of cross-validation in image analysis. SIAM Journal on Applied Mathematics, 52(1):292–313, February 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Caviglia:1992:RAW

[2887] Giacomo Caviglia and Angelo Morro.

Ray analysis of wave propagation in nonhomogeneous viscoelastic fluids. SIAM Journal on Applied Mathematics, 52(2):315–326, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tew:1992:MSE

[2888] R. H. Tew and J. R. Ockendon. Multiple scales expansions near critical rays. SIAM Journal on Applied Mathematics, 52(2):327–336, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hillion:1992:NDP

[2889] Pierre Hillion. A note on the derivation of paraxial equation in nonhomogeneous media. SIAM Journal on Applied Mathematics, 52(2):337–346, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lindemann:1992:SHP

[2890] Ina Lindemann. Stability of hyperbolic partial differential equations with random loads. SIAM Journal on Applied Mathematics, 52(2):347–367, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Volpert:1992:TDC

[2891] Vit. A. Vol′pert, Vl. A. Vol′pert, S. P. Davtyan, I. N. Megrabova, and N. F. Surkov. Two-dimensional combustion modes in condensed flow. SIAM Journal on Applied Mathematics, 52 (2):368–383, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garbey:1992:QPW

[2892] Marc Garbey, Hans G. Kaper, Gary K. Leaf, and Bernard J. Matkowsky. Quasi-periodic waves and the transfer of stability in condensed-phase surface combustion. SIAM Journal on Applied Mathematics, 52(2):384–395, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bayliss:1992:NDC

[2893] Alvin Bayliss and Bernard J. Matkowsky. Nonlinear dynamics of cellular flames. SIAM Journal on Applied Mathematics, 52(2):396–415, April 1992. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sinay:1992:AAD

[2894] L. Sinay and F. A. Williams. An analytical approach to the description of nonadiabatic cellular flames near extinction. SIAM Journal on Applied Mathematics, 52(2):416–427, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Giordana:1992:NSI

[2895] Claudia Giordana, Marina Mochi, and Francesco Zirilli. The numerical solution of an inverse problem for a class of one-dimensional diffusion equations with piecewise constant coefficients. SIAM Journal on Applied Mathematics, 52(2):428–441, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dobson:1992:CRM

[2896] David C. Dobson. Convergence of a reconstruction method for the inverse conductivity problem. SIAM Journal on Applied Mathematics, 52 (2):442–458, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Faridani:1992:LT

[2897] Adel Faridani, Erik L. Ritman, and Kennan T. Smith. Local tomography. SIAM Journal on Applied Mathematics, 52(2):459–484, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Szmolyan:1992:ASP

[2898] P. Szmolyan. Analysis of a singularly perturbed traveling wave problem. SIAM Journal on Applied Mathematics, 52(2):485–493, April 1992.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bosley:1992:AIT

[2899] D. L. Bosley and J. Kevorkian. Adiabatic invariance and transient resonance in very slowly varying oscillatory Hamiltonian systems. SIAM Journal on Applied Mathematics, 52 (2):494–527, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hsu:1992:AMT

[2900] S. B. Hsu and P. Waltman. Analysis of a model of two competitors in a chemostat with an external inhibitor. SIAM Journal on Applied Mathematics, 52 (2):528–540, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Simon:1992:RNS

[2901] Carl P. Simon and John A. Jacquez. Reproduction numbers and the stability of equilibria of Si models for heterogeneous populations. SIAM Journal on Applied Mathematics, 52(2):541–576, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Donoho:1992:SRL

[2902] D. L. Donoho and B. F. Logan. Signal recovery and the large sieve. SIAM Journal on Applied Mathematics, 52 (2):577–591, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Margolis:1992:ENR

[2903] Stephen B. Margolis. Erratum: “New Routes to Quasi-Periodic Combustion of Solids and High-Density Fluids Near Resonant Hopf Bifurcation Points” [SIAM J. Appl. Math. 51 (1991), no. 3, 693–726, MR 92d:80005]. SIAM Journal on Applied Mathematics, 52(2):592, April 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [2816].

Novick-Cohen:1992:MPA

[2904] Amy Novick-Cohen. On a minimization problem arising in wetting. SIAM Journal on Applied Mathematics, 52 (3):593–613, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:1992:SSS

[2905] Sheng Wang and Yisong Yang. Symmetric superconducting states in thin films. SIAM Journal on Applied Mathematics, 52(3):614–629, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Britton:1992:NMM

[2906] N. F. Britton. A new method for mass transport problems with an application to haemofilters. SIAM Journal on Applied Mathematics, 52(3):630–650, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barshinger:1992:EFA

[2907] Richard Barshinger. The electrostatic field about a thin oblate dielectric body of revolution. SIAM Journal on Applied Mathematics, 52(3):651–675, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lim:1992:RWS

[2908] C. C. Lim, J. M. Pimbley, C. Schmeiser, and D. W. Schwendeman. Rotating waves for semiconductor inverter rings. SIAM Journal on Applied Mathematics, 52(3):676–690, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pimbley:1992:CCD

[2909] J. M. Pimbley and J. D. Cole. Charge carrier dynamics in alpha particle tracks. SIAM Journal on Applied Mathematics, 52(3):691–698, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fang:1992:IPM

[2910] Wei Fu Fang and Ellis Cumberbatch. Inverse problems for metal oxide semiconductor field-effect transistor contact resistivity. SIAM Journal on Applied Mathematics, 52(3):699–709, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maass:1992:IRT

[2911] Peter Maass. The interior Radon transform. SIAM Journal on Applied Mathematics, 52(3):710–724, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zayed:1992:IEP

[2912] E. M. E. Zayed. An inverse eigenvalue problem for an arbitrary, multiply connected, bounded domain in R^{3} with impedance boundary conditions. SIAM Journal on Applied Mathematics, 52(3):725–729, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Meyer:1992:EAN

[2913] R. E. Meyer and M. C. Shen. On exponential asymptotics for nonseparable wave equations. III. Approximate spectral bands of periodic potentials on strips. SIAM Journal on Applied Mathematics, 52(3):730–745, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Forest:1992:FPL

[2914] M. Gregory Forest, S.-P. Sheu, and Amarendra Sinha. Frequency and phase locking of spatially periodic perturbed Sine-Gordon breather trains. SIAM Journal on Applied Mathematics, 52(3):746–761, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Neto:1992:DDD

[2915] F. Moura Neto. Defect dynamics in a dissipative system. SIAM Journal on Applied Mathematics, 52(3):762–791, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Aston:1992:BRW

[2916] P. J. Aston, A. Spence, and W. Wu. Bifurcation to rotating waves in equations with O(2)-symmetry. SIAM Journal on Applied Mathematics, 52(3):792–809, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Curtis:1992:NSA

[2917] W. D. Curtis, R. O. Fox, and K. Halasi.

Numerical stability analysis of a class of functional differential equations. SIAM Journal on Applied Mathematics, 52 (3):810–834, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huang:1992:SBM

[2918] Wen Zhang Huang, Kenneth L. Cooke, and Carlos Castillo-Chavez. Stability and bifurcation for a multiplegroup model for the dynamics of HIV/ AIDS transmission. SIAM Journal on Applied Mathematics, 52(3):835–854, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Aiello:1992:AMR

[2919] Walter G. Aiello, H. I. Freedman, and J. Wu. Analysis of a model representing stage-structured population growth with state-dependent time delay. SIAM Journal on Applied Mathematics, 52 (3):855–869, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fuller:1992:DQE

[2920] William Roger Fuller. A discussion of the quasi-Euler–Lagrange equation. SIAM Journal on Applied Mathematics, 52(3):870–882, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:1992:PSQ

[2921] Charles Knessl and Charles Tier. A processor-shared queue that models switching times: Normal usage. SIAM Journal on Applied Mathematics, 52 (3):883–899, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hajela:1992:FTN

[2922] D. Hajela. On faster than Nyquist signaling: Further estimations on the minimum distance. SIAM Journal on Applied Mathematics, 52(3):900–907, June 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bloch:1992:DRW

[2923] A. M. Bloch and Y. Kodama. Dispersive regularization of the Whitham equation for the Toda lattice. SIAM Journal on Applied Mathematics, 52 (4):909–928, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Coffey:1992:PSA

[2924] Mark W. Coffey. Phase space analysis for the direct algebraic method for nonlinear evolution and wave equations. SIAM Journal on Applied Mathematics, 52(4):929–945, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chikwendu:1992:MSS

[2925] S. C. Chikwendu and C. V. Easwaran.

Multiple-scale solution of initialboundary value problems for weakly nonlinear wave equations on the semiinfinite line. SIAM Journal on Applied Mathematics, 52(4):946–958, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

E:1992:HSC

[2926] Weinan E. Homogenization of scalar conservation laws with oscillatory forcing terms. SIAM Journal on Applied Mathematics, 52(4):959–972, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Delaurens:1992:DPM

[2927] F. Delaurens and F. J. Mustieles. A deterministic particle method for solving kinetic transport equations: The semiconductor Boltzmann equation case. SIAM Journal on Applied Mathematics, 52(4):973–988, August 1992. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bennaton:1992:GSS

[2928] J. Freitas Bennaton. General solutions for semilinear elliptic equations with electric-thermal coupling. SIAM Journal on Applied Mathematics, 52 (4):989–997, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fowler:1992:TSC

[2929] A. C. Fowler, I. Frigaard, and S. D. Howison. Temperature surges in current-limiting circuit devices. SIAM Journal on Applied Mathematics, 52 (4):998–1011, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Paulson:1992:EME

[2930] Kevin Paulson, William Breckon, and Michael Pidcock. Electrode modelling in electrical impedance tomography. SIAM Journal on Applied Mathematics, 52(4):1012–1022, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Somersalo:1992:EUE

[2931] Erkki Somersalo, Margaret Cheney, and David Isaacson. Existence and uniqueness for electrode models for electric current computed tomography. SIAM Journal on Applied Mathematics, 52(4):1023–1040, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Benedetto:1992:VLV

[2932] D. Benedetto and M. Pulvirenti. From vortex layers to vortex sheets. SIAM Journal on Applied Mathematics, 52 (4):1041–1056, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Borgers:1992:DLF

[2933] Christoph Borgers, Claude Greengard, and Enrique Thomann. The diffusion limit of free molecular flow in thin plane channels. SIAM Journal on Applied Mathematics, 52(4):1057–1075, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:1992:SSC

[2934] Feng Wang, Carl L. Gardner, and David G. Schaeffer. Steady-state computations of granular flow in an axisymmetric hopper. SIAM Journal on Applied Mathematics, 52(4):1076–1088, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fogelson:1992:CMP

[2935] Aaron L. Fogelson. Continuum models of platelet aggregation: Formulation and mechanical properties. SIAM Journal on Applied Mathematics, 52 (4):1089–1110, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

DiCarlo:1992:REA

[2936] Antonio Di Carlo, Morton E. Gurtin, and Paolo Podio-Guidugli. A regularized equation for anisotropic motionby-curvature. SIAM Journal on Applied Mathematics, 52(4):1111–1119, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Healey:1992:LRO

[2937] Timothy J. Healey. Large rotatory oscillations of transversely isotropic rods: Spatio-temporal symmetry-breaking bifurcation. SIAM Journal on Applied Mathematics, 52(4):1120–1135, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Retzloff:1992:MMS

[2938] David G. Retzloff, P. C.-H. Chan, Carmen Chicone, and Ira Papick. Maximal multiplicity for sequential bifurcations of a first-order reaction occurring in continuous stirred tank reactors coupled in series. SIAM Journal on Applied Mathematics, 52(4): 1136–1147, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ghil:1992:NHP

[2939] M. Ghil and G. Wolansky. Non-Hamiltonian perturbations of integrable systems and resonance trapping. SIAM Journal on Applied Mathematics, 52(4):1148–1171, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Perko:1992:GAB

[2940] L. M. Perko. A global analysis of the Bogdanov–Takens system. SIAM Journal on Applied Mathematics, 52 (4):1172–1192, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Faridani:1992:ELT

[2941] Adel Faridani, Erik L. Ritman, and Kennan T. Smith. Examples of local tomography. SIAM Journal on Applied Mathematics, 52(4):1193–1198, August 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Embid:1992:SAE

[2942] Pedro Embid, John Hunter, and Andrew Majda. Simplified asymptotic equations for the transition to detonation in reactive granular materials. SIAM Journal on Applied Mathematics, 52(5):1199–1237, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mancebo:1992:WNT

[2943] Francisco J. Mancebo and José M.

Vega. Weakly nonuniform thermal effects in a porous catalyst: Asymptotic models and local nonlinear stability of the steady states. SIAM Journal on Applied Mathematics, 52 (5):1238–1259, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Isaacson:1992:NRS

[2944] Eli Isaacson and Blake Temple. Nonlinear resonance in systems of conservation laws. SIAM Journal on Applied Mathematics, 52(5):1260–1278, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Merchant:1992:PIR

[2945] G. J. Merchant, R. J. Braun, K. Brattkus, and S. H. Davis. Pulsatile instability in rapid directional solidification: Strongly nonlinear analysis. SIAM Journal on Applied Mathematics, 52 (5):1279–1302, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brattkus:1992:NSU

[2946] K. Brattkus and D. I. Meiron. Numerical simulations of unsteady crystal growth. SIAM Journal on Applied Mathematics, 52(5):1303–1320, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Castaneda:1992:NVP

[2947] P. Ponte Castañeda. A new variational principle and its application to nonlinear heterogeneous systems. SIAM Journal on Applied Mathematics, 52 (5):1321–1341, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ilic:1992:CMB

[2948] M. Ilic. A class of moving boundary problems arising in drying processes. SIAM Journal on Applied Mathematics, 52(5):1342–1359, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hansen:1992:DL

[2949] Erik B. Hansen. On drying of laundry. SIAM Journal on Applied Mathematics, 52(5):1360–1369, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Keener:1992:CS

[2950] James P. Keener. The core of the spiral.

SIAM Journal on Applied Mathematics, 52(5):1370–1390, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barcilon:1992:IFTa

[2951] Victor Barcilon, D.-P. Chen, and R. S.

Eisenberg. Ion flow through narrow membrane channels: Part I. SIAM Journal on Applied Mathematics, 52 (5):1391–1404, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barcilon:1992:IFTb

[2952] Victor Barcilon, D.-P. Chen, and R. S.

Eisenberg. Ion flow through narrow membrane channels: Part II. SIAM Journal on Applied Mathematics, 52 (5):1405–1425, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ford:1992:RBC

[2953] Roseanne M. Ford and Peter T. Cummings. On the relationship between cell balance equations for chemotactic cell populations. SIAM Journal on Applied Mathematics, 52(5):1426–1441, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See errata [3043].

Jager:1992:DDP

[2954] Eva Jäger and Lee A. Segel. On the distribution of dominance in populations of social organisms. SIAM Journal on Applied Mathematics, 52 (5):1442–1468, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stensby:1992:LDP

[2955] John Stensby. Lock detection in phaselocked loops. SIAM Journal on Applied Mathematics, 52(5):1469–1475, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gaunersdorfer:1992:TAH

[2956] Andrea Gaunersdorfer. Time averages for heteroclinic attractors. SIAM Journal on Applied Mathematics, 52 (5):1476–1489, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:1992:XRT

[2957] Ling Wang. The X-ray transform and its inversion for the series expansion basis functions in three-dimensional tomography. SIAM Journal on Applied Mathematics, 52(5):1490–1499, October 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Easwaran:1992:TTD

[2958] C. V. Easwaran and S. L. Kokal. A theory for transient discharge in tubes with applications to a new type of viscometer. SIAM Journal on Applied Mathematics, 52(6):1501–1508, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Young:1992:FEV

[2959] G. W. Young, S. I. Hariharan, and R. Carnahan. Flow effects in a vertical CVD reactor. SIAM Journal on Applied Mathematics, 52(6):1509–1532, December 1992. CODEN SMJMAP.

ISSN 0036-1399 (print), 1095-712X (electronic).

Kim:1992:NSM

[2960] Hyun S. Kim and Andrea Prosperetti.

Numerical simulation of the motion of rigid spheres in potential flow. SIAM Journal on Applied Mathematics, 52(6):1533–1562, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jacobson:1992:DC

[2961] Dmitry Jacobson and John H. Maddocks. On the dynamics of chains. SIAM Journal on Applied Mathematics, 52(6):1563–1583, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hu:1992:RCP

[2962] Ji Shan Hu and Martin Kruskal. Reflection coefficient problems for weakly nonlinear wave equations. SIAM Journal on Applied Mathematics, 52(6): 1584–1596, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Angell:1992:CBC

[2963] T. S. Angell and A. Kirsch. The conductive boundary condition for Maxwell’s equations. SIAM Journal on Applied Mathematics, 52(6): 1597–1610, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fang:1992:ISD

[2964] Wei Fu Fang and Kazufumi Ito. Identifiability of semiconductor defects from LBIC images. SIAM Journal on Applied Mathematics, 52(6):1611–1626, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pernarowski:1992:PTM

[2965] M. Pernarowski, R. M. Miura, and J. Kevorkian. Perturbation techniques for models of bursting electrical activity in pancreatic β-cells. SIAM Journal on Applied Mathematics, 52(6): 1627–1650, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baer:1992:SHB

[2966] S. M. Baer and T. Erneux. Singular Hopf bifurcation to relaxation oscillations II. SIAM Journal on Applied Mathematics, 52(6):1651–1664, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ermentrout:1992:SPS

[2967] G. Bard Ermentrout. Stable periodic solutions to discrete and continuum arrays of weakly coupled nonlinear oscillators. SIAM Journal on Applied Mathematics, 52(6):1665–1687, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Muratori:1992:LHF

[2968] Simona Muratori and Sergio Rinaldi. Low- and high-frequency oscillations in three-dimensional food chain systems. SIAM Journal on Applied Mathematics, 52(6):1688–1706, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dohtani:1992:OCH

[2969] Akitaka Dohtani. Occurrence of chaos in higher-dimensional discretetime systems. SIAM Journal on Applied Mathematics, 52(6):1707–1721, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kalbaugh:1992:OSA

[2970] D. V. Kalbaugh. Optimal search among false contacts. SIAM Journal on Applied Mathematics, 52(6): 1722–1750, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Browne:1992:GIS

[2971] Sid Browne, E. G. Coffman, Jr., E. N.

Gilbert, and Paul E. Wright. The gated infinite-server queue: Uniform service times. SIAM Journal on Applied Mathematics, 52(6):1751–1762, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pevzner:1992:MAC

[2972] Pavel A. Pevzner. Multiple alignment, communication cost, and graph matching. SIAM Journal on Applied Mathematics, 52(6):1763–1779, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Murota:1992:RIS

[2973] Kazuo Murota and Kiyohiro Ikeda.

On random imperfections for structures of regular-polygonal symmetry. SIAM Journal on Applied Mathematics, 52(6):1780–1803, December 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hoff:1993:SCP

[2974] David Hoff and M. Khodja. Stability of coexisting phases for compressible van der Waals fluids. SIAM Journal on Applied Mathematics, 53(1):1–14, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reddy:1993:POS

[2975] Satish C. Reddy, Peter J. Schmid, and Dan S. Henningson. Pseudospectra of the Orr–Sommerfeld operator. SIAM Journal on Applied Mathematics, 53 (1):15–47, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Guttmann:1993:SPF

[2976] A. J. Guttmann and C. J. Thompson. Subsonic potential flow and the transonic controversy. SIAM Journal on Applied Mathematics, 53(1):48–59, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Power:1993:SKI

[2977] Henry Power and B. Febres de Power. Second-kind integral equation formulation for the slow motion of a particle of arbitrary shape near a plane wall in a viscous fluid. SIAM Journal on Applied Mathematics, 53(1):60–70, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhao:1993:CRS

[2978] Xiao Hua Zhao, Keng Huat Kwek, Ji Bin Li, and Ke Lei Huang. Chaotic and resonant streamlines in the ABC flow. SIAM Journal on Applied Mathematics, 53(1):71–77, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Impey:1993:BAC

[2979] M. D. Impey, D. S. Riley, and A. A.

Wheeler. Bifurcation analysis of cellular interfaces in unidirectional solidification of a dilute binary mixture. SIAM Journal on Applied Mathematics, 53 (1):78–95, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nesi:1993:CPP

[2980] V. Nesi. On the G-closure in the polycrystalline problem. SIAM Journal on Applied Mathematics, 53(1):96–127, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shkadinsky:1993:TFT

[2981] K. G. Shkadinsky, G. V. Shkadinskaya, B. J. Matkowsky, and V. A. Volpert. Two-front traveling waves in filtration combustion. SIAM Journal on Applied Mathematics, 53(1):128–140, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dassios:1993:DTT

[2982] George Dassios and Zafiria Rigou. On the density of traction traces in scattering of elastic waves. SIAM Journal on Applied Mathematics, 53(1): 141–153, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hohlfeld:1993:SCI

[2983] R. G. Hohlfeld, J. I. F. King, T. W. Drueding, and G. v. H. Sandri. Solution of convolution integral equations by the method of differential inversion. SIAM Journal on Applied Mathematics, 53(1):154–167, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:1993:MMS

[2984] Xinfu Chen, Avner Friedman, and Li Shang Jiang. Mathematical modeling of semiconductor lasers. SIAM Journal on Applied Mathematics, 53 (1):168–186, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Busenberg:1993:MAL

[2985] Stavros Busenberg, Wei Fu Fang, and Kazufumi Ito. Modeling and analysis of laser-beam-induced current images in semiconductors. SIAM Journal on Applied Mathematics, 53(1): 187–204, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carroll:1993:OOC

[2986] Joseph E. Carroll and Roland H. Lamberson. The owl’s odyssey. A continuous model for the dispersal of territorial species. SIAM Journal on Applied Mathematics, 53(1):205–218, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cantrell:1993:SPI

[2987] Robert Stephen Cantrell and Chris Cosner. Should a park be an island? SIAM Journal on Applied Mathematics, 53(1):219–252, February 1993.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vanvreeswijk:1993:SSF

[2988] Carl Vanvreeswijk and L. F. Abbott.

Self-sustained firing in populations of integrate-and-fire neurons. SIAM Journal on Applied Mathematics, 53(1): 253–264, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vakakis:1993:MLC

[2989] Alexander F. Vakakis and Cetin Cetinkaya. Mode localization in a class of multidegree-of-freedom nonlinear systems with cyclic symmetry. SIAM Journal on Applied Mathematics, 53 (1):265–282, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Loparo:1993:LER

[2990] Kenneth A. Loparo and Xiangbo Feng.

Lyapunov exponent and rotation number of two-dimensional linear stochastic systems with telegraphic noise. SIAM Journal on Applied Mathematics, 53 (1):283–300, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Drmota:1993:ELD

[2991] M. Drmota and U. Schmid. Exponential limiting distributions in queueing systems with deadlines. SIAM Journal on Applied Mathematics, 53 (1):301–318, February 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Xin:1993:RFL

[2992] J. Xin, A. Peirce, J. Chadam, and P. Ortoleva. Reactive flows in layered porous media. II. The shape stability of the reaction interface. SIAM Journal on Applied Mathematics, 53 (2):319–339, April 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosen:1993:SNS

[2993] Dan Rosen and Donald E. Cormack. Singular and near singular integrals in the BEM: A global approach. SIAM Journal on Applied Mathematics, 53 (2):340–357, April 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Natterer:1993:SFB

[2994] F. Natterer. Sampling in fan beam tomography. SIAM Journal on Applied Mathematics, 53(2):358–380, April 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Steinberg:1993:LRT

[2995] B. Z. Steinberg and Ehud Heyman. The local radon transform and variable scale resolution with application to time-dependent wave radiation. SIAM Journal on Applied Mathematics, 53 (2):381–400, April 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kozlov:1993:EDP

[2996] Serguei M. Kozlov and Andrei L. Piatnitski. Effective diffusion for a parabolic operator with periodic potential. SIAM Journal on Applied Mathematics, 53(2):401–418, April 1993.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baker:1993:RNS

[2997] Daniel R. Baker. Reducing nonlinear systems of transport equations to laplace’s equation. SIAM Journal on Applied Mathematics, 53(2):419–439, April 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moses:1993:DRE

[2998] Harry E. Moses and Reese T. Prosser.

Decay of the relative error in the formation of acoustic bullets. SIAM Journal on Applied Mathematics, 53 (2):440–446, April 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Howison:1993:FTL

[2999] S. D. Howison and R. J. Lawrence. Fluorescent transfer of light in dyed materials. SIAM Journal on Applied Mathematics, 53(2):447–458, April 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gaver:1993:SAM

[3000] Donald P. Gaver, John A. Morrison, and Rogerio Silveira. Service-adaptive multitype repairman problems. SIAM Journal on Applied Mathematics, 53 (2):459–470, April 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:1993:DAH

[3001] J. A. Morrison. Diffusion approximation for head-of-the-line processor sharing for two parallel queues. SIAM Journal on Applied Mathematics, 53 (2):471–490, April 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tan:1993:FCP

[3002] Xiaoming Tan and Charles Knessl. A finite capacity PS queue which models switching times. SIAM Journal on Applied Mathematics, 53(2):491–554, April 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Xie:1993:TBE

[3003] Shisheng Xie and Charles Knessl. On the transient behavior of the Erlang loss model: Heavy usage asymptotics. SIAM Journal on Applied Mathematics, 53(2):555–599, April 1993. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Weisenborn:1993:AAO

[3004] A. J. Weisenborn and B. I. M. ten Bosch. Analytical approach to the Oseen drag on a sphere at infinite Reynolds number. SIAM Journal on Applied Mathematics, 53(3):601–620, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

An:1993:LHE

[3005] Lian Jun An. Loss of hyperbolicity in elastic-plastic material at finite strains. SIAM Journal on Applied Mathematics, 53(3):621–654, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kriegsmann:1993:MHD

[3006] Gregory A. Kriegsmann. Microwave heating of dispersive media. SIAM Journal on Applied Mathematics, 53 (3):655–669, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smyshlyaev:1993:HFD

[3007] Valery P. Smyshlyaev. The highfrequency diffraction of electromagnetic waves by cones of arbitrary cross sections. SIAM Journal on Applied Mathematics, 53(3):670–688, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Du:1993:MAP

[3008] Qiang Du, Max D. Gunzburger, and Janet S. Peterson. Modeling and analysis of a periodic Ginzburg–Landau model for Type-II superconductors. SIAM Journal on Applied Mathematics, 53(3):689–717, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Budd:1993:BPD

[3009] Chris Budd, Bill Dold, and Andrew Stuart. Blowup in a partial differential equation with conserved first integral. SIAM Journal on Applied Mathematics, 53(3):718–742, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bai:1993:NCH

[3010] Fenghsan Bai, Alastair Spence, and Andrew M. Stuart. The numerical computation of heteroclinic connections in systems of gradient partial differential equations. SIAM Journal on Applied Mathematics, 53(3):743–769, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ward:1993:SLP

[3011] Michael J. Ward and Joseph B. Keller. Strong localized perturbations of eigenvalue problems. SIAM Journal on Applied Mathematics, 53(3):770–798, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ward:1993:SLE

[3012] Michael J. Ward, William D. Henshaw, and Joseph B. Keller. Summing logarithmic expansions for singularly perturbed eigenvalue problems. SIAM Journal on Applied Mathematics, 53(3):799–828, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chaderjian:1993:RPA

[3013] Bruce J. Chaderjian and Kenneth P. Bube. Recovery of perturbations in an acoustic medium with attenuation from several plane wave responses. SIAM Journal on Applied Mathematics, 53(3):829–846, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Colton:1993:CIE

[3014] David Colton and Peter Monk. On a class of integral equations of the first kind in inverse scattering theory. SIAM Journal on Applied Mathematics, 53(3):847–860, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smolen:1993:PBM

[3015] Paul Smolen, David Terman, and John Rinzel. Properties of a bursting model with two slow inhibitory variables. SIAM Journal on Applied Mathematics, 53(3):861–892, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Beaty:1993:DSR

[3016] M. G. Beaty and M. M. Dodson. The distribution of sampling rates for signals with equally wide, equally spaced spectral bands. SIAM Journal on Applied Mathematics, 53(3):893–906, June 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Solovyov:1993:RSF

[3017] S. E. Solovyov, V. A. Volpert, and S. P.

Davtyan. Radially symmetric flow of reacting liquid with changing viscosity. SIAM Journal on Applied Mathematics, 53(4):907–914, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bergman:1993:HSF

[3018] David J. Bergman. Hierarchies of Stieltjes functions and their application to the calculation of bounds for the dielectric constant of a two-component composite medium. SIAM Journal on Applied Mathematics, 53(4):915–930, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boersma:1993:SIE

[3019] J. Boersma and E. Danicki. On the solution of an integral equation arising in potential problems for circular and elliptic disks. SIAM Journal on Applied Mathematics, 53(4):931–941, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ikehata:1993:IPP

[3020] Masaru Ikehata. An inverse problem for the plate in the Love–Kirchhoff theory. SIAM Journal on Applied Mathematics, 53(4):942–970, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dewynne:1993:MRF

[3021] J. N. Dewynne, A. C. Fowler, and P. S. Hagan. Multiple reaction fronts in the oxidation-reduction of iron-rich uranium ores. SIAM Journal on Applied Mathematics, 53(4):971–989, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bates:1993:DNC

[3022] Peter W. Bates and Paul C. Fife. The dynamics of nucleation for the Cahn– Hilliard equation. SIAM Journal on Applied Mathematics, 53(4):990–1008, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Booty:1993:ATW

[3023] M. R. Booty, R. Haberman, and A. A. Minzoni. The accommodation of traveling waves of Fisher’s type to the dynamics of the leading tail. SIAM Journal on Applied Mathematics, 53 (4):1009–1025, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hsu:1993:SRD

[3024] Sze Bi Hsu and P. Waltman. On a system of reaction-diffusion equations arising from competition in an unstirred chemostat. SIAM Journal on Applied Mathematics, 53(4):1026–1044, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Holden:1993:SPT

[3025] Lisa Holden and Thomas Erneux. Slow passage through a Hopf bifurcation: from oscillatory to steady state solutions. SIAM Journal on Applied Mathematics, 53(4):1045–1058, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Li:1993:NTB

[3026] Ji Bin Li and Jian Ming Zhang. New treatment on bifurcations of periodic solutions and homoclinic orbits at high r in the Lorenz equations. SIAM Journal on Applied Mathematics, 53 (4):1059–1071, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grenander:1993:SPI

[3027] Ulf Grenander and Daniel MacRae Keenan. On the shape of plane images. SIAM Journal on Applied Mathematics, 53(4):1072–1094, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kushner:1993:HTA

[3028] Harold J. Kushner and L. Felipe Martins. Heavy traffic analysis of a data transmission system with many independent sources. SIAM Journal on Applied Mathematics, 53(4):1095–1122, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Adan:1993:AMQ

[3029] I. J. B. F. Adan, J. Wessels, and W. H. M. Zijm. Analysing multiprogramming queues by generating functions. SIAM Journal on Applied Mathematics, 53(4):1123–1131, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yang:1993:CST

[3030] Yongzhi Yang and Charles Knessl. Conditional sojourn time moments in the finite capacity GI/M/1 queue with processor-sharing service. SIAM Journal on Applied Mathematics, 53(4): 1132–1193, August 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Norris:1993:WPL

[3031] Andrew N. Norris. Waves in periodically layered media: A comparison of two theories. SIAM Journal on Applied Mathematics, 53(5): 1195–1209, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kuske:1993:SEO

[3032] R. Kuske, Z. Schuss, I. Goldhirsch, and S. H. Noskowicz. Schrödinger’s equation on a one-dimensional lattice with weak disorder. SIAM Journal on Applied Mathematics, 53(5): 1210–1252, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schiavone:1993:OPT

[3033] P. Schiavone and C. Constanda. Oscillation problems in thin plates with transverse shear deformation. SIAM Journal on Applied Mathematics, 53 (5):1253–1263, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goldberg:1993:STD

[3034] Irwin S. Goldberg and Robert T.

Folk. Solutions to time-dependent pure-end-condition problems of elasticity: Pressure-step wave propagation and end-resonance effects. SIAM Journal on Applied Mathematics, 53(5): 1264–1292, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pember:1993:NMH

[3035] Richard B. Pember. Numerical methods for hyperbolic conservation laws with stiff relaxation I. spurious solutions. SIAM Journal on Applied Mathematics, 53(5):1293–1330, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Newton:1993:RFI

[3036] Paul K. Newton. Rapidly forced initial value problems. SIAM Journal on Applied Mathematics, 53(5): 1331–1351, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bossavit:1993:BVP

[3037] Alain Bossavit. Boundary value problems with symmetry and their approximation by finite elements. SIAM Journal on Applied Mathematics, 53 (5):1352–1380, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kirby:1993:SWR

[3038] James T. Kirby and Changhoon Lee. Short waves in a rotating, shallow tank with bathymetry: A model equation in the mild-slope approximation. SIAM Journal on Applied Mathematics, 53 (5):1381–1400, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klein:1993:RBC

[3039] Rupert Klein and D. Scott Stewart. The relation between curvature, rate state-dependence, and detonation velocity. SIAM Journal on Applied Mathematics, 53(5):1401–1435, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Frankel:1993:RPB

[3040] Paul Frankel and Tim Kiemels. Relative phase behavior of two slowly coupled oscillators. SIAM Journal on Applied Mathematics, 53(5):1436–1446, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Thieme:1993:HMI

[3041] Horst R. Thieme and Carlos Castillo-Chavez. How may infection-age-dependent infectivity affect the dynamics of HIV/AIDS? SIAM Journal on Applied Mathematics, 53(5): 1447–1479, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arino:1993:CAM

[3042] Ovide Arino and Marek Kimmel. Comparison of approaches to modeling of cell population dynamics. SIAM Journal on Applied Mathematics, 53(5): 1480–1504, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ford:1993:ERB

[3043] R. M. Ford and P. T. Cummings. Errata: “On the relationship between cell balance equations for chemotactic cell populations” [SIAM J. Appl. Math. 52 (1992), no. 5, 1426–1441, MR 93h:92018]. SIAM Journal on Applied Mathematics, 53(5):1505, October 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [2953].

Dhia:1993:MAG

[3044] Anne-Sophie Bonnet-Ben Dhia and Patrick Joly. Mathematical analysis of guided water waves. SIAM Journal on Applied Mathematics, 53(6): 1507–1550, December 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kim:1993:SFM

[3045] J. S. Kim and F. A. Williams. Structures of flow and mixture-fraction fields for counterflow diffusion flames with small stoichiometric mixture fractions. SIAM Journal on Applied Mathematics, 53(6):1551–1566, December 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vazquez-Espi:1993:ESC

[3046] Carlos Vázquez-Espí and Amable Liñán. The effect of square corners on the ignition of solids. SIAM Journal on Applied Mathematics, 53(6): 1567–1590, December 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Marchant:1993:MHM

[3047] T. R. Marchant and N. F. Smyth. Microwave heating of materials with nonohmic conductance. SIAM Journal on Applied Mathematics, 53(6): 1591–1612, December 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Luke:1993:VUB

[3048] Jonathan H. C. Luke. A variational upper bound on the renormalized mean sedimentation speed in concentrated suspensions of identical randomly arranged spheres. SIAM Journal on Applied Mathematics, 53(6): 1613–1635, December 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Santosa:1993:FOC

[3049] Fadil Santosa and Michael Vogelius. First-order corrections to the homogenized eigenvalues of a periodic composite medium. SIAM Journal on Applied Mathematics, 53(6):1636–1668, December 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [3183].

Rubinstein:1993:FIN

[3050] Jacob Rubinstein, Peter Sternberg, and Joseph B. Keller. Front interaction and nonhomogeneous equilibria for tristable reaction-diffusion equations. SIAM Journal on Applied Mathematics, 53(6):1669–1685, December 1993.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eyre:1993:SCH

[3051] David J. Eyre. Systems of Cahn– Hilliard equations. SIAM Journal on Applied Mathematics, 53(6):1686–1712, December 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sherratt:1993:CGC

[3052] Jonathan A. Sherratt. Cellular growth control and travelling waves of cancer. SIAM Journal on Applied Mathematics, 53(6):1713–1730, December 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rundell:1993:DDR

[3053] William Rundell. Determining the death rate for an age-structured population from census data. SIAM Journal on Applied Mathematics, 53(6): 1731–1746, December 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kimura:1993:PIP

[3054] Tsuyoshi Kimura and Takashi Suzuki.

A parabolic inverse problem arising in a mathematical model for chromatography. SIAM Journal on Applied Mathematics, 53(6):1747–1761, December 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ram:1993:IEP

[3055] Yitshak M. Ram. Inverse eigenvalue problem for a modified vibrating system. SIAM Journal on Applied Mathematics, 53(6):1762–1775, December 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dohtani:1993:ESQ

[3056] Akitaka Dohtani. Existence of stable quasi-periodic solutions and persistence of business fluctuations. SIAM Journal on Applied Mathematics, 53 (6):1776–1784, December 1993. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vanden-Broeck:1994:FNI

[3057] J. M. Vanden-Broeck and E. O. Tuck. Flow near the intersection of a free surface with a vertical wall. SIAM Journal on Applied Mathematics, 54(1):1–13, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/24083.

Yang:1994:SMS

[3058] Joseph Yang and Toshi Kubota. The steady motion of a symmetric, finite core size, counterrotating vortex pair. SIAM Journal on Applied Mathematics, 54(1):14–25, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24091.

vanHerwaarden:1994:SPD

[3059] Onno A. Van Herwaarden. Spread of pollution by dispersive groundwater flow. SIAM Journal on Applied Mathematics, 54(1):26–41, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Romero:1994:LHP

[3060] L. A. Romero. Low or high Peclet number flow past a sphere in a saturated porous medium. SIAM Journal on Applied Mathematics, 54(1):42–71, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23714.

Please:1994:LBC

[3061] C. P. Please, P. S. Hagan, and D. W. Schwendeman. Light-off behavior of catalytic converters. SIAM Journal on Applied Mathematics, 54(1):72–92, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/22684.

Dgaygui:1994:ABC

[3062] Kebir Dgaygui and Patrick Joly. Absorbing boundary conditions for linear gravity waves. SIAM Journal on Applied Mathematics, 54(1):93–131, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/23432.

Matkowsky:1994:SGC

[3063] B. J. Matkowsky and V. Volpert.

Spiral gasless condensed phase combustion. SIAM Journal on Applied Mathematics, 54(1):132–146, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/22925.

Bayliss:1994:TWC

[3064] Alvin Bayliss and Bernard J. Matkowsky. From traveling waves to chaos in combustion. SIAM Journal on Applied Mathematics, 54(1):147–174, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/23292.

Schmeiser:1994:MTB

[3065] Christian Schmeiser. A model for the transient behaviour of long-channel MOSFETs. SIAM Journal on Applied Mathematics, 54(1):175–194, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/22575.

Ikeda:1994:PST

[3066] Tsutomu Ikeda and Yasumasa Nishiura. Pattern selection for two breathers. SIAM Journal on Applied Mathematics, 54(1):195–230, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/23725.

Dockery:1994:EST

[3067] Jack D. Dockery and Roger Lui. Existence and stability of traveling wave solutions for a population genetic model via singular perturbations. SIAM Journal on Applied Mathematics, 54(1):231–248, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23938.

Lange:1994:SPAe

[3068] Charles G. Lange and Robert M.

Miura. Singular perturbation analysis of boundary-value problems for differential-difference equations. V. small shifts with layer behavior. SIAM Journal on Applied Mathematics, 54 (1):249–272, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lange:1994:SPAf

[3069] Charles G. Lange and Robert M.

Miura. Singular perturbation analysis of boundary value problems for differential-difference equations. VI. Small shifts with rapid oscillations. SIAM Journal on Applied Mathematics, 54(1):273–283, February 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rusak:1994:NSS

[3070] Z. Rusak. Novel similarity solutions of the sonic small-disturbance equation with applications to airfoil transonic aerodynamics. SIAM Journal on Applied Mathematics, 54(2):285–308, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25005.

Romero:1994:MNN

[3071] L. A. Romero. On the monotonicity of the Nusselt number as a function of the Peclet number. SIAM Journal on Applied Mathematics, 54(2):309–316, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23758.

Gartland:1994:ASL

[3072] Eugene C. Gartland, Jr. Analysis of a simplified liquid crystal shear-flow model. SIAM Journal on Applied Mathematics, 54(2):317–332, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/24297.

Fannjiang:1994:CED

[3073] Albert Fannjiang and George Papanicolaou. Convection enhanced diffusion for periodic flows. SIAM Journal on Applied Mathematics, 54(2):333–408, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/23678.

Gardner:1994:QHM

[3074] Carl L. Gardner. The quantum hydrodynamic model for semiconductor devices. SIAM Journal on Applied Mathematics, 54(2):409–427, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/24042.

Lipton:1994:OBP

[3075] Rob Lipton and Jim Northrup. Optimal bounds on the in-plane shear moduli for orthotropic elastic composites. SIAM Journal on Applied Mathematics, 54(2):428–442, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/23598.

Faris:1994:TDR

[3076] William G. Faris and Woody J.

Tsay. Time delay in random scattering. SIAM Journal on Applied Mathematics, 54(2):443–455, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/23392.

Ellermeyer:1994:CCG

[3077] S. F. Ellermeyer. Competition in the chemostat: Global asymptotic behavior of a model with delayed response in growth. SIAM Journal on Applied Mathematics, 54(2):456–465, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/22522.

Chiu:1994:PMP

[3078] Chichia Chiu and Frank C. Hoppensteadt. A particle method for population waves. SIAM Journal on Applied Mathematics, 54(2):466–477, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23612.

Ermentrout:1994:IPP

[3079] G. B. Ermentrout and N. Kopell.

Inhibition-produced patterning in chains of coupled nonlinear oscillators. SIAM Journal on Applied Mathematics, 54(2):478–507, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/23196.

Knessl:1994:DMN

[3080] Charles Knessl. On the distribution of the maximum number of broken machines for the repairman problem. SIAM Journal on Applied Mathematics, 54(2):508–547, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/23372.

March:1994:BLR

[3081] Peter March and Timo Seppäläinen. Bounds for least relative vacancy in a simple mosaic process. SIAM Journal on Applied Mathematics, 54(2):548– 558, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/23360.

Inselberg:1994:MLR

[3082] Alfred Inselberg and Bernard Dimsdale. Multidimensional lines. I. Representation. SIAM Journal on Applied Mathematics, 54(2):559–577, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/21689.

Inselberg:1994:MLI

[3083] Alfred Inselberg and Bernard Dimsdale. Multidimensional lines II: Proximity and applications. SIAM Journal on Applied Mathematics, 54(2):578– 596, April 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/21690.

Or:1994:DTT

[3084] A. C. Or. The dynamics of a Tippe top. SIAM Journal on Applied Mathematics, 54(3):597–609, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/23512.

Budd:1994:BSP

[3085] C. J. Budd, J. W. Dold, and A. M.

Stuart. Blow-up in a system of partial differential equations with conserved first integral. Part II: Problems with convection. SIAM Journal on Applied Mathematics, 54(3):610–640, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23213.

Chen:1994:LSA

[3086] Zhangxin Chen. Large-scale averaging analysis of single phase flow in fractured reservoirs. SIAM Journal on Applied Mathematics, 54(3):641–659, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23520.

Rabbani:1994:SEF

[3087] M. G. Rabbani and James W. Warner.

Shortcomings of existing finite element formulations for subsurface water pollution modeling and its rectification: One-dimensional case. SIAM Journal on Applied Mathematics, 54(3):660–673, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/24114.

Rabbani:1994:DFF

[3088] M. Ghulam Rabbani. Directformulation finite element (DFFE) method for groundwater flow modeling: Two-dimensional case. SIAM Journal on Applied Mathematics, 54(3):674– 687, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/24118.

Spector:1994:SNP

[3089] M. D. Spector and T. Miloh. Stability of nonlinear periodic internal waves in a deep stratified fluid. SIAM Journal on Applied Mathematics, 54(3):688–707, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/24411.

An:1994:EME

[3090] Lian Jun An and Anthony Peirce. The effect of microstructure on elasticplastic models. SIAM Journal on Applied Mathematics, 54(3):708–730, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/23849.

Pasa:1994:ENE

[3091] Gelu I. Paşa. Estimations for nonsymmetric effective coefficients. SIAM Journal on Applied Mathematics, 54(3):731–737, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/22836.

Baginski:1994:COP

[3092] Frank E. Baginski. The computation of one-parameter families of bifurcating elastic surfaces. SIAM Journal on Applied Mathematics, 54(3):738–773, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/22479.

Powell:1994:RLB

[3093] J. A. Powell, E. M. Wright, and J. V.

Moloney. Reflection of localized beams from a nonlinear absorbing interface. SIAM Journal on Applied Mathematics, 54(3):774–788, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Byrne:1994:SSC

[3094] Helen Byrne and John Norbury. Stable solutions for a catalytic converter. SIAM Journal on Applied Mathematics, 54(3):789–813, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/21787.

Pernarowski:1994:FSB

[3095] M. Pernarowski. Fast subsystem bifurcations in a slowly varying Lienard system exhibiting bursting. SIAM Journal on Applied Mathematics, 54(3):814–832, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/23449.

Steinberg:1994:FPS

[3096] Y. Steinberg, B. Z. Bobrovsky, and Z. Schuss. Fixed-point smoothing of scalar diffusions I: An asymptotically optimal smoother. SIAM Journal on Applied Mathematics, 54(3):833–853, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/18890.

Knessl:1994:PSQ

[3097] Charles Knessl and Charles Tier. A processor-shared queue that models switching times: Heavy usage asymptotics. SIAM Journal on Applied Mathematics, 54(3):854–875, June 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gegick:1994:PCA

[3098] M. Gegick and G. W. Young. Plasma carburization of an axisymmetric steel sample. SIAM Journal on Applied Mathematics, 54(4):877–906, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23234.

Lewicki:1994:LTE

[3099] Pawel Lewicki. Long time evolution of wavefronts in random media. SIAM Journal on Applied Mathematics, 54(4):907–934, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24905.

Coker:1994:SRA

[3100] David A. Coker and W. Brent Lindquist. Structural reliability analysis for one-dimensional, two-phase miscible flow. SIAM Journal on Applied Mathematics, 54(4):935–953, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/23943.

Chancelier:1994:ACP

[3101] J.-Ph. Chancelier, M. Cohen De Lara, and F. Pacard. Analysis of a conservation PDE with discontinuous flux: a model of settler. SIAM Journal on Applied Mathematics, 54(4):954–995, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24952.

Forest:1994:DSV

[3102] M. Gregory Forest and Q. Wang.

Dynamics of slender viscoelastic free jets. SIAM Journal on Applied Mathematics, 54(4):996–1032, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23676.

Haaker:1994:DAO

[3103] T. I. Haaker and A. H. P. Van der Burgh. On the dynamics of aeroelastic oscillators with one degree of freedom. SIAM Journal on Applied Mathematics, 54(4):1033–1047, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24480.

Brown:1994:CSF

[3104] David L. Brown, M. Gregory Forest, Brian J. Miller, and N. Anders Petersson. Computation and stability of fluxons in a singularly perturbed sine-Gordon model of the Josephson junction. SIAM Journal on Applied Mathematics, 54(4):1048–1066, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/23393.

Fang:1994:RSD

[3105] Wei Fu Fang and Kazufumi Ito. Reconstruction of semiconductor doping profile from laser-beam-induced current image. SIAM Journal on Applied Mathematics, 54(4):1067–1082, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/23480.

Schwartz:1994:SHP

[3106] Ira B. Schwartz and Thomas Erneux. Subharmonic hysteresis and period doubling bifurcations for a periodically driven laser. SIAM Journal on Applied Mathematics, 54(4):1083–1100, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/23070.

Frankel:1994:SPD

[3107] M. Frankel, V. Roytburd, and G. Sivashinsky. A sequence of period doublings and chaotic pulsations in a free boundary problem modeling thermal instabilities. SIAM Journal on Applied Mathematics, 54(4):1101–1112, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/23072.

Smith:1994:CSL

[3108] Hal L. Smith and Paul Waltman. Competition for a single limiting resource in continuous culture: the variableyield model. SIAM Journal on Applied Mathematics, 54(4):1113–1131, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24534.

Gaines:1994:RGS

[3109] J. G. Gaines and T. J. Lyons. Random generation of stochastic area integrals. SIAM Journal on Applied Mathematics, 54(4):1132–1146, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23570.

Mei:1994:AAQ

[3110] Jing Dong Mei and Charles Tier.

Asymptotic approximations for a queueing network with multiple classes. SIAM Journal on Applied Mathematics, 54(4):1147–1180, August 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vazquez-Espi:1994:IRS

[3111] Carlos Vázquez-Espí and Amable Liñán. Ignition of a rectangular solid by an external heat flux. SIAM Journal on Applied Mathematics, 54(5):1181–1202, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/25069.

McLaughlin:1994:RSS

[3112] Joyce R. McLaughlin, Peter L. Polyakov, and Paul E. Sacks. Reconstruction of a spherically symmetric speed of sound. SIAM Journal on Applied Mathematics, 54(5):1203–1223, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/23821.

Chaderjian:1994:UTL

[3113] Bruce J. Chaderjian. A uniqueness theorem for a lossy inverse problem in reflection seismology. SIAM Journal on Applied Mathematics, 54(5):1224–1249, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/24115.

Knill:1994:DAI

[3114] Oliver Knill. Diagonalization of Abel’s integral operator. SIAM Journal on Applied Mathematics, 54(5):1250–1253, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/25090.

Nunez:1994:LEI

[3115] Manuel Núñez. Localized eigenmodes of the induction equation. SIAM Journal on Applied Mathematics, 54(5):1254–1267, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/23479.

Gong:1994:MSS

[3116] Lianger Gong and Samuel S. Shen.

Multiple supercritical solitary wave solutions of the stationary forced Korteweg–de Vries equation and their stability. SIAM Journal on Applied Mathematics, 54(5):1268–1290, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23351.

Balmforth:1994:CPT

[3117] N. J. Balmforth, G. R. Ierley, and E. A. Spiegel. Chaotic pulse trains. SIAM Journal on Applied Mathematics, 54(5):1291–1334, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/24742.

Falcone:1994:LSV

[3118] M. Falcone, T. Giorgi, and P. Loreti.

Level sets of viscosity solutions: Some applications to fronts and rendez-vous problems. SIAM Journal on Applied Mathematics, 54(5):1335–1354, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23306.

Ei:1994:DIC

[3119] S.-I. Ei and E. Yanagida. Dynamics of interfaces in competition-diffusion systems. SIAM Journal on Applied Mathematics, 54(5):1355–1373, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24734.

Sherratt:1994:EPP

[3120] Jonathan A. Sherratt. On the evolution of periodic plane waves in reaction-diffusion systems of λ-ω type. SIAM Journal on Applied Mathematics, 54(5):1374–1385, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24374.

Paullet:1994:ESW

[3121] Joseph Paullet, Bard Ermentrout, and William Troy. The existence of spiral waves in an oscillatory reaction-diffusion system. SIAM Journal on Applied Mathematics, 54(5):1386–1401, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/24918.

Belair:1994:SBE

[3122] Jacques Bélair and Sue Ann Campbell. Stability and bifurcations of equilibria in a multiple-delayed differential equation. SIAM Journal on Applied Mathematics, 54(5):1402–1424, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24885.

Choudhury:1994:TIC

[3123] S. Roy Choudhury. Turing instability in competition models with delay I: Linear theory. SIAM Journal on Applied Mathematics, 54(5):1425–1450, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/24724.

Choudhury:1994:ASS

[3124] S. Roy Choudhury. Analysis of spatial structure in a predator-prey model with delay II: Nonlinear theory. SIAM Journal on Applied Mathematics, 54(5):1451–1467, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24725.

Barbagallo:1994:AEW

[3125] Roberto Barbagallo, Marina Mochi, and Francesco Zirilli. Asymptotic expansion of the waiting time distribution of two models of a closed processor-sharing system: Heavy usage. SIAM Journal on Applied Mathematics, 54(5):1468–1491, October 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24797.

Tew:1994:GAD

[3126] R. H. Tew. Geometrical aspects of the diffraction of critically incident bulk waves and Gaussian beams. SIAM Journal on Applied Mathematics, 54(6):1493–1520, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24852.

Bonilla:1994:GEI

[3127] Luis L. Bonilla, Francisco J. Higuera, and Stephanos Venakides. The gunn effect: Instability of the steady state and stability of the solitary wave in long extrinsic semiconductors. SIAM Journal on Applied Mathematics, 54(6):1521–1541, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/23655.

Dobson:1994:RSA

[3128] David C. Dobson and Fadil Santosa. Resolution and stability analysis of an inverse problem in electrical impedance tomography: Dependence on the input current patterns. SIAM Journal on Applied Mathematics, 54(6):1542–1560, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/23759.

Velazquez:1994:DCT

[3129] J. J. L. Velázquez. On the dynamics of a closed thermosyphon. SIAM Journal on Applied Mathematics, 54(6):1561–1593, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/24678.

Margolis:1994:RMI

[3130] Stephen B. Margolis. Resonant mode interactions and the bifurcation of combustion-driven acoustic oscillations in resonance tubes. SIAM Journal on Applied Mathematics, 54(6):1594–1633, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/24255.

Reddy:1994:PCD

[3131] Satish C. Reddy and Lloyd N. Trefethen. Pseudospectra of the convection-diffusion operator. SIAM Journal on Applied Mathematics, 54(6):1634–1649, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/24698.

Fowler:1994:GSF

[3132] Andrew C. Fowler and William B.

Krantz. A generalized secondary frost heave model. SIAM Journal on Applied Mathematics, 54(6):1650–1675, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25255.

Gardner:1994:NSU

[3133] Carl L. Gardner and David G. Schaeffer. Numerical simulation of uniaxial compression of a granular material with wall friction. SIAM Journal on Applied Mathematics, 54(6):1676–1692, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25172.

Clarkson:1994:ANM

[3134] Peter A. Clarkson and Elizabeth L.

Mansfield. Algorithms for the nonclassical method of symmetry reductions. SIAM Journal on Applied Mathematics, 54(6):1693–1719, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Paullet:1994:SRW

[3135] Joseph E. Paullet and G. Bard Ermentrout. Stable rotating waves in two-dimensional discrete active media. SIAM Journal on Applied Mathematics, 54(6):1720–1744, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25068.

Knessl:1994:EAS

[3136] Charles Knessl and Craig Steven Peters. Exact and asymptotic solutions for the time-dependent problem of collective ruin I. SIAM Journal on Applied Mathematics, 54(6):1745–1767, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25057.

Tan:1994:HTA

[3137] Xiaoming Tan and Charles Knessl. Heavy traffic asymptotics for a gated, infinite-server queue with uniform service times. SIAM Journal on Applied Mathematics, 54(6):1768–1779, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/24967.

Newton:1994:VRS

[3138] Nigel J. Newton. Variance reduction for simulated diffusions. SIAM Journal on Applied Mathematics, 54(6):1780–1805, December 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/23622.

Borgers:1995:TIS

[3139] Christoph Börgers and Edward W. Larsen. The transversely integrated scalar flux of a narrowly focused particle beam. SIAM Journal on Applied Mathematics, 55(1):1–22, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25229.

Ida:1995:DLC

[3140] M. P. Ida and M. J. Miksis. Dynamics of a lamella in a capillary tube. SIAM Journal on Applied Mathematics, 55(1):23–57, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25613.

Bechtel:1995:TEH

[3141] S. E. Bechtel, K. D. Bolinger, J. Z.

Cao, and M. G. Forest. Torsional effects in high-order viscoelastic thinfilament models. SIAM Journal on Applied Mathematics, 55(1):58–99, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23957.

Rosakis:1995:EAR

[3142] Phoebus Rosakis. An equal area rule for dissipative kinetics of propagating strain discontinuities. SIAM Journal on Applied Mathematics, 55(1): 100–123, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/
25088.

Grimshaw:1995:WNS

[3143] Roger Grimshaw and Nalini Joshi.

Weakly nonlocal solitary waves in a singularly perturbed Korteweg–de Vries equation. SIAM Journal on Applied Mathematics, 55(1):124–135, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/24382.

An:1995:WNA

[3144] Lian Jun An and Anthony Peirce. A weakly nonlinear analysis of elasto-plastic-microstructure models. SIAM Journal on Applied Mathematics, 55(1):136–155, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/25532.

Chapman:1995:LDA

[3145] S. Jonathan Chapman, Qiang Du, and Max D. Gunzburger. On the Lawrence– Doniach and anisotropic Ginzburg– Landau models for layered superconductors. SIAM Journal on Applied Mathematics, 55(1):156–174, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25683.

Gasser:1995:DDW

[3146] I. Gasser and P. Szmolyan. Detonation and deflagration waves with multistep reaction schemes. SIAM Journal on Applied Mathematics, 55(1):175–191, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/24477.

Ratnanather:1995:STB

[3147] J. T. Ratnanather and P. G. Daniels. Solution of the thermal boundary layer equations in regions of flow reversal. SIAM Journal on Applied Mathematics, 55(1):192–204, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23861.

Panfilov:1995:DDS

[3148] Alexander V. Panfilov and James P.

Keener. Dynamics of dissipative structures in reaction-diffusion equations. SIAM Journal on Applied Mathematics, 55(1):205–219, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/22910.

Pinsky:1995:SCE

[3149] Paul F. Pinsky. Synchrony and clustering in an excitatory neural network model with intrinsic relaxation kinetics. SIAM Journal on Applied Mathematics, 55(1):220–241, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25737.

Grebert:1995:RPL

[3150] Benoît Grébert and Ricardo Weder.

Reconstruction of a potential on the line that is A priori known on the half line. SIAM Journal on Applied Mathematics, 55(1):242–254, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25465.

Bube:1995:URD

[3151] Kenneth P. Bube. Uniqueness of reflector depths and characterization of the slowness null space in linearized seismic reflection tomography. SIAM Journal on Applied Mathematics, 55(1):255–266, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/24116.

Dewynne:1995:NAR

[3152] J. N. Dewynne and P. Wilmott. A note on average rate options with discrete sampling. SIAM Journal on Applied Mathematics, 55(1):267–276, February 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cole:1995:P

[3153] Julian D. Cole, Donald A. Drew, Joseph E. Flaherty, Bernard J. Matkowsky, and Lu Ting. Preface. SIAM Journal on Applied Mathematics, 55(2):277–278, April 1995. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Grote:1995:ENB

[3154] Marcus J. Grote and Joseph B. Keller. Exact nonreflecting boundary conditions for the time dependent wave equation. SIAM Journal on Applied Mathematics, 55(2):280–297, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/26926. Perturba-

tion methods in physical mathematics (Troy, NY, 1993).

Balmforth:1995:IRR

[3155] N. J. Balmforth, L. N. Howard, and E. A. Spiegel. Instability of rapidly rotating polytropes. SIAM Journal on Applied Mathematics, 55(2):298–331, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26929. Perturba-

tion methods in physical mathematics (Troy, NY, 1993).

Laforgue:1995:SLM

[3156] Jacques G. L. Laforgue and Robert E.

O’Malley, Jr. Shock layer movement for Burgers’ equation. SIAM Journal on Applied Mathematics, 55(2):332–347, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26928. Per-

turbation methods in physical mathematics (Troy, NY, 1993).

Cohen:1995:SFM

[3157] Donald S. Cohen, Andrew B. White, Jr., and Thomas P. Witelski. Shock formation in a multidimensional viscoelastic diffusive system. SIAM Journal on Applied Mathematics, 55(2):348–368, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26933. Per-

turbation methods in physical mathematics (Troy, NY, 1993).

Lu:1995:SCS

[3158] Ting Lu. On-surface conditions for structural acoustic interactions in moving media. SIAM Journal on Applied Mathematics, 55(2):369–389, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26921. Perturba-

tion methods in physical mathematics (Troy, NY, 1993).

Burridge:1995:AEI

[3159] Robert Burridge. Asymptotic evaluation of integrals related to timedependent fields near caustics. SIAM Journal on Applied Mathematics, 55 (2):390–409, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/

26924. Perturbation methods in physical mathematics (Troy, NY, 1993).

Cole:1995:LPE

[3160] Julian D. Cole. Limit process expansions and homogenization. SIAM Journal on Applied Mathematics, 55 (2):410–424, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/

26923. Perturbation methods in physical mathematics (Troy, NY, 1993).

Ward:1995:ILS

[3161] Michael J. Ward and Luis G. Reyna. Internal layers, small eigenvalues, and the sensitivity of metastable motion. SIAM Journal on Applied Mathematics, 55(2):425–445, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26925. Perturba-

tion methods in physical mathematics (Troy, NY, 1993).

Kevorkian:1995:WNW

[3162] J. Kevorkian, J. Yu, and L. Wang.

Weakly nonlinear waves for a class of linearly unstable hyperbolic conservation laws with source terms. SIAM Journal on Applied Mathematics, 55(2):446–484, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26931. Perturba-

tion methods in physical mathematics (Troy, NY, 1993).

Bernoff:1995:MFE

[3163] A. J. Bernoff, R. Kuske, B. J.

Matkowsky, and V. Volpert. Mean field effects for counterpropagating traveling wave solutions of reaction-diffusion systems. SIAM Journal on Applied Mathematics, 55(2):485–519, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26930. Perturba-

tion methods in physical mathematics (Troy, NY, 1993).

Adjerid:1995:HOF

[3164] Slimane Adjerid, Mohammed Aiffa, and Joseph E. Flaherty. High-order finite element methods for singularly perturbed elliptic and parabolic problems. SIAM Journal on Applied Mathematics, 55(2):520–543, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26934. Perturba-

tion methods in physical mathematics (Troy, NY, 1993).

Hoppensteadt:1995:SPS

[3165] Frank C. Hoppensteadt. Singular perturbation solutions of noisy systems. SIAM Journal on Applied Mathematics, 55(2):544–551, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/26922. Perturba-

tion methods in physical mathematics (Troy, NY, 1993).

Childress:1995:TBM

[3166] Stephen Childress. Turbulent Baker’s maps. SIAM Journal on Applied Mathematics, 55(2):552–563, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/26932. Perturba-

tion methods in physical mathematics (Troy, NY, 1993).

Ludwig:1995:TSH

[3167] Donald Ludwig. A theory of sustainable harvesting. SIAM Journal on Applied Mathematics, 55(2):564–575, April 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/26927. Per-

turbation methods in physical mathematics (Troy, NY, 1993).

Weisenborn:1995:ODS

[3168] A. J. Weisenborn and B. I. M. ten Bosch. On the Oseen drag on a sphere. SIAM Journal on Applied Mathematics, 55(3):577–592, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/26052.

Ribaric:1995:ETM

[3169] Marijan Ribarič and Luka Šušteršič.

Expansions in terms of moments of time-dependent, moving charges and currents. SIAM Journal on Applied Mathematics, 55(3):593–624, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/24197.

Isaacson:1995:CGM

[3170] Eli Isaacson and Blake Temple. Convergence of the 2 × 2 Godunov method for a general resonant nonlinear balance law. SIAM Journal on Applied Mathematics, 55(3):625–640, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/24071.

Arkeryd:1995:BMB

[3171] Leif Arkeryd and Reinhard Illner.

The Broadwell model in a box: strong L^{1}-convergence to equilibrium. SIAM Journal on Applied Mathematics, 55(3):641–650, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/23240.

Holden:1995:MPC

[3172] Helge Holden, Nils Henrik Risebro, and Aslak Tveito. Maximum principles for a class of conservation laws. SIAM Journal on Applied Mathematics, 55(3):651–661, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/25598.

Edwards:1995:UMB

[3173] D. A. Edwards and D. S. Cohen. A unusual moving boundary condition arising in anomalous diffusion problems. SIAM Journal on Applied Mathematics, 55(3):662–676, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25405.

Ring:1995:ICB

[3174] Wolfgang Ring. Identification of a core from boundary data. SIAM Journal on Applied Mathematics, 55(3):677–706, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25630.

Sharafutdinov:1995:ETI

[3175] Vladimir A. Sharafutdinov. On emission tomography of inhomogeneous media. SIAM Journal on Applied Mathematics, 55(3):707–718, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/24434.

Murthy:1995:NDI

[3176] A. S. Vasudeva Murthy. A note on the differential inversion method of Hohlfeld et al. SIAM Journal on Applied Mathematics, 55(3):719–722, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25498.

Rosen:1995:CAG

[3177] Dan Rosen and Donald E. Cormack. The continuation approach:

A general framework for the analysis and evaluation of singular and nearsingular integrals. SIAM Journal on Applied Mathematics, 55(3):723–762, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24325.

Hsu:1995:GSC

[3178] Sze Bi Hsu and Tzy Wei Huang.

Global stability for a class of predatorprey systems. SIAM Journal on Applied Mathematics, 55(3):763–783, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25320.

Chen:1995:TDB

[3179] Thomas M. Chen and Vijay K.

Samalam. Time-dependent behavior of fluid buffer models with Markov input and constant output rates. SIAM Journal on Applied Mathematics, 55(3):784–799, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/24010.

Hall:1995:EMS

[3180] Peter Hall. On the effect of measuring a self-similar process. SIAM Journal on Applied Mathematics, 55(3):800–808, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23656.

Simonian:1995:ADE

[3181] Alain Simonian. Asymptotic distribution of exit times for smallnoise diffusions. SIAM Journal on Applied Mathematics, 55(3):809–826, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/24460.

Chambolle:1995:ISV

[3182] Antonin Chambolle. Image segmentation by variational methods: Mumford and shah functional and the discrete approximations. SIAM Journal on Applied Mathematics, 55(3):827–863, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25713.

Santosa:1995:EPF

[3183] Fadil Santosa and Michael Vogelius. Erratum to the paper: “First-order corrections to the homogenized eigenvalues of a periodic composite medium” [SIAM J. Appl. Math. 53 (1993), no. 6, 1636–1668, MR 94h:35188]. SIAM Journal on Applied Mathematics, 55 (3):864, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [3049].

Siegel:1995:SSF

[3184] Michael Siegel. A study of singularity formation in the Kelvin– Helmholtz instability with surface tension. SIAM Journal on Applied Mathematics, 55(4):865–891, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/26265.

Vanden-Broeck:1995:CSG

[3185] Jean-Marc Vanden-Broeck and Touvia Miloh. Computations of steep gravity waves by a refinement of Davies–Tulin’s approximation. SIAM Journal on Applied Mathematics, 55(4):892–903, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26240.

Luke:1995:FSI

[3186] C. J. Luke and P. A. Martin. Fluidsolid interaction: Acoustic scattering by a smooth elastic obstacle. SIAM Journal on Applied Mathematics, 55(4):904–922, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/25902.

Wang:1995:NOR

[3187] Meng Wang and David R. Kassoy.

Nonlinear oscillations in a resonant gas column: An initial-boundary-value study. SIAM Journal on Applied Mathematics, 55(4):923–951, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24290.

Romero:1995:LHP

[3188] Louis A. Romero. Low or high Peclet number flow past a prolate spheroid in a saturated porous medium. SIAM Journal on Applied Mathematics, 55(4):952–974, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/25006.

Romero:1995:FCP

[3189] Louis A. Romero. Forced convection past a slender body in a saturated porous medium. SIAM Journal on Applied Mathematics, 55(4):975–985, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25155.

Klemm:1995:MVA

[3190] Richard A. Klemm. Models of vortices in anisotropic superconductors. SIAM Journal on Applied Mathematics, 55(4):986–1015, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/26386.

Watt:1995:ADM

[3191] Simon D. Watt and Anthony J. Roberts. The accurate dynamic modelling of contaminant dispersion in channels. SIAM Journal on Applied Mathematics, 55(4):1016–1038, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25797.

Edwards:1995:CFS

[3192] David A. Edwards. Constant front speed in weakly diffusive non-fickian systems. SIAM Journal on Applied Mathematics, 55(4):1039–1058, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25903.

Willms:1995:ESS

[3193] Allan R. Willms. An exact solution of Stikker’s nonlinear heat equation. SIAM Journal on Applied Mathematics, 55(4):1059–1073, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24917.

Franklin:1995:ERP

[3194] Joel Franklin and Wyeth Bair. The effect of a refractory period on the power spectrum of neuronal discharge. SIAM Journal on Applied Mathematics, 55(4):1074–1093, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/25869.

Majda:1995:RBP

[3195] George Majda and Mu Sheng Wei. Relationships between a potential and its scattering frequencies. SIAM Journal on Applied Mathematics, 55(4):1094– 1116, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/23118.

Lebovitz:1995:DBH

[3196] Norman R. Lebovitz and Adriana I.

Pesci. Dynamic bifurcation in Hamiltonian systems with one degree of freedom. SIAM Journal on Applied Mathematics, 55(4):1117–1133, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/22090.

Ocken:1995:RCO

[3197] Stanley Ocken. Recognizing convergent orbits of discrete dynamical systems. SIAM Journal on Applied Mathematics, 55(4):1134–1160, August 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25757.

Frigaard:1995:DSF

[3198] Ian A. Frigaard. The dynamics of spray-formed billets. SIAM Journal on Applied Mathematics, 55(5):1161–1203, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/25756.

Ogami:1995:SUC

[3199] Yoshifumi Ogami and Angela Y. Cheer. Simulations of unsteady compressible fluid motion by an interactive cored particle method. SIAM Journal on Applied Mathematics, 55(5):1204–1226, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/25546.

Weisenborn:1995:ODI

[3200] A. J. Weisenborn and B. I. M. ten Bosch. The Oseen drag at infinite Reynolds number. SIAM Journal on Applied Mathematics, 55(5):1227–1232, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/27347.

Chapman:1995:SFT

[3201] S. Jonathan Chapman. Superheating field of type II superconductors. SIAM Journal on Applied Mathematics, 55(5):1233–1258, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25476.

Chapman:1995:MFM

[3202] S. Jonathan Chapman. A mean-field model of superconducting vortices in three dimensions. SIAM Journal on Applied Mathematics, 55(5):1259–1274, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/26366.

Chapman:1995:MVT

[3203] S. Jonathan Chapman and G. Richardson. Motion of vortices in type II superconductors. SIAM Journal on Applied Mathematics, 55(5):1275–1296, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26387.

Kopell:1995:SSF

[3204] Nancy Kopell and Michael Landman. Spatial structure of the focusing singularity of the nonlinear Schrödinger Equation: A geometrical analysis. SIAM Journal on Applied Mathematics, 55(5):1297–1323, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26238.

Kirsch:1995:FES

[3205] Andreas Kirsch and Peter Monk. A finite element/spectral method for approximating the time-harmonic Maxwell system in r^{3}. SIAM Journal on Applied Mathematics, 55(5): 1324–1344, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/
25989.

Dassios:1995:EHF

[3206] George Dassios and Zafiria Rigou. Elastic Herglotz functions. SIAM Journal on Applied Mathematics, 55(5): 1345–1361, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/
25599.

Miller:1995:TRA

[3207] Richard Miller, James Vesenka, and Eric Henderson. Tip reconstruction for the atomic force microscope. SIAM Journal on Applied Mathematics, 55(5):1362–1371, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25547.

Booth:1995:UPF

[3208] Victoria Booth and Thomas Erneux. Understanding propagation failure as a slow capture near a limit point. SIAM Journal on Applied Mathematics, 55(5):1372–1389, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26494.

Layton:1995:DNM

[3209] H. E. Layton, E. Bruce Pitman, and Mark A. Knepper. A dynamic numerical method for models of the urine concentrating mechanism. SIAM Journal on Applied Mathematics, 55(5):1390–1418, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25286.

Dickinson:1995:TEI

[3210] Richard B. Dickinson and Robert T.

Tranquillo. Transport equations and indices for random and biased cell migration based on single cell properties. SIAM Journal on Applied Mathematics, 55(5):1419–1454, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/23733.

Rusak:1995:TFA

[3211] Zvi Rusak. Transonic flow around optimum critical airfoils. SIAM Journal on Applied Mathematics, 55(5):1455–1467, October 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/27421.

Daalhuis:1995:SPM

[3212] A. B. Olde Daalhuis, S. J. Chapman, J. R. King, J. R. Ockendon, and R. H. Tew. Stokes phenomenon and matched asymptotic expansions. SIAM Journal on Applied Mathematics, 55(6):1469–1483, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/26176.

Kropinski:1995:HAN

[3213] M. C. A. Kropinski, Michael J. Ward, and Joseph B. Keller. A hybrid asymptotic-numerical method for low Reynolds number flows past a cylindrical body. SIAM Journal on Applied Mathematics, 55(6):1484–1510, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/27107.

Almgren:1995:CST

[3214] Robert Almgren. Crystalline Saffman– Taylor fingers. SIAM Journal on Applied Mathematics, 55(6):1511–1535, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/25990.

Jin:1995:NIS

[3215] Shi Jin. Numerical integrations of systems of conservation laws of mixed type. SIAM Journal on Applied Mathematics, 55(6):1536–1551, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/26837.

Sainsaulieu:1995:TWS

[3216] Lionel Sainsaulieu. Traveling waves solution of convection-diffusion systems whose convection terms are weakly nonconservative: Application to the modeling of two-phase fluid flows. SIAM Journal on Applied Mathematics, 55(6):1552–1576, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26829.

Stockie:1995:SAI

[3217] John M. Stockie and Brian T. R.

Wetton. Stability analysis for the immersed fiber problem. SIAM Journal on Applied Mathematics, 55(6):1577–1591, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26701.

Fitt:1995:CPM

[3218] Alistair D. Fitt, Amanda D. Kelly, and Colin P. Please. Crack propagation models for rock fracture in a geothermal energy reservoir. SIAM Journal on Applied Mathematics, 55(6):1592–1608, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26024.

Mahadevan:1995:PFT

[3219] L. Mahadevan and Joseph B. Keller.

Periodic folding of thin sheets. SIAM Journal on Applied Mathematics, 55(6):1609–1624, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26341.

Bonilla:1995:OEG

[3220] Luis L. Bonilla and Francisco J.

Higuera. The onset and end of the Gunn effect in extrinsic semiconductors. SIAM Journal on Applied Mathematics, 55(6):1625–1649, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/19945.

Bose:1995:SAP

[3221] Amitabha Bose. Symmetric and antisymmetric pulses in parallel coupled nerve fibres. SIAM Journal on Applied Mathematics, 55(6):1650–1674, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/26239.

Kuang:1995:ADT

[3222] Yang Kuang and Joseph W.-H. So. Analysis of a delayed two-stage population model with space-limited recruitment. SIAM Journal on Applied Mathematics, 55(6):1675–1696, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25283.

Tan:1995:ERR

[3223] Wai Yuan Tan, Si Chin Tang, and Sho Rong Lee. Effects of randomness of risk factors on the HIV epidemic in homosexual populations. SIAM Journal on Applied Mathematics, 55(6):1697–1723, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25795.

Colton:1995:EFF

[3224] David Colton and Rainer Kress. Eigenvalues of the far field operator for the Helmholtz equation in an absorbing medium. SIAM Journal on Applied Mathematics, 55(6):1724–1735, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25611.

Hasanov:1995:IPE

[3225] Alemdar Hasanov. An inverse problem for an elastoplastic medium. SIAM Journal on Applied Mathematics, 55(6):1736–1752, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26749.

Zharnitsky:1995:GDN

[3226] V. Zharnitsky. The geometrical description of the nonlinear dynamics of a multiple pendulum. SIAM Journal on Applied Mathematics, 55(6):1753–1763, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/25660.

Chow:1995:DDN

[3227] Shui-Nee Chow and Wen Xian Shen.

Dynamics in a discrete Nagumo equation: Spatial topological chaos. SIAM Journal on Applied Mathematics, 55(6):1764–1781, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26175.

Galias:1995:DTN

[3228] Zbigniew Galias. On a discretetime nonlinear system associated with the second-order digital filter.

SIAM Journal on Applied Mathematics, 55(6):1782–1799, December 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/25185.

Drazin:1996:NSB

[3229] P. G. Drazin and Y. Tourigny. Numerical study of bifurcations by analytic continuation of a function defined by a power series. SIAM Journal on Applied Mathematics, 56(1):1–18, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27243.

Khuri:1996:BSS

[3230] S. A. Khuri. Biorthogonal series solution of Stokes flow problems in sectorial regions. SIAM Journal on Applied Mathematics, 56(1):19–39, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mezic:1996:MED

[3231] Igor Mezić, John F. Brady, and Stephen Wiggins. Maximal effective diffusivity for time-periodic incompressible fluid flows. SIAM Journal on Applied Mathematics, 56(1):40–56, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/27044.

Roe:1996:NEM

[3232] P. L. Roe and D. S. Balsara. Notes on the eigensystem of magnetohydrodynamics. SIAM Journal on Applied Mathematics, 56(1):57–67, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27084.

Figotin:1996:BGSa

[3233] Alex Figotin and Peter Kuchment.

Band-gap structure of spectra of periodic dielectric and acoustic media. I. Scalar model. SIAM Journal on Applied Mathematics, 56(1):68–88, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26385.

Pawell:1996:DET

[3234] Angela Pawell and Klaus-Dieter Krannich. Dissolution effects in transport in porous media. SIAM Journal on Applied Mathematics, 56(1):89–118, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/24951.

Courtemanche:1996:DER

[3235] Marc Courtemanche, James P. Keener, and Leon Glass. A delay equation representation of pulse circulation on a ring in excitable media. SIAM Journal on Applied Mathematics, 56(1):119–142, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25933.

Kirschner:1996:DMA

[3236] Denise E. Kirschner. A diffusion model for AIDS in a closed, heterosexual population: Examining rates of infection. SIAM Journal on Applied Mathematics, 56(1):143–166, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/23911.

Katsevich:1996:PT

[3237] Alexander I. Katsevich and Alexander G. Ramm. Pseudolocal tomography. SIAM Journal on Applied Mathematics, 56(1):167–191, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/26611.

Katsevich:1996:AIG

[3238] Alexander I. Katsevich and Alexander G. Ramm. Approximate inverse geophysical scattering on a small body. SIAM Journal on Applied Mathematics, 56(1):192–218, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/26344.

Aktosun:1996:ISS

[3239] Tuncay Aktosun. Inverse Schrödinger scattering on the line with partial knowledge of the potential. SIAM Journal on Applied Mathematics, 56(1):219–231, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/27399.

Ram:1996:IEP

[3240] Yitshak M. Ram and Sylvan Elhay. An inverse eigenvalue problem for the symmetric tridiagonal quadratic pencil with application to damped oscillatory systems. SIAM Journal on Applied Mathematics, 56(1):232–244, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26700.

Belair:1996:FSD

[3241] Jacques Belair, Sue Ann Campbell, and P. van den Driessche. Frustration, stability, and delay-induced oscillations in a neural network model. SIAM Journal on Applied Mathematics, 56(1):245–255, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/27452.

Lewicki:1996:BEM

[3242] Pawel Lewicki, Robert Burridge, and Maarten V. de Hoop. Beyond effective medium theory: Pulse stabilization for multimode wave propagation in high-contrast layered media. SIAM Journal on Applied Mathematics, 56(1):256–276, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27366.

Khasminskii:1996:AES

[3243] R. Z. Khasminskii, G. Yin, and Q. Zhang. Asymptotic expansions of singularly perturbed systems involving rapidly fluctuating Markov chains. SIAM Journal on Applied Mathematics, 56(1):277–293, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/25993.

Ballester:1996:AIS

[3244] C. Ballester, V. Caselles, and M. González. Affine invariant segmentation by variational method. SIAM Journal on Applied Mathematics, 56(1):294–325, February 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/26702.

Russo:1996:KTBa

[3245] Giovanni Russo and Peter Smereka. Kinetic theory for bubbly flow I: Collisionless case. SIAM Journal on Applied Mathematics, 56(2):327–357, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/26056.

Russo:1996:KTBb

[3246] Giovanni Russo and Peter Smereka. Kinetic theory for bubbly flow. II. Fluid dynamic limit. SIAM Journal on Applied Mathematics, 56(2):358–371, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/26057.

Gomez:1996:RVI

[3247] Juan C. Gomez and Vianey Villamizar. Response of a viscous incompressible fluid cylinder to an incident plane compressional elastic wave. SIAM Journal on Applied Mathematics, 56(2):372–387, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/26428.

Diehl:1996:CLP

[3248] Stefan Diehl. A conservation law with point source and discontinuous flux function modelling continuous sedimentation. SIAM Journal on Applied Mathematics, 56(2):388–419, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/24242.

Bosley:1996:IMP

[3249] D. L. Bosley. An improved matching procedure for transient resonance layers in weakly nonlinear oscillatory systems. SIAM Journal on Applied Mathematics, 56(2):420–445, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26677.

Barcilon:1996:SPA

[3250] Victor Barcilon. Singular perturbation analysis of the Fokker–Planck equation: Kramers’ underdamped problem. SIAM Journal on Applied Mathematics, 56(2):446–479, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/25511.

Olives:1996:CCE

[3251] Juan Olives. A combined capillarity and elasticity problem for a thin plate. SIAM Journal on Applied Mathematics, 56(2):480–493, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26316.

Castillo-Chavez:1996:CEG

[3252] Carlos Castillo-Chavez, Wenzhang Huang, and Jia Li. Competitive exclusion in gonorrhea models and other sexually transmitted diseases. SIAM Journal on Applied Mathematics, 56(2):494–508, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/25419.

Gomulkiewicz:1996:SGI

[3253] Richard Gomulkiewicz and Jay H. Beder. The selection gradient of an infinite-dimensional trait. SIAM Journal on Applied Mathematics, 56 (2):509–523, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/
25576.

Lamm:1996:AIP

[3254] Patricia K. Lamm. Approximation of ill-posed Volterra problems via predictor-corrector regularization methods. SIAM Journal on Applied Mathematics, 56(2):524–541, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/27449.

Chavent:1996:WNI

[3255] G. Chavent and K. Kunisch. On weakly nonlinear inverse problems. SIAM Journal on Applied Mathematics, 56(2):542–572, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26744.

Antyufeev:1996:XRT

[3256] Victor S. Antyufeev and Anatoliy N.

Bondarenko. X-ray tomography in scattering media. SIAM Journal on Applied Mathematics, 56(2):573–587, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26639.

Hettlich:1996:UIC

[3257] F. Hettlich. Uniqueness of the inverse conductive scattering problem for time-harmonic electromagnetic waves. SIAM Journal on Applied Mathematics, 56(2):588–601, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27382.

Nakamura:1996:NTI

[3258] Gen Nakamura and Kazumi Tanuma.

A nonuniqueness theorem for an inverse boundary value problem in elasticity. SIAM Journal on Applied Mathematics, 56(2):602–610, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26630.

Kutz:1996:SPN

[3259] J. Nathan Kutz and William L. Kath.

Stability of pulses in nonlinear optical fibers using phase-sensitive amplifiers. SIAM Journal on Applied Mathematics, 56(2):611–626, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/27760.

Grorud:1996:ALE

[3260] Axel Grorud and Denis Talay. Approximation of Lyapunov exponents of nonlinear stochastic differential equations. SIAM Journal on Applied Mathematics, 56(2):627–650, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/22771.

Lopes:1996:OBI

[3261] Artur Lopes and Roberto Markarian. Open billiards: Invariant and conditionally invariant probabilities on Cantor sets. SIAM Journal on Applied Mathematics, 56(2):651–680, April 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/27943.

Bertozzi:1996:SSF

[3262] Andrea L. Bertozzi. Symmetric singularity formation in lubricationtype equations for interface motion. SIAM Journal on Applied Mathematics, 56(3):681–714, June 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/27197.

Bryan:1996:IPT

[3263] Kurt Bryan and Lester F. Caudill, Jr. An inverse problem in thermal imaging. SIAM Journal on Applied Mathematics, 56(3):715–735, June 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/27782.

Sylvester:1996:LSH

[3264] John Sylvester, Dale Winebrenner, and Fred Gylys-Colwell. Layer stripping for the Helmholtz equation. SIAM Journal on Applied Mathematics, 56(3):736–754, June 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28025.

Berenstein:1996:IGH

[3265] Carlos A. Berenstein and Enrico Casadio Tarabusi. Integral geometry in hyperbolic spaces and electrical impedance tomography. SIAM Journal on Applied Mathematics, 56(3):755–764, June 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/27734.

Vance:1996:MLG

[3266] R. W. C. Vance. Matrix Lie group-theoretic design of coupled linear optical waveguide devices. SIAM Journal on Applied Mathematics, 56(3):765–782, June 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26992.

David:1996:AMM

[3267] Guy David. C^{1}-arcs for minimizers of the Mumford–Shah functional. SIAM Journal on Applied Mathematics, 56(3):783–888, June 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/27607.

Maree:1996:SPT

[3268] G. J. M. Marée. Slow passage through a pitchfork bifurcation. SIAM Journal on Applied Mathematics, 56(3):889–918, June 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25739.

Colonius:1996:SMM

[3269] Fritz Colonius, F. Javier de la Rubia, and Wolfgang Kliemann. Stochastic models with multistability and extinction levels. SIAM Journal on Applied Mathematics, 56(3):919–945, June 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/27770.

Baker:1996:ICC

[3270] C. R. Baker and I.-F. Chao. Information capacity of channels with partially unknown noise I. finite-dimensional channels. SIAM Journal on Applied Mathematics, 56(3):946–963, June 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/24985.

Dawson:1996:LTA

[3271] C. N. Dawson, C. J. Van Duijn, and R. E. Grundy. Large time asymptotics in contaminant transport in porous media. SIAM Journal on Applied Mathematics, 56(4):965–993, August 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/27629.

Bruaset:1996:SPV

[3272] Are Magnus Bruaset and Bjørn Fredrik Nielsen. On the stability of pressure and velocity computations for heterogeneous reservoirs. SIAM Journal on Applied Mathematics, 56(4):994–1007, August 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26669.

Saichev:1996:DFB

[3273] A. I. Saichev and W. A. Woyczynski.

Density fields in Burgers and KdV– Burgers turbulence. SIAM Journal on Applied Mathematics, 56(4):1008–1038, August 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26647.

Dahm:1996:IMS

[3274] Werner J. A. Dahm, Gretar Tryggvason, and Mei Zhuang. Integral method solution of time-dependent strained diffusion-reaction layers with multistep kinetics. SIAM Journal on Applied Mathematics, 56(4):1039–1059, August 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26584.

Du:1996:HKL

[3275] Qiang Du and Paul Gray. High-kappa limits of the time-dependent Ginzburg– Landau model. SIAM Journal on Applied Mathematics, 56(4):1060–1093, August 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28050.

Mathar:1996:PCD

[3276] Rudolf Mathar and Jürgen Mattfeldt. Pulse-coupled decentral synchronization. SIAM Journal on Applied Mathematics, 56(4):1094–1106, August 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/27813.

Ermentrout:1996:RWI

[3277] G. Bard Ermentrout and John Rinzel. Reflected waves in an inhomogeneous excitable medium. SIAM Journal on Applied Mathematics, 56(4):1107–1128, August 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/27679.

Ola:1996:EIP

[3278] Petri Ola and Erkki Somersalo. Electromagnetic inverse problems and generalized Sommerfeld potentials. SIAM Journal on Applied Mathematics, 56(4):1129–1145, August 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/28394.

Sun:1996:RSM

[3279] Zhiming Sun. Reconstruction of source and medium parameters via wavesplitting and Green function equations. SIAM Journal on Applied Mathematics, 56(4):1146–1163, August 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/25058.

Kim:1996:STP

[3280] Jeong-Hoon Kim. Stochastic turning point problem in a one-dimensional refractive random multilayer. SIAM Journal on Applied Mathematics, 56 (4):1164–1180, August 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/
27781.

Dobson:1996:RBI

[3281] David C. Dobson and Fadil Santosa.

Recovery of blocky images from noisy and blurred data. SIAM Journal on Applied Mathematics, 56(4):1181–1198, August 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27560.

Caselles:1996:WBC

[3282] V. Caselles and C. Sbert. What is the best causal scale space for threedimensional images? SIAM Journal on Applied Mathematics, 56(4):1199–1246, August 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26935.

Ward:1996:MBS

[3283] Michael J. Ward. Metastable bubble solutions for the Allen–Cahn equation with mass conservation. SIAM Journal on Applied Mathematics, 56(5):1247–1279, October 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28291.

Babovsky:1996:DLF

[3284] Hans Babovsky. Diffusion limits for flows in thin layers. SIAM Journal on Applied Mathematics, 56(5):1280–1294, October 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/27082.

Short:1996:UGE

[3285] Mark Short and J. William Dold. Unsteady gasdynamic evolution of an induction domain between a contact surface and a shock wave. I: Thermal runaway. SIAM Journal on Applied Mathematics, 56(5):1295–1316, October 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27630.

Cox:1996:OTC

[3286] Stephen M. Cox. The onset of thermal convection between poorly conducting horizontal boundaries in the presence of a shear flow. SIAM Journal on Applied Mathematics, 56(5):1317–1328, October 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28159.

Ikehata:1996:LDN

[3287] Masaru Ikehata. The linearization of the Dirichlet-to-Neumann map in the anisotropic Kirchhoff–Love plate theory. SIAM Journal on Applied Mathematics, 56(5):1329–1352, October 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27043.

Cox:1996:PCD

[3288] Steven J. Cox and Michael L. Overton.

Perturbing the critically damped wave equation. SIAM Journal on Applied Mathematics, 56(5):1353–1362, October 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27740.

Tajdari:1996:MSP

[3289] Mohammad Tajdari. The moving singularities of the perturbation expansion of the classical Kepler problem. SIAM Journal on Applied Mathematics, 56(5):1363–1378, October 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/26784.

Shen:1996:LLH

[3290] Wenxian Shen. Lifted lattices, hyperbolic structures, and topological disorders in coupled map lattices. SIAM Journal on Applied Mathematics, 56(5):1379–1399, October 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28267.

Namachchivaya:1996:MLE

[3291] N. Sri Namachchivaya, H. J. Van Roessel, and M. M. Doyle. Moment Lyapunov exponent for two coupled oscillators driven by real noise. SIAM Journal on Applied Mathematics, 56(5):1400–1423, October 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28138.

Mair:1996:SIE

[3292] Bernard A. Mair and Frits H. Ruymgaart. Statistical inverse estimation in Hilbert scales. SIAM Journal on Applied Mathematics, 56(5):1424–1444, October 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/26447.

Bailly:1996:PGW

[3293] F. Bailly, J. F. Clouet, and J. P. Fouque. Parabolic and Gaussian white noise approximation for wave propagation in random media. SIAM Journal on Applied Mathematics, 56(5):1445–1470, October 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/28024.

Knessl:1996:EAS

[3294] Charles Knessl and Craig Steven Peters. Exact and asymptotic solutions for the time-dependent problem of collective ruin II. SIAM Journal on Applied Mathematics, 56(5):1471–1521, October 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/27564.

Schwendeman:1996:FDA

[3295] D. W. Schwendeman. A front dynamics approach to curvature-dependent flow. SIAM Journal on Applied Mathematics, 56(6):1523–1538, December 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28048.

Choksi:1996:CLD

[3296] Rustum Choksi. The conservation law ∂_{y}u + ∂_{x}\sqrt{1 − u^{2}} = 0 and deformations of fibre-reinforced materials. SIAM Journal on Applied Mathematics, 56(6):1539–1560, December 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27153.

Figotin:1996:BGSb

[3297] A. Figotin and P. Kuchment. Bandgap structure of spectra of periodic dielectric and acoustic media. II. Two-dimensional photonic crystals. SIAM Journal on Applied Mathematics, 56(6):1561–1620, December 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28523.

Elgin:1996:SCS

[3298] John N. Elgin and Xuesong Wu. Stability of cellular states of the Kuramoto– Sivashinsky equation. SIAM Journal on Applied Mathematics, 56(6): 1621–1638, December 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/
26368.

Brosilow:1996:NST

[3299] Benjamin J. Brosilow, Roseanne M.

Ford, Sten Sarman, and Peter T. Cummings. Numerical solution of transport equations for bacterial chemotaxis: Effect of discretization of directional motion. SIAM Journal on Applied Mathematics, 56(6):1639–1663, December 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28202.

Bendali:1996:ETC

[3300] A. Bendali and K. Lemrabet. The effect of a thin coating on the scattering of a time-harmonic wave for the Helmholtz equation. SIAM Journal on Applied Mathematics, 56(6):1664–1693, December 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28182.

Blackmore:1996:GFD

[3301] Denis Blackmore and Jack G. Zhou. A general fractal distribution function for rough surface profiles. SIAM Journal on Applied Mathematics, 56(6):1694–1719, December 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/28312.

Yagasaki:1996:MTS

[3302] Kazuyuki Yagasaki. The Melnikov theory for subharmonics and their bifurcations in forced oscillations. SIAM Journal on Applied Mathematics, 56(6):1720–1765, December 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/28131.

Khasminskii:1996:ASS

[3303] R. Z. Khasminskii and G. Yin. Asymptotic series for singularly perturbed Kolmogorov–Fokker–Planck equations. SIAM Journal on Applied Mathematics, 56(6):1766–1793, December 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27008.

Khasminskii:1996:TDS

[3304] R. Z. Khasminskii and G. Yin. On transition densities of singularly perturbed diffusions with fast and slow components. SIAM Journal on Applied Mathematics, 56(6):1794–1819, December 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28290.

Wang:1997:ISS

[3305] C. Y. Wang, Q. Du, M. Miklavčič, and C. C. Chang. Impulsive stretching of a surface in a viscous fluid. SIAM Journal on Applied Mathematics, 57(1):1–14, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28205.

Nielsen:1997:ASP

[3306] Bjørn Fredrik Nielsen and Aslak Tveito. On the approximation of the solution of the pressure equation by changing the domain. SIAM Journal on Applied Mathematics, 57(1):15–33, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28086.

Ding:1997:SBT

[3307] Zhonghai Ding and Dimitris C.

Lagoudas. Solution behavior of the transient heat transfer problem in thermoelectric shape memory alloy actuators. SIAM Journal on Applied Mathematics, 57(1):34–52, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28366.

Rybka:1997:CMU

[3308] Piotr Rybka. A crystalline motion: Uniqueness and geometric properties. SIAM Journal on Applied Mathematics, 57(1):53–72, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27990.

Sneyd:1997:PCW

[3309] James Sneyd and Jonathan Sherratt. On the propagation of calcium waves in an inhomogeneous medium. SIAM Journal on Applied Mathematics, 57(1):73–94, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28603.

Garnier:1997:HPT

[3310] Josselin Garnier. Homogenization in a periodic and time-dependent potential. SIAM Journal on Applied Mathematics, 57(1):95–111, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28200.

Eckhaus:1997:PFS

[3311] Wiktor Eckhaus and Rachel Kuske. Pattern formation in systems with slowly varying geometry. SIAM Journal on Applied Mathematics, 57(1): 112–152, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/
27753.

Alvarez:1997:IQU

[3312] Luis Alvarez and Julio Esclarín. Image quantization using reaction-diffusion equations. SIAM Journal on Applied Mathematics, 57(1):153–175, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27758.

Olver:1997:IGE

[3313] Peter J. Olver, Guillermo Sapiro, and Allen Tannenbaum. Invariant geometric evolutions of surfaces and volumetric smoothing. SIAM Journal on Applied Mathematics, 57(1):176–194, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26631.

Liu:1997:UIP

[3314] Jijun Liu and Yuanming Wang. On uniqueness of an inverse problem for a 1-D wave equation from transmission data. SIAM Journal on Applied Mathematics, 57(1):195–204, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27606.

Balmforth:1997:PDU

[3315] N. J. Balmforth, G. R. Ierley, and R. Worthing. Pulse dynamics in an unstable medium. SIAM Journal on Applied Mathematics, 57(1):205–251, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/28408.

Terman:1997:PSN

[3316] David Terman and Euiwoo Lee. Partial synchronization in a network of neural oscillators. SIAM Journal on Applied Mathematics, 57(1):252–293, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/27892.

Qian:1997:DCO

[3317] Min Qian, Shu Zhu, and Wen-Xin Qin. Dynamics in a chain of overdamped pendula driven by constant torques. SIAM Journal on Applied Mathematics, 57(1):294–305, February 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/27875.

Short:1997:MLS

[3318] Mark Short. Multidimensional linear stability of a detonation wave at high activation energy. SIAM Journal on Applied Mathematics, 57(2):307–326, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28810.

Capasso:1997:ARD

[3319] V. Capasso and R. E. Wilson. Analysis of a reaction-diffusion system modeling man- environment-man epidemics. SIAM Journal on Applied Mathematics, 57(2):327–346, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28468.

Lipton:1997:RRB

[3320] Robert Lipton. Reciprocal relations, bounds, and size effects for composites with highly conducting interface. SIAM Journal on Applied Mathematics, 57(2):347–363, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29118.

Bao:1997:VAM

[3321] Gang Bao. Variational approximation of Maxwell’s equations in biperiodic structures. SIAM Journal on Applied Mathematics, 57(2):364–381, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27940.

Kriegsmann:1997:CEM

[3322] G. A. Kriegsmann. Cavity effects in microwave heating of ceramics. SIAM Journal on Applied Mathematics, 57(2):382–400, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/27866.

Forbes:1997:CCD

[3323] Lawrence K. Forbes, Stuart Crozier, and David M. Doddrell. Calculating current densities and fields produced by shielded magnetic resonance imaging probes. SIAM Journal on Applied Mathematics, 57(2):401–425, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28311.

Kuruc:1997:EEL

[3324] Alvin Kuruc. Efficient estimation of linear functionals in emission tomography. SIAM Journal on Applied Mathematics, 57(2):426–452, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/28320.

Athanassoulis:1997:EAL

[3325] Gerasimos A. Athanassoulis and Vassilis G. Papanicolaou. Eigenvalue asymptotics of layered media and their applications to the inverse problem. SIAM Journal on Applied Mathematics, 57(2):453–471, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/28360.

Brodsky:1997:AIE

[3326] Mikhail Brodsky. An asymptotical inversion of the eigenfrequencies for a three-dimensional problem with spherical symmetry. SIAM Journal on Applied Mathematics, 57(2):472–484, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28453.

Ogawa:1997:SPN

[3327] Toshiyuki Ogawa and Hiromasa Suzuki. On the spectra of pulses in a nearly integrable system. SIAM Journal on Applied Mathematics, 57 (2):485–500, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/
28878.

Dimitrov:1997:PDP

[3328] Boyan Dimitrov, Stefanka Chukova, and David Green, Jr. Probability distributions in periodic random environment and their applications. SIAM Journal on Applied Mathematics, 57(2):501–517, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26384.

Cohen:1997:NAR

[3329] Albert Cohen and Jean-Pierre D’Ales.

Nonlinear approximation of random functions. SIAM Journal on Applied Mathematics, 57(2):518–540, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/27915.

Abrahams:1997:SWH

[3330] I. David Abrahams. On the solution of Wiener–Hopf problems involving noncommutative matrix kernel decompositions. SIAM Journal on Applied Mathematics, 57(2):541–567, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28767.

Chakrabarti:1997:DGA

[3331] A. Chakrabarti and A. J. George. Diagonalizable generalized Abel integral operators. SIAM Journal on Applied Mathematics, 57(2):568–575, April 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/27967.

Trevino:1997:AAT

[3332] C. Treviño, A. Espinoza, and F. Méndez. Asymptotic analysis of the transient conjugate heat transfer process between two forced counterflowing streams. SIAM Journal on Applied Mathematics, 57(3):577–596, June 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28448.

Impey:1997:CNM

[3333] M. D. Impey, P. Grindrod, H. Takase, and K. J. Worgan. A capillary network model for gas migration in lowpermeability media. SIAM Journal on Applied Mathematics, 57(3):597–608, June 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28995.

Park:1997:QPS

[3334] J.-H. Park and J. W. Jerome. Qualitative properties of steady-state Poisson– Nernst–Planck systems: Mathematical study. SIAM Journal on Applied Mathematics, 57(3):609–630, June 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/27980.

Barcilon:1997:QPS

[3335] V. Barcilon, D-P Chen, R. S. Eisenberg, and J. W. Jerome. Qualitative properties of steady-state Poisson–Nernst–Planck systems: Perturbation and simulation study. SIAM Journal on Applied Mathematics, 57(3):631–648, June 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/31214.

Frigaard:1997:SPB

[3336] I. A. Frigaard and O. Scherzer. Spraying the perfect billet. SIAM Journal on Applied Mathematics, 57(3):649–682, June 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28713.

Levine:1997:SRD

[3337] Howard A. Levine and Brian D. Sleeman. A system of reaction diffusion equations arising in the theory of reinforced random walks. SIAM Journal on Applied Mathematics, 57(3):683–730, June 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29110.

Dai:1997:RCD

[3338] Liyi Dai. Rate of convergence for derivative estimation of discrete-time Markov chains via finite-difference approximation with common random numbers. SIAM Journal on Applied Mathematics, 57(3):731–751, June 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/27173.

Maier:1997:LEL

[3339] Robert S. Maier and Daniel L. Stein. Limiting exit location distributions in the stochastic exit problem. SIAM Journal on Applied Mathematics, 57(3):752–790, June 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/27175.

Tan:1997:IRAa

[3340] Xiaoqian Tan and Charles Knessl. Integral representations and asymptotics for infinite- and finite-capacity queues described by the unfinished work I. SIAM Journal on Applied Mathematics, 57(3):791–823, June 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/27673.

Tan:1997:IRAb

[3341] Xiaoqian Tan and Charles Knessl. Integral representations and asymptotics for infinite- and finite-capacity queues described by the unfinished work. II. SIAM Journal on Applied Mathematics, 57(3):824–870, June 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/27680.

Hsieh:1997:PTV

[3342] Din-Yu Hsieh and Xiao-Ping Wang. Phase transition in van der Waals fluid. SIAM Journal on Applied Mathematics, 57(4):871–892, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29516.

Hui:1997:HOC

[3343] W. H. Hui and Yuanping He. Hyperbolicity and optimal coordinates for the three-dimensional supersonic Euler equations. SIAM Journal on Applied Mathematics, 57(4):893–928, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28271.

Hall:1997:MIA

[3344] Philip Hall and Demetrios T. Papageorgiou. On the modulational instability of O(1) amplitude waves in supersonic boundary layers. SIAM Journal on Applied Mathematics, 57(4):929– 958, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/28576.

DiLellio:1997:SBF

[3345] J. A. DiLellio and W. E. Olmstead.

Shear band formation due to a thermal flux inhomogeneity. SIAM Journal on Applied Mathematics, 57(4):959–971, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28959.

Knio:1997:VFO

[3346] Omar M. Knio and Lu Ting. Vortical flow outside a sphere and sound generation. SIAM Journal on Applied Mathematics, 57(4):972–981, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/29397.

Otto:1997:VFO

[3347] Felix Otto. Viscous fingering: An optimal bound on the growth rate of the mixing zone. SIAM Journal on Applied Mathematics, 57(4):982–990, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/29438.

Diehl:1997:DSS

[3348] Stefan Diehl. Dynamic and steadystate behavior of continuous sedimentation. SIAM Journal on Applied Mathematics, 57(4):991–1018, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/29010.

Wolkowicz:1997:GAB

[3349] Gail S. K. Wolkowicz and Huaxing Xia. Global asymptotic behavior of a chemostat model with discrete delays. SIAM Journal on Applied Mathematics, 57(4):1019–1043, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28731.

Othmer:1997:ABC

[3350] Hans G. Othmer and Angela Stevens. Aggregation, blowup, and collapse: The ABC’s of taxis in reinforced random walks. SIAM Journal on Applied Mathematics, 57(4):1044–1081, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28897.

Hyman:1997:BCS

[3351] James M. Hyman and Jia Li. Behavior changes in SIS STD models with selective mixing. SIAM Journal on Applied Mathematics, 57(4):1082–1094, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29412.

Faridani:1997:LTI

[3352] Adel Faridani, David V. Finch, Erik L.

Ritman, and Kennan T. Smith. Local tomography. II. SIAM Journal on Applied Mathematics, 57(4):1095–1127, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28635.

Katsevich:1997:LTG

[3353] Alexander I. Katsevich. Local tomography for the generalized radon transform. SIAM Journal on Applied Mathematics, 57(4):1128–1162, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29184.

Goldsztein:1997:DHS

[3354] Guillermo H. Goldsztein, Fernando Broner, and Steven H. Strogatz. Dynamical hysteresis without static hysteresis: Scaling laws and asymptotic expansions. SIAM Journal on Applied Mathematics, 57(4):1163–1187, August 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29073.

Marchesin:1997:OCW

[3355] Dan Marchesin, Bradley J. Plohr, and Stephen Schecter. An organizing center for wave bifurcation in multiphase flow models. SIAM Journal on Applied Mathematics, 57(5):1189–1215, October 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28068.

Rosencrans:1997:TDC

[3356] Steve Rosencrans. Taylor dispersion in curved channels. SIAM Journal on Applied Mathematics, 57(5):1216–1241, October 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/26990.

Short:1997:CCI

[3357] Mark Short. On the critical conditions for the initiation of a detonation in a nonuniformly perturbed reactive fluid. SIAM Journal on Applied Mathematics, 57(5):1242–1280, October 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28469.

Wolkowicz:1997:CCD

[3358] Gail S. K. Wolkowicz, Huaxing Xia, and Shigui Ruan. Competition in the chemostat: A distributed delay model and its global asymptotic behavior. SIAM Journal on Applied Mathematics, 57(5):1281–1310, October 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28984.

Cantrell:1997:CMC

[3359] Robert Stephen Cantrell and James R.

Ward, Jr. On competition-mediated coexistence. SIAM Journal on Applied Mathematics, 57(5):1311–1327, October 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29236.

Kichenassamy:1997:PMP

[3360] Satyanad Kichenassamy. The Perona– Malik paradox. SIAM Journal on Applied Mathematics, 57(5):1328–1342, October 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/29558.

Bonilla:1997:ABI

[3361] L. L. Bonilla, J. A. Carrillo, and J. Soler. Asymptotic behavior of an initial-boundary value problem for the Vlasov–Poisson–Fokker–Planck system. SIAM Journal on Applied Mathematics, 57(5):1343–1372, October 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29154.

Rikte:1997:EUC

[3362] Sten Rikte. Existence, uniqueness, and causality theorems for wave propagation in stratified, temporally dispersive, complex media. SIAM Journal on Applied Mathematics, 57(5):1373–1389, October 1997.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/27919.

Tassa:1997:HTD

[3363] Tamir Tassa. Homogenization of twodimensional linear flows with integral invariance. SIAM Journal on Applied Mathematics, 57(5):1390–1405, October 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/29982.

Booth:1997:NTB

[3364] Victoria Booth, Thomas W. Carr, and Thomas Erneux. Near-threshold bursting is delayed by a slow passage near a limit point. SIAM Journal on Applied Mathematics, 57(5):1406–1420, October 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/29510.

Huveneers:1997:MAE

[3365] R. J. A. G. Huveneers and F. Verhulst. A metaphor for adiabatic evolution to symmetry. SIAM Journal on Applied Mathematics, 57(5):1421–1442, October 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28956.

Holden:1997:BLE

[3366] Lars Holden. The Buckley–Leverett equation with spatially stochastic flux function. SIAM Journal on Applied Mathematics, 57(5):1443–1454, October 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27307.

Gaines:1997:VSS

[3367] J. G. Gaines and T. J. Lyons. Variable step size control in the numerical solution of stochastic differential equations. SIAM Journal on Applied Mathematics, 57(5):1455–1484, October 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28651.

Cruywagen:1997:BPF

[3368] Gerhard C. Cruywagen, Philip K.

Maini, and James D. Murray. Biological pattern formation on twodimensional spatial domains: A nonlinear bifurcation analysis. SIAM Journal on Applied Mathematics, 57(6):1485–1509, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29790.

Gani:1997:SQH

[3369] J. Gani, S. Yakowitz, and M. Blount.

The spread and quarantine of HIV infection in a prison system. SIAM Journal on Applied Mathematics, 57(6):1510–1530, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28323.

Toth:1997:ELE

[3370] János Tóth, Genyuan Li, Herschel Rabitz, and Alison S. Tomlin. The effect of lumping and expanding on kinetic differential equations. SIAM Journal on Applied Mathematics, 57(6):1531–1556, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29329.

Friedman:1997:AMC

[3371] Avner Friedman and Chaocheng Huang. Averaged motion of charged particles in a curved strip. SIAM Journal on Applied Mathematics, 57(6):1557–1587, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28023.

Bonilla:1997:PGP

[3372] Luis L. Bonilla, Manuel Kindelan, Miguel Moscoso, and Stephanos Venakides. Periodic generation and propagation of traveling fronts in DC voltage biased semiconductor superlattices. SIAM Journal on Applied Mathematics, 57(6):1588–1614, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28888.

Lucas:1997:IES

[3373] S. K. Lucas, R. Sipcic, and H. A. Stone. An integral equation solution for the steady-state current at a periodic array of surface microelectrodes. SIAM Journal on Applied Mathematics, 57(6):1615–1638, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29252.

Conca:1997:HPS

[3374] Carlos Conca and Muthusamy Vanninathan. Homogenization of periodic structures via Bloch decomposition. SIAM Journal on Applied Mathematics, 57(6):1639–1659, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29474.

Nevard:1997:HRB

[3375] J. Nevard and J. B. Keller. Homogenization of rough boundaries and interfaces. SIAM Journal on Applied Mathematics, 57(6):1660–1686, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29108.

MacCamy:1997:LFE

[3376] R. C. MacCamy. Low frequency expansions for two-dimensional interface scattering problems. SIAM Journal on Applied Mathematics, 57(6):1687–1701, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/27988.

Hillion:1997:DWF

[3377] Pierre Hillion. Diffraction and Weber functions. SIAM Journal on Applied Mathematics, 57(6):1702–1715, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27959.

Sheen:1997:AEF

[3378] Dongwoo Sheen. Approximation of electromagnetic fields: Part I. continuous problems. SIAM Journal on Applied Mathematics, 57(6):1716–1736, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28923.

Elayyan:1997:IDP

[3379] Alaeddin Elayyan and Victor Isakov. On an inverse diffusion problem. SIAM Journal on Applied Mathematics, 57(6):1737–1748, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28873.

Zhao:1997:WSL

[3380] Shiying Zhao, Grant Welland, and Ge Wang. Wavelet sampling and localization schemes for the Radon transform in two dimensions. SIAM Journal on Applied Mathematics, 57(6):1749–1762, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29057.

Maes:1997:FQC

[3381] Stéphane H. Maes. Fast quasicontinuous wavelet algorithms for analysis and synthesis of one-dimensional signals. SIAM Journal on Applied Mathematics, 57(6):1763–1801, December 1997. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28801.

tenBosch:1998:VPD

[3382] B. I. M. ten Bosch and A. J.

Weisenborn. On a variational principle for the drag in linear hydrodynamics. SIAM Journal on Applied Mathematics, 58(1):1–14, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29455.

Smoller:1998:SSC

[3383] Joel Smoller and Blake Temple. Shockwave solutions in closed form and the Oppenheimer–Snyder limit in general reality. SIAM Journal on Applied Mathematics, 58(1):15–33, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29793.

Nie:1998:SHF

[3384] Qing Nie and Fei Ran Tian. Singularities in Hele–Shaw flows. SIAM Journal on Applied Mathematics, 58(1):34– 54, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/29792.

Lipton:1998:HCF

[3385] Robert Lipton. Heat conduction in fine scale mixtures with interfacial contact resistance. SIAM Journal on Applied Mathematics, 58(1):55–72, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29515.

Costantini:1998:NAF

[3386] C. Costantini, B. Pacchiarotti, and F. Sartoretto. Numerical approximation for functionals of reflecting diffusion processes. SIAM Journal on Applied Mathematics, 58(1):73–102, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/29104.

Berger:1998:BAP

[3387] Jorge Berger and Jacob Rubinstein. Bifurcation analysis for phase transitions in superconducting rings with nonuniform thickness. SIAM Journal on Applied Mathematics, 58(1):103–121, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29729.

Cheng:1998:MIE

[3388] Hongwei Cheng and Leslie Greengard. A method of images for the evaluation of electrostatic fields in systems of closely spaced conducting cylinders. SIAM Journal on Applied Mathematics, 58(1):122–141, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29761.

Bertram:1998:PDS

[3389] Richard Bertram and Arthur Sherman. Population dynamics of synaptic release sites. SIAM Journal on Applied Mathematics, 58(1):142–169, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29791.

Ruan:1998:GSC

[3390] Shigui Ruan and Xue-Zhong He.

Global stability in chemostat-type competition models with nutrient recycling. SIAM Journal on Applied Mathematics, 58(1):170–192, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29924. See erratum

[3585].

Dai:1998:CRG

[3391] Guoren Dai and Moxun Tang. Coexistence region and global dynamics of a harvested predator-prey system. SIAM Journal on Applied Mathematics, 58(1):193–210, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27579.

Chi:1998:AMM

[3392] Chia-Wei Chi, Sze-Bi Hsu, and Lih-Ing Wu. On the asymmetric May– Leonard model of three competing species. SIAM Journal on Applied Mathematics, 58(1):211–226, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27206.

vandenDriessche:1998:TDC

[3393] P. van den Driessche and M. L. Zeeman. Three-dimensional competitive Lotka–Volterra systems with no periodic orbits. SIAM Journal on Applied Mathematics, 58(1):227–234, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/29476.

Ocken:1998:CCA

[3394] Stanley Ocken. Convergence criteria for attracting cycles of Newton’s method. SIAM Journal on Applied Mathematics, 58(1):235–244, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28846.

Khasminskii:1998:MLE

[3395] R. Khasminskii and N. Moshchuk. Moment Lyapunov exponent and stability index for linear conservative system with small random perturbation. SIAM Journal on Applied Mathematics, 58(1):245–256, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/29589.

Knessl:1998:HTA

[3396] Charles Knessl and Charles Tier. Heavy traffic analysis of a markovmodulated queue with finite capacity and general service times. SIAM Journal on Applied Mathematics, 58(1):257–323, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/29385.

March:1998:APM

[3397] Riccardo March and Piero Barone. Application of the Padé method to solving the noisy trigonometric moment problem: Some initial results. SIAM Journal on Applied Mathematics, 58(1):324–343, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/26874.

Ryzhov:1998:BNT

[3398] Oleg S. Ryzhov, Julian D. Cole, and Norman D. Malmuth. A bluntnosed thin body in hypersonic flow. SIAM Journal on Applied Mathematics, 58(2):345–369, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29949.

Charalambopoulos:1998:ASL

[3399] A. Charalambopoulos, G. Dassios, and M. Hadjinicolaou. An analytic solution for low-frequency scattering by two soft spheres. SIAM Journal on Applied Mathematics, 58(2):370–386, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30408.

Fennemore:1998:WFO

[3400] George Fennemore and Jack X. Xin.

Wetting fronts in one-dimensional periodically layered soils. SIAM Journal on Applied Mathematics, 58(2):387–427, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29274.

Gray:1998:CTI

[3401] L. J. Gray and Glaucio H. Paulino. Crack tip interpolation, revisited. SIAM Journal on Applied Mathematics, 58(2):428–455, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/27916.

Ida:1998:DTFa

[3402] M. P. Ida and M. J. Miksis. The dynamics of thin films I: General theory. SIAM Journal on Applied Mathematics, 58(2):456–473, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/28955.

Ida:1998:DTFb

[3403] M. P. Ida and M. J. Miksis. The dynamics of thin films II: Applications. SIAM Journal on Applied Mathematics, 58(2):474–500, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/29117.

Borcea:1998:NAT

[3404] Liliana Borcea and George C. Papanicolaou. Network approximation for transport properties of high contrast materials. SIAM Journal on Applied Mathematics, 58(2):501–539, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/30189.

Zhang:1998:LSI

[3405] Xiaojing Zhang and Alexander L.

Frenkel. Large-scale instability of generalized oscillating Kolmogorov flows. SIAM Journal on Applied Mathematics, 58(2):540–564, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/30527.

Younes:1998:CED

[3406] Laurent Younes. Computable elastic distances between shapes. SIAM Journal on Applied Mathematics, 58 (2):565–586, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/
28768.

Chapman:1998:MHV

[3407] S. J. Chapman and G. Richardson. Motion and homogenization of vortices in anisotropic type II superconductors. SIAM Journal on Applied Mathematics, 58(2):587–606, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28268.

Vries:1998:ACM

[3408] Gerda de Vries and Robert M. Miura.

Analysis of a class of models of bursting electrical activity in pancreatic β-cells. SIAM Journal on Applied Mathematics, 58(2):607–635, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30159.

Canic:1998:RPT

[3409] Suncica Canic and Barbara Lee Keyfitz. Riemann problems for the two-dimensional unsteady transonic small disturbance equation. SIAM Journal on Applied Mathematics, 58(2):636–665, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/30087.

Potthast:1998:CUI

[3410] Roland Potthast. On a concept of uniqueness in inverse scattering for a finite number of incident waves. SIAM Journal on Applied Mathematics, 58(2):666–682, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/30457.

Figotin:1998:SPC

[3411] A. Figotin and P. Kuchment. Spectral properties of classical waves in high-contrast periodic media. SIAM Journal on Applied Mathematics, 58(2):683–702, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/29724.

Anderssen:1998:GBE

[3412] Robert S. Anderssen, Ivan G. Götz, and Karl-Heinz Hoffmann. The global behavior of elastoplastic and viscoelastic materials with hysteresis-type state equations. SIAM Journal on Applied Mathematics, 58(2):703–723, April 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30088.

Kirschner:1998:SAM

[3413] Denise Kirschner and Jeff Morgan. Solution of the asymmetric mixing problem. SIAM Journal on Applied Mathematics, 58(3):725–736, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/27752.

Gobbert:1998:AAM

[3414] Matthias K. Gobbert and Christian A.

Ringhofer. An asymptotic analysis for a model of chemical vapor deposition on a microstructured surface. SIAM Journal on Applied Mathematics, 58(3):737–752, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28467.

Piccioni:1998:VPA

[3415] Mauro Piccioni, Sergio Scarlatti, and Alain Trouvé. A variational problem arising from speech recognition. SIAM Journal on Applied Mathematics, 58(3):753–771, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29253.

Bergman:1998:LCD

[3416] David J. Bergman. A limiting case of diffusion in a composite medium. SIAM Journal on Applied Mathematics, 58(3):772–779, June 1998.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/30260.

Gardner:1998:ATE

[3417] Carl L. Gardner and Christian Ringhofer. Approximation of thermal equilibrium for quantum gases with discontinuous potentials and application to semiconductor devices. SIAM Journal on Applied Mathematics, 58(3):780–805, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/30390.

Schult:1998:TWN

[3418] D. A. Schult, A. Bayliss, and B. J. Matkowsky. Traveling waves in natural counterflow filtration combustion and their stability. SIAM Journal on Applied Mathematics, 58(3):806–852, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30475.

Ahmed:1998:MAD

[3419] N. U. Ahmed and H. Harbi. Mathematical analysis of dynamic models of suspension bridges. SIAM Journal on Applied Mathematics, 58(3):853–874, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30869.

Lin:1998:DRD

[3420] Zhongyan Lin. On the determination of radially dependent Lamé coefficients. SIAM Journal on Applied Mathematics, 58(3):875–903, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/30303.

Varley:1998:SDS

[3421] Eric Varley and Brian R. Seymour. A simple derivation of the N-soliton solutions to the Korteweg–de Vries equation. SIAM Journal on Applied Mathematics, 58(3):904–911, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/30327.

Hyman:1998:MEI

[3422] James M. Hyman and Jia Li. Modeling the effectiveness of isolation strategies in preventing STD epidemics. SIAM Journal on Applied Mathematics, 58(3):912–925, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/30561.

Colton:1998:LSM

[3423] David Colton and Peter Monk. A linear sampling method for the detection of leukemia using microwaves. SIAM Journal on Applied Mathematics, 58(3):926–941, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/30800.

Cherepanov:1998:TDC

[3424] G. P. Cherepanov. Two-dimensional convective heat/mass transfer for low Prandtl and any Peclet numbers.

SIAM Journal on Applied Mathematics, 58(3):942–960, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/28231.

Baginski:1998:PSM

[3425] Frank Baginski, William Collier, and Tami Williams. A parallel shooting method for determining the natural shape of a large scientific balloon. SIAM Journal on Applied Mathematics, 58(3):961–974, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/29235.

Maybank:1998:BOT

[3426] S. J. Maybank. Bearings-only tracking in the plane. SIAM Journal on Applied Mathematics, 58(3):975–998, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/29086.

Ovaskainen:1998:EMA

[3427] O. Ovaskainen and J. Pitkäranta. An energy method approach to the problem of elastic strip. SIAM Journal on Applied Mathematics, 58(3):999–1021, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/27057.

Ammari:1998:LFB

[3428] H. Ammari, M. Laouadi, and J.- C. Nédélec. Low frequency behavior of solutions to electromagnetic scattering problems in chiral media. SIAM Journal on Applied Mathematics, 58(3):1022–1042, June 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31043.

Bruno:1998:ONG

[3429] W. J. Bruno, F. Sun, and D. C. Torney. Optimizing nonadaptive group tests for objects with heterogeneous priors. SIAM Journal on Applied Mathematics, 58(4):1043–1059, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30506.

Alessandrini:1998:ICP

[3430] Giovanni Alessandrini and Edi Rosset. The inverse conductivity problem with one measurement: Bounds on the size of the unknown object. SIAM Journal on Applied Mathematics, 58(4):1060–1071, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30646.

Rodriguez-Bernal:1998:DIC

[3431] Aníbal Rodríguez-Bernal and Erik S.

Van Vleck. Diffusion induced chaos in a closed loop thermosyphon. SIAM Journal on Applied Mathematics, 58(4):1072–1093, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/30418.

Zounes:1998:TCQ

[3432] Randolph S. Zounes and Richard H. Rand. Transition curves for the quasiperiodic Mathieu equation. SIAM Journal on Applied Mathematics, 58 (4):1094–1115, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/article/
                                        30387.

Lenci:1998:SCM

[3433] Stefano Lenci. On the suppression of chaos by means of bounded excitations in an inverted pendulum. SIAM Journal on Applied Mathematics, 58(4):1116–1127, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/30665.

Bobylev:1998:MTE

[3434] A. V. Bobylev and G. Spiga. On a model transport equation with inelastic scattering. SIAM Journal on Applied Mathematics, 58(4):1128–1137, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31180.

Degond:1998:MNW

[3435] P. Degond. A model of near-wall conductivity and its application to plasma thrusters. SIAM Journal on Applied Mathematics, 58(4):1138–1162, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/30089.

Sun:1998:OTA

[3436] S. M. Sun. On the oscillatory tails with arbitrary phase shift for solutions of the perturbed Korteweg–de Vries equation. SIAM Journal on Applied Mathematics, 58(4):1163–1177, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29921.

Sneyd:1998:TWB

[3437] James Sneyd, Paul D. Dale, and Alastair Duffy. Traveling waves in buffered systems: Applications to calcium waves. SIAM Journal on Applied Mathematics, 58(4):1178–1192, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30507.

Cohen:1998:CDR

[3438] Donald S. Cohen and Thomas Erneux.

Controlled drug release asymptotics. SIAM Journal on Applied Mathematics, 58(4):1193–1204, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29326.

Rinaldi:1998:LPI

[3439] Sergio Rinaldi. Laura and Petrarch:

An intriguing case of cyclical love dynamics. SIAM Journal on Applied Mathematics, 58(4):1205–1221, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/30592. Access to

the full text of this article is free.

He:1998:LFD

[3440] Xue zhong He. The Lyapunov functionals for delay Lotka–Volterra-type models. SIAM Journal on Applied Mathematics, 58(4):1222–1236, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/29511.

Schneider:1998:LWL

[3441] Guido Schneider. The long wave limit for a Boussinesq equation. SIAM Journal on Applied Mathematics, 58(4):1237–1245, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/28794.

Chang:1998:GSR

[3442] Hsueh-Chia Chang, Evgeny A. Demekhin, and Evgeny Kalaidin. Generation and suppression of radiation by solitary pulses. SIAM Journal on Applied Mathematics, 58(4):1246–1277, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31528.

Fosdick:1998:MNC

[3443] Roger L. Fosdick and Darren E. Mason. On a model of nonlocal continuum mechanics part I: Existence and regularity. SIAM Journal on Applied Mathematics, 58(4):1278–1306, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31480.

Rivera:1998:MCP

[3444] Jaime E. Muñoz Rivera and Reinhard Racke. Multidimensional contact problems in thermoelasticity. SIAM Journal on Applied Mathematics, 58(4):1307–1337, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/31306.

Cox:1998:LWR

[3445] Stephen M. Cox. Long-wavelength rotating convection between poorly conducting boundaries. SIAM Journal on Applied Mathematics, 58(4):1338–1364, August 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31400.

Canic:1998:NSR

[3446] Suncica Canic and Dragan Mirkovic.

A numerical study of Riemann problems for the two-dimensional unsteady transonic small disturbance equation. SIAM Journal on Applied Mathematics, 58(5):1365–1393, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30884.

Kim:1998:PBT

[3447] Sun-Chul Kim. On Prandtl–Batchelor theory of a cylindrical eddy: Asymptotic study. SIAM Journal on Applied Mathematics, 58(5):1394–1413, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30328.

Zalipaev:1998:ROA

[3448] V. Zalipaev. Ray optical analysis of a plane wave scattering by an echelette diffraction grating. SIAM Journal on Applied Mathematics, 58(5):1414–1449, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29361.

Benamou:1998:WES

[3449] J.-D. Benamou and Y. Brenier. Weak existence for the semigeostrophic equations formulated as a coupled Monge–Ampère/transport problem. SIAM Journal on Applied Mathematics, 58(5):1450–1461, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29411.

Ivansson:1998:CLF

[3450] Sven Ivansson. A class of low-frequency modes in laterally homogeneous fluidsolid media. SIAM Journal on Applied Mathematics, 58(5):1462–1508, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30225.

Luong:1998:QIC

[3451] Bruno Luong and Fadil Santosa. Quantitative imaging of corrosion in plates by eddy current methods. SIAM Journal on Applied Mathematics, 58(5):1509–1531, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/31335.

Chillan:1998:PFG

[3452] J. Chillan and J. P. Fouque. Pressure fields generated by acoustical pulses propagating in randomly layered media. SIAM Journal on Applied Mathematics, 58(5):1532–1546, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31078.

Bonnans:1998:MSV

[3453] J. Frédéric Bonnans. Mathematical study of very high voltage power networks II: The AC power flow problem. SIAM Journal on Applied Mathematics, 58(5):1547–1567, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28932.

Crisan:1998:CBP

[3454] Dan Crisan, Jessica Gaines, and Terry Lyons. Convergence of a branching particle method to the solution of the Zakai equation. SIAM Journal on Applied Mathematics, 58(5):1568–1590, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30737.

Druskin:1998:UIP

[3455] Vladimir Druskin. On the uniqueness of inverse problems from incomplete boundary data. SIAM Journal on Applied Mathematics, 58(5):1591–1603, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/29829.

Clouet:1998:DAT

[3456] J.-F. Clouet. Diffusion approximation of a transport process in random media. SIAM Journal on Applied Mathematics, 58(5):1604–1621, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31271.

Lacey:1998:MRD

[3457] A. A. Lacey, J. R. Ockendon, and J. Sabina. Multidimensional reaction diffusion equations with nonlinear boundary conditions. SIAM Journal on Applied Mathematics, 58(5):1622–1647, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                        siam.org/sam-bin/dbq/article/30812.

Anita:1998:OHP

[3458] S. Anita, M. Iannelli, M.-Y. Kim, and E.-J. Park. Optimal harvesting for periodic age-dependent population dynamics. SIAM Journal on Applied Mathematics, 58(5):1648–1666, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/30118.

Pernarowski:1998:FSS

[3459] M. Pernarowski. Fast and slow subsystems for a continuum model of bursting activity in the pancreatic islet. SIAM Journal on Applied Mathematics, 58(5):1667–1687, October 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/30458.

Harabetian:1998:RIP

[3460] Eduard Harabetian and Stanley Osher. Regularization of ill-posed problems via the level set approach. SIAM Journal on Applied Mathematics, 58(6):1689–1706, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29079.

Giacomin:1998:PSD

[3461] Giambattista Giacomin and Joel L.

Lebowitz. Phase segregation dynamics in particle systems with long range interactions II: Interface motion. SIAM Journal on Applied Mathematics, 58(6):1707–1729, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31304.

Vakakis:1998:GLD

[3462] Alexander F. Vakakis and Gary Salenger. Generation of localization in a discrete chain with periodic boundary conditions: Numerical and analytical results. SIAM Journal on Applied Mathematics, 58(6):1730–1747, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31635.

Figotin:1998:MDM

[3463] Alexander Figotin and Abel Klein.

Midgap defect modes in dielectric and acoustic media. SIAM Journal on Applied Mathematics, 58(6):1748–1773, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/32053.

Chandler-Wilde:1998:URS

[3464] Simon N. Chandler-Wilde and Bo Zhang. A uniqueness result for scattering by infinite rough surfaces. SIAM Journal on Applied Mathematics, 58(6):1774–1790, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/30972.

Schaeffer:1998:SMS

[3465] David G. Schaeffer and Michael Shearer. A simple model for stress fluctuations in plasticity with application to granular materials. SIAM Journal on Applied Mathematics, 58(6):1791–1807, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/30974.

Chapman:1998:MSV

[3466] S. J. Chapman and D. R. Heron. The motion of superconducting vortices in thin films of varying thickness. SIAM Journal on Applied Mathematics, 58(6):1808–1825, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/30733.

Charach:1998:SFS

[3467] Chaim Charach and Paul C. Fife. Solidification fronts and solute trapping in a binary alloy. SIAM Journal on Applied Mathematics, 58(6):1826–1851, December 1998.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/30606.

Wang:1998:ENS

[3468] Xiao-Ping Wang. On the existence of nontrivial solutions for a scalar “light-nematic” system and its Frederiks transition threshold. SIAM Journal on Applied Mathematics, 58(6):1852–1861, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29922.

Lerman:1998:ATL

[3469] G. Lerman and Z. Schuss. Asymptotic theory of large deviations for Markov chains. SIAM Journal on Applied Mathematics, 58(6):1862–1877, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30055.

vandenDriessche:1998:GAD

[3470] P. van den Driessche and Xingfu Zou. Global attractivity in delayed Hopfield neural network models. SIAM Journal on Applied Mathematics, 58(6):1878–1890, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32121.

Neubauer:1998:RCR

[3471] Andreas Neubauer and Otmar Scherzer.

Regularization for curve representations: Uniform convergence for discontinuous solutions of ill-posed problems. SIAM Journal on Applied Mathematics, 58(6):1891–1900, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/32145.

Schielen:1998:MES

[3472] R. Schielen and A. Doelman. Modulation equations for spatially periodic systems: Derivation and solutions. SIAM Journal on Applied Mathematics, 58(6):1901–1930, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/30318.

Zhang:1998:ASI

[3473] Bo Zhang and Simon N. Chandler-Wilde. Acoustic scattering by an inhomogeneous layer on a rigid plate. SIAM Journal on Applied Mathematics, 58(6):1931–1950, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31269.

Porsching:1998:MAO

[3474] T. A. Porsching, C. A. Hall, and T. L. Bennett. Minimax approximation of optical profiles. SIAM Journal on Applied Mathematics, 58(6):1951–1968, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/30973.

Garnier:1998:ATS

[3475] Josselin Garnier. Asymptotic transmission of solitons through random media. SIAM Journal on Applied Mathematics, 58(6):1969–1995, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31857.

Greenberg:1998:ASF

[3476] J. M. Greenberg and D. R. Owen. Antiplane shear flows in viscoplastic solids exhibiting isotropic and kinematic hardening. SIAM Journal on Applied Mathematics, 58(6):1996–2023, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31494.

Kirsch:1998:CFE

[3477] Andreas Kirsch and Peter Monk. Corrigendum: A finite element/spectral method for approximating the timeharmonic Maxwell system in ℜ^{3}. SIAM Journal on Applied Mathematics, 58(6):2024–2028, December 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/32768.

Greensite:1999:SOA

[3478] Fred Greensite. Second-order approximation of the pseudoinverse for operator deconvolutions and families of illposed problems. SIAM Journal on Applied Mathematics, 59(1):1–16, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31585.

Littman:1999:SPS

[3479] Walter Littman and Bo Liu. On the spectral properties and stabilization of acoustic flow. SIAM Journal on Applied Mathematics, 59(1):17–34, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31410.

Chen:1999:PEA

[3480] Kevin C. Chen, Roseanne M. Ford, and Peter T. Cummings. Perturbation expansion of Alt’s cell balance equations reduces to Segel’s one-dimensional equations for shallow chemoattractant gradients. SIAM Journal on Applied Mathematics, 59(1):35–57, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30128.

Maritz:1999:RAC

[3481] Milton F. Maritz, Gabor T. Herman, and Conway Yee. Recovery of the absorption coefficient from diffused reflected light using a discrete diffusive model. SIAM Journal on Applied Mathematics, 59(1):58–71, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/27288.

Levermore:1999:GMC

[3482] C. David Levermore and William J. Morokoff. The Gaussian moment closure for gas dynamics. SIAM Journal on Applied Mathematics, 59(1):72–96, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29923.

Damon:1999:GST

[3483] James Damon. Generic structure of two-dimensional images under Gaussian blurring. SIAM Journal on Applied Mathematics, 59(1):97–138, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/31803.

Brown:1999:ATD

[3484] Russell M. Brown, Peter Perry, and Zhongwei Shen. The additive turbulent decomposition for the two-dimensional incompressible Navier–Stokes equations: Convergence theorems and error estimates. SIAM Journal on Applied Mathematics, 59(1):139–155, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/31161.

Buonomo:1999:PSV

[3485] A. Buonomo. The periodic solution of Van Der Pol’s equation. SIAM Journal on Applied Mathematics, 59(1):156–171, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/31979.

Kohn:1999:SMP

[3486] Robert V. Kohn and Thomas D. Little. Some model problems of polycrystal plasticity with deficient basic crystals. SIAM Journal on Applied Mathematics, 59(1):172–197, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/32001.

vanderHeijden:1999:SCL

[3487] G. H. M. van der Heijden, A. R. Champneys, and J. M. T. Thompson. The spatial complexity of localized buckling in rods with noncircular cross section. SIAM Journal on Applied Mathematics, 59(1):198–221, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/30683.

Muller:1999:OVP

[3488] Johannes Müller. Optimal vaccination patterns in age-structured populations. SIAM Journal on Applied Mathematics, 59(1):222–241, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/29327.

Angell:1999:AMO

[3489] T. S. Angell, R. E. Kleinman, and B. Vainberg. Asymptotic methods for an optimal antenna problem. SIAM Journal on Applied Mathematics, 59(1):242–260, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/30407.

Villarroel:1999:SIP

[3490] Javier Villarroel. On the solution to the inverse problem for the Toda chain. SIAM Journal on Applied Mathematics, 59(1):261–285, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30090.

Papageorgiou:1999:SCJ

[3491] Demetrios T. Papageorgiou and Oscar Orellana. Study of cylindrical jet breakup using one-dimensional approximations of the Euler equations. SIAM Journal on Applied Mathematics, 59(1):286–317, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28823.

Hulshof:1999:ADF

[3492] Josephus Hulshof and John R. King.

Analysis of a Darcy flow model with a dynamic pressure saturation relation. SIAM Journal on Applied Mathematics, 59(1):318–346, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30886.

Brothers:1999:APA

[3493] Jerry D. Brothers and Richard Haberman. Accurate phase after slow passage through subharmonic resonance. SIAM Journal on Applied Mathematics, 59(1):347–364, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30703.

Fowler:1999:FSC

[3494] Andrew C. Fowler and Xin she Yang.

Fast and slow compaction in sedimentary basins. SIAM Journal on Applied Mathematics, 59(1):365–385, February 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28737.

Alimov:1999:HEP

[3495] M. Alimov, K. Kornev, and G. Mukhamadullina. Hysteretic effects in the problems of artificial freezing. SIAM Journal on Applied Mathematics, 59(2):387–410, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31378.

Li:1999:GAB

[3496] Bingtuan Li. Global asymptotic behavior of the chemostat: General response functions and different removal rates. SIAM Journal on Applied Mathematics, 59(2):411–422, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31100.

Eliazar:1999:RTG

[3497] Iddo Eliazar and Uri Yechiali. Randomly timed gated queueing systems. SIAM Journal on Applied Mathematics, 59(2):423–441, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/29728.

Kress:1999:IOS

[3498] Rainer Kress and William Rundell. Inverse obstacle scattering using reduced data. SIAM Journal on Applied Mathematics, 59(2):442–454, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31659.

Cahn:1999:TWS

[3499] John W. Cahn, John Mallet-Paret, and Erik S. Van Vleck. Traveling wave solutions for systems of ODEs on a two-dimensional spatial lattice. SIAM Journal on Applied Mathematics, 59(2):455–493, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/31270.

Morrison:1999:RAA

[3500] John A. Morrison, K. G. Ramakrishnan, and Debasis Mitra. Refined asymptotic approximations to loss probabilities and their sensitivities in shared unbuffered resources. SIAM Journal on Applied Mathematics, 59(2):494–513, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/32647.

Wang:1999:SSA

[3501] C.-W. Wang and Z. Rusak. Similarity solutions of ϕ_{x}^{3}ϕ_{xx} = ϕ_{ỹỹ} with applications to transonic aerodynamics of dense gases. SIAM Journal on Applied Mathematics, 59(2):514–528, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31482.

Merzon:1999:HFA

[3502] A. E. Merzon and P. N. Zhevandrov. High-frequency asymptotics of edge waves on a beach of nonconstant slope. SIAM Journal on Applied Mathematics, 59(2):529–546, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31785.

Powell:1999:CCM

[3503] James A. Powell, Tyler McMillen, and Peter White. Connecting a chemotactic model for mass attack to a rapid integro-difference emulation strategy. SIAM Journal on Applied Mathematics, 59(2):547–572, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31345.

Ballyk:1999:ERM

[3504] Mary Ballyk, Le Dung, Don A. Jones, and Hal L. Smith. Effects of random motility on microbial growth and competition in a flow reactor. SIAM Journal on Applied Mathematics, 59(2):573–596, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32534.

Borcea:1999:AAQ

[3505] Liliana Borcea. Asymptotic analysis of quasi-static transport in high contrast conductive media. SIAM Journal on Applied Mathematics, 59(2):597–635, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32134.

Birnir:1999:FDA

[3506] Björn Birnir and Höskuldur Ari Hauksson. A finite dimensional attractor of the Moore–Greitzer PDE model. SIAM Journal on Applied Mathematics, 59(2):636–650, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/32552.

Chen:1999:SSE

[3507] Shuping Chen, Kangsheng Liu, and Zhuangyi Liu. Spectrum and stability for elastic systems with global or local Kelvin–Voigt damping. SIAM Journal on Applied Mathematics, 59(2):651–668, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29201.

Sylvester:1999:LNI

[3508] John Sylvester and Dale P. Winebrenner. Linear and nonlinear inverse scattering. SIAM Journal on Applied Mathematics, 59(2):669–699, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31977.

Edwards:1999:ACA

[3509] D. A. Edwards and D. A. French.

Asymptotic and computational analysis of large shear deformations of a thermoplastic material. SIAM Journal on Applied Mathematics, 59(2):700–724, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32241.

Labreuche:1999:PIR

[3510] Christophe Labreuche. Purely imaginary resonant frequencies for a lossy inhomogeneous medium. SIAM Journal on Applied Mathematics, 59(2):725–742, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/32856.

Zhukov:1999:MTP

[3511] Mikhail Yu. Zhukov, Sergey V. Ermakov, and Pier Giorgio Righetti. Modeling of transport processes in the presence of substance-locking effects. SIAM Journal on Applied Mathematics, 59(2):743–776, April 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/30208.

Hu:1999:NAI

[3512] Jishan Hu. Nonuniformity of adiabatic invariance for nonlinear oscillators. SIAM Journal on Applied Mathematics, 59(3):777–786, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/32758.

Geigant:1999:IMO

[3513] Edith Geigant, Karina Ladizhansky, and Alexander Mogilner. An integrodifferential model for orientational distributions of F -actin in cells. SIAM Journal on Applied Mathematics, 59(3):787–809, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/30953.

Luo:1999:LTB

[3514] Tao Luo, Roberto Natalini, and Zhouping Xin. Large time behavior of the solutions to a hydrodynamic model for semiconductors. SIAM Journal on Applied Mathematics, 59(3):810–830, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31216.

Braasch:1999:NSS

[3515] Peter Braasch, Gerhard Rein, and Jesenko Vukadinovic. Nonlinear stability of stationary plasmas—an extension of the energy-Casimir method. SIAM Journal on Applied Mathematics, 59(3):831–844, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/32906.

Biler:1999:GES

[3516] Piotr Biler and Wojbor A. Woyczynski. Global and exploding solutions for nonlocal quadratic evolution problems. SIAM Journal on Applied Mathematics, 59(3):845–869, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31344.

McCabe:1999:ETW

[3517] P. M. McCabe, J. A. Leach, and D. J.

Needham. The evolution of travelling waves in fractional order autocatalysis with decay. I. permanent form travelling waves. SIAM Journal on Applied Mathematics, 59(3):870–899, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31259.

Granzow:1999:DPS

[3518] Glen D. Granzow and Hermann Riecke. Double phase slips and spatiotemporal chaos in a model for parametrically excited standing waves. SIAM Journal on Applied Mathematics, 59(3):900–919, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31386.

Kawaguchi:1999:CTW

[3519] Satoshi Kawaguchi and Masayasu Mimura. Collision of travelling waves in a reaction-diffusion system with global coupling effect. SIAM Journal on Applied Mathematics, 59(3):920–941, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/30664.

Class:1999:LFC

[3520] Andreas Class, Alvin Bayliss, and Bernard J. Matkowsky. Localized flickering cellular flames. SIAM Journal on Applied Mathematics, 59(3):942–964, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31784.

Helsing:1999:IEM

[3521] Johan Helsing and Gunnar Peters. Integral equation methods and numerical solutions of crack and inclusion problems in planar elastostatics. SIAM Journal on Applied Mathematics, 59(3):965–982, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/33293.

Klar:1999:HMMa

[3522] Axel Klar and Raimund Wegener.

A hierarchy of models for multilane vehicular traffic I: Modeling. SIAM Journal on Applied Mathematics, 59(3):983–1001, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/32694.

Klar:1999:HMMb

[3523] Axel Klar and Raimund Wegener. A hierarchy of models for multilane vehicular traffic II: Numerical investigations. SIAM Journal on Applied Mathematics, 59(3):1002–1011, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32695.

Hanke:1999:EAE

[3524] Martin Hanke and Otmar Scherzer.

Error analysis of an equation error method for the identification of the diffusion coefficient in a quasilinear parabolic differential equation. SIAM Journal on Applied Mathematics, 59(3):1012–1027, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/33162.

Lafitte:1999:DHFa

[3525] Olivier D. Lafitte. Diffraction in the high frequency regime by a thin layer of dielectric material I: The equivalent impedance boundary condition. SIAM Journal on Applied Mathematics, 59(3):1028–1052, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/30715.

Lafitte:1999:DHFb

[3526] Olivier D. Lafitte. Diffraction in the high frequency regime by a thin layer of dielectric material II: the trace of the wave in the shadow of the obstacle. SIAM Journal on Applied Mathematics, 59(3):1053–1079, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/30716.

Gregory:1999:THS

[3527] R. Douglas Gregory, Thomas I. Milac, and Frederic Y. M. Wan. A thick hollow sphere compressed by equal and opposite concentrated axial loads: An asymptotic solution. SIAM Journal on Applied Mathematics, 59(3):1080–1097, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31573.

Borrelli:1999:SVP

[3528] A. Borrelli and M. C. Patria. Saint– Venant’s principle for a piezoelectric body. SIAM Journal on Applied Mathematics, 59(3):1098–1111, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31932.

Shadwick:1999:ECI

[3529] B. A. Shadwick, John C. Bowman, and P. J. Morrison. Exactly conservative integrators. SIAM Journal on Applied Mathematics, 59(3):1112–1133, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/28931.

Edwards:1999:SDD

[3530] D. A. Edwards. Skinning during desorption of polymers: An asymptotic analysis. SIAM Journal on Applied Mathematics, 59(3):1134–1155, June 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31106.

Pilyugin:1999:CUC

[3531] Sergei S. Pilyugin and Paul Waltman. Competition in the unstirred chemostat with periodic input and washout. SIAM Journal on Applied Mathematics, 59(4):1157–1177, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32395.

Georgiou:1999:DLS

[3532] Ioannis T. Georgiou and Ira B.

Schwartz. Dynamics of large scale coupled structural/mechanical systems: A singular perturbation/proper orthogonal decomposition approach. SIAM Journal on Applied Mathematics, 59(4):1178–1207, August 1999.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/29980.

Bornemann:1999:SLQ

[3533] Folkmar A. Bornemann and Christof Schütte. On the singular limit of the quantum-classical molecular dynamics model. SIAM Journal on Applied Mathematics, 59(4):1208–1224, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31883.

Du:1999:SGL

[3534] Qiang Du. Studies of a Ginzburg– Landau model for d-wave superconductors. SIAM Journal on Applied Mathematics, 59(4):1225–1250, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/32990.

Goldshtein:1999:SUM

[3535] V. Gol’dshtein, A. C. McIntosh, and A. Zinoviev. A spatially uniform model of self-ignition due to combustible fluid leakage in insulation materials: The effect of initial conditions. SIAM Journal on Applied Mathematics, 59(4):1251–1266, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/31044.

Obnosov:1999:PHS

[3536] Yurii V. Obnosov. Periodic heterogeneous structures: New explicit solutions and effective characteristics of refraction of an imposed field. SIAM Journal on Applied Mathematics, 59(4):1267–1287, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/31477.

Holmes:1999:DBP

[3537] Philip Holmes and J. Nathan Kutz.

Dynamics and bifurcations of a planar map modeling dispersion managed breathers. SIAM Journal on Applied Mathematics, 59(4):1288–1302, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33003.

Chow:1999:TAN

[3538] Shui-Nee Chow, Bo Deng, and Mark J.

Friedman. Theory and application of a nongeneric heteroclinic loop bifurcation. SIAM Journal on Applied Mathematics, 59(4):1303–1321, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/29454.

Ammari:1999:SME

[3539] Habib Ammari, Chiraz Latiri-Grouz, and Jean-Claude Nédélec. Scattering of Maxwell’s equations with a Leontovich boundary condition in an inhomogeneous medium: A singular perturbation problem. SIAM Journal on Applied Mathematics, 59(4):1322–1334, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32486.

Sanchez:1999:CBE

[3540] Antonio L. Sánchez, Amable Liñán, and Forman A. Williams. Chainbranching explosions in mixing layers. SIAM Journal on Applied Mathematics, 59(4):1335–1355, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/32648.

Knowles:1999:UEI

[3541] Ian Knowles. Uniqueness for an elliptic inverse problem. SIAM Journal on Applied Mathematics, 59(4):1356–1370, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/32778.

Iosilevskii:1999:AAT

[3542] G. Iosilevskii, Ya. A. Iosilevskii, and A. Rosen. Asymptotic aerodynamic theory of oscillating rotary wings in axial flight. SIAM Journal on Applied Mathematics, 59(4):1371–1412, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28735.

Kress:1999:FFO

[3543] Rainer Kress and Lassi Päivärinta. On the far field in obstacle scattering. SIAM Journal on Applied Mathematics, 59(4):1413–1426, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/33225.

vanHorssen:1999:PMBa

[3544] W. T. van Horssen. A perturbation method based on integrating factors. SIAM Journal on Applied Mathematics, 59(4):1427–1443, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/30915.

vanHorssen:1999:PMBb

[3545] W. T. van Horssen. A perturbation method based on integrating vectors and multiple scales. SIAM Journal on Applied Mathematics, 59(4):1444–1467, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32686.

Lindstrom:1999:DBC

[3546] Torsten Lindström. Dependencies between competition and predation—and their consequences for initial value sensitivity. SIAM Journal on Applied Mathematics, 59(4):1468–1486, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32085.

Lehman:1999:ATD

[3547] Brad Lehman and Steven P. Weibel.

Averaging theory for delay difference equations with time-varying delays. SIAM Journal on Applied Mathematics, 59(4):1487–1506, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32567.

Cortesani:1999:NAN

[3548] Guido Cortesani and Rodica Toader. Nonlocal approximation of nonisotropic free-discontinuity problems. SIAM Journal on Applied Mathematics, 59(4):1507–1519, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/32769.

Weidman:1999:ISF

[3549] P. D. Weidman and V. Calmidi. Instantaneous Stokes flow in a conical apex aligned with gravity and bounded by a stress-free surface. SIAM Journal on Applied Mathematics, 59(4):1520–1531, August 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33410.

Kang:1999:ICP

[3550] Hyeonbae Kang and Jin Keun Seo. Inverse conductivity problem with one measurement: Uniqueness of balls in R^{3}. SIAM Journal on Applied Mathematics, 59(5):1533–1539, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32459.

Plank:1999:DLV

[3551] Manfred Plank. On the dynamics of Lotka–Volterra equations having an invariant hyperplane. SIAM Journal on Applied Mathematics, 59(5):1540–1551, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/31946.

Pilyugin:1999:SCW

[3552] Sergei S. Pilyugin and Paul Waltman. The simple chemostat with wall growth. SIAM Journal on Applied Mathematics, 59(5):1552–1572, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32618.

Conca:1999:FHM

[3553] Carlos Conca and Fernando Lund.

Fourier homogenization method and the propagation of acoustic waves through a periodic vortex array. SIAM Journal on Applied Mathematics, 59(5):1573–1581, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/32268.

Jonsson:1999:ECR

[3554] Elias Jonsson. Electrical conductivity reconstruction using nonlocal boundary conditions. SIAM Journal on Applied Mathematics, 59(5):1582–1598, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/32777.

Duterte:1999:NMS

[3555] Jean Duterte and Patrick Joly. A numerical method for surface waves in a cylindrically perturbed elastic half-space. part 1: Construction and analysis. SIAM Journal on Applied Mathematics, 59(5):1599–1635, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/31802.

Chen:1999:ITR

[3556] Pei Li Chen and Vadim Kagansky. The influence of timing on radiation damage to cell populations. SIAM Journal on Applied Mathematics, 59(5):1636–1666, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/30805.

Calsina:1999:GDO

[3557] Àngel Calsina and Joan Saldaña. Global dynamics and optimal life history of a structured population model. SIAM Journal on Applied Mathematics, 59(5):1667–1685, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/33123.

Budhiraja:1999:SNS

[3558] Amarjit Budhiraja and Paul Dupuis. Simple necessary and sufficient conditions for the stability of constrained processes. SIAM Journal on Applied Mathematics, 59(5):1686–1700, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/33022.

Belenkaya:1999:USO

[3559] L. Belenkaya, S. Friedlander, and V. Yudovich. The unstable spectrum of oscillating shear flows. SIAM Journal on Applied Mathematics, 59(5):1701–1715, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/32757.

Serra:1999:AAG

[3560] Stefano Serra and Cristina Tablino Possio. Analytical analysis of the Gavrilov–Guckenheimer bifurcation unfolding in the case of a proportional-integral controlled CSTR. SIAM Journal on Applied Mathematics, 59(5):1716–1744, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/29489.

Athanasiadis:1999:ESH

[3561] C. Athanasiadis, P. A. Martin, and I. G. Stratis. Electromagnetic scattering by a homogeneous chiral obstacle: Boundary integral equations and low-chirality approximations. SIAM Journal on Applied Mathematics, 59(5):1745–1762, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33633.

Klibanov:1999:NSP

[3562] Michael V. Klibanov and Thomas R.

Lucas. Numerical solution of a parabolic inverse problem in optical tomography using experimental data. SIAM Journal on Applied Mathematics, 59(5):1763–1789, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33417.

Castillo-Chavez:1999:CEC

[3563] Carlos Castillo-Chavez, Wenzhang Huang, and Jia Li. Competitive exclusion and coexistence of multiple strains in an SIS STD model.

SIAM Journal on Applied Mathematics, 59(5):1790–1811, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/32586.

Cook:1999:SSA

[3564] L. Pamela Cook, Elsa Newman, Scott Rimbey, and Gilberto Schleiniger. Sonic and subsonic axisymmetric nozzle flows. SIAM Journal on Applied Mathematics, 59(5):1812–1824, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/32735.

Lukkassen:1999:NRS

[3565] Dag Lukkassen. A new reiterated structure with optimal macroscopic behavior. SIAM Journal on Applied Mathematics, 59(5):1825–1842, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/32008.

Kupferman:1999:OSP

[3566] Raz Kupferman and Alexandre J. Chorin. One-sided polarization and renormalization flow in the Kosterlitz–Thouless phase transition. SIAM Journal on Applied Mathematics, 59(5):1843–1866, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/33963.

Saez:1999:DPP

[3567] E. Sáez and E. González-Olivares. Dynamics of a predator-prey model.

SIAM Journal on Applied Mathematics, 59(5):1867–1878, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/31845.

Nakamura:1999:LST

[3568] Gen Nakamura, Kazumi Tanuma, and Gunther Uhlmann. Layer stripping for a transversely isotropic elastic medium. SIAM Journal on Applied Mathematics, 59(5):1879–1891, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33716.

Berlyand:1999:HHM

[3569] L. Berlyand and E. Khruslov. Homogenization of harmonic maps and superconducting composites. SIAM Journal on Applied Mathematics, 59 (5):1892–1916, October 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/
32852.

Knessl:1999:TTQa

[3570] Charles Knessl and Charles Tier.

Two tandem queues with general renewal input I: Diffusion approximation and integral representations. SIAM Journal on Applied Mathematics, 59(6):1917–1959, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32846.

Knessl:1999:TTQb

[3571] Charles Knessl and Charles Tier.

Two tandem queues with general renewal input II: Asymptotic expansions for the diffusion model. SIAM Journal on Applied Mathematics, 59(6):1960–1997, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/32847.

Siegel:1999:ISR

[3572] Michael Siegel. Influence of surfactant on rounded and pointed bubbles in two-dimensional Stokes flow. SIAM Journal on Applied Mathematics, 59(6):1998–2027, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32743.

Hazard:1999:SFM

[3573] Anne-Sophie Bonnet-Ben Dhia Christophe Hazard and Stephanie Lohrengel. A singular field method for the solution of Maxwell’s equations in polyhedral domains. SIAM Journal on Applied Mathematics, 59(6):2028–2044, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32338.

Ammari:1999:CME

[3574] H. Ammari, K. Hamdache, and J.- C. Nédélec. Chirality in the Maxwell equations by the dipole approximation. SIAM Journal on Applied Mathematics, 59(6):2045–2059, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33416.

Ostrov:1999:VSC

[3575] Daniel N. Ostrov. Viscosity solutions and convergence of monotone schemes for synthetic aperture radar shape-from-shading equations with discontinuous intensities. SIAM Journal on Applied Mathematics, 59(6):2060–2085, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/32717.

Almgren:1999:SOP

[3576] Robert F. Almgren. Second-order phase field asymptotics for unequal conductivities. SIAM Journal on Applied Mathematics, 59(6):2086–2107, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/33002.

Cox:1999:MBG

[3577] Steven J. Cox and David C. Dobson. Maximizing band gaps in two-dimensional photonic crystals. SIAM Journal on Applied Mathematics, 59(6):2108–2120, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/33845.

Rossokhataya:1999:MMG

[3578] Nataliya Rossokhataya, Vladimir Makarov, Victor Rossokhaty, and Ivan Gavrilyuk. Mathematical model of the graded-band-gap semiconductor structure with high internal quantum efficiency. SIAM Journal on Applied Mathematics, 59(6):2121–2138, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/28800.

Albert:1999:SWS

[3579] John P. Albert, Jerry L. Bona, and Juan Mario Restrepo. Solitarywave solutions of the Benjamin equation. SIAM Journal on Applied Mathematics, 59(6):2139–2161, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/32168.

Kodama:1999:WEO

[3580] Yuji Kodama. The Whitham equations for optical communications: Mathematical theory of NRZ. SIAM Journal on Applied Mathematics, 59(6):2162–2192, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/32815.

Izhikevich:1999:WCQ

[3581] Eugene M. Izhikevich. Weakly connected quasi-periodic oscillators, FM interactions, and multiplexing in the brain. SIAM Journal on Applied Mathematics, 59(6):2193–2223, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/33062.

Katsevich:1999:CBL

[3582] Alexander Katsevich. Cone beam local tomography. SIAM Journal on Applied Mathematics, 59(6):2224–2246, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33604.

Dai:1999:LCA

[3583] Longxiang Dai and Robert M. Miura.

A lattice cellular automata model for ion diffusion in the brain-cell microenvironment and determination of tortuosity and volume fraction. SIAM Journal on Applied Mathematics, 59(6):2247–2273, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32394.

Fontelos:1999:SBS

[3584] M. A. Fontelos and J. J. L. Velázquez.

On some breakup and singularity formation mechanisms for inviscid liquid jets. SIAM Journal on Applied Mathematics, 59(6):2274–2300, December 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33206.

Ruan:2006:EGS

[3585] Shigui Ruan and Xue-Zhong He. Erratum: “Global Stability in Chemostat-Type Competition Models with Nutrient Recycling”. SIAM Journal on Applied Mathematics, 66(6):2204–2205, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [3390].